A Complete Bibliography of Publications in Journal of Computational and Applied Mathematics (2010–2014)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

22 March 2018 Version 1.03

Title word cross-reference

(A, η, m) [PY10b]. (I + S_{m}) [Mor10b]. (m, k) [dJMPPC14]. (M, N) [dJP13, dJMPPC14]. (P, Q, η) [JWW10]. (R, S) [YGHH14]. 0 [ARF14, GSM11, Lay13]. 1 [ARF14, BY14, CZZ14, CLP10, GSM11, Lay13, LWXW12, TY11, ZCC^{+}13]. 11 [NBBKV10]. 12 [DY13]. 2 [AHLJ14, ADL^{+}10a, Bab10, BG11, BGIN10, BKB14, BJ12, BHMW13, CPP13, CvdMS12, DLG14, DD14b, FGEM10, JMS10b, Kha10, LEKT10, MS13b, NTPVNHLX14, OS10, UBSG12, YZ10b]. 25 [CWY11]. 2p [CHMT10a]. 3 [BSPD14, BCD^{+}10, CPP13, CDQ14, Col13, GCY13, HKL11, JMS10a, JJE13, KSA14, Lia14, PDE^{+}13, PRT10, QYZZ14, QZX^{+}11, Rah14, SBFC10, SSK^{+}13, ZD10]. 32 × 32 [SA14]. 4 [DDG^{+}13b, Jia10]. 7 [NBBKV10]. [−1, 1] [CBGVPP10]. [−∞, 0] [BGVHN13]. ^{+} [MDVB13]. _{2}F_{1} [DJJ13]. A [GPHAM10, GPHAS14, Yeu14]. A_{j} [Esc14b]. α [AHK14b, BHBP10, NTRNXB10, YZ10b]. αβ [Akt14a]. AX = B [HD11, LsGHZ10]. AXB + CY D = E [SC10c]. AXB = C [yP10b]. B [pLjH10, LP10c]. β [FCS12]. c [El 11, RPS^{+}11]. C^{0} [GG13b]. C^{1} [BGIN10, BBK^{+}14, KLK13, NRS11, SSR10]. C^{2} [DY13]. C_{0} [Mia10]. C_{p} [AÖ11]. C^{r} [FGPP13]. H^{2} [AHK14a]. PD [PD13]. Q_{1} [DLP13]. χ^{2} [Mah11]. D [LTW10, Yan10a, BLO11, BR11a, CG11b]. δ [MRD13]. Erlang(n) [ZZ13]. F_{2} [BR14a]. F_{D} [Zho11a]. g [ENSCS12]. G^{1} [HwK11, JŽ11, MSWY12]. G^{2} [CP13b, HwK11, Lu13a]. G^{3} [CC12, Lu14]. G^{k} [JJK^{+}13]. GARCH(1, 1) [BÜ14]. GCROT(m, k) [MZL14]. H [BL10a, BG14b, Che14b, HMS11, WH12b, GZ14b]. H(·, ·) [jTW14]. H(div) [CW14b]. H(div) [CHX11]. H^{1} [SM14b]. H_{0}^{2} [KN10a]. H_{+} [ZY14b]. h_{2} [BGKVW10]. Hcurl [DDG13a]. Hdiv [DDG13a]. hp [CL11e, Gia12, Gia14, GZ14a, Gwi13, HHT12, KSA14, MSP12, Sch14c, SDK10, SCDA10, SACS10, Wih11, PSKA14, SKK14]. I [Dom10]. I^{(−1)} [HX14a]. K [XCY12, EX14, JGH^{+}13, KD10, KPW10, NTNBV14, YE14]. k + 1 [XSH10]. {K, s + 1} [LRT13]. k − 0 [MA14a]. l = 2 [GD13]. L_{2}^{(m)}(0, 1) [SH11]. L^{∗}a^{∗}b^{∗} [RMTO11]. L^{1} [ADL^{+}10b]. L^{2} [BS11b, VMMM10]. L^{∞} [HS12]. L^{p} [SWT12]. l_{1} [XHmDW14]. L_{2}^{(m)}(0, 1) [SHN13]. L_{p} [SYİ11]. Lip(ω(t), p) [SS14c]. LlogL(log log L)(S^{N}) [AA12]. LQR [HKO11]. LU [GWX10, WH12a, WH12b]. M [dL14b, ML10, Nad11]. M(a, b; z) [LP10c]. F_{2} [Har11]. F_{5^{5n}} [CÖ11a]. F_{7^{7n}} [CÖ11a]. R^{2} [DL14a]. R^{4} [HwKB14]. N [SCRG12, ZD10, LT11, LWZ13, Nad11, SH10, Yeu14]. n: G [EX14]. O(\sqrt{n}L) [FF14]. ω [KD10]. P [AGM12, DP14b, SASH11, CZZ14, GKLZ12, GZ14b, KLL11a, KLM14, Kuz10, LHC14, Obe13, WZ10d, Zie14]. P_{∗} [KC11]. P_{0} [ZLL10b]. P_{1} [FKNS11]. P_{k}Q_{l} [AB10]. ϕ [QCK10]. pq [EI14]. q [AB11a, ÁNA14, Arv10b, BLO11, HCQ14, LW10a, Mah11, YE14]. QR [AMMS12, MPS10, VVM10]. R^{2} [BP13]. R^{3} [BP13]. S [KB13, ADCC12, DSY^{+}11, DSZ11, TS10]. s, L [TS10]. S_{N} [RW10]. SL(2, R) [Liu12]. SO(2, n)/S(O(2) × O(n)) [El 11]. T [DCJ13a, GH10a, SW13, TCD11, WZY^{+}14]. θ [SLD13, ZFH10, ZW14]. U(a, b; z) [LP10c]. V [WCZ11]. W [Fuk13d, KW11, RS10a]. W (L^{p}, ξ(t)) [SS14b]. X + A^{∗}X^{−α}A = Q [Zha13, WL10c]. ξ [JK10b]. y′′ = f(x, y) [DEP11]. y′′ = f(y(t)) [DP14b]. y^{2} = x^{3} − c [Kır11]. z [LP10c]. Z_{p} [AÖ11].

- [Yan10a, KD10]. -accretive [PY10b, jTW14]. -adaptive [Kuz10, Gia12, Gia14, HHT12, Wih11]. -analogue [HCQ14]. -binomial [YE14]. -body [SCRG12]. -carrier [ZD10]. -Ceságro [AB11a]. -circulants [ENSCS12]. -class [SS14c]. -coherent [dJP13, dJMPPC14]. -constrained [Lu13a]. -continuity [CP13b]. -D [Col13, DLG14]. -DG [SKK14]. -dimensional [LWZ13]. -FEM [PSKA14, SKK14]. -finite [FGPP13]. -flame [WCZ11]. -fold [Jia10]. -function [El 11, HMS11, MRD13, ZLL10b]. -functions [Fuk13d]. -gap [LTW10]. -geometric [YE14]. -Hahn [Arv10b]. -iteration [FF14]. -Jacobi [LW10a]. -Laguerre [BLO11]. -Laplace [ML10, Obe13]. -Laplacian [KLL11a, KLM14, WZ10d]. -linear [Har11]. -matrices [WH12b]. -matrix [GH10a, dL14b, ZY14b]. -means [RPS^{+}11, XCY12]. -Meixner [FCS12]. -meshes [DCJ13a, Sch14c, SW13, TCD11, WZY^{+}14]. -method [ZFH10]. -methods [SLD13, KW11, RS10a]. -model [SWT12]. -monotone [pLjH10]. -nonconforming [FKNS11]. -nonexpansive [QCK10]. -norm [LHC14, BGKVW10]. -operator [PD13]. -orthogonal [BR11a, CG11b]. -orthogonality [BLO11]. -out-of- [EX14]. -palindromic [BG14b]. -point [CWY11]. -potent [LRT13]. -projection [KN10a]. -Racah [ÁNA14]. -reflexive [JWW10]. -regular [BS11b]. -regularizer [XHmDW14]. -ROCK [KB13]. -sector [Zie14]. -semigroups [Mia10]. -singularities [Esc14b]. -stability [VMMM10]. -stable [AGM12, ADL^{+}10b, DP14b, GPHAM10, GPHAS14, SASH11].

-stage [Kha10, NBBKV10]. -statistical [AB11a, Akt14a]. -step [ADCC12, LT11, VVM10, XSH10]. -step-ahead [ZCC^{+}13]. -strict [KPW10, JK10b]. -symmetric [BR11a, CG11b, YGHH14]. -Szász [Mah11]. -th [GKLZ12, JGH^{+}13, Zie14]. -th-order [SH10]. -theory [AHK14b]. -Toeplitz [ENSCS12]. -transpose [Yeu14]. -up [MDVB13]. -version [CZZ14]. -wavelet [AHK14a]. -weighted [Che14b].

1 [WYP10, ZH10a]. 10.1016/j.cam.2009.08.041 [DC10]. 143 [Gün10].

2-norms [UM14]. 2011 [Sol12]. 224 [LHL11, TSC10, ZJ12]. 228 [Wan11c]. 230 [Wan11c]. 231 [Dai11, LB10]. 233 [Asl10, BAdS^{+}12, WYO11, vDvFZ13]. 236 [DS13c, XSV13]. 262 [DD14b]. 266 [ÁD15]. 271 [CDKN15].

360 [WYO11].

4th [LLSW11].

60th [MYZ10]. 61 [Kre11].

70th [ARV10a].

8-stage [NTNBV14].

A. [Khr10]. ablation [MV12b]. ablation-type [MV12b]. abrasive [KS12b]. abrupt [KBKR14]. abscissae [Not14]. absolute [HHZ11, Liu11c]. absorbing [Bab10]. absorption [Dra13]. abstract [RT12]. Accelerated [CHMT10a, NLL13, And10, BY12, GGAVRC14, Ism14, Wan11f, WSS14, ZY14b]. accelerated-time [GGAVRC14]. accelerating [CCS11, LEVG14]. Acceleration [ENS14, KSRM12, DZW14, Kuz12, Zho12]. accelerators [GPVL10]. accretive [PY10b, jTW14]. accuracy [BG13, CS14, Cie14, CyYwY11, HMA12, Lia14, MT10a, Oou13, TTB14, ZTC11]. Accurate [Cha10a, JGH^{+}13, Rum13a, Rum13b, Wri10, ZD10, vGALL10, ADGP13, CD10, CL11a, CHH10, DR14b, ESM13b, EN14, GP11, Kho10, KN10a, MV12b, WCH11, pWqSmZ14]. Acoustic [GMZ^{+}10, AV10b, BP13, DLG14, DDV^{+}11, HLLW13, Lia14, PRT10, RL10, TA13, VVD^{+}14, WUB10]. acoustics [AM10b, DMMP10, HLSX11, HNPX11]. acousto [ILEE13, MR10, Mön10]. acousto-elastic [Mön10]. acousto-gravitational [MR10]. acousto-structural [ILEE13]. ACSP [SBM14]. active [Uty10b, Uty10a, ZTP14]. activity [LL13a]. actuators [APO14]. adaptation [AZ11, BFB14, Dea11, HK14, KM10]. adapted [SW12a]. Adaptive [CGRS11, DQST14, FN10, KK14, PC12a, Prz10, SDK10, TCD11, BS13b, BFHK12, BL10b, BHPS12, CZWD14, CGS12, CC11a, CC14a, DME^{+}12, EN14, FPP14, FiMO10, FV10, Gia12, Gia14, GH14b, GAS14b, GCC14, HHT12, HK13a, HJS14, Ism14, IG12, KKR11, KS14d, KBKR14, KMLT10, KMP14, KSA14, Kuz10, Li10, LMY11, MKS12, Mar11, NÖ10, ÖW14, RW10, SLP11, SCDA10, SACS10, SG12, SKK14, WCH11, Wih11, XN14, ZTL13]. Addendum [CDKN15, Pir14]. addition [Vel10]. Additive [LCI11, BS11b, BB12c, GLHCLPN10, Sla14, Wik10]. addressing [HDT11]. ADI [JK13, SBL12]. adjacency [MHH^{+}14]. Adjoint [MMP11, SLW14]. Adjoint-based [MMP11]. adjustable [FJLL10]. adjusting [ZZC10]. ADM [AF10]. admissible [BSV10]. admittance [Pag10]. Adomian [Eba11, Mah14]. adsorption [MHL11]. Advanced [FHMZ10, WZL11]. Advances [AGK^{+}14, VAV11, AB14, BK14c, FFM^{+}11]. advancing [CFMR12]. advection [Bay14, BB12a, GVM^{+}12, GZ11, GN11, GPHAPR14, HML14, MPS11, PMC11, RRMD13, SRP^{+}13, TR11, VS14, ZW10]. advection-dispersion [HML14, ZW10]. advection-dominated [SRP^{+}13]. aeroacoustics [NCCF^{+}10]. aeroelastic [DAW11]. aerosol [CFL^{+}14]. aesthetic [MSWY12]. affected [GLHCLPN10]. affine [EF14, Liu14a, NS13]. affinely [UM14]. age [LSW11, mZkPkL11]. age-dependent [mZkPkL11]. age-structured [LSW11]. Aggregate [TBK14]. aggregating [DLSV14]. aggregation [PM11b]. aggregation-disaggregation [PM11b]. agricultural [WDG^{+}12]. ahead [ZCC^{+}13]. AHSS [CJ10]. air [WCZ11, ZS11, HSH12]. air-vapor-heat [ZS11]. aircraft [ZGZB13]. airfoil [NdCDCFM10]. airfoils [MSP12]. al. [SPLHCB14]. algebra [HV11, Spe11]. Algebraic [CR11b, BY11, BS11a, BSS14, BSS13, BHJ11, BL11, CK13, CLS14, CA14, GPHAS14, GG14, HSH12, KX13, LM12a, LwCLW13, LM14, NOS14, PC12a, PV10, PGH13, Pul14, SA14, VS14, WLD14, XN14, YL11a, YZ11a, YLD13, ZW11a]. algebraic-trigonometric [YL11a]. algebraically [CDJP13, IJ14]. algebras [But12, CGRLR11]. Algorithm [MD13, AMMS12, AMMS14, AZ11, AG10a, AGT12, AR10a, And13, ARF14, BGRS12, BR14b, BSBA14, BL14, BFMS11, CHL11, CC14b, CC14a, CSSX14a, CSSX14b, CLS10, Cin12, Cin14, CRF14, DBHV10, DP11, DW10, DKH^{+}10, FF14, FB12, GLH12, GCNL14, Gen12, GG14, HKK10, Har11, HMA12, HK12, HDT11, HTM14, HL10b, HXSF14, HXW11, IHRA11, IA13, Iid12, JK10a, JWW14, KKR11, KC11, KSLD11, KBKR14, KDG^{+}13, KR14, KG12, Lay13, LCXL11, xLwCL10, LZH14, LZI10, LFY11, LJ12, LPLX10, LPGX10, LG11, LJ13, LZWJ11, MAK^{+}14, MB11, MY14, McN10, Mir14, PFLZ11, Prz10, QZX^{+}11, RPS^{+}11, SMD14, SBFC10, Str12, TT12, TJL11, jTW14, TOF14, WZD10, WSWD11, WW10c, WK10, XZL11, XSW12, XMC11, YWL14, ZSG^{+}14, ZH10c, ZW12, ZLL10b, ZZG14]. algorithm-based [MAK^{+}14]. Algorithms [BC10, CW13, FHMZ10, Gra11, HS11a, HV14, LRT13, Rum13a, Rum13b, YKL11, AEWMES11, And10, AuI13, BMS11, BFS14, CSY10, DC11, EY10, GH14a, GLHCLPN10, GZ11, HLLW13, JKV11, KK14, KR10, LEVG12, LH11b, LdHA10, MRT13b, MRT13a, Mol10, MY13, NACLR12, OORO12, PWZ11, SS14d, WW10a, Xu11b]. All-angle-negative-refraction [FGEM10]. all-at-once [YK14]. alleviate [GCY13]. Allmaras [DH13]. allocation [EC13, GCNL14, HW10, Iid12, LNJ^{+}10, LNB^{+}12]. allowing [JY13]. alloy [KS12b]. Almost [Alz10, Lan12, MSG11, Rie13, CL11c, CGL10, CDJP13, DN13, EO14, FS12a, RK11]. along [KR10]. alpha [Ayd14, NTRNXB10, NTPVNHLX14]. alpha-FEM [NTPVNHLX14]. alternately [WZL11]. Alternating [AMRS12, Bog14b, CC14a, PL10b, ZBC13, Cui10]. alternative [CDTV11, NTRNXB10]. alveolar [MP10b]. always [LZI10]. ambiguity [PP14]. American [FS10, HL14, LL13a, ZYL11]. American-style [HL14]. AMF [BGPPRW14, GPHAPR14]. AMF- [BGPPRW14]. AMLI [BMN10]. among [ZZ12]. Ampère [Nei14]. analogue [HCQ14, RT12]. analyses [NTPVNHLX14]. Analysis [ÁD15, BDD^{+}10, BP13, BL11, CR12, CZZ14, CD11, DC11, GH14a, HLWM10, IO14b, Jeo10, KO13, KLM14, MVMG12, MFL13, MR13, RSK14, WZ11b, YZ11b, ZY11b, eMA14, AMÇ13, AM10c, AC14a, AMR13, AR11, AAD13, BNY14, BAdS^{+}10, BAdS^{+}12, BGLV11, BM12, Cao13, Car11, CLB11, CLW14, CL13b, CSC10, ÇDE14, CC11c, CGS11, CF11, CRS12, CFL^{+}14, DCG^{+}14, DDV^{+}11, FD13, FM11b, FM11a, FGO12, FRXY13, Gau10, GUBS13, GCC14, GCY13, Gre10, GH10b, HKL11, HP10, HNPX11, HXSF14, HL11, HGW12, HS12, Hua14b, HYY14, JJL11, JPX11, KS12a, Kho11, KAS13, Kro10, KS12b, LF10, Li11a, LSXZ11, LLL12, LHLC13, LZ13a, LZ12a, LZ10, LFY11, LvDtTB^{+}13, dL14b, LO14b, MLZ14, MPS11, MJ14, OS10, PW12, PL10a, PTP10, RBK11, Ran14, RBD11, RBD14, RMTO11, RSGL14, Sez10, SRA10, SOL11]. analysis [SR13, SS14e, TL13, TS12, TR11, VMMM10, WYY^{+}12, WS13a, WW14, Wei11, WH12a, XL11, Xu11a, YWW10, YG14, YW10, ZH10a, yZ11c, ZY11a, ZBW11, ZCC^{+}13, ZLW14a, ZC10c, ZTP14, ÁD14]. Analytic [LVCC13, CFF12, Fat11b, MZ14, PS14a, Ste10b, ZFY11]. Analytical [CFdAR11, FFM^{+}11, KS12a, LTFL10, N’G10, Eba11, GXW14, KZ13, WA14, YFWH10]. Analytics [FdOdS12]. Andersen [TYCL14, Zha14c]. angle [FGEM10]. anisotropic [ANS11, BFB14, Bir11, Cao14, HK14, LL10, SSW14, Sun14, Wil13, YY11, ZC10c]. annealing [AG10a]. annuity [LMG13]. annular [ABHMW11]. Anonymous [APS14]. Ant [GPVL10, SCRB^{+}14]. antenna [ZDP^{+}14]. anti [HY11, lLfD10, SBM14, Spa12, WZ10b]. anti-counting [SBM14]. anti-Gaussian [Spa12]. anti-periodic [WZ10b]. anti-reflexive [HY11, lLfD10]. antireflective [DM12]. any [FZ11, bJfCwH14]. API [SPC14]. Apostol [Pré10]. apparent [DL14a]. Appell [CL10c, CL11f, ÖY14, YNO10]. Appl [ÁD15, Asl10, BAdS^{+}12, CDKN15, Dai11, DC10, DD14b, DS13c, Gün10, Kre11, LB10, LHL11, TSC10, XSV13, ZJ12, vDvFZ13]. applicability [AHK14b]. applicable [SB10]. Application [BGP14, CGS12, EDP14, FWBSBGA11, KKPV11, KWW11, LGMA13, NS10, NOS14, NR14b, PMC11, Ren10, SSW14, SBFC10, SH10, TKW12, AAM14, ASST10, AÖ11, AKT12, AKT14b, Bak10, BSZ13, BSKL12, CPP13, CR14, DH14, DJ10, GZ14a, GZ14b, Iid12, JGH^{+}13, LF10, LvDtTB^{+}13, MJ11, MHH^{+}14, N’G10, Prz13, SLP11, UBSG12, VDG11, WDG^{+}12, WHLC11, AY13, DGC^{+}13]. application-specific [MHH^{+}14]. Applications [GR13, Gra11, ÁECS12, ABG11, BBP14, BDJ^{+}11, BCM10, BHJ13, BJR^{+}12, BLTW10, CG11b, CGRS11, DCG^{+}14, FS12b, FB12, GW14a, HL10b, KT13, LLG14, LCI11, Ma10, Nai11, Nei14, NH10, PD13, PY10b, QLL11, RP11a, RWRS14, Ste10b, Van11a, XC14, YSZ11, YL11b, YSW13]. Applied [AGK^{+}14, AMT11, ADG11, ABÇ^{+}11, AR10b, GH14a, HVV11b, OCU11, Pit13, ZTL13, ZFH10]. Applying [BKV10, ILEE13]. approach [ABG11, AÖY10, Bal14, BW14, BS13b, BDV10, BDR10, Bir11, BL13a, BKN^{+}13, BNRZ12, BHPS12, CS12, CCS11, Che14b, CC11b, CL10c, DM10a, Dax10, DdS14, DDDSC14, DL14a, DN13, FN12, FJLL10, FM14, GK14, GG12, HM10, HHHS14, JKS13, KG14, Kho11, KS10b, KL11, KV12, Kuz14, LLG14, LZT13, LF11, LSXZ11, NF14, RP11a, RSS14, SK14, SRP^{+}13, TKW12, Wan12b, XJ11, Yan13, YK14]. approaches [XCY12, Yua12]. approximant [IA13, PGGCGF12]. approximants [CBGVPP11, PGGCGF14]. Approximate [CP13b, JKS13, KF11, YFWH10, ZFY11, yZ11c, ADRM12, AHH13, BH11, BDSG10, BGL13, BL13a, FPGMG11, HHM12, HR12, LPLX10, MZ14, Prz10, RSS14, WL11b, Xie11, ZYC13]. approximated [DM12]. Approximating [EL11, EALA11, FS12b, Hu12, Nar14, XSH10, XH13, HS11a, YKL11]. Approximation [BB14a, BTSA14, CGM10, Dra10, DGW10, FGPP13, GR13, KP14c, SS14b, AA12, Akt14a, AWF11, Ant11, BGIN13, BS13a, BBD^{+}11, BDL11, BSJ14, BSV10, BDR11, BR14c, BP13, BM12, CGCG^{+}13, Cao14, CW11b, Col13, CHZ10, CCZ12, CC12, DDV10, Dos14, Dra13, DLWZ14, DLP13, Fat13, Flo11, FV10, FN10, FSB10, Gat14, GY11, GSNA13, GG13b, HMA12, Hau10, HK14, HX13, HY11, JS10, JW12, Kho10, wKK12, KZ13, KL11, KLL11b, KLM14, KL10b, Lan12, LGMA13, LP12, Li14, LHC14, Lu11, Lu13b, MA13, MPS11, MU14, MV12a, MJ11, MD10b, MML11, NV11, NH10, NS14, Odi11, OA14, PRM11, Pré10, PL10c, Prz13, RST14, SSW14, San11, SBL12, Sla14, SS14c, SS14e, TW14, UM14, WWC14, WGZL14, WH10, WH13, YGHH14, ZZ11a, ZWM11, ZL13]. Approximations [FPP13, SWY14, AMPP11, Ain10, ÁD10a, AC14c, ACH11, BG10a, CL10b, CHH10, Dax10, GN11, HwK11, KSTW14, KSA14, KM10, Kuz13, LZ12a, MD10a, Naf14, Rie13, RVN13, SBGB10, VDG11, ZD11, ZLG12, dFGAN11]. Arbitrary [KSA14, BBKS10, CG11a, DMMP10, ESM13c, HMA12, Hu12, Lui10, PS11c, SSP13]. Arbitrary-level [KSA14]. arc [Nar14]. Archimedean [Nai11]. Architecting [FLKW14]. architecture [SP11]. architectures [DCG^{+}14]. arcs [CHZ10, HwK11]. area [AV10a, LL10, RP11b]. argument [AAY10, Fuk11, Fuk12, Fuk13a]. arguments [LL13b, Wan10, WZL11, Wan12a]. arid [WDG^{+}12]. arising [BM11a, GCH14, GKL12b, xLwCL10, LZL11a, MA13, MVMG12, SSF^{+}11, YLD13]. arithmetic [AÖÖ14, CMR11b]. arithmetics [RP11a]. Armijo [Zho13]. Arnoldi [BHJ11, BJRS12, BNRZ12, CFJ^{+}11, DJ10, GN14, WW10c]. Arnoldi-based [BHJ11]. Arnoldi-extrapolation [WW10c]. arrangement [SH12]. arrangements [WLZ11, WSJ14]. arrays [GMZ^{+}10]. arrivals [YWW10]. Arslan [HV13]. arteries [CPP13, JMS10a]. artificial [ARF14, CRF14, GLH12]. Asian [DDV10, DPV10, FPP13, HS11b, Mar11]. Aspects [Gra11, HDL^{+}13, LM12a]. assess [BDT^{+}13]. Assessing [LWH^{+}11]. Assessment [ZH10b]. Asset [JM12, HL14, LS10]. assets [LZ13e, RSK14]. assignment [YL14]. assisted [BPS12, KL12, LZWJ11, WCH11]. associated [ADGP11, ADLT13, BGVHN13, CLN14, CBGVPP11, DS10a, DHS11, HV11, HHZ11, HY11, Hua14b, dL14b, PMMS14, YGHH14, dAGR10]. associative [ZW11b]. assortment [KSK11]. assumptions [bJfCwH14]. asymmetric [LZT13, MB11]. Asymptotic [CC11c, DM10b, HGW13, Kuz13, LW14, LM13b, LP10b, Ste10a, Tan10, TS14, WX12, XhW11, AC14d, BV10, BCHL11, BK12, BE11, Che13a, yGZ13, KL14, LF11, Sch11, Tur12, ZyG14, Wan11c, WYO11]. asymptotically [CSY10, HZ11, KPW10, WZA10, XSH10, Xu11a]. Asymptotics [ADMF10, BCGMB10, KKK12, MGZ10, Par10]. atmosphere [MR10]. atmospheres [KCBA13]. atmospheric [BB10a, BB12b, FWBSBGA11, SH13a]. atom [MK10]. atomic [Ave10, CFGE10]. atoms [Bra14b, KS10a, TT14a]. attacks [Tez14]. Attracting [PR11]. attractor [YZ10b]. augmentation [Cao13]. Augmented [AD11b, BPS12, CZL14, CCS11, Che14b, CRF14, RMF11, ZHCW12]. AUSM [MDVB13]. authentication [NACLR12, SPLHCB14]. Author [Ano10a, Ano11a, Ano11b]. authors [BDR10]. auto [AMPP11]. auto-regressive [AMPP11]. autocorrelation [EI14]. automated [HTM14]. Automatic [GN14, RG12, EALA11, SL14a]. autonomous [KP14a]. autoregressive [BA10, ZWZ11]. auxiliary [DW10]. average [ADL^{+}10b, CMR11b, Li14, PVP13]. averaging [CN10]. AVF [Cie14]. avoid [TKW12]. axial [Jia10, KFR13]. axially [LL10]. Axis [NMBG14, KL11, OG10]. axisymmetric [CAT13, KD10, Li11a].

B [ARV10a, ABIS10, BLR10, EGES14, GW14d, He12, MMS13, MJG14, OGM14, RG11, Woź14, XhW11, XH13]. B-spline [EGES14, MMS13, RG11, Woź14, XH13]. B-splines [ABIS10, BLR10, GW14d, He12, MJG14, OGM14, XhW11]. backordering [TNN10]. backscattering [HW12]. Backward [XLCC10, AØ14a, BT12, Che10, Fad10, FLLT10, GP10, LPDN13, MSG11, Wan11b, Xu11b, ZFM14, ZRW11]. backward-in-time [GP10]. bacterial [COW10]. Bagley [CK14a]. balance [AMT11, BB10a, LMW^{+}11]. balanced [CBB13, HHM12, HGW13, PPG14, QW10, RZD14, YC11]. BAM [BV10, DL13, LF11]. Banach [But12, DW10, EGON10, EHRV13, GP12, pLjH10, jTW14, WZA10, WWL10, WSWD11, ZFG10, ZAWA13]. band [BGM10, DdS14, Gia12, GMZ^{+}10, KA10]. banded [CV10a, KS13]. Bandlimited [Gro13]. bandwidth [Iid12]. bankruptcy [WZ11b]. bar [GRV11, JL11]. barrier [EGBS12, ESC14a, HS10b, Hui10, Jun13, LTZ10, LL11c, MT10c, MY13, TTB14]. barriers [Jun13, MS13c, Xu11b]. barycentric [BK14c]. Barzilai [XSW12]. base [LLG14]. based [ADA13, AHS14, AZ11, AA14b, ASST10, AMR13, AM11, AA14c, AKT12, AKT14b, ACH11, AuIK14, BKGMA10, BCM10, BCR11, BDMP13, BHJ11, BL14, CSL^{+}11, Cai13, CCG^{+}13a, CCS11, CR14, ÇDE14, CDKN14, CDKN15, CT11, CRF14, CN10, DDV^{+}11, DBHV10, DS12a, DAE13, DKU14, EGBS12, EDAP10, EX14, FJLL10, FLHS^{+}13, FV13, GW14d, GWKE14, GWX10, HWG10, HCQ14, HZ14, HL10a, HH13, HH14, HLFP14, HLF14, HL13b, HK13b, HwKB14, ILMS10, IHRA11, ISN10, JKS13, KKR11, KNR12, KWW11, KMG12, Kuz10, Kuz14, KS12c, Lam11, LMT13, LWH^{+}11, LZ11, LSXZ11, LY13a, Liu11a, Liu12, LMG13, LZ13d, LP13, LG11, Lu11, LJ13, LXSC11, MS11a, MAK^{+}14, MMP11, MD10a, MVC10, MA11, NY12, NMS10, Nei14, NTNBV14, PS11a, PWZ11, PS11b, PS14b, PHS14, PO11, PZ13, QYZZ14, RMF11, RYM^{+}14, SCRB^{+}14, SPLHCB14]. based [SL14a, SXZ11, SHZ10, SGG11, SH12, uIAAF14, Śmi11, SACS10, VBP^{+}10, VL13, VB14, WLL^{+}11, WSK^{+}11, WLD14, Wih11, WSJ14, pWqSmZ14, XMSS11, XL12, YL11c, ZSK^{+}11, ZY14a, ZY14b, ZTL13, ZZBX14, ZCX14, ZZ11b, ZLG12, Zim13, BM11a, DHSW14]. Bases [GJ13, PS11a, PS13, BLM11, BwKB^{+}14, BGL13, DDM10, Woź13]. basic [MJ10a]. basing [Yan10b]. basis [AE11a, AM14a, AAD13, BLM11, BL10a, Boy10, CGCG^{+}13, DS13b, FND10, GK14, JGK14, KG13, KS14b, LYD13, MMS13, MM11b, NRS10, NH13, SW10b, SWY13]. Baskakov [SAB14]. basket [DDV10, DPV10, KZ13, SH13b]. batch [WYY^{+}12]. Bauer [SW12b]. Baum [MD10c]. Bayesian [DCGS13, LWH^{+}11, NA11]. BB [HHL12]. BCS [ZZ12]. be [Kni10]. beam [AC14a, BS11b, CF11, HKO11, JWY14]. BEC [KHFC11]. Beddington [DF14]. bee [GLH12]. behavior [CCDS14, Ery13a, GZL14, LL14b, Pag10, Ste10a, Tan10, Wan11c, Wan13]. behaviors [yGZ13, ZyG14]. behaviour [DM10b]. Bell [GGH^{+}14]. belonging [SS14b, SS14c]. Beltrami [Dou12]. BEM [CR12, HLSX11]. Ben [Ano10b]. Bénard [GD13]. benchmark [MKS12]. bending [LGM14]. beneath [BKMG11]. benefits [JWY11]. Bermudan [LLK14]. Bernoulli [BE11, GQ14b, Khr10, MM10, Pré10]. Bernstein [Beh14, HCQ14, LK13, Spe11, WL10d, YBD11]. Berrut [BDH11, Zha14b].

Bessel [Che13a, CA14, HS10c, Vel11, XXH14]. best [BISS10, GLH12, HwK11, Nie13b, Van10]. Beta [CZ14, FPG12, ÖÖA11, TT11]. Better [BLM11, DDG^{+}13b]. between [AM10b, CDQ14, CFdAR11, DKU14, EY11, RSK14, RVPOR^{+}11, Zha10a, ZY10, ZY11a, ZW11a]. Bézier [AHR14, ALM11, AM11, CYZ^{+}11, DP10a, DP11, HCQ14, Hu12, HX13, Lu11, Lu13a, SR11, Spe11, WM12a, WZLP11, WL10d, HwK11, Lu14]. Bézout [Bel13]. BFGS [VL13]. Bi [Jia10, FGO11, Jia11]. Bi-frames [Jia10, Jia11]. bi-infinite [FGO11]. biarcs [BBK^{+}14]. bias [MD13]. bibliometric [AB14]. BiCG [AS10a]. BiCGStab [Yeu14]. BiCGStabt [Yeu14]. BiCR [AS10a]. bicriterion [EC13]. bidiagonalization [NY14]. bidirectional [ZW11b]. BIE [Fat11b, MS13b, SSP13]. Bifurcation [Gre10, KAS13, AMÇ13, BSS13, CYJS14, CHRZ13, CGS13, CSZ11b, DL13, SLD13, WW11b]. bifurcations [BFG^{+}14]. big [LW10a]. biharmonic [ALM11, Lam11, LLCL11, RCM^{+}12, Seg12]. bilateral [YYL^{+}10]. bilevel [CG10a, CG12, LCW10]. bilinear [Bay14, LZ12b, LY13a]. Binary [CCZ12, ARF14, EY11, SA14]. Bingham [DA10]. binomial [KPsQ11, MY13, TT14b, YE14]. binomials [Tyr10]. bio [COW10]. bio-clogging [COW10]. biological [ACF^{+}14, JPX11, RSFS14]. bioluminescence [GCH14]. biomagnetic [TTSB14]. biomedical [KWXZ14]. bioremediation [COW10]. Biorthogonal [BGGG^{+}11, Sid14, CCCM13, Lam11]. biparametric [DP14c, GK11]. bipartite [BFS14]. biperiodic [Tau13]. Biquadratic [YL11c, DCJ13a, WG10b]. biquintic [CW11a, CW12, CW14a]. Birkhoff [NBBKV10, CLLX11, DAE13, LCXL11, NTNBV14]. Birkhoff-type [DAE13].

Birthday [ARV10a, MYZ10]. bisection [HKK10, XMC11]. bits [Tez14]. bivariate [BGIN10, BK13, BSZ13, CDD12, GJ13, KG14, LC11, WZD10, ZL13, YG13]. Black [MKK14, CL11a]. bladed [ZTL13]. blended [YL11a]. blending [DS10a]. blind [FN12, SZJ12]. Bloch [Tau13, YLTB14]. Block [BHJ11, CZ12b, DMB10, FJP14, BMMR11, BY12, Beh14, BNY14, Bog14a, Cao13, CLS14, CV10a, DDDSC14, Dos14, DSY^{+}11, HL13b, Kok11, KMG12, Liu11a, LCI11, MMS10, MZL14, NL10, QY10, SBGB10, WH12a, WH12b, WHL13, WL14a, WT14, ZYW14]. block-circulant [SBGB10]. block-grid [Dos14]. blocked [ZGZW10]. blocks [Bai13, CV10a]. blood [BDT^{+}13, CPP13, JMS10a]. blossoming [AHSN13, ILMS10, LMT13]. Blow [AS10b, FDH10, ZLC10]. Blow-up [AS10b, FDH10, ZLC10]. blowflies [Alz10, CL11c, DN13]. blur [JMV13]. Board [Ano11c]. body [CDK10, MT14b, SSP13, SCRG12, WLL^{+}11]. Böhmer [DS13a]. Boltzmann [BCD^{+}10, HBE14]. bonds [EF14, ZYL11]. bone [FM11b, MP10b]. Boolean [LCN12, UÇD14, XCZ^{+}11]. boosted [LMM^{+}10b, MM11c]. bootstrapping [GS14, HHHS14]. Borwein [XSW12]. bottled [SCRB^{+}14]. bottled-products [SCRB^{+}14]. bound [AG10a, Che13b, DdS14, Gün10, LTW10, LLX14, Wan11d, YSQ11, Zha14b]. bound-constrained [DdS14]. boundaries [BDSG10, DDV^{+}11, LDGY10]. Boundary [CLS12, Fat13, Gus14, MMMG10, NSZ10, Tar10, Wan10, XZ14, Zha11a, ADG11, AU14b, AM11, Bab10, BG11, BDL10, Bar11, BDSG10, BK10, BB14b, Cai10a, CDTV11, CS10, CZ12b, DHR12, Dea11, DM12, DR14b, Dou12, DJ14, Eba11, ESM13a, Fat11a, For10, GC11, Gen12, HWG10, HHH^{+}13, HS10a, HBE14, HLFP14, HVV11b, HTH10, IO14a, Jeo10, JX10, JJE13, JTV14, KLK13, KP11b, KLM14, Kur10, LL11a, LS11, LW13, LM12b, LZ12a, LL13b, LB11, LC10, MY11, MV14, MTJ11, ML10, Mui10, NS10, PTH10, PW12, PT12, PP10a, QZX^{+}11, RG11, RE10, RU13, Rua13, RR10, SD10, SZ10, SZS11, SP11, Tau13, VS13b, VT14, WZ10b, qW11, Wan12a, WWG10, WZ10d, YC10, ZSD10, ZFG10, ZD10, ZZBX14, AGM12]. boundary-layer [VT14]. boundary-value [KP11b, SP11]. bounded [RRMD13, TYCL14, Uty10a]. boundedness [bJfCwH14, LM13b]. bounds [Ain10, FW10, HC14, KLNS11, LL11d, MA11, PS14a, PR12b, SL12, SP10, Yan10b]. Boussinesq [ZBW11]. box [KP11a, ZWG11]. box-splines [KP11a]. brain [LP10a, RPS^{+}11, VBP^{+}10]. Bratu [Moh13]. break [EDAP10]. breakdown [WYP10]. breaking [BSS13, HdJL^{+}13, SKAM11]. Bregman [BL13b, KO13]. bridge [AHM13]. brief [NSZ10]. Brownian [ÁE10, ÁE11b, BB11a, VAZS13, YÖVAYOW14]. Broyden [CC13b]. BS [MS12]. BSDEs [FJ10, RE10]. Bubble [Bay14, DBHV10, Sta11, GGS12]. bubbles [DGC^{+}13, Naf14]. Buchberger [WZD10]. buckles [WL11a]. build [DdS14]. built [KW10]. built-in [KW10]. Burgers [AF10, YK11, YK14]. Burnside [Qi14]. BVMs [BIT11]. BVPs [ABHMW11, FJ14]. BWENO [CCZ12].

C [SKK14]. cable [CLB11]. Cahn [ZYC13]. Calculating [YAKA11, BGRS12]. calculation [Hos11, WL11b]. calculations [Bra12, Bra14b, GKL12b, HJ11, HZ14]. calculus [MMF10]. calibration [BDMP13, GS14, LS10, MRAM10, MD10c]. call [FS10]. CAM [Wan11c, WYO11, BBD^{+}11]. Camassa [CLS10, DI12, FiMO10, Mat10]. can [Kni10, Prz10, RKS^{+}13]. cancer [GPVL10]. canonical [ZZG10, ZWG11]. capable [NO12]. capillary [AJK12, MS13c, PC12b]. capital [mZkPkL11]. caps [FTTT13]. Caputo [JD13]. carbon [SRA10]. carbonation [RMP^{+}13]. cardinal [LD10]. Carlitz [LW14]. Carlo [GPVL10, KZ13, LM13a]. carriage [LGR11]. carrier [ZD10]. Cartesian [BP13]. CAS [RLGGGÁM14]. Cascadic [RS10b]. case [AKT12, DM10b, FKL10, LP10b, PC12b]. cases [CW14a]. catastrophes [BGS12]. categorical [PM11a]. Cauchy [ELA10, EL11, EALA11, KKPV11, LWY10, PM13, QXW10, RBD11, RBD14, WX10, Xio10, XZH14, ZMZ13, ZW10]. Cauchy-type [EALA11]. caused [DGI^{+}10]. cavitating [Gas14a]. cavitation [LL11b]. cavity [DA14, DS11]. CCA [AHS14]. CCR [LJE^{+}10]. CDS [GUO14, Nai11]. cell [AV10a, ADL^{+}10b, CHMD14, Jeo10, RSGL14, Sun14, SS14e]. cell-average [ADL^{+}10b]. cell-centered [RSGL14]. cellular [AAY10]. cement [KBKR14, VB14]. cement-based [VB14]. Censored [Hos11, Ism14, LWH^{+}11]. censoring [AHAH11, ADA13, BP11a, Fer14, WSS14]. centered [HHLW14, RSGL14]. centerless [HV11]. centers [Bra12]. central [CBB13, Li12]. centralities [Rom11]. centrality [ABN10, CFdAR11]. Centralized [LNJ^{+}10, LNB^{+}12, NACLR12]. certain [AP14b, BGM10, Dra10, Dra13, Gat14, GSK11, LQSL11, SP10, Ste10b, WX12]. certified [GCC14]. Ceságro [AB11a]. CEV [LRZ14, TTB14]. CG [AK14d, DW11]. CG1 [BKB14]. CGSTAB [TS10]. chain [HXSF14, SB10]. challenges [GKB10]. Chandra [El 11]. Changeable [SW10b, SWY13]. changeover [YTFP11]. changes [KBKR14, LS10, RMTO11]. changing [KC10, MY11, NV11, ZC10b]. channel [Gas14a, HL13a, PKH14, TTSB14]. channelized [GW14a]. Chaotic [HL10c, MT14a, WC10]. Chaotification [XHZ12]. characterisation [Kro10]. Characteristic [HBE14, AÖÖ14, CFL^{+}14, Fu10, HZ14, LSW11, WM12b]. characteristic-based [HZ14]. characteristics [BB11b, ORP14]. Characterization [Ant11, LS10, Ars11, Ars14, BCR13, CG11b, GRZP10, SCRB^{+}14]. characterization-based [SCRB^{+}14]. characterize [SK14]. Characterizing [BAdS^{+}10, BAdS^{+}12]. Chebyshev [AHSN13, AEG^{+}11, BDD^{+}10, Boy13, DDM10, KXH13, KHKM14, LD10, Mui10, Not14, OP12, SWY14, WC11, Zha10b]. Chebyshev-like [BDD^{+}10]. Checking [LLW14]. chemical [BKN^{+}13, DGI^{+}10, WC10]. chemistry [GZ11]. chemotaxis [ABS11b, SSK^{+}13]. CHIMERE [GVM^{+}12]. Chinese [KS14a]. Chirp [TT14a]. Chirp-wave [TT14a]. Chlodovsky [TT11]. Chlodowsky [BTSA14]. Cho [SPLHCB14]. Choice [ZW14, BFLP12, HPR12, MT14b, PW13, PP14, ZFM14]. choices [CJY10]. Cholesky [FV13, Liu11a]. choosing [Wan12b]. Christoffel [Ste10a]. chronic [ÇDE14]. cipher [BNY14]. circle [CBGVPP10, CBGVPP11, GHM10, HV14]. Circles [Kim12]. circulant [SBGB10]. circulants [ENSCS12]. circular [HV11, HwK11, LDGY10, PAD10]. circumferential [AC14a]. city [GGAVRC14]. claim [AAK10, Sch14d, TT14b, ZY11a, ZwGhX14]. claims [CP13a, GW14c, LB14, hXZ11, ZZ13]. clamped [BySZZ13, BDyS14]. class [AM14a, ÁECS12, ADL^{+}10b, AP14b, BH11, BSS14, CS12, CL11c, CR14, CM10d, DH11, DAE13, DXL10, DN13, Dom14, EZ14, FBB14, FP13a, FKL10, GW14b, GPS13, JY13, JTV14, KP14a, KC10, Li14, LL13b, LZ14b, LP13, LYD13, LY13b, pLjH10, Mah14, MJ10a, MA11, ML10, NLO13, OG10, PW12, PZ13, SS14b, SS14c, SS14d, SZ11, WCW10, WS13a, YL11a, YKL10, YD11, mZkPkL11, ZW12]. classes [Hua14a, JZ10]. classical [BR11a, CG11b, FPP12, HXSF14, KMLT10, MPS10, MBS10]. classification [LMM^{+}10a, MVC10, MSP12, QTS14]. classifiers [ZTS11]. Clenshaw [Dom14]. climate [BBD^{+}11, FLEa10]. clinical [GPVL10]. clogging [COW10]. close [KG12, SH12, WSJ14]. close-packed [SH12]. close-packing [KG12]. Closed [JR10b, SRWZ14, BC12, BG10b, FB12, MT12, Owa12]. Closed-form [SRWZ14, Owa12]. closer [Yan10b]. closure [CR14]. cloud [Bra14a]. clustering [FV13]. CMARS [ABÇ^{+}11, YÖVAYOW14]. CMMSE [AVA13]. CMMSE10 [GVM^{+}12]. coarse [CDKN14, CDKN15, DZW14, XN14]. coast [MKK14]. coastal [WDG^{+}12]. coating [KFR13]. coaxial [ZZC11]. code [AS12a]. codes [CGRLR11, GM14b]. codifferential [TBK14]. codimension [DD14b, DD14a]. Coefficient [MM10, CCG^{+}13a, DY11, LD10, LPDN13, LBM12, Str12, VS13a, ZWZ11]. coefficients [BY14, BBD13, BVV10, CG12, CK11, DC10, DCZ10, DXL10, GG10, HKL11, KH10, LVCC13, MS13a, NF14, NV11, PVP13, SH11, YSW12, Zha10a]. cognitive [MHH^{+}14]. Cohen [LF11]. coherent [GJCL11, dJP13, dJMPPC14]. cohesive [BDMP13]. coincidence [SR11, WZLP11]. collections [KS14b]. collision [GÖ14b]. Collocation [ASS12, AC14a, BB11c, Cai10a, CA12, CS14, CT11, CDJP13, DMR13, EDP14, FLHL14, GPHAM10, GPHAS14, IO14a, IR11, JM13, KH11, LZDW14, OMS10, PDP13, PT11a, PT11b, PT12, PT14, Pul14, RG11, RD10, WCW10, WZF11, WW14, YZ11b]. colony [GLH12, GPVL10, SCRB^{+}14]. colour [RMTO11]. combination [Ger12, WUB10]. combinations [AMPR10]. Combined [LLS14, MHL11, BDR10, LJE^{+}10, LYD13, Sha14, SM14b, VdKPB10]. combined-oriented [LJE^{+}10]. Combining [CWQ11, HR12, MR11]. Comment [Asl10, KRS11, NK10, TSC10]. Comments [ZJ12]. Common [Büy12, ĆRR^{+}10, PD13, CKM14, HL10b, WL11b, XSH10]. commutative [KB11]. Comonotonic [VDG11]. comonotonicity [LMG13]. comonotonicity-based [LMG13]. Comp [BAdS^{+}12]. compact [CCG13b, Cui10, Cui14b, DF12, DFH14, LLS14, Lia14, MZ14, PN11, PLN13, SBGB10, Wan11e]. compactly [CCCM13, CCH10]. companion [DDM11]. company [HD14]. Comparative [LT13]. Comparison [AM10b, BCM10, BL10b, GSK11, KL14, Li10, SH13a, WYP10, ZYC13, BGPPRW14, DD14b, DD14a, JK13, LEKT10, Mui10, NK10, PM11a, Seg12, yS05]. Comparisons [CDQ14, CPY10, LO14b]. competing [WSS14]. competition [AALMLM14, GLLM12, LZT13]. competitive [Emm11]. complement [Yan10b]. complementarity [CPY10, DH11, EK12, HCWsO13, KC11, LJCL11, RX10, SZ11, ZW12, ZY14b]. complementary [AW10]. Complete [KMS14, QGG10, Fuk13a, Fuk13b, HWG10, HD14, HHT12, LvDtTB^{+}13, NMBG14, Qi14, RVPOR^{+}11]. completely [AAK10]. completeness [MJG14]. completion [NGAZ11, ZC10a]. Complex [ASM13, AV10a, CRS12, DH14, DJ14, FV14, HC14, KS14c, LK13, MMX10, Rom11, TM11, WH12a, ZW11b]. complex-valued [ZW11b]. Complexity [PHS14, AÖÖ14, BKTY14, Dra10, EI14, SXZ11, SZJ12, YLD13]. compliance [FM11a]. component [ABN10, PGGCGF12, ZBZ10]. Componentwise [Hua14b]. composite [BHMW13, GH14b, HL10b, OAHR11, VB14, WDZ10]. composites [PSKA14]. composition [BDMR10, EO14, LZSL11]. compound [CP13a, CO13, GL10, LL11c, TT14b, hXZ11, ZYL10]. Comprehensive [HwKB14]. compressed [CSSX14a, CSSX14b]. compressible [BSKL12, FHPH^{+}13, GKT10, GKT11, HJ11, HHM12, PKH14, XZ14]. compression [AMR13, LFWL14, WSK^{+}11, YSZ11]. compressive [ZSK^{+}11]. comprising [CH11]. Comput [ÁD15, Asl10, CDKN15, Dai11, DC10, DD14b, DS13c, Gün10, Kre11, LB10, LHL11, TSC10, XSV13, ZJ12, vDvFZ13]. Computable [Hu14]. Computation [BJR^{+}12, GZ13, KXH13, PVP13, PR12b, SC10a, Xie12, BFG^{+}14, CM10a, DZW14, DC11, Ehr11a, EDAP10, FGO11, Fuk11, Fuk12, Fuk13a, Fuk13b, Fuk13d, LCN12, LEVG14, Oou13, SST12, SC10d, WZ10c, qWhY11, WLH12, WH13, Win14, ZY11b]. Computational [AGK^{+}14, AELH10, CM13, DST10, FHMZ10, GR12, HDL^{+}13, LM12a, ORP14, SKKS12, TT14b, VAV11, WK10, AM10b, And10, BKTY14, BDT^{+}13, BL11, DCG^{+}14, EX14, FN12, FM11a, GCNL14, GKB10, GSGN11b, LM12b, LC10, NCCF^{+}10, RRMD13, AVA13]. Computations [OP12, ADGP13, ÁD14, ÁD15, BG11, DR14c, FLHS^{+}13, GH10a, GGS12, JY13, RVN13, SGG11]. compute [DL14a, VVV10, pWqSmZ14]. computer [DSL12, KL12, WCH11]. computers [DHSW14]. Computing [Bel13, BSV12, Ery14, Fer14, FV14, FS12a, HHH^{+}13, KMG12, Liu12, LHF11, Tau13, Van10, ZH11, AMMS10, CKM14, Cin12, Cin14, Gia12, GQ14a, HHHS14, Hua14c, IA13, LHY10, LZ13c, MMSW13, NY14, PR12a, PS11b, PS14b, RB13, SS14d, SP11, WZ12, WW10c, WM12b, XCZ^{+}11, Zho12, ZLG12, ZZG14]. COMSOL [YK11]. concatenated [HSY13]. concentrated [Cha11a]. concentration [RMP^{+}13]. concentration-dependent [RMP^{+}13]. concentrations [DGI^{+}10]. concept [AJK12]. concepts [SL14b]. concerning [HD11]. concomitants [YG13, YG14]. concrete [RMP^{+}13, VSB12, XSV12, XSV13]. Condition [SXZ11, AH12a, CLW14, CLN14, CXZ14, Fer11, FGO12, Gün10, KNY10, RE10, TA13, YC10, ZSD10, ZD10, ZZBX14]. Conditional [BEJ14, BK13]. conditionally [PS13]. conditioned [BNRZ12]. Conditioning [ADGP13, TKW12]. Conditions [WZLP11, ABCP14, AS10b, Bab10, BY11, BBB10, BDL10, BDSG10, BB14b, CDTV11, CR11b, Dea11, DM12, DR14b, HWG10, HBE14, HLFP14, HM11, KP11b, Kur10, LL11a, LB11, MT10b, MV14, NY12, NS10, PD13, RA10, SR11, SZ10, Sza10, Tar10, ZFG10, Zha11c]. conducting [ATS14, CR12]. conduction [LZ14a, MVMG12, SHZ12, Sza10, WJJ14]. conduction-convection [SHZ12]. conductivity [DS12b, DS13c, SRP^{+}13]. conductor [MA13]. conductor-like [MA13]. cone [SL14b]. cone-invexity [SL14b]. cones [HHZ11]. Conference [SKS14, LBCL10]. confidence [Fer14]. configuration [BDT^{+}13, MT14b]. Configurations [CG10b]. confined [XZ14]. confluent [LP10c]. conformal [HQR13, LDGY10]. conforming [CHX11, CW14b, FL10a, HKK10, HP10, LVD^{+}14]. congruential [Wik10]. conic [CWQ11, Hu12, JLZZ10]. conics [CW13]. conjecture [HX14a]. Conjugate [HQR13, NY12, And10, And13, BKGMA10, BK14a, CXH09, HK12, JW12, JWW14, KNY10, LP13, SCFR14, SS14b, XSW12, YWL14, ZJ12, ZZG14]. connected [ABN10, LDGY10, Vit10, ZMZ13]. connection [CK11, FMP14, MBS10]. connectivity [ASG11]. conservation [AGJ13, BGP11, HK13a, KH11, Kuz14, MSZ13, SR13, YNO10]. Conservative [MM12b, NNO12, YMS10, BK14b, BI12, BGP14, GVM^{+}12, IG12, RVDM10, ZBW11]. conserved [ZML14]. conserving [Mat10, RG12]. considering [SSW14]. consistency [Fad10, SLW14]. consistent [ADLS14, KW10, Zha14a]. constancy [WYO11]. constant [AAY10, AAY11, BDH11, ÇU14, GL10, KN10a, ZSD10, Zha14b]. constants [AC14c, PS14c]. constitutive [KP10a, LP10a]. Constrained [HX13, WL10d, AG10a, AG10b, CDV10, DdS14, EG11, Hu12, JW10b, KA14, Kuz12, LPGX10, Lu13a, MRS10, MPC10, MT14b, Owa12, Ren10, YSQ11]. Constraint [BM11a, ZS13, CDPP10, yP10b, SY11, TNN10, YGHH14]. constraints [AHR14, CPP11, HHX14, LZL11b, LQSL11, LLW14, Mol10, Naz11, PNCC10, QW10]. Constructing [CZ14, CGK10, ML13, ZCS11, ABIS10, KKK14, MJ10b, Ran14]. Construction [BCR13, BDR11, GZ14a, GZ14b, ILMS10, IJ14, JR10a, KPBN13, KPV14, LMT13, PP10b, Woź13, Woź14, YSZ11, CCCM13, CGS13, CM10c, DDG13a, EO14, HP13, LO14b, SA14]. constructions [Esc14b, LRT13]. Constructive [FO10]. contact [CF11, DKH^{+}10, FM11b, FM11a, GPKV10, Gwi13, HKL11, JMS10b, Kok11, NRP13, QYZZ14, Ren10, SBFC10, WLL^{+}11]. contact-impact [GPKV10]. container [HTM14]. containing [BSJ14]. contest [AALMLM14]. context [RVN13]. contingent [Sch14d]. contingent-claim [Sch14d]. continuation [BKV10, CCJC13, CFF12, GP12, JKS13, LG11, WC11, ZFY11, ZW12]. Continuity [Jun13, ZZ11a, CP13b, WZY^{+}14]. Continuous [BY12, BÜ14, BM12, CSY10, CLW14, CL14a, EZ10, FJ10, GP12, LwCLW13, LYP13, MS13a, Min10, PS11c, PMMS14, Sch14d, Tan14b, WZL11, Wih11, Xu11b, YZ11b, Zha10a, ZM14, dO10]. Continuous-time [BY12, BÜ14, LwCLW13, Sch14d, dO10]. continuum [Mee12, MP10b]. contour [DL14a]. contraction [JK10b]. contractions [KPW10]. Contractivity [PAJ10]. contrast [AU14a, WB10, Wil13]. contrasts [BG10a]. contribution [Mor10a]. control [AA14a, AØ14a, AK14a, AVMVMV10, AP14a, AB13, BBP14, BK14b, BCR11, BAdS^{+}10, BAdS^{+}12, CK13, CY11, CL11e, DR14a, EG11, Fu10, GPVL10, GY11, GPHAS14, HKO11, KFR13, KW11, LYD13, MX10, MMP11, PL11, RA10, SYK14, Tan14b, TW14, Uty10b, Uty10a, WK14, XCY10, XL11, YK11, YK14, ZSG^{+}14, ZTL13, Zhu14]. control-type [RA10]. Controllability [RHS11, AM10b, CD11, MZ14, Mön10]. controlled [NF14]. controlling [DM10a, YWW10]. convection [AK14a, ANS11, BS13a, BB11b, BKB14, CL11b, CGS11, Cui14b, ESV10, Fu10, GD13, GTS14, HS12, Jeo10, KMHM14, LE10, MS11a, SYK14, SHZ12, SBGB10, TY11, WZ10c, Wan11f, XL11]. Convective [KD10]. Convergence [ÁECS12, AD11b, BBMY11, BJ12, CA12, CLW14, DS12b, DS13c, FMP14, FPP14, GCC14, HP10, JW10a, JKV11, KPKK12, LZ12a, LZ10, LMYJ12, Lui10, MLZ14, PM11b, Sha14, TL13, Wei11, ZHCW12, ZY14b, ZC10c, AMMS10, AMMS12, AMMS14, AB11a, Akt14a, ÁD14, ÁD15, AR11, AEG^{+}11, AH12a, AH13, AK14d, BY11, BSS14, BBS14, BT12, CB13, CPP11, CHMT10b, CTV11, CTV13, Fer11, FGO12, GKLZ12, GSNA13, GSNDB14, GP12, Gwi13, HML14, HJS14, HXW11, JK10b, Jia12, JW10b, KH10, Kuz12, Lan12, LT11, LW11, MS13a, MNQ14, Mel10, MK11, PT11a, PS11c, PC12b, QCK10, RA10, Rie13, RT12, Sch11, SYİ11, SLP11, SSR10, WW14, Wik10, pYXPH10, Yan13, YZ10b, ZLW14a, Zho13, ZC13]. Convergent [FKL13, Bog11, CG13, GK11, bJfCwH14, LZI10, MRD13, Mir14, RK11, YPL10]. Convex [CCG12, BGRS12, BEJ14, CL11e, GWKE14, HHX14, LMCW10, LV14, LZ13c, LPLX10, LPGX10, LCW10, SY11, YWL14, ZC13]. Convexity [Zha11c, VCA14]. Convolution [OSR13, BGRS12, HLFP14, Kho10, LP10b, MR12]. convolutional [RST14]. convolutions [KLL11b]. Cooperative [GPO14]. coordinate [Car11, CDS10, ZTC14]. coordinates [DGG12, PRT10, RMTO11, ZD10]. COPD [ÇDE14]. copolymer [XZ14]. copulas [Nai11]. cord [WL11a]. core [SKAM11]. corners [AV10b]. Cornu [CC12, Lu13b, Nar14]. corrected [KG12, VS14]. correction [BL14, Jun13, KPKK12, Kuz12, fLyHZ10, RVDM10, SHZ12, WZ14]. corrections [ÁE10, XN14]. corrector [FF14, KKR11, LXY13, TLNE10]. correlated [GD10]. correlation [CW14c, DCGS13, LS14, VAZS13]. correspondence [ZW11a]. Corrigendum [ÁD15, BAdS^{+}12, Dai11, DD14b, Kre11, vDvFZ13]. corrupted [JWW14]. cosine [Boy11]. cost [DCG^{+}14]. cost-effectiveness [DCG^{+}14]. costs [AV12]. Cotes [LZY10, LWY10]. Coulomb [Bra12, HKL11, PTP10]. countably [WWG10]. countercurrent [EASS13]. counterpart [GGSP10]. Counting [UÇD14, SBM14]. coupled [AF10, AC14d, AK11, BBS14, CHX11, CW14b, CyY10, CyYwY11, FP13b, HLSX11, ILEE13, Jan14, KMD12, NRS10, Naf14, NTPVNHLX14, RK11, SB10, SRP^{+}13, WC10, Wan11d, YMB14, ZZ12, ZY11b, ZX10]. Coupling [FND10, CMR11a, KP14b, MDVB13, PDE^{+}13, SKAM11]. coupon [ZYL11]. Cournot [GLLM12]. covered [DS11]. covolume [YJ11]. Cox [WW12, XYW14]. Coxeter [BFS14]. CP [LG11]. CQP [ABÇ^{+}11]. crack [AC14a, MMMG10]. cracks [AV10b]. Cramér [Zhu14]. cranes [HTM14]. Crank [CS10, JK10c]. credibility [Pit13, ZZC10]. credit [ABÇ^{+}11, FM14, Zhu14]. criteria [AK14b, BDMP13, Li10, LB09, LB10, LGZ10, MAK^{+}14, ZCH10, ZTS11]. criterion [AEGB10, BFLP12, Cha11a, Gos11, Hu14, Rah14]. Criterions [LWZ^{+}14]. critical [BL13a, CHRZ13, GD13, QLL11]. cross [AB10, FV10, HHM12, LdHA10]. cross-grid [AB10]. cross-section [HHM12]. cross-validation [LdHA10]. Crouzeix [BMN10]. Cryptanalysis [SPLHCB14, BKTY14]. cryptographically [SA14]. cryptography [AA14b]. cryptosystem [AK14c]. crystals [CvdMS12, FGEM10, Gia12, HHLW14]. CSCS [ZZ11b]. CSCS-based [ZZ11b]. CSIE [De 14]. Cubature [DS10a, GSV11, ML13]. cube [RP11b, XhW11]. cubed [YC11]. cubed-sphere [YC11]. Cubic [GW10, BGIN10, BBK^{+}14, CHRZ13, DM10a, EK11, HHLW14, HwK11, JŽ11, wKK12, Koç11, KPR11, KRS11, MMS13, SR11, SH12, VCA14, WM12a, WZLP11, Zha10b]. cubics [Str11]. CUDA [SP11]. cue [LZSL11]. curl [BL10a]. Currency [LWN13]. current [CR12]. Curtis [Dom14]. Curve [CM10c, CSL^{+}11, Flo11, Jia11, KPR11, LZ13b, LHF11, ZM13]. curved [AC14a]. curves [AHR14, AHM14, Alc14, BL13a, BR13, CW11b, CGS13, CZ14, CK14b, GM14b, HCQ14, Hu12, HX13, HwKB14, JR10b, JR14, Kır11, KLK13, Lu11, Lu13a, Lu14, Nie13b, PV10, RSS14, SF12, SR11, WZLP11, XH13, YWY^{+}14]. curvilinear [LMCW10]. customers [WK10]. cut [WW10a]. cutting [KSK11]. cycles [CGS13, PR11]. cyclic [XSW12]. cylinder [BKMG11, GPKV10]. cylinders [Nie13a]. cylindrical [DGG12, MKD12, PAD10, PRT10]. cytoplasmatic [CHMD14].

[YZ10b, AHLJ14, ADL^{+}10a, Bab10, BG11, BSPD14, BY14, BCD^{+}10, BKB14, BJ12, BHMW13, CZZ14, CPP13, CLP10, CDQ14, Col13, CvdMS12, DLG14, DDG^{+}13b, FGEM10, GCY13, HKL11, JMS10a, JJE13, JMS10b, KSA14, LEKT10, LWXW12, Lia14, MS13b, NTPVNHLX14, OS10, PDE^{+}13, PRT10, QYZZ14, QZX^{+}11, Rah14, SBFC10, SSK^{+}13, TY11, UBSG12, ZD10]. DAEs [GPHAM10, SL14a]. Dai [And10, XWW10]. Dam [EDAP10]. damage [LCCK13, MP10b]. damage/repair [MP10b]. damped [CYJS14, GM13, QC14, RTL14, Wan12b, YL14]. damping [HVV^{+}11a, JK10c, fLyHZ10, Zha11b]. Darbo [AAM14]. Darcy [CHX11, CW14b, EZ10, LTX12, Naf14]. Darcy-flux [EZ10]. data [AB11b, AGT12, ÁAG^{+}13, AC14c, BR13, CCG^{+}13a, CD10, CGRS11, DDG^{+}13b, FPP14, FZ11, GW12, GW14d, GPO14, IO14b, Ism14, Kho11, Kni10, KLK13, KL10b, LNB^{+}12, MM10, MD10a, MM11c, Nie13a, PM11a, RBD11, RBD14, RVN13, SW10a, SGZ10, ZL13]. data-sparse [RVN13]. databases [PVM^{+}14]. dates [CMR11b]. Daubechies [PM13]. DC [TBK14]. DEA [LJMG11]. DEA-efficient [LJMG11]. deal [RKS^{+}13]. DeAngelis [DF14]. death [JWY11]. debit [Zhu14]. deblurring [XCQ14, ZZBX14]. debt [ÇU14]. decay [PJ12]. decision [AÖY10, FJLL10, KWW11]. deck [AHM13]. decoding [ZSK^{+}11]. decomposed [GLWF13]. decomposing [Man11]. Decomposition [FLLT10, SB10, BR14a, BKN^{+}13, BL13b, CY11, CFH14, Eba11, FV13, GNN11, HL13b, HK13b, KNR12, KR14, KMP14, LPLX10, LPGX10, MS11a, Mah14, MPS10, MMX10, PL10a, PAJ10, QW10, RK11, RSFS14, SLP11, Ste10b, SM10, SM14b, XN14, XCQ14, Yan10a, Zim13]. decomposition-based [HL13b]. decompositions [HR10]. deconvolution [DM12, FN12, HLF14, MM11c, ZZBX14]. decoupled [CM12]. decreasing [Zha10b]. dedicated [LLG14, MYZ10]. Dedication [ARV10a]. deductibles [HW10]. Default [HD14, ABÇ^{+}11, SH13b, YYL^{+}10, Van11b]. Defect [AKT12, AKT14b, SHZ12]. Defect-based [AKT12, AKT14b]. defect-correction [SHZ12]. defective [FPG12, San11, yZ11c]. defects [CG10b]. deferred [BL14]. defined [DHS11, GCC14, Mia10]. definite [AGT13, BLTW10, CJ10, Che13b, EGST11, GT09, LHL11, PS13, ZZ12]. definiteness [LWZ^{+}14]. definition [KAYS14]. Deflated [NL10, GZ13]. deflation [pWqSmZ14]. deformation [VBP^{+}10]. Degasperis [MM12b]. degenerate [AST14, ABS11b, Cha11b, DY11, FDH10, FN10, Naz11, YP10a]. degenerative [CL14b]. degraded [JMV13]. degree [BC12, DDM11, FZ11, HMA12, Hu12, IA13, JJK^{+}13, Kır11, KPV14, Lu13a, ML13, SW10b, SWY13, WL11b, WL10d]. degrees [PGGCGF14]. Dehesa [MYZ10]. Delay [HGW12, RBK11, AØ14a, AAY11, AMÇ13, Alz10, Bar11, CZ13, CCDS14, DBZ10, GXW14, GZL14, GU10, Gre10, Hua14a, KS12a, LF10, xLwCL10, LSXZ11, LB09, LW10b, LB10, LYZ11, MJ11, MN10, MFL13, NBZ13, PJ12, RHS11, SI11, SLD13, Tan10, WWL10, WWG10, WM12b, YZ11b, ZCH10, Zha11b]. Delay-dependent [HGW12]. delay-differential [SLD13]. delay-induced [Gre10]. delay-partitioning [LSXZ11]. delayed [CYJS14, CP13a, GW14c, HLWM10, LSXZ11, LWZ13, RRMD13, WF10, WZ10a, WW11b, hXZ11, ZZ13]. delayed-claims [GW14c]. delays [BV10, CL11c, DL13, DF14, FFM^{+}11, GLHCLP11, LF11, RBK11, Wan11c, WYO11, Wei11, XJ11]. Delta [AH14b, BSJ14]. demand [HD13, Wu10]. demands [SJA12]. Demiclosed [pYXPH10]. demicontractive [CM10a]. dendritic [Dou12]. denoising [Bir11, PY10a, YSZ11]. denominator [IA13]. dense [FV10, KG12, Rum13a, Rum13b, TA13]. density [GA12, MK10, ZZ13]. dependence [BG13, BDMR10, BKV10, DLSV14, Nai11, Zha10a, ZY10, ZY11a, Zha14c]. dependent [BY12, BAK11, BSJ14, CR12, CGS11, DS14, DBZ10, EGES14, EGON10, GS14, HWG10, HKL11, HL10a, HL11, HGW12, JJL11, JW10a, KR14, KT13, LD10, LPDN13, LB14, MJ11, MSZ13, NPTH09, PPG14, PR12b, RMP^{+}13, RBK11, Sch14a, Sch14b, SYK14, SOL11, TSC10, TR11, WF10, Wan11c, WYO11, YY11, YLS12, YF14, mZkPkL11, ZD11, ZwGhX14, dAGR10]. depending [WLZ14]. deposition [BKN^{+}13]. depth [DI12]. dequantization [HM10]. Derivation [GLHCLPN10]. Derivative [CFH14, CXH09, CN11, CHMT13, GP12, JM12, JGH^{+}13, KAYS14, LLS14, NBZ13, PFLZ11, WLZ14, YPL10, ZJ12]. derivative-free [CXH09, CHMT13, NBZ13, PFLZ11, YPL10, ZJ12]. Derivative-optimized [CFH14]. derivatives [BDD^{+}10, CW11b, CCH12, DFG^{+}10, Dos14, GSNA13, HS11a, HMA12, JD13, MX10, NÖ10, Xie12]. Deriving [YY11]. descending [MS12]. descent [LZ11, NY12, DW11]. describe [BBP14]. describing [FHPH^{+}13, PC12b]. description [KSRM12]. descriptions [DDV^{+}11]. Design [DKU14, AY13, GU10, LV14, Pro12, Prz13, RMTO11, ZBZ10, ZGZB13]. detect [MD10a]. Detecting [AHM14, AGT12, LM12a]. Detection [GST12, Alc14, BR13, CRS12, KBKR14, Wri10, ZZP13]. deteriorating [GU10]. determinant [Cin12, Cin14]. determinantal [CL10c]. Determination [CM10b, HTLI14, VS13a, ZSD10, MSW14]. Determining [CGRLR11]. Deterministic [Rie13]. deTestSet [MCS12]. detrended [BAdS^{+}10, BAdS^{+}12]. Development [Bak10, GWKE14, Li11b, MHH^{+}14, SKK14]. deviating [LL13b, Wan10, Wan12a]. deviation [Liu11c, PP14]. deviations [Sez10]. Dew [BSS13]. DFLU [AGJ13]. DFT [OS10]. DG [SKK14, XL11, ZL10]. DG2 [BKB14]. DGFEM [FHPH^{+}13, RW10]. DGTD [LVD^{+}14]. diagonal [Büy12, KR10, PZJ12]. diagonalization [ISN10]. diagonally [DP14a, GS12, McN10]. diagrams [AHSN13]. dielectric [CDK10]. dielectrics [CC13c]. diffeomorphisms [JY13]. difference [AV10a, AGT12, AWF11, ADLS14, Bak10, BGLV11, BJ12, BB12a, BK12, CZWD14, CCDS14, CXL10, CL11a, CWY11, CGL10, CR11a, CET14, CID10, DLG14, DXL10, DF12, DFH14, FDD10, FD13, FJ14, FKL13, FWBSBGA11, For10, GS10, GUBS13, GN11, GSGN11a, GSNA13, GWZ11, HMA12, HMVPAR11, Hu14, HS12, KH10, Lak13, LLS14, lLzS11, Lia14, LGZ10, MV12b, MM12b, NNO12, Obe13, PRT10, PMC11, SKN11, UBSG12, Wan11a, Wan11f, Wan11d, WC14, Yar10, ZBW11, ZTC14, ZD10, ZWZ11]. difference/control [ZTC14]. Differences [SPC14, AS12a, BB11a]. different [GI10]. differentiability [NRP13, dO10]. differentiable [Ant11, FFM^{+}11]. Differential [Ali11b, BNY14, LP10a, AØ14a, AAY11, AMÇ13, AM11, AC14b, BY11, Bak10, BS11a, BSS14, Bar11, BT12, Beh14, BGL13, BP12, BSS13, BJ14, BR11b, CK13, CZ13, CA12, Cas10, CY11, CZ12b, CC14b, DS12a, DO12, DXL10, DZZ14, ESM13a, EZ14, EDP14, FID11, FKL13, FFM^{+}11, FLLT10, FYM10, GXW14, GH14a, GP11, GAS14b, GPHAS14, HHHS14, HGW13, HXSF14, HGW12, Hua12, Hua14a, JM11, JM13, JD13, KS12a, KG13, KKN11, KA12, KLNS11, KSLD11, KB11, KB12, KB13, LD10, LSD11, LM12a, Lan10, LLS14, LF10, LS11, LXY13, LVCC13, LTFL10, LLSW11, LM14, LB09, LW10b, LB10, LYZ11, LJ12, LM13b, LL13b, LHL13, MLZ14, MSG11, MS13a, MJ11, MN10, MFL13, MR13, NS10, NOS14, NBZ13, NÖ10, Odi11, OAHR11, ÖY14, PJ12, PT11a, PT11b, PT12, PT14, PWZ11, Pul14]. differential [RHS11, SH10, SC10b, uIAAF14, Sla14, SZS11, Ste10b, SP11, SLD13, TM11, Tan10, Tar10, Tez14, TS12, TS14, VS14, Wan10, WZS10, WZ10b, WWL10, WL10a, Wan11e, Wan11c, WYO11, WZL11, Wan12a, WLZ14, WWG10, WM12b, WT14, Xu11b, YKL10, YW10, ZAT11, ZFG10, Zha11a, ZAWA13, ZF14, ZC10b, ZRW11, ZW14]. differential-algebraic [BY11, BS11a, BSS14, BSS13, CK13, GPHAS14, LM14, NOS14, VS14]. differentially [HL13a]. differentials [Ali11b]. Differentiation [yLGP11, AA14c, BBD13, MY14, RMR11, SL14a, WW10b, ZCC^{+}13]. diffraction [BDL10, DC10, DCZ10]. diffuse [PHA14]. diffuse-interface [PHA14]. diffusion [AK14a, ANS11, AD10b, APW14b, BM10, BV10, BS13a, Bay14, BS13b, BFB14, BSZ13, BKB14, BR11b, CP13a, CHMD14, CL11b, CGL10, CGS11, CG13, CLMR14, Cui14b, DV13, Dra13, DLP13, ESV10, FYA^{+}14, FdOS12, Fu10, GZ11, GN11, GPHAPR14, HMS11, HMVPAR11, HGW13, HL14, HS12, JS10, Jeo10, JWY11, KMHM14, KPsQ11, KR14, LP12, LCZ11, LWZ13, LJ11, MS11a, MPS11, Pop10, PMC11, RK11, RRMD13, SYK14, Sto11b, SBGB10, TTB14, TY11, TR11, VS13a, VS13b, VS14, WZ10c, WZ10a, Wan11f, WD14, XL11, XGL12, XZH14, ZC10c, ZLC10]. diffusion-chemistry [GZ11]. diffusion-convection [Wan11f]. diffusion-convection-reaction [AK14a, ANS11, SYK14]. diffusion-reaction [GPHAPR14, RRMD13]. diffusion-wave [Sto11b]. diffusions [LM13a]. dilation [CG11a, GPS13].

Dimension [KL10b, BF10, Kho10, LZ14a, SF12]. Dimension-splitting [KL10b]. dimensional [APO14, ABS11a, AE11a, AM14a, AAD13, AuIK14, BMMR11, BDSG10, BDL11, BSJ14, BL11, CYJ13, Cha11a, CSZ11a, CGM10, CK14b, CGK10, CyY10, Cui14a, DA10, DD11, DKH^{+}10, Dou12, EASS13, Fat11a, Fat11b, Fat13, FKL13, GH10a, GW10, Gas14a, GMRK14, HdJL^{+}13, HTH10, HS12, HHLW14, JW10a, JV10, KSK11, KSLD11, KHFC11, LWZ13, MD10b, Moh13, NLO13, NRP13, NMBG14, PR12a, PC12b, Sch14c, SC10a, SWY14, SLK14, WZ10c, WG10b, WLL^{+}11, ZML14, ZTC11, ZTC14]. dimensionality [WWW14]. Dimensions [BwKB^{+}14, DCJ13a, BOR14, Bra14b, CC11a, GJ13, GZ14a, GZ14b, LC11]. DIMSIMs [IJ14]. Dirac [MRD13]. Direct [AHH13, SMMD11, BHPS12, Cai13, CV10a, ESM13b, KMD12, MMF10, MM11a]. directed [Rum13b]. Direction [Cui10, Bog14b, CC14a, ZBC13]. directional [BISS10, EY10, vGALL10]. directions [NY12]. Dirichlet [CDKN15, CDKN14, FPP14, SSP12, Sun14]. Dirichlet-to-Neumann [CDKN15, CDKN14]. Dirichlet/Robin [Sun14]. DIRK [Ran14]. disaggregation [PM11b]. disc [FMP14]. discharges [VMAVGCM13]. discontinuities [MD10a, Rua13]. discontinuity [AGT12, BR13, Wri10]. Discontinuous [KSTW14, Sta11, AJK12, BFDP13, BY14, CMR11a, CC13c, CN10, DL12, ESV10, EJ10, FL10a, Gia12, Gia14, GH14b, GG10, HX14b, JWY14, Kuz10, LM13a, Li11b, LMY11, LMYJ12, Mar11, MA14b, MWWY14, MJ14, MR13, SW12a, Sch14c, Wan11a, YJ11]. discount [HW10]. discounted [hBW10, CGW14, GW14c, hXZ11, ZwGhX14]. Discrepancy [BB14a]. Discrete [Min10, WZF11, AMV13, ASST10, AH14b, AAD14, BAK11, BS10a, BCM10, hBW10, BISS10, BK13, BHJ11, BM12, BGKVW10, CW11a, CW12, CGW14, CW14a, DS10a, DR14c, Ery13a, Gos11, GWKE14, GG13b, HS10a, HR10, HR12, JS10, Jun13, KKPV11, KP14c, LB14, Ma10, MRD13, MT10c, MRS10, Nei14, PR12a, PC12b, RT12, SLW14, SJA12, SKAM11, Sza10, TYCL14, WZ12, Wan13, WL14b, YZ10a, YLS12]. Discrete-time [WZF11, BHJ11, LB14]. discretization [AK14a, AK11, BS13c, CAT13, DD11, JR10a, KMLT10, NTNBV14, PSRV10, RMP^{+}13, SYK14, SHZ12, SCDA10, Spr14, VVD^{+}14]. discretizations [BB10b, FD13, MSG11, MMP11, RCM^{+}12, WS13b]. discretized [CK14a, DMR13, SI11, WHL13]. disease [ÇDE14, Ven10]. disjoint [JKV11]. Disk [BSV10, BC12, BSV12, DKU14, OCU11, WW11a]. disks [ZTL13]. dispatch [BSBA14]. Dispersion [GPKV10, HML14, Lia14, MD13, WZMC11, WC14, ZD14, ZW10]. dispersions [HHP^{+}10]. dispersive [ÁD11a, BKL14, DLG14, MOG10]. displacement [BySZZ13, BDyS14]. Dissecting [KDG^{+}13]. dissipation [BBKS10, CM10b]. dissipative [BG11, HLM10, MF14]. Dissipativity [LW10b, Hua14a]. distance [FS12b, SC10d, WGZL14]. distances [EY11]. Distributed [SYK14, AB11b, BV10, Mau10, PVM^{+}14, ZH11, ZZG14]. distribution [AHAH11, ADA13, Ars11, Ars14, GSK11, Ism14, KKK12, LQYL11, LCN12, MD13, MRAM10, MKK14, Nad11, NACLR12, OB14, PP14, PR12b, SCRB^{+}14, Tor14, WS13a, WSS14, YG13, YG14]. distributional [LA11]. distributions [AAK10, AKA14, AELH10, ABN10, BK13, Col13, Ery13b, IR11, Isl14, Tur12, XCZ^{+}11, YE14]. div [KP14c]. div-rot [KP14c]. divergence [HP13, PM11a]. divergence-free [HP13]. diversification [RSK14]. divided [BB11a, GSGN11a, GSNA13]. dividend [CP13a, GL10, LL11c, TYCL14, YSW13, Zha14c, Zhu14, ZwGhX14]. dividends [CGW14, WZ11b]. divisibility [GM14b]. divisions [EK12]. divisor [CKM14, WL11b]. divisors [RSS14]. DLASQ [AMMS10]. DNA [ÁECS12]. DNS [YESD14]. do [AMPR10]. doi [DC10]. Domain [BHMW13, Spr14, SM14b, XCQ14, BG11, BKN^{+}13, BFP11, CDTV11, CR12, CY11, CHH10, CvdMS12, DDV^{+}11, FL10a, KR14, KMP14, MS11a, Man11, PAJ10, RK11, SLP11, SM10, TKW12, XN14, XW12, Yan10a, ZFY11]. domain-decomposing [Man11]. domains [ABHMW11, AV10a, AV10b, AD11b, AAD13, DMB10, GPZ10, JKV11, Li11a, Lui10, LDGY10, MNQ14, SSP12, SWY14, SSR10, Uty10a, ZAWA13]. dominant [MV12a, McN10, SMD14]. dominated [Jeo10, SRP^{+}13]. Donoho [Gos11]. double [BG10a, EJJ11, Fuk11, Fuk12, Fuk13a, GRV11, HZR14, KA10, MHL11, MT10c]. double-band [KA10]. double-bar [GRV11]. doubling [GG14, xLwCL10, LwCLW13, WLD14]. Doubly [KW10, Liu14a]. down [WSK^{+}11]. down-sampling [WSK^{+}11]. downdating [Liu11a]. dqds [AMMS14]. DR [NLL13]. Drazin [CID10, LCI11, MMSW13, YL11b]. Drazin-inverse [MMSW13]. DRBEM [ATS14, GTS14]. drift [BT12, VAZS13]. driven [AG10a, Lan12, RE10, Sch14c]. Droplet [GGS12]. drops [BSS13]. Drucker [PTP10]. drug [FdOdS12]. DSS [KS14a]. DSSY [PSS13]. DtN [BDSG10, HNPX11]. Dual [Yua12, BFP11, Che14b, CSSX14a, CSSX14b, CLS10, CFdAR11, FO10, Hui10, KMLT10, Lam14, LL11b, LL11d, SKAM11, WL10d, Woź13, Woź14, ZZG10, ZWG11, Zie14]. dual-core [SKAM11]. dual-parametric [LL11b]. duality [GD14, PY10a, SL14b]. duals [Gro13]. ductile [VB14]. ducts [HHM12]. due [Cha11a]. Duffing [DBZ10]. Dunkl [CG11b, CV10b]. Dunkl-classical [CG11b]. dyadic [CGS12]. Dynamic [APW14b, EX14, AZ11, BBS14, BAAS11, BGLV11, BHBP10, CFL^{+}14, GZL14, GUBS13, LH11a, MKK14, NTPVNHLX14, NA11, PC12b, Ren10, WHLC11, WX12, WL14b, Zha11b]. Dynamical [GH10b, AMV13, AG10b, AP10, BGS12, BHJ13, BDMR10, BGKVW10, FHPH^{+}13, FGO11, LL14b, Naz11, Sch14c, SDK10, VS13b]. Dynamically [ADLS14, FDD10, SW12a]. Dynamics [GHR10, Reb12, SI11, AA14c, BFB14, BB10a, DGC^{+}13, HBE14, HLWM10, SC10a].

each [Miy12]. ear [ILEE13]. earth [MR10]. earth-atmosphere [MR10]. eccentric [ASG11]. Ecoepidemic [Ven10]. Economic [GR12, GU10, AA14c, BSBA14, GLLM12]. economies [BOR^{+}12]. eddy [CR12, GJCL11, WCZ11, FDD10]. EDEM [ABHMW11]. edge [BFS14, YAKA11]. edge-bipartite [BFS14]. Editor [KN10b]. Editorial [Ano11c]. EEG [Che14a, DCGS13]. effect [BSBA14, HD13, Ku13, Mee12, OB14, Vel10]. Effective [HP13, CPY10]. effectiveness [DCG^{+}14]. effects [BKMG11, IO14b, KS10a, LGR11, NPTH09, PC12b, TSC10, XZ14]. Efficiency [DAH11, GSGN11b, Kho11]. Efficient [Alc14, BG11, CÖ11a, DSL12, Dra13, GCNL14, GGLR10, GÖ14a, HLM10, KX13, KB11, KZ13, KH13, LLK14, LM12b, LM14, LdHA10, MT13, NRS10, SSP12, SW12a, TTB14, XXH14, AAK10, AS12b, AEG^{+}11, Bal14, BJ14, BIT11, CC14b, Chu10, CwYyY11, DLG14, DDV^{+}11, Dha14, EDAP10, FN12, GSM11, GPHAM10, Har11, KKR11, KSLD11, lLzS11, LJMG11, LC10, MX10, NR14a, pWqSmZ14, LJMG11]. Efficiently [HS10b]. EFG [GCY13]. eigendata [ZBC13]. Eigenfunction [BHMW13]. eigenpair [Xie12]. eigenparameter [BAK11]. eigenparameter-dependent [BAK11]. eigenproblem [HY11, XLCC10, YGHH14]. eigenproblems [CLN14, DJ10, KDG^{+}13]. eigensensitivity [yZ11c]. eigensolution [KF11]. eigenspace [MV12a, Nau11]. eigenstructure [CL13b, YL14]. Eigenvalue [Bai13, HHLW14, RVN13, WL14a, AMMS12, AM14b, BN14, CL11d, CL13b, DJ10, HDT11, ISN10, JWL13, xLwCL10, lLfD10, LZI10, MY11, MVV14, Miy10a, Miy12, Nau11, NL10, OA14, PN11, PLN13, QT13, QC14, QW10, QY10, SBCD10, SG12, TL13, WTM12, YD11, ZBC13]. eigenvalues [BGM10, BCC^{+}13, Che13b, CL13a, CM10d, FV14, Liu12, Miy10a, RB13, ZZ12]. eigenvector [Rom11]. eigenvector-like [Rom11]. eigenvectors [KMG12]. eigenvibrations [BG10a]. eighth [CHMT10b, CTV11, WL10b]. eighth-order [CTV11, WL10b]. eKdVB [TA10]. Elastic [AC14a, BDL10, GPKV10, GST12, ILEE13, Mön10, OS10, WL11a]. elasticity [CRM10, Fat11a, Fat13, KL10a, Lam14, LL11b, LZ12b, LL10, NSZ10, Vel10, YC10]. elastodynamic [BFP11, VVD^{+}14]. elastodynamics [YY11]. Elastoplastic [ZTP14]. elastostatic [HKL11]. elastostatics [Gwi13, LHLC13]. electric [JS10, Vel10]. electrical [GKL^{+}12a, HHT12, ZDP^{+}14]. electricity [HT14]. electrode [HHT12]. electrodeposition [SBL12]. electrodynamics [DHR12]. electroencephalogram [Che14a]. electromagnetic [Bab10, BB10b, CDK10, DS11, GH10a, Jan14, LVD^{+}14, SW12a, VVD^{+}14, ZSD10]. electromagnetics [AC14d, CHH10, FL10a, PS10]. electromechanical [PSKA14]. electromigration [NS14]. electron [GPVL10, TT14a]. Electrostatic [HW12]. Electrostatics [Bra14a]. Element [MNQ14, SKS14, AMT11, ADG11, Ain10, AR10b, AD11b, BCM10, BW14, BS13a, Bay14, BB11b, BG14a, BFP11, Bra12, Bra14b, BySZZ13, BGP14, CHL11, Cao14, CM13, CC11a, CL10b, CHX11, CL11b, CL11e, CW14b, CvdMS12, DV13, DHR12, DGG12, DDG13a, Dek10, DGC^{+}13, DJ14, DLP13, DS12b, DS13c, FKNS11, FRB13, FPGMG11, FLHS^{+}13, Fu10, FLKW14, GW14a, GK14, GW10, GW14b, GPZ10, GY11, GMRK14, GWKE14, GH10b, GZ14a, GZ14b, GKT10, GKT11, HL10a, HJS14, HK11, Jeo10, JL11, JM11, JW10a, JKS10, JJE13, KL10a, KS14d, KP14c, KP10b, KS12b, Lam11, Lam14, LVD^{+}14, Li11a, LL11a, LGM14, LL11b, LLSW11, LZ12a, LLX14, LXSC11, MLZ14, MHL11, MSP12, Nei14, NTRNXB10, NV11, NS14, OS10, OA14, PSS13, PHA14, PL10a, PTSK13, QZX^{+}11, RMP^{+}13, RWRS14, RR10, SSR10, SACS10, SG12, SSK^{+}13]. element [SRP^{+}13, SM14b, SS14e, SLK14, TCD11, TR11, TKW12, TTSB14, VMS12, VdKPB10, VS14, WG10b, WLL^{+}11, Wan13, WY13, WMGM14, WRS14, XCY10, YC10, YL11c, Yua12, pZxY12, ZC10c, ZD11, ZLC11, ZX10, dAGR10, dFGAN11]. element-global [SRP^{+}13]. element/level [DGC^{+}13]. elementary [Cin12, Cin14]. Elements [HP10, AB10, AM14b, BL10a, CRM10, Dou12, FGPP13, JJE13, LZ12b, MA14a, Mön10, NRP13, RW14, Rua13, Tau13]. elimination [ADGP11, ADGP12]. ellipse [MT10b]. ellipsoidal [fLxX12]. Elliptic [CN11, Ali11a, AD11b, BG14a, Bog14a, CZZ14, CY11, CMR11a, DS13a, Dek10, Fuk11, Fuk12, Fuk13a, Fuk13b, Fuk13c, GW10, GY11, JKV11, Kır11, Kur10, LY13a, LMY11, LMYJ12, Min10, MJ14, PTH10, SSP12, SG12, TCD11, WY13, XZW14, YJ11, pZxY12, Zho11b]. elliptical [KH13]. ellipticity [BB10a, CXZ14]. embedded [PR10, Sch14a, Sch14b]. embedding [CL13a, Kır11, SC10b]. embeddings [Wih11]. EMC [ZDP^{+}14]. empirical [CFH14, HK13b, MD13, Wik10]. enclosure [KN10a, Miy10a, Miy10b, Miy12]. enclosures [GTS14]. encryption [AHS14]. end [CS14, HS11a]. end-point [CS14, HS11a]. Endpoint [Boy13]. endpoints [Ste10a]. energetic [ADG11]. Energy [BCMR12, ZML14, AHR14, BFGP10, CG11a, CDV10, FLHS^{+}13, GD10, HM10, JŽ11, RG12, RBD11, RBD14, SCFR14]. energy-based [FLHS^{+}13]. Energy-conserved [ZML14]. energy-like [RBD11, RBD14]. energy-momentum [RG12]. Energy-preserving [BCMR12]. enforcing [Kuz14]. Engineering [AVA13, GR12, SKS14, VAV11]. Engquist [Wan11a]. Enhanced [VBP^{+}10, BSKL12, BL14, LNJ^{+}10, SCRG12, Zim13]. enhancement [Dou12, LA10]. enriched [Naf14]. Ensemble [GW14a, HV11]. enterprise [LLL12]. entries [KR10]. Entropy [LQYL11, AO14b, ADMF10, BNY14, JPX11, TLNE10]. envelopment [Kho11]. environment [PTSK13, ZZC11]. environmental [VMAVGCM13]. environments [AÖY10, NACLR12]. epidemic [AELH10, SI11]. epidemiological [ADLS14]. epidemiology [KAS13]. Equal [Naf14, RP11b, Ros11]. equalities [ZH10c]. equality [AG10b, QH10]. equation [ABS11a, AV10a, AMÇ13, AU14a, AT10, AS12b, AM11, Asl10, Ass09, AK11, BPS12, BG11, BGGG^{+}11, BGL13, BSS13, BKV10, Cai10b, CCG13b, CL11a, CK14a, CCJC13, CLB11, CWY11, CL14a, CLS10, CN11, CR11a, Cui10, Cui14b, DLG14, DDM10, DGG12, DY11, DCJ13b, Dos14, DBZ10, DLP13, EK11, Fad10, Fat11b, FiMO10, FRXY13, GVM^{+}12, GN11, GG13a, GG13b, GP10, HLM10, HHH^{+}13, HTLI14, HD11, Hau10, HML14, HC11, HX14b, JK10c, JW10a, JWY14, JMS10b, KS10b, KHKM14, KKPV11, KHFC11, KP11b, Kre95, Kre11, Kud10, KRS11, KMLT10, KMP14, KT13, LD10, Lam11, LPDN13, LsGHZ10, LZL11a, LWXW12, LZDW14, LXX14, Lia14, LVCC13, LY13a, LZ14a, dL14b, LP13, LGZ10, MA13, MMS13, MM10, Mat10, MN10, MD10b, MM12b, MM11b, MS11c, MA11, MS13b, NSZ10, Nei14, OS10, PM13, PDP13]. equation [Pas11, PP10a, yP10b, PSRV10, PMC11, QXW10, RGJ10, RD10, RT12, Ros11, RRMD13, SC10c, Sie11, SD10, SRWZ14, SM10, SM14b, SKN11, TR11, TLHU10, WTM12, WL10c, WZ10c, WZF11, WLWH11, WYO11, WC11, Wan11b, Wan12a, WLZ14, WC14, Xio10, XGL12, XZH14, Xu11a, YMS10, YNO10, YK11, YK14, YBD11, YLD13, ZSX11, ZBW11, ZTC11, ZYC13, ZMZ13, Zha13, ZD14, ZW10]. Equations [YY11, AAM14, AØ14a, AMT11, AAY11, AK14a, AF10, AE11a, ASST10, ASS12, AST14, AM14a, ÁECS12, ABG11, AS10b, AH10b, AH13, AAD13, AAD14, AKT12, AKT14b, AuI13, AuIK14, BKGMA10, BDD^{+}10, BH11, BMMR11, BY11, BY12, BAK11, Bak10, BS10a, BS11a, BSS14, Bar13, Bar11, BT12, BS13a, Bay14, BS13b, Beh14, BB11c, BUG^{+}13, BS10b, BL10a, BPGP11, BCC12, BR14b, BKB14, BKL14, BHJ11, BJRS12, BB10a, BB12a, BB14b, BJ14, BK12, BR11b, Cai10a, Cai13, CK13, CZ13, CBB13, CA12, Cas10, CGS12, CCDS14, CI13, CY11, CL10b, CZ12b, CC14b, CXH09, CGM10, CDS10, CHMT10a, CHMT12, CHMT13, CO11b, CDK10, CyY10, Cui14a, CFL^{+}14, DSC09, Dai11, DSN14, DS13a, DV13, DS12a, DS14, DO12, DD11, Dha14, DXL10, DMR13, Dra13, DZZ14, DS10b, DS12b]. equations [DS13c, DI12, ESM13a, EZ14, EDP14, EHR10, EHR11b, FID11, FP13a, FMP14, FKNS11, FKL13, FYA^{+}14, FFM^{+}11, FWBSBGA11, FLLT10, FYM10, FN10, Fu10, GS10, GZ10, GXW14, GH14a, GW10, GZL14, GP11, Gat14, GAS14b, GPHAS14, GSGN11b, GSGN11a, GSNDB14, GM10, GM13, GKL12b, GG14, HHHS14, HMS11, HL10a, HMVPAR11, HSY13, HM11, HHZ11, HGW13, HNK10, HL11, HGW12, Hua12, HS12, Hua14a, IR11, IK11, IJMV10, Jeo10, JX10, JM11, JM13, JD13, KC10, KS12a, KH11, KG13, KKN11, KA12, KLNS11, KSLD11, Koç11, KB11, KB12, KB13, LSD11, Lak13, Lam14, LM12a, Lan10, LLS14, LF10, Li11b, LH11a, LS11, LLCL11, LWW13, LZ13a, LwCLW13, LXY13, lLzS11, LTFL10, LY10, LLSW11, LM14, LB09, LW10b, LB10, LW11, LYZ11, LJ11, LJ12, LM13b, LL13b, LHL13, Liu14a, LZ14b, LL14b, LEVG12]. equations [LEVG14, LH11b, Luo10, LFWL14, MLZ14, Mah14, MZ14, MA14a, MSG11, MS13a, MT14a, MM11a, MJ11, MML11, MFL13, MK11, MR13, MMMG10, N’G10, NS10, NOS14, NLO13, NNO12, NLL13, NBZ13, NÖ10, Obe13, Odi11, OMS10, OAHR11, OPPV10, ÖY14, PM10, PSS13, PJ12, PT11a, PT11b, PT12, PT14, PWZ11, PFLZ11, PS11b, PS14b, PGH13, Pop10, Pul14, RK11, Ren10, RST14, San11, SYK14, Sha14, SH10, uIAAF14, Sla14, Śmi11, SZS11, Spr14, Ste10b, SP11, SLD13, SM11, TM11, Tan10, Tar10, TP10, TCD11, TY11, TS14, TA10, VMMM10, Wan10, WL10c, WZS10, WZ10b, WL10b, WG10b, WWL10, WL10a, Wan11e, Wan11f, Wan11d, Wan11c, WZL11, WX12, Wan13, WD14, WLD14, WWC14, WW14, WWG10, Wei11, WZ11a, WM12b, WHL13, WT14, Wu14, XL11, Xu11b, XL12, YT12, YPL10]. equations [Yan11, YFWH10, YKL10, YZ11b, YC11, YSW12, YW10, YLTB14, Yun11, Yun12, ZAT11, ZML14, ZH10b, ZFG10, Zha11a, Zha11b, ZJ12, ZAWA13, ZTC14, ZY14a, ZF14, ZC10b, ZGZW10, ZM14, ZCS11, ZRW11, ZZ11b, ZD11, ZLC11, ZX10, ZW14, dFGAN11]. equidistant [BDH11]. equilibria [ÁD10a, GRV11, GLLM12]. equilibrium [BFG^{+}14, CB13, HLC11, JK10b, KPW10, Li14, LZL11b, OG10, QCK10, RP10, She10, WG10a, WSWD11]. equity [FM14]. equivalence [Lan10, Zha10a]. Equivalent [FP13a, Kur10, WL10c]. ERES [CKM14]. ERKN [YFZ14]. Erlang [EC13]. erosion [KS12b]. Erratum [DC10, Gün10, LB10, LHL11]. Error [ÁD10a, BGP11, Boy10, CW11a, CW12, CCH12, DSC09, Dai11, HNPX11, KG13, LTW10, LHLC13, LW13, LvDtTB^{+}13, MA11, PS14a, PSRV10, Sch14c, SP10, Spa12, Tan14b, Wan11e, XCY10, XL11, pZxY12, AW10, Ain10, ANS11, AM14b, AAD13, AKT12, AKT14b, BG14a, BFP11, CS10, CZWD14, CM13, Che10, CL11e, CG14, CW14b, CET14, CN10, DP10a, DP11, EJJ11, ESV10, FW10, FM11a, Gia12, GY11, HK14, HJS14, Hua14b, KL10a, KPKK12, KN10a, KLNS11, KP10b, Ku13, KW10, KW11, LPDN13, LGM14, lLzS11, LY13a, LLX14, MVC10, MJ14, NV11, NÖ10, OCU11, SW10a, Seg12, SR13, Wan11d, WK14, XSH10, ZL10, ZCC^{+}13, ZWZ11, dFGAN11]. Error-driven [Sch14c]. errors [CH14, XLCC10, pYXPH10]. Esscher [NR14b]. essential [LHL13]. essentially [CCZ12]. establish [ZF14]. estimate [BGP11, CW14b, DH10, HJS14, KG13, OCU11, SXZ11, XZH14, ZL10]. Estimates [EJJ11, Ma10, PS14c, Sch11, AM14b, Bai13, BG14a, CS10, CL11e, CW12, CCH12, DSC09, Dai11, ESV10, GY11, KN10a, KKN11, Ku13, LPDN13, LGM14, lLzS11, LY13a, Lui10, NV11, NÖ10, PSRV10, Rom11, Seg12, Spa12, Sta11, TA13, WL14a, XCY10, pZxY12]. Estimating [ADA13, ÁAG^{+}13, BA10, RLGGGÁM14, MRAM10]. Estimation [AO14b, Isl14, Tor14, YÖVAYOW14, ANS11, BFP11, CCG^{+}13a, CZWD14, CG14, CET14, DJM10, GLHCLPN10, GLHCLP11, GSK11, KBKR14, KW10, KW11, LWH^{+}11, LW13, MD13, RSFS14, WK14]. estimations [dFGAN11]. estimator [CN10, DAH11, Gia12, LLX14, WWW14]. estimators [AKT12, AKT14b, BAU13, CM13, KL10a, KP10b, NR14a, PM11a, Pit13]. ETD [VMMM10]. ethylene [WCZ11]. ethylene-air [WCZ11]. Euclidean [SF12]. Euler [BT12, CK14b, DGG12, Fad10, GQ14a, HL10a, LXY13, MSG11, MM10, Pré10, Rza12, Tan14b, Wei11]. Eulerian [BBKS10, He12, XhW11]. European [LL13a]. evaluate [LL13a]. Evaluating [ZYL11, BDR10]. Evaluation [FdLW^{+}11, MOA10, SW11, WL11a, AAK10, BOR14, Che13a, CA14, DP10a, DP11, ELA10, GKL^{+}12a, HTH10, HS11b, JGH^{+}13, JJE13, KS14b, LLL^{+}11, MS13b, QZX^{+}11, WX10, XXH14, Yan13]. evaluations [Fuk13d, WCH11]. event [NR14a]. Evolution [PM10, Ali11b, HXSF14, MZ14, Mor10a, Wan13]. Evolutionary [Bak10, MAK^{+}14, KPBN13, LTX12, LJ13, LZWJ11]. evolving [YESD14]. Exact [CK14a, Hui10, LLL^{+}11, LHC14, Nad11, Asl10, Ass09, EK11, KAE^{+}11, Kud10, KRS11, LLT^{+}14, LZ13c, LL14b]. example [MVC10, Ran14]. examples [AP14a]. exceedance [Tur12]. exchange [SS14e]. exchangeable [BP11b, Ery13b]. excitation [Reu10]. excited [BS11b, DC11, MK10]. executive [KWXZ14]. Existence [BK10, BB11c, CL14a, KC10, KLL11a, NPTH09, TSC10, Yan11, ZFG10, HCWsO13, N’G10, Nie13b, SL12, WWG10, ZC10b]. Exit [YSW13]. Exp [Asl10, Ass09, Kud10]. Exp-function [Asl10, Ass09, Kud10]. Expanded [JL11]. Expanding [AHK14b]. expansion [Ali11a, HNK10, LW14, PWZ11, TT14a]. expansions [BE11, HWG10, KXH13, LP10b, MJ10b, Par10, Rza12]. Expected [CGW14, hXZ11, ZwGhX14]. expensive [LZWJ11]. experiment [RMTO11]. experimental [Dea11, GKL^{+}12a, RYM^{+}14]. experiments [FM11a, WCH11, ZCC^{+}13].

ExpertRank [Zho12]. Explicit

[BB12a, CCCM13, Esc14b, Fat11a, GM10, GQ14a, Lu13a, LO14a, LO14c, SWY13, BDMR10, Cai10b, CCG13b, El 11, FG14, GPHAS14, GM13, GH10b, GQ14b, HLC11, HYY14, HS11b, Jan14, Kha10, KW10, LM14, LL14b, LT13, Lu14, NTNBV14, NBZ13, PR10, PGH13, SCRG12, SM11, WK14]. Exponential [AC14b, Gon10, Hua12, LH11a, LO14b, mZkPkL11, BB11a, CZ13, CS14, CR11b, Cui14b, HZR14, KLNS11, LT13, LO14a, LO14c, Luo10, MSG11, NH10, SGG11, TY11, WS13a, XC14, XJ11, ZLG12, ZW14]. Exponentially [CIPS14, DEP11, DP14a, HVV11b, HV12, Par10, MRD13, PR10]. Exponentially-fitted [CIPS14, HV12]. Exponentials [CL10a]. expressed [Liu14a]. expression [El 11]. expressions [Fat11a]. Extended [BHMW13, BRZ10, OR14, AKA14, BB12a, CDPP10, CPP11, GG13b, KS10b, KK14, MVV14, PHA14, YFZ14]. Extended-Domain-Eigenfunction [BHMW13]. Extending [AH10b]. Extension [ÖÖA11, EHR10, KO13, SACS10, WZ12]. extensions [AH14a, GZ14a, GZ14b]. exterior [OCU11, RCM^{+}12]. external [Reu10]. extra [LD10]. extraction [HCM14]. Extragradient [NASNY11, WHL10]. extraordinarily [Bal14]. extrapolation [Bal14, BRZ10, CHL11, CG13, Mar11, Wan11e, WW10c, ZFH10]. Extremal [WZS10, FKL10]. extreme [EY11, vDvFZ09, vDvFZ13].

face [HHLW14]. face-centered [HHLW14]. faces [LJMG11]. factor [ADGP12, CG11a, GÖ14a, Sch11, VTLF11]. factorization [AC14b, BFLP12, GG12, Liu11a, MS11b, Mir14, SSF^{+}11, WH12a, WH12b]. factorizations [AGT13, KMG12]. factors [ADGP11, HW10]. failure [ADA13, CL14b, WSS14, XSV12, XSV13]. failures [EY11]. families [EZ10, HC14, PPH10, Zie14]. family [AGM12, AAT11, AKA14, ADL^{+}10a, CC13b, CXH09, GK11, GKLZ12, GPHAM10, GSNDB14, GHR10, HPR12, HLC11, HL10b, KAS13, SYİ11, TP10, WSWD11, ZDZ11, ZJ12]. Fast [BC10, BR14a, BDL11, CD10, CC14a, DM12, ESM13b, EY10, Fuk13a, HLFP14, JX10, KS14b, Kho10, KR10, LFWL14, Miy10a, Miy10b, Oou13, OORO12, VTLF11, WZ10c, WD14, AH13, BDD14, BOR14, BFMS11, Cai10a, Cai13, Cin14, Fuk11, Fuk12, Fuk13d, GM14a, GCH14, HZR14, JY13, LCXL11, McN10, SSP13, SGG11, Str10, WJJ14]. Faster [BOR14]. FBSDEs [Zha10a]. feasibility [CCG12]. feasible [JW10b]. FEAST [KDG^{+}13]. features [BCR11]. feedback [CYJS14, HL10c, IO14b, WW11b]. Fekete [BSV12]. FEM [NTRNXB10, BMN10, CR12, CFL^{+}14, FPP14, Gwi13, HLSX11, HNPX11, JKV11, KSA14, MKS12, NTPVNHLX14, PSKA14, RW14, Seg12, SDK10, SCDA10, SACS10, SKK14, VBP^{+}10]. FEM-based [VBP^{+}10]. FEMTEC [SKS14, Sol12]. fence [NdCDCFM10]. fermentation [WYY^{+}12]. Fermi [Boy13, KHKM14, PDP13, SP11]. ferromagnetic [FLHS^{+}13]. few [ML13]. Feynman [DHR12]. FFT [MPS10, PO11, Tan14b]. fiber [VB14]. fiber-reinforced [VB14]. Fiedler [pWqSmZ14]. field [CGCG^{+}13, CR14, FND10, JS10, Jan14, LGMA13, MKD12, Zha14a]. field/circuit [Jan14]. fields [AÖÖ14, AA14b, ÇMT14, CGS13, DD14b, DD14a, GM14b, KP14c, WZ12, ZSD10, ZZC11]. fifth [LC10]. fifth-order [LC10]. Fike [SW12b]. Filippov [DD14b, BHJ13, BJ13, DD14a]. Filled [LY10, WLWH11]. Filling [BFGP10]. film [KS14c, WL11a]. Filon [Dom14, Mel10]. filter [DJM10, ISN10, LV14]. filtering [CFF12, GNN11, KK14]. filters [BL10b, CCS10, CH13, Han11]. finance [DH14]. financial [NÖ10, YLS12]. Finding [LJMG11, YT12, DP14c, PPH10, SCFR14]. Fine [KA10]. Finite [AV10a, AMT11, ABS11b, BGLV11, Bra12, Bra14b, DHR12, Dek10, FJ14, GMRK14, GKT10, HP10, HXSF14, JKS10, KS12b, Li11a, MNQ14, Nei14, Obe13, SPC14, SKS14, SS14e, TTSB14, VMS12, Ain10, AÖÖ14, AA14b, AS12a, AWF11, AR10b, ADLS14, AD11b, AB10, AM14b, BS10a, BCM10, BW14, BS13a, Bay14, BB11b, BS10b, BL10a, BG14a, BJ12, BGP11, BFP11, BB12a, Boy10, BySZZ13, BGP14, CHL11, CZZ14, CZWD14, Cao14, CM13, CXL10, CL11a, ÇMT14, CC11a, CL10b, CHX11, CL11b, CWY11, CL11e, CG14, CXZ14, CW14b, CMR11a, CGL10, CR11a, CET14, CvdMS12, DLG14, DV13, DGG12, DDG13a, DXL10, DGC^{+}13, DLP13, DS12b, DS13c, DF12, DFH14, EZ10, FDD10, FD13, FKNS11, FKL13, FRB13, FPGMG11, For10, FGPP13, FLHS^{+}13, Fu10, FLKW14, GW14a, GK14, GW10, GW14b, GVF^{+}13, GUBS13]. finite [GI10, GN11, GY11, GM14b, GZ14a, GZ14b, GKT11, HS11a, HMA12, HL10a, HML14, HMVPAR11, HL10b, HS12, HK11, JL11, JM11, JW10a, KH10, KL10a, KS14d, KP14c, KP10b, Lak13, Lam11, Lam14, LL11a, LL11b, Lia14, LLSW11, LZ12a, LZ12b, LY13a, LLX14, LMY11, LMYJ12, LXSC11, MLZ14, MHL11, MV12b, MM12b, MSZ13, MJ14, NTRNXB10, NV11, NNO12, NCCF^{+}10, NS14, ÖY14, PSS13, PHA14, PL10a, PTSK13, PRT10, PSRV10, PMC11, RMP^{+}13, Rua13, RWRS14, RR10, Sid14, SSR10, SACS10, SG12, SSK^{+}13, SRP^{+}13, SM14b, SLK14, SKN11, TCD11, TR11, TKW12, UBSG12, VdKPB10, VS14, WG10b, Wan11a, WLL^{+}11, Wan11f, WZ12, Wan13, WY13, WMGM14, WRS14, WDZ10, WC14, XCY10, YC11, YL11c, Yua12, pZxY12, ZZS11, ZTC14, ZY14a, ZD10, Zho11a, ZC10c, ZD11, ZLC11, dAGR10, dFGAN11]. Finite-difference [BGLV11, HMVPAR11, HS12, Lak13, MV12b, PRT10]. finite-element [TKW12, WLL^{+}11, dFGAN11]. finite-part [WDZ10]. finite-range [Sid14]. finite-source [GI10]. finitely [Mol10, Yun12]. firm [LZT13]. First [MTJ11, Vel11, ADA13, BSZ13, BGL13, Cas10, Dea11, DDV^{+}11, DY11, Dom11, EJ10, Fuk13b, GPHAS14, KB13, KP10b, LL13b, MMS10, uIAAF14, SB10, SZS11, TM11, Tar10, WZS10, WZ10b, ZL10, Zha11a, ZWZ11]. first-failure [ADA13]. First-order [MTJ11, Dea11, DY11, KP10b, LL13b, MMS10, SB10, WZ10b, ZL10, ZWZ11]. fish [ARF14, CRF14, RMF11]. Fisher [Dom10, GG13b, HM10, KS10b]. fit [DKU14]. fitted [CIPS14, DEP11, DP14a, HVV11b, HV12, PR10, WT14]. Fitting [Nie13a, EY14, KV12, PY10a, Str11, Str12]. Fitzpatrick [XZL11]. five [LO14b]. Fixed [BGL13, Iid12, Luo10, She10, ZW11b, Bar11, CB13, CDTV11, CM10a, ĆRR^{+}10, HLC11, HL10b, JK10b, KPW10, LYZ11, MMS13, PD13, RKS^{+}13, Van10, VAR11, XSH10, ZF14]. fixed-point [Bar11]. flageolet [Fie13]. flame [GJCL11, WCZ11]. flat [FWBSBGA11, KD10]. Flexible [ZG14, AHM13, BP11a, WRS14]. flight [BBP14]. floating [OORO12]. floating-point [OORO12]. flops [GWX10]. Floquet [Tau13]. Flow [AVMVMV10, APO14, AHLJ14, AJK12, ATS14, BCD^{+}10, BSKL12, BHMW13, BJ13, BGP14, CPP13, CL11d, CDPP10, CPP11, CDS10, DA14, DA10, DQST14, EDAP10, FWBSBGA11, GGAVRC14, Gas14a, GTS14, HDF^{+}10, HJ11, HZ14, HL13a, HdJL^{+}13, JMS10a, LE10, MDVB13, NO12, PR12a, PDE^{+}13, PPG14, SSW14, SSP13, SLK14, SM11, TTSB14, WTM12, YMB14, ZZC11, dAGR10]. flows [eMA14, BW14, BG10b, DMMP10, DGC^{+}13, GK14, GH14b, GMRK14, GKT10, GKT11, HHM12, IG12, KAE^{+}11, MS13c, ORP14, PDE^{+}13, PKH14, RVDM10, ZY11b]. fluctuating [HD13]. fluctuation [BAdS^{+}10, BAdS^{+}12]. fluid [AHLJ14, CPP13, CDQ14, DBHV10, DA10, DGC^{+}13, FHPH^{+}13, HL13a, HBE14, JMS10a, KS14d, KLM14, NTPVNHLX14, PP10b, Rah14, SSW14, SM14a, SSP13, TTSB14, YZ11a]. fluid-solid [NTPVNHLX14]. fluid-structure [CPP13, CDQ14, DBHV10, JMS10a, PP10b, YZ11a]. fluids [Bab10]. flux [AGJ13, EZ10, EDAP10, GZ10, KM10, KG12, Kuz12, LvDtTB^{+}13, MV14, MSZ13, SSW14, VS14, Wan11a]. flux-corrected [KG12]. flux-limited [KM10]. flux-limiter [GZ10]. foam [BCD^{+}10]. Fock [Bra14b]. Fokker [CR14, AB13]. fold [Jia10]. force [ZBZ10]. forced [NTRNXB10, TA10]. forecasts [Wu10]. Foreword [AL13]. form [Ku13, Owa12, SRWZ14, Spe11, VL13, WHL13]. formalism [CC11c]. formation [BKMG11]. forming [ZGZB13]. forms [DDM11]. formula [GQ14b, Hua14c, LA11, PS14a, Qi14, SHN13, Tan14b]. formulae [ÇU14, GQ14a, MO10, ML13, SC10b, SP10, Spa12]. formulas [FS12a, MT10a, MS12, SH11, Sid14, Xie11, ZCC^{+}13]. formulation [CDTV11, CSSX14a, CSSX14b, CAT13, FND10, HMVPAR11, HCM14, Jan14, JWY11, LP10a, LXSC11]. formulations [FP13a, JMS10b, KA14, MS13b]. forward [AØ14a, FLLT10, HHT12, HL14, WL10a]. forward-backward [AØ14a, FLLT10]. forward-path [HL14]. four [AS12b, BISS10, Flo11, LTX12, RB13, WZ10d]. four-directional [BISS10]. four-point [Flo11, WZ10d]. four-step [AS12b]. fourfold [BK13]. Fourier [AH14a, AA12, BC10, Cai10a, Cai13, CSZ11a, CV10b, HNPX11, HZR14, JX10, KKPV11, MPS10, MD10a, PWZ11, PMMS14, RSGL14, SW10a, SS14b, Tan14b]. Fourier-collocation [Cai10a]. fourth [AR10b, CI13, Cha10a, CHMT10b, DLG14, GK11, HS10a, HVV11b, Koç11, Mui10, PFLZ11, PRM11, RK11, ZX10, SKS14]. fourth-order [AR10b, DLG14, HS10a, HVV11b, Koç11, PFLZ11, PRM11, ZX10]. fourth-step [GK11]. fractal [BM11b, NS13, VCA14]. fractals [CGM10]. fraction [FPG12]. fractional [APJ10, Cui14b, DO12, DXL10, EDP14, FS14, FRXY13, GP11, GD14, HLM10, HMS11, HML14, JM11, JD13, KS12a, KG13, KAYS14, LXX14, LP10a, LYD13, LY13b, MLZ14, MZ14, MFL13, NS10, Odi11, PT11a, PT11b, PT12, PT14, PMMS14, SKN11, Tan14b, Wan12a, WD14, WLZ14, XGL12, XZH14, Yan11, YÖVAYOW14, YLTB14, ZAWA13, ZD14, ZW10]. fractional-order [HLM10, NS10]. fracture [PR12a]. fragmentation [PVM^{+}14]. frame [CG11a]. frames [BDJ^{+}11, CCH10, Gro13, Jia10, Jia11, YSZ11]. framework [Ain10, AB13, CMR11b, EF14, GWKE14, HS10b, Kro10, LS14, NR14b, RWRS14, Van11b, VS14, ZZBX14]. Fredholm [AE11a, AM14a, AuI13, BH11, BMMR11, Beh14, BGL13, CC14b, DCJ13b, DZZ14, EHR11b, LSD11, LWXW12, Ma10, MA14a, OMS10, OA14, OPPV10, WWC14, WW14]. free [eMA14, AWF11, BDD^{+}10, BKMG11, BHMW13, CRM10, CXH09, CHMT13, DdS14, FYA^{+}14, For10, HHH^{+}13, HP13, HKO11, LM12b, LM14, LFY11, MJ10a, ML10, NTRNXB10, NBZ13, PFLZ11, QY10, Sla14, YPL10, YC10, ZJ12, Zha13]. frequencies [Pag10]. frequency [BG10a, CLP10, CvdMS12, DDV^{+}11, IO14b, Lak13, LMCW10, TA13, TKW12]. Fresnel [KP11b]. Friction [FRB13, CCG^{+}13a, DKH^{+}10, GRZP10, Gwi13, HKL11]. frictional [SBFC10]. front [BFHK12, KKK14]. Frozen [GSGN11a]. FSAI [BM11a, FJP14]. FSAI-based [BM11a]. fuel [Sun14, SS14e]. Full [CCS10, CDQ14, ZW12]. full-Newton [ZW12]. Fully [DMR13, JS10, ASK11, BGLV11, DS13a, FKL13, GG13b, HJ11, Nei14, Wan13, YMB14]. Function [Vel11, Ali11a, AE11a, Asl10, Ass09, hBW10, BLM11, BSJ14, Boy11, Boy13, BFMS11, Büy12, CHL14, DS13b, Ehr11a, EGBS12, El 11, EL11, Fuk13d, GW14c, GK11, GKLZ12, GKL12b, HQR13, HMS11, HMVPAR11, Isl14, JLZ14, Kud10, Kuz13, Law14, LY10, LL11d, LG11, MRD13, MGZ10, Mol10, PM10, PM13, Par10, PAD10, PHS14, PR12b, Qi14, Rza12, VCA14, Wan11a, WLWH11, WCH11, Wan12b, WGZL14, hXZ11, XMSS11, Zha11c, ZZG10, ZLL10b, ZWG11, Zie14, ZwGhX14]. function-space [Law14]. Functional [LMM^{+}10a, RMTO11, AD10b, BPGP11, BCC12, CL14a, DF14, FYM10, GW12, HM11, LF10, LTFL10, LHL13, LMM^{+}10b, MRD13, MS11c, Owa12, PJ12, RHS11, Tar10, TLHU10, WZS10, WF10, Yan11, YZ11b, YW10, Zha11a, ZC10b]. functional-discrete [MRD13]. Functionally [WT14]. Functionally-fitted [WT14]. Functionals [BB11a, HM10]. Functions [BVV10, Vel11, AH14a, AA12, AAY11, AMV13, AM14a, AAD13, BMMR11, BGRS12, Beh14, Boy10, BL10c, BGVHN13, BE11, CGCG^{+}13, CAA10, Che13a, CS14, CZ14, CD11, CT10, CO11b, DST10, DHS11, EY14, FW10, FND10, FS12b, FKL10, Fuk13d, Gau10, GA12, GPS13, GZ13, HHX14, HV12, JKS13, JGK14, JZ10, JW12, JLZ14, KPA14, KG13, KS14b, LD10, LTW10, LY13a, LP10c, MMS13, MRAM10, MJ10a, MJ10b, MM12a, MK10, MSZ13, NH13, NS13, Not14, OGM14, ÖÖA11, PS14a, PS11c, PL10c, Prz10, QGG10, RW14, San11, SW10b, SWY13, SWY14, Sid14, SS14b, SS14c, Ste10a, UÇD14, Van11a, WX10, WW11a, Woź14, XC14, ZWM11, Zho11a, ZRK^{+}12]. fundamental [AC14b, LLCL11, LHLC13, OCU11, YY11]. Further [HMS11, Hos11, LHY10, WM12a]. Fuzzy [AS11, Liu11b, ASK11, FS14, FJLL10, KWW11, LQYL11, MS11c, Mol10, MOA10, MHH^{+}14, NF14, NR14b, QLL11, RPS^{+}11, Sto11a, WW10a]. FVM [NdCDCFM10]. FVS [DM10a].

G [WYP10, ZH10a]. Gabor [BDJ^{+}11]. gain [VAZS13]. Galerkin [ADG11, AAD14, BFDP13, CCJC13, CMR11a, CC13c, CN10, DS14, ESV10, EJ10, FL10a, Fat11b, FND10, GW14b, Gia12, Gia14, GH14b, GAS14b, GG13b, HX14b, JX10, KSTW14, Ku13, Kuz10, Kuz12, Lan12, Li11b, MS11a, MA14a, Mar11, Mat10, MA14b, MD10b, MWWY14, MR13, Pul14, Ros11, SW12a, Sch14c, Sta11, SM10, SM14b, WC11, WY13, WMGM14, XZW14, Zho11b]. GAM [ABÇ^{+}11]. games [AA14c, BP12, GPO14]. Gamma [Rza12, Vel11, BE11, Nad11, ÖÖA11, Qi14]. Gander [EHR10]. Gap [FW10, LTW10]. gaps [Gia12, GMZ^{+}10]. garment [LMW^{+}11, WLL^{+}11]. gas [AJK12, BL10b, BSKL12, GKT10, GKT11, HBE14]. gases [Col13]. Gauss [BSZ13, CT11, CIPS14, EJJ11, FGO12, FS10, KN10b, LYD13, ZY14a]. Gaussian [AE11a, Boy10, JWY14, Kni10, SP10, Spa12]. GBi [TS10]. GCR [AIA10]. GCRO [NLL13]. gear [BSPD14]. Gegenbauer [BCGMB10, ESM13b, ESM13c, ESM13a]. gene [LCN12, XCZ^{+}11]. General [IJMV10, ABIS10, AMÇ13, BLR10, BJ14, BF10, CB13, CK13, CL14a, DP14b, Emm11, EX14, Fuk11, Fuk12, Fuk13a, Fuk13c, HL14, KH10, LB14, LJE^{+}10, MWWY14, PJ12, QYZZ14, QZX^{+}11, VVV10]. generalisation [WM12a]. generalised [CC12]. Generalization [DR14a, HCWsO13, ASM13, AAM14, ADCC12, ÁE10, NH13, PM10]. Generalizations [VL13, ÁE11b, DI12, Pir13, Pir14]. Generalized [AY13, FBB14, FPG12, HCQ14, MNQ14, PL10c, RE10, Vel11, ZyG14, AM10a, AF10, Ali11b, Ali11a, ÁNA14, Ars11, BJRS12, BHBP10, BGP14, CS12, CJY10, CYJ13, Cao13, CL11a, CK11, CZL14, CP13b, Che13b, CR11a, Cui10, DW10, DP12, DP14c, FJP14, FS12b, GUBS13, HKK10, HHZ11, HY11, ISN10, KÖ11, KDG^{+}13, LTW10, LZL11a, LLP12, Li14, lLfD10, LHY10, Liu12, LG11, Lu13b, LdHA10, MS11b, Miy10a, Miy12, MM11b, Nie13b, NNO12, PY10b, PC11, PMMS14, PMC11, PZ13, QT13, QY10, Reu10, Ros11, SD13, SSP12, SAB14, She10, Śmi11, TLHU10, UBSG12, VDG11, Yan10b, ZHCW12, ZG14, ZYW14, ZZG14, dAGR10]. generalized- [BHBP10]. generate [NY12]. generated [APO14, CGM10, DJ14, HZ11, MKD12]. Generating [BDV10]. Generation [BL10a, AM11, RSV12, SZ10, SP11, XCY12]. generator [BAAS11]. generators [BDR11, BR14c, BFMS11, Che14a, DSL12, FJ10, FO10, Har11, Wik10, YSZ11]. generic [DDV^{+}11, GK11, Van11b]. Genetic [MB11, AEWMES11, HTM14, LGR11, TOF14]. geodesy [JKS13]. Geom [YTFP11]. geomechanics [BM11a, LCCK13, YMB14]. Geometric [AHR14, CR11a, CH11, HS11b, Mor10a, DDV^{+}11, GGLR10, YE14]. geometries [CDS10]. Geometry [SWT12, VdKPB10]. geophysics [EN14]. geosciences [CGRS11]. Gerber [GW14c, JZ10, ZY11a]. Geronimus [GHM10, HV14]. GEW [Ros11]. GFDM [BUG^{+}13]. GH [AK14c]. GH-public [AK14c]. GI [YTFP11]. GI/Geom/1/N/MWV [YTFP11]. giant [ABN10]. Gini [HH13, HH14]. Given [Van11b, WM12b]. glacier [CDTV11]. Global [BV10, BG14a, CSZ11b, KW11, LF11, TOF14, ZZG10, ZWG11, AMMS12, AG10a, AJK12, Ant11, BBD^{+}11, BJRS12, BK12, DJ10, FW10, GLH12, HKK10, HXW11, JW10b, KPBN13, KW10, KS12c, LZ13c, LJ13, PPG14, RMF11, SSW14, SRP^{+}13, WF10, WK14, ZDZ11, ZG14]. globally [DD11, Liu14b, MS13a, YPL10]. glycerol [WYY^{+}12]. GMRES [KMS14, SW11]. Goal [CG14, GH14b, KM10]. Goal-oriented [CG14, GH14b, KM10]. Godunov [KAE^{+}11]. Gompertz [Tor14, YG13, YG14]. good [LL11d, SA14]. Google [CSC10]. Gordon [Cui10, LL14b, MD10b, RGJ10]. Goursat [Oou13]. governed [AK14a, CY11, Fu10, XL11]. GPGPU [FPGMG11]. GPU [DHSW14, FLKW14, HK12, KOV11]. GPU-based [DHSW14]. GPUs [GWKE14, WJJ14]. graded [Cui14a, DLP13, Wu14]. Gradient [CCG^{+}13a, LGMA13, MMSW13, ZZP13, Zim13, ADCC12, And10, And13, BKGMA10, BK14a, BEJ14, CY11, CXH09, CN10, HK12, HC11, JW12, JWW14, KNY10, KS14d, LP13, NY12, Rah14, Rua13, SCFR14, SGZ10, XWW10, XSW12, YSQ11, YWL14, ZJ12]. Gradient-based [CCG^{+}13a]. Gradient-enhanced [Zim13]. gradients [IK11, KS12c]. Gragg [ARV10a, GGSP10]. Gram [Liu11a, LZ13d]. granite [RMTO11]. granular [SBFC10]. graph [ABN10, KF11]. graphs [BL13a, BFS14, Car11, YAKA11]. gravel [HdJL^{+}13]. gravitational [MR10]. gravity [eMA14, BB14b]. great [Khr10]. greater [CG10b, SF12]. greatest [CKM14, WL11b]. Green [Ehr11a, GKL12b, HMVPAR11, PAD10]. Grey [LLL12, KWW11]. Grid [DFG^{+}10, AB10, BFHK12, BHPS12, CL10b, CGRS11, DZW14, DQST14, Dos14, JKV11, Li12, Spr14, Wan13, XN14]. grids [AV10a, Bay14, BS13b, CGS12, CC11a, DFH14, EZ10, FJ14, FWBSBGA11, GGLR10, HZ14, Lam14, LE10, NMBG14, NCCF^{+}10, RSGL14, RP11b]. Gross [CCJC13, KHFC11, WC11]. Grossberg [LF11]. Grossberg-type [LF11]. ground [KL12]. group [eMA14, CCCM13, Liu12, MT14b, ZTS11, ZCC^{+}13]. groups [BFS14, HDL^{+}13, NH10]. Growth [ADGP11, ADGP12, AS10b, COW10, FJ10, HHH^{+}13, Nie13b]. GS [KN10b]. GSOR [AIA10]. GSVD [DR14c]. Guaranteed [ESV10, Ain10, JWY11, LMG13]. Guided [VBP^{+}10, KS14c]. gust [AHM13]. gyroscopic [QC14, YD12].

Haar [AuI13, AuIK14, uIAAF14].

haemodynamics [CDQ14]. Hahn [Arv10b]. Halanay [Bak10]. Halanay-type [Bak10]. half [BGVHN13, Fuk11, Fuk12, Fuk13a, HL13b, LM14, OP12, WM12b, Zha11b, ZX10]. half-explicit [LM14]. half-linear [Zha11b]. half-quadratic [HL13b]. half-range [OP12]. Halley [GS12]. Halpern [WSWD11]. Halpern-type [WSWD11]. HAM [Odi11]. Hamiltonian [ÁD12, BIT11, BB12c, BN14, CSZ11a, GH14a, Gus14, HZ11, Mat10, QT13]. Hamiltonian-conserving [Mat10]. Hamiltonian/skew [QT13]. Hamiltonian/skew-Hamiltonian [QT13]. Hammerstein [AST14, ABG11, BCC12, MA14a]. hand [DL12, DSY^{+}11, HD11, MZL14, ZDZ11, ZG14]. handling [KG12]. hanging [KSA14]. Hard [KMD12]. Hard-coupled [KMD12]. hardening [BSPD14, ZTP14]. Hardy [Kuz13]. Harish [El 11]. Harish-Chandra [El 11]. Harmonic [ALM11, Ave10, Car11, KPA14, AP14b, Ayd14, BDL10, CYJS14, DC10, DMMP10, DCZ10, DJ10, JKS13, Mön10, MMMG10, NY14, WHL13]. harmonics [Ave10, Fie13]. harmony [JK10a, Lay13]. harsh [FLEa10]. Hartmann [KNR12]. Hartree [Bra14b]. harvesting [LCZ11, LYP13, Reb12]. hash [SPLHCB14]. hash-based [SPLHCB14]. Hausdorff [RSS14, SS14b]. having [Dou12, Pas11]. HBT [LLG14]. heat [BPS12, HTLI14, KD10, LBM12, LXX14, LMW^{+}11, LZ14a, MV14, MVMG12, SDK10, Sza10, Wan11b, WJJ14, XW12, YF14, ZS11, ZD10, ZLC10]. heat-flux [MV14]. heated [HL13a]. Heath [VTLF11]. heating [KMD12, MKD12]. heaving [MSP12]. heavy [NR14a]. hedging [AH14b, DPV10, EF14, JWY11, LTZ10, NÖ10]. height [MKK14]. helical [CDS10]. helium [DC11]. Helmholtz [CDKN15, ABHMW11, CWY11, CDKN14, EY10, GG13a, HTH10, KOV11, LEKT10, PAD10, QXW10, WC14, Xio10, ZMZ13]. hemi [WSWD11]. hemi-relatively [WSWD11]. hemitropic [NSZ10]. Herglotz [ZMZ13]. Hermes [MKS12]. Hermes2D [SKK14]. Hermite [ÁE10, ÁE11b, BBK^{+}14, BC10, CK11, CW11a, CM10c, JJK^{+}13, KÖ11, KLK13, MS12, MSWY12, ML14, MS11b, NBBKV10, NTNBV14, PL10c, Rua13, SMDMY10, SW13, SGZ10, ZyG14]. Hermitian [AGT13, Bai13, BGM10, DJ10, EGST11, JWL13, NOS14, PW13, PZ13, QT13, ZM14]. Hessenberg [EGST11, MVV14, MT12, MRT13b, ZG14]. Hessian [AR10a, MML11, Nei14]. Heterogeneous [DCG^{+}14, AJK12, BUG^{+}13, BHBP10, CDKN14, CDKN15, GK14, GPZ10, KA14, KSRM12, MR10, MS13c, YMB14]. Heuristics [LEVG14, SCRB^{+}14]. hexagonal [CH11]. hexagonally [Mee12]. hexahedral [FKNS11, Lam14]. hierarchic [SACS10]. Hierarchical [JPX11, BwKB^{+}14, DDG13a, GJ13, Kuz10, MJG14, SAY13, TCD11, ZZS11]. hierarchically [AALMLM14, TOF14]. Higdon [Dea11]. Higdon-type [Dea11]. High [ADLT13, Bab10, CPP13, CHH10, Cui10, DH14, DF12, DFH14, GM13, JM11, LMCW10, LP12, MOG10, MM11c, RS10a, WB10, ZZS11, AU14a, AD11b, BG10a, BDL11, BKTY14, BDR11, BR14c, CCG13b, CR14, CyYwY11, Cui14b, DDG13a, EO14, EN14, EHR11b, FL10a, yGZ13, HZ13, IO14b, KSLD11, LEKT10, LO14a, LO14c, Oou13, Pag10, PP10b, SBGB10, TTB14, TY11, WW10b, Wil13, ZTC11]. high-contrast [AU14a, Wil13]. high-dimensional [BDL11, KSLD11]. high-field [CR14]. High-order [CPP13, CHH10, DF12, DFH14, GM13, JM11, LP12, MOG10, RS10a, CCG13b, Cui14b, FL10a, HZ13, SBGB10, TY11]. High-resolution [Bab10, MM11c, BKTY14]. Higher [BS13a, BOS12, CXZ14, Li12, MGZ10, RCM^{+}12, SL14b, Tan14a, ZAT11, BL10a, BHL11, BK12, CW11b, CG13, HK14, JR10a, Lia14, LW11, Lui10, MY11, MHL11, NRP13, NCCF^{+}10, RW10, SM14a, SWY14, Sie11, uIAAF14, TS10, VVD^{+}14, qWhY11, WZ10d, Yun11, ZCS11, ZL13, AKT14b]. higher-dimensional [SWY14]. Higher-order [CXZ14, Li12, SL14b, Tan14a, ZAT11, BL10a, BK12, Lia14, LW11, MY11, NCCF^{+}10, RW10, SM14a, Sie11, TS10, WZ10d, ZCS11, AKT14b]. highest [Lak13]. highest-frequency [Lak13]. Highly [WRS14, Dom14, EN14, GK14, GG13a, KX13, KAE^{+}11, SGG11, YMB14]. highly-oscillatory [Dom14]. Hilbert [BM11b, Cai13, CB13, CFH14, DHS11, HS11a, WG10a, XXH14, YKL11]. Hilbert-type [Cai13]. Hilliard [ZYC13]. HITS [Zho12]. hitting [WW12]. HIV [HLWM10, ZLW14b]. HJB [ZSX11]. hodograph [BBK^{+}14, KL11, SF12]. Hoffman [Che13b]. Hölder [CLW14, GP12]. holding [HD14]. holes [BFGP10]. Holland [CC11c]. Holling [AD10b]. Holm [CLS10, DI12, FiMO10, Mat10]. Holm-like [DI12]. homoclinic [HZ11]. homogeneous [BB11c, El 11, JL11]. Homogenization [RL10, GKB10, SKKS12]. Homotopy [BCC^{+}13, GC11, DCJ13b, LHF11, SY11, SBCD10, ZYC13]. Hopf [AMÇ13, CYJS14, CGRLR11, DL13, SLD13, WW11b, XL12]. horizon [AØ14a]. horizontal [DA14]. horn [WUB10]. horn-lens [WUB10]. hot [ZZC11]. Householder [pWqSmZ14]. HSS [LHL11, CJ10, CLW14, GT09, Hua14c, LLP12]. Huang [CFH14]. huge [CGRS11]. human [BAdS^{+}10, BAdS^{+}12, WLL^{+}11]. hunting [Ven10]. Hurst [YÖVAYOW14]. Huxley [AF10]. Hybrid [CW11b, SZJ12, AS10a, BW14, Beh14, CZWD14, CXL10, CGS11, DEP11, Dha14, Ism14, JK10b, KMHM14, KPW10, Lay13, LZL11b, MSG11, NMS10, RD10, SAY13, SASH11, TW14, WSS14, YZ11a, ZTC14, ZGZW10]. hydraulic [SRP^{+}13]. hydrogen [KS10a]. hydrogen-like [KS10a]. hydrostatic [BB12b]. hyperbolic [CL10b, CP13b, ĆRR^{+}10, CRF14, CyY10, CwYyY11, CyYwY11, EJ10, GZ10, HK13a, JR10a, JV10, KP14b, KM10, LZ13d, PC11, Ren10, TLNE10, YY11, ZL10]. hyperchaotic [LYP10, PL11]. hypercomplex [CGK10]. hypercubes [DDM10]. hyperelastic [CR11a, Kro10, RG12]. hyperelastic-rod [CR11a]. hypergeometric [AT10, DJJ13, LP10c, MGZ10, ÖÖA11]. hyperplane [WLZ11]. hyperplanes [LJMG11]. Hypersingular [FMP14, Kre95, Kre11, Mah14, Yan13]. hyperspherical [Ave10]. Hypersurfaces [Esc14b]. hyperthermia [SM14a]. hyperviscoelastic [LP10a]. hypothesis [AH10b]. hysteresis [FLHS^{+}13].

IAM [AGK^{+}14]. ICACM [AGK^{+}14]. IDEA [BNY14]. ideal [Col13]. ideals [ZW11a]. idempotent [CID10]. identical [SB10]. identification [LBM12, MM10, WYY^{+}12, YP10a]. identify [PGGCGF14]. identifying [DY11]. identities [GQ14b, XC14]. IDR [DSY^{+}11, DSZ11, TS10]. IDRstab [AAI14]. IGAOR [LJCL11]. II [AHAH11, AKT14b, BFP11, CXZ14, CHH10, FKL10, Ism14, JS10, KO13, LWH^{+}11, NH13, Rum13b, SHZ12]. III [AD10b]. IIIA [AK11]. IIIB [AK11]. ill [BJRS12, BNRZ12, CHL14, DR14c, HPR12, HR10, HR12, Liu14b, LFWL14, MRS10, Ram10, RS10b, TKW12]. ill-conditioned [BNRZ12]. ill-conditioning [TKW12]. ill-posed [BJRS12, CHL14, DR14c, HPR12, HR10, HR12, Liu14b, LFWL14, MRS10, Ram10, RS10b]. illumination [ZZP13]. Image [DDG^{+}13b, VBP^{+}10, ZSK^{+}11, AMRS12, AMR13, BL13b, BEJ14, CC11b, DDE14, HL13b, LGMA13, LZSL11, PY10a, WSK^{+}11, XCQ14, YSZ11, ZZSW11, ZZBX14]. Image-Guided [VBP^{+}10]. images [CCS11, CRS12, JMV13, ZZP13]. imaginary [ZRK^{+}12]. imaging [ACF^{+}14]. imbibition [EASS13]. IMEX [HHKK14]. Immersed [LLSW11, BY14, LZ12b, LZ14a]. immersion [XHZ12]. immiscible [MS13c]. impact [DFG^{+}10, GPKV10, KS12b, NRP13, VMAVGCM13]. impatient [WK10]. impedance [GKL^{+}12a, HHT12, HLFP14, PAD10, ZSD10]. impinging [KD10]. Implementation [GVM^{+}12, GZ11, HCM14, ABHMW11, BDJ^{+}11, BIT11, CCG13b, GGLR10, KOV11, LTZ10, MR11]. implementations [CFMR12]. Implicit [CSY10, GKT11, HJ11, SM11, BLKZ10, DP14a, HL10a, KW11, LL13a, LT13, MS13a, MF14, MMS10, MT13, NNB13, SLW14, Wei11, Xie11, YESD14, ZAT11, Cui10]. Implicit-explicit [SM11]. implicit/ explicit [HL10a]. implicitly [BB11c, NY14]. implied [GS14, LS14]. Imprecise [Sto11a, MB11]. Improbable [Tez14]. Improved [AH12a, MML11, WZ12, WZ11a, ZD14, KP10a, KD10, LSXZ11, MA14a, Mee12, Owa12, Sch14a, Sch14b, uIAAF14, WHLC11, WH13, Xie11, Yan10a, Zha13, Zha14b, ZZG14]. improvement [Kha10, MDVB13, RLGGGÁM14]. Improvements [PNCC10, Hos11, PS14b]. Improving [AEGB10, Cie14, VMAVGCM13]. impulse [AP14a]. impulses [HZ11, LGZ10, Tan10]. Impulsive [WZ10b, Alz10, DF14, Gus14, LS11, LL13b, NPTH09, RHS11, SZS11, Tar10, TSC10, WZS10, Yan11, ZFG10, Zha11a, ZF14]. incline [Ehr11a]. including [DCGS13, FM11b, MRD13, PC12b]. inclusions [GMZ^{+}10, pLjH10, PY10b, RHS11, jTW14]. income [GW14c, LB14, ZY10]. incomes [ZwGhX14]. Incomplete [LZ13d, Vel11, AGT13, DH10, FV13, Fuk11, Fuk12, Fuk13c, SSF^{+}11]. incompressible [AHLJ14, BSB14, FWBSBGA11, GH14b, HZ14, Lam14, MS13c, ORP14, PKH14, SSP13, ZH10b]. incorporated [ZTS11]. incorporating [HD13]. incorporation [DDG^{+}13b]. Increasing [BGIN13, CTV13]. incubation [Ven10]. indecomposable [CGRLR11]. indefinite [AGT13, Bai13, LZ13d, MV12a, WL14a]. independent [AMV13, PDE^{+}13, SACS10]. Indeterminate [AC14c]. Index [Ano10a, Ano11a, Ano11b, ADA13, ASG11, CHL14, GSGN11b, IO14b, LWH^{+}11, Nai11]. indexed [EF14]. indexes [AB14]. indices [Ery14, YAKA11, YP10a]. indirect [LEVG12]. individuals [AELH10]. induced [Gre10, HDL^{+}13, JY13]. Induction [MKD12, BSPD14, DKU14, MVMG12]. induction-conduction [MVMG12]. inequalities [AW10, BLKZ10, CW11a, FW10, GCC14, HHL12, LZ11, Ma10, MR12, Mol10, NASNY11, NNB13, PS14c, ZH10c, Zha11c, ZLL10b]. inequality [CB13, CV10b, DW10, JW10b, LK13, LTW10, LLW14, WG10a, YKL10, ZLL10a]. inertia [WZMC11]. inertial [ZBZ10].

Inexact [Bog14a, GPHAPR14, AR11, AK14b, BLKZ10, CJY10, FGO12, GS12, HPR12, Jia12, RA10, RX10, SS14a, SW11, WLD14, ZYW14]. Inf [CXZ14]. Inf-sup [CXZ14]. infected [AELH10]. Inference [AHAH11, Ism14, WSS14, DCGS13]. infimal [BGRS12]. infimum [VAZS13]. Infinite [AØ14a, CCDS14, CGM10, FJ14, FGO11, GM14a, HLC11, HC14, IK13, KHKM14, LL13a, LF10, LS11, LLT^{+}14, NH13, Oou13, PJ12, RHS11, WWG10, Zha11c]. Infinity [Obe13, BL13b, yGZ13]. inflation [EF14]. inflation-indexed [EF14]. Influence [GRZP10, LCCK13, ÁD12]. information [AEGB10, AR10a, CFGE10, DH10, DCGS13, Dom10, HD14, HM10, KS10a, LZT13, Wri10, ZM13]. infrared [LA10]. ingots [MKD12]. inhomogeneous [FPP14, Kro10]. Initial [MCS12, AGM12, AU14b, BS11a, CXL10, DAE13, EGH12, KP14a, KPKK12, KKN11, Mui10]. initial-value [DAE13, KP14a]. injector [ZZC11]. Inner [MK10, AB11b]. Innovative [BB10b, PDE^{+}13]. input [DR14a, Kho11, LLL12, Zho11b]. input-output [LLL12]. input-to-state [DR14a]. inspired [Lay13, PGGCGF12, PKH14]. instability [YESD14]. instantaneous [GU10]. instruments [Fie13]. insulating [HW12]. insurance [Emm11, Sez10, YLS12]. insurer [LRZ14]. integer [CG11a, YT12]. integrable [Lan12]. Integral [Qi14, SZS11, ABS11a, AAM14, AE11a, ASST10, ASS12, AST14, AM14a, ÁECS12, AAD13, AAD14, AuI13, AuIK14, BH11, BMMR11, Bar13, BB11c, BGGG^{+}11, BPGP11, BCC12, Boy11, BI12, BL11, Cai10a, Cai13, CZ12b, CGM10, CDK10, DSN14, De 14, DMR13, DCJ13b, ESM13a, EALA11, EHR11b, FMP14, Fat13, Fuk11, Fuk12, Fuk13a, Fuk13c, HTLI14, HNK10, IR11, IK11, IJMV10, JX10, JMS10b, KC10, KKPV11, Kre95, Kre11, LWXW12, LZ13a, LP10a, LFWL14, Mah14, MA13, MMS13, MA14a, Moh13, MMMG10, NSZ10, NLO13, OMS10, OA14, Oou13, OPPV10, PN11, PLN13, PM13, Pas11, Pir13, Pir14, PGH13, RD10, RVN13, SYİ11, uIAAF14, TM11, Tar10, WWC14, WW14, WDZ10, Wu14, YZ11b, YSW12, ZFG10]. integral-algebraic [BL11]. integrals [BG11, CN11, CIPS14, Dom14, EJJ11, ELA10, EL11, Fat11a, Fuk13b, HTH10, JJE13, KXH13, KX13, LZY10, LWY10, LP10b, Prz10, Prz13, QZX^{+}11, Sid14, WX10, Yan13]. integrands [CA14, KXH13]. integrated [AÖY10, BDV10, HTM14, KFR13]. Integrating [AEWMES11, CCG13b]. integration [AALMLM14, ASST10, AS12b, BSB14, BHBP10, DHR12, ESM13c, ESM13a, GP11, GPHAPR14, GH10b, JX10, KA12, LCCK13, Li12, LM14, yLGP11, MM12a, MT14b, NS13, Not14, Prz13, SL14a, SASH11, VVD^{+}14, WW10b]. integrations [ZY14a]. integrator [KLNS11]. integrators [BBP14, BP12, Gon10, KHFC11, KV12, LT13, RG12]. integro [Beh14, BGL13, CA12, CZ12b, CC14b, DS12a, DZZ14, ESM13a, EDP14, FID11, JM13, LSD11, LX12, LW10b, LL13b, MR13, NS10, SH10, uIAAF14, SZS11, Wan10, WZ10b, WLZ14, ZFG10, Zha11a, ZAWA13]. integro-differential [Beh14, BGL13, CA12, CZ12b, CC14b, DS12a, DZZ14, ESM13a, EDP14, FID11, LSD11, LW10b, LL13b, MR13, NS10, SH10, uIAAF14, SZS11, Wan10, WZ10b, WLZ14, ZFG10, Zha11a, ZAWA13]. integrodifferential [Yan11]. intelligent [KFR13]. intensity [HZR14]. intensive [DCG^{+}14]. interaction [APO14, CPP13, CDQ14, DBHV10, FHPH^{+}13, Hui10, JMS10a, LVD^{+}14, Mön10, NTPVNHLX14, PP10b, YZ11a]. interactions [FLEa10]. interactive [LJE^{+}10, WJJ14]. interclaim [ZY11a, ZwGhX14]. interclaim-dependent [ZwGhX14]. Interest [CRS12, AH14b, BÜ14, GL10, YYL^{+}10, Zhu14]. Interface [GN11, BY14, BJ12, DBHV10, KP11b, LZ12b, LZ14a, Lui10, MJ14, MMMG10, PTH10, PHA14, PTSK13, SGO14, XZW14, ZS11, Zho11b]. interfaces [Dek10, KS14d, LZ14a]. interfield [BHBP10]. Interior [EGBS12, KC11, Ku13, MWWY14, ADG11, BOS12, DH13, GQL14, GG13b, HCM14, HLSX11, LL11d, LG11, Més12, MRS10, PHS14, ZLL10a, ZW12]. Interior-point [EGBS12, LL11d, MRS10, PHS14]. intermediate [CSY10]. intermittent [EN14, HL10c]. internal [ZZC11]. International [SKS14, ABÇ^{+}11]. Interoperable [KWXZ14]. interpolant [Bar13, BDH11]. interpolants [ABIS10, ASST10, BGIN13, DS10a, ILMS10, LMT13, Van10]. interpolated [Tan14a]. interpolating [BCR13, MSWY12]. Interpolation [CK14b, AW10, AB11b, BISS10, BGIN10, BBK^{+}14, BLM11, BBD13, BC10, BK14c, BM11b, BDR11, BR14c, BF10, CD10, CLLX11, CW11a, CW12, CCH12, CW14a, CT10, CL11f, CBGVPP10, DS13b, DAE13, FZ11, GW12, GW14d, GSV11, HK14, JJK^{+}13, JR14, KLK13, KPR11, LX12, LCXL11, LB11, MS12, ML14, NRS11, OR14, PTH10, QYZZ14, SW13, VCA14, WW11a, XZL11, Zha14b, ZL13]. interpolations [BCC12]. interpolatory [ADLT13, BCR11, CCS10, DGW10, Flo11, JŽ11, KP13, LLL^{+}11]. Interproximate [LZ13b]. interruption [ZH10a]. intersection [HwKB14, LHF11]. intersections [Kim12]. interval [Boy10, CG12, CG11a, CIPS14, CBGVPP10, ELA10, EL11, GPO14, HD11, HDT11, KHKM14, LWW13, Liu11c, LLW14, OORO12, PS11c, WWG10, Yun12, ZCH10, ZD13]. interval-valued [Liu11c]. intervals [FJ14, LS11, MY11, NH13, Sch11]. Introduction [BDM^{+}14, CI13, FHMZ10]. intrusion [WDG^{+}12]. intuitionistic [MS11c]. invariant [NF14]. invariants [ÁD12, BI12]. inventory [GU10, HD13, SJA12]. Inverse [BJR^{+}12, LBM12, MVV14, Che10, CL11d, DY11, ENS14, FLHL14, FPGMG11, Fuk13b, GCH13, GCH14, HS10c, HY11, IK13, KS13, KKN11, Kro10, LL14a, lLfD10, LCI11, MMSW13, PS11b, PS14b, QT13, QC14, QY10, pWqSmZ14, XGL12, XW12, YL11b, YF14, YGHH14, YD11, YD12, ZFY11, ZMZ13, ZBC13]. inverse-free [QY10]. inverses [LHY10, MT12, SS14d]. inversion [Fuk13c, MRT13a, MT10a, NOS14, NA11, Tyr10, Zha13]. inversion-free [Zha13]. invert [GWX10]. invertibility [CID10]. inverting [MRT13b]. Investigating [BCHL11]. investigation [EASS13, GI10, LTFL10, NH13, ZLC10]. investigations [DAW11]. investment [AV12, CO13, HD13, LRZ14, XYW14]. investors [AÖY10]. invexity [SL14b]. inviscid [eMA14, SSP13]. involving [BCGMB10, BHMW13, DDDSC14, fLxX12, ML10, PY10b, QGG10, jTW14, XC14]. ionic [PSKA14]. ionization [BFHK12]. irreducible [LZI10]. irregularly [AB11b]. ISE30 [IO14b]. Ishikawa [PS11c]. Isogeometric [MPS11]. Isospectral [CL11d]. isothermal [BSKL12, KCBA13]. Isotropic [MP10b, AC14a, ALM11, AM11, Bir11, Mau10, SZ10]. isotropy [BFS14]. Issue [FHMZ10, Gra11, BDM^{+}14, MYZ10]. issues [FPGMG11, FLLT10]. Itô [Prz10, Prz13, ZAT11]. items [GU10]. iterated [FBB14]. iterates [Bog11]. iteration [AU14b, BBS14, BPGP11, BL13b, Cha10b, CJ10, CyY10, CwYyY11, ENS14, FF14, GT09, HVV10, HC11, Hua14c, Jia12, LL11d, LFWL14, LHL11, PZ13, SSP13, SH10, Sun14, WHLC11, WSWD11, WLD14, pWqSmZ14, ZC10a, ZZ11a, Zha14a, ZGZW10, ZY14b, ZM14, ZZ11b]. iteration-by-subdomain [Sun14]. iterations [FBB14, GKLZ12, Zie14, PS11c].

Iterative [CZ12a, CM10a, MD10a, PS11b, Yun12, ABG11, AH12b, CJY10, CSY10, CL14a, CTV11, CTV13, CHMT13, EGON10, EHR11b, GSGN11b, GSNA13, GSNDB14, GHR10, HSH12, HL10b, HL13b, bJfCwH14, Koç11, KPW10, LsGHZ10, LHY10, NK10, PR12a, yP10b, PS14b, PM11b, SAY13, SC10c, SG12, yS05, WL10b, WG10a, WZA10, Wan11f, Wan12a, WL14b, XSH10, pYXPH10, YLD13, YL14, ZSX11, ZZ11a]. Ito [OAHR11]. iTraxx [Nai11]. IVPs [DL12, SASH11].

J [ÁD15, Asl10, BAdS^{+}12, CDKN15, Dai11, DC10, DD14b, DS13c, Gün10, Kre11, LB10, LHL11, TSC10, XSV13, ZJ12, vDvFZ13]. j.cam.2009.08.041 [DC10]. Jacobi [Ali11a, LW10a, yLGP11, XN14]. Jacobian [AWF11, FYA^{+}14, Gat14]. Jacobian-free [AWF11, FYA^{+}14]. Jakimovski [BTSA14]. Janowski [KPA14]. Jarrow [VTLF11]. JAVA [AÖ11]. Jean [Khr10]. jerk [MT14a]. Jesús [MYZ10]. jet [KD10, KS12b]. Joint [ZZ13, DKU14, fLxX12]. Josephson [VMS12]. journals [BDR10]. Jump [Wei11, APW14b, BR11b, CLMR14, HGW13, HZR14, JWY11, KPsQ11, YSW13]. jump-diffusion [APW14b, BR11b, CLMR14, HGW13, JWY11, KPsQ11]. jumps [BBMY11, HS11b, LYZ11, OGM14, PO11, TW14, YZ10a, ZYL10, mZkPkL11]. junctions [VMS12].

Kalman [KK14]. Kantorovich

[ABCP14, AH10b, AR11, AK14b, Zha11c]. Kantorovich-type [AR11]. Kapteyn [Dom11]. Kawahara [Asl10, Ass09, Kud10]. KdV [DI12, Sie11, YNO10]. KdVB [TA10]. kennel [LL11d]. Kepler [GWKE14]. Kernel [AY13, BM11b, BLR10, AST14, BH11, CGK10, DZZ14, EGBS12, EY10, GA12, GC11, HM11, JY13, LW13, MM11b, PM13, PS11a, PHS14, RD10, SD13, SYİ11, Wu10, YSW12, Yan13]. kernel-based [PS11a]. kernels [AAD14, CC14b, JM13, JMV13, PS13, PGH13, Wu14]. key [AA14b, AK14c, NACLR12]. keyboard [MKK14]. kiln [KBKR14]. Kind [Vel11, ABS11a, ASS12, AAD13, BGGG^{+}11, CN11, FZ11, Fuk11, Fuk12, Fuk13a, GQ14a, HNK10, IK11, KHKM14, LWXW12, LZ13a, LL11d, Mah14, OMS10, PM13, SP11, XL12, YZ11b]. kinds [Fuk13b, XLCC10]. kinematic [ZTP14]. kinetic [BG10b, HM10, WYY^{+}12]. kinetics [SB10]. Kirchhoff [BySZZ13, BDyS14, GH10b, RT12]. Kirsch [BFLP12]. Klein [RGJ10]. knapsack [ARF14]. knot [Sla14]. knowledge [ZTS11]. knowledge-incorporated [ZTS11]. Kolmogorov [CHRZ13, FN10, GG13b, YFWH10]. Kolomogrov [KS10b]. Koornwinder [FPP12]. Korovkin [Akt14a]. Korteweg [HX14b, RTL14]. Kraft [SH13a]. Krall [FPPX10]. Krall-type [FPPX10]. Kronecker [CCS10]. Krylov [AMRS12, AWF11, BGPPRW14, FYA^{+}14, IHRA11, KOV11, Liu14a, NL10, QY10]. Krylov-methods [BGPPRW14]. Kummer [DST10]. Kuramoto [GP10, VMMM10]. Kutta [FG14, LO14c, SCRG12, BR11b, BB12c, But12, CFMR12, DP14a, DR14a, FBB14, GPHAM10, HVV10, HHKK14, HSY13, KKR11, Kha10, KB11, LL14a, LO14a, MK11, NTNBV14, PR10, Xie11]. Kuznetsov [NNO12].

L [DA14, DMB10]. L-shaped [DA14, DMB10]. laden [GKT10, GKT11]. lag [Bak10, SCRG12]. Lagrange [Ger12, NRS11, OR14]. Lagrangian [BBKS10, CCS11, Che14b, CRF14, RMF11, RYM^{+}14, XMC11]. Laguerre [BLO11, CIPS14, DM10b, FS10, SMMD11, qW11, ZX10]. Lambert [Fuk13d]. Lanczos [Kni10, NY14, TL13]. Lanczos-type [TL13]. Landau [BPS12, JK10c, PS14c].

Landweber [ZC10a]. Laplace [AA12, BFDP13, ÇU14, Dos14, Dou12, KOV11, Li11a, MT10a, ML10, NA11, Obe13]. Laplacian [AM10b, BL13b, CCG13b, KF11, KLL11a, KLM14, WZ10d]. Large [BKTY14, GJCL11, LV13, WCZ11, AG10b, BGPPRW14, BFG^{+}14, BHJ11, BGKVW10, CD10, CXH09, DHSW14, DdS14, DSL12, Dra13, DS11, DJ10, FV10, GGH^{+}14, GG13a, KR10, LwCLW13, LZ13d, LP10c, MRS10, NL10, PR12a, SW12a, Sez10, YPL10, ZJ12, ZGZW10, ZM14, FDD10]. Large-Eddy [FDD10]. large-order [DSL12]. Large-scale [BKTY14, LV13, AG10b, CXH09, DdS14, GGH^{+}14, LwCLW13, YPL10, ZJ12]. largest [LZI10]. latent [DJM10]. lateral [PDE^{+}13]. Lattice [BCD^{+}10, CMR11b, Har11, HBE14, HHLW14, LZ13e, SH12, SL12, Tan14a, WSJ14]. Lattices [JKK^{+}10, Vit10, BLR10, CH11, GM14a, KP11a, SKAM11]. Laurent [BHJN10, DMGVPN10]. Lauricella [MGZ10, Zho11a]. law [BGP11]. laws [AGJ13, HK13a, LP10a, MSZ13, SR13, YNO10]. Lax [CET14, Lan10, Li12]. layer [CP13a, CHH10, GQL14, GD10, RL10, SZ10, Sun14, TA13, VT14, YESD14]. layered [KP11b]. layers [CHH10, Li11b]. layout [ZBZ10]. leading [Bai13]. learning [BGS12, LMM^{+}10a, LMM^{+}10b, NA11, QTS14, SH13a, SGZ10]. Least [BSV10, GLHCLP11, Ku13, WWC14, XHmDW14, AG10b, CC11a, HPL^{+}14, JKS13, KH11, KNY10, KP10b, LV13, LZ12a, LZ13d, LEVG12, MPC10, Nie13b, SD13, TR11, Tor14]. Least-squares [BSV10, GLHCLP11, Ku13, CC11a, KH11, KP10b, LZ12a, MPC10, Nie13b, TR11]. Lebesgue [BDH11, BSV12, CA12, PS14c, Prz13, RSV12, Zha14b]. Lebesgue-like [RSV12]. Left [BLTW10, lLfD10, ZZ12]. Left-definite [BLTW10, ZZ12]. leg [LW10b, YW10]. Legendre [CT11, DSN14, DHS11, EJJ11, LSD11, LYD13, NLO13, Odi11, PN11, PLN13, ZX10]. lemma [LHLL10, YNO10]. lemniscate [Boy11]. length [AHR14, EI14]. lengths [SMDMY10]. lens [WUB10]. Leon [XSV13, XSV12]. less [HJ11]. Letter [KN10b]. level [BW14, BK14b, CG10a, GW14a, GGS12, HPR12, KSA14, Kuz14, LFWL14, PM11b, TNN10, XN14, ZM13, ZY14a, BHMW13]. Levenberg [Zho13]. leverage [IO14b]. Leviatan [BTSA14]. Levin [Luo10]. Levinger [AM10a]. Levy [NR14b, BM10, EY14, IK13, LL13a, LS10, RE10, ZRW11]. library [KWXZ14, MKS12, SKK14]. lid [DA14]. Lidstone [AW10, CDD12]. Lie [eMA14, Liu12, MT14b, NH10]. Lie-group [Liu12]. life [Ism14]. lifetime [ADA13, Ery13a, LWH^{+}11]. Lifshitz [BPS12, JK10c]. light [BCM10, NA11]. like [AMMS12, AH13, BDD^{+}10, DW10, DI12, FGO12, FWBSBGA11, GST12, KS10a, MA13, MMX10, PY10b, RSV12, RBD11, RBD14, Rom11, Zha14a]. likelihood [BAU13, BA10, MD13, Tor14]. Limit [PL10a, PTP10, ZWZ11, ÁNA14, BRZ10, CR14, CGS13, KG12, LFY11]. limited [KM10, VL13, Wri10]. limited-memory [VL13]. limiter [GZ10, Kuz10]. limits [HW10, LLL^{+}11]. Line [BI12, KK13, AV10a, BGVHN13, CBGVPP11, CWQ11, MBS10, OR14, Zho13, ZX10]. Linear [CG10a, CG12, EI14, IO14a, LZ12b, AS10a, AAI14, AMÇ13, AMPR10, AS11, ASK11, ÁECS12, AS12b, AIA10, AC14b, AKT12, AKT14b, BBP14, BY11, BS10a, BGRS12, BB11c, BK14c, BP12, BKTY14, BJ14, BNRZ12, Büy12, CK13, CS10, Cao14, CLS14, CI13, CRM10, Dea11, DH11, DS12a, Dra10, DJ14, DSY^{+}11, DS12b, DS13c, DR14c, EGBS12, EK12, Fad10, FJ10, Fat11a, Fat13, FJP14, FV10, FSB10, GW12, GLHCLPN10, GLHCLP11, GPVL10, GP11, GG10, GT09, Gwi13, HZ11, Han11, HHX14, Har11, HD11, HR10, HR12, Hu14, HS12, Hua14b, IK11, IJMV10, JS10, JWW14, KS12a, Kho10, KL10a, KKN11, KN10b, LLS14, LJCL11, LLP12, LHLC13, LW13, LWW13, LL11c, LJ12, LHL13, Liu14a, Liu14b, LLW14, LC10, LHL11, MHL11, Man11, MJ10b, MZL14, MM11a, Mol10, MPC10]. linear [MMMG10, NF14, NMS10, NLL13, PW12, PT11b, PT12, PHS14, Pul14, RST14, Rum13a, Rum13b, SA14, SSP12, SC10c, SZJ12, SCDA10, SSF^{+}11, TL13, WWL10, WH10, WH13, WL14b, YT12, YSW12, ZTS11, ZDZ11, Zha11b, ZG14, ZYW14, Zha11c, ZGZW10, ZY14b, ZTP14]. linear-growth [FJ10]. linear-phase [Han11]. linear-quadratic [BP12, CK13]. Linear/linear [IO14a]. Linearity [Kuz12]. Linearity-preserving [Kuz12]. Linearizability [CHRZ13]. linearization [CK11, GSM11]. linearized [APJ10, CC14a, DJM10, HX14b, IHRA11, Wan11d]. linearly [KW11, MF14]. linearly-implicit [KW11]. Liouville [BAK11, Cha10a, CO11b, CSZ11b, GCH13, JTV14, Liu12, MRD13, RB13, Reu10, ZM10, ZZ12]. Lipschitz [CSY10, KS12c, MS13a, SS14b]. Lipschitz-continuous [CSY10]. Lipschitzian [pYXPH10]. liquid [FGEM10]. liquidity [JM12]. Little [BLO11]. Littlewood [Kuz13]. Liu [DAH11]. Liu-type [DAH11]. LMI [LF11]. loading [MMMG10]. loads [WHLC11]. Lobatto [AK11]. Local [EJ10, Fer11, FGO12, KV12, NRS11, PPG14, RSGL14, Rom11, SLP11, WK14, AMV13, AH12a, AH13, AKT12, AKT14b, BDSG10, BCR13, CHRZ13, CXZ14, DCGS13, DS14, DS10b, FPP13, GSNA13, GSNDB14, GM10, GM13, GLLM12, HKK10, HX14b, KMD12, Li12, MD10b, RA10, RT12, SSW14, ZZP13, ZY14a, ZLC11]. Local-global [PPG14]. Localized [LWW13, FDH10, Lak13, SKAM11]. Locally [APJ10, FKNS11, KF11, ZTC11]. locking [CRM10, GCY13, HKO11, YC10]. locking-free [CRM10, HKO11, YC10]. LOD [Wan11e]. log [AP14b, Ayd14, KH13, LTZ10, MSWY12, PRT10]. log-aesthetic [MSWY12]. log-barrier [LTZ10]. log-elliptical [KH13]. log-harmonic [AP14b, Ayd14]. logarithmic [AAD14, Dom14]. logarithms [CL10a, WZ12]. logistic [MSW14, Pir13, Pir14]. logistic-normal [Pir13, Pir14]. logistics [SCRB^{+}14]. Long [LTX12, SC10d, Wan13, BG11, BBD^{+}11, Cai10b, EN14, GPKV10, HHP^{+}10, Hui10, JW10a, MM11b]. long-range [Hui10]. Long-time [Wan13, BG11]. long-wave [Cai10b]. longitudinal [JL11]. lookback [KPsQ11]. loss [EC13, GSK11, VAZS13, ZY10, Van11b]. Loss-Given-Default [Van11b]. Lossless [AMR13]. Lotka [TSC10, NPTH09]. Love [Pas11]. Low [BG10a, BB12c, DLWZ14, YLD13, AÖÖ14, AM14b, BC12, CFMR12, CLP10, DLG14, Dax10, DMMP10, EI14, Gat14, JJK^{+}13, JWW14, KPV14, LZH14, MU14, MDVB13, TA13, UM14, WH10, WH13]. low-frequency [CLP10, TA13]. low-Mach [DMMP10]. low-order [AM14b]. low-rank [Dax10, Gat14, JWW14, MU14, UM14]. lower [LLX14, WZS10, WLZ14, XMSS11]. LP [XCY12]. LQP [LZ11]. LQP-based [LZ11]. Lü [PL11]. lumped [pZxY12]. lumping [ZD11]. Lundberg [Zhu14]. Lupas [HCQ14]. Lyapunov [AAY11, Büy12, PWZ11, YKL10]. Lyapunov-type [YKL10].

M [KKK12, KKK12, WYP10, ZH10a]. M/G/1 [WYP10, ZH10a]. M/M/m [KKK12]. Macaulay [BDD14]. Mach [DMMP10, MDVB13]. machine [BGS12, LMM^{+}10a, LMM^{+}10b, PC12b, QTS14, TNN10, Wu10]. machines [NGAZ11]. machining [KS12b]. macro [ZTP14]. macro-zones [ZTP14]. magnetic [MKD12, SM14a]. magneto [TKW12]. magneto-quasistatics [TKW12]. magnetostatic [KCBA13]. magnets [MKD12]. magnitude [OB14]. magnitudes [Wei11]. Magnus [BP12]. majorant [AH12a, Fer11, FGO12]. Majorizing [AH12b, CAA10, EGH12]. making [FJLL10, HZ14, KWW11]. malignant [Kni10]. management [FM14, SKAM11]. manifold [PTSK13]. manifolds [ABCP14, FB12]. Mann [PS11c]. Manning [CCG^{+}13a]. manufacture [Wu10]. many [CKM14, Esc14b, Mol10, WWG10, Yun12]. map [MHH^{+}14, NH10, SSP12]. mapped [SWY14]. mapping [AC14a, Ayd14, DL14a, LDGY10]. mapping-collocation [AC14a]. mappings [AP14b, CSY10, CM10a, ĆRR^{+}10, CGK10, HQR13, HLC11, HL10b, JK10b, KPA14, QCK10, She10, WG10a, WZA10, WSWD11, XSH10, pYXPH10, ZYG10]. maps [SC10a]. marginal [fLxX12]. market [Emm11, GS14]. Markoff [Khr10]. Markov [APW14b, BSZ13, CGW14, HS10b, HXSF14, Rie13]. Markov-switching [APW14b, HS10b]. Markovian [EY11, GLHCLP11, LXY13]. Marquardt [Zho13]. Marshall [AKA14, OB14]. martingale [NR14b, ZWZ11]. martingales [Lan12]. Maruyama [CZ13, LXY13, MSG11, TS14]. masks [YSZ11]. masonry [SKKS12]. Mass [KH11, BK14b, Kuz14, YBD11, pZxY12, ZD11]. matched [Li11b, PTH10]. matching [DDV10, GS14, XCY12, ZyG14]. material [GRZP10, Kro10, ZSD10]. materials [BSB14, CC13c, DC10, DCZ10, FdOS12, KSRM12, NSZ10, ZS11, ZTP14]. Math [ÁD15, Asl10, BAdS^{+}12, CDKN15, Dai11, DC10, DD14b, DS13c, Gün10, Kre11, LB10, LHL11, TSC10, XSV13, ZJ12, vDvFZ13]. Mathematical [AVA13, CHMD14, Fie13, FLEa10, LLG14, MR10, SRA10, VAV11, AHLJ14, DH14, KSK11, KPsQ11, LGR11, LZL11b, RYM^{+}14]. Mathematics [AGK^{+}14]. Mathieu [ZRK^{+}12]. MATLAB [AS12a, HSH12, MR11, BGPPRW14, BFG^{+}14]. matrices [AS10a, ADGP12, ADGP13, Bai13, Bel13, BGM10, BCC^{+}13, BVV10, CJY10, Cao13, CL10a, Cin12, Cin14, CM10d, DDDSC14, DR14b, ESM13a, EGST11, FV14, HY11, Hua14b, JWW10, JWW14, KA10, KF11, KS14b, KS13, KMG12, KMS14, fLyHZ10, lLfD10, LCI11, MVV14, MT12, MRT13b, MRT13a, MVC10, PTP10, PW13, QT13, VVV10, VVM10, WH12a, WH12b, XLCC10, YL11b, YGHH14, YBD11]. Matrix [Cas10, CKM14, RB13, AMMS10, AB11a, AWF11, AC14b, BMMR11, BDD14, BR14a, BJRS12, BL10c, CL13b, CFJ^{+}11, CO11b, CH13, Dax10, DdS14, DCGS13, DDM11, DH11, DLWZ14, DP10b, EGST11, FKL10, GH10a, Gat14, GZ11, GG12, GKLZ12, GZ13, Gün10, GKL12b, HX14a, HD11, KR10, LRT13, LsGHZ10, LHL10, LZH14, LZ14b, dL14b, LYD13, MV12a, MMX10, Mir14, NH10, OORO12, yP10b, PWZ11, PGGCGF12, SST12, SC10c, SRWZ14, VVV10, WL10c, WH10, WH13, Win14, XCZ^{+}11, Yar10, ZC10a, yZ11c, Zha13, ZD14, Zha11c, ZY14b, ZLG12, Zie14, PGGCGF14]. matrix-free [DdS14]. matrix-valued [FKL10, Yar10]. max [XMC11]. max-bisection [XMC11]. Maximal [ASK11, ADGP12, BN14, ZYG10]. Maximal- [ASK11]. maximization [BOR^{+}12, LQYL11, Liu11b]. maximize [FTTT13]. maximizing [LRZ14, Zha14a]. Maximum [CS10, lLzS11, MD13, WW10a, Wan11d, AA14a, AELH10, BAU13, BA10, BFLP12, Min10, NO12, Tor14, VAZS13]. Maxwell [DS13c, AC14d, BL10a, BKL14, Cui14a, DS14, DS10b, DS12b, GM10, HL11, Li11b, WHL13, ZML14]. MCMC [Law14]. MDG [AAD14]. Mean [CW14c, Hua14a, PP14, TS12, BF10, CZ13, EY11, Hua12, Liu11c, PO11, SW10a, TS14, ZW14]. mean-absolute [Liu11c]. Mean-square [TS12, CZ13, TS14, ZW14]. Mean-variance [CW14c]. means [HH13, HH14, RPS^{+}11, SS14b, XCY12]. measure [DLSV14, IK13]. measured [fLyHZ10]. Measurement [KG14, GRZP10, LD10, LS14, LZSL11]. measurements [GLHCLPN10]. measures [ABN10, BBB10, CFdAR11, CBGVPP11, GHM10, KS10a, Liu11b, NR14b, SMMD11, ZZC10, dJP13]. mechanical [CA12, KA14, NRS10, VSB12, WL11a, WLL^{+}11]. mechanics [KLM14, Vel10]. media [AH10a, AJK12, BUG^{+}13, BGLV11, BKL14, CC13c, DQST14, EASS13, GK14, GPZ10, LE10, LM13a, MHL11, MOG10, MS13c, PDE^{+}13, RP10, SSW14, SBFC10, SRP^{+}13, TA13, YMB14]. medial [KL11]. median [SST12]. medium [ATS14, COW10, KP11b, PSRV10, Rah14]. Meixner [FCS12]. Mellin [FS10, LP10b]. membrane [SS14e]. membranes [CHMD14, Kro10]. Memoriam [Ano10b]. memory [BS10a, DP12, DP14c, VL13, Wan13, YLD13, ZW11b]. Menke [MBS10]. merging [Lu14]. meromorphic [JLZ14]. Mesh [ZD11, BISS10, BFB14, BFS14, FiMO10, FRB13, HKK10, HJ11, HK14, KM10, Kva11, Li10, RW10, RKS^{+}13, RZD14]. Mesh-dependent [ZD11]. mesh-less [HJ11]. meshes [BSV10, Cao14, CBB13, Cui14a, DCJ13a, Dou12, DLP13, GKU14, HS12, JKS10, Lam14, LQSL11, MWWY14, NTRNXB10, SW12a, Sch14c, SW13, SDK10, TCD11, Wu14, WZY^{+}14, YZ11a]. meshfree [AV10b, BBD13, MMP11]. meshing [RWRS14]. Meshless [MD10b, AAD13, AAD14, DS12a, DS14, FND10, KL10b]. Message [AK14c]. meta [CC13c]. meta-materials [CC13c]. metaheuristics [LEVG14]. metal [PSKA14]. metallic [BCD^{+}10, LVD^{+}14]. metamaterials [Li11b]. metamodel [LZWJ11]. metamodel-assisted [LZWJ11]. methane [GJCL11]. Method [ASM13, AAY11, DDV^{+}11, MNQ14, Sta11, ABIS10, ABS11a, ADG11, ACF^{+}14, Ali11a, AHK14a, AU14b, AC14a, AWF11, AHH13, AS12b, Ant11, AV10b, AD11b, AIA10, AH12a, AHK14b, Asl10, AAD13, AAD14, Ass09, AK14d, AuIK14, BMMR11, BS11a, BSS14, Bar11, BFDP13, BT12, BA10, BFLP12, BY14, BS10b, BG14a, BPGP11, BCC12, BKB14, BLKZ10, BJ12, BDT^{+}13, BGP11, BJRS12, BEJ14, BFP11, BySZZ13, BDyS14, BCC^{+}13, BGP14, Cai10a, CM12, Cai13, CHL11, CZZ14, CZWD14, CCG13b, CA12, CGS12, CL11a, Cha10b, CY11, CCJC13, CZL14, CL10b, CL11d, CL11b, CSZ11a, CLW14, CFF12, CV10a, CMR11a, CKM14, CFJ^{+}11, CC13c, Chu10, Cie14, CGS11, CG13, CvdMS12, CT10, CHMT10b, CRF14, Cui10, CyY10, DSC09, Dai11, DW11, DS13a, DV13, DR14a, DS12a, DAE13, DS14, DO12, Dha14, DCJ13b, DMB10, Dos14]. method [DGC^{+}13, DJ14, DH13, DSY^{+}11, DSZ11, DZZ14, DJ10, DS12b, DS13c, Eba11, EGES14, ENS14, ESM13b, EG11, EDP14, EGH12, EHRV13, FL10a, FLHL14, FMP14, Fat11b, FiMO10, Fer11, For10, FV13, Fu10, FLKW14, GZ10, GW10, GCH13, GW14b, GKU14, GPVL10, Gat14, GUBS13, GC11, GQL14, Gia14, GCY13, GS12, GH10b, yGlJ11, yGZ13, GZ14a, GZ14b, GP12, HPL^{+}14, HQR13, HHX14, Har11, HJ11, HZ14, HL10a, HLC11, HK13a, HML14, HH13, HH14, HBE14, HM11, HHZ11, HL14, HL10c, HYY14, HX14b, ISN10, IK13, IR11, JLZZ10, JWL13, bJfCwH14, JW12, JW10b, JJE13, KH10, KH11, KFR13, KMHM14, KKK14, KHKM14, KPsQ11, KOV11, Koç11, KN10b, Kok11, KS14d, KP10b, Kud10, Lam11, Lam14, LV14, LL13a, LVD^{+}14, fLyHZ10, LsGHZ10, LZ11, LL11a, LT11, LLCL11, LLP12, LHLC13, LW13, LZ13a, Li14, LL11b, LY10]. method [LY13a, LLT^{+}14, LLX14, LTZ10, LMY11, LMYJ12, LMG13, LZ13c, LZ13e, LZ14a, Liu14b, LZ14b, LJE^{+}10, Lu14, Lui10, LC10, MX10, MS11a, Mah14, MRD13, MMF10, Mar11, MSP12, MMS10, MZL14, MN10, MV12b, MD10c, MML11, MSZ13, MRS10, MPC10, MWWY14, NS10, NTRNXB10, NLL13, NBZ13, NY14, NH10, NNB13, NS14, OCU11, ORP14, Pag10, PTH10, PDP13, PHA14, PL10a, PS11b, PS14b, PR11, Pov14, PMC11, QXW10, QYZZ14, QY10, RP10, RK11, RD10, Ren10, Reu10, RBD11, RBD14, RMF11, Ros11, RX10, RR10, SAY13, SCFR14, SD13, SS14a, SSP13, SW12a, Sch14c, SL14a, Sha14, SH10, SY11, SC10c, SGG11, SASH11, SHZ12, SLP11, uIAAF14, SB10, Śmi11, SSR10, SG12, Ste10b, Sto11b, SSK^{+}13, yS05, SZ11, SM14b, Sun14, SGO14, SKN11, TM11, TR11, TBK14, UBSG12, VL13]. method [VS14, WTM12, WW10b, WCW10, WG10b, Wan11e, WHLC11, qW11, Wan11f, WLWH11, WC11, WCH11, WZ12, Wan13, WY13, WWC14, WMGM14, WW14, qWhY11, WZ11a, WH10, WH13, WM12b, pWqSmZ14, WL14b, XZW14, XWW10, Xie11, Xio10, XL11, XW12, XZH14, XCZ^{+}11, XMSS11, XCQ14, YPL10, YJ11, YC11, YSW12, YF14, YÖVAYOW14, YP10a, YSQ11, YL14, YY10, pZxY12, ZLL10a, ZL10, ZFY11, yZ11c, ZHCW12, ZMZ13, Zha13, ZyG14, ZYW14, ZTC14, ZD14, ZFM14, ZY14a, Zha10b, ZBC13, ZW10, ZYL11, Zho13, ZC13, ZCX14, ZLC11, ZLG12, ZFH10, ZW14, dAGR10, BG13, BHMW13, NK10]. method-flux [VS14]. methodology [PGGCGF14]. Methods [AVA13, QLL11, SPC14, SKS14, VAV11, AMRS12, AS10a, AMT11, AM10b, AK14a, AT10, ASST10, ASS12, AST14, ABÇ^{+}11, ADCC12, ÁD14, ÁD15, ABG11, ABCP14, ABS11b, AY13, AR11, AEG^{+}11, AH12b, AH13, AK14b, APJ10, AKT12, AKT14b, BKGMA10, BK14a, BDD^{+}10, BY11, BY12, BBKS10, BCM10, BK14b, Bay14, BGPPRW14, BSB14, BHL11, BDMR10, Bog14a, BKL14, BHJ11, BMN10, Boy13, BJ14, BOS12, BCMR12, BI12, BL14, BR11b, BL11, BB12c, BHBP10, But12, CDTV11, CFMR12, CJY10, CZ13, CCG12, CM13, CLS12, CC11a, CJ10, CAA10, CPY10, CHX11, CZ12b, CC13b, CG14, CXZ14, CXH09, CR14, CM10a, CT11, CH14, CDJP13, CHMT10a, CTV11, CHMT12, CHMT13, CN10, Cui14a, DEP11, DP14a, DP14b, DSN14, De 14, Dek10, DA10, DL12, DXL10, DMR13, DQST14, DP14c, EO14, EGBS12]. methods [EHR11b, EJ10, FID11, FBB14, FKNS11, FKL13, FYA^{+}14, FGO12, FND10, FG14, GS10, GR12, GN14, GSK11, GK11, GPZ10, Gia12, GH14b, GAS14b, GPHAM10, GPHAPR14, GPHAS14, GSGN11b, GSNA13, GSNDB14, GM10, GM13, GT09, HX14a, HSH12, HHL12, HHKK14, HVV11b, HV12, HGW13, HNK10, HGW12, Hua12, HL13b, Hua14a, Hua14c, IJMV10, JK10b, Jeo10, JK13, JX10, JL11, JM11, JM13, JD13, JWY14, JKS10, KKR11, KS12a, Kha10, KL10a, KPKK12, KNY10, KA14, KB11, KB12, KB13, KK13, KSTW14, Ku13, KL14, KW10, KW11, Kuz10, Kuz14, Kva11, Lak13, LTX12, Li11b, LJCL11, LXY13, LZDW14, LM12b, LLSW11, LM14, LZ10, LW10b, LHY10, Liu11a, LL11d, LP13, LO14a, LO14b, LO14c, LHF11, LFWL14, LHL11, MLZ14, MS13a, MD10a, Més12, MM11a, MMSW13, MA14b, MFL13, MJ14]. methods [Mui10, NY12, NSZ10, Nei14, NTNBV14, NL10, NASNY11, NÖ10, Obe13, OMS10, OS10, OAHR11, OPPV10, PN11, PLN13, PR10, PT11a, PT11b, PT14, yP10b, PW13, PPH10, PRM11, PHS14, PAJ10, PZ13, Pul14, PM11b, QCK10, RS10a, Ran14, RB13, RS10b, RZD14, RA10, RSGL14, Sch14a, Sch14b, SLW14, SH13b, SBL12, SBCD10, SLD13, TL13, TP10, TCD11, Tor14, WL10b, WHL10, WWL10, WL10a, WZF11, Wan11b, WD14, WLD14, WA14, Wei11, Win14, WT14, WZ14, XCY10, XCY12, XL12, YZ11a, YFZ14, YW10, YL11c, Yun11, Yun12, ZSX11, ZML14, ZDZ11, ZJ12, ZG14, ZLW14a, ZYG10, ZGZW10, ZY14b, ZM14, ZCS11, Zho11b, ZC10c, ZZ11b, ZX10, AGM12]. metric [ĆRR^{+}10, CFdAR11]. metrics [HK14]. METU [AGK^{+}14]. MHD [ATS14, Rah14]. MIB [XZW14]. Micchelli [PS14a]. micro [APO14, Sun14, ZD10]. micro-porous [Sun14]. micro-shock [APO14]. microindentation [GRZP10]. micromagnetism [CM10b]. Micromechanical [BDMP13, LCCK13]. Micromechanical-based [BDMP13]. microstructures [ACF^{+}14]. microwave [WB10]. mid [DDV^{+}11]. mid-frequency [DDV^{+}11]. middle [ILEE13]. midpoint [WDZ10]. mildly [CZ12a]. mill [SH13a]. Milstein [OAHR11]. Mimetic [SPC14, VVD^{+}14, BB10b]. MIMO [BGKVW10]. min [ZWM11]. min/max [ZWM11]. Mindlin [WLL^{+}11]. Minimal [BC12, AHR14, ASK11, BFGP10, DSZ11, FS12a, IA13, LMYJ12, PGGCGF14, SCFR14]. minimax [GD14, Naz11]. minimization [CC14a, FB12, LZH14, LHC14, LPLX10, Owa12, RBD11, RBD14, SAY13, SS14a, YKL11]. minimize [NGAZ11]. minimizers [Ant11]. Minimizing [BBB10, PVM^{+}14, HHX14, WC14]. Minimum [CG11a, MVC10, Dax10, JWY11, LHC14, MD13, PW13, SW10a]. Minimum-energy [CG11a]. mining [AC14c]. MINRES [Jia12]. Mirakjan [Mah11]. Mises [ZTP14]. missing [ÁAG^{+}13]. mistuned [ZTL13]. Mitscherlich [Nie13b]. Mixed [BM11a, MT14a, AM14a, AB10, AM14b, BP11b, BGL13, BLKZ10, BFP11, CM12, CB13, CPY10, CW14b, DW10, DCJ13b, DS12b, DS13c, FFM^{+}11, FND10, GSM11, GTS14, HK11, JL11, KAE^{+}11, Kur10, Lam11, Lam14, LLS14, LF11, LTFL10, LWZ13, PL10a, PY10b, RMP^{+}13, RBK11, SZS11, SM14b, WZ10a, WHL13, WZY^{+}14, XCY10, YJ11, dFGAN11]. Mixed-mode [MT14a]. mixed-type [LTFL10]. mixture [GMRK14, HW10]. mKdVB [TA10]. ML [Yeu14]. MLPG [MD10b]. MMOC [CL11b]. MMOCAA [CL11b]. mobile [RYM^{+}14]. modal [OS10]. mode [CFH14, HK13b, MT14a]. Model [AG10b, AHAH11, AALMLM14, AHLJ14, Alz10, ABS11b, AC14d, hBW10, BÜ14, BA10, BSB14, BAAS11, BBD^{+}11, BGKVW10, BJ13, CSL^{+}11, CM12, CR12, CP13a, CL11c, CO13, Che14b, CGW14, CSC10, DMMP10, DME^{+}12, DN13, DL13, DH13, DF14, Ery13a, FCS12, FGO11, FLHS^{+}13, GL10, GW14c, GÖ14a, GD10, GLLM12, HHT12, HdJL^{+}13, HD13, HLWM10, HW10, HL11, HZR14, ILEE13, Ism14, JMS10a, JLZZ10, KMD12, KSK11, KP10a, KD10, KMLT10, KMP14, LGR11, LQYL11, LRZ14, LSW11, LL11c, LZ13e, LB14, LJE^{+}10, MA13, MOG10, MKK14, Mee12, MP10b, MR10, MVMG12, Naz11, NO12, NPTH09, PY10a, PM11a, PHA14, PPG14, PSKA14, RMP^{+}13, SMD14, SH13b, SI11, SWT12, SRA10, SOL11, SSF^{+}11, Sun14, SS14e, Sza10, TYCL14, TSC10, TT14b, VTLF11, WLL^{+}11, Wan12b, hXZ11, XSV12, XSV13, XZ14, XYW14, YZ10a, YYL^{+}10]. model [YLS12, Yua12, YZ10b, ZZSW11, ZYL10, ZY10, ZY11a, ZZS11, Zha14c, ZD10, ZZ13, Zhu14, ZwGhX14]. Modeling [AVMVMV10, Sez10, APO14, BSPD14, BÜ14, CCG12, CHMD14, CLP10, EX14, GR12, GUO14, HDF^{+}10, HHH^{+}13, HT14, KPsQ11, LA10, LCCK13, MR10, PTP10, RYM^{+}14, VBP^{+}10, WDG^{+}12, WL11a, ZM13, Zim13]. Modelling [Nai11, FLEa10, ILEE13, LP10a, PDE^{+}13, vGALL10]. Models [WW10a, AMPP11, ÁD11a, AP10, ADLS14, AELH10, BCM10, BCHL11, BHL11, Bir11, BKTY14, BB12b, BGP14, CC11b, CDQ14, CLMR14, CH11, DJM10, DDG^{+}13b, DAH11, DF12, DFH14, FPG12, Fer14, Fie13, FPP13, GCNL14, GMRK14, GS14, HS11b, JK13, KAS13, LL13a, LEVG12, LEVG14, LNJ^{+}10, LS10, LYP13, MOA10, MHH^{+}14, NRS10, PDE^{+}13, PGGCGF12, PPG14, Sch14d, Ven10, WF10, WA14, WS13b, WSS14, ZH10b, ZCX14]. Modern [ABÇ^{+}11]. modes [BKMG11, Lak13, fLyHZ10, SKAM11, Tau13]. modification [And10, KLK13, Nau11]. Modifications [LW11, CHMT10b, HH13, HH14]. Modified [BAU13, GCH13, Liu11a, RPS^{+}11, Vel11, WHL13, ABHMW11, AC14a, BKGMA10, BK14a, BA10, BLKZ10, Cai10b, CYJ13, CLW14, DW11, DD11, DCJ13b, FLHL14, FS10, GW14b, GG14, HXW11, KF11, LP13, Mah14, MML11, Wan11f, WSWD11, WMGM14, XWW10, XSW12, XZH14, pYXPH10, YWL14, Zho13, ZC13]. modular [RP11a]. module [GZ11]. modulus [ZY14b]. modulus-based [ZY14b]. moisture [LMW^{+}11, SDK10]. molecules [Bra14b]. Möller [WZD10]. mollification [YF14]. MOLP [LJE^{+}10, LJMG11]. Moment [DDV10, BHJN10, CR14, FKL10, GA12, GS14, LM13b, XCY12]. moment-closure [CR14]. moments [ACH11, Han11, Tor14, ZH11]. momentum [KH11, RG12]. Monge [Nei14]. monitoring [BL10b, SH13a]. Monolithic [SCDA10]. Monotone [EZ14, Wan12a, AAK10, Bog11, Bog14a, Bog14b, CSY10, EGON10, LZ11, pLjH10, SZ11, Wan11f, YPL10, Zha10a, ZYG10]. monotonic [WF10, WS13b]. Monotonicity [MP10a, LWZ13, QGG10, Qi14, WZ10a]. Monte [GPVL10, KZ13, LM13a]. Moore [PS11b, PS14b]. Morley [BySZZ13, GH10b, LGM14]. Morozov [Wan12b]. Morton [VTLF11]. motion [ÁE10, ÁE11b, BM10, BB11a, DA14, JL11, PW12, VAZS13, YÖVAYOW14]. motions [JJK^{+}13, KPV14, SGO14]. Mountain [GPO14, TT12]. movement [Dou12].

Moving [PP10a, CS10, DQST14, FiMO10, FRB13, GTS14, HS12, MLZ14, RKS^{+}13, RZD14, SD13, ZS11]. MPCC [MR11]. MPRP [LT11]. MR3144710 [GZ14b]. MRI [RPS^{+}11]. MSSOR [ZHCW12]. Mulders [Har11]. Multi [GKB10, GD10, HVV^{+}11a, Wan12b, AG10a, CP13a, Cha11a, CCH10, CSZ11a, FO10, Gen12, GÖ14a, GWZ11, HL14, KS14b, KKK14, KAS13, KS12b, LZT13, LVD^{+}14, LFWL14, MAK^{+}14, PLN13, PT11b, PM11b, SBFC10, SY11, SB10, WCH11, WL10d, WK10, ZBZ10]. multi-accurate [WCH11]. multi-asset [HL14]. multi-component [ZBZ10]. multi-contact [SBFC10]. multi-criteria [MAK^{+}14]. multi-degree [WL10d]. multi-dimensional [Cha11a]. multi-element [LVD^{+}14]. multi-factor [GÖ14a]. multi-firm [LZT13]. Multi-layer [GD10, CP13a]. multi-level [LFWL14, PM11b]. multi-objective [KKK14, SY11]. Multi-parameter [Wan12b]. multi-particle [KS12b]. multi-point [Gen12]. multi-projection [PLN13]. Multi-scale [GKB10]. multi-server [WK10]. multi-species [SB10]. Multi-stage [HVV^{+}11a]. multi-start [AG10a]. multi-step [GWZ11]. multi-strain [KAS13]. multi-symplectic [CSZ11a]. multi-term [PT11b]. multi-tree [KS14b]. multi-wavelets [CCH10]. multi-window [FO10]. multibody [DKH^{+}10, ILEE13, WA14]. multicast [NACLR12]. multicore [MM11a]. Multidimensional [LE10, AB13, EC13, HZ14, Tan14a, XW12]. multifactor [HT14]. multiframe [FN12, JMV13]. Multigrid [Cui14a, CC11b, DZW14, EG11, GGLR10, KMHM14, KOV11, LZ10, RSGL14, WZ10c, YZ11a]. multilevel [CM12, CLS14, GG12, RS10b, Wil13]. multimesh [SCDA10]. Multimodal [Pag10]. multiobjective [Pov14]. multiphase [SGO14, YMB14].

Multiphysics [YK11]. Multiple

[DZW14, LZSL11, AHH13, Arv10b, BBS14, BK10, BA10, Bra12, CG10a, CYJS14, CL11c, CDS10, CMR11b, CSZ11b, DSL12, DL13, DSY^{+}11, DJ10, HHHS14, HD11, KÖ11, MZL14, TA13, VVM10, WYO11, WWG10, WLH12, Win14, Yun11, ZTS11, ZDZ11, ZG14, ZC10b, ZCS11]. Multiple-cue [LZSL11]. multiplication [OORO12]. multiplications [CÖ11a]. multiplicative [CC14a, GLHCLPN10, LZ13c]. multiplicity [CGS13]. multiplier [LK13]. multipliers [ZBC13]. multiply [LDGY10]. multipoint [DP12, DP14c, GK11, SSW14]. multiprecision [BR14b]. Multiproduct [TNN10]. multiquadratic [KG13]. multiquadric [AM14a, FZ11, GW14d]. multirate [KL14]. Multiresolution [ZM13, ADL^{+}10b, ADL^{+}10a, AM10c, ADLT13, AMR13, CGS12, CCS10, Jia10, Jia11]. Multiscale [VSB12, Bay14, BKN^{+}13, BGP14, DZW14, GW14a, GK14, KA14, Wil13]. multiserver [GI10]. multishift [AMMS12]. multisplitting [DH11, LJCL11]. multistage [LTZ10]. multistep [CH14, MF14, MMS10, WWL10]. multisymplectic [Cai10b]. multitherapy [ZLW14b]. multitude [SSW14]. multivariable [ASS12]. Multivariate [BBD13, GR13, LS10, RMR11, WLZ11, WGZL14, AÖ11, CLLX11, CCH10, CC13a, DLSV14, Isl14, Kho10, LCXL11, ML14, ÖW14, Spe11, TT11, XZL11, YSQ11, ZW11a]. multiwavelets [AM10c, CCCM13, LSD11]. Müntz [AHSN13, Ste10a]. mutual [SPLHCB14]. MWV [YTFP11].

N [YTFP11]. name [GUO14]. nano [GKL12b]. nanofluids [GTS14]. nanometric [LLG14]. nanoparticles [LVD^{+}14]. nanosecond [APO14]. nanotori [ASG11]. nanotube [SRA10]. nanotubes [ASG11, CH11]. Nash [Li14]. Natural [CFGE10, BB11b]. nature [WW10a]. Navier [Dai11, CAT13, CDS10, DSC09, DD11, HL10a, HdJL^{+}13, KH11, ZH10b, ZY14a, ZLC11, dFGAN11]. NCAR [BBD^{+}11]. NCLab [Bra14a]. NCP [YP10a]. Near [GD13, ÁD10a, AMR13, BISS10, Van10]. near-best [BISS10, Van10]. near-lossless [AMR13]. Near-onset [GD13]. nearest [Rum13a]. nearly [AV10a, BSB14, JJE13, Lam14, QZX^{+}11, WZ14]. Nédélec [BL10a]. needed [FSB10]. negative [EY14, FGEM10, KC10]. negativity [Sza10]. neighborhood [FF14]. Nesterov [CSSX14b, CSSX14a]. nesting [MOG10]. net [XMC11]. nets [CD11]. network [CFdAR11, CRS12, DL13, Iid12, LCW10, MOA10, NO12, RLGGGÁM14]. networks [AAY10, BV10, CLS12, LF10, LF11, LSXZ11, LCN12, PR12a, RSFS14, Rom11, XCZ^{+}11]. Neumann [CDKN15, CDKN14, RE10, SSP12, qW11, ZD10]. neural [AAY10, BV10, DL13, LF10, LF11, LSXZ11, LCW10, MOA10, NO12]. neuroscience [BSZ13]. Neurosurgery [VBP^{+}10]. Neutral [WLZ14, DF14, GMZ^{+}10, Hu14, LB09, LW10b, LB10, LYZ11, RW10, RBK11, RHS11, SLD13, WWL10]. neutral-particle [RW10]. Nevanlinna [BL10c, BC12, LB11, WW11a]. Neville [ADGP11, ADGP12]. Newton [AR10a, AWF11, AH10b, AR11, AH12a, AH13, AHK14b, AK14b, BHL11, CAA10, CC13b, CLW14, CT10, DA10, ENS14, EGH12, FYA^{+}14, Fer11, FGO12, Gat14, HPL^{+}14, HH13, HH14, HHZ11, KO13, LZY10, LWY10, LZ14b, LH11b, MPC10, PR11, Pov14, RP10, RA10, RX10, Śmi11, SZ11, WLD14, Xie11, YP10a, ZSX11, ZW12, ZC13]. Newton-like [AH13]. Newton-type [CAA10, RA10]. Newtonian [MT14a, JMS10a, ZY11b]. next [SP11]. nice [Gro13]. Nicholson [Alz10, CL11c, DN13]. Nicolson [CS10, JK10c]. Nikishin [PL10c]. NIST [MKS12]. No [PS10]. Node [DGG12, Dou12]. Node-pair [DGG12]. nodes [BDH11, CG10b, ESM13c, KSA14]. noise [BS11b, BB12c, CC14a, HPR12, IK13, LZ13c, Sla14]. noises [GLHCLPN10]. noisy [CCG^{+}13a, JMV13, RBD11, RBD14, SW10a]. Non [AM10c, BSKL12, KP13, LYP13, MRT13a, Uty10a, AMÇ13, ÁNA14, AD11b, AAD13, BBP14, BS10a, BT12, BCR11, BCR13, CBB13, CV10a, ĆRR^{+}10, CR11b, CO11b, CCZ12, Dea11, DS12a, DJ10, DS10b, DS12b, DS13c, DFH14, FL10a, GU10, GJCL11, HPL^{+}14, HP10, JMS10a, KP14a, KB11, LVD^{+}14, LM14, LG11, MHL11, MP10a, MS13a, MOG10, PW13, SSR10, SKKS12, Sza10, WH10, WH13, pYXPH10, Yar10, ZLL10a, ZW12]. non- [BBP14]. non-autonomous [KP14a]. non-banded [CV10a]. non-commutative [KB11]. non-conforming [FL10a, HP10, LVD^{+}14]. non-continuous [LYP13]. non-globally [MS13a]. non-Hermitian [DJ10, PW13]. non-instantaneous [GU10]. non-interior [LG11, ZW12]. non-interior-point [ZLL10a]. Non-isothermal [BSKL12]. non-linear [AMÇ13, BS10a, DS12a, DS12b, DS13c, MHL11, WH10, WH13]. non-Lipschitzian [pYXPH10]. non-local [DS10b]. non-negativity [Sza10]. non-Newtonian [JMS10a]. non-oscillatory [CBB13, CCZ12]. non-polynomial [MP10a]. non-premixed [GJCL11]. non-rectangular [AAD13]. non-reflecting [Dea11]. non-self-mappings [ĆRR^{+}10]. non-selfadjoint [CO11b, Yar10]. non-singularity [HPL^{+}14]. Non-smooth [LYP13, AD11b, BT12, SSR10]. non-standard [ÁNA14]. Non-stationary [Uty10a, CR11b, SKKS12]. non-stiff [LM14]. Non-symbolic [MRT13a]. Non-uniform [AM10c, KP13, BCR11, BCR13, DFH14, MOG10]. nonclassical [MMF10]. noncompact [JM13]. nonconforming [CM13, CRM10, FKNS11, HJS14, KL10a, LLX14, PSS13, YC10, ZC10c, ZLC11]. nonconvex [CS12, NASNY11]. nondegenerate [FKL10]. nondiagonal [DM10b]. nondifferentiable [GD14]. nonequilibrium [FYA^{+}14]. nonexpansive [CSY10, HLC11, HL10b, QCK10, She10, WG10a, WZA10, WSWD11, XSH10]. nonhierarchically [GLWF13]. nonhomogeneous [Kur10, SRWZ14, TT14b]. nonidentically [ZH11]. nonisothermal [Sun14]. Nonlinear [AH10a, CyYwY11, GRV11, ILEE13, KCBA13, KNY10, SM14a, Uty10b, VS13b, ZAWA13, AEWMES11, eMA14, ABS11a, AA14a, AM14a, ÁD11a, ADL^{+}10b, ADL^{+}10a, ADLT13, AH13, APJ10, AAD13, AK11, AuI13, AuIK14, BK14a, BDD^{+}10, Bab10, BCHL11, Bar13, BK10, BSB14, Beh14, BS11b, BGL13, Bog14a, Bog14b, BS13c, BGP11, BK12, CCG13b, CZ12a, Cha11a, Cha11b, CL10b, CLB11, CXH09, CN11, CHMT10a, CHMT12, CTV13, CHMT13, CyY10, CwYyY11, CFL^{+}14, DJM10, DS13a, DR14a, DAW11, Dha14, DMR13, DCJ13b, Dra10, Dra13, DGW10, EK11, EZ14, EDP14, EHR11b, FDH10, FD13, FKL13, GS10, GZ10, GK14, GXW14, GZL14, GC11, Gen12, GSGN11b, GSGN11a, GSNDB14, GKL12b, Gus14, HSY13, HC11, KS14c, KC11, KRS11, KMP14, LL11a, LL14a, LL11b, LSW11, LY10, LB09, LB10, LW11, LZ14b, LGZ10, Ma10]. nonlinear [MY11, MMS13, MA14a, MKK14, N’G10, Nei14, NLO13, Odi11, PT14, PFLZ11, PAJ10, Prz13, PZ13, RBK11, RGJ10, RD10, RE10, RX10, Sha14, SZJ12, uIAAF14, SD10, SC10d, Ste10b, SR13, TM11, Tan10, TP10, Vel10, Wan10, WL10c, WL10b, WWL10, WZF11, Wan11d, WLWH11, Wan12a, WLZ14, WWG10, qWhY11, XJ11, YPL10, YW10, Yun11, Yun12, ZCH10, ZJ12, ZCS11, ZZ11b, SKAM11]. nonlinearity [EK11, KRS11, Sie11, ZC10b]. nonlocal [HMVPAR11, NS10, SD10, Yan11]. nonmagnetic [MKD12]. Nonmonotone [FB12, CWQ11, JLZZ10, SS14a, TJL11, ZH10c, Zho13, ZCX14]. nonnegative [CL11d, GG12, LZI10, LLW14, Mir14, ZBC13]. nonnormal [Isl14]. nonoverlapping [SLP11]. nonparallel [QTS14]. nonself [WZA10]. nonself-mappings [WZA10]. nonseparable [ADL^{+}10a]. nonsmooth [LTW10, LPGX10, Śmi11, TBK14, YWL14]. nonstandard [ADLS14, AS10b, GPZ10, JTV14, MO10]. nonstationary [BDJ^{+}11]. nonsymmetric [AS10a, BMS11, CJ10, DSY^{+}11, FJP14, GG10, GG14, LHL10, YLD13]. nontrivial [BB11c]. nonuniform [BS13b]. Noor [PS11c]. norm [BC12, BB14a, BGKVW10, CS10, Dax10, Gün10, LZH14, lLzS11, LHC14, Wan11d]. normal [FM11a, FV14, KMS14, MD13, Pir13, Pir14, ZM14]. normalized [Beh14]. normed [MS11c]. norms [UM14]. north [KG14]. north-south [KG14]. Note [WF10, Zha10b, AAT11, ADGP12, BIT11, Cao13, CM10d, Ehr11a, ENSCS12, GUBS13, GLLM12, GWZ11, KN10b, Kud10, Lu13b, UBSG12, WWW14, Yar10, ZYG10]. Notes [XWW10, ÇU14]. notice [DS13c, Wan11c, WYO11, XSV13]. notion [Ram10]. Nov [Ano10b]. Novel [BHBP10, BK13, BAAS11, BKTY14, VMMM10, qWhY11, XJ11, ZZSW11]. NP [RVPOR^{+}11]. NP-complete [RVPOR^{+}11]. NRBCs [LEKT10]. NS [YZ10b, YZ10b]. nuclear [LZH14]. Number [ZD13, AELH10, BAAS11, Che14a, DSL12, DMMP10, FSB10, Gün10, Har11, MDVB13, SXZ11, TA13, Wik10, ZZ13]. numbers [CLN14, GG13a, GQ14a, GQ14b, LEKT10, XhW11, XC14]. Numeric [BFS14, BFMS11]. Numerical [ABHMW11, AALMLM14, AE11a, AM14a, AJK12, AV12, BPS12, BSZ13, BPGP11, BFG^{+}14, BSJ14, CDTV11, CPY10, CLB11, CLP10, CDJP13, CF11, CFL^{+}14, DS13a, DDM10, DA10, DS11, EASS13, ELA10, FM11b, FYM10, FRXY13, Fuk13b, Fuk13c, GVF^{+}13, GI10, Gra11, GKL12b, HDF^{+}10, HdJL^{+}13, HTH10, HS10c, HL11, HLLW13, IK13, IK11, JX10, JWY11, JV10, KA14, LSD11, LYD13, LDGY10, MT10a, MM12a, MKK14, MT10c, Miy12, ML10, MS13c, NH13, NS13, NLO13, NdCDCFM10, PW12, PT14, PKH14, RGJ10, RTL14, RBD11, RBD14, SH13b, SBL12, SZ10, Sto11b, SR13, SKN11, TR11, VB14, WW10b, Xu11b, XL12, YZ11a, YLTB14, ZBW11, ZZC11, APO14, AHLJ14, ACF^{+}14, ÁD10a, ÁD11a, ÁD12, AS12a, AP10, AV10b, AAD14, AC14c, AuI13, AuIK14, BMMR11, Bab10, BBMY11, BCC12, BL13a]. numerical [BKV10, BM12, CK13, CCJC13, CDV10, CKM14, CC11c, CGS11, Col13, De 14, DS12a, DAE13, Dek10, DL14a, DME^{+}12, DL12, DS10b, DI12, Eba11, FP13a, FFM^{+}11, FLLT10, GXW14, Gau10, Gen12, GQL14, HQR13, HGW12, Hua12, JJE13, JR10a, JTV14, KS12a, KKK14, KS10b, KA12, KP10a, Kre95, Kre11, Kro10, KLM14, LCCK13, LZ13a, LVCC13, LTFL10, LM13b, LY13b, LJ13, LO14b, LXSC11, MMF10, MA13, MS13a, MN10, MV14, MVMG12, MFL13, Mui10, MT14b, NS14, Odi11, OAHR11, PTP10, Pas11, PGH13, QZX^{+}11, RCM^{+}12, RMR11, Rie13, SSP12, SASH11, Sid14, uIAAF14, SS14d, SLK14, Tan14b, TW14, TS12, VAR11, VT14, WL11a, WZMC11, WC11, WZL11, Xu11a, YMS10, mZkPkL11, ZS11, ZCC^{+}13, ZLC10]. numerics [FdOdS12]. Numerov [GCH13]. NURBS [YWY^{+}14]. NVIDIA [GWKE14]. Nyström [AST14, Bar13, FG14, OPPV10, YFZ14, DP14b, SCRG12].

O [HL13a]. object [SPC14]. object-oriented [SPC14]. objective [KSK11, KKK14, Mol10, SY11]. objectives [CG10a]. objects [PS10, RKS^{+}13]. obligation [ÇU14]. oblique [Gas14a]. Obrechkoff [SASH11]. observations [GLHCLP11, HTLI14, JWW14]. obstacle [BySZZ13, BDyS14, RE10]. obstacles [HW12]. obstructive [ÇDE14]. obtain [CHMT13]. Obtaining [MPS10, Ain10, RRMD13, ZCS11]. Occasion [ARV10a, MYZ10]. occupancy [BDL11]. occurrences [TT14b]. odd [DDM11, YKL10]. odd-order [YKL10]. ODE [BGPPRW14, KK14, NBBKV10, VAR11, WS13b]. ODEs [DL12, Dra10, HwKB14, IHRA11, LT13, MMS10, Sch14a, Sch14b]. off [CH14, GGS12]. offset [KL11]. Olkin [AKA14, OB14]. Oman [WDG^{+}12]. once [YK14]. One [PC12b, BM12, CC13b, FKL13, GW10, GST12, HS12, JLZ14, KSK11, Kır11, KS12c, LW10b, MOG10, MML11, Moh13, Mol10, NBBKV10, SC10a, WWW14, YW10, ZTC11]. One-dimensional [PC12b, FKL13, GW10, HS12, KSK11, Moh13, SC10a, ZTC11]. one-leg [LW10b, YW10]. one-point-based [KS12c]. one-step [NBBKV10, WWW14]. one-way [MOG10]. onset [GD13]. onto [CW13, CGK10, KD10]. Ooura [Tan14b]. open [LA10, MCS12, NGAZ11]. open-source [LA10]. OpenMP [MM11a]. operation [CKM14, ZLG12]. operational [BMMR11, LYD13, YBD11]. operations [AÖÖ14]. Operator [BKL14, AMT11, CCG13b, CDKN14, CDKN15, DSC09, Dai11, FZ11, HCQ14, HC11, JK13, Kur10, Li11a, MSE10, ML10, PN11, PLN13, PD13, PAD10, Sha14, TA13, TT11, WZF11, WZ10d, Zha14b]. operator-splitting [DSC09, Dai11]. operators [AB11b, ADLT13, BB12b, BDR11, BR14c, BTSA14, CDD12, CID10, De 14, GSNA13, HZ13, HS10c, Hu14, KKN11, pLjH10, Mah11, MJ10b, Min10, PY10b, RVN13, SYİ11, SAB14, SC10b, jTW14, Yan11, Yar10]. Opial [MR12]. OPSFA [Van11a]. OPSFA-10 [Van11a]. optical [AH10a, BCM10, CDPP10, CPP11]. Optimal [BGIN10, BK14b, CC11a, CZL14, CDPP10, CPP11, ESM13c, Gal13, HW10, KP14a, LRZ14, Prz13, SH11, TW14, WZ11b, ZSG^{+}14, ZL10, AA14a, AØ14a, AK14a, AVGCMVM10, AV12, BBP14, BGKVW10, CK13, CY11, CWY11, CL11e, CFF12, CL14b, CHMT10b, CTV11, CHMT13, EG11, Fer14, Fu10, GCNL14, GY11, HJS14, Hua14c, KN10a, fLyHZ10, Liu14b, LYD13, MMP11, MNQ14, NF14, SYK14, SHN13, Sla14, SSR10, TYCL14, TP10, VDG11, XCY10, XL11, YK11, YK14, YL11c, ZLW14b, Zhu14, ZwGhX14]. Optimality [GAS14b, FPP14, LLW14]. optimally [GK11, Prz10]. optimisation [LLG14]. Optimization [GLWF13, HVV10, KPR11, KT13, YWW10, ZDP^{+}14, AEWMES11, AZ11, AEGB10, AG10a, ADRM12, ABÇ^{+}11, And10, And13, Ant11, BDV10, BEJ14, CCS11, Che14b, Chu10, CWQ11, DW11, DdS14, EGBS12, ESM13b, GLH12, GKU14, GGH^{+}14, GS12, HTM14, HK13b, Iid12, JLZZ10, JW10b, KKK14, KK13, Kuz14, KS12c, Lay13, LV14, LP13, LMM^{+}10b, Lu13a, LJ13, LZWJ11, MAK^{+}14, Mol10, NY12, PNCC10, PHS14, Pov14, RMF11, TOF14, TBK14, VMAVGCM13, WB10, WUB10, WK10, Wu10, XSW12, YSQ11, ZSG^{+}14, ZZC10, ZSK^{+}11, ZC13, ZZG10, ZWG11, ZGZB13]. optimization-based [Kuz14]. Optimized [HHKK14, JR10a, CFH14, FDD10, HVV^{+}11a, Liu14a, Lui10, LZSL11]. optimizing [MY14]. Optimum [MRAM10, YTFP11]. Option [BHPS12, CLMR14, HZR14, YY10, Bal14, DDV10, DF12, DFH14, Gon10, JK13, KPsQ11, KZ13, LWN13, LZ13e, NR14b, PO11, Tan14a]. optional [WYP10]. options [CMR11b, DPV10, ESC14a, FPP13, FS10, HS10b, HL14, HS11b, Jun13, LL13a, LLK14, fLxX12, LMG13, Mar11, MT10c, MY13, PZJ12, TTB14, ZYL11]. orbits [ÁD12]. Order [BP11b, AGM12, AHLJ14, ABG11, ADLT13, ABCP14, AS12a, AR10b, Ant11, AD11b, AM14b, AM11, Ars11, AKT14b, Ayd14, BDD^{+}10, BGIN13, BS13a, BK13, BB11b, BGL13, BL10a, BG14a, BHL11, BK12, BOS12, Büy12, CCG13b, Cas10, CXL10, CPP13, CI13, Cha10a, CRM10, CL10b, CLB11, CXZ14, CLS10, CHH10, CGL10, CG13, CDQ14, CHMT10a, CTV11, CTV13, Cui10, Cui14b, CFL^{+}14, DSC09, Dai11, DLG14, DDG13a, Dea11, DO12, DY11, DSL12, DH14, Dos14, DF12, DFH14, DGW10, EO14, EN14, EZ14, Ery13b, EHR11b, EJ10, FL10a, FKL13, Flo11, GZL14, GC11, GK11, GPHAM10, GSNA13, GM13, yGZ13, GD14, HLM10, HMA12, HS10a, HK14, HH13, HH14, HZ13, HVV11b, HSY13, HHZ11, ILEE13, JM11, JR10a, Koç11, KB11, KB12, KB13, KP10b, LLS14, LP12, LS11, Li12, LSW11, lLzS11]. order [Lia14, LLSW11, LZ12a, LB09, LB10, LW11, LL13b, LGZ10, LO14a, LO14b, LO14c, Lui10, LC10, MY11, MHL11, MOG10, MGZ10, MMS10, MTJ11, Mui10, Naf14, NS10, Nei14, NBBKV10, NTNBV14, NRP13, NBZ13, NCCF^{+}10, ÖY14, PP10b, PFLZ11, PRM11, RCM^{+}12, RW10, RS10a, RK11, RB13, Ren10, Rua13, SMD14, SM14a, SH10, SASH11, Sie11, SB10, SZS11, SL14b, SBGB10, Tan14a, Tan10, TS10, Tar10, TP10, TY11, TKW12, VVD^{+}14, WZS10, WZ10c, WZ10b, WL10b, WX12, WY13, WLZ14, qWhY11, WZ10d, XMSS11, YE14, YKL10, YG13, Yun11, ZAT11, ZFG10, ZL10, Zha11a, ZH11, ZZS11, Zha11b, ZF14, ZWZ11, ZCS11, Zho13, ZL13, ZX10, dJMPPC14]. ordered [DMGVPN10]. orders [BDR11, BR14c, CHMT10b, uIAAF14, WW10b, WZY^{+}14]. ordinary [BJ14, KKN11, SP11, WT14]. ordinates [BCM10]. orientational [BKL14]. oriented [CG14, GH14b, KM10, LJE^{+}10, SPC14, ZGZW10]. ornamental [LMM^{+}10a]. Ornstein [Sch14d, WZ11b]. Orr [WTM12]. Orthogonal [BVV10, GHM10, LZDW14, Van11a, AMPR10, ADMF10, Arv10b, BR11a, BHJN10, BCGMB10, BLTW10, BGVHN13, CL10a, CG11b, CH13, Dom10, DM10b, FPPX10, FPP12, Han11, HV14, LW10a, MGZ10, MJ10a, MBS10, Mir14, RSFS14, WZF11, Zim13, dJP13, dJMPPC14, vDvFZ09, vDvFZ13]. Orthogonality [CBGVPP10, DMGVPN10, BLO11, BDF^{+}10]. orthogonalization [BDD14]. orthonormal [LYD13, SWY14]. orthotropic [HKL11]. oscillating [RSFS14]. Oscillation [Zha11b, LB09, LB10]. oscillations [BKMG11, Gre10, MT14a, WZL11]. oscillator [CYJS14, WW11b]. oscillators [Che13a, FG14, WC10]. Oscillatory [GZL14, LGZ10, AH14a, AS12b, CBB13, CA14, CCZ12, Dom14, FFM^{+}11, FYM10, KXH13, KX13, LWXW12, LP10b, SGG11, WX10, Wu14, XXH14]. Osher [Wan11a]. OSMOSIS [LA10]. Ostrovsky [YMS10]. Ostrowski [Koç11]. other [BM12, SL14b]. outer [MK10, SS14d]. output [LLL12]. ovalling [HDL^{+}13]. over-penalized [BOS12]. over-posed [MM10]. Overlapping [GPZ10, KR14, Li12, NMBG14]. overrelaxation [BY12]. overtaking [BJ13]. overview [AA14b].

Package [MCS12, GGSP10, HSH12]. packed [Mee12, SH12, WSJ14]. Packing [FTTT13, KG12]. Padé [PGGCGF14, AF10, ACH11, IA13, PGGCGF12, Pré10, PL10c, GKLZ12, Zie14]. PageRank [WW10c]. PageRanking [CSC10]. pair [Che13b, DGG12]. pairs [CFJ^{+}11, PR10, PD13, SCRG12, dJP13, dJMPPC14]. Palindromic [DDM11, BG14b]. paper [HV13, PC12b]. PaperRank [AB14]. papers [AVA13, BDR10]. parabolic [AS10b, APJ10, Bog11, Bog14b, BSS13, BJ12, BSJ14, BMN10, Cha11a, Cha11b, CG13, CyY10, CwYyY11, CyYwY11, DY11, FID11, FDH10, FP13a, FN10, GW10, GW14b, GAS14b, LPDN13, LJ12, LO14a, LO14c, LXSC11, Min10, PAJ10, RBD14, SD10, SM14b, VS13a, Wan11e, Yan10a, YBD11, YL11c, ZFM14, ZD11]. parabolic-hyperbolic [CwYyY11, CyYwY11]. paraboloid [CP13b, PC11]. paradigm [WZD10]. Parallel [AMV13, HK12, Kva11, MS11a, MM11a, SM10, WZ14, Yan10a, YESD14, BDV10, BM11a, BHBP10, DHSW14, FGO11, GG13a, HHX14, KSLD11, KW10, KW11, LVD^{+}14, Man11, NGAZ11, NMS10, YC11, ZDP^{+}14, ZSG^{+}14, ZZG14]. Parallelization [Gra11, FPGMG11]. Parameter [WS13b, BG13, BFLP12, BKV10, CC13b, CC14a, FRXY13, GN14, GSK11, HPR12, KT13, Pas11, PPG14, RSFS14, Sch14a, Sch14b, Wan12b, qWhY11, WK10, YÖVAYOW14, ZFM14]. parameter-dependent [KT13, PPG14, Sch14a, Sch14b]. parameterization [BBD^{+}11, CH13, MMF10]. parameterizations [BL13a, YSZ11]. Parameterized [LZL11a, BSS14, CJY10, WCW10, ZYW14]. parameters [BA10, BDMP13, CZL14, CM10b, GS14, HS10a, Hua14c, KPR11, LQYL11, Liu11b, MJ10a, SOL11, TJL11, Tor14, WL11a, Wan12b, WYY^{+}12, ZRK^{+}12]. Parametric [SMD14, AS12b, CS12, LWW11, LL11b]. parametrization [RSS14]. parametrized [Alc14]. parareal [BL14, GH14a, LJ12]. paraxial [AC14d]. Pareto [AVGCMVM10, Fer14, KKK14, Nad11].

Pareto-optimal [AVGCMVM10]. Part [CHH10, NH13, Rum13b, WDZ10, AKT12, AKT14b, Rum13a]. Partial [CH13, WC10, AR10a, AM11, CY11, DXL10, FKL13, GAS14b, HD14, JM11, KLNS11, LD10, LLS14, LJ12, Ste10b, TM11, TNN10, VS14, Xie12, Yan11, ZBC13, ZRW11]. partially [DJM10, Ism14]. particle [AEWMES11, AZ11, GKT10, GKT11, KS12b, RW10, WZMC11, Wu10]. particle-laden [GKT10, GKT11]. particulate [GMRK14]. partition [BDyS14, HTH10]. partitioned [DBHV10]. partitioning [KS12c, LSXZ11]. partitions [JKK^{+}10, NRS11, Vit10]. partly [DS11]. parts [HZ13]. Pascal [ADGP13]. pass [TT12]. passage [BSZ13]. passing [Rah14]. passive [ZTL13]. past [Gas14a]. patch [Cao14]. patches [CZ14, FGPP13]. path [AHH13, HL14, KV12]. pathogenesis [HLWM10]. Pathwise [KLNS11]. patient [BDT^{+}13]. patient-specific [BDT^{+}13]. pattern [MVC10]. patterns [GD13, KSK11, SBL12]. payments [GU10]. PC [CT11]. PDE [ALM11, KP14b, KPBN13, SACS10]. PDE-independent [SACS10]. PDEs [AD11b, AC14c, CSZ11a, GPHAPR14, HSY13, Nei14, RCM^{+}12, SSP12]. PDIEs [RE10]. Peaks [OGM14]. Pearson [LP12]. Peer [Sch14a, Sch14b, KW10, KW11, SLW14, WK14]. PEM [Sun14]. penalized [BOS12]. penalty [hBW10, BOS12, DH13, GW14c, GG13b, HL10a, JZ10, KA14, LLT^{+}14, MWWY14, SHZ10, TJL11, hXZ11, XMSS11, ZYL11, ZwGhX14]. [MSP12]. antisymmetric [YSZ11]. circuit [Jan14]. control [ZTC14]. Darcy [CM12]. explicit [HL10a]. finite [DGG12]. front [BW14]. level [DGC^{+}13]. linear [IO14a]. max [ZWM11]. quotient [VVV10]. repair [MP10b]. Robin [Sun14]. skew-Hamiltonian [QT13]. Penrose [PS11b, PS14b]. pentadiagonal [Cin12, McN10]. perfectly [Li11b]. perforated [RL10]. Performance [LA10, ZH10a, ADA13, BFB14, BL10b, LWH^{+}11, LT13, NdCDCFM10, SCRG12, SW11]. period [YTFP11]. Periodic [BS10a, CN11, DF14, HZ11, ÁD11a, ÁD12, Alz10, CYJS14, CL11c, CLN14, CW14a, CFJ^{+}11, DN13, DBZ10, GW14d, Gre10, KC10, KCBA13, MN10, NPTH09, PWZ11, SS14c, TSC10, WZ10b, Xu11a, Yan13]. periodically [TA10]. periods [CHRZ13]. permanence [WF10]. Permanent [EX14, MKD12]. Permanent-based [EX14]. permissible [GU10]. permittivity [KS14c]. Permutations [ÇMT14]. Perron [Yan10b]. Perturbation [dL14b, AT10, AS12a, COW10, DCJ13b, GC11, HHP^{+}10, Ku13, LCN12, SBCD10, WCW10, XCZ^{+}11]. perturbation-reproducing [GC11]. perturbations [BN14, RBK11, XCZ^{+}11, ZCH10]. perturbed [Bog11, Bog14b, CXL10, CP13a, CGL10, CG13, FG14, GL10, GQL14, KR14, LL11c, RK11, RU13, jTW14, WA14, ZYL10, ZY11a, Zha14c, ZC10c]. Petrov [DS14, MA14a, MD10b, Ros11]. Petrovskii [KS10b]. Petviashvili [ÁD15, ÁD14]. Pexiderized [MS11c]. PGSCM [AGM12]. PH [KLK13]. Phase [EY14, AK14d, BW14, DQST14, EASS13, Han11, HdJL^{+}13, MV14, MS13c, PW12, PPG14, SCRG12, Sun14, SS14e, ZH11]. phase-lag [SCRG12]. Phase-type [EY14, ZH11]. phenomena [Sto11b]. phi [PM11a]. phi-divergence [PM11a]. phononic [FGEM10]. photonic [CvdMS12, Gia12, HHLW14]. physics [Ste10b]. Picard [FBB14]. Pick [BC12, LB11, WW11a]. Piecewise [AW10, PT12, Rie13, AAY10, AAY11, BFMS11, DY13, FSB10, IHRA11, LWW11, WZL11, YT12, ZW11a]. piecewise-linear [YT12]. piecewise-linearized [IHRA11]. piezoelectric [ZTL13]. pilot [AC14c]. pinch [GGS12]. pinch-off [GGS12]. pipe [ATS14]. pipeline [FLKW14]. pipelining [MM11a]. pipes [BG10b]. PIRK [CT11]. PIRK-type [CT11]. Piskunov [KS10b]. Pitaevskii [CCJC13, KHFC11, WC11]. Pitchfork [DL13]. pivoting [ADGP11]. Pizetti [LA11]. placing [KR10]. Planar [JŽ11, MSWY12, AV10b, BFHK12, CRM10, CGS13, LZ12b, RYM^{+}14, WM12a, YC10]. Planck [AB13, CR14]. plane [AHM14, Ehr11a, Gwi13, HWG10, WM12b]. planning [FPG12]. plant [FLEa10]. plasma [APO14, LvDtTB^{+}13]. plasmonic [Gia14]. plate [BOR14, Gas14a, KD10, LGM14, WHLC11, ZDP^{+}14]. plates [BySZZ13, BDyS14, GH10b, NdCDCFM10]. platform [RYM^{+}14]. platforms [HTM14]. PML [BUG^{+}13, BP13, CWY11, HL11, WC14]. pneumococcal [LGR11]. POD [AHLJ14, AHLJ14, LXSC11]. Poincaré [MR12]. point [BMS11, Bai13, BBB10, Bar11, BGL13, BM11a, BSBA14, CB13, CYJ13, Cao13, CDN14, CAA10, CWY11, CS14, ĆRR^{+}10, CRS12, DJM10, DFG^{+}10, Eba11, EGBS12, FdLW^{+}11, Flo11, Gen12, GQL14, GST12, GÖ14b, HS11a, HCM14, Iid12, JK10b, KC11, KPW10, KS12c, LHL10, LZL11a, LS11, LZ12a, LL11d, LC10, MMS13, Mau10, MSWY12, Més12, MTJ11, MRS10, OORO12, PD13, PL10b, PHS14, QYZZ14, RG11, She10, SHZ10, TP10, WZ10d, WL14a, ZLL10a, ZS13, ZLW14a, ZF14]. point-like [GST12]. points [AGT12, BL13a, BSV12, CW13, CM10a, CT11, HLC11, HL10b, KG14, KL10b, LLL^{+}11, LYZ11, Luo10, MRAM10, MBS10, ML13, RSV12, VAR11, qWhY11, XSH10, YL11c, ZW11b]. pointwise [Ku13]. Poisson [CR14, BCMR12, CM13, CP13a, CO13, DDM10, Fat11b, FWBSBGA11, GL10, GM14a, LL11c, LYZ11, WG10b, hXZ11, ZYL10, mZkPkL11]. Poisson-like [FWBSBGA11]. Pol [WW11b]. Polak [YWL14]. polar [AC14b]. polarization [BKL14]. polarized [KP11b]. polarized-radiation [KP11b]. polaron [Vel10]. pole [SMD14, Van10, WW11a]. poles [BGVHN13]. policies [GI10]. policy [CLS12, CL14b, GU10, HW10, WYP10]. Pollard [WZ12]. pollutant [DGI^{+}10, SH13a]. pollution [Ku13]. Polyak [YWL14]. polygamma [QGG10]. polygon [Dos14, SSP12]. polygonal [BFGP10, BFP11, Li11a, MNQ14, MWWY14]. polygons [GSV11, LMCW10, YWY^{+}14]. polyharmonic [BDR11, BR14c]. polyhedral [CH11, GWKE14, MWWY14, ÖW14, PS10, RW14]. polymer [PSKA14]. polymer-metal [PSKA14]. Polynomial [BCR11, CC13a, GSV11, ADMF10, BCR13, Bel13, BBD13, BL10a, BR14b, BSV10, BFMS11, CW11b, CR11b, CHZ10, CC12, DDM10, DMGVPN10, FZ11, Ger12, GSM11, GZ14a, GZ14b, HX13, HYY14, JGH^{+}13, LWW11, Lu11, Lu13b, MP10a, Odi11, PT12, PRM11, RVPOR^{+}11, SW10a, SW13, SXZ11, Str11, Str12, WLH12, Win14, XH13]. Polynomial-based [BCR11]. polynomialization [Boy11]. polynomially [Alc14]. polynomials [AAT11, AMPR10, ÁE10, ÁE11b, ÁNA14, Arv10b, Ave10, Beh14, BLO11, BR11a, BHJN10, BCGMB10, BLTW10, BG14b, BE11, Cas10, CK11, CG11b, CKM14, CL10c, DY13, DDM11, Dom10, DJJ13, DM10b, FPPX10, FPP12, GHM10, GQ14a, GHR10, He12, HC14, HV14, KÖ11, LW14, LW10a, yLGP11, MGZ10, MBS10, NLO13, OP12, ÖY14, Pré10, RP11a, SMDMY10, SMMD11, Sid14, Spe11, TS10, Van11a, WC11, WL11b, WL10d, YBD11, ZD13, dJP13, dJMPPC14, vDvFZ09, vDvFZ13]. population [AALMLM14, AMT11, ÁAG^{+}13, LSW11, TOF14, ZZS11]. poroelastic [CLP10, ZTC14]. poroelasticity [BY14]. porosity [KMLT10, MHL11, RMP^{+}13]. porous [AJK12, BGLV11, COW10, DQST14, EASS13, LE10, MHL11, MS13c, PTP10, PDE^{+}13, PSRV10, RP10, Rah14, SSW14, SRP^{+}13, Sun14, YMB14]. Portfolio [ZZC10, AMPP11, CW14c, fLxX12, Liu11c, PP14, RSK14, VDG11]. posed [BJRS12, CR12, CHL14, DR14c, HPR12, HR10, HR12, Liu14b, LFWL14, MM10, MRS10, Ram10, RS10b]. posedness [PW12]. Positive [CL11c, LS11, WZ10d, AGT13, BK10, BFS14, CJ10, DN13, DLWZ14, EZ10, EZ14, EGST11, GT09, KC10, KLL11a, LWZ^{+}14, LHL11, PS13, RRMD13, SYİ11, SH11, Wan13, WWG10, ZF14, ZC10b]. positivity [Kha10, MT13, SSK^{+}13, ZZS11]. positivity-preserving [SSK^{+}13, ZZS11]. possible [Vel10]. Post [MR13]. Post-processing [MR13]. posterior [LLX14]. posteriori [ANS11, AM14b, BG14a, CZWD14, CM13, CL11e, CG14, CW14b, CET14, CN10, ESV10, FM11a, Gia12, HJS14, KL10a, KKN11, KP10b, LY13a, MJ14, NV11, NÖ10, Seg12, Sta11, ZFM14, dFGAN11]. postponed [SJA12]. postural [BAdS^{+}10, BAdS^{+}12]. Potapov [BL10c]. potent [LRT13]. Potential [MS13b, MRD13]. potentials [Fat13, JWY14]. Potra [CHMT10b]. powder [KFR13]. Powell [BFGP10, DY13, FGPP13, LMT13, SCFR14]. power [AH14a, Ars14, Ma10, pWqSmZ14]. powered [CZWD14]. pp [Wan11c, WYO11]. Practical [Zho12, Hua14c]. Prager [PTP10]. pre [PM11a]. pre-test [PM11a]. precession [CM10b]. Precise [Fuk11, Fuk12, Fuk13d]. precondition [JX10]. Preconditioned [SCFR14, AIA10, Cao13, FYA^{+}14, HK12, HL13b, KOV11, KN10b, LHL10, PZ13, QY10]. preconditioner [CYJ13, FJP14, GW14a, GG10, PZJ12, PL10b, PW13, WL14a]. Preconditioners [BM11a, DDE14, AU14a, CDN14, DdS14, FdLW^{+}11, GPZ10, LZ13d, Mor10b, SHZ10, SSF^{+}11, SBGB10, WHL13, XN14, ZS13]. preconditioning [AM10b, CJY10, ENSCS12, FWBSBGA11, FPGMG11, Kur10, LZL11a, Spr14]. predator [AD10b, DF14, JJL11, LCZ11, LYP13, NPTH09, TSC10, WF10]. predict [BGS12]. prediction [ABÇ^{+}11]. predictions [KD10]. predictor [FF14, KKR11, LXY13]. Preface [Ano10c, APW^{+}14a, BFD^{+}10, BBMSC10, CCQ^{+}11, DG11, FL10b, GVV10, GGBA11, GVV12, NHVD10, SKSS10, Sol12, VV14, WWLY11, WWGM14]. premium [XYW14]. premiums [ZY10]. premixed [GJCL11, WCZ11]. preprocessing [WZD10]. prescribed [ÇMT14, CL11d]. presence [ÁAG^{+}13]. Present [Tez14]. Presentation [ZCC^{+}13]. Preservation [SLD13, ÁD12, HD13, Sza10]. preserve [BI12, Mah11]. preserves [Nau11, VAR11]. Preserving [VCA14, AP10, BBP14, BCMR12, CLS10, DR14a, GPS13, GSNA13, GG14, HHKK14, HLLW13, JWL13, wKK12, KHFC11, Kuz12, fLyHZ10, xLwCL10, MP10a, MT13, NTNBV14, RRMD13, SSK^{+}13, SGO14, WH10, ZZS11]. pressing [PC12b]. Pressure [LL11a, MDVB13, AJK12, CAT13, Rah14, RVDM10]. pressure-correction [RVDM10]. Pressure-velocity [MDVB13]. pressures [AJK12]. pressurised [BG10b]. prey [AD10b, DF14, JJL11, LCZ11, LYP13, NPTH09, Reb12, TSC10, WF10]. prey-predator [AD10b, LCZ11]. price [GUO14]. prices [DDV10, GD10, HT14]. Pricing [ÇU14, DPV10, EF14, Emm11, NÖ10, YYL^{+}10, Bal14, BHPS12, CMR11b, CLMR14, DF12, DFH14, ESC14a, FM14, Gon10, HS10b, HL14, HZR14, JK13, KZ13, LWN13, LLK14, LZ13e, MX10, MY13, NR14b, PZJ12, PO11, Tan14a, TTB14, YY10]. primal [Lam14, LL11d, SHZ10]. primal-based [SHZ10]. primal-dual [LL11d]. primary [HL13a]. Principal [CO11b, LWY10, WX10]. principle [AA14a, DW10, Wan12b, pYXPH10]. principles [Min10]. prior [FPG12]. priori [BFP11, KN10a]. probabilistic [LCN12, PM10, XCZ^{+}11]. probabilities [AAK10, ACH11, CFGE10, HD14, KH13]. probability [LCN12, Sto11a, XCZ^{+}11, YLS12]. Problem [MCS12, AA14a, AVGCMVM10, AP14a, AB14, AB10, AM14b, AV12, AC14d, BDL10, BFDP13, BOR^{+}12, BY14, BSZ13, BSJ14, BC12, BEJ14, BFP11, BySZZ13, BDyS14, CDTV11, CR12, CS10, CM13, Cha11a, Cha10b, Che10, CHX11, CL11d, Che14b, CDV10, CAT13, CSC10, CF11, CL11f, DHSW14, DCGS13, DDDSC14, DY11, EGES14, EK12, EC13, FM11b, FM11a, FKL10, GPKV10, GA12, GD13, GSM11, GCH14, HHH^{+}13, HDT11, HS10c, HLSX11, HNPX11, HY11, Jan14, JL11, JV10, KL10a, KLL11a, Kok11, KP11b, LZT13, LTX12, LV14, LCXL11, LQYL11, LL11a, LRZ14, Liu11c, LB11, LCW10, MHL11, MB11, Miy12, Moh13, Naf14, Nau11, NO12, PN11, PC12b, Pro12, QTS14, QXW10, QC14, QW10, Ram10, RE10, RBD11, RBD14, RU13, RR10, SSW14, SK14, SP11, TL13, Uty10b, Uty10a]. problem [VS13a, VDG11, VS13b, VT14, WZ10b, WZ10d, Xio10, XL11, XGL12, XW12, XZH14, YF14, YGHH14, YD11, YD12, ZLL10a, Zha11a, ZMZ13, ZW10, ZY14b]. Problems [DBZ10, AEWMES11, AGM12, ADG11, AMMS12, AK14a, ADRM12, AU14b, AS12a, AHH13, AS12b, AR10b, Ant11, AV10b, ANS11, APJ10, AKT14b, ARF14, BMS11, BBP14, BS11a, BBS14, Bar11, BK10, BLM11, BB11b, BM11a, BC10, BG14a, BGLV11, Bog11, Bog14a, Bog14b, BJ12, BHJN10, BJR^{+}12, BP13, BB10b, BI12, BB12c, BN14, CB13, CK13, CYJ13, CZZ14, CDN14, CXL10, Cha11b, Cha10a, CPY10, CL11b, CL11e, CHL14, CMR11a, CGS11, CG13, CDQ14, CDKN14, CDKN15, CSZ11b, DH11, DAE13, Dek10, DC10, DCZ10, DW10, DMB10, DKH^{+}10, DJ10, DR14c, Eba11, EGON10, ESM13a, EG11, EN14, ESV10, EGH12, FBB14, FdLW^{+}11, FLHL14, FS14, FP13b, For10, Fu10, GCH13, GW14b, GVF^{+}13, GC11, Gen12, GQL14, GM14a, GGH^{+}14, GY11, GLWF13, GS12, yGZ13, Gus14, Gwi13, HHT12, HPR12, HKL11]. problems [HS10a, HLC11, HR10, HR12, HVV11b, HTH10, HCWsO13, IO14a, ISN10, JK10a, JK10b, JWW10, JWL13, JR10a, JKV11, JMS10b, JTV14, KP14a, KSK11, KMHM14, KKK14, KC11, KPKK12, KKN11, KNY10, KWW11, KLM14, KPW10, KR14, Kur10, KM10, LEKT10, Lay13, LTW10, xLwCL10, LHL10, LJCL11, LZL11a, LS11, LW13, LGM14, LL14a, Li14, lLfD10, LM12b, LZ12a, LZ12b, LLT^{+}14, LMY11, LMYJ12, Liu12, LL13b, LZ13d, LYD13, LY13b, LO14a, LO14c, LXSC11, LC10, MY11, MS11a, MKS12, MVV14, MRD13, MMF10, MPS11, MV12b, MV14, Miy10a, MTJ11, Mol10, MS13b, ML10, MRS10, MPC10, MJ14, Mui10, Naz11, NTPVNHLX14, NV11, NL10, NCCF^{+}10, OCU11, OA14, OG10, PTH10, PW12, PLN13, PNCC10, PT12, PP10a, PL10b, PAJ10, QT13, QCK10, QY10, RG11, RB13, RS10b, Ren10, Reu10, Rua13, RX10, RVPOR^{+}11]. problems [SAY13, SSW14, SBFC10, Sch14c, Seg12, SCRG12, She10, SHZ10, SWY14, SHZ12, SB10, SDK10, SCDA10, SG12, SZS11, SSK^{+}13, SZ11, SBGB10, SLK14, Tan14a, Tar10, TKW12, VS14, Wan10, WG10a, WCW10, qW11, WSWD11, Wan12a, WY13, WWG10, qWhY11, Wil13, WL14a, XZW14, XCY10, XMC11, Yan10a, YJ11, YZ11a, YKL11, YK11, YSW13, YL11c, pZxY12, ZFG10, ZL10, ZM10, ZZ12, ZS13, ZFM14, ZLW14a, ZW12, ZBC13, Zho11b, ZC10c, ZLC10]. procedure [DDG13a, Li12, Xie12]. procedures [BRZ10, GN11, MS11a, SM10, SM14b, WZ11b, Yan10a]. Procesi [MM12b]. process [BSZ13, CFJ^{+}11, KFR13, LWN13, PO11, RE10, WZ11b, WW12, pYXPH10, ZWZ11]. process-integrated [KFR13]. Processes [Rie13, AB13, EY14, FJ10, GHR10, HL14, JZ10, KWXZ14, KSRM12, LP12, NR14b, PC12b, SKKS12, YSW13, ZRW11]. processing [AY13, DDG^{+}13b, Jia10, Jia11, LGMA13, MR13]. Procrustes [JWW10, QW10]. Product [ASST10, Not14, Wu10, AB11b, AWF11, AA14c, BFLP12, BwKB^{+}14, GP11, GJ13, KS14b, LRZ14, MJG14, Nad11, NH10, Tyr10, VVV10, VL13]. product/quotient [VVV10]. production [HD13, TNN10]. production-inventory [HD13]. products [KF11, LWH^{+}11, LA11, MR12, Pop10, SCRB^{+}14]. Professor [MYZ10]. profile [CDTV11, PM10]. profit [BOR^{+}12, Liu11b]. program [AÖ11, LPGX10]. programming [APW14b, CG12, FF14, GD14, LLT^{+}14, LLW14, LCW10, PTSK13, SY11, SK14, YT12, ZTS11, dO10]. programs [CG10a, CS12, LTZ10, LZL11b, YWL14]. progressive [AHAH11, ADA13, BP11a, Fer14, LWH^{+}11, LTZ10, WSS14]. progressively [Ism14]. project [AÖY10]. Projected [HC11, ZC10a, YSQ11]. projecting [CW13]. projection [CCG12, CY11, DSN14, HHL12, ISN10, JK10b, KN10a, LsGHZ10, LL11a, NRS10, PN11, PLN13, QCK10, RP11b, UM14]. projections [CLN14]. Projector [SL14a, TT12]. prolate [BDSG10]. Proof [HX14a]. propagation [ADG11, ÁD10a, AV10b, BFHK12, CC13c, HVV^{+}11a, HHP^{+}10, Sch14c, UBSG12, VVD^{+}14]. Proper [RSFS14, CLLX11, yGZ13, Zim13]. Properties [BR14c, Arv10b, BCHL11, BB12b, ÇMT14, CKM14, DST10, FDH10, KÖ11, KO13, LHL10, MT13, SHZ10, WX12, XhW11, ZDP^{+}14, Zho11a]. property [FZ11, LHL13, Sza10]. proportion [ÇU14]. proportional [AV12, CO13, YZ11b]. proportions [ÁAG^{+}13]. Proposals [Law14]. Prothero [Ran14]. protocol [SPLHCB14, SBM14]. proton [SS14e]. proved [Akt14a]. provinces [ÇDE14]. provisioning [VDG11]. proximal [HHX14, SS14a]. proximity [CD11, Mee12]. Pseudo [SC10b, BAAS11, BGVHN13, Che14a, DSL12, FP13a, JK10b, KPW10, NOS14, YG14, YG13]. pseudo-contraction [JK10b]. pseudo-contractions [KPW10]. Pseudo-differential [SC10b]. pseudo-Gompertz [YG14, YG13]. pseudo-inversion [NOS14]. pseudo-orthogonal [BGVHN13]. pseudo-parabolic [FP13a]. pseudo-random [Che14a, DSL12]. pseudoconvex [SK14]. pseudorandom [Har11]. Pseudospectral [AT10, yGlJ11, CSZ11a, KHKM14]. PSF [DM12]. Pták [CHMT10b]. public [AA14b, AK14c]. PUFEM [LEKT10]. pulmonary [ÇDE14]. pulp [SH13a]. pulse [BMMR11, Beh14, SI11]. pure [Dos14, YC10]. purely [JL11, ZRK^{+}12]. pursuit [SZJ12]. put [ZYL11]. Pythagorean [BBK^{+}14, KL11, SF12]. Pythagorean-hodograph [BBK^{+}14, SF12].

qd [AGT12]. QP [LFY11]. QP-free

[LFY11]. QRUP [GGSP10]. QTT [RST14]. Quading [LQSL11]. Quadratic [QC14, AHR14, ARF14, BBP14, BGRS12, BP12, CS12, CK13, DS10a, EHR10, GHR10, HL13b, xLwCL10, LT11, dL14b, Lu13a, LCW10, MS11c, Naz11, Owa12, XCY10, YD11, ZCX14]. quadratic-transformable [CS12]. Quadrature [MS12, CA12, CR14, CIPS14, Dek10, ESM13c, EJJ11, EALA11, FS10, GCC14, HZ13, HLFP14, KX13, Kni10, LZ13a, LYD13, MO10, Mel10, PS14a, SH11, SHN13, SH12, Sid14, Śmi11, SP10, Spa12, WSJ14]. quadrature-based [CR14, Śmi11]. Quadratures [BGVHN13, CBGVPP10, LLK14]. quadrilateral [FKNS11, Jia10, Lam14, LE10, PSS13]. quadrilaterals [yGlJ11]. quadrotor [BBP14]. Qualitative [HKL11, JJL11]. quality [AVMVMV10]. Qualocation [RU13]. quantification [HLF14]. quantify [SH13b]. Quantile [AMPP11, Hos11, JWY11, KG14]. Quantum [ÁE11b, MPS10, ÁE10, CFGE10, Lay13, LJ13, Vel10]. quartic [HwK11, LX12, Str10]. Quasi [ABN10, EZ10, GW12, MNQ14, Pov14, ABIS10, ASST10, AR10a, AELH10, BISS10, BGIN10, BGIN13, BGRS12, BHL11, BDR11, BR14c, CCH12, CC13b, CF11, DS10a, DSZ11, DS10b, FJ14, FPP14, FZ11, GW14d, HPL^{+}14, HJS14, ILMS10, KO13, KW10, LMT13, MS12, PY10b, QXW10, QCK10, VL13, WZA10, XSH10, ZY11b]. quasi- [QCK10]. quasi-consistent [KW10]. quasi-interpolants [ABIS10, ASST10, BGIN13, DS10a, ILMS10, LMT13]. Quasi-interpolation [GW12, BISS10, BGIN10, BDR11, BR14c, CCH12, FZ11, GW14d, MS12]. quasi-linear [BGRS12]. quasi-minimal [DSZ11]. Quasi-Newton [Pov14, AR10a, BHL11, CC13b, KO13]. quasi-Newton-least [HPL^{+}14]. quasi-Newtonian [ZY11b]. quasi-nonexpansive [WZA10, XSH10]. Quasi-optimal [MNQ14, HJS14]. quasi-optimality [FPP14]. Quasi-positive [EZ10]. quasi-product [VL13]. quasi-reversibility [QXW10]. quasi-static [CF11, DS10b]. Quasi-stationary [ABN10, AELH10]. quasi-uniform [FJ14]. quasi-variational-like [PY10b]. quasilinear [BG14a, FP13b, VT14]. quasistatics [TKW12]. quaternary [EI14]. quaternion [JWL13]. Quenching [Cha11b, BS13b, Cha11a]. queue [KKK12, YTFP11, ZH10a]. queueing [CLS12, YWW10]. queues [WYP10]. Quintic [CW12, CW14a, CW11a, CC12, DY13, LMT13, Lu13b, NRS11]. quotient [AGT12, BE11, Jia12]. quotient-difference [AGT12]. quotients [Zha14a].

R [MCS12]. Rabinovich [LYP10]. Racah [ÁNA14]. radial [AE11a, AM14a, AAD13, BLM11, Boy10, CGCG^{+}13, DS13b, FND10, JGK14, KG13, KFR13, KLL11a, MK10, QYZZ14, RLGGGÁM14]. Radiation [HWG10, BBD^{+}11, DME^{+}12, FYA^{+}14, KP11b, MT10b, VdKPB10]. radiofrequency [LLG14]. radiographs [HLF14]. radiosity [DHSW14]. railway [RLGGGÁM14]. Random [Hui10, AO14b, BP11b, BAAS11, Che14a, DSL12, Ery13b, GW14c, GAS14b, KM11, LQYL11, LZ13a, LVCC13, Nad11, OSR13, SRP^{+}13, Wei11, Wik10, ZH11, ZWZ11, Zho11b, ZwGhX14]. randomized [WYP10]. randomly [Mau10, SL12]. range [CLP10, Hui10, OP12, Sid14, WS13b]. rank [AZ11, BB12c, CS12, CCS10, Dax10, DLWZ14, Gal13, Gat14, JWW14, LZH14, LHC14, MU14, MML11, UM14, VVM10, WH10, WH13]. rank-one [MML11]. rank-two [CS12]. ranked [BAU13]. ranks [Tyr10]. Raphson [Gat14]. rapid [LWXW12, SKK14]. rare [NR14a]. rate [AH14b, BBMY11, BBS14, BÜ14, PS11c, SSR10, TNN10, XYW14, YYL^{+}10, YTFP11]. rates [DGI^{+}10, MNQ14, SB10, TYCL14]. ratio [JJL11, PS14c, WF10]. ratio-dependent [JJL11, WF10]. Rational [Boy13, BVV10, CBGVPP11, RSS14, AHM14, Ali11a, BK14c, BC12, BDH11, BNRZ12, BL10c, BGVHN13, CW11b, DP10a, DP11, FKL10, GK11, HX13, IO14a, JJK^{+}13, JR10b, KHKM14, KPV14, LLL^{+}11, Lu11, NH13, Van10, VCA14, WCW10, WW11a, XZL11, XH13, XL12, Zha14b, ZLG12]. ratios [JK10a]. Raviart [BMN10]. ray [HLF14, VdKPB10]. Rayleigh [CFJ^{+}11, Jia12, LWH^{+}11, Xie12, Zha14a]. Razumikhin [PJ12]. Razumikhin-type [PJ12]. Reaching [AK14d]. reacting [RVDM10]. Reaction [FdOS12, AK14a, ANS11, AD10b, BV10, BS13a, BS13b, BFB14, CHMD14, CGL10, CG13, Dra13, DLP13, ESV10, GPHAPR14, HMS11, HMVPAR11, KR14, LWZ13, LJ11, Pop10, RK11, RRMD13, SYK14, SB10, WZ10a, Wan11f, ZD14, ZC10c, ZLC10]. reactions [CHMD14]. reactive [Emm11, RP10, SB10]. Real [DJJ13, BVV10, BGVHN13, CBGVPP11, HDT11, LWZ^{+}14, MA13, MT10a, MBS10, OR14, OG10, Str12]. realization [FO10]. reblurring [DDE14]. recognition [MVC10]. recombining [LLK14]. reconstructing [Ger12]. reconstruction [ADLT13, CSL^{+}11, GKL^{+}12a, HSH12, KNR12, OGM14, PS10, WHLC11]. record [WS13a, YG14]. recording [BPS12]. Recovering [CS14, JWW14]. recovery [BSKL12, BR13, Cao14, CCG12, CN10, Gos11, HYY14, LZH14]. rectangular [AIA10, AAD13, Bay14, CBB13, GJCL11, WZ11a]. recurrence [Cas10, DST10, DMGVPN10, Kni10]. recurrent [LF10]. Recursive [MV12a, BDD14, CLS14, DSL12, MY14, PWZ11]. redistribution [KL10b]. Reduced [BBKS10, AHLJ14, CDQ14, DDM10, GK14, LXSC11, MPC10, OCU11]. Reduced-dissipation [BBKS10]. reducible [ABN10]. Reducing [CH14]. reduction [AG10b, BMS11, BSB14, BGKVW10, GPVL10, Har11, ILEE13, Lu13a, NRS10, PPG14, QLL11, SMD14, WS13b, WL10d]. reductions [LL14b]. Refinable [DY13, GPS13, KS14b]. refined [NY14]. refinement [BG14b, CFJ^{+}11, Li10, MA11, RW10]. refinements [HKK10]. reflected [RE10, Xu11b]. reflecting [Dea11]. Reflection [Mau10]. Reflectionless [LL10]. reflexive [HY11, JWW10, lLfD10]. refraction [FGEM10]. refuge [Reb12]. Regime [MD10c, CLMR14, DS10b, ESC14a, JWY11, LZ13e, YY10, Zhu14]. regime-switching [CLMR14, JWY11, LZ13e, Zhu14]. region [CWQ11, JLZZ10, LJ13, OCU11, WDG^{+}12, Zha10b, ZCX14]. region-based [LJ13]. Regions [GKLZ12, LDGY10]. registration [CC11b, LGMA13]. regression [DAH11, GCNL14, MOA10, MSW14, Nie13b, ÖW14, Pit13, XHmDW14]. Regressions [Hos11]. regressive [AMPP11]. regular [BS11b, GM14a, Hua14b, SSP12]. Regularity [Fad10, Rah14, LMYJ12]. Regularization [GP10, LPDN13, LXX14, Més12, BG13, CHL14, Chu10, DR14b, HPR12, JKS13, LV13, LL14a, MR11, QXW10, WHLC11, Wan11b, Wan12b, Xio10, XZH14, YF14, ZFM14, ZW10]. Regularized [BG13, Cai10b, MM11b, SWT12, ZC13, ZLL10b]. regularizer [XHmDW14]. regularizing [ENSCS12]. regulation [AVMVMV10, LZT13]. Reichel [GGSP10]. reinforced [VB14]. reinsurance [LRZ14]. reinsurance-investment [LRZ14]. reinsurer [LRZ14]. Reissner [WLL^{+}11]. REIT [RSK14]. related [AS12b, CGS13, FLLT10, GGH^{+}14, KPA14, Qi14, SL14b, Zha11c]. relation [Mol10]. Relations [ZZ12, Cas10, DST10]. relationships [CFdAR11]. relative [ÁD10a, ÁD12]. relatively [WSWD11]. Relativistic [KS10a]. relaxation [BY11, BS11a, BSS14, JD13, Kho11, Kok11, LJ11, LJ12, LZ13c]. relaxed [PD13, PR11]. release [FdOdS12]. Reliability [WS13a, Ery14, FPG12, OSR13]. reliable [AAI14, SBFC10]. Reliably [WM12b]. Remainder [KS14a, Bel13, Qi14]. remapping [BBKS10]. Remarks [De 14, HV13]. remodeling [FM11b, MP10b]. removal [CC14a, LZ13c]. renewal [hBW10, JZ10, San11, ZZ13]. repairable [Ery14]. reparameterization [HX13]. Repeated [Bal14]. replacement [CL14b]. Representation [Ger12, CKM14, DFG^{+}10, FJ10, HV11, IK13, MT12, WSK^{+}11]. Representations [vDvFZ09, vDvFZ13, MP10a, Qi14, SWY13, Zho11a]. Reprint [CSSX14b, GZ14b, LO14c, Sch14b]. Reproducing [BM11b, DZZ14, GC11, LW13, MM11b, SD13, YSW12]. reproduction [CC13a, CR11b, FZ11]. research [GKL12b]. reservoir [DDG^{+}13b]. reset [CMR11b]. Residual [BMS11, ANS11, Che13b, DSZ11, HVV10, LZ10, PW13, RA10, ZZG14]. resistance [KMD12]. Resolution [NCCF^{+}10, AH14a, Bab10, BKTY14, MM11c, VTLF11, Wri10]. resonance [MTJ11, ZM10]. resonant [TA10]. resource [LNJ^{+}10, LNB^{+}12]. respect [ÇDE14]. Response [PS10, AHM13, AD10b, DF14, PVM^{+}14, TL13]. responses [WF10]. restart [LT11]. restarted [NY14]. restarting [GZ13]. Restoration [CCS11, AMRS12, AVMVMV10, BEJ14, DDDSC14, DDE14, HL13b, JMV13]. restricted [HR10]. restrictions [LMM^{+}10b]. result [CPP11, EHR10]. resultant [WH10, WH13]. Results [HD11, Ayd14, CSZ11b, EX14, Fad10, GPHAM10, Hui10, Kha10, LTW10, LHY10, LCI11, SL12, TT14b, Wik10]. retarded [LHL13, WZL11]. retargeting [LZSL11]. RETRACTED [Beh14, DS12b, XSV12]. Retraction [DS13c, Wan11c, WYO11, XSV13]. retrial [KKK12]. return [PP14, XYW14]. returns [Liu11c]. reusable [WRS14]. Revelations [DP10b]. reversed [WZS10]. reversibility [QXW10]. reverting [PO11]. review [NSZ10, Pit13]. revoke [APS14]. reweighted [LZH14]. RFID [SPLHCB14, SBM14]. rho [WZ12]. Ribière [YWL14]. Riccati [BHJ11, GG14, LwCLW13, LVCC13, PWZ11, WLD14, YLD13]. Richards [KMLT10, KMP14]. Richardson [CG13, ZFH10]. Riemann [HHM12, KAE^{+}11, MA14b, Rza12]. Riemannian [ABCP14, BSBA14, PTSK13]. Riesz [ZD14]. right [DL12, DSY^{+}11, HD11, LWH^{+}11, lLfD10, MZL14, WM12b, ZDZ11, ZG14]. right-hand [DL12, DSY^{+}11, HD11, MZL14, ZDZ11, ZG14]. rigid [MT14b]. rings [KFR13]. Risk [ÇDE14, DH10, BBB10, hBW10, CP13a, CO13, CGW14, CW14c, FM14, GL10, GW14c, GSK11, JZ10, LL11c, LB14, SH13b, TT14b, WW12, hXZ11, XYW14, YZ10a, YYL^{+}10, YLS12, ZYL10, ZY10, ZY11a, Zha14c, ZZ13, ZwGhX14]. risks [DLSV14, HW10, KG14, KH13, PR12b, YLS12]. risky [AÖY10]. Ritz [CFJ^{+}11, Xie12]. river [AVMVMV10]. RKDG [SR13]. RMARS [ÖW14]. Robinson [Ran14]. Robust [AU14a, BMN10, GY11, GG13a, fLxX12, Wil13, ZCH10, CL11a, CC11b, DP11, ESV10, FRB13, KR14, LS14, LdHA10, PP14, Pit13, Spr14, WZ10c]. Robustification [ÖW14]. robustness [KO13]. ROCK [KB12, KB13]. rod [CR11a]. Roe [TLNE10]. role [KH10]. rolling [KFR13]. ROM [AHLJ14]. ROM-POD [AHLJ14]. root [DP12, DP14c, GKLZ12, KP11a, Owa12, PPH10, Yan10b, Zie14]. root-finding [PPH10]. root-solvers [DP12]. rooting [Str12]. roots [WLH12, Win14, WM12b, Yun11, Yun12, ZZ11a, ZCS11]. Rosenbrock [BSB14, Gon10, GPHAPR14, KL14, LO14b]. Rosenbrock-type [GPHAPR14]. rot [KP14c]. rotary [KBKR14]. rotating [GRV11, HL13a]. rotation [MKD12]. round [CH14]. round-off [CH14]. rounding [Rum13a, Rum13b]. roundtrip [PVM^{+}14].

Row [GG10, PGGCGF14, Ran14]. RQEA [LJ13]. RSTLS [DM12]. ruin [AAK10, ACH11, YLS12, ZZ13]. rule [BFLP12, Car11, CIPS14, DS10a, LWY10, LYD13, WDZ10, ZFM14]. rules [Dom14, GP11, HPR12, LZY10, Not14, SH12, SL12, WSJ14]. Runge [LO14c, BR11b, BB12c, But12, CFMR12, DP14a, DR14a, FBB14, FG14, GPHAM10, HVV10, HHKK14, HSY13, KKR11, Kha10, KB11, LL14a, LO14a, MK11, NTNBV14, PR10, SCRG12, Xie11]. Running [DP10a, DP11]. Russell [LNJ^{+}10].

S [KB12]. S-ROCK [KB12]. Sabin [BFGP10, DY13, FGPP13, LMT13, SCFR14]. Sabin- [DY13]. saddle [BMS11, Bai13, BBB10, BM11a, CYJ13, Cao13, CDN14, FdLW^{+}11, LHL10, LZL11a, PL10b, SHZ10, WL14a, ZS13, ZLW14a]. Sakurai [ISN10]. salesman [MB11]. saliency [LZSL11]. Saltwater [WDG^{+}12]. Sample [Li14, Lu11, MSW14, KK13, MD13]. Sample-based [Lu11]. samples [LWH^{+}11, PVP13]. Sampling [HLF14, AO14b, BAU13, BK13, Prz13, WSK^{+}11]. Sampling-based [HLF14]. Sánchez [MYZ10]. satisfying [Cas10, FZ11, San11]. saturated [BGLV11]. saturation [Boy10, Fad10]. Saving [GWX10]. scalable [DKH^{+}10, GG13a]. Scalar [KP14b, BGP11, CC13a, LB11, MSZ13, PGGCGF12, SLP11, SR13]. scale [AG10b, BOR^{+}12, BKTY14, BHJ11, BGKVW10, CXH09, DHSW14, DdS14, EY14, ESM13b, FGO11, GKB10, GGH^{+}14, LV13, LwCLW13, MA11, NL10, PM13, SW12a, YPL10, ZH10b, ZJ12]. scaled [BK14a]. scales [DHS11, GZL14, LH11a, WZMC11, ZM10, Zha11b]. scaling [CC13b, GG10, Kho10, LZ10, MMS13, OGM14]. Scattered [ZL13, BR13, CD10, FZ11, GST12]. scatterers [GST12, Mau10]. scattering [BP13, CDK10, DS11, GH10a, HNPX11, Kre95, Kre11, LMCW10, NA11, TA13]. Scenario [XCY12, BDV10]. SCFT [XZ14]. Schatten [LHC14]. Schauder [BGL13]. Scheduling [NGAZ11, HTM14]. scheme [ADA13, AHS14, ÁD11a, And10, BPS12, BDD14, BB11b, Bog14b, BG10b, Cai10b, CS10, CBB13, CCG13b, CXL10, CWY11, CGL10, CET14, Col13, CwYyY11, CyYwY11, Cui14b, DLG14, DS10b, DF12, EO14, EALA11, Fad10, FDD10, GW14d, GSGN11a, HL10a, HKO11, HK13b, IG12, JS10, JK10c, JW10a, JR14, KAE^{+}11, KLNS11, KBKR14, KPW10, KS12c, LLS14, LCCK13, Li12, Lia14, LvDtTB^{+}13, MHL11, Mat10, MT13, MDVB13, NCCF^{+}10, RMP^{+}13, RVDM10, RT12, SLP11, TY11, VMMM10, WG10a, Wan11a, Wan11d, WLD14, WC14, XSH10, ZBW11, ZTC11]. schemes [ADL^{+}10b, ADL^{+}10a, ADLT13, ADLS14, BK13, BB12a, CI13, CC13a, CL11b, CR11a, CCS10, CD11, CTV13, CM10c, DSC09, Dai11, DGG12, DH11, DH14, DFH14, DGW10, EZ10, EN14, FP13a, FJ14, GWZ11, GKT11, HMVPAR11, Kuz12, LE10, LM13a, LSW11, lLzS11, MF14, MS11b, MM12b, MA11, MK11, NNO12, SM11, Tan14a, TS12, TS14, WZA10, YMS10, YLD13, ZAT11, ZZS11, ZD10]. Schmidt [Liu11a, LZ13d]. Schmidt-based [Liu11a, LZ13d]. Scholes [CL11a]. Schröder [PPH10]. Schrödinger [KRS11, AS12b, AKT12, AKT14b, AK11, BKV10, CCG13b, CN11, EK11, JW10a, JWY14, Lak13, SKAM11, WZF11, Wan11d, ZBW11]. Schrödinger-type [Lak13]. Schur [BL10c]. Schwarz [CY11, GPZ10, KR14, Lui10, XN14]. Science [AVA13, VAV11]. Sciences [SKS14]. scientific [BDR10]. scrapped [TNN10]. screening [MA13]. scroll [BFB14]. Sea [MKK14]. Search [BJ14, CDJP13, CWQ11, DSL12, JK10a, KK13, Lay13, NY12, WTM12, Zho13]. searching [YTFP11]. Seberry [AHS14]. secant [AEG^{+}11, BKGMA10, Gal13, KNY10, LP13, MML11, NY12]. secant-type [AEG^{+}11]. secant-updates [Gal13]. Second [CRM10, DSC09, Dai11, LSW11, ABS11a, AGM12, ASS12, AS12a, AAD13, BDD^{+}10, BB11b, BGGG^{+}11, BG14a, Cai10a, CXL10, CL10b, CFL^{+}14, Dom11, Dos14, EZ14, FKL13, Fuk11, Fuk13b, GC11, GQ14a, GP12, GD14, HSY13, HHZ11, HNK10, IK11, JR10a, KHKM14, KLL11a, KB11, KB12, LLS14, LS11, LZ13a, lLzS11, LZ12a, LB09, LB10, Mah14, Nei14, OMS10, PM13, Ren10, Rua13, SASH11, Tan10, WYP10, WY13, XL12, ZFG10, Zha11b, ZF14, Zho13, ZX10]. second- [ZX10]. Second-order [CRM10, DSC09, Dai11, LSW11, AGM12, BG14a, CXL10, CL10b, CFL^{+}14, EZ14, GD14, HSY13, LLS14, LS11, lLzS11, LZ12a, LB09, LB10, SASH11, Tan10, WY13, ZFG10, Zha11b]. section [AVMVMV10, HHM12, PC12b]. sections [Hu12]. sector [GKLZ12, Zie14]. secular [BR14b]. secure [AHS14, NACLR12]. securities [JM12]. security [SBM14]. segmentation [BFMS11, RPS^{+}11, ZZSW11]. segments [CC12, FSB10, Lu13b]. Seidel [KN10b]. seismic [BUG^{+}13, DDG^{+}13b, MKK14, MR10, UBSG12]. Selected [AVA13]. selection [CC14a, CW14c, fLxX12, Liu11c, VDG11]. selective [Ven10]. self [ĆRR^{+}10, FiMO10, KKR11, Zha14a]. self-adaptive [FiMO10, KKR11]. self-consistent-field-like [Zha14a]. selfadjoint [CO11b, Yar10]. Semi [BSS14, Fat11b, WA14, ZLW14a, CGW14, Jan14, KP10a, LLT^{+}14, LJ12, PP14, PGH13, RT12, Wei11, ZAT11, Zha11c]. Semi-analytic [Fat11b]. Semi-analytical [WA14]. Semi-convergence [BSS14, ZLW14a]. semi-deviation [PP14]. semi-discrete [RT12]. semi-explicit [Jan14, PGH13]. semi-implicit [Wei11, ZAT11]. semi-infinite [LLT^{+}14, Zha11c]. semi-linear [LJ12]. semi-Markov [CGW14]. semi-solid [KP10a]. semidefinite [DLWZ14, FF14]. semidirect [NH10]. semigroup [MZ14]. semigroups [Mia10]. semilinear [Bog11, Mui10, MK11, VS13a, Wan13, pZxY12]. Semilocal [GP12, AR11, AEG^{+}11]. semiparametric [DAH11]. semiseparable [BVV10, VVV10]. Semismooth [ZSX11, DA10, SZ11]. sense [CSY10, Col13]. sensing [CSSX14a, CSSX14b, ZSK^{+}11]. sensitivities [Sch14a, Sch14b]. Sensitivity [Kni10, WYY^{+}12, AHLJ14, DGI^{+}10, SBCD10, YWW10]. sensor [DCGS13]. separable [BDMR10, YT12]. separated [ZZ12]. separation [SZJ12]. separator [MAK^{+}14]. sequence [Bel13, ZWZ11]. sequences [AMPR10, AH12b, BL10c, DMGVPN10, EI14, EY11, ENSCS12, EGH12, bJfCwH14, Khr10, NLL13]. series [AA12, BCHL11, Boy11, Boy13, Che13a, CV10b, DJM10, Dom11, HNK10, IK13, MJ11, SL14a, SS14b]. server [EC13, WYP10, WK10]. servers [WK10]. service [TNN10, WYP10, YTFP11]. Set [BHMW13, MCS12, ADRM12, BAU13, BW14, BK14b, DW10, DGC^{+}13, FW10, FJLL10, GCC14, KWW11, Kuz14, LH11a, fLxX12, ÖY14, SK14, Sto11a, TJL11]. set-valued [ADRM12, DW10, FW10]. set/ front [BW14]. sets [CD10, CKM14, CGRS11, Fer14, FB12, GGS12, GÖ14b]. setting [GN14, LJE^{+}10]. several [CT10, FPPX10, HM11]. SFDEs [BBMY11]. Shack [KNR12]. shaft [DKU14]. shallow [BS10b, CBB13, DV13, YC11]. Shannon [Wan11b]. Shape [LMM^{+}10b, WUB10, DM10a, GPS13, JY13, wKK12, LBM12]. shape-preserving [GPS13, wKK12]. shaped [DA14, DMB10]. shaping [AHSN13]. Sharing [KS14a, JLZ14].

Sharma [PS14a]. sharp [IK11, PR12b, Rua13, SW12b]. shear [WTM12, YESD14]. shearlet [Gro13]. shearlet-type [Gro13]. sheet [TKW12]. sheets [KMD12]. shell [GCY13, WLL^{+}11]. Shepard [CDD12]. Shift [CDN14, AMMS14, CL13b, GWX10, VVM10, VBP^{+}10]. shift-and-invert [GWX10]. Shift-splitting [CDN14]. shifted [AM10b, KOV11, SL12, SSF^{+}11, Tor14]. shifted-Laplacian [AM10b]. shifting [CKM14, DC10, DCZ10, SY11]. shifts [EGST11, NY14]. ship [LA10]. Shiu [GW14c, JZ10, ZY11a]. shock [APO14, Ery13a, OB14]. shocks [IG12]. shooting [AHH13]. shops [NGAZ11]. short [Ehr11a]. shortest [AHH13]. shot [IK13]. shrink [DKU14]. shrinkage [PM11a]. shrinking [QCK10]. shunt [ZTL13]. side [DL12, SZ10]. sided [SKN11, YZ10a, ZYL10]. sides [DSY^{+}11, GNN11, HD11, MZL14, ZDZ11, ZG14]. sideways [LXX14]. SiGeC [LLG14]. sigma [DJM10]. sign [DC10, DCZ10, Hua14b, KC10, MY11, NV11, ZC10b]. sign-changing [KC10, MY11, ZC10b]. sign-shifting [DC10, DCZ10]. signal [AY13, CCG12, DDDSC14, Gos11]. signal-restoration [DDDSC14]. signals [Che14a, GLHCLP11, JPX11]. signature [Ery14]. signatures [LA10]. signed [WGZL14]. significant [MM12a]. silo [HDL^{+}13]. similarity [AHM14]. simple [And13, BS10b, Col13, FRB13, JW10b, KZ13, LGR11, LL11d, TJL11, YWY^{+}14, ZCX14]. simpler [CGS11]. Simplex [BSV12]. simplexes [WSJ14]. simplicial [JKK^{+}10, Vit10]. simplification [CGRS11]. Simplified [DR14c, ARF14, DL13]. simply [Vit10]. Simpson [CC11c]. simulated [AG10a]. simulating [ÁD12, BHMW13, HDL^{+}13, KAE^{+}11, KS14d, LM13a]. Simulation [BFHK12, DGC^{+}13, IG12, Bab10, BDV10, BSKL12, DA10, FRB13, GGAVRC14, GPVL10, GVF^{+}13, GÖ14a, GMRK14, GWKE14, GJCL11, GKT10, HdJL^{+}13, KWXZ14, KM11, KP10a, KH13, LVD^{+}14, LMW^{+}11, LvDtTB^{+}13, MKK14, MVMG12, MS13c, NR14a, NdCDCFM10, NMBG14, PW12, SLK14, VMAVGCM13, VB14, WL11a, WCZ11, YZ11a, YLTB14, ZZC11, FDD10]. simulation-optimization [BDV10, VMAVGCM13]. simulations [ÁD11a, AJK12, AP10, BW14, BCD^{+}10, Bra14a, DBHV10, yGZ13, HLLW13, KP14b, LXSC11, MDVB13, MR13, RVDM10, SW12a, VMS12, VSB12, WRS14, Xu11b]. simulator [WJJ14]. simultaneous [BG14b, LEVG12, LEVG14, PRM11]. Sinc [MA13, OMS10, PDP13, HX14a]. Sinc-collocation [OMS10, PDP13]. sine [Cui10, MD10b]. sine-Gordon [Cui10, MD10b]. single [AHM13, GUO14, PPG14, TNN10, TA13, VTLF11, WW11a, qWhY11, YWW10]. single-layer [TA13]. single-machine [TNN10]. single-phase [PPG14]. singly [DP14a, MK10]. singly-excited [MK10]. Singular [MMX10, Zhu14, AMMS10, AMMS14, ASST10, AS12a, AIA10, BH11, Bar13, BK10, Cai13, CA12, Che10, CL13a, CC14b, CA14, DMR13, Eba11, ELA10, EL11, EALA11, Gat14, HR10, JX10, JJE13, KXH13, KNR12, KLM14, LSD11, LHLL10, LLP12, LM12b, ML10, NY14, OMS10, PV10, PGH13, Prz10, QZX^{+}11, RD10, Ren10, WCW10, qWhY11, Wu14, WZ14, ZS13, ZLW14a, ZC10b, ZC13]. Singular-value-like [MMX10]. singularities [Boy13, CS14, Dom14, Esc14b, HS11a, KP14a, KX13, OPPV10, WWG10, Yar10]. Singularity [LHLL10, DZZ14, HPL^{+}14, ML14, Pag10, RR10]. singularly [Bog11, Bog14b, CXL10, CGL10, CG13, GQL14, KR14, RK11, RU13, WA14, ZC10c]. SIR [SI11]. SIS [AELH10]. situations [GPO14]. Sivashinsky [GP10, VMMM10]. six [MJ10a]. sixteenth [GK11]. sixteenth-order [GK11]. Sixth [HH14, CLS10, WZ10c]. Sixth-order [HH14, CLS10, WZ10c]. size [AALMLM14, KKK12, KA12, KK13, MD13, MSW14, ZZS11]. size-structured [AALMLM14, ZZS11]. sizes [HHL12, ZY11a, ZwGhX14]. skew [CL10a, PZ13, RRMD13, SC10c, ZM14]. skew-Hermitian [PZ13, ZM14]. skew-symmetric [CL10a]. skewness [ZZP13]. skin [Mee12, ZGZB13]. slab [Mau10]. slabs [Tau13]. slicing [XhW11]. sliding [DD14b, DD14a]. slip [LL11a]. slope [HdJL^{+}13, Kuz10, LFY11]. SLP [HXW11]. Smale [AHK14b]. small [BS13a, Bra14b, ESM13b, JŽ11, JLZ14, Kni10, MM12a, Par10, Pas11]. small-scale [ESM13b]. smallest [NY14]. smart [GGAVRC14]. Smooth [CHZ10, AD11b, BT12, CS14, Dek10, DL14a, LYP13, SS14a, SSR10, ZWM11]. smoother [XN14]. smoothers [PC12a]. Smoothing [YP10a, DR14b, HVV10, LM12b, LG11, LH11b, LdHA10, MT13, QYZZ14, RX10, XMSS11, ZLL10a, ZH10c, ZLL10b]. smoothing-type [LH11b, ZH10c, ZLL10b]. smoothness [SR13]. snake [CSL^{+}11]. snl [MK10]. snow [NdCDCFM10]. Sobolev [BDF^{+}10, BCGMB10, DHS11, DM10b, SHN13, Wih11, dJP13, dJMPPC14]. Sobolev-type [DM10b]. SOCCP [LH11b]. soft [FJLL10, KWW11, QH10]. Software [MU14, LA10, MCS12, VMAVGCM13]. solar [DME^{+}12]. solectron [Vel10]. solid [GCY13, KP10a, NTPVNHLX14]. solid-shell [GCY13]. solids [PTP10]. solitary [SC10d]. Solodov [ZYG10]. solute [RP10]. Solution [AU14b, AP14a, Beh14, PM13, AE11a, AM14a, AS12a, AV12, AAD14, ATS14, AuI13, AuIK14, BH11, BS13b, BSZ13, BGL13, BPGP11, COW10, CL11c, CDV10, DS13a, DDM10, DS12a, DAE13, Dos14, Dra10, Dra13, DS11, DR14c, EGES14, FLLT10, GG13a, GTS14, GKL12b, HLM10, HHT12, HHP^{+}10, HKL11, HS10c, HS12, IK11, JV10, JTV14, KP14a, KHKM14, KLL11a, KSLD11, KA14, Kre95, Kre11, KT13, LD10, LSD11, LWXW12, LZ13d, Liu14a, LYD13, MMF10, MA13, MMS13, MMSW13, MN10, MV14, Moh13, Mön10, ML10, NLO13, NPTH09, OAHR11, Owa12, PT14, PKH14, Ram10, RGJ10, RG11, RST14, Rum13a, Rum13b, SSP12, Sha14, uIAAF14, SK14, SKN11, TSC10, VT14, WD14, WWC14, Xu11a, YY11]. solution-dependent [HKL11]. solutions [AF10, AS11, ASK11, ADRM12, AVGCMVM10, Alz10, AS10b, Asl10, Ass09, BS10a, BBMY11, BK10, BB11c, BC12, BHJN10, BKV10, CYJS14, CCDS14, CZ12a, CLS12, Cha10a, CCJC13, CL11e, CL14a, CN11, Chu10, DN13, DBZ10, DF14, EK11, EZ14, FKL10, GXW14, HZ11, HHH^{+}13, HHHS14, HMS11, HR12, JWY11, KC10, KCBA13, KL12, Ku13, Kud10, KRS11, LS11, LLCL11, LHLC13, LWW13, LVCC13, LWZ13, LM13b, LZ13c, LL14b, MT14a, MA14b, MJ11, Miy10b, MR13, N’G10, OCU11, yP10b, PV10, RRMD13, She10, SC10c, SGG11, Sie11, SRWZ14, Tan10, TA10, WZS10, WZ10a, Wan11c, WC11, WZL11, WX12, WWG10, WZ10d, YT12, Yan11, YFWH10, YZ11b, Zha10a, ZFG10, mZkPkL11, ZF14, ZC10b, ZRW11, ZC13]. Solvability [HSY13, LH11b, ZM10, AAM14, DJ14]. solve [CV10a, CDS10, Gat14, LV14, LW13, Liu14b, MR11, PMC11, RBD11, RBD14, SBFC10, ZBW11]. solved [GA12, Lak13, Seg12]. solvent [MA13]. solver [Cai13, CLS14, GVM^{+}12, GM14a, GCH14, HHM12, KAE^{+}11, KOV11, Man11, NMS10, NBBKV10, PP10b, SM14a, Str10, TLNE10, VAR11, YMB14]. Solvers [MCS12, DP12, HVV^{+}11a, HHLW14, KK14, SKK14, Wil13]. solves [Koç11]. Solving [AGT13, Bar11, BR14b, DHSW14, ESM13a, EHR11b, FLHL14, FV10, GGH^{+}14, JK10a, KMP14, LWW11, LwCLW13, MKS12, MM11b, QW10, SSW14, SP11, Str11, YK11, AEWMES11, ABS11a, AS10a, AAI14, AT10, ASS12, ÁECS12, AHH13, AH10b, AH13, AAD13, BDD^{+}10, BMMR11, BBP14, BGGG^{+}11, BC10, BP12, Bog14a, BJ12, CS12, CS10, Cha10b, CSZ11a, CZ12b, CC14b, CSSX14a, CSSX14b, CHMT12, CTV13, CHMT13, CyY10, DH11, DO12, Dha14, DCJ13b, DZZ14, EK12, EDP14, FG14, GZ10, GSGN11b, GSGN11a, GSNDB14, IHRA11, IR11, JK10b, JLZZ10, JX10, KS10b, KPKK12, KKPV11, LsGHZ10, LLP12, LW11, LZ14b, LY13b, LFWL14, LCW10, Mah14, MRD13, MA14a, McN10, MMS10, MM11a, Mol10, Mui10, Naz11, NO12, NOS14, NLL13, NBZ13, NASNY11, PLN13, PDP13, PT11a, Reu10, RP11a, Ros11, SH10]. solving [SXZ11, SB10, Śmi11, Ste10b, Sto11b, TP10, TY11, WL10b, WCW10, qW11, Wei11, XW12, XMC11, YPL10, YSW12, YBD11, Yun11, Yun12, ZH10c, ZS13, ZMZ13, Zha13, ZFM14, ZD10, ZLL10b, ZLC11]. Some [Ayd14, BK13, BHJ13, BB12b, BCGMB10, GQ14b, HV13, HHL12, KÖ11, MO10, Śmi11, Ste10b, Sza10, XC14, Yan10b, ÁD11a, Arv10b, Bar13, CFMR12, CGRLR11, GG10, GSGN11b, IR11, JMS10b, Ma10, NH10, PS14c, QGG10, QT13, QW10, Rza12, Sid14, TM11]. Sommeijer [Ano10b]. Sommerfeld [WTM12]. SOR [NK10, CZL14, yS05]. sorption [RP10]. sound [Uty10b, Uty10a]. source [Cha11a, DS10b, GI10, HTLI14, HCM14, LA10, MKK14, MCS12, SZJ12, SM11, XW12, YF14]. sources [FDH10, ZLC10]. south [KG14]. SP [PS11c]. SP-iterations [PS11c]. space [BL13a, CLLX11, CDKN14, CDKN15, DXL10, El 11, FJ10, HML14, KP14c, Law14, LPGX10, MLZ14, MSZ13, MM11b, MS11c, MS13b, MT14b, PSS13, SYİ11, SYK14, SH11, SHN13, SKN11, WWL10, Wan11a, WD14, XHmDW14, ZAWA13, ZD14]. space-fractional [SKN11, WD14]. space-time [MS13b, SYK14]. Spaces [BM11b, AHSN13, AB11b, BMMR11, BDR11, BR14c, CB13, CA12, ĆRR^{+}10, DDG13a, DCJ13a, DW10, DHS11, EGON10, EHRV13, FO10, GP12, HK11, JY13, LC11, pLjH10, PS11a, SF12, SACS10, jTW14, WG10a, WZA10, WSWD11, WZY^{+}14, YKL11, ZFG10]. spacetime [MA14b]. spacing [AV10a]. Spalart [DH13]. Sparre [TYCL14, Zha14c]. Sparse [FV13, NA11, WSK^{+}11, AAI14, BDL11, BHPS12, CCG12, CLS14, DFG^{+}10, GZ11, GGSP10, KR10, Man11, MVC10, NMS10, RVN13, ZGZW10, ZM14]. Sparsity [GKL^{+}12a]. Spatial [KLL11b, Bal14, BBK^{+}14, CC11a, ÇDE14, GKU14, YF14]. spatial-dependent [YF14]. SPDE [KH10]. SPDEs [Lan12]. Special [FHMZ10, Gra11, MYZ10, YFZ14, BDM^{+}14, CM10d, Gau10, Van11a]. species [FLEa10, SB10]. specific [BDT^{+}13, MHH^{+}14, VMAVGCM13]. specification [YBD11]. specified [UÇD14]. spectra [Ave10, KA10]. Spectral [CC13b, CAT13, FD13, GK14, JM13, MY13, OA14, SHZ10, TA13, ZW10, AGT12, BFS14, Boy13, Cai13, CCJC13, CLN14, DSN14, FID11, FMP14, FV13, GPZ10, Gün10, yGZ13, HS10c, KH11, LHL10, LP13, MSP12, Mön10, NOS14, OS10, WCW10, qW11, WC11, WM12b, XWW10, Yan13, Yar10, YSQ11, ZML14, ZyG14, ZX10]. spectral-Galerkin [CCJC13, WC11]. Spectral-scaling [CC13b]. spectral/ [MSP12]. spectrally [EY14]. spectroscopic [MM11c]. spectrum [BAK11, Cao13, LHL13, UÇD14]. speech [MVC10]. speed [Law14]. speedily [Koç11]. sphere [CD10, CGK10, LK13, RP11b, WSJ14, YC11]. spheres [LA11]. spherical [Car11, DGG12, FTTT13, GD13, ILMS10, RP11b, ZTC14, ZD10]. spherically [PTP10]. spheroid [BDSG10]. spider [BSS13]. spiral [CHZ10, CC12, Gau10, Lu13b, WM12a]. spirals [CK14b, MSWY12, Nar14]. Spline [DDV^{+}11, PT11b, PC11, WZY^{+}14, ASST10, BGIN13, CW12, CP13b, CW14a, DM10a, DS10a, DCJ13a, EGES14, FO10, ILMS10, IO14a, JKS13, JR14, LMT13, LX12, LC11, LZDW14, LdHA10, MMS13, MD10a, MS12, ÖW14, PT11a, PT14, RG11, SW10b, SWY13, Sla14, VCA14, WZF11, WGZL14, Woź14, XH13, ZW11a]. Spline-based [DDV^{+}11]. spline-type [FO10]. splines [ABIS10, BOR14, BCR13, BwKB^{+}14, BLR10, BDR11, BR14c, CW12, GW14d, GJ13, He12, JŽ11, KP13, KP11a, wKK12, KPR11, KPBN13, KP14c, Kva11, MPS11, MJG14, Nar14, NRS11, OGM14, SW13, WLZ11, XhW11, YL11a, ZWM11, ZL13]. Split [BL13b, WL10a, BT12, CYJ13]. Split-step [WL10a, BT12]. Splitting [BDMR10, AMT11, AKT12, AKT14b, BS13b, BKL14, CHL11, CDN14, DSC09, Dai11, DH14, EO14, EDAP10, HHX14, JK13, KL10b, LTX12, LZ14a, MSE10, PAJ10, PZ13, SGG11, WL14b, ZML14, ZY14b, ZM14]. spot [HT14]. spread [DPV10]. Spreading [SMDMY10, SMMD11]. Spurious [Lak13]. SQCQP [TJL11]. SQP [bJfCwH14, JW10b, MR11]. Square [DR14b, BSV12, CZ13, FS12a, HP13, Hua12, Hua14a, Lan12, SW10a, SD13, TS12, TS14, ZW14]. squared [ZZG14]. squares [AG10b, BSV10, CC11a, GLHCLP11, HPL^{+}14, JKS13, KH11, KNY10, KP10b, Ku13, LV13, LZ12a, LZ13d, LEVG12, MPC10, Nie13b, TR11, Tor14, WWC14, XHmDW14]. ST [FdLW^{+}11]. Stability [AAY10, GXW14, HML14, LF10, LSXZ11, LHL13, MS11c, MK11, SLW14, TLHU10, WWL10, WZL11, WH12a, WH12b, XZH14, YW10, ZFH10, BV10, BB12a, BK12, CZ13, CK14a, CL14a, DR14a, DH13, GRV11, HHKK14, HV12, HGW13, Hu14, HGW12, Hua12, HS12, Hua14a, KH10, KS12a, KL14, LTX12, LC11, LH11a, LF11, Lia14, LYZ11, LFY11, Luo10, MSG11, MS13a, NTNBV14, NPTH09, PJ12, RBK11, TSC10, TS12, TS14, VMMM10, VAR11, WF10, XJ11, ZCH10, mZkPkL11, ZD11, ZW14, GPHAM10]. stabilization [Bay14, KFR13, MF14, TS10, VS14]. Stabilized [Sta11, AM14b, FKNS11, Lam11, LL11a, RYM^{+}14, ZLC11]. Stable [AB10, EN14, OSR13, AGM12, ADL^{+}10b, ADL^{+}10a, BM10, CFF12, CDJP13, DP14b, DS13b, GPHAM10, GPHAS14, Gos11, IJ14, KM11, SASH11, ZFY11]. stage [BS11a, GPHAS14, HVV^{+}11a, HDF^{+}10, Kha10, LCZ11, LEVG12, NBBKV10, NTNBV14, ZGZW10]. stage-structure [LCZ11]. staggered [BBKS10, CC13c, MK11, Spr14]. stagnation [KMS14]. staircase [Dos14]. standard [ÁNA14]. star [Car11]. Stark [Gos11]. starlike [AP14b, Ayd14, KPA14]. start [AG10a]. startup [WYP10]. state [DJM10, DR14a, DBZ10, EX14, HHH^{+}13, Koç11, KL12, LCN12, NPTH09, SS14e, TSC10, TA10, Wan11c, WYO11, XCZ^{+}11]. state-dependent [DBZ10, Wan11c, WYO11]. states [CL14b, DC11, MK10]. static [CF11, DS10b, MKD12]. stationary [AELH10, ABN10, hBW10, CR11b, EGON10, GKT11, SKKS12, Uty10a, ZLC11]. statistical [AB11a, Akt14a, CDPP10, CPP11, WZMC11]. statistics [Ars11, BP11b, BK13, Ery13b, Tur12, YG13, YG14, ZH11]. steady [HHH^{+}13, KM10, LCN12, SS14e, TA10, XCZ^{+}11]. steady-state [HHH^{+}13, LCN12, XCZ^{+}11]. steam [HDF^{+}10]. steel [MVMG12]. Stefan [MV12b, MV14]. Steffensen [CHMT12, EHRV13]. Steiner [LEVG14]. Stenger [HX14a]. stenosed [TTSB14]. Step [ASM13, KA12, ADCC12, AS12b, AH13, APJ10, BT12, CZ13, CCG13b, CT11, CDJP13, CTV11, DEP11, DDV^{+}11, GK11, GSNDB14, GWZ11, HHL12, Ism14, KKR11, KW11, LT11, MMS10, NBBKV10, NTNBV14, PY10a, Sha14, SASH11, TS14, VVM10, WL10a, WWW14, WK14, XSH10, pYXPH10, ZCC^{+}13, ZW12]. step-stress [Ism14]. steplength [XSW12]. Steplengths [WHL10]. stepping [EJ10, GM10, GM13, MSE10]. steps [BDF^{+}10]. stepsize [GPHAS14]. Stiefel [FB12]. Stieltjes [BHJN10, GA12, Mia10]. stiff [BGPPRW14, BL14, FBB14, GPHAM10, GPHAS14, IHRA11, LM14, LT13, MCS12, SM11]. stiffness [fLyHZ10]. Stir [FRB13]. Stirling [GQ14a, GQ14b, XC14]. Stochastic [AA14a, BÜ14, HT14, LZL11b, Pul14, Zho11b, AØ14a, AH14b, AB13, BV10, BT12, BB12c, CZ13, DF12, DFH14, FN10, GXW14, GÖ14a, Hau10, HGW13, HGW12, Hua12, Hua14a, HZR14, HS11b, Isl14, JJL11, JW12, JMV13, Kho11, KLNS11, KB11, KB12, KB13, KLL11b, Lan10, LF10, LF11, LXY13, Li14, LTZ10, LYZ11, LM13b, LNB^{+}12, Luo10, MSG11, MS13a, MJ11, MSZ13, NBZ13, NÖ10, OAHR11, PJ12, PO11, Prz13, RHS11, Sla14, SOL11, TNN10, TW14, TS12, TS14, VTLF11, Van11b, WL10a, Wei11, WZ11a, Xu11b, XYW14, ZAT11, ZY10, mZkPkL11, ZRW11, ZW14]. Stokes [Dai11, AU14a, AB10, AM14b, BY12, BHMW13, CM12, CHX11, CW14b, CAT13, CDS10, DSC09, DD11, FLHL14, FKNS11, HHH^{+}13, HL10a, HdJL^{+}13, KH11, Lam14, LTX12, LZL11a, LL11a, Naf14, PSS13, Spr14, ZH10b, ZY14a, ZLC11, dAGR10, dFGAN11]. Stokes/Darcy [CM12]. Stokeslets [BG13]. Stolarsky [HH13, HH14]. stone [LMM^{+}10a]. storage [CFMR12, HTM14]. Storjohann [Har11]. strain [GVF^{+}13, JŽ11, KAS13]. Strang [EO14]. strangeness [LM14]. strangeness-free [LM14]. strategies [AAI14, ADGP11, KP14b, KA12]. strategy [AMMS14, CP13a, DSZ11, GL10, GCC14, GWX10, LT11, MR11, PY10a, TYCL14, Wih11, Zha14c, ZTL13, ZLW14b, ZwGhX14]. Stratonovich [KB11, KB12]. streamline [DV13, MS11a]. Streamwise [BKMG11]. stress [AHAH11, Ism14, PNCC10, YL11c]. stresses [BDT^{+}13]. stretch [ZGZB13]. stretch-forming [ZGZB13]. strict [Ant11, JK10b, KPW10]. string [BG10a]. stringed [Fie13]. strip [JW10a, ZFY11]. Strong [BT12, CB13, JK10b, KB13, LXY13, MS13a, NTNBV14, QCK10, BHJN10, HHKK14, IG12, LLW14, RR10]. Strong-stability-preserving [NTNBV14]. Strongly [GPHAS14, ABN10, AK11, GPHAM10]. Structural [ZGZB13, BHBP10, GUO14, ILEE13, NTRNXB10, PNCC10, YL14]. Structure [BBP14, AP10, BB12b, CPP13, CL11d, CDQ14, DBHV10, DFG^{+}10, GG14, HLLW13, JMS10a, JWL13, KHFC11, xLwCL10, LCZ11, Nai11, PP10b, WH10, YZ11a, ZGZW10]. structure-oriented [ZGZW10]. structure-preserving [GG14, HLLW13, JWL13, KHFC11, xLwCL10]. Structured [BN14, Win14, AALMLM14, AR10a, EZ10, GS12, KNY10, LE10, Li10, LZ11, LSW11, MU14, TOF14, UM14, VVM10, WH13, XLCC10, ZZS11]. structures [Bab10, DA14, FHPH^{+}13, GJCL11, LM12a, SKKS12]. studies [HLLW13, RTL14]. Study [Mor10b, XSV12, XSV13, dAGR10, AHLJ14, AD11b, AC14c, BFS14, CI13, Col13, DI12, HwKB14, JK13, JR10a, KL12, ORP14, SBCD10, TTSB14, WZMC11, ZS11]. Studying [DGI^{+}10]. Sturm [BAK11, Cha10a, CO11b, CSZ11b, GCH13, JTV14, Liu12, MRD13, RB13, Reu10, ZM10, ZZ12]. style [HL14]. sub [CYZ^{+}11]. sub-surfaces [CYZ^{+}11]. subdiffusion [LZDW14]. subdivision [BCR11, CC13a, CCS10, CR11b, CD11, CM10c, DGW10, Flo11, KP13, LLL^{+}11, LZ13b, MS11b, MA11]. subdomain [Sun14]. subgradient [BSBA14]. subgrid [ZH10b]. subgrid-scale [ZH10b]. subgridding [WC14]. subject [CDPP10, LD10, YBD11]. subjected [AHM13]. sublinear [HZ11]. submatrix [YGHH14]. subproblems [LTX12]. Subspace [GCY13, HR10, AMRS12, CCS11, ENS14, IHRA11, Liu14a, NL10, PSS13, QY10, WZ14]. Subspace-restricted [HR10]. subspaces [Gos11]. substencils [Ger12]. substitution [XL12]. substitutions [LM12b]. subsystems [BBS14]. Successive [LsGHZ10, BY12, BCC12]. sufficient [Sza10]. Sugiura [ISN10]. suitable [ZZG14]. suite [MKS12, NACLR12]. sum [CID10, DS10a, GK11, HHX14, JK10a, Vel11, XHmDW14, YL11b, Zha14a, Zho11a]. sum-of-ratios [JK10a]. Summation [HZ13]. Summation-by-parts [HZ13]. sums [Ery13b, EGST11, KH13]. sup [CXZ14]. Super [Kho11, BM10, GS12]. super-diffusion [BM10]. Super-efficiency [Kho11]. Supercloseness [DLP13, FP13b]. supercompact [AM10c]. superconductors [JS10]. Superconvergence [HYY14, HX14b, HK11, LZY10, PN11, BG14a, Cao14, LWY10, WDZ10, ZL10, ZC10c]. Superconvergent [AST14, WG10b, GW10]. Superfast [RST14]. superlinear [AK14d, JW10b, LPGX10]. superlinearly [bJfCwH14]. Superquadratic [AMMS10, AMMS14]. Supersingular [GM14b, LZY10, Yan13]. supersymmetric [LWZ^{+}14]. supervised [SH13a]. Supplement [KB11]. support [AÖY10, MM12a, QTS14, Wu10]. supported [CCCM13, CCH10, CBGVPP11]. Supraconvergence [FP13b]. supremum [VAZS13]. sure [Lan12, MSG11, Rie13]. Surface [CHMD14, eMA14, AM11, BKMG11, BHMW13, CYZ^{+}11, CDK10, CM10c, FGEM10, HLLW13, Jia10, KS14d, LBM12, MT10b, Mor10a, NRP13, PY10a, SZ10, vGALL10]. surfaces [ALM11, BFGP10, CYZ^{+}11, DM10a, DP10a, DP11, FS12b, HCQ14, JR10b, KPBN13, LLL^{+}11, MP10a, SCFR14, WL10d]. surrogate [WCH11, Zim13]. surrogate-assisted [WCH11]. Survey [RVPOR^{+}11, CM13, DL12, FS14]. survival [YG14]. suspensions [KG12]. SVD [VVV10]. swap [YYL^{+}10]. swaps [SH13b]. swarm [AEWMES11, AZ11, ARF14, CRF14, RMF11, Wu10]. Sweeping [PV10]. swirl [ZZC11]. switched [Büy12, ZSG^{+}14]. switching [AA14a, APW14b, CLMR14, ESC14a, HS10b, JWY11, LXY13, LZ13e, MD10c, YY10, Zhu14]. Sylvester [KT13, WH10, WH13, ZM14]. Symbol [DDDSC14]. symbolic [BL13a, MRT13a, SC10a]. symbolic-numerical [BL13a]. symbols [CR11b]. Symmetric [Han11, KP11a, SH12, WSJ14, AMMS12, AS11, ASK11, AS12b, BR11a, BOS12, BCC^{+}13, CL10a, CCCM13, CZL14, CG11b, CL13a, CH14, DH13, DLWZ14, EO14, EJ10, FV14, LL10, MJ10a, McN10, MML11, QY10, RSV12, RT12, SC10c, Sid14, SL14b, WLD14, WH12a, YY11, YSZ11, YGHH14, YL14]. symmetric/antisymmetric [YSZ11]. symmetrically [TTSB14]. symmetries [Alc14]. Symmetry [LL14b, SOL11, SKAM11, BSS13, HL13a, Jia10, Jia11, fLyHZ10, PM11a, RRMD13, ZYC13]. symmetry-preserving [RRMD13]. Symplectic [FG14, KHFC11, CSZ11a]. symplecticity [BB12c]. synchronization [HL10c, WC10]. synchronous [DH11]. synthetic [PRT10]. system [AD10b, BAdS^{+}10, BAdS^{+}12, BL14, BJ13, Büy12, CXL10, CHRZ13, CW14b, CR14, CL14b, Col13, CyY10, CwYyY11, CyYwY11, DW10, EC13, FDH10, FV10, GRV11, HZ11, HwKB14, ILEE13, JJL11, KFR13, KS14b, LCZ11, LYP10, Liu14a, Liu14b, Man11, MM11a, N’G10, NF14, NMS10, PL11, PY10b, RK11, RTL14, SCRB^{+}14, Sez10, SJA12, TNN10, jTW14, TW14, WL11a, Wan11c, WA14, WK10, Wu10, XHZ12, XZ14, YY11, YWW10, YSW12, YZ10b, ZH10c, mZkPkL11, ZYW14, ZTC14, ZD10, ZGZW10, ZLL10b, ZZ11b, dAGR10]. systems [AS10a, AA14a, AGJ13, AAM14, AAI14, AMV13, AF10, AS11, ASK11, AG10b, AIA10, AC14b, AGT13, BDD^{+}10, BS13a, BGPPRW14, BFB14, BGGG^{+}11, BGS12, BHJ13, BFG^{+}14, BDMR10, BMN10, Bra12, BNRZ12, BCMR12, BL11, BGKVW10, BHBP10, CA12, CLS14, CZ12a, CZL14, CHMD14, CJ10, CXH09, CV10a, CGL10, CHMT10a, CTV13, Dea11, DR14a, DAW11, Dha14, DJ14, DSY^{+}11, Ery14, EX14, EJ10, FO10, FHPH^{+}13, FJP14, GW14a, GC11, GI10, GPHAM10, GG10, GSGN11b, GSGN11a, GT09, Gus14, HD11, HLC11, HK13a, HM11, Hu14, Hua12, Hua14b, Hua14a, ILEE13, KA12, KA14, KN10b, LWW11, Lam11, xLwCL10, LLP12, LWZ13, LT13, LHL11, MF14, McN10, MZL14, Miy10b, MT14b, NOS14, NLL13, PL10c, PZ13, QC14, RBK11, RYM^{+}14, RP11a, Rum13a, Rum13b, SXZ11, SSF^{+}11, TS12, TLNE10]. systems [Wan10, WZ10a, WSWD11, WZ14, WL14b, XJ11, YT12, YD12, YL14, ZSG^{+}14, ZCH10, ZDZ11, ZJ12, ZHCW12, ZG14]. Szász [Mah11]. Szeged [YAKA11]. Szego [CGK10].

T1 [RPS^{+}11]. T2 [RPS^{+}11]. T2-weighted [RPS^{+}11]. tabular [PGGCGF14]. tag [SPLHCB14]. Tail [KKK12, KH13]. tails [NR14a]. tangencies [FTTT13]. tangential [GNN11, Mor10a]. tangents [LLL^{+}11]. Target [LJE^{+}10, LL10, PP14]. Taylor [Dom11, HNK10, MS11b, MJ11, NBBKV10, Prz10, SL14a, WW14, ZAT11]. Taylor-series [HNK10]. tearing [NMS10]. tearing-based [NMS10]. technique [ÁECS12, CZWD14, ÇU14, CL13b, CHMT13, DBHV10, DAE13, EGON10, FRB13, GSM11, LY13b, LFWL14, MMS13, RW10, RRMD13, Wan12a]. techniques [AELH10, AC14c, BGL13, Cao14, CWQ11, DDE14, GPVL10, LZ10, LMM^{+}10a, LMM^{+}10b, Més12, PR12a, SH13a, SLK14, WB10, ZF14, ZTL13]. technology [HD13]. telegraph [FRXY13]. telephone [WL11a]. Tempered [BM10, KM11]. temporal [AHM13, WZMC11]. temporally [YESD14]. temporally-evolving [YESD14]. Tension [MPS11, DM10a, KS14d, Kva11]. Tensor [Tyr10, BwKB^{+}14, BHL11, CL13a, GJ13, KS14b, Kho10, LZI10, MD13, MJG14, Spe11]. tensor-product [BwKB^{+}14, GJ13, MJG14, Tyr10]. tensors [LWZ^{+}14, ZZP13]. term [And13, Cas10, DST10, EGBS12, MY11, PT11b, SD10, WLZ14]. terminals [HTM14]. terms [BV10, GQ14a, HLM10, HMS11, Rah14, SM11, XSH10]. terrain [CGRS11, ZM13]. Test [MCS12, PS10, FPG12, GÖ14b, Ism14, PM11a]. testing [ZDP^{+}14]. tests [AÖ11, VSB12]. tetrahedra [SH12]. tetrahedral [JKS10, NRS11]. textile [ZS11]. TFETI [DKH^{+}10]. th [GKLZ12, JGH^{+}13, SH10, Zie14]. their [BSB14, BFS14, DH14, GK11, HV12, LW14, Ma10, MBS10, QLL11, UÇD14, VAR11]. Theodorus [Gau10]. Theorem [GGH^{+}14, KS14a, AAM14, FJ10, HCWsO13, Khr10, Lan10, NK10, SW12b, yS05]. theorems [Akt14a, CB13, ĆRR^{+}10, JK10b, PD13, PJ12]. Theoretical [Wik10, OCU11]. theories [ZF14]. Theory [BDJ^{+}11, GR13, AHK14b, Bak10, BS13c, BLTW10, CFGE10, GG14, KWW11, KS14d, Kre95, Kre11, LMW^{+}11, NSZ10, WLL^{+}11, YFZ14, YLD13, ZWZ11]. therapy [GPVL10]. thermal [VdKPB10]. Thermally [ZY11b]. thermo [NRS10]. thermo-mechanical [NRS10]. thermoacoustic [HHP^{+}10]. thermoelasticity [SCDA10]. thermoviscoelastic [CF11]. theta [CI13, Hua14a, ZW14]. thickness [LLG14]. thin [BOR14, CHH10, CC11c, KMD12, TKW12, WL11a]. thin-sheet [TKW12]. Third [ABG11, ABCP14, HH13, AM11, BDD^{+}10, Büy12, CN11, CGL10, Fuk12, Fuk13a, GZL14, LGZ10, WX12]. Third-order [ABG11, ABCP14, HH13, GZL14, LGZ10, WX12]. thixoforming [KP10a]. Thomas [Boy13, KHKM14, PDP13]. Three [BDSG10, CTV11, GH10a, NMBG14, AÖÖ14, And13, Bra14b, CG10b, Cas10, CGS13, CGK10, DST10, DD11, EX14, Fat11a, Fat11b, Fat13, FND10, GMRK14, HS10a, HdJL^{+}13, HHLW14, LS11, MTJ11, PW12, PL10c, Sch14c, SLK14, TP10, WZ10c, WLL^{+}11, pYXPH10, ZTC14]. Three-dimensional [BDSG10, NMBG14, CGK10, Fat11a, Fat11b, Fat13, GMRK14, HdJL^{+}13, Sch14c, SLK14, WZ10c, WLL^{+}11]. three-field [FND10]. three-phase [PW12]. three-point [LS11, MTJ11, TP10]. three-state [EX14]. Three-step [CTV11]. three-term [And13, DST10]. threshold [GL10, KBKR14, Reb12, Zha14c]. Tight [CCH10, YSZ11]. Tikhonov [BFLP12, BEJ14, CHL14, Chu10, GN14, HPR12, LV13, Wan12b, XW12]. Tikhonov-type [Chu10, XW12]. Time [DC10, DCZ10, JGK14, MSE10, Mön10, SHZ12, AK14a, AH14b, BY12, Bak10, BCHL11, BG11, hBW10, BÜ14, BSB14, BDMR10, BSJ14, BHJ11, CFMR12, CR12, CK13, CL11c, CHH10, CGS11, CG13, DJM10, DS14, DD11, DMMP10, EGES14, EGON10, EN14, Ery13a, EJ10, FL10a, FGO11, GGAVRC14, GZL14, GPHAPR14, GM10, GM13, GS14, GH10b, GP10, HWG10, HL10a, HD13, HL11, JM11, JW10a, JR10a, KKR11, KR14, KMLT10, LD10, LTX12, LPDN13, xLwCL10, LH11a, Li12, LwCLW13, LSW11, LB14, LS10, MJ11, MSZ13, MS13b, NGAZ11, NF14, NTNBV14, PVM^{+}14, RBK11, RG12, RRMD13, Sch14d, SI11, SYK14, SJA12, SOL11, TR11, VVD^{+}14, WZF11, Wan13, WW12, WHL13, XJ11, YY11, YLS12, YTFP11, ZCH10, ZM10, Zha11b, ZZ13, ZW10, dAGR10, dO10]. time-delay [xLwCL10]. time-delayed [RRMD13]. time-dependent [BY12, BSJ14, CR12, CGS11, DS14, EGES14, GS14, HWG10, HL10a, HL11, JW10a, KR14, LD10, LPDN13, LB14, MJ11, SYK14, SOL11, TR11, YY11, dAGR10]. time-discretization [NTNBV14]. time-domain [BG11, CHH10, FL10a]. time-fractional [JM11]. Time-harmonic [Mön10, DMMP10, WHL13]. time-scale [FGO11]. time-space [MSZ13]. time-step-based [KKR11]. time-stepping [GM10, GM13]. time-varying [CK13, CL11c, RBK11, XJ11, ZCH10]. times [MB11, ZY11a]. timesteps [DH14]. Timoshenko [HKO11]. tissue [LP10a]. titanium [KS12b]. TM [KS14c]. TM-waves [KS14c]. Toeplitz [LHL11, BGM10, CJ10, CV10a, Cin12, Cin14, DDDSC14, ENSCS12, GT09, McN10, ZZ11b]. together [ZFH10]. tomography [BCM10, GKL^{+}12a, GCH14, HHT12, WB10]. Toolkit [SPC14]. tools [DDG^{+}13b, HSH12]. topography [YC11]. Topological [AP10, DA14, LQSL11]. topology [DL14a, PNCC10, WB10, WUB10, ZGZB13]. torus [Kim12]. torus-torus [Kim12]. Torvik [CK14a]. total [Bir11, LV13, LL14a, NGAZ11, SWT12]. trace [APS14, AA14b]. tracing [VdKPB10]. tracking [BW14, DBHV10, FdOdS12]. tract [PKH14]. traction [YC10]. traffic [BJ13, GGAVRC14]. training [MVC10]. trajectory [ESM13b]. transaction [AV12]. transcendence [Zho11a]. transcendental [Fuk13d]. transcription [ESM13b]. transduction [PSKA14]. transfer [ÁECS12, KP11b, KD10, LBM12, LMW^{+}11, SDK10, ZD10]. Transfinite [BF10]. Transform [BC10, MPS10, ÇU14, CFH14, HS11a, HZR14, KKPV11, KL11, MT10a, NS10, NA11, PMMS14, SW10a, Tan14b, ZD14]. transformable [CS12]. Transformation [ZF14, AM10a, CDS10, HV14, SGG11]. transformations [CCCM13, Che13a, Fuk11, Fuk12, Fuk13a, GHM10, RVPOR^{+}11, SA14]. transformed [NR14b]. transforms [CA14, FS10, Mia10, XXH14]. transient [AHM13, BM10, Bay14, CLS12, CLP10, KAE^{+}11, RVDM10, SM14a, SDK10]. transition [HL13a]. Transmission [CHH10, AK14c, HLSX11, JV10, LGR11, Mau10, NV11, RL10]. transonic [APO14]. Transparent [BDL10, SZ10]. transport [BCM10, CGS12, GG14, KSRM12, KG12, MHL11, PC12b, RP10, RW10, SB10, SRP^{+}13, Sun14, SS14e, SKKS12, Vel10, VS14, YLD13, ZS11]. transpose [SRWZ14, Yeu14]. transposition [DP10b]. transverse [SZ10]. Tranter [FMP14]. trapezoidal [Car11]. travel [MB11]. Traveling [LCZ11, LWZ13, Sie11, WZ10a, ÁD14, ÁD15, MB11, ZLC10]. travelling [ÁD11a]. travelling-wave [ÁD11a]. treating [MVMG12]. treatment [AVGCMVM10, CK13, Eba11, Fat11b, FFM^{+}11, FYM10, Odi11, PNCC10, Pas11]. treatments [PGH13]. Tree [BBD^{+}11, KS14b, KPsQ11, LEVG14, MY13, XCY12, YFZ14, YY10]. trees [BDV10, BDL11]. Trefftz [FLHL14, KSTW14, LLCL11, RW14]. Trends [GKB10]. Tresca [Gwi13]. Tri [PZJ12, Liu14b]. Tri-diagonal [PZJ12]. tri-vector [Liu14b]. trial [LY13a, RW14]. triangle [BSV10, RSV12]. Triangular [CYZ^{+}11, ALM11, AM11, CZ14, GGLR10, HK14, JJE13, KA10, LQSL11, NTRNXB10, RSGL14, WL10d, WL14a, YC10]. triangulated [FS12b]. triangulation [BGIN10]. triangulations [BFGP10, DM10a, DY13, SCFR14]. triboelectrostatic [MAK^{+}14]. Tricomi [LW14]. Tridiagonal [FWBSBGA11, LL13a, BCC^{+}13, KMG12, MVV14, MRT13a, WH12a, WH12b]. Trigonometric [HP10, SS14c, EGBS12, GW14d, JR10b, JR14, PHS14, TLHU10, Wu14, YL11a]. trinomial [YY10]. trip [LQYL11]. triples [MMX10, WK14]. triplets [NY14]. Troesch [Cha10b]. Trotter [Pop10]. truncated [GA12]. truncation [LL10, PPG14, ZFM14]. trust [CWQ11, JLZZ10, ZCX14]. trust-region [JLZZ10]. Tseng [ZYG10]. tsunami [MKK14]. tumor [HHH^{+}13]. Turbine [NMBG14, BL10b, HDF^{+}10]. turbulence [BFB14, DH13]. turbulent [ORP14, WCZ11, YESD14]. Turing [SBL12]. Turkey [ÇDE14]. Turkish [MKK14]. turning [GQL14]. TVD [GZ10]. twelve [FTTT13]. twisted [KMG12]. Two [APO14, AB11b, BKGMA10, BK14a, BS11a, BL11, CSL^{+}11, CT11, DH11, Gas14a, GNN11, LW10a, LEVG12, LYP13, Mol10, PS14b, XN14, ABS11a, AAT11, AF10, AE11a, AM14a, AH13, AAD13, AuIK14, BDD^{+}10, BMMR11, BW14, BOR14, BSJ14, BDF^{+}10, CS12, CZ13, CCG13b, CC11a, CAA10, CPY10, CL10b, CSZ11a, CL14b, CK14b, CDJP13, CyY10, Cui14a, CID10, DEP11, De 14, DA10, DQST14, DKH^{+}10, Dou12, DGC^{+}13, Eba11, EASS13, FPP12, FRXY13, GW14c, GSK11, GK11, GSNDB14, GZ14a, GZ14b, HdJL^{+}13, HTH10, Hua14a, JK10b, JZ10, JW10a, JV10, JKV11, Jun13, KSK11, KHFC11, KW11, LZ12a, LZ13e, Lu14, LC10, MSWY12, MD10b, MV14, MA11, MS13c, NLO13, NRP13, PY10a, PR12a, RW10, RG11, SR11, Sch11, Sha14, SASH11, Sun14, SS14e, SKN11, TS14, WYP10, WG10b]. two [WZLP11, Wan13, WK14, WL11b, XLCC10, Xu11b, YZ10a, YSZ11, YL11b, ZML14, ZYL10, ZYW14, ZY14a, ZGZW10]. two-by-two [ZYW14]. Two-dimensional [APO14, BL11, Gas14a, ABS11a, AE11a, AM14a, AAD13, AuIK14, BMMR11, BSJ14, CSZ11a, CK14b, CyY10, Cui14a, DA10, DKH^{+}10, Dou12, HTH10, JW10a, JV10, KHFC11, NLO13, NRP13, PR12a, WG10b]. two-fluid [DGC^{+}13]. two-grid [CL10b, JKV11, Wan13]. Two-level [XN14, ZY14a]. two-mesh [RW10]. two-objective [KSK11]. two-parameter [FRXY13, GSK11]. two-phase [BW14, DQST14, EASS13, HdJL^{+}13, MV14, MS13c, Sun14, SS14e]. two-point [CAA10, Eba11, LZ12a, LC10, MSWY12, RG11]. two-scale [MA11]. two-sided [SKN11, YZ10a, ZYL10]. Two-stage [LEVG12, ZGZW10]. two-step [AH13, CZ13, CCG13b, CDJP13, DEP11, GSNDB14, KW11, PY10a, Sha14, SASH11, TS14, WK14]. Two-step-by-two-step [CT11]. two-types [CL14b]. Two-variable [LW10a, AAT11, GK11]. Two-view [CSL^{+}11]. type [Akt14a, AT10, ÁD14, ÁD15, AR10b, ANS11, AD10b, AR11, AEG^{+}11, AK14b, Bak10, BFGP10, BGIN10, BLO11, BTSA14, Cai13, CAA10, Che13b, Chu10, ĆRR^{+}10, CT11, CHMT12, Dea11, DAE13, Dom11, DBZ10, DM10b, DAH11, EY14, ELA10, EL11, EALA11, EHR11b, FO10, FPPX10, FGPP13, FG14, FN10, GST12, GPHAPR14, Gro13, KKR11, Lak13, Lan12, LWH^{+}11, LW10a, LF11, LL14a, LTFL10, LH11b, Ma10, MV12b, NRS11, OPPV10, OB14, PM13, PJ12, PGGCGF12, QLL11, RCM^{+}12, RA10, Sch14d, SAB14, SZS11, SZ11, TL13, WHL10, Wan11a, WSWD11, WZL11, XW12, YKL10, Yan10a, ZH10c, ZH11, Zho13, ZLL10b, AHAH11, Ism14, WSS14, JS10]. type- [BGIN10]. type-2 [QLL11]. type-6 [NRS11]. Type-I [WSS14]. Type-II [AHAH11, Ism14, JS10]. types [CL14b, EK11, GW14c, KRS11, LL14b, LYP13].

Uhlenbeck [Sch14d, WZ11b]. ultra [FGEM10, HCM14]. ultra-refraction [FGEM10]. ultraconvergence [LZY10]. ultrahigh [WWW14]. unbounded [CIPS14, SWY14, XW12, ZAWA13]. uncertain [AÖY10, DCGS13, TLNE10]. Uncertainty [DAW11, CV10b, HLF14, KSRM12, fLxX12, ÖW14]. unconstrained [And10, And13, CWQ11, DW11, GS12, JLZZ10, KK13, LP13, NY12, SS14a, WLWH11, XSW12]. Undamped [BS11b, YD12]. underdetermined [Miy10b]. underlying [LZ13e]. underwater [MKK14]. undetermined [BBD13]. undisturbed [Tez14]. unfitted [NS14]. uni [vGALL10]. uni-directional [vGALL10]. unified [ABG11, FM14, KL11, NNB13, SAY13, Yan13]. Uniform [Bog11, GÖ14b, RP11b, ZWM11, AV10a, AM10c, Ars11, BCR11, BCR13, Bog14b, CXZ14, DFH14, FJ14, Jia11, KP13, MOG10, XH13, YZ10b]. uniformly [CG13, RK11]. Unifying [TA10, Bir11]. unilateral [Gwi13, Kok11]. Uniqueness [JLZ14, BB11c, KL12, Sha14, Zha10a]. unit [BC12, CGK10, CBGVPP10, CBGVPP11, GHM10, HV14, WW11a]. unitary [HV11]. unity [BDyS14, HTH10]. universum [QTS14]. unknown [KP14a, LD10, VS13a]. unloaded [BDT^{+}13]. unreliable [WK10]. unsaturated [PDE^{+}13]. unstable [AB10, Hu14]. unsteady [BG10b, HDF^{+}10, MDVB13, PKH14, SHZ12, TY11, YK11, YK14, ZY14a]. unstructured [BFHK12, Dou12, EZ10, HZ14, LE10, NCCF^{+}10]. untangling [GKU14]. until [ZZ13]. update [AAI14, VL13]. updates [CC13b, Gal13, KO13]. Updating [SSF^{+}11, Che14b, KSRM12, Yua12, YL14]. upon [Vel10]. upper [CG10a, DA14, Gün10, KA10, WZS10, YG14, Zha14b]. Upscaling [BFDP13]. upwind [BGP11, LE10]. Urysohn [DSN14]. usage [BSB14]. Use [VdKPB10, GPVL10, LD10, LFY11]. Using [BGS12, CDS10, EK12, Gat14, MA14a, VVV10, WCH11, AHR14, AO14b, AEGB10, AE11a, AM14a, AR10a, AC14a, AIA10, AY13, AHK14b, Asl10, AAD13, Ass09, AC14c, AGT13, AuI13, BMMR11, BAU13, BFDP13, Beh14, BC10, BL13a, BAdS^{+}10, BAdS^{+}12, CGCG^{+}13, CHL11, CM10c, CRS12, DDM10, DDV^{+}11, DGC^{+}13, ESM13b, ESM13a, FND10, GGS12, GUBS13, GA12, GST12, GMRK14, GSNA13, GS12, Har11, HdJL^{+}13, HX13, HZR14, IO14b, KXH13, KG13, KPsQ11, KA14, KP14c, KD10, Kud10, LSD11, Lam14, fLyHZ10, LLCL11, LLK14, LM12b, LMM^{+}10a, LJE^{+}10, LJMG11, MMF10, MMP11, MJ10b, MVC10, MSP12, MM11a, MHH^{+}14, NLO13, NTRNXB10, NMBG14, NÖ10, PY10a, PM13, RW10, RMTO11, RST14, Rum13b, SSW14, SW12a, SL14a, SYK14, SC10d, SP11, SKN11, Tez14, TOF14, Tor14, WB10, WTM12]. using [WC11, WS13b, WL10d, Wri10, Wu10, WM12b, XN14, XCY12, YT12, YÖVAYOW14, YL14, ZM13, ZGZB13]. utilities [LRZ14]. Uzawa [BSS14, Kok11, ZYW14, ZLW14a].

vacation [CLS12, GI10, YWW10, YTFP11, ZH10a]. vacations [ZH10a]. vaccination [SI11]. validation [BKTY14, LA10, LdHA10, RYM^{+}14]. valuation [FS10, JM12, LMG13, Mar11, MT10c, Sch14d]. Value [AGM12, DH10, MCS12, AU14b, BS11a, Bar11, BK10, BF10, CXL10, CLS12, Che10, CZ12b, DHR12, DAE13, Eba11, ESM13a, EGH12, GSK11, GC11, Gen12, Gus14, HS10a, HR10, HVV11b, HHZ11, IO14a, JTV14, KP14a, KPKK12, KNR12, KKN11, KP11b, KLM14, LS11, LW13, LZ12a, LWY10, LL13b, LC10, MY11, MMX10, MTJ11, Mui10, PT12, PP10a, QLL11, RG11, RU13, Rua13, RR10, SZS11, SP11, Tar10, VS13b, Wan10, WX10, WZ10b, qW11, Wan12a, WWG10, WZ10d, ZFG10, Zha11a]. value-at-risk [GSK11]. valued [ADRM12, DW10, FW10, FKL10, Liu11c, MR12, Sto11a, Yar10, ZW11b, ZLG12]. values [AMMS10, AMMS14, CL13a, CS14, LHLL10, UÇD14, WS13a]. valve [BSBA14]. valve-point [BSBA14]. vapor [BKN^{+}13, ZS11]. Variable [UM14, AAT11, AIA10, CLS14, CLB11, DXL10, DJ14, DI12, GK11, HHM12, KK13, LW10a, LM12b, MMS10, Wei11, XYW14, XL12, YSW12]. variable-order [CLB11]. variables [BP11b, BDF^{+}10, CT10, Ery13b, FPPX10, FPP12, Nad11, OSR13, QLL11, ZH11]. Variance [SL12, CW14c, GPVL10, Isl14, MX10]. variant [AAI14, CLS12, DSZ11, TT11, Yeu14]. variants [AS10a, ADCC12, CT10]. variate [MX10]. variates [KM11]. variation [Bir11, LL14a, MMF10, SWT12]. Variational [AK14a, GKU14, JMV13, AU14b, BPGP11, BLKZ10, CB13, Cha10b, CC11b, DW10, FW10, FS14, FLHS^{+}13, GS10, HHL12, HCM14, KV12, KPBN13, KP14c, LTW10, LZ11, LY13b, pLjH10, NASNY11, NNB13, PY10b, SH10, SGO14, jTW14, WG10a, ZZSW11, ZLL10a]. variational-like [DW10]. variations [BB14a, DGI^{+}10]. varieties [ZW11a]. varying [CK13, CL11c, JGK14, RBK11, XJ11, ZCH10]. VBARMS [CLS14]. Vector [CGCG^{+}13, dO10, AWF11, Ant11, BCHL11, CGS13, Chu10, DD14b, DD14a, EK12, EDAP10, KP14c, Liu14b, MR12, QTS14, SST12, Wu10, pWqSmZ14, ZLG12]. vector-valued [MR12]. vectorial [MSE10]. vectors [Zho12]. vehicles [HTM14]. velocity [BBKS10, CAT13, MDVB13]. Verhulst [Nie13b]. Verification [WDG^{+}12, MN10]. version [CZZ14, LK13, SW12b]. versions [Śmi11]. versus [LP10a]. vertex [Kuz10]. vertex-based [Kuz10]. Vertical [NMBG14, BB12b, PDE^{+}13, PVM^{+}14, SZ10]. very [KN10a, MM12a, MWWY14, Oou13, Pas11]. vesicle [DGC^{+}13]. vessels [BDT^{+}13]. VHF [ACF^{+}14]. via [Akt14a, AG10b, Bel13, BBD13, BHL11, CCCM13, Che13a, Eba11, Ery14, ENSCS12, Fat11b, FV14, HNPX11, JWW14, KP13, KZ13, LHC14, LYD13, PS14c, PPG14, RP11a, RP11b, SCDA10, XHZ12, Yan11, ZWG11]. vibration [NTRNXB10]. vibrations [HDL^{+}13]. ViennaMesh [RWRS14]. ViennaX [WRS14]. view [BCM10, CSL^{+}11]. VIM [Odi11]. Virasoro [HV11]. viscoelastic [FdOS12, dAGR10]. viscoelasticity [FM11a]. viscoplastic [GVF^{+}13]. viscosity [HK13a]. viscous [DV13, HJ11]. visibility [YWY^{+}14]. vivo [BDT^{+}13]. Vlasov [AC14d, CR14]. VMS [JKS10]. vocal [PKH14]. void [NS14]. Voigt [PM10]. volatility [BÜ14, DF12, DFH14, FPP13, GÖ14a, HZR14, HS11b, IO14b, MD10c, PO11, Sch14d, SOL11, VTLF11]. Volterra [TSC10, AM14a, AuI13, BMMR11, BS10a, Bar13, BB11c, BGGG^{+}11, BGL13, CCDS14, CZ12b, DMR13, DCJ13b, FID11, IK11, IJMV10, JM13, LZ13a, Luo10, Ma10, MR13, NLO13, NPTH09, PGH13, RD10, RST14, WWC14, WW14, WZ11a, Wu14, Xu11a, YZ11b, YSW12]. Volterra-integro [JM13]. Volume [Ano10a, Ano11a, Ano11b, CDK10, ABS11b, AD11b, BS10b, BGP11, CHL11, CZZ14, CL10b, CL11b, CG14, CXZ14, CMR11a, DGG12, EZ10, Fat13, GW10, HML14, LY13a, LMY11, LMYJ12, LXSC11, MHL11, MSZ13, MJ14, NCCF^{+}10, PSRV10, SLK14, SGO14, WG10b, YC11, YL11c, ZZS11, ZTC14, ZY14a]. volume-preserving [SGO14]. volume/ finite [DGG12]. volumetric [GCY13]. Vortex [Pro12, BKMG11]. vorticity [CAT13]. Vries [HX14b, RTL14].

w.r.t [KS14b]. Wake [MSP12]. walk [Hui10, SRP^{+}13]. Wallis [PS14c]. walls [GTS14]. Walsh [UÇD14]. wastewater [AVGCMVM10, VMAVGCM13]. water [AVMVMV10, AJK12, BS10b, BG10b, CBB13, DV13, Ehr11a, KS12b, YC11]. water-gas [AJK12]. Wave [DDV^{+}11, HP10, APO14, AV10a, ADG11, ÁD11a, ÁD14, ÁD15, AV10b, Bab10, BG11, BFB14, BBD^{+}11, BUG^{+}13, BFHK12, BB14b, Cai10b, CI13, CC13c, CR11a, DLG14, DC10, DCZ10, GG13a, GM13, HLM10, HWG10, Hau10, HLLW13, LEKT10, LVD^{+}14, lLzS11, Lia14, LWZ13, MKK14, MM11b, MS13b, MK11, OCU11, OS10, PP10a, Sch14c, Sie11, SC10d, Sto11b, SM10, SKN11, TT14a, UBSG12, VVD^{+}14, WZ10a, WZF11, ZTC11, ZTC14]. Waveform [JD13, LJ11, BY11, BS11a, BSS14, LJ12]. wavefront [KNR12]. wavefronts [LCZ11]. wavefunction [ZMZ13]. wavefunctions [TT14a]. waveguide [BDL10, PAD10]. waveguides [Gia14, Pag10]. wavelength [Ehr11a, HHP^{+}10]. Wavelet [Jia11, AHK14a, AuIK14, CG11a, CH13, Gos11, HK13a, PVP13, PMMS14, Wan11b, Wu10, YSZ11]. wavelets [AY13, AuI13, CCH10, Han11, HP13, uIAAF14]. waves [AH10a, ACF^{+}14, CLP10, Ehr11a, FGEM10, GPKV10, GST12, HWG10, HdJL^{+}13, KS14c, MR10, RG12, SC10d, vGALL10, LBCL10]. way [BM12, MOG10]. Weak [Hau10, KB12, YWY^{+}14, AH12a, CB13, FN10, GW14b, HCM14, KB11, OPPV10, WY13, WMGM14]. Weakening [BG13]. Weaker [AK14b]. weakly [ASST10, BH11, Bar13, BSV10, BOS12, CC14b, DZZ14, LSD11, NRS10, OMS10, PGH13, PZ13, RD10, Wu14, ZZ11b]. webs [BSS13]. Weibull [AHAH11, ADA13, GSK11, Ism14, WSS14]. weight [EL11, GM14b, MVC10, Not14, Sid14]. Weighted [DXL10, AEGB10, Che14b, CDV10, CCZ12, EGST11, FS12b, wKK12, MU14, RPS^{+}11, SS14b, UM14]. weighting [GK11]. weights [BCGMB10, Kni10, MO10]. Welch [MD10c]. Welding [FRB13]. well [CBB13, HHM12, PW12, YC11]. well-balanced [CBB13, HHM12, YC11]. well-posedness [PW12]. Wendroff [CET14, Li12]. WENO [ZZS11]. West [Nie13b]. Wheeler [DHR12]. wheels [BSPD14]. which [BI12, Law14, Mah11, Vit10]. Whittaker [Che13a]. wide [FF14]. width [Ros11]. Wielandt [Che13b, LHLL10]. Wielandt-type [Che13b]. Wiener [NH13, XL12, YAKA11]. Wilkinson [AMMS12]. Wilkinson-like [AMMS12]. William [ARV10a]. Wilson [HJS14]. Wind [NMBG14, HDL^{+}13, SSF^{+}11]. window [FO10, VMS12]. wire [CC11c]. wires [Mee12]. Wishart [LWN13]. within [WK14]. without [Fuk13d, JK10c, bJfCwH14, KSK11, dO10]. working [TJL11, YWW10, YTFP11, ZH10a]. world [BM12]. Wright [Par10].

X [HLF14]. X-ray [HLF14].

Yakubovich [SRWZ14]. Yakubovich-transpose [SRWZ14]. yardstick [LZT13]. years [BM12]. yield [AMPR10]. yielding [LL11d]. Young [AHSN13]. Yuan [XWW10, And10, XWW10].

Zakai [Lan12]. Zakharov [NNO12]. Zero [WW11b, HC14, MGZ10, ZYL11]. zerocoupon [ZYL11]. Zero-Hopf [WW11b]. Zeros [BHJN10, BG14b, DJJ13, LW14, PRM11, ZD13, vDvFZ09, vDvFZ13]. Zeta [Rza12]. Zheng [AHS14]. Zienkiewicz [AR10b]. Zienkiewicz-type [AR10b]. zone [BDMP13]. zones [ZTP14].

References

Ahmedov:2012:AFF

[AA12] Anvarjon A. Ahmedov and Norashikin Abdul Aziz. The approximation of functions from LlogL(log log L)(S^{N}) by Fourier–Laplace series. Journal of Computational and Applied Mathematics, 236(15):3827–3834, September 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                                       www.sciencedirect.com/
                                                       science/article/pii/S037704271100464X.

Abushov:2014:SMP

[AA14a] Qurban Abushov and Charkaz Aghayeva. Stochastic maximum principle for nonlinear optimal control problem of switching systems. Journal of Computational and Applied Mathematics, 259 (part B)(??):371–376, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                                       www.sciencedirect.com/
                                                       science/article/pii/S0377042713003087.

Akyildiz:2014:OTB

[AA14b] Ersan Akyildiz and Muhammad Ashraf. An overview of trace based public key cryptography over finite fields. Journal of Computational and Applied Mathematics, 259 (part B)(??):599–621, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005669.

Askar:2014:DEG

[AA14c] S. S. Askar and Ahmad Alshamrani. The dynamics of economic games based on product differentiation. Journal of Computational and Applied Mathematics, 268(??):135–144, October 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714001277.

Assari:2013:MMS

[AAD13] Pouria Assari, Hojatollah Adibi, and Mehdi Dehghan. A meshless method for solving nonlinear twodimensional integral equations of the second kind on non-rectangular domains using radial basis functions with error analysis. Journal of Computational and Applied Mathematics, 239(??):72–92, February 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003743.

Assari:2014:MDG

[AAD14] Pouria Assari, Hojatollah Adibi, and Mehdi Dehghan. A meshless discrete Galerkin (MDG) method for the numerical solution of integral equations with logarithmic kernels. Journal of Computational and Applied Mathematics, 267(??):160– 181, September 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704271400065X.

Alvarez:2013:EPP

[ÁAG^{+}13] E. Álvarez, A. Arcos, S. González, J. F. Muñoz, and M. Rueda. Estimating population proportions in the presence of missing data. Journal of Computational and Applied Mathematics, 237(1):470–476, January 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271200252X.

Aihara:2014:VIR

[AAI14] Kensuke Aihara, Kuniyoshi Abe, and Emiko Ishiwata. A variant of IDRstab with reliable update strategies for solving sparse linear systems. Journal of Computational and Applied Mathematics, 259 (part A)(??):244–258, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713004408.

Albrecher:2010:EER

[AAK10] Hansjörg Albrecher, Florin Avram, and Dominik Kortschak. On the efficient evaluation of ruin probabilities for completely monotone claim distributions. Journal of Computational and Applied Mathematics, 233(10):2724–2736, March 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709007626.

Abia:2014:NIH

[AALMLM14] L. M. Abia, O. Angulo, J. C.

López-Marcos, and M. A. López-Marcos. Numerical integration of a hierarchically size-structured population model with contest competition. Journal of Computational and Applied Mathematics, 258(??):116– 134, March 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713004536.

Aghajani:2014:GDT

[AAM14] A. Aghajani, R. Allahyari, and M. Mursaleen. A generalization of Darbo’s theorem with application to the solvability of systems of integral equations. Journal of Computational and Applied Mathematics, 260(??): 68–77, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713004949.

Aktas:2011:NFT

[AAT11] Rabia Aktas, Abdullah Altin, and Fatma Tasdelen. A note on a family of two-variable polynomials. Journal of Computational and Applied Mathematics, 235(16):4825–4833, June 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006217.

Akhmet:2010:SCN

[AAY10] M. U. Akhmet, D. Arugaslan, and E. Yilmaz. Stability in cellular neural networks with a piecewise constant argument. Journal of Computational and Applied Mathematics, 233(9):2365– 2373, March 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709007195.

Akhmet:2011:MLF

[AAY11] M. U. Akhmet, D. Arugaslan, and E. Yilmaz. Method of Lyapunov functions for differential equations with piecewise constant delay. Journal of Computational and Applied Mathematics, 235(16):4554– 4560, June 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704271000141X.

Armentano:2010:SUC

[AB10] María G. Armentano and Jordi Blasco. Stable and unstable cross-grid P_{k}Q_{l} mixed finite elements for the Stokes problem. Journal of Computational and Applied Mathematics, 234(5):1404– 1416, July 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710000981.

Aktuglu:2011:CMS

[AB11a] Hüseyin Aktuğlu and Şerife Bekar. q-Ceságro matrix and q-statistical convergence. Journal of Computational and Applied Mathematics, 235(16):4717–4723, June 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710004565.

Allasia:2011:TIO

[AB11b] Giampietro Allasia and Cesare Bracco. Two interpolation operators on irregularly distributed data in inner product spaces. Journal of Computational and Applied Mathematics, 235(7):1763– 1774, February 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042710002335.

Annunziato:2013:FPC

[AB13] M. Annunziato and A. Borzì. A Fokker–Planck control framework for multidimensional stochastic processes. Journal of Computational and Applied Mathematics, 237(1):487–507, January 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712002543.

Amodio:2014:RAB

[AB14] Pierluigi Amodio and Luigi Brugnano. Recent advances in bibliometric indexes and the PaperRank problem. Journal of Computational and Applied Mathematics, 267(??):182–194, September 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714001046.

Alp:2011:CGC

[ABÇ^{+}11] Özge Sezgin Alp, Erkan Büyükbebeci, Aysegül Iscanoğlu Çekiç, Fatma Yerlikaya Özkurt, Pakize Taylan, and Gerhard-Wilhelm Weber. CMARS and GAM & CQP — modern optimization methods applied to international credit default prediction. Journal of Computational and Applied Mathematics, 235(16):4639– 4651, June 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710002591.

Amat:2014:TOM

[ABCP14] S. Amat, S. Busquier, R. Castro, and S. Plaza. Third-order methods on Riemannian manifolds under Kantorovich conditions. Journal of Computational and Applied Mathematics, 255(??):106–121, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713002094.

Amat:2011:TOI

[ABG11] S. Amat, S. Busquier, and J. M. Gutiérrez. Thirdorder iterative methods with applications to Hammerstein equations: a unified approach. Journal of Computational and Applied Mathematics, 235(9):2936– 2943, March 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710006618.

Aarao:2011:NIE

[ABHMW11] J. Aarão, B. H. Bradshaw-Hajek, S. J. Miklavcic, and D. A. Ward. Numerical implementation of the EDEM for modified Helmholtz BVPs on annular domains. Journal of Computational and Applied Mathematics, 235(5):1342– 1353, January 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710004607.

Abbadi:2010:GMC

[ABIS10] A. Abbadi, D. Barrera, M. J. Ibáñez, and D. Sbibih. A general method for constructing quasi-interpolants from B-splines. Journal of Computational and Applied Mathematics, 234(4):1324– 1337, June 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704271000021X.

Avrachenkov:2010:QSD

[ABN10] Konstantin Avrachenkov, Vivek Borkar, and Danil Nemirovsky. Quasi-stationary distributions as centrality measures for the giant strongly connected component of a reducible graph. Journal of Computational and Applied Mathematics, 234(11):3075–3090, October 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710000725.

Abdou:2011:MST

[ABS11a] M. A. Abdou, A. A. Badr, and M. B. Soliman. On a method for solving a twodimensional nonlinear integral equation of the second kind. Journal of Computational and Applied Mathematics, 235(12):3589–3598, April 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711000926.

Andreianov:2011:FVM

[ABS11b] Boris Andreianov, Mostafa Bendahmane, and Mazen Saad. Finite volume methods for degenerate chemotaxis model. Journal of Computational and Applied Mathematics, 235(14):4015– 4031, May 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042711001117.

Amireghbali:2014:EAC

[AC14a] Aydin Amireghbali and Demirkan Coker. Elastic analysis of a circumferential crack in an isotropic curved beam using the modified mappingcollocation method. Journal of Computational and Applied Mathematics, 264(??): 131–138, July 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042714000090.

Arnal:2014:EPF

[AC14b] Ana Arnal and Fernando Casas. Exponential polar factorization of the fundamental matrix of linear differential systems. Journal of Computational and Applied Mathematics, 268(??):168– 178, October 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042714001344.

Assous:2014:ICN

[AC14c] F. Assous and J. Chaskalovic.

Indeterminate constants in numerical approximations of PDEs: a pilot study using data mining techniques. Journal of Computational and Applied Mathematics, 270(??):462–470, November 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713006882.

Assous:2014:PAM

[AC14d] F. Assous and J. Chaskalovic.

A paraxial asymptotic model for the coupled Vlasov–Maxwell problem in electromagnetics. Journal of Computational and Applied Mathematics, 270(??): 369–385, November 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042713007103.

Ala:2014:NMI

[ACF^{+}14] Guido Ala, Pietro Cassarà, Elisa Francomano, Salvatore Ganci, Giuseppe Caruso, and Pio D. Gallo. A numerical method for imaging of biological microstructures by VHF waves. Journal of Computational and Applied Mathematics, 259 (part B)(??):805–814, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713003646.

Avram:2011:MBP

[ACH11] F. Avram, D. F. Chedom, and A. Horváth. On moments based Padé approximations of ruin probabilities. Journal of Computational and Applied Mathematics, 235(10):3215–3228, March 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711000094.

Alvarez:2010:EPN

[ÁD10a] J. Álvarez and A. Durán. Error propagation in numerical approximations near relative equilibria. Journal of Computational and Applied Mathematics, 234 (12):3373–3386, October 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042710002463.

Apreutesei:2010:PPR

[AD10b] Narcisa Apreutesei and Gabriel Dimitriu. On a prey-predator reaction– diffusion system with Holling type III functional response. Journal of Computational and Applied Mathematics, 235(2):366–379, November 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710003304.

Alvarez:2011:NSP

[ÁD11a] J. Álvarez and A. Durán.

A numerical scheme for periodic travelling-wave simulations in some nonlinear dispersive wave models. Journal of Computational and Applied Mathematics, 235(7):1790–1797, February 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710004127.

Aoki:2011:AHO

[AD11b] Yasunori Aoki and Hans De Sterck. Augmented high order finite volume element method for elliptic PDEs in non-smooth domains: Convergence study. Journal of Computational and Applied Mathematics, 235(17):5177– 5187, July 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711002512.

Alvarez:2012:INP

[ÁD12] J. Álvarez and A. Durán.

On the influence of numerical preservation of invariants when simulating Hamiltonian relative periodic orbits. Journal of Computational and Applied Mathematics, 236(12):2954– 2961, June 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004936.

Alvarez:2014:PTM

[ÁD14] J. Álvarez and A. Durán. Petviashvili type methods for traveling wave computations: I. Analysis of convergence. Journal of Computational and Applied Mathematics, 266(??):39–51, August 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714000363.

See corrigendum [ÁD15].

Alvarez:2015:CPT

[ÁD15] J. Álvarez and A. Durán. Corrigendum to “Petviashvili type methods for traveling wave computations: I. Analysis of convergence” [J. Comput. Appl. Math. 266 (2014) 39– 51]. Journal of Computational and Applied Mathematics, 277(??):215–216, March 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714004002.

See [ÁD14].

Ahmadi:2013:ELP

[ADA13] Mohammad Vali Ahmadi, Mahdi Doostparast, and Jafar Ahmadi. Estimating the lifetime performance index with Weibull distribution based on progressive first-failure censoring scheme. Journal of Computational and Applied Mathematics, 239(??):93–102, February 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003706.

Alvarez-Dios:2012:GSV

[ADCC12] J. A. Alvarez-Dios, J. C.

Cabaleiro, and G. Casal. A generalization of s-step variants of gradient methods. Journal of Computational and Applied Mathematics, 236(12):2938–2953, June 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712000398.

Aimi:2011:EGB

[ADG11] A. Aimi, M. Diligenti, and C. Guardasoni. On the energetic Galerkin boundary element method applied to interior wave propagation problems. Journal of Computational and Applied Mathematics, 235(7):1746– 1754, February 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710000828.

Alonso:2011:GFP

[ADGP11] Pedro Alonso, Jorge Delgado, Rafael Gallego, and Juan Manuel Peña. Growth factors of pivoting strategies associated with Neville elimination. Journal of Computational and Applied Mathematics, 235(7):1755– 1762, February 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709007432.

Alonso:2012:NMM

[ADGP12] Pedro Alonso, Jorge Delgado, Rafael Gallego, and Juan Manuel Peña. A note on matrices with maximal growth factor for Neville elimination. Journal of Computational and Applied Mathematics, 236(12):2971– 2974, June 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711003931.

Alonso:2013:CAC

[ADGP13] Pedro Alonso, Jorge Delgado, Rafael Gallego, and Juan Manuel Peña. Conditioning and accurate computations with Pascal matrices. Journal of Computational and Applied Mathematics, 252(??):21– 26, November 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711006212.

Amat:2010:FSN

[ADL^{+}10a] S. Amat, K. Dadourian, J. Liandrat, J. Ruiz, and J. C. Trillo. A family of stable nonlinear nonseparable multiresolution schemes in 2D. Journal of Computational and Applied Mathematics, 234(4):1277–1290, June 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709006724.

Amat:2010:CSN

[ADL^{+}10b] S. Amat, K. Dadourian, J. Liandrat, J. Ruiz, and J. C. Trillo. On a class of L^{1}-stable nonlinear cell-average multiresolution schemes. Journal of Computational and Applied Mathematics, 234(4):1129–1139, June 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709003513.

Anguelov:2014:DCN

[ADLS14] R. Anguelov, Y. Dumont, J. M.-S. Lubuma, and M. Shillor. Dynamically consistent nonstandard finite difference schemes for epidemiological models. Journal of Computational and Applied Mathematics, 255(??):161–182, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713002380.

Amat:2013:HON

[ADLT13] S. Amat, K. Dadourian, J. Liandrat, and J. C. Trillo. High order nonlinear interpolatory reconstruction operators and associated multiresolution schemes. Journal of Computational and Applied Mathematics, 253(??):163–180, December 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271300112X.

Aptekarev:2010:AOP

[ADMF10] A. I. Aptekarev, J. S. Dehesa, and A. Martinez-Finkelshtein. Asymptotics of orthogonal polynomial’s entropy. Journal of Computational and Applied Mathematics, 233(6):1355–1365, January 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704270900140X.

Alonso-Duran:2012:ASS

[ADRM12] María Alonso-Durán and Luis Rodríguez-Marín. On approximate solutions in set-valued optimization problems. Journal of Computational and Applied Mathematics, 236(17):4421–4427, November 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712001720.

Alvarez-Estrada:2010:BMQ

[ÁE10] R. F. Álvarez-Estrada.

Brownian motion, quantum corrections and a generalization of the Hermite polynomials. Journal of Computational and Applied Mathematics, 233(6):1453–1461, January 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709001472.

Alipanah:2011:NST

[AE11a] Amjad Alipanah and Shahrokh Esmaeili. Numerical solution of the two-dimensional Fredholm integral equations using Gaussian radial basis function. Journal of Computational and Applied Mathematics, 235(18):5342– 5347, July 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709008000.

Alvarez-Estrada:2011:QBM

[ÁE11b] R. F. Álvarez-Estrada. Quantum Brownian motion and generalizations of Hermite polynomials. Journal of Computational and Applied Mathematics, 236(1): 7–18, August 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710000646.

Alvarez-Estrada:2012:TIT

[ÁECS12] Ramón F. Álvarez-Estrada, Gabriel F. Calvo, and Hélia Serrano. A transfer integral technique for solving a class of linear integral equations: Convergence and applications to DNA. Journal of Computational and Applied Mathematics, 236(15):3561– 3571, September 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711002160.

Argyros:2011:SCE

[AEG^{+}11] I. K. Argyros, J. A. Ezquerro, J. M. Gutiérrez, M. A. Hernández, and S. Hilout. On the semilocal convergence of efficient Chebyshev–secanttype methods. Journal of Computational and Applied Mathematics, 235(10):3195– 3206, March 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711000069.

Aladag:2010:IWI

[AEGB10]  Çagdas Hakan Aladag, Erol

Egrioğlu, Suleyman Gunay, and Murat A. Basaran. Improving weighted information criterion by using optimization. Journal of Computational and Applied Mathematics, 233(10):2683– 2687, March 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704270900747X.

Artalejo:2010:MNI

[AELH10] J. R. Artalejo, A. Economou, and M. J. Lopez-Herrero. The maximum number of infected individuals in SIS epidemic models: Computational techniques and quasi-stationary distributions. Journal of Computational and Applied Mathematics, 233(10):2563–2574, March 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709007341.

Abd-El-Wahed:2011:IPS

[AEWMES11] W. F. Abd-El-Wahed, A. A. Mousa, and M. A. El-Shorbagy. Integrating particle swarm optimization with genetic algorithms for solving nonlinear optimization problems. Journal of Computational and Applied Mathematics, 235(5):1446– 1453, January 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710004796.

Alharbi:2010:APS

[AF10] Abir Alharbi and E. S. Fahmy. ADM–Padé solutions for generalized Burgers and Burgers–Huxley systems with two coupled equations. Journal of Computational and Applied Mathematics, 233(8):2071–2080, February 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709006645.

Ali:2010:SAD

[AG10a] M. M. Ali and M. N.

Gabere. A simulated annealing driven multi-start algorithm for bound constrained global optimization. Journal of Computational and Applied Mathematics, 233(10):2661–2674, March 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709007444.

An:2010:MRL

[AG10b] Yu’e An and Chuanqing Gu.

Model reduction for largescale dynamical systems via equality constrained least squares. Journal of Computational and Applied Mathematics, 234(8):2420–2431, August 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710001470.

Adimurthi:2013:DFS

[AGJ13] Adimurthi, G. D. Veerappa Gowda, and Jérôme Jaffré. The DFLU flux for systems of conservation laws. Journal of Computational and Applied Mathematics, 247(??):102–123, August 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713000174.

Akyildiz:2014:RAA

[AGK^{+}14] Ersan Akyildiz, Ömer L. Gebizlioğlu, Bülent Karasözen, Ömür Ugur, and Gerhard Wilhelm Weber. Recent advances in applied and computational mathematics: ICACM–IAM– METU. Journal of Computational and Applied Mathematics, 259 (part B)(??): 327–328, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042713005773.

Aceto:2012:PFS

[AGM12] Lidia Aceto, Paolo Ghelardoni, and Cecilia Magherini. PGSCM: a family of P - stable Boundary Value Methods for second-order initial value problems. Journal of Computational and Applied Mathematics, 236(16):3857–3868, October 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712001525.

Allouche:2012:DDP

[AGT12] Hassane Allouche, Noura Ghanou, and Khalid Tigma. Detecting discontinuity points from spectral data with the quotient-difference (qd) algorithm. Journal of Computational and Applied Mathematics, 236(9):2406–2424, March 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711005978.

Avron:2013:SHP

[AGT13] Haim Avron, Anshul Gupta, and Sivan Toledo. Solving Hermitian positive definite systems using indefinite incomplete factorizations. Journal of Computational and Applied Mathematics, 243(??):126–138, May 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712005043.

Ablowitz:2010:NWO

[AH10a] Mark J. Ablowitz and Theodoros P. Horikis. Nonlinear waves in optical media. Journal of Computational and Applied Mathematics, 234(6):1896–1903, July 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709005172.

Argyros:2010:ENK

[AH10b] Ioannis K. Argyros and Saïd Hilout. Extending the Newton–Kantorovich hypothesis for solving equations. Journal of Computational and Applied Mathematics, 234(10):2993–3006, September 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042710002116.

Argyros:2012:ILC

[AH12a] Ioannis K. Argyros and Saïd Hilout. Improved local convergence of Newton’s method under weak majorant condition. Journal of Computational and Applied Mathematics, 236(7):1892– 1902, January 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042711005681.

Argyros:2012:MSI

[AH12b] Ioannis K. Argyros and Saïd Hilout. Majorizing sequences for iterative methods. Journal of Computational and Applied Mathematics, 236(7):1947–1960, January 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711005735.

Argyros:2013:LCF

[AH13] I. K. Argyros and S. Hilout.

On the local convergence of fast two-step Newton-like methods for solving nonlinear equations. Journal of Computational and Applied Mathematics, 245(??): 1–9, June 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271200533X.

Adcock:2014:RPF

[AH14a] Ben Adcock and Daan Huybrechs. On the resolution power of Fourier extensions for oscillatory functions. Journal of Computational and Applied Mathematics, 260(??):312–336, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005244.

Angelini:2014:DHD

[AH14b] Flavio Angelini and Stefano Herzel. Delta hedging in discrete time under stochastic interest rate. Journal of Computational and Applied Mathematics, 259 (part B)(??):385–393, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713003208.

Abdel-Hamid:2011:IPS

[AHAH11] Alaa H. Abdel-Hamid and Essam K. Al-Hussaini. Inference for a progressive stress model from Weibull distribution under progressive Type-II censoring. Journal of Computational and Applied Mathematics, 235(17):5259–5271, July 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711002925.

An:2013:DMS

[AHH13] P. T. An, N. N. Hai, and T. V. Hoai. Direct multiple shooting method for solving approximate shortest path problems. Journal of Computational and Applied Mathematics, 244(??): 67–76, May 15, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042712004839.

Alm:2014:WM

[AHK14a] Daniel Alm, Helmut Harbrecht, and Ulf Krämer. The H^{2}-wavelet method. Journal of Computational and Applied Mathematics, 267(??):131–159, September 2014. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042714000582.

Argyros:2014:EAN

[AHK14b] Ioannis K. Argyros, Saïd Hilout, and Sanjay K. Khattri. Expanding the applicability of Newton’s method using Smale’s αtheory. Journal of Computational and Applied Mathematics, 261(??):183–200, May 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713006195.

Akkari:2014:MNS

[AHLJ14] N. Akkari, A. Hamdouni, E. Liberge, and M. Jazar. A mathematical and numerical study of the sensitivity of a reduced order model by POD(ROM-POD), for a 2D incompressible fluid flow. Journal of Computational and Applied Mathematics, 270(??):522–530, November 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713006572.

Amandolese:2013:TTR

[AHM13] X. Amandolese, P. Hémon, and S. Manzoor. The transient temporal response of a flexible bridge deck subjected to a single gust. Journal of Computational and Applied Mathematics, 246 (??):153–160, July 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042712003172.

Alcazar:2014:DSR

[AHM14] Juan Gerardo Alcázar, Carlos Hermoso, and Georg Muntingh. Detecting similarity of rational plane curves. Journal of Computational and Applied Mathematics, 269(??):1–13, October 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714001538.

Ahn:2014:GCQ

[AHR14] Young Joon Ahn, Christoph Hoffmann, and Paul Rosen. Geometric constraints on quadratic Bézier curves using minimal length and energy. Journal of Computational and Applied Mathematics, 255(??):887–897, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271300349X.

Ak:2014:ICS

[AHS14] Murat Ak, Turgut Hanoymak, and Ali Aydin Selçuk. IND–CCA secure encryption based on a Zheng– Seberry scheme. Journal of Computational and Applied Mathematics, 259 (part B) (??):529–535, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042713003403.

Ait-Haddou:2013:CBM

[AHSN13] Rachid Ait-Haddou, Yusuke Sakane, and Taishin Nomura. Chebyshev blossoming in Müntz spaces: Toward shaping with Young diagrams. Journal of Computational and Applied Mathematics, 247(??):172– 208, August 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713000289.

Aoto:2010:VPG

[AIA10] Daisuke Aoto, Emiko Ishiwata, and Kuniyoshi Abe. A variable preconditioned GCR() method using the GSOR method for singular and rectangular linear systems. Journal of Computational and Applied Mathematics, 234(3):703– 712, June 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710000142.

Ainsworth:2010:FOG

[Ain10] Mark Ainsworth. A framework for obtaining guaranteed error bounds for finite element approximations. Journal of Computational and Applied Mathematics, 234(9):2618–2632, September 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710000415.

Amaziane:2012:NSW

[AJK12] Brahim Amaziane, Mladen Jurak, and Ana Zgaljić Keko. Numerical simulations of water-gas flow in heterogeneous porous media with discontinuous capillary pressures by the concept of global pressure. Journal of Computational and Applied Mathematics, 236(17):4227– 4244, November 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042712002233.

Aydin:2011:LII

[AK11] A. Aydin and B. Karasözen. Lobatto IIIA–IIIB discretization of the strongly coupled nonlinear Schrödinger equation. Journal of Computational and Applied Mathematics, 235(16):4770– 4779, June 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710005303.

Akman:2014:VTD

[AK14a] Tugba Akman and Bülent Karasözen. Variational time discretization methods for optimal control problems governed by diffusionconvection-reaction equations. Journal of Computational and Applied Mathematics, 272(??):41–56, December 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714002350.

Argyros:2014:WKT

[AK14b] I. K. Argyros and S. K.

Khattri. Weaker Kantorovich type criteria for inexact Newton methods. Journal of Computational and Applied Mathematics, 261(??):103–117, May 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713006018.

Ashraf:2014:MTG

[AK14c] Muhammad Ashraf and Baris Bülent Kirlar. Message transmission for GHpublic key cryptosystem. Journal of Computational and Applied Mathematics, 259 (part B)(??):578–585, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713005499.

Axelsson:2014:RSC

[AK14d] O. Axelsson and J. Karátson. Reaching the superlinear convergence phase of the CG method. Journal of Computational and Applied Mathematics, 260(??):244– 257, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713005451.

Alshangiti:2014:NFM

[AKA14] Arwa M. Alshangiti, M. Kayid, and B. Alarfaj. A new family of Marshall– Olkin extended distributions. Journal of Computational and Applied Mathematics, 271(??):369–379, December 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042714002192.

Auzinger:2012:DBL

[AKT12] Winfried Auzinger, Othmar Koch, and Mechthild Thalhammer. Defect-based local error estimators for splitting methods, with application to Schrödinger equations, Part I: the linear case. Journal of Computational and Applied Mathematics, 236(10):2643–2659, April 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712000027.

Aktuglu:2014:KTA

[Akt14a]  Hüseyin Aktuğlu.  Ko-

rovkin type approximation theorems proved via αβ-statistical convergence. Journal of Computational and Applied Mathematics, 259 (part A)(??):174–181, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713002860.

Auzinger:2014:DBL

[AKT14b] Winfried Auzinger, Othmar Koch, and Mechthild Thalhammer. Defect-based local error estimators for splitting methods, with application to Schrödinger equations, Part II. Higher-order methods for linear problems. Journal of Computational and Applied Mathematics, 255(??):384–403, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713002495.

Allgower:2013:F

[AL13] Eugene Allgower and James Liu. Foreword. Journal of Computational and Applied Mathematics, 254(??):1, December 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713001337.

Alcazar:2014:EDS

[Alc14] Juan Gerardo Alcázar. Efficient detection of symmetries of polynomially parametrized curves. Journal of Computational and Applied Mathematics, 255 (??):715–724, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042713003397.

Ali:2011:NGJ

[Ali11a]  Ahmad T. Ali.  New

generalized Jacobi elliptic function rational expansion method. Journal of Computational and Applied Mathematics, 235(14):4117–4127, May 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711001257.

Ali:2011:DEG

[Ali11b] M. M. Ali. Differential evolution with generalized differentials. Journal of Computational and Applied Mathematics, 235(8):2205– 2216, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042710005844.

Arnal:2011:PTB

[ALM11] A. Arnal, A. Lluch, and J. Monterde. PDE triangular Bézier surfaces: Harmonic, biharmonic and isotropic surfaces. Journal of Computational and Applied Mathematics, 235(5): 1098–1113, January 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042710004188.

Alzabut:2010:APS

[Alz10] Jehad O. Alzabut. Almost periodic solutions for an impulsive delay Nicholson’s blowflies model. Journal of Computational and Applied Mathematics, 234(1):233– 239, May 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709008279.

Adam:2010:GLT

[AM10a] M. Adam and J. Maroulas. The generalized levinger transformation. Journal of Computational and Applied Mathematics, 233(11):3018– 3029, April 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042709007985.

Airaksinen:2010:CBS

[AM10b] Tuomas Airaksinen and Sanna Mönkölä. Comparison between the shifted-Laplacian preconditioning and the controllability methods for computational acoustics. Journal of Computational and Applied Mathematics, 234(6):1796– 1802, July 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042709005056.

Amat:2010:NUM

[AM10c] Sergio Amat and María Moncayo. Non-uniform multiresolution analysis with supercompact multiwavelets. Journal of Computational and Applied Mathematics, 235(1):334– 340, November 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704271000289X.

Arnal:2011:TOP

[AM11] A. Arnal and J. Monterde.

A third order partial differential equation for isotropic boundary based triangular Bézier surface generation. Journal of Computational and Applied Mathematics, 236(2):184–195, August 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711003530.

Almasieh:2014:NSC

[AM14a] H. Almasieh and J. Nazari Meleh. Numerical solution of a class of mixed two-dimensional nonlinear Volterra–Fredholm integral equations using multiquadric radial basis functions. Journal of Computational and Applied Mathematics, 260(??):173–179, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005220.

Armentano:2014:PEE

[AM14b] María G. Armentano and Verónica Moreno. A posteriori error estimates of stabilized low-order mixed finite elements for the Stokes eigenvalue problem. Journal of Computational and Applied Mathematics, 269(??):132–149, October 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042714001769.

Akkocaoglu:2013:HBA

[AMÇ13] Hande Akkocaoğlu, Hüseyin Merdan, and Canan Çelik. Hopf bifurcation analysis of a general non-linear differential equation with delay. Journal of Computational and Applied Mathematics, 237(1):565–575, January 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712002713.

Aishima:2010:SCD

[AMMS10] Kensuke Aishima, Takayasu Matsuo, Kazuo Murota, and Masaaki Sugihara. Superquadratic convergence of DLASQ for computing matrix singular values. Journal of Computational and Applied Mathematics, 234(4):1179–1187, June 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709004221.

Aishima:2012:WLM

[AMMS12] Kensuke Aishima, Takayasu Matsuo, Kazuo Murota, and Masaaki Sugihara. A Wilkinson-like multishift QR algorithm for symmetric eigenvalue problems and its global convergence. Journal of Computational and Applied Mathematics, 236(15):3556–3560, September 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711001890.

Aishima:2014:SSS

[AMMS14] Kensuke Aishima, Takayasu Matsuo, Kazuo Murota, and Masaaki Sugihara. A shift strategy for superquadratic convergence in the dqds algorithm for singular values. Journal of Computational and Applied Mathematics, 257(??):132–143, February 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713004263.

Ahcan:2011:QAA

[AMPP11] Ales Ahcan, Igor Masten, Saso Polanec, and Mihael Perman. Quantile approximations in autoregressive portfolio models. Journal of Computational and Applied Mathematics, 235(8):1976–1983, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710005388.

Alfaro:2010:WDL

[AMPR10] Manuel Alfaro, Francisco Marcellán, Ana Peña, and M. Luisa Rezola. When do linear combinations of orthogonal polynomials yield new sequences of orthogonal polynomials? Journal of Computational and Applied Mathematics, 233(6):1446– 1452, January 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042709001460.

Arandiga:2013:LNL

[AMR13] Francesc Aràndiga, Pep Mulet, and Vicent Renau. Lossless and nearlossless image compression based on multiresolution analysis. Journal of Computational and Applied Mathematics, 242(??):70– 81, April 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712004682.

Abad:2012:AKS

[AMRS12] J. O. Abad, S. Morigi, L. Reichel, and F. Sgallari. Alternating Krylov subspace image restoration methods. Journal of Computational and Applied Mathematics, 236(8):2049– 2062, February 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711005097.

Ahmed:2011:FEM

[AMT11] Naveed Ahmed, Gunar Matthies, and Lutz Tobiska. Finite element methods of an operator splitting applied to population balance equations. Journal of Computational and Applied Mathematics, 236(6):1604–1621, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711005048.

Aledo:2013:PDD

[AMV13] Juan A. Aledo, S. Martinez, and Jose C. Valverde. Parallel discrete dynamical systems on independent local functions. Journal of Computational and Applied Mathematics, 237(1):335– 339, January 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712002373.

Alvarez-Nodarse:2014:LGN

[ÁNA14] R. Álvarez-Nodarse and R. Sevinik Adigüzel. On the limit of generalized nonstandard q-Racah polynomials. Journal of Computational and Applied Mathematics, 259 (part B)(??): 877–886, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042713004317.

Andrei:2010:NAC

[And10] Neculai Andrei. New accelerated conjugate gradient algorithms as a modification of Dai–Yuan’s computational scheme for unconstrained optimization. Journal of Computational and Applied Mathematics, 234(12):3397–3410, October 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710002566.

Andrei:2013:STT

[And13] Neculai Andrei. A simple three-term conjugate gradient algorithm for unconstrained optimization. Journal of Computational and Applied Mathematics, 241 (??):19–29, March 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042712004219.

Anonymous:2010:AIV

[Ano10a] Anonymous. Author index volume 233 (2010). Journal of Computational and Applied Mathematics, 233(12): 3189–3199, April 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042710000932.

Anonymous:2010:MBS

[Ano10b] Anonymous. In memoriam Ben Sommeijer: (Nov. 4, 1949–Nov. 19, 2009). Journal of Computational and Applied Mathematics, 233 (9):2422, March 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042709008115.

Anonymous:2010:P

[Ano10c]  Anonymous.  Preface.

Journal of Computational and Applied Mathematics, 235(4):879–881, December 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710004929.

Anonymous:2011:AIVa

[Ano11a] Anonymous. Author index volume 235 (2011). Journal of Computational and Applied Mathematics, 235(12): 3691–3701, April 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704271100166X.

Anonymous:2011:AIVb

[Ano11b] Anonymous. Author index volume 235 (2011). Journal of Computational and Applied Mathematics, 235(18): 5441–5456, July 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042711003839.

Anonymous:2011:EB

[Ano11c]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 235(11):i, April 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711001300.

Apel:2011:PEE

[ANS11] Thomas Apel, Serge Nicaise, and Dieter Sirch. A posteriori error estimation of residual type for anisotropic diffusion-convection-reaction problems. Journal of Computational and Applied Mathematics, 235(8):2805– 2820, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042710006503.

Antczak:2011:CVS

[Ant11] Tadeusz Antczak. Characterization of vector strict global minimizers of order 2 in differentiable vector optimization problems under a new approximation method. Journal of Computational and Applied Mathematics, 235(17):4991–5000, July 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711002238.

Arslan:2011:JPM

[AÖ11] Güvenç Arslan and Ilknur Özmen. A JAVA program for the multivariate Z_{p} and C_{p} tests and its application. Journal of Computational and Applied Mathematics, 235(16):4729–4735, June 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271000484X.

Agram:2014:IHO

[AØ14a] Nacira Agram and Bernt Øksendal. Infinite horizon optimal control of forwardbackward stochastic differential equations with delay. Journal of Computational and Applied Mathematics, 259 (part B)(??):336–349, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713002549.

Al-Omari:2014:EEU

[AO14b] Amer Ibrahim Al-Omari. Estimation of entropy using random sampling. Journal of Computational and Applied Mathematics, 261(??): 95–102, May 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713006006.

Akleylek:2014:AOF

[AÖÖ14] S. Akleylek, F. Özbudak, and C. Özel. On the arithmetic operations over finite fields of characteristic three with low complexity. Journal of Computational and Applied Mathematics, 259 (part B)(??):546–554, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713004160.

Armaneri:2010:IDS

[AÖY10] Özgür Armaneri, Güzin Özdagoğlu, and Özgür Yalçinkaya. An integrated decision support approach for project investors in risky and uncertain environments. Journal of Computational and Applied Mathematics, 234(8):2530–2542, August 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710001664.

Anguelov:2010:TSP

[AP10] Roumen Anguelov and Evgenija D. Popova. Topological structure preserving numerical simulations of dynamical models. Journal of Computational and Applied Mathematics, 235(2):358– 365, November 15, 2010.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042710003286.

Amami:2014:SEI

[AP14a] Rim Amami and Monique Pontier. Solution examples of an impulse control problem. Journal of Computational and Applied Mathematics, 259 (part B)(??): 695–710, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042713002343.

Aydogan:2014:CCS

[AP14b] Melike Aydoğan and Yasar Polatoğlu. A certain class of starlike log-harmonic mappings. Journal of Computational and Applied Mathematics, 270(??):506–509, November 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713006626.

Arraras:2010:LLF

[APJ10] A. Arrarás, L. Portero, and J. C. Jorge. Locally linearized fractional step methods for nonlinear parabolic problems. Journal of Computational and Applied Mathematics, 234(4):

1117–1128, June 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042709003501.

Abdollahzadeh:2014:TDN

[APO14] M. Abdollahzadeh, J. C.

Páscoa, and P. J. Oliveira. Two-dimensional numerical modeling of interaction of micro-shock wave generated by nanosecond plasma actuators and transonic flow. Journal of Computational and Applied Mathematics, 270(??):401– 416, November 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713007036.

Ak:2014:ATR

[APS14] Murat Ak, Serdar Pehlivanoğlu, and Ali Aydin Selçuk. Anonymous trace and revoke. Journal of Computational and Applied Mathematics, 259 (part B)(??):586–591, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005633.

Arnold:2014:P

[APW^{+}14a] Martin Arnold, Helmut Podhaisky, Rüdiger Weiner, Jason Frank, and Willem Hundsdorfer. Preface. Journal of Computational and Applied Mathematics, 262 (??):1–2, May 15, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042713005761.

Azevedo:2014:DPM

[APW14b] N. Azevedo, D. Pinheiro, and G.-W. Weber. Dynamic programming for a Markov-switching jumpdiffusion. Journal of Computational and Applied Mathematics, 267(??):1–19, September 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042714000491.

Amini:2010:NSQ

[AR10a] Keyvan Amini and Ashraf Ghorbani Rizi. A new structured quasi-Newton algorithm using partial information on Hessian. Journal of Computational and Applied Mathematics, 234(3):805– 811, June 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710000579.

Andreev:2010:ZTF

[AR10b] A. B. Andreev and M. R.

Racheva. A zienkiewicztype finite element applied to fourth-order problems. Journal of Computational and Applied Mathematics, 235(2):348–357, November 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710003274.

Argyros:2011:KTS

[AR11] Ioannis K. Argyros and Hongmin Ren. Kantorovichtype semilocal convergence analysis for inexact Newton methods. Journal of Computational and Applied Mathematics, 235(9):2993– 3005, March 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710006709.

Azad:2014:SBA

[ARF14] Md. Abul Kalam Azad, Ana Maria A. C. Rocha, and Edite M. G. P. Fernandes. A simplified binary artificial fish swarm algorithm for 0– 1 quadratic knapsack problems. Journal of Computational and Applied Mathematics, 259 (part B)(??): 897–904, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0377042713005074.

Arslan:2011:CUD

[Ars11] G. Arslan. On a characterization of the uniform distribution by generalized order statistics. Journal of Computational and Applied Mathematics, 235(16):4532– 4536, June 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S037704271000138X.

Arslan:2014:NCP

[Ars14] G. Arslan. A new characterization of the power distribution. Journal of Computational and Applied Mathematics, 260(??):99– 102, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713005232.

Ammar:2010:DWB

[ARV10a] Greg Ammar, Lothar Reichel, and Marc Van Barel. Dedication to William B. Gragg on the occasion of his 70th birthday. Journal of Computational and Applied Mathematics, 233(5):1191, January 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704270900171X.

Arvesu:2010:SPH

[Arv10b] J. Arvesú. On some properties of q-Hahn multiple orthogonal polynomials. Journal of Computational and Applied Mathematics, 233(6):1462–1469, January 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709001484.

Abe:2010:BVH

[AS10a] Kuniyoshi Abe and Gerard L. G. Sleijpen. BiCR variants of the hybrid BiCG methods for solving linear systems with nonsymmetric matrices. Journal of Computational and Applied Mathematics, 234(4):985– 994, June 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709002088.

Antontsev:2010:BSP

[AS10b] S. Antontsev and S. Shmarev.

Blow-up of solutions to parabolic equations with nonstandard growth conditions. Journal of Computational and Applied Mathematics, 234(9):2633–2645, September 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710000300.

Allahviranloo:2011:FSS

[AS11] T. Allahviranloo and S. Salahshour. Fuzzy symmetric solutions of fuzzy linear systems. Journal of Computational and Applied Mathematics, 235(16):4545–4553, June 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710001408.

Amodio:2012:FDM

[AS12a] Pierluigi Amodio and Giuseppina Settanni. A finite differences MATLAB code for the numerical solution of second order singular perturbation problems. Journal of Computational and Applied Mathematics, 236(16):3869–3879, October 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712001719.

Anastassi:2012:PSL

[AS12b] Z. A. Anastassi and T. E. Simos. A parametric symmetric linear four-step method for the efficient integration of the Schrödinger equation and related oscillatory problems. Journal of Computational and Applied Mathematics, 236(16):3880– 3889, October 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712001355.

Ashrafi:2011:ECI

[ASG11] A. R. Ashrafi, M. Saheli, and M. Ghorbani. The eccentric connectivity index of nanotubes and nanotori. Journal of Computational and Applied Mathematics, 235(16):4561–4566, June 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710001421.

Allahviranloo:2011:MMS

[ASK11] T. Allahviranloo, S. Salahshour, and M. Khezerloo. Maximaland minimal symmetric solutions of fully fuzzy linear systems. Journal of Computational and Applied Mathematics, 235(16):4652–4662, June 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271000275X.

Aslan:2010:CNE

[Asl10] Ismail Aslan. Comment on: “New exact solutions for the Kawahara equation using Exp-function method” [J. Comput. Appl. Math. 233 (2009) 97–102]. Journal of Computational and Applied Mathematics, 234(12):3213–3215, October 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710002347.

See [Ass09].

Abreu:2013:GCS

[ASM13] Rafael Abreu, Daniel Stich, and José Morales. On the generalization of the complex step method. Journal of Computational and Applied Mathematics, 241(??): 84–102, March 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042712004207.

Assas:2009:NES

[Ass09] Laila M. B. Assas. New exact solutions for the Kawahara equation using exp-function method. Journal of Computational and Applied Mathematics, 233(2):97–102, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709004038.

See comment [Asl10].

Allouch:2012:CMS

[ASS12] C. Allouch, P. Sablonnière, and D. Sbibih. Collocation methods for solving multivariable integral equations of the second kind. Journal of Computational and Applied Mathematics, 236(17):4494–4512, November 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712001884.

Allouch:2010:PIM

[ASST10] C. Allouch, P. Sablonnière, D. Sbibih, and M. Tahrichi. Product integration methods based on discrete spline quasi-interpolants and application to weakly singular integral equations. Journal of Computational and Applied Mathematics, 233(11):2855–2866, April 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704270900778X.

Allouch:2014:SND

[AST14] C. Allouch, D. Sbibih, and M. Tahrichi. Superconvergent Nyström and degenerate kernel methods for Hammerstein integral equations. Journal of Computational and Applied Mathematics, 258(??):30– 41, March 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713004305.

Alici:2010:PMS

[AT10] H. Alici and H. Taseli. Pseudospectral methods for solving an equation of hypergeometric type with a perturbation. Journal of Computational and Applied Mathematics, 234(4):1140–1152, June 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709003525.

Aydin:2014:DSM

[ATS14] S. Han Aydin and M. Tezer-Sezgin. A DRBEM solution for MHD pipe flow in a conducting medium. Journal of Computational and Applied Mathematics, 259 (part B)(??):720–729, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713002732.

Aksoylu:2014:RPH

[AU14a] Burak Aksoylu and Zuhal Unlu. Robust preconditioners for the highcontrast Stokes equation.

Journal of Computational and Applied Mathematics, 259 (part B)(??):944–954, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005608.

Altintan:2014:SIB

[AU14b] D. Altintan and Ö. Ugur.

Solution of initial and boundary value problems by the variational iteration method. Journal of Computational and Applied Mathematics, 259 (part B)(??): 790–797, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042713003567.

Aziz:2013:NAN

[AuI13] Imran Aziz and Siraj ul Islam. New algorithms for the numerical solution of nonlinear Fredholm and Volterra integral equations using Haar wavelets. Journal of Computational and Applied Mathematics, 239(??):333–345, February 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003536.

Aziz:2014:NMB

[AuIK14] Imran Aziz, Siraj ul Islam, and Fawad Khan. A new method based on Haar wavelet for the numerical solution of two-dimensional nonlinear integral equations. Journal of Computational and Applied Mathematics, 272(??):70–80, December 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042714002398.

Acosta:2010:FDG

[AV10a] Sebastian Acosta and Vianey Villamizar. Finite difference on grids with nearly uniform cell area and line spacing for the wave equation on complex domains. Journal of Computational and Applied Mathematics, 234(6):1970– 1979, July 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042709005263.

Antunes:2010:MNM

[AV10b] Pedro R. S. Antunes and Svilen S. Valtchev. A meshfree numerical method for acoustic wave propagation problems in planar domains with corners and cracks. Journal of Computational and Applied Mathematics, 234(9):2646–2662, September 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271000035X.

Arregui:2012:NSO

[AV12] Iñigo Arregui and Carlos Vázquez. Numerical solution of an optimal investment problem with proportional transaction costs. Journal of Computational and Applied Mathematics, 236(12):2923– 2937, June 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712000404.

Alonso:2013:SPC

[AVA13] Pedro Alonso and Jesús Vigo-Aguiar. Selected papers on Computational and Mathematical Methods in Science and Engineering (CMMSE). Journal of Computational and Applied Mathematics, 252(??):1–2, November 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713001283.

Avery:2010:HPH

[Ave10] John Scales Avery. Harmonic polynomials, hyperspherical harmonics, and atomic spectra. Journal of Computational and Applied Mathematics, 233(6):1366– 1379, January 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709001411.

Alvarez-Vazquez:2010:POS

[AVGCMVM10] L. J. Alvarez-Vázquez, N. García-Chan, A. Martínez, and M. E. Vázquez-Méndez. Pareto-optimal solutions for a wastewater treatment problem. Journal of Computational and Applied Mathematics, 234(7):2193– 2201, August 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704270900555X.

Alvarez-Vazquez:2010:FRW

[AVMVMV10] L. J. Alvarez-Vázquez, A. Martínez, M. E. Vázquez-Méndez, and M. A. Vilar. Flow regulation for water quality restoration in a river section: Modeling and control. Journal of Computational and Applied Mathematics, 234(4):1267–1276, June 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709006700.

Agarwal:2010:PCL

[AW10] Ravi P. Agarwal and Patricia J. Y. Wong. Piecewise complementary Lidstone interpolation and error inequalities. Journal of Computational and Applied Mathematics, 234(8):2543– 2561, August 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710001871.

An:2011:FDA

[AWF11] Heng-Bin An, Ju Wen, and Tao Feng. On finite difference approximation of a matrix–vector product in the Jacobian-free Newton–Krylov method. Journal of Computational and Applied Mathematics, 236(6):1399–1409, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004730.

Arandiga:2013:GWD

[AY13] Francesc Aràndiga and Dionisio F. Yáñez. Generalized wavelets design using kernel methods. Application to signal processing. Journal of Computational and Applied Mathematics, 250(??):1–15, October 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713000885.

Aydogan:2014:SRS

[Ayd14] Melike Aydoğan. Some results on a starlike logharmonic mapping of order alpha. Journal of Computational and Applied Mathematics, 256(??):77–82, January 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271300352X.

Akbari:2011:RBP

[AZ11] Reza Akbari and Koorush Ziarati. A rank based particle swarm optimization algorithm with dynamic adaptation. Journal of Computational and Applied Mathematics, 235 (8):2694–2714, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042710006382.

Bayrak:2010:EPM

[BA10] Özlem Türker Bayrak and Aysen D. Akkaya. Estimating parameters of a multiple autoregressive model by the modified maximum likelihood method. Journal of Computational and Applied Mathematics, 233(8):1763–1772, February 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709006359.

Behnia:2011:NDM

[BAAS11] S. Behnia, A. Akhavan, A. Akhshani, and A. Samsudin. A novel dynamic model of pseudo random number generator. Journal of Computational and Applied Mathematics, 235(12):3455–3463, April 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711000793.

Baboolal:2010:HRN

[Bab10] S. Baboolal. High-resolution numerical simulation of 2D nonlinear wave structures in electromagnetic fluids with absorbing boundary conditions. Journal of Computational and Applied Mathematics, 234(6):1710–1716, July 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709004944.

Blazquez:2010:CHP

[BAdS^{+}10] M. Teresa Blázquez, Marta Anguiano, Fernando Arias de Saavedra, Antonio M. Lallena, and Pedro Carpena. Characterizing the human postural control system using detrended fluctuation analysis. Journal of Computational and Applied Mathematics, 233(6):1478–1482, January 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709001502.

See corrigendum [BAdS^{+}12].

Blazquez:2012:CCH

[BAdS^{+}12] M. Teresa Blázquez, Marta Anguiano, Fernando Arias de Saavedra, Antonio M. Lallena, and Pedro Carpena. Corrigendum to “Characterizing the human postural control system using detrended fluctuation analysis” [J. Comp. Appl. Math. 233 (2010) 1478– 1482]. Journal of Computational and Applied Mathematics, 236(9):2605, March 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711005759.

See [BAdS^{+}10].

Bai:2013:EES

[Bai13]  Zhong-Zhi Bai.  Eigen-

value estimates for saddle point matrices of Hermitian and indefinite leading blocks. Journal of Computational and Applied Mathematics, 237(1):295– 306, January 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712002087.

Baker:2010:DAH

[Bak10] Christopher T. H. Baker.

Development and application of halanay-type theory: Evolutionary differential and difference equations with time lag. Journal of Computational and Applied Mathematics, 234(9):2663–2682, September 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710000312.

Bairamov:2011:SED

[BAK11] Elgiz Bairamov, Yelda Aygar, and Turhan Koprubasi. The spectrum of eigenparameter-dependent discrete Sturm–Liouville equations. Journal of Computational and Applied Mathematics, 235(16):4519– 4523, June 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709008541.

Ballestra:2014:RSE

[Bal14] Luca Vincenzo Ballestra. Repeated spatial extrapolation: an extraordinarily efficient approach for option pricing. Journal of Computational and Applied Mathematics, 256(??):83–91, January 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713003774.

Bartoszewski:2011:SBV

[Bar11]  Z. Bartoszewski.  Solv-

ing boundary value problems for delay differential equations by a fixed-point method. Journal of Computational and Applied Mathematics, 236(6):1576–1590, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711005000.

Baratella:2013:NIS

[Bar13] Paola Baratella. A Nyström interpolant for some weakly singular nonlinear Volterra integral equations. Journal of Computational and Applied Mathematics, 237(1): 542–555, January 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0377042712002592.

Balci:2013:MML

[BAU13] Sibel Balci, Aysen D.

Akkaya, and B. Emre Ulgen. Modified maximum likelihood estimators using ranked set sampling. Journal of Computational and Applied Mathematics, 238(??):171–179, January 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003524.

Baysal:2014:BMS

[Bay14]  Onur Baysal.  Bubble

and multiscale stabilization of bilinear finite element methods for transient advection–diffusion equations on rectangular grids. Journal of Computational and Applied Mathematics, 259 (part B)(??):641–650, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271300143X.

Bourchtein:2010:EBE

[BB10a] Andrei Bourchtein and Ludmila Bourchtein. On ellipticity of balance equations for atmospheric dynamics. Journal of Computational and Applied Mathematics, 234(4):1017–1026, June 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709002155.

Brezzi:2010:IMD

[BB10b] Franco Brezzi and Annalisa Buffa. Innovative mimetic discretizations for electromagnetic problems. Journal of Computational and Applied Mathematics, 234(6):1980–1987, July 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709005275.

Baxter:2011:FEB

[BB11a] B. J. C. Baxter and R. Brummelhuis. Functionals of exponential Brownian motion and divided differences. Journal of Computational and Applied Mathematics, 236(4):424–433, September 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042711003335.

Benitez:2011:SOC

[BB11b] M. Benítez and A. Bermúdez. A second order characteristics finite element scheme for natural convection problems. Journal of Computational and Applied Mathematics, 235(11):3270–3284, April 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711000082.

Benitez:2011:EUN

[BB11c] R. Benítez and V. J. Bolós.

Existence and uniqueness of nontrivial collocation solutions of implicitly linear homogeneous Volterra integral equations. Journal of Computational and Applied Mathematics, 235(12):3661– 3672, April 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711001142.

Bourchtein:2012:EFD

[BB12a] Andrei Bourchtein and Ludmila Bourchtein. Explicit finite difference schemes with extended stability for advection equations. Journal of Computational and Applied Mathematics, 236(15):3591– 3604, September 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711002226.

Bourchtein:2012:SPV

[BB12b] Andrei Bourchtein and Ludmila Bourchtein. Some properties of the vertical structure operators for hydrostatic atmospheric models. Journal of Computational and Applied Mathematics, 236(15):3624–3635, September 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711002949.

Burrage:2012:LRR

[BB12c] Kevin Burrage and Pamela M. Burrage. Low rank Runge– Kutta methods, symplecticity and stochastic Hamiltonian problems with additive noise. Journal of Computational and Applied Mathematics, 236(16):3920–3930, October 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712001240.

Bouchot:2014:DNA

[BB14a] Jean-Luc Bouchot and Frank Bauer. Discrepancy norm: Approximation and variations. Journal of Computational and Applied Mathematics, 272(??):162–179, December 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714002489.

Bourchtein:2014:BCG

[BB14b] Andrei Bourchtein and Ludmila Bourchtein. On boundary conditions for the gravity wave equations. Journal of Computational and Applied Mathematics, 259 (part A)(??):312–326, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003676.

Balbas:2010:MMR

[BBB10] Alejandro Balbás, Beatriz Balbás, and Raquel Balbás. Minimizing measures of risk by saddle point conditions. Journal of Computational and Applied Mathematics, 234(10):2924–2931, September 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710001998.

Belochitski:2011:TAL

[BBD^{+}11] Alexei Belochitski, Peter Binev, Ronald DeVore, Michael Fox-Rabinovitz, Vladimir Krasnopolsky, and Philipp Lamby. Tree approximation of the long wave radiation parameterization in the NCAR CAM global climate model. Journal of Computational and Applied Mathematics, 236(4):447–460, September 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711003992.

Belley:2013:MPI

[BBD13] J.-M. Belley, P. E. Belley, and A. G. R. Day. Multivariate polynomial interpolation and meshfree differentiation via undetermined coefficients. Journal of Computational and Applied Mathematics, 239(??):415– 430, February 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712003767.

Bastl:2014:HIS

[BBK^{+}14] Bohumír Bastl, Michal Bizzarri, Marjeta Krajnc, Miroslav Lávička, Kristýna Slabá, Zbyněk Šír, Vito Vitrih, and Emil Žagar. C^{1} Hermite interpolation with spatial Pythagorean-hodograph cubic biarcs. Journal of Computational and Applied Mathematics, 257(??):65– 78, February 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713004123.

Bailey:2010:RDR

[BBKS10] David Bailey, Markus Berndt, Milan Kucharik, and Mikhail Shashkov. Reduced-dissipation remapping of velocity in staggered arbitrary Lagrangian– Eulerian methods. Journal of Computational and Applied Mathematics, 233(12): 3148–3156, April 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270900630X.

Berry:2010:P

[BBMSC10] Michael W. Berry, Dario A.

Bini, Nicola Mastronardi, and Stefano Serra-Capizzano. Preface. Journal of Computational and Applied Mathematics, 234(11):3073–3074, October 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710000816.

Bao:2011:CRN

[BBMY11] Jianhai Bao, Björn Böttcher, Xuerong Mao, and Chenggui Yuan. Convergence rate of numerical solutions to SFDEs with jumps. Journal of Computational and Applied Mathematics, 236(2):119–131, August 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711003153.

Bader:2014:SPI

[BBP14] Philipp Bader, Sergio Blanes, and Enrique Ponsoda. Structure preserving integrators for solving (non-)linear quadratic optimal control problems with applications to describe the flight of a quadrotor. Journal of Computational and Applied Mathematics, 262(??):223– 233, May 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713005165.

Bartel:2014:CRD

[BBS14] Andreas Bartel, Markus Brunk, and Sebastian Schöps. On the convergence rate of dynamic iteration for coupled problems with multiple subsystems. Journal of Computational and Applied Mathematics, 262(??):14– 24, May 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713003750.

Berriochoa:2010:ASH

[BC10] Elías Berriochoa and Alicia Cachafeiro. Algorithms for solving Hermite interpolation problems using the fast Fourier transform. Journal of Computational and Applied Mathematics, 235(4):882–894, December 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709004300.

Bolotnikov:2012:NPP

[BC12] Vladimir Bolotnikov and Stephen P. Cameron. The Nevanlinna–Pick problem on the closed unit disk: Minimal norm rational solutions of low degree. Journal of Computational and Applied Mathematics, 236(13): 3123–3136, July 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712000660.

Bica:2012:ANM

[BCC12] A. M. Bica, M. Curila, and S. Curila. About a numerical method of successive interpolations for functional Hammerstein integral equations. Journal of Computational and Applied Mathematics, 236(7):2005–2024, January 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711005802.

Brockman:2013:HME

[BCC^{+}13] Philip Brockman, Timothy Carson, Yun Cheng, T. M. Elgindi, K. Jensen, X. Zhoun, and M. B. M. Elgindi. Homotopy method for the eigenvalues of symmetric tridiagonal matrices. Journal of Computational and Applied Mathematics, 237(1):644–653, January 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712003329.

Beugre:2010:LBF

[BCD^{+}10] Djomice Beugre, Sbastien Calvo, Grard Dethier, Michel Crine, Dominique Toye, and Pierre Marchot. Lattice Boltzmann 3D flow simulations on a metallic foam. Journal of Computational and Applied Mathematics, 234(7):2128–2134, August 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709005482.

Bracciali:2010:SAS

[BCGMB10] Cleonice F. Bracciali, Laura Castaño-García, and Juan J. Moreno-Balcázar. Some asymptotics for Sobolev orthogonal polynomials involving Gegenbauer weights. Journal of Computational and Applied Mathematics, 235(4):904–915, December 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710003043.

Banicescu:2011:IAP

[BCHL11] Ioana Banicescu, Ricolindo L. Cariño, Jane L. Harvill, and John Patrick Lestrade. Investigating asymptotic properties of vector nonlinear time series models. Journal of Computational and Applied Mathematics, 236(3):411–421, September 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004134.

Balima:2010:CLT

[BCM10] O. Balima, A. Charette, and D. Marceau. Comparison of light transport models based on finite element and the discrete ordinates methods in view of optical tomography applications. Journal of Computational and Applied Mathematics, 234(7):2259–2271, August 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709005627.

Brugnano:2012:EPM

[BCMR12] L. Brugnano, M. Calvo, J. I. Montijano, and L. Rández. Energy-preserving methods for Poisson systems. Journal of Computational and Applied Mathematics, 236(16):3890–3904, October 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712001008.

Beccari:2011:PBN

[BCR11] Carolina Beccari, Giulio Casciola, and Lucia Romani. Polynomial-based non-uniform interpolatory subdivision with features control. Journal of Computational and Applied Mathematics, 235(16):4754–4769, June 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710005066.

Beccari:2013:CCN

[BCR13] Carolina Vittoria Beccari, Giulio Casciola, and Lucia Romani. Construction and characterization of nonuniform local interpolating polynomial splines. Journal of Computational and Applied Mathematics, 240(??): 5–19, March 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712002609.

Babajee:2010:ATC

[BDD^{+}10] D. K. R. Babajee, M. Z.

Dauhoo, M. T. Darvishi, A. Karami, and A. Barati. Analysis of two Chebyshevlike third order methods free from second derivatives for solving systems of nonlinear equations. Journal of Computational and Applied Mathematics, 233(8):2002–2012, February 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709006591.

Batselier:2014:FRO

[BDD14] Kim Batselier, Philippe Dreesen, and Bart De Moor. A fast recursive orthogonalization scheme for the Macaulay matrix. Journal of Computational and Applied Mathematics, 267(??):20– 32, September 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042714000636.

Bracciali:2010:NSS

[BDF^{+}10] Cleonice F. Bracciali, Antonia M. Delgado, Lidia Fernández, Teresa E. Pérez, and Miguel A. Piñar. New steps on Sobolev orthogonality in two variables. Journal of Computational and Applied Mathematics, 235(4):916–926, December 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710003924.

Bos:2011:LCB

[BDH11] Len Bos, Stefano De Marchi, and Kai Hormann. On the Lebesgue constant of Berrut’s rational interpolant at equidistant nodes. Journal of Computational and Applied Mathematics, 236(4):504–510, September 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711001816.

Balazs:2011:TIA

[BDJ^{+}11] P. Balazs, M. Dörfler, F. Jaillet, N. Holighaus, and G. Velasco. Theory, implementation and applications of nonstationary Gabor frames. Journal of Computational and Applied Mathematics, 236(6):1481–1496, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004900.

Baronian:2010:TBC

[BDL10] V. Baronian, A. S. Bonnet-Ben Dhia, and E. Lunéville. Transparent boundary conditions for the harmonic diffraction problem in an elastic waveguide. Journal of Computational and Applied Mathematics, 234(6): 1945–1952, July 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042709005238.

Binev:2011:FHD

[BDL11] Peter Binev, Wolfgang Dahmen, and Philipp Lamby. Fast high-dimensional approximation with sparse occupancy trees. Journal of Computational and Applied Mathematics, 235(8):2063– 2076, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042710005613.

Bini:2014:ISI

[BDM^{+}14] Dario Bini, Lieven De Lathauwer, Nicola Mastronardi, Marc Van Barel, Raf Vandebril, and Paul Van Dooren. Introduction to the special issue. Journal of Computational and Applied Mathematics, 272(??):275, December 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042714001307.

Blal:2013:MBC

[BDMP13] Nawfal Blal, Loïc Daridon, Yann Monerie, and Stéphane Pagano. Micromechanicalbased criteria for the calibration of cohesive zone parameters. Journal of Computational and Applied Mathematics, 246(??):206– 214, July 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712004712.

Blanes:2010:SCM

[BDMR10] Sergio Blanes, Fasma Diele, Carmela Marangi, and Stefania Ragni. Splitting and composition methods for explicit time dependence in separable dynamical systems. Journal of Computational and Applied Mathematics, 235(3):646–659, December 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710003705.

Bini:2010:CAE

[BDR10] Dario A. Bini, Gianna M.

Del Corso, and F. Romani. A combined approach for evaluating papers, authors and scientific journals. Journal of Computational and Applied Mathematics, 234(11):3104–3121, October 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710000749.

Bozzini:2011:CGQ

[BDR11] Mira Bozzini, Nira Dyn, and Milvia Rossini. Construction of generators of quasiinterpolation operators of high approximation orders in spaces of polyharmonic splines. Journal of Computational and Applied Mathematics, 236(4):557–564, September 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711003888.

Barucq:2010:TDA

[BDSG10] H. Barucq, R. Djellouli, and A. Saint-Guirons. Threedimensional approximate local DtN boundary conditions for prolate spheroid boundaries. Journal of Computational and Applied Mathematics, 234(6):1810– 1816, July 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S037704270900507X.

Bols:2013:CMA

[BDT^{+}13] J. Bols, J. Degroote, B. Trachet, B. Verhegghe, P. Segers, and J. Vierendeels. A computational method to assess the in vivo stresses and unloaded configuration of patient-specific blood vessels. Journal of Computational and Applied Mathematics, 246(??):10– 17, July 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712004815.

Beraldi:2010:GST

[BDV10] Patrizia Beraldi, Francesco De Simone, and Antonio Violi. Generating scenario trees: a parallel integrated simulation-optimization approach. Journal of Computational and Applied Mathematics, 233(9):2322–2331, March 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709006955.

Brenner:2014:PUM

[BDyS14] Susanne C. Brenner, Christopher B. Davis, and Li yeng Sung. A partition of unity method for the displacement obstacle problem of clamped Kirchhoff plates. Journal of Computational and Applied Mathematics, 265(??):3–16, August 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713004883.

Buric:2011:BPA

[BE11] Tomislav Burić and Neven Elezović. Bernoulli polynomials and asymptotic expansions of the quotient of gamma functions. Journal of Computational and Applied Mathematics, 235(11):3315–3331, April 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711000562.

Behiry:2014:RSN

[Beh14]  S. H. Behiry.  RE-

TRACTED: Solution of nonlinear Fredholm integrodifferential equations using a hybrid of block pulse functions and normalized Bernstein polynomials. Journal of Computational and Applied Mathematics, 260(??): 258–265, April 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042713004913.

Bouhamidi:2014:CGT

[BEJ14] A. Bouhamidi, R. Enkhbat, and K. Jbilou. Conditional gradient Tikhonov method for a convex optimization problem in image restoration. Journal of Computational and Applied Mathematics, 255(??):580–592, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713003099.

Belhaj:2013:CPR

[Bel13] Skander Belhaj. Computing the polynomial remainder sequence via Bézout matrices. Journal of Computational and Applied Mathematics, 250(??):244–255, October 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713001167.

Bruvoll:2010:TMV

[BF10] Solveig Bruvoll and Michael S. Floater. Transfinite mean value interpolation in general dimension. Journal of Computational and Applied Mathematics, 233(7):1631– 1639, February 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709000879.

Belhamadia:2014:PAM

[BFB14] Youssef Belhamadia, André Fortin, and Yves Bourgault. On the performance of anisotropic mesh adaptation for scroll wave turbulence dynamics in reaction–diffusion systems. Journal of Computational and Applied Mathematics, 271(??):233–246, December 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714001873.

Baker:2010:P

[BFD^{+}10] Christopher T. H. Baker, Neville J. Ford, Teresa Diogo, Pedro Lima, and Adélia Sequeira. Preface. Journal of Computational and Applied Mathematics, 234(9):2617, September 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710000221.

Barucq:2013:ULP

[BFDP13] H. Barucq, T. Chaumont Frelet, J. Diaz, and V. Péron. Upscaling for the Laplace problem using a discontinuous Galerkin method. Journal of Computational and Applied Mathematics, 240(??):192–203, March 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271200235X.

Bindel:2014:NCB

[BFG^{+}14] D. Bindel, M. Friedman, W. Govaerts, J. Hughes, and Yu. A. Kuznetsov. Numerical computation of bifurcations in large equilibrium systems in Matlab. Journal of Computational and Applied Mathematics, 261(??):232–248, May 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005785.

Barrera:2010:FPH

[BFGP10] D. Barrera, M. A. Fortes, P. González, and M. Pasadas. Filling polygonal holes with minimal energy surfaces on Powell–Sabin type triangulations. Journal of Computational and Applied Mathematics, 234(4):1058–1068, June 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709002672.

Benkhaldoun:2012:SPI

[BFHK12] Fayssal Benkhaldoun, Jaroslav Fort, Khaled Hassouni, and Jan Karel. Simulation of planar ionization wave front propagation on an unstructured adaptive grid. Journal of Computational and Applied Mathematics, 236(18):4623–4634, December 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712001707.

Bazan:2012:MPC

[BFLP12] Fermín S. V. Bazán, J. B.

Francisco, Koung Hee Leem, and G. Pelekanos. A maximum product criterion as a Tikhonov parameter choice rule for Kirsch’s factorization method. Journal of Computational and Applied Mathematics, 236(17): 4264–4275, November 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042712002099.

Butler:2011:FSA

[BFMS11] Jon T. Butler, C. L.

Frenzen, Njuguna Macaria, and Tsutomu Sasao. A fast segmentation algorithm for piecewise polynomial numeric function generators. Journal of Computational and Applied Mathematics, 235(14):4076–4082, May 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271100121X.

Boulaajine:2011:PEE

[BFP11] L. Boulaajine, M. Farhloul, and L. Paquet. A priori error estimation for the dual mixed finite element method of the elastodynamic problem in a polygonal domain, II. Journal of Computational and Applied Mathematics, 235(5):1288– 1310, January 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710004528.

Bocian:2014:NMA

[BFS14] Rafal Bocian, Mariusz Felisiak, and Daniel Simson. Numeric and mesh algorithms for the Coxeter spectral study of positive edge-bipartite graphs and their isotropy groups. Journal of Computational and Applied Mathematics, 259 (part B)(??):815–827, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003579.

Babych:2010:LHF

[BG10a] Natalia Babych and Yuri Golovaty. Low and high frequency approximations to eigenvibrations of string with double contrasts. Journal of Computational and Applied Mathematics, 234 (6):1860–1867, July 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042709005135.

Bourdarias:2010:KSU

[BG10b] C. Bourdarias and S. Gerbi.

A kinetic scheme for unsteady pressurised flows in closed water pipes. Journal of Computational and Applied Mathematics, 234(7):2098–2105, August 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709005457.

Banjai:2011:ELT

[BG11] Lehel Banjai and Volker Gruhne. Efficient longtime computations of timedomain boundary integrals for 2D and dissipative wave equation. Journal of Computational and Applied Mathematics, 235(14):4207– 4220, May 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042711001452.

Barrero-Gil:2013:WAD

[BG13] A. Barrero-Gil. Weakening accuracy dependence with the regularization parameter in the Method of Regularized Stokeslets. Journal of Computational and Applied Mathematics, 237(1):672–679, January 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712003366.

Bi:2014:GSP

[BG14a] Chunjia Bi and Victor Ginting. Global superconvergence and a posteriori error estimates of the finite element method for second-order quasilinear elliptic problems. Journal of Computational and Applied Mathematics, 260(??): 78–90, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713004974.

Brugiapaglia:2014:SRZ

[BG14b] Simone Brugiapaglia and Luca Gemignani. On the simultaneous refinement of the zeros of Hpalindromic polynomials.

Journal of Computational and Applied Mathematics, 272(??):293–303, December 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713002859.

Berenguer:2011:BSS

[BGGG^{+}11] M. I. Berenguer, D. Gámez, A. I. Garralda-Guillem, M. Ruiz Galán, and M. C. Serrano Pérez. Biorthogonal systems for solving Volterra integral equation systems of the second kind. Journal of Computational and Applied Mathematics, 235(7):1875– 1883, February 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710004085.

Barrera:2010:OBC

[BGIN10] D. Barrera, A. Guessab, M. J. Ibáñez, and O. Nouisser. Optimal bivariate C^{1} cubic quasi-interpolation on a type-2 triangulation. Journal of Computational and Applied Mathematics, 234 (4):1188–1199, June 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042709004397.

Barrera:2013:IAO

[BGIN13] D. Barrera, A. Guessab, M. J. Ibáñez, and O. Nouisser. Increasing the approximation order of spline quasiinterpolants. Journal of Computational and Applied Mathematics, 252(??):27– 39, November 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S037704271300054X.

Bunse-Gerstner:2010:NOM

[BGKVW10] A. Bunse-Gerstner, D. Kubalińska, G. Vossen, and D. Wilczek. h_{2}-norm optimal model reduction for large scale discrete dynamical MIMO systems. Journal of Computational and Applied Mathematics, 233(5):1202–1216, January 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709001824.

Berenguer:2013:FPT

[BGL13] M. I. Berenguer, D. Gámez, and A. J. López Linares. Fixed point techniques and Schauder bases to approximate the solution of the first order nonlinear mixed Fredholm– Volterra integro-differential equation. Journal of Computational and Applied Mathematics, 252(??):52– 61, November 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712003846.

Boal:2011:FDA

[BGLV11] N. Boal, F. J. Gaspar, F. J. Lisbona, and P. N. Vabishchevich. Finitedifference analysis of fully dynamic problems for saturated porous media. Journal of Computational and Applied Mathematics, 236 (6):1090–1102, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042711004365.

Bottcher:2010:IEC

[BGM10] A. Böttcher, S. M. Grudsky, and E. A. Maksimenko. Inside the eigenvalues of certain Hermitian Toeplitz band matrices. Journal of Computational and Applied Mathematics, 233(9):2245– 2264, March 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709006888.

Bouche:2011:EEU

[BGP11] Daniel Bouche, Jean-Michel Ghidaglia, and Frédéric P.

Pascal. Error estimate for the upwind finite volume method for the nonlinear scalar conservation law. Journal of Computational and Applied Mathematics, 235(18):5394–5410, July 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711003220.

Bush:2014:ACG

[BGP14] Lawrence Bush, Victor Ginting, and Michael Presho. Application of a conservative, generalized multiscale finite element method to flow models. Journal of Computational and Applied Mathematics, 260(??):395– 409, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005505.

Beck:2014:CAK

[BGPPRW14] S. Beck, S. González-Pinto, S. Pérez-Rodríguez, and R. Weiner. A comparison of AMF- and Krylov-methods in Matlab for large stiff ODE systems. Journal of Computational and Applied Mathematics, 262(??):292– 303, May 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005153.

Bayon:2012:QLA

[BGRS12] L. Bayón, J. M. Grau, M. M. Ruiz, and P. M. Suárez. A quasi-linear algorithm for calculating the infimal convolution of convex quadratic functions. Journal of Computational and Applied Mathematics, 236(12):2990–2997, June 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711001889.

Berwald:2012:UML

[BGS12] Jesse Berwald, Tomás Gedeon, and John Sheppard. Using machine learning to predict catastrophes in dynamical systems. Journal of Computational and Applied Mathematics, 236(9):2235–2245, March 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711005760.

Bultheel:2013:QAP

[BGVHN13] Adhemar Bultheel, Pablo González-Vera, Erik Hendriksen, and Olav Njåstad. Quadratures associated with pseudo-orthogonal rational functions on the real half line with poles in [−∞, 0]. Journal of Computational and Applied Mathematics, 237(1):589–602, January 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712002798.

Babolian:2011:ASC

[BH11] E. Babolian and A. Arzhang Hajikandi. The approximate solution of a class of Fredholm integral equations with a weakly singular kernel. Journal of Computational and Applied Mathematics, 235(5):1148–1159, January 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710004231.

Bursi:2010:NGM

[BHBP10] O. S. Bursi, L. He, A. Bonelli, and P. Pegon. Novel generalized-α methods for interfield parallel integration of heterogeneous structural dynamic systems. Journal of Computational and Applied Mathematics, 234(7):2250–2258, August 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709005615.

Bouhamidi:2011:BAB

[BHJ11] A. Bouhamidi, M. Heyouni, and K. Jbilou. Block Arnoldi-based methods for large scale discrete-time algebraic Riccati equations. Journal of Computational and Applied Mathematics, 236(6):1531–1542, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004961.

Biak:2013:SAF

[BHJ13] Martin Biák, Tomás Hanus, and Drahoslava Janovská. Some applications of Filippov’s dynamical systems. Journal of Computational and Applied Mathematics, 254(??):132–143, December 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713001428.

Bonan-Hamada:2010:ZOL

[BHJN10] C. Bonan-Hamada, W. B.

Jones, and O. Njåstad. Zeros of orthogonal Laurent polynomials and solutions of strong Stieltjes moment problems. Journal of Computational and Applied Mathematics, 235(4):895– 903, December 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042710003675.

Biglari:2011:NQN

[BHL11] Fahimeh Biglari, Malik Abu Hassan, and Wah June Leong. New quasi-Newton methods via higher order tensor models. Journal of Computational and Applied Mathematics, 235(8):2412– 2422, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042710006072.

Bradshaw-Hajek:2013:CLS

[BHMW13] B. H. Bradshaw-Hajek, S. J. Miklavcic, and D. A. Ward. A composite Level Set and Extended-Domain-Eigenfunction Method for simulating 2D Stokes flow involving a free surface. Journal of Computational and Applied Mathematics, 237(1):389–402, January 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712002440.

Bungartz:2012:OPD

[BHPS12] Hans-Joachim Bungartz, Alexander Heinecke, Dirk Pflüger, and Stefanie Schraufstetter. Option pricing with a direct adaptive sparse grid approach. Journal of Computational and Applied Mathematics, 236(15):3741– 3750, September 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711005036.

Brugnano:2012:LIM

[BI12] Luigi Brugnano and Felice Iavernaro. Line integral methods which preserve all invariants of conservative problems. Journal of Computational and Applied Mathematics, 236(16):3905– 3919, October 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712001549.

Birkholz:2011:UAI

[Bir11] Harald Birkholz. A unifying approach to isotropic and anisotropic total variation denoising models. Journal of Computational and Applied Mathematics, 235(8):2502–2514, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006199.

Barrera:2010:NBD

[BISS10] D. Barrera, M. J. Ibáñez, P. Sablonnière, and D. Sbibih. On near-best discrete quasiinterpolation on a fourdirectional mesh. Journal of Computational and Applied Mathematics, 233(6):1470– 1477, January 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042709001496.

Brugnano:2011:NEI

[BIT11] Luigi Brugnano, Felice Iavernaro, and Donato Trigiante. A note on the efficient implementation of Hamiltonian BVMs. Journal of Computational and Applied Mathematics, 236(3):375–383, September 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711004171.

Bojovic:2012:CFD

[BJ12] Dejan R. Bojović and Bosko S. Jovanović. Convergence of a finite difference method for solving 2D parabolic interface problems. Journal of Computational and Applied Mathematics, 236(15):3605–3612, September 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711002366.

Buric:2013:TFM

[BJ13] Lubor Buric and Vladimír Janovský. A traffic flow model with overtaking as a Filippov system. Journal of Computational and Applied Mathematics, 254(??):55– 64, December 15, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713001416.

Bras:2014:SEG

[BJ14] M. Braś and Z. Jackiewicz.

Search for efficient general linear methods for ordinary differential equations. Journal of Computational and Applied Mathematics, 262(??):180–192, May 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003762.

Jian:2014:SCS

[bJfCwH14] Jin bao Jian, Qiao fang Chen, and Zong wen Huang. A superlinearly convergent SQP method without boundedness assumptions on any of the iterative sequences. Journal of Computational and Applied Mathematics, 263(??):115– 128, June 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713006717.

Bouhamidi:2012:IPC

[BJR^{+}12] Abderrahman Bouhamidi, Khalide Jbilou, Ronny Ramlau, Lothar Reichel, Hassane Sadok, and Fiorella Sgallari. Inverse problems: Computation and applications. Journal of Computational and Applied Mathematics, 236(8):2047–2048, February 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711005085.

Bouhamidi:2012:GGA

[BJRS12] A. Bouhamidi, K. Jbilou, L. Reichel, and H. Sadok. A generalized global Arnoldi method for ill-posed matrix equations. Journal of Computational and Applied Mathematics, 236(8):2078– 2089, February 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042711005103.

Baxley:2010:EMP

[BK10] John V. Baxley and Kristen E. Kobylus. Existence of multiple positive solutions of singular nonlinear boundary value problems. Journal of Computational and Applied Mathematics, 234(9):2699–2708, September 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710000336.

Braverman:2012:GAS

[BK12] E. Braverman and B. Karpuz.

On global asymptotic stability of nonlinear higher-order difference equations. Journal of Computational and Applied Mathematics, 236 (11):2803–2812, May 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704271200026X.

Bayramoglu:2013:SND

[BK13] Ismihan Bayramoğlu (Bairamov)

and Gülder Kemalbay. Some novel discrete distributions under fourfold sampling schemes and conditional bivariate order statistics. Journal of Computational and Applied Mathematics, 248(??):1–14, August 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713000435.

Babaie-Kafaki:2014:TMS

[BK14a] Saman Babaie-Kafaki. Two modified scaled nonlinear conjugate gradient methods. Journal of Computational and Applied Mathematics, 261(??):172–182, May 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713006183.

Basting:2014:OCM

[BK14b] Christopher Basting and Dmitri Kuzmin. Optimal control for mass conservative level set methods. Journal of Computational and Applied Mathematics, 270(??):343–352, November 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713007139.

Berrut:2014:RAL

[BK14c] Jean-Paul Berrut and Georges Klein. Recent advances in linear barycentric rational interpolation. Journal of Computational and Applied Mathematics, 259 (part A)(??):95–107, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042713001714.

Bittl:2014:CDM

[BKB14] Melanie Bittl, Dmitri Kuzmin, and Roland Becker. The CG1–DG2 method for convection–diffusion equations in 2D. Journal of Computational and Applied Mathematics, 270(??):21– 31, November 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042714001484.

Babaie-Kafaki:2010:TNC

[BKGMA10] Saman Babaie-Kafaki, Reza Ghanbari, and Nezam Mahdavi-Amiri. Two new conjugate gradient methods based on modified secant equations. Journal of Computational and Applied Mathematics, 234(5):1374– 1386, July 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710000658.

Bokil:2014:OSM

[BKL14] V. A. Bokil, O. A. Keefer, and A. C.-Y. Leung. Operator splitting methods for Maxwell’s equations in dispersive media with orientational polarization. Journal of Computational and Applied Mathematics, 263(??): 160–188, June 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S037704271300678X.

Bozkaya:2011:SOC

[BKMG11] Canan Bozkaya, Serpil Kocabiyik, L. A. Mironova, and O. I. Gubanov. Streamwise oscillations of a cylinder beneath a free surface: Free surface effects on vortex formation modes. Journal of Computational and Applied Mathematics, 235(16):4780–4795, June 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710005315.

Bogers:2013:MDD

[BKN^{+}13] J. Bogers, K. Kumar, P. H. L. Notten, J. F. M. Oudenhoven, and I. S. Pop. A multiscale domain decomposition approach for chemical vapor deposition. Journal of Computational and Applied Mathematics, 246 (??):65–73, July 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S037704271200444X.

Bogdanov:2014:LSH

[BKTY14] Andrey Bogdanov, Elif Bilge Kavun, Elmar Tischhauser, and Tolga Yalçin. Largescale high-resolution computational validation of novel complexity models in linear cryptanalysis. Journal of Computational and Applied Mathematics, 259 (part B)(??):592–598, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005645.

Broeckhove:2010:ANC

[BKV10] Jan Broeckhove, Przemysław Klosiewicz, and Wim Vanroose. Applying numerical continuation to the parameter dependence of solutions of the Schrödinger equation. Journal of Computational and Applied Mathematics, 234(4):1238–1248, June 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709004622.

Bergot:2010:GHO

[BL10a] M. Bergot and P. Lacoste. Generation of higherorder polynomial basis of Nédélec H (curl) finite elements for Maxwell’s equations. Journal of Computational and Applied Mathematics, 234(6):1937–1944, July 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709005226.

Borguet:2010:CAF

[BL10b] S. Borguet and O. Léonard. Comparison of adaptive filters for gas turbine performance monitoring. Journal of Computational and Applied Mathematics, 234(7):2202–2212, August 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709005561.

Bultheel:2010:SNP

[BL10c] Adhemar Bultheel and Andreas Lasarow. Schur– Nevanlinna–Potapov sequences of rational matrix functions. Journal of Computational and Applied Mathematics, 235(4):927– 949, December 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042709007936.

Bulatov:2011:TDI

[BL11] M. V. Bulatov and P. M. Lima. Two-dimensional integral-algebraic systems: Analysis and computational methods. Journal of Computational and Applied Mathematics, 236(2):132– 140, August 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711003232.

Bizzarri:2013:SNA

[BL13a] Michal Bizzarri and Miroslav Lávicka. A symbolicnumerical approach to approximate parameterizations of space curves using graphs of critical points. Journal of Computational and Applied Mathematics, 242(??):107–124, April 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712004414.

Bonamy:2013:SBI

[BL13b] C. Bonamy and C. Le Guyader. Split Bregman iteration and infinity Laplacian for image decomposition. Journal of Computational and Applied Mathematics, 240(??):99–110, March 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712002981.

Bu:2014:EPA

[BL14] Sunyoung Bu and June-Yub Lee. An enhanced parareal algorithm based on the deferred correction methods for a stiff system. Journal of Computational and Applied Mathematics, 255(??):297– 305, January 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713002550.

Bnouhachem:2010:MII

[BLKZ10] Abdellah Bnouhachem, Min Li, Mohamed Khalfaoui, and Sheng Zhaohan. A modified inexact implicit method for mixed variational inequalities. Journal of Computational and Applied Mathematics, 234 (12):3356–3365, October 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704271000244X.

Beatson:2011:BBR

[BLM11] R. K. Beatson, J. Levesley, and C. T. Mouat. Better bases for radial basis function interpolation problems. Journal of Computational and Applied Mathematics, 236(4):434–446, September 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711003669.

BenCheikh:2011:OLL

[BLO11] Y. Ben Cheikh, I. Lamiri, and A. Ouni. d-orthogonality of Little q-Laguerre type polynomials. Journal of Computational and Applied Mathematics, 236(1):74–84, August 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711001361.

Bozzini:2010:KBS

[BLR10] Mira Bozzini, Licia Lenarduzzi, and Milvia Rossini. Kernel B-splines on general lattices. Journal of Computational and Applied Mathematics, 233(7):1620–1630, February 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709000867.

Bruder:2010:LDT

[BLTW10] Andrea Bruder, Lance L.

Littlejohn, Davut Tuncer, and R. Wellman. Leftdefinite theory with applications to orthogonal polynomials. Journal of Computational and Applied Mathematics, 233(6):1380–1398, January 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709001423.

Baeumer:2010:TSL

[BM10] Boris Baeumer and Mark M. Meerschaert. Tempered stable Lévy motion and transient super-diffusion. Journal of Computational and Applied Mathematics, 233 (10):2438–2448, March 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042709007250.

Bergamaschi:2011:FBP

[BM11a] Luca Bergamaschi and Ángeles Martínez. FSAIbased parallel mixed constraint preconditioners for saddle point problems arising in geomechanics. Journal of Computational and Applied Mathematics, 236 (3):308–318, September 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042711004201.

Bouboulis:2011:RKH

[BM11b] P. Bouboulis and M. Mavroforakis. Reproducing kernel Hilbert spaces and fractal interpolation. Journal of Computational and Applied Mathematics, 235(12):3425– 3434, April 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711000768.

Brugnano:2012:YNA

[BM12] Luigi Brugnano and Francesca Mazzia. 40 years of numerical analysis: “Is the discrete world an approximation of the continuous one or is it the other way around?”. Journal of Computational and Applied Mathematics, 236(16):3855– 3856, October 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712001732.

Babolian:2011:NMS

[BMMR11] E. Babolian, K. Maleknejad, M. Mordad, and B. Rahimi. A numerical method for solving Fredholm–Volterra integral equations in twodimensional spaces using block pulse functions and an operational matrix. Journal of Computational and Applied Mathematics, 235(14):3965–3971, May 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710005947.

Boyanova:2010:RAM

[BMN10] Petia Boyanova, Svetozar Margenov, and Maya Neytcheva.

Robust AMLI methods for parabolic Crouzeix– Raviart FEM systems. Journal of Computational and Applied Mathematics, 235(2):380–390, November 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710003298.

Bacuta:2011:RRA

[BMS11] Constantin Bacuta, Brendan McCracken, and Lu Shu. Residual reduction algorithms for nonsymmetric saddle point problems. Journal of Computational and Applied Mathematics, 235(6):1614–1628, January 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710004930.

Butta:2014:SMP

[BN14] Paolo Buttà and Silvia Noschese. Structured maximal perturbations for Hamiltonian eigenvalue problems. Journal of Computational and Applied Mathematics, 272(??):304–312, December 15, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713002173.

Brezinski:2012:RAA

[BNRZ12] C. Brezinski, P. Novati, and M. Redivo-Zaglia. A rational Arnoldi approach for ill-conditioned linear systems. Journal of Computational and Applied Mathematics, 236(8):2063–2077, February 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711005115.

Biryukov:2014:DEA

[BNY14] Alex Biryukov, Jorge Nakahara, Jr., and Hamdi Murat Yildirim. Differential entropy analysis of the IDEA block cipher. Journal of Computational and Applied Mathematics, 259 (part B)(??):561–570, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003981.

Boglaev:2011:UCM

[Bog11] Igor Boglaev. Uniform convergent monotone iterates for semilinear singularly perturbed parabolic problems. Journal of Computational and Applied Mathematics, 235(12):3541–3553, April 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711000859.

Boglaev:2014:IBM

[Bog14a] Igor Boglaev. Inexact block monotone methods for solving nonlinear elliptic problems. Journal of Computational and Applied Mathematics, 269(??):109–117, October 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042714001782.

Boglaev:2014:UMA

[Bog14b]  Igor Boglaev.  A uni-

form monotone alternating direction scheme for nonlinear singularly perturbed parabolic problems. Journal of Computational and Applied Mathematics, 272(??):148–161, December 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042714002416.

Bayon:2012:PMP

[BOR^{+}12] L. Bayón, J. A. Otero, M. M. Ruiz, P. M. Suárez, and C. Tasis. The profit maximization problem in economies of scale. Journal of Computational and Applied Mathematics, 236(12): 3065–3072, June 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712000428.

Beatson:2014:FFE

[BOR14] R. K. Beatson, W. E. Ong, and I. Rychkov. Faster fast evaluation of thin plate splines in two dimensions. Journal of Computational and Applied Mathematics, 261(??):201–212, May 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713006249.

Brenner:2012:HOW

[BOS12] Susanne C. Brenner, Luke Owens, and Li-Yeng Sung. Higher order weakly overpenalized symmetric interior penalty methods. Journal of Computational and Applied Mathematics, 236 (11):2883–2894, May 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042712000362.

Boyd:2010:ESG

[Boy10] John P. Boyd. Error saturation in Gaussian radial basis functions on a finite interval. Journal of Computational and Applied Mathematics, 234(5):1435–1441, July 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710001019.

Boyd:2011:NSC

[Boy11] John P. Boyd. New series for the cosine lemniscate function and the polynomialization of the lemniscate integral. Journal of Computational and Applied Mathematics, 235(8):1941–1955, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710005352.

Boyd:2013:RCS

[Boy13] John P. Boyd. Rational Chebyshev series for the Thomas–Fermi function: Endpoint singularities and spectral methods. Journal of Computational and Applied Mathematics, 244(??): 90–101, May 15, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042712005080.

Bairamov:2011:FPC

[BP11a] Ismihan Bairamov and Safar Parsi. On flexible progressive censoring. Journal of Computational and Applied Mathematics, 235(16):4537–4544, June 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710001391.

Bairamov:2011:OSM

[BP11b] Ismihan Bairamov and Safar Parsi. Order statistics from mixed exchangeable random variables. Journal of Computational and Applied Mathematics, 235(16): 4629–4638, June 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042710002384.

Blanes:2012:MIS

[BP12] Sergio Blanes and Enrique Ponsoda. Magnus integrators for solving linear-quadratic differential games. Journal of Computational and Applied Mathematics, 236(14):3394– 3408, August 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712001252.

Bramble:2013:ACP

[BP13] James H. Bramble and Joseph E. Pasciak. Analysis of a Cartesian PML approximation to acoustic scattering problems in R^{2} and R^{3}. Journal of Computational and Applied Mathematics, 247(??):209–230, August 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713000149.

Biazar:2011:NSF

[BPGP11] Jafar Biazar, Mehdi Gholami Porshokouhi, Behzad Ghanbari, and Mohammad Gholami Porshokouhi. Numerical solution of functional integral equations by the variational iteration method. Journal of Computational and Applied Mathematics, 235(8):2581–2585, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006266.

Banas:2012:NSA

[BPS12] L’ubomír Banas, Andreas Prohl, and Marián Slodicka. Numerical scheme for augmented Landau– Lifshitz equation in heat assisted recording. Journal of Computational and Applied Mathematics, 236(18):4775– 4787, December 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712000817.

BenCheikh:2011:SCO

[BR11a] Y. Ben Cheikh and N. Ben Romdhane. On d-symmetric classical d-orthogonal polynomials. Journal of Computational and Applied Mathematics, 236(1):85–93, August 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711001579.

Buckwar:2011:RKM

[BR11b] Evelyn Buckwar and Martin G. Riedler. Runge– Kutta methods for jumpdiffusion differential equations. Journal of Computational and Applied Mathematics, 236(6):1155–1182, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004407.

Bozzini:2013:DRD

[BR13] Mira Bozzini and Milvia Rossini. The detection and recovery of discontinuity curves from scattered data. Journal of Computational and Applied Mathematics, 240(??):148–162, March 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271200249X.

Bertolazzi:2014:FMD

[BR14a] Enrico Bertolazzi and Anna Rimoldi. Fast matrix decomposition in F_{2}. Journal of Computational and Applied Mathematics, 260(??): 519–532, April 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042713005700.

Bini:2014:SSP

[BR14b] Dario A. Bini and Leonardo Robol. Solving secular and polynomial equations: a multiprecision algorithm. Journal of Computational and Applied Mathematics, 272(??):276–292, December 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271300232X.

Bozzini:2014:PGQ

[BR14c] Mira Bozzini and Milvia Rossini. Properties of generators of quasiinterpolation operators of high approximation orders in spaces of polyharmonic splines. Journal of Computational and Applied Mathematics, 267(??):96–106, September 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714000570.

Braun:2012:FEC

[Bra12] Moritz Braun. Finite element calculations for systems with multiple Coulomb centers. Journal of Computational and Applied Mathematics, 236(18):4840–4845, December 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712000623.

Braun:2014:ESC

[Bra14a] Moritz Braun. Electrostatics simulations on the cloud with NCLab. Journal of Computational and Applied Mathematics, 270(??):188– 197, November 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042714000788.

Braun:2014:FEH

[Bra14b] Moritz Braun. Finite element Hartree–Fock calculations in three dimensions for atoms and small molecules. Journal of Computational and Applied Mathematics, 270(??):100– 108, November 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042714000880.

Brezinski:2010:EPE

[BRZ10] C. Brezinski and M. Redivo-Zaglia. Extended procedures for extrapolation to the limit. Journal of Computational and Applied Mathematics, 235(3):631– 645, December 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042710003699.

Baker:2010:PSN

[BS10a] Christopher T. H. Baker and Yihong Song. Periodic solutions of nonlinear discrete Volterra equations with finite memory. Journal of Computational and Applied Mathematics, 234(9):2683–2698, September 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710000233.

Benkhaldoun:2010:SFV

[BS10b] Fayssal Benkhaldoun and Mohammed Seaïd. A simple finite volume method for the shallow water equations. Journal of Computational and Applied Mathematics, 234(1):58–72, May 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709008139.

Bao:2011:TSW

[BS11a] Wendi Bao and Yongzhong Song. Two stage waveform relaxation method for the initial value problems of differential-algebraic equations. Journal of Computational and Applied Mathematics, 236(6):1123–1136, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004389.

Belinskiy:2011:UNB

[BS11b] B. P. Belinskiy and H. Schurz.

Undamped nonlinear beam excited by additive L^{2}- regular noise. Journal of Computational and Applied Mathematics, 235(17):5284– 5306, July 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711002962.

Bause:2013:HOF

[BS13a] Markus Bause and Kristina Schwegler. Higher order finite element approximation of systems of convection– diffusion–reaction equations with small diffusion. Journal of Computational and Applied Mathematics, 246 (??):52–64, July 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042712002956.

Beauregard:2013:ASA

[BS13b] Matthew A. Beauregard and Qin Sheng. An adaptive splitting approach for the quenching solution of reaction–diffusion equations over nonuniform grids. Journal of Computational and Applied Mathematics, 241(??):30– 44, March 15, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042712004244.

Bohmer:2013:NDT

[BS13c] Klaus Böhmer and Robert Schaback. A nonlinear discretization theory. Journal of Computational and Applied Mathematics, 254(??):204–219, December 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713001374.

Becker:2014:RMT

[BSB14] Urs Becker, Bernd Simeon, and Michaek Burger. On Rosenbrock methods for the time integration of nearly incompressible materials and their usage for nonlinear model reduction. Journal of Computational and Applied Mathematics, 262(??):333–345, May 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005955.

Borckmans:2014:RSA

[BSBA14] Pierre B. Borckmans, S. Easter Selvan, Nicolas Boumal, and P.-A. Absil. A Riemannian subgradient algorithm for economic dispatch with valve-point effect. Journal of Computational and Applied Mathematics, 255(??):848–866, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003464.

Bojovic:2014:NAT

[BSJ14] Dejan R. Bojović, Bratislav V.

Sredojević, and Bosko S. Jovanović. Numerical approximation of a twodimensional parabolic timedependent problem containing a delta function. Journal of Computational and Applied Mathematics, 259 (part A)(??):129–137, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713001891.

Bottcher:2012:NIC

[BSKL12] N. Böttcher, A.-K. Singh, O. Kolditz, and R. Liedl. Non-isothermal, compressible gas flow for the simulation of an enhanced gas recovery application. Journal of Computational and Applied Mathematics, 236(18):4933–4943, December 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711005838.

Barglik:2014:MIH

[BSPD14] Jerzy Barglik, Albert Smalcerz, Roman Przylucki, and Ivo Dolezel. 3D modeling of induction hardening of gear wheels. Journal of Computational and Applied Mathematics, 270(??):231–240, November 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042714000478.

Bohmer:2013:DDS

[BSS13] Klaus Böhmer, Karl-Heinz Schild, and Bernhard A.

Schmitt. Dew drops on spider webs: a symmetry breaking bifurcation for a parabolic differentialalgebraic equation. Journal of Computational and Applied Mathematics, 254(??): 99–115, December 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042713001350.

Bao:2014:SCP

[BSS14] Wendi Bao, Yongzhong Song, and Shuqian Shen. Semi-convergence of parameterized Uzawa waveform relaxation method for a class of differential-algebraic equations. Journal of Computational and Applied Mathematics, 258(??):67– 77, March 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713004585.

Bos:2010:LSP

[BSV10] Len Bos, Alvise Sommariva, and Marco Vianello. Leastsquares polynomial approximation on weakly admissible meshes: Disk and triangle. Journal of Computational and Applied Mathematics, 235(3):660–668, December 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710003717.

Briani:2012:CFL

[BSV12] Matteo Briani, Alvise Sommariva, and Marco Vianello. Computing Fekete and Lebesgue points: Simplex, square, disk. Journal of Computational and Applied Mathematics, 236(9):2477– 2486, March 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711006133.

Benedetto:2013:FPP

[BSZ13] Elisa Benedetto, Laura Sacerdote, and Cristina Zucca. A first passage problem for a bivariate diffusion process: Numerical solution with an application to neuroscience when the process is Gauss–Markov. Journal of Computational and Applied Mathematics, 242(??): 41–52, April 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712004402.

Bastani:2012:SCS

[BT12] Ali Foroush Bastani and Mahdieh Tahmasebi. Strong convergence of split-step backward Euler method for stochastic differential equations with non-smooth drift. Journal of Computational and Applied Mathematics, 236(7):1903–1918, January 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271100570X.

Buyukyazici:2014:ACT

[BTSA14] Ibrahim Büyükyazici, Hande Tanberkan, Sevilay Kirci Serenbay, and Çigdem Atakut. Approximation by Chlodowsky type Jakimovski– Leviatan operators. Journal of Computational and Applied Mathematics, 259 (part A)(??):153–163, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042713002070.

Bayraci:2014:SIR

[BÜ14] Selçuk Bayraci and Gazanfer Ünal. Stochastic interest rate volatility modeling with a continuoustime GARCH(1, 1) model. Journal of Computational and Applied Mathematics, 259 (part B)(??):464–473, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271300561X.

Benito:2013:GPS

[BUG^{+}13] J. J. Benito, F. Ureña, L. Gavete, E. Salete, and A. Muelas. A GFDM with PML for seismic wave equations in heterogeneous media. Journal of Computational and Applied Mathematics, 252(??): 40–51, November 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042712003299.

Butcher:2012:RKM

[But12] J. C. Butcher. Runge–Kutta methods and Banach algebras. Journal of Computational and Applied Mathematics, 236(16):3931–3936, October 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712001379.

Buyukkoroglu:2012:CDL

[Büy12] Taner Büyükköroğlu. Common diagonal Lyapunov function for third order linear switched system. Journal of Computational and Applied Mathematics, 236(15):3647–3653, September 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711003360.

Balasubramaniam:2010:GAS

[BV10] P. Balasubramaniam and C. Vidhya. Global asymptotic stability of stochastic BAM neural networks with distributed delays and reaction–diffusion terms. Journal of Computational and Applied Mathematics, 234(12):3458–3466, October 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710002736.

Bultheel:2010:ORF

[BVV10] A. Bultheel, P. Van gucht, and M. Van Barel. Orthogonal rational functions with real coefficients and semiseparable matrices. Journal of Computational and Applied Mathematics, 233(5):1192–1201, January 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709001733.

Basting:2014:HLS

[BW14] Steffen Basting and Martin Weismann. A hybrid level set/front tracking approach for finite element simulations of two-phase flows. Journal of Computational and Applied Mathematics, 270(??):471–483, November 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713006870.

Berdinsky:2014:DBH

[BwKB^{+}14] Dmitry Berdinsky, Tae wan Kim, Cesare Bracco, Durkbin Cho, Bernard Mourrain, Min jae Oh, and Sutipong Kiatpanichgij. Dimensions and bases of hierarchical tensor-product splines. Journal of Computational and Applied Mathematics, 257(??):86–104, February 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271300424X.

Bai:2011:CCW

[BY11] Zhong-Zhi Bai and Xi Yang. On convergence conditions of waveform relaxation methods for linear differential-algebraic equations. Journal of Computational and Applied Mathematics, 235(8):2790–2804, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710006497.

Bai:2012:CTA

[BY12] Zhong-Zhi Bai and Xi Yang. Continuous-time accelerated block successive overrelaxation methods for timedependent Stokes equations. Journal of Computational and Applied Mathematics, 236(13):3265–3285, July 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712000866.

Bean:2014:IIM

[BY14] Maranda Bean and Son-Young Yi. An immersed interface method for a 1D poroelasticity problem with discontinuous coefficients. Journal of Computational and Applied Mathematics, 272(??):81–96, December 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714002453.

Brenner:2013:MFE

[BySZZ13] Susanne C. Brenner, Li yeng Sung, Hongchao Zhang, and Yi Zhang. A Morley finite element method for the displacement obstacle problem of clamped Kirchhoff plates. Journal of Computational and Applied Mathematics, 254(??):31–42, December 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713001088.

Caraus:2012:CCM

[CA12] Iurie Caraus and Feras M. Al Faqih. Convergence of the collocation method and the mechanical quadrature method for systems of singular integro-differential equations in Lebesgue spaces. Journal of Computational and Applied Mathematics, 236(15):3796–3804, September 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042712000374.

Chen:2014:EBT

[CA14] Ruyun Chen and Congpei An. On evaluation of Bessel transforms with oscillatory and algebraic singular integrands. Journal of Computational and Applied Mathematics, 264(??): 71–81, July 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042714000235.

Chen:2010:MFT

[CAA10] Jinhai Chen, Ioannis K. Argyros, and Ravi P. Agarwal. Majorizing functions and two-point Newtontype methods. Journal of Computational and Applied Mathematics, 234(5):1473– 1484, July 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710001081.

Cai:2010:FFC

[Cai10a] Haotao Cai. A fast Fouriercollocation method for second boundary integral equations. Journal of Computational and Applied Mathematics, 234(1):165–173, May 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709008206.

Cai:2010:MES

[Cai10b] Jiaxiang Cai. A multisymplectic explicit scheme for the modified regularized long-wave equation. Journal of Computational and Applied Mathematics, 234(3): 899–905, June 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710000701.

Cai:2013:FSH

[Cai13] Haotao Cai. A fast solver for the Hilbert-type singular integral equations based on the direct Fourier spectral method. Journal of Computational and Applied Mathematics, 250(??):83–95, October 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271300068X.

Cao:2013:NSA

[Cao13] Zhi-Hao Cao. A note on spectrum analysis of augmentation block preconditioned generalized saddle point matrices. Journal of Computational and Applied Mathematics, 238(??):109– 115, January 15, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042712003469.

Cao:2014:SPR

[Cao14] Weiming Cao. On the superconvergence patch recovery techniques for the linear finite element approximation on anisotropic meshes. Journal of Computational and Applied Mathematics, 265(??):33–51, August 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271300486X.

Carlson:2011:HAS

[Car11]  Robert Carlson.  Har-

monic analysis for star graphs and the spherical coordinate trapezoidal rule. Journal of Computational and Applied Mathematics, 235(8):2077–2089, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710005625.

Castro:2010:MPS

[Cas10] Mirta M. Castro. Matrix polynomials satisfying first order differential equations and three term recurrence relations. Journal of Computational and Applied Mathematics, 233(6):1491– 1498, January 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709001526.

Chorfi:2013:SDA

[CAT13] Nejmeddine Chorfi, Nahla Abdellatif, and Sihem Trabelsi. Spectral discretization of the axisymmetric vorticity, velocity and pressure formulation of the Navier–Stokes problem. Journal of Computational and Applied Mathematics, 241(??):1–18, March 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712004050.

Cai:2013:SWC

[CB13] Gang Cai and Shangquan Bu. Strong and weak convergence theorems for general mixed equilibrium problems and variational inequality problems and fixed point problems in Hilbert spaces. Journal of Computational and Applied Mathematics, 247(??):34–52, August 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713000228.

Capilla:2013:NWB

[CBB13] M. T. Capilla and A. Balaguer-Beser. A new wellbalanced non-oscillatory central scheme for the shallow water equations on rectangular meshes. Journal of Computational and Applied Mathematics, 252(??): 62–74, November 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042713000538.

Cruz-Barroso:2010:OIQ

[CBGVPP10] Ruymán Cruz-Barroso, Pablo González-Vera, and Francisco Perdomo-Pío. Orthogonality, interpolation and quadratures on the unit circle and the interval [−1, 1]. Journal of Computational and Applied Mathematics, 235(4):966–981, December 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709008292.

Cruz-Barroso:2011:RAA

[CBGVPP11] Ruymán Cruz-Barroso, Pablo González-Vera, and Francisco Perdomo-Pío. Rational approximants associated with measures supported on the unit circle and the real line. Journal of Computational and Applied Mathematics, 236(1):107–117, August 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711003141.

Chang:2011:OAG

[CC11a] Shin-Perng Chang and Tsu-Fen Chen. Optimal adaptive grids of least-squares finite element methods in two spatial dimensions. Journal of Computational and Applied Mathematics, 235(13):3817– 3824, May 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711000392.

Chumchob:2011:RMA

[CC11b] Noppadol Chumchob and Ke Chen. A robust multigrid approach for variational image registration models. Journal of Computational and Applied Mathematics, 236(5):653–674, October 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711003621.

Claeys:2011:ANA

[CC11c] X. Claeys and F. Collino. Asymptotic and numerical analysis for Holland and Simpson’s thin wire formalism. Journal of Computational and Applied Mathematics, 235(15):4418–4438, June 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711001853.

Cross:2012:QPA

[CC12] Benjamin Cross and Robert J. Cripps. G^{3} quintic polynomial approximation for generalised Cornu spiral segments. Journal of Computational and Applied Mathematics, 236(13):3111– 3122, July 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712000453.

Charina:2013:PRM

[CC13a] Maria Charina and Costanza Conti. Polynomial reproduction of multivariate scalar subdivision schemes. Journal of Computational and Applied Mathematics, 240(??):51–61, March 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042712002488.

Chen:2013:SSQ

[CC13b] Zixin Chen and Wanyou Cheng. Spectral-scaling quasi-Newton methods with updates from the one parameter of the Broyden family. Journal of Computational and Applied Mathematics, 248(??):88–98, August 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713000423.

Chung:2013:SDG

[CC13c] Eric T. Chung and Patrick Ciarlet, Jr. A staggered discontinuous Galerkin method for wave propagation in media with dielectrics and meta-materials. Journal of Computational and Applied Mathematics, 239(??):189– 207, February 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042712004062.

Chen:2014:FLA

[CC14a] Dai-Qiang Chen and Li-Zhi Cheng. Fast linearized alternating direction minimization algorithm with adaptive parameter selection for multiplicative noise removal. Journal of Computational and Applied Mathematics, 257(??):29– 45, February 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713004172.

Chen:2014:EAS

[CC14b] Zhong Chen and Xue Cheng. An efficient algorithm for solving Fredholm integro-differential equations with weakly singular kernels. Journal of Computational and Applied Mathematics, 257(??):57– 64, February 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713004238.

Cen:2013:ECS

[CCCM13] Yi-Gang Cen, Li-Hui Cen, Xiao-Fang Chen, and Zhen-Jiang Miao. Explicit construction of symmetric compactly supported biorthogonal multiwavelets via group transformations. Journal of Computational and Applied Mathematics, 244(??):49– 66, May 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712005110.

Castro:2014:ABS

[CCDS14] Airton Castro, Claudio Cuevas, Filipe Dantas, and Herme Soto. About the behavior of solutions for Volterra difference equations with infinite delay. Journal of Computational and Applied Mathematics, 255(??):44–59, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713002197.

Carmi:2012:CFM

[CCG12] Avishy Carmi, Yair Censor, and Pini Gurfil. Convex feasibility modeling and projection methods for sparse signal recovery. Journal of Computational and Applied Mathematics, 236(17):4318– 4335, November 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712001410.

Calo:2013:GBE

[CCG^{+}13a] Victor M. Calo, Nathan Collier, Matthias Gehre, Bangti Jin, Hany Radwan, and Mauricio Santillana. Gradient-based estimation of Manning’s friction coefficient from noisy data. Journal of Computational and Applied Mathematics, 238(??):1–13, January 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712003263.

Caplan:2013:TSH

[CCG13b] R. M. Caplan and R. Carretero-González. A two-step high-order compact scheme for the Laplacian operator and its implementation in an explicit method for integrating the nonlinear Schrödinger equation. Journal of Computational and Applied Mathematics, 251(??):33–46, October 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713001180.

Charina:2010:TFC

[CCH10] Maria Charina, Charles K. Chui, and Wenjie He. Tight frames of compactly supported multivariate multi-wavelets.

Journal of Computational and Applied Mathematics, 233(8):2044–2061, February 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709006621.

Chen:2012:EEQ

[CCH12] Zhixiang Chen, Feilong Cao, and Jinjie Hu. Error estimates of quasiinterpolation and its derivatives. Journal of Computational and Applied Mathematics, 236(13):3137–3146, July 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712000672.

Chang:2013:SGC

[CCJC13] S.-L. Chang, H.-S. Chen, B.-W. Jeng, and C.-S. Chien. A spectral-Galerkin continuation method for numerical solutions of the Gross–Pitaevskii equation. Journal of Computational and Applied Mathematics, 254(??):2–16, December 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713001118.

Chan:2011:P

[CCQ^{+}11] C. K. Chan, Xiaojun Chen, Liqun Qi, Colin Rogers, Zhong ci Shi, and Ian Sloan. Preface. Journal of Computational and Applied Mathematics, 235(13):3703–3704, May 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711000240.

Conti:2010:FRI

[CCS10] Costanza Conti, Mariantonia Cotronei, and Tomas Sauer. Full rank interpolatory subdivision schemes: Kronecker, filters and multiresolution. Journal of Computational and Applied Mathematics, 233(7):1649– 1659, February 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042709000892.

Chen:2011:RIB

[CCS11] Dai-Qiang Chen, Li-Zhi Cheng, and Fang Su. Restoration of images based on subspace optimization accelerating augmented Lagrangian approach. Journal of Computational and Applied Mathematics, 235 (8):2766–2774, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042710006461.

Crnkovic:2012:BWE

[CCZ12] Bojan Crnković and Nelida Crnjarić-Zic. Binary weighted essentially nonoscillatory (BWENO) approximation. Journal of Computational and Applied Mathematics, 236(9):2431– 2451, March 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711005991.

Cavoretto:2010:FAI

[CD10] Roberto Cavoretto and Alessandra De Rossi. Fast and accurate interpolation of large scattered data sets on the sphere. Journal of Computational and Applied Mathematics, 234(5):1505– 1521, July 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710001299.

Conti:2011:ASS

[CD11] Costanza Conti and Nira Dyn. Analysis of subdivision schemes for nets of functions by proximity and controllability. Journal of Computational and Applied Mathematics, 236(4):461–475, September 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271100327X.

Caira:2012:BSL

[CDD12] R. Caira, F. Dell’Accio, and F. Di Tommaso. On the bivariate Shepard–Lidstone operators. Journal of Computational and Applied Mathematics, 236(7):1691– 1707, January 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042711005322.

Ciftci:2014:RAB

[ÇDE14] Sezgin Çiftçi, Sebnem Düzgün, and B. Burçak Basbug Erkan. Risk analysis based on spatial analysis of chronic obstructive pulmonary disease (COPD) with respect to provinces in Turkey. Journal of Computational and Applied Mathematics, 259 (part B)(??): 413–424, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042713004287.

Conte:2013:NSA

[CDJP13] Dajana Conte, Raffaele D’Ambrosio, Zdzisław Jackiewicz, and Beatrice Paternoster. Numerical search for algebraically stable two-step almost collocation methods. Journal of Computational and Applied Mathematics, 239(??):304–321, February 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003342.

Costabel:2010:VSI

[CDK10] M. Costabel, E. Darrigrand, and E. H. Koné. Volume and surface integral equations for electromagnetic scattering by a dielectric body. Journal of Computational and Applied Mathematics, 234(6):1817–1825, July 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709005081.

Conen:2014:CSH

[CDKN14] Lea Conen, Victorita Dolean, Rolf Krause, and Frédéric Nataf. A coarse space for heterogeneous Helmholtz problems based on the Dirichlet-to-Neumann operator. Journal of Computational and Applied Mathematics, 271(??):83–99, December 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/S0377042714001800.

See addendum [CDKN15].

Conen:2015:ACS

[CDKN15] Lea Conen, Victorita Dolean, Rolf Krause, and Frédéric Nataf. Addendum to “A coarse space for heterogeneous Helmholtz problems based on the Dirichletto-Neumann operator” [J. Comput. Appl. Math. 271 (2014) 83–99]. Journal of Computational and Applied Mathematics, 290(??):670– 674, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042715002502. See

[CDKN14].

Cao:2014:SSP

[CDN14] Yang Cao, Jun Du, and Qiang Niu. Shift-splitting preconditioners for saddle point problems. Journal of Computational and Applied Mathematics, 272(??):239– 250, December 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042714002623.

Clarysse:2010:OEO

[CDPP10] Patrick Clarysse, Bertrand Delhay, Martine Picq, and Jérôme Pousin. Optimal extended optical flow subject to a statistical constraint. Journal of Computational and Applied Mathematics, 234(4):1291–1302, June 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704270900692X.

Colciago:2014:CBR

[CDQ14] C. M. Colciago, S. Deparis, and A. Quarteroni. Comparisons between reduced order models and full 3D models for fluid-structure interaction problems in haemodynamics. Journal of Computational and Applied Mathematics, 265(??):120– 138, August 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713005049.

Cookson:2010:UCT

[CDS10] A. N. Cookson, D. J.

Doorly, and S. J. Sherwin. Using coordinate transformation of Navier–Stokes equations to solve flow in multiple helical geometries. Journal of Computational and Applied Mathematics, 234(7):2069–2079, August 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709005421.

Calvo:2011:NMF

[CDTV11] N. Calvo, J. Durany, R. Toja, and C. Vázquez. Numerical methods for a fixed domain formulation of the glacier profile problem with alternative boundary conditions. Journal of Computational and Applied Mathematics, 235(5):1394– 1411, January 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042710004759.

Chesnokov:2010:NSC

[CDV10] Andrey Chesnokov, Karl Deckers, and Marc Van Barel. A numerical solution of the constrained weighted energy problem. Journal of Computational and Applied Mathematics, 235(4):950–965, December 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709008073.

Collins:2014:PEE

[CET14] J. B. Collins, Don Estep, and Simon Tavener. A posteriori error estimation for the Lax–Wendroff finite difference scheme. Journal of Computational and Applied Mathematics, 263(??):299– 311, June 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713007085.

Copetti:2011:NAQ

[CF11] M. I. M. Copetti and J. R.

Fernández. Numerical analysis of a quasi-static contact problem for a thermoviscoelastic beam. Journal of Computational and Applied Mathematics, 235(14):4165– 4173, May 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711001415.

Criado:2011:ARB

[CFdAR11] R. Criado, J. Flores, A. García del Amo, and M. Romance. Analytical relationships between metric and centrality measures of a network and its dual. Journal of Computational and Applied Mathematics, 235(7):1775–1780, February 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710002086.

Cheng:2012:OFM

[CFF12] Hao Cheng, Chu-Li Fu, and Xiao-Li Feng. An optimal filtering method for stable analytic continuation. Journal of Computational and Applied Mathematics, 236(9):2582–2589, March 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711006303.

Carrera:2010:NAP

[CFGE10] Edmundo M. Carrera, Nelson Flores-Gallegos, and Rodolfo O. Esquivel. Natural atomic probabilities in quantum information theory. Journal of Computational and Applied Mathematics, 233(6):1483–1490, January 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709001514.

Chu:2014:DOE

[CFH14] Peter C. Chu, Chenwu Fan, and Norden Huang. Derivative-optimized empirical mode decomposition for the Hilbert–Huang transform. Journal of Computational and Applied Mathematics, 259 (part A)(??): 57–64, March 15, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713001738.

Chu:2011:RRM

[CFJ^{+}11] Eric King-Wah Chu, Hung-Yuan Fan, Zhongxiao Jia, Tiexiang Li, and Wen-Wei Lin. The Rayleigh–Ritz method, refinement and Arnoldi process for periodic matrix pairs. Journal of Computational and Applied Mathematics, 235(8):2626– 2639, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704271000631X.

Cui:2014:NAS

[CFL^{+}14] Ming Cui, Kai Fu, Dong Liang, Yu Cheng, and Wenqia Wang. Numerical analysis of the secondorder characteristic FEM for nonlinear aerosol dynamic equations. Journal of Computational and Applied Mathematics, 261(??):48– 61, May 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005980.

Calvo:2012:SNL

[CFMR12] M. Calvo, J. M. Franco, J. I.

Montijano, and L. Rández. On some new low storage implementations of time advancing Runge– Kutta methods. Journal of Computational and Applied Mathematics, 236(15):3665– 3675, September 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042711003906.

Calvete:2010:LBP

[CG10a] Herminia I. Calvete and Carmen Galé. Linear bilevel programs with multiple objectives at the upper level. Journal of Computational and Applied Mathematics, 234(4):950–959, June 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042708006353.

Carnicer:2010:CND

[CG10b] J. M. Carnicer and C. Godés. Configurations of nodes with defects greater than three. Journal of Computational and Applied Mathematics, 233(7):1640–1648, February 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709000880.

Cao:2011:MEW

[CG11a] Chunhong Cao and Xieping Gao. Minimum-energy wavelet frame on the interval with arbitrary integer dilation factor.

Journal of Computational and Applied Mathematics, 235(8):1885–1896, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271000659X.

Cheikh:2011:CDC

[CG11b] Y. Ben Cheikh and M. Gaied.

A characterization of Dunklclassical d-symmetric dorthogonal polynomials and its applications. Journal of Computational and Applied Mathematics, 236(1):49–64, August 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711001555.

Calvete:2012:LBP

[CG12] Herminia I. Calvete and Carmen Galé. Linear bilevel programming with interval coefficients. Journal of Computational and Applied Mathematics, 236(15):3751– 3762, September 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711005450.

Clavero:2013:HOU

[CG13] C. Clavero and J. L. Gracia.

A higher order uniformly convergent method with Richardson extrapolation in time for singularly perturbed reaction–diffusion parabolic problems. Journal of Computational and Applied Mathematics, 252(??): 75–85, November 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712002336.

Chen:2014:GOP

[CG14] Qingshan Chen and Max Gunzburger. Goal-oriented a posteriori error estimation for finite volume methods. Journal of Computational and Applied Mathematics, 265(??):69–82, August 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005487.

Cabrera:2013:VFA

[CGCG^{+}13] Daniel A. Cervantes Cabrera, Pedro González-Casanova, Christian Gout, L. Héctor Juárez, and L. Rafael Reséndiz. Vector field approximation using radial basis functions. Journal of Computational and Applied Mathematics, 240(??):163– 173, March 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712002932.

Constales:2010:CTD

[CGK10] Denis Constales, Dennis Grob, and Rolf Sören Kraußhar. Constructing three-dimensional mappings onto the unit sphere with the hypercomplex Szegő kernel. Journal of Computational and Applied Mathematics, 233(11):2884–2901, April 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704270900781X.

Clavero:2010:ATO

[CGL10] C. Clavero, J. L. Gracia, and F. J. Lisbona. An almost third order finite difference scheme for singularly perturbed reaction–diffusion systems. Journal of Computational and Applied Mathematics, 234(8):2501–2515, August 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710001597.

Chitescu:2010:AID

[CGM10] I. Chitescu, H. Georgescu, and R. Miculescu. Approximation of infinite dimensional fractals generated by integral equations. Journal of Computational and Applied Mathematics, 234(5):1417–1425, July 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710000993.

Cuadra:2011:DAI

[CGRLR11] J. Cuadra, J. M. García-Rubira, and J. A. López-Ramos. Determining all indecomposable codes over some Hopf algebras. Journal of Computational and Applied Mathematics, 235(7):1833–1839, February 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710004115.

Coll:2011:ASH

[CGRS11] Narcís Coll, Marité Guerrieri, Maria-Cecilia Rivara, and J. Antoni Sellarès. Adaptive simplification of huge sets of terrain grid data for geosciences applications. Journal of Computational and Applied Mathematics, 236(6):1410–1422, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004845.

Clavero:2011:SAH

[CGS11] C. Clavero, J. L. Gracia, and M. Stynes. A simpler analysis of a hybrid numerical method for time-dependent convection– diffusion problems. Journal of Computational and Applied Mathematics, 235(17): 5240–5248, July 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042711002822.

Castro:2012:AMM

[CGS12] Douglas A. Castro, Sônia M.

Gomes, and Jorge Stolfi. An adaptive multiresolution method on dyadic grids: Application to transport equations. Journal of Computational and Applied Mathematics, 236(15):3636– 3646, September 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711003165.

Cherkas:2013:CBC

[CGS13] Leonid Cherkas, Alexander Grin, and Klaus R. Schneider. On the construction of bifurcation curves related to limit cycles of multiplicity three for planar vector fields. Journal of Computational and Applied Mathematics, 237(1):654– 662, January 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042712003330.

Chen:2014:EDD

[CGW14] Mi Chen, Junyi Guo, and Xueyuan Wu. Expected discounted dividends in a discrete semi-Markov risk model. Journal of Computational and Applied Mathematics, 266(??):1–17, August 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042714000545.

Cox:2011:GPM

[CH11] Barry J. Cox and James M. Hill. Geometric polyhedral models for nanotubes comprising hexagonal lattices. Journal of Computational and Applied Mathematics, 235(13):3943–3952, May 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711000513.

Cotronei:2013:PPO

[CH13] Mariantonia Cotronei and Matthias Holschneider. Partial parameterization of orthogonal wavelet matrix filters. Journal of Computational and Applied Mathematics, 243(??):113–125, May 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712005092.

Console:2014:RRE

[CH14] Paola Console and Ernst Hairer. Reducing round-off errors in symmetric multistep methods. Journal of Computational and Applied Mathematics, 262(??):217– 222, May 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003695.

Chanane:2010:ASF

[Cha10a] Bilal Chanane. Accurate solutions of fourth order Sturm–Liouville problems. Journal of Computational and Applied Mathematics, 234(10):3064–3071, September 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710002219.

Chang:2010:VIM

[Cha10b] Shih-Hsiang Chang. A variational iteration method for solving Troesch’s problem. Journal of Computational and Applied Mathematics, 234(10):3043–3047, September 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710002153.

Chan:2011:QCM

[Cha11a] C. Y. Chan. A quenching criterion for a multidimensional parabolic problem due to a concentrated nonlinear source. Journal of Computational and Applied Mathematics, 235(13):3724– 3727, May 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711000288.

Chan:2011:QND

[Cha11b] W. Y. Chan. Quenching for nonlinear degenerate parabolic problems. Journal of Computational and Applied Mathematics, 235(13): 3831–3840, May 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042711000410.

Chen:2010:BEI

[Che10] Xiao Shan Chen. A backward error for the inverse singular value problem. Journal of Computational and Applied Mathematics, 234(8):2450–2455, August 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710001445.

Chen:2013:EBT

[Che13a] Ruyun Chen. On the evaluation of Bessel transformations with the oscillators via asymptotic series of Whittaker functions. Journal of Computational and Applied Mathematics, 250(??):107– 121, October 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713000952.

Chen:2013:HWT

[Che13b]  Xiao Shan Chen.  A

Hoffman–Wielandt-type residual bound for generalized eigenvalues of a definite pair. Journal of Computational and Applied Mathematics, 237(1):208–214, January 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271200307X.

Chen:2014:EES

[Che14a] Guangyi Chen. Are electroencephalogram (EEG) signals pseudo-random number generators? Journal of Computational and Applied Mathematics, 268(??):1–4, October 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271400123X.

Chen:2014:ALD

[Che14b] Mei-Xiang Chen. An augmented Lagrangian dual optimization approach to the H-weighted model updating problem. Journal of Computational and Applied Mathematics, 268(??):111– 120, October 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042714001228.

Chun:2010:HOA

[CHH10] S. Chun, H. Haddar, and J. S. Hesthaven. Highorder accurate thin layer approximations for timedomain electromagnetics, part II: Transmission layers. Journal of Computational and Applied Mathematics, 234(8):2587–2608, August 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710001809.

Cao:2011:AUF

[CHL11] Yong Cao, Xiaoming He, and Tao Lü. An algorithm using the finite volume element method and its splitting extrapolation. Journal of Computational and Applied Mathematics, 235(13):3734–3742, May 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711000306.

Cheng:2014:IFT

[CHL14] Jin Cheng, Bernd Hofmann, and Shuai Lu. The index function and Tikhonov regularization for ill-posed problems. Journal of Computational and Applied Mathematics, 265(??):110– 119, August 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713004901.

Chaudhry:2014:SRC

[CHMD14] Q. A. Chaudhry, M. Hanke, R. Morgenstern, and K. Dreij. Surface reactions on the cytoplasmatic membranes — mathematical modeling of reaction and diffusion systems in a cell. Journal of Computational and Applied Mathematics, 262(??):244– 260, May 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713004810.

Cordero:2010:AMO

[CHMT10a] Alicia Cordero, José L. Hueso, Eulalia Martínez, and Juan R. Torregrosa. Accelerated methods for order 2p for systems of nonlinear equations. Journal of Computational and Applied Mathematics, 233(10):2696– 2702, March 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042709007493.

Cordero:2010:NMP

[CHMT10b] Alicia Cordero, José L. Hueso, Eulalia Martínez, and Juan R. Torregrosa. New modifications of Potra– Pták’s method with optimal fourth and eighth orders of convergence. Journal of Computational and Applied Mathematics, 234 (10):2969–2976, September 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710002062.

Cordero:2012:STM

[CHMT12] Alicia Cordero, José L. Hueso, Eulalia Martínez, and Juan R. Torregrosa. Steffensen type methods for solving nonlinear equations. Journal of Computational and Applied Mathematics, 236(12):3058–3064, June 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710005030.

Cordero:2013:NTO

[CHMT13] Alicia Cordero, José L.

Hueso, Eulalia Martínez, and Juan R. Torregrosa. A new technique to obtain derivative-free optimal iterative methods for solving nonlinear equations. Journal of Computational and Applied Mathematics, 252(??):95–102, November 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712001586.

Chen:2013:LLB

[CHRZ13] Xingwu Chen, Wentao Huang, Valery G. Romanovski, and Weinian Zhang. Linearizability and local bifurcation of critical periods in a cubic Kolmogorov system. Journal of Computational and Applied Mathematics, 245(??): 86–96, June 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712005341.

Chuong:2010:TTR

[Chu10] Thai Doan Chuong. Tikhonovtype regularization method for efficient solutions in vector optimization. Journal of Computational and Applied Mathematics, 234(3):761– 766, June 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710000440.

Chen:2011:CFE

[CHX11] Yumei Chen, Feiteng Huang, and Xiaoping Xie. H(div) conforming finite element methods for the coupled Stokes and Darcy problem. Journal of Computational and Applied Mathematics, 235(15):4337–4349, June 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711001531.

Cripps:2010:SPA

[CHZ10] R. J. Cripps, M. Z. Hussain, and S. Zhu. Smooth polynomial approximation of spiral arcs. Journal of Computational and Applied Mathematics, 233(9):2227–2234, March 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709006864.

Chabassier:2013:ISF

[CI13] Juliette Chabassier and Sébastien Imperiale. Introduction and study of fourth order theta schemes for linear wave equations. Journal of Computational and Applied Mathematics, 245(??): 194–212, June 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713000150.

Cvetkovic-Ilic:2010:DID

[CID10] Dragana S. Cvetković-Ilić and Chun Yuan Deng. The Drazin invertibility of the difference and the sum of two idempotent operators. Journal of Computational and Applied Mathematics, 233(8):1717–1722, February 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709006517.

Cieslinski:2014:IAA

[Cie14] Jan L. Cieśliński. Improving the accuracy of the AVF method. Journal of Computational and Applied Mathematics, 259 (part A)(??): 233–243, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042713004135.

Cinkir:2012:EAC

[Cin12] Zubeyir Cinkir. An elementary algorithm for computing the determinant of pentadiagonal Toeplitz matrices. Journal of Computational and Applied Mathematics, 236(9):2298–2305, March 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711005875.

Cinkir:2014:FEA

[Cin14] Zubeyir Cinkir. A fast elementary algorithm for computing the determinant of Toeplitz matrices. Journal of Computational and Applied Mathematics, 255(??):353–361, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713002896.

Conte:2014:EFG

[CIPS14] Dajana Conte, Liviu Gr. Ixaru, Beatrice Paternoster, and Giuseppe Santomauro. Exponentially-fitted Gauss– Laguerre quadrature rule for integrals over an unbounded interval. Journal of Computational and Applied Mathematics, 255(??): 725–736, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042713003385.

Chen:2010:HAI

[CJ10] Fang Chen and Yao-Lin Jiang. On HSS and AHSS iteration methods for nonsymmetric positive definite Toeplitz systems. Journal of Computational and Applied Mathematics, 234(8):2432– 2440, August 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710001469.

Cao:2010:NCP

[CJY10] Yang Cao, Mei-Qun Jiang, and Lin-Quan Yao. New choices of preconditioning matrices for generalized inexact parameterized iterative methods. Journal of Computational and Applied Mathematics, 235(1):263– 269, November 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710003444.

Chaggara:2011:LCC

[CK11] Hamza Chaggara and Wolfram Koepf. On linearization and connection coefficients for generalized Hermite polynomials. Journal of Computational and Applied Mathematics, 236(1): 65–73, August 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711001403.

Campbell:2013:NTL

[CK13] Stephen L. Campbell and Peter Kunkel. On the numerical treatment of linearquadratic optimal control problems for general linear time-varying differentialalgebraic equations. Journal of Computational and Applied Mathematics, 242(??): 213–231, April 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712004372.

Cermak:2014:EDS

[CK14a] Jan Cermák and Tomás Kisela. Exact and discretized stability of the Bagley–Torvik equation. Journal of Computational and Applied Mathematics, 269(??):53–67, October 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714001575.

Connor:2014:ITD

[CK14b] Dale Connor and Lilia Krivodonova. Interpolation of two-dimensional curves with Euler spirals. Journal of Computational and Applied Mathematics, 261(??): 320–332, May 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713006286.

Christou:2014:MRS

[CKM14] D. Christou, N. Karcanias, and M. Mitrouli. Matrix representation of the shifting operation and numerical properties of the ERES method for computing the greatest common divisor of sets of many polynomials. Journal of Computational and Applied Mathematics, 260(??):54– 67, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713004767.

Cardoso:2010:ESS

[CL10a] João R. Cardoso and F. Silva Leite. Exponentials of skew-symmetric matrices and logarithms of orthogonal matrices. Journal of Computational and Applied Mathematics, 233(11):2867– 2875, April 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042709007791.

Chen:2010:TGM

[CL10b] Chuanjun Chen and Wei Liu. A two-grid method for finite volume element approximations of secondorder nonlinear hyperbolic equations. Journal of Computational and Applied Mathematics, 233(11):2975– 2984, April 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042709007900.

Costabile:2010:DAA

[CL10c] F. A. Costabile and E. Longo. A determinantal approach to Appell polynomials. Journal of Computational and Applied Mathematics, 234(5):1528–1542, July 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042710001317.

Cen:2011:RAF

[CL11a] Zhongdi Cen and Anbo Le. A robust and accurate finite difference method for a generalized Black– Scholes equation. Journal of Computational and Applied Mathematics, 235(13):3728– 3733, May 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271100029X.

Chen:2011:MMS

[CL11b] Chuanjun Chen and Wei Liu. The MMOC and MMOCAA schemes for the finite volume element method of convection– diffusion problems. Journal of Computational and Applied Mathematics, 236 (6):1423–1434, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042711004857.

Chen:2011:PAP

[CL11c] Wei Chen and Bingwen Liu.

Positive almost periodic solution for a class of Nicholson’s blowflies model with multiple time-varying delays. Journal of Computational and Applied Mathematics, 235(8):2090–2097, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710005637.

Chen:2011:IFM

[CL11d] Xi Chen and Dali Liu.

Isospectral flow method for nonnegative inverse eigenvalue problem with prescribed structure. Journal of Computational and Applied Mathematics, 235(14): 3990–4002, May 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042711000781.

Chen:2011:PEE

[CL11e] Yanping Chen and Yijie Lin. A posteriori error estimates for hp finite element solutions of convex optimal control problems. Journal of Computational and Applied Mathematics, 235(12):3435– 3454, April 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704271100077X.

Costabile:2011:AIP

[CL11f] F. A. Costabile and E. Longo. The Appell interpolation problem. Journal of Computational and Applied Mathematics, 236(6):1024– 1032, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042711003694.

Chen:2013:TSV

[CL13a] Zhen Chen and Linzhang Lu. A tensor singular values and its symmetric embedding eigenvalues. Journal of Computational and Applied Mathematics, 250(??):217– 228, October 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713001222.

Chiang:2013:EST

[CL13b] Chun-Yueh Chiang and Matthew M. Lin. The eigenvalue shift technique and its eigenstructure analysis of a matrix. Journal of Computational and Applied Mathematics, 253(??):235–248, December 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713002100.

Chen:2014:ECS

[CL14a] Jingmin Chen and Lin Li.

Existence of continuous solutions and stability for a general iterative functional equation. Journal of Computational and Applied Mathematics, 260(??):509– 518, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005736.

Cheng:2014:ORP

[CL14b] Guo Qing Cheng and Ling Li. An optimal replacement policy for a degenerative system with two-types of failure states. Journal of Computational and Applied Mathematics, 261(??):139– 145, May 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713006201.

Chen:2011:NAV

[CLB11] Chang-Ming Chen, F. Liu, and K. Burrage. Numerical analysis for a variableorder nonlinear cable equation. Journal of Computational and Applied Mathematics, 236(2):209–224, August 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711003554.

Chai:2011:PIS

[CLLX11] Junjie Chai, Na Lei, Ying Li, and Peng Xia. The proper interpolation space for multivariate Birkhoff interpolation. Journal of Computational and Applied Mathematics, 235(10):3207– 3214, March 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711000070.

Costabile:2014:OPU

[CLMR14] Massimo Costabile, Arturo Leccadito, Ivar Massabó, and Emilio Russo. Option pricing under regimeswitching jump-diffusion models. Journal of Computational and Applied Mathematics, 256(??):152–167, January 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003907.

Chen:2014:CNS

[CLN14] Xiao Shan Chen, Wen Li, and Michael K. Ng. On condition numbers of the spectral projections associated with periodic eigenproblems. Journal of Computational and Applied Mathematics, 272(??):417–429, December 15, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713003178.

Chiavassa:2010:NMT

[CLP10] Guillaume Chiavassa, Bruno Lombard, and Joël Piraux. Numerical modeling of 1D transient poroelastic waves in the low-frequency range. Journal of Computational and Applied Mathematics, 234(6):1757–1765, July 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709005007.

Chiu:2010:SOD

[CLS10] Pao-Hsiung Chiu, Long Lee, and Tony W. H. Sheu. A sixth-order dual preserving algorithm for the Camassa– Holm equation. Journal of Computational and Applied Mathematics, 233(11):2767– 2778, April 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S037704270900764X.

Chan:2012:BVM

[CLS12] Raymond H. Chan, Spike T. Lee, and Hai-Wei Sun. Boundary value methods for transient solutions of queueing networks with variant vacation policy. Journal of Computational and Applied Mathematics, 236(16):3948– 3955, October 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042712000386.

Carpentieri:2014:VVB

[CLS14] Bruno Carpentieri, Jia Liao, and Masha Sosonkina. VBARMS: a variable block algebraic recursive multilevel solver for sparse linear systems.

Journal of Computational and Applied Mathematics, 259 (part A)(??):164–173, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713002227.

Chen:2014:CAM

[CLW14] Minhong Chen, Rongfei Lin, and Qingbiao Wu. Convergence analysis of the modified Newton–HSS method under the Hölder continuous condition. Journal of Computational and Applied Mathematics, 264(??):115– 130, July 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713007206.

Chidume:2010:IMC

[CM10a] Charles E. Chidume and Stefan Maruster. Iterative methods for the computation of fixed points of demicontractive mappings. Journal of Computational and Applied Mathematics, 234(3):861–882, June 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710000634.

Cimrak:2010:DPD

[CM10b] I. Cimrák and V. Melicher. Determination of precession and dissipation parameters in micromagnetism. Journal of Computational and Applied Mathematics, 234(7):2239–2249, August 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709005603.

Costantini:2010:CSC

[CM10c] Paolo Costantini and Carla Manni. Curve and surface construction using Hermite subdivision schemes. Journal of Computational and Applied Mathematics, 233(7):1660–1673, February 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709000909.

Cuminato:2010:NES

[CM10d] J. A. Cuminato and S. Mc-Kee. A note on the eigenvalues of a special class of matrices. Journal of Computational and Applied Mathematics, 234(9):2724–2731, September 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710000427.

Cai:2012:MDM

[CM12] Mingchao Cai and Mo Mu.

A multilevel decoupled method for a mixed Stokes/ Darcy model. Journal of Computational and Applied Mathematics, 236(9):2452– 2465, March 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711006108.

Carstensen:2013:CSP

[CM13] C. Carstensen and C. Merdon. Computational survey on a posteriori error estimators for nonconforming finite element methods for the Poisson problem. Journal of Computational and Applied Mathematics, 249(??):74– 94, September 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713000137.

Chidyagwai:2011:CFV

[CMR11a] Prince Chidyagwai, Ilya Mishev, and Béatrice Rivière. On the coupling of finite volume and discontinuous Galerkin method for elliptic problems. Journal of Computational and Applied Mathematics, 235(8):2193– 2204, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042710005832.

Costabile:2011:PAA

[CMR11b] Massimo Costabile, Ivar Massabó, and Emilio Russo. On pricing arithmetic average reset options with multiple reset dates in a lattice framework. Journal of Computational and Applied Mathematics, 235(17):5307– 5325, July 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711002986.

Cesmelioglu:2014:PFF

[ÇMT14] Ayça Çesmelioğlu, Wilfried Meidl, and Alev Topuzoğlu. Permutations of finite fields with prescribed properties. Journal of Computational and Applied Mathematics, 259 (part B)(??):536–545, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713003804.

Creuse:2010:PEE

[CN10] Emmanuel Creusé and Serge Nicaise. A posteriori error estimator based on gradient recovery by averaging for discontinuous Galerkin methods. Journal of Computational and Applied Mathematics, 234 (10):2903–2915, September 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710001858.

Chow:2011:PSD

[CN11] K. W. Chow and T. W.

Ng. Periodic solutions of a derivative nonlinear Schrödinger equation: Elliptic integrals of the third kind. Journal of Computational and Applied Mathematics, 235(13):3825–3830, May 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711000409.

Cenk:2011:EM

[CÖ11a] Murat Cenk and Ferruh Özbudak. Efficient multiplications in F_{5^{5n}} and F_{7^{7n}}. Journal of Computational and Applied Mathematics, 236(2):177–183, August 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711003396.

Coskun:2011:PFN

[CO11b] Cafer Coskun and Murat Olgun. Principal functions of non-selfadjoint matrix Sturm–Liouville equations. Journal of Computational and Applied Mathematics, 235(16):4834–4838, June 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006540.

Chen:2013:CPR

[CO13] Xu Chen and Hui Ou. A compound Poisson risk model with proportional investment. Journal of Computational and Applied Mathematics, 242(??):248– 260, April 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712004670.

Colombeau:2013:SNS

[Col13] M. Colombeau. A simple numerical scheme for the 3-D system of ideal gases and a study of approximation in the sense of distributions. Journal of Computational and Applied Mathematics, 248(??):15–30, August 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713000186.

Chapwanya:2010:PSB

[COW10] Michael Chapwanya, Stephen O’Brien, and J. F. Williams. A perturbation solution for bacterial growth and bioremediation in a porous medium with bio-clogging. Journal of Computational and Applied Mathematics, 234(9):2709–2723, September 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710000282.

Chadjiconstantinidis:2013:PDC

[CP13a] Stathis Chadjiconstantinidis and Apostolos D. Papaioannou. On a perturbed by diffusion compound Poisson risk model with delayed claims and multi-layer dividend strategy. Journal of Computational and Applied Mathematics, 253(??):26– 50, December 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713000848.

Chen:2013:ASC

[CP13b] Juanjuan Chen and Fengfu Peng. Approximate spline of G^{2}-continuity on a generalized hyperbolic paraboloid. Journal of Computational and Applied Mathematics, 248(??):99–117, August 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713000575.

Clarysse:2011:OEO

[CPP11] Patrick Clarysse, Martine Picq, and Jérôme Pousin. Optimal extended optical flow and statistical constraints: a result of convergence. Journal of Computational and Applied Mathematics, 235(7):1840–1848, February 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710003018.

Chabannes:2013:HOF

[CPP13] Vincent Chabannes, Gonçalo Pena, and Christophe Prud’homme. High-order fluid-structure interaction in 2D and 3D application to blood flow in arteries. Journal of Computational and Applied Mathematics, 246(??):1– 9, July 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712004256.

Chen:2010:NCT

[CPY10] Jein-Shan Chen, Shaohua Pan, and Ching-Yu Yang. Numerical comparisons of two effective methods for mixed complementarity problems. Journal of Computational and Applied Mathematics, 234(3):667– 683, June 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710000087.

Cohen:2011:GFD

[CR11a] David Cohen and Xavier Raynaud. Geometric finite difference schemes for the generalized hyperelastic-rod wave equation. Journal of Computational and Applied Mathematics, 235(8):1925– 1940, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042710005273.

Conti:2011:ACN

[CR11b] Costanza Conti and Lucia Romani. Algebraic conditions on non-stationary subdivision symbols for exponential polynomial reproduction. Journal of Computational and Applied Mathematics, 236(4):543–556, September 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711001725.

Camano:2012:AFB

[CR12] Jessika Camaño and Rodolfo Rodríguez. Analysis of a FEM–BEM model posed on the conducting domain for the time-dependent eddy current problem. Journal of Computational and Applied Mathematics, 236(13):3084– 3100, July 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712000416.

Cheng:2014:CQB

[CR14] Yongtao Cheng and James A. Rossmanith. A class of quadrature-based momentclosure methods with application to the Vlasov– Poisson–Fokker–Planck system in the high-field limit. Journal of Computational and Applied Mathematics, 262(??):384–398, May 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042713005943.

Costa:2014:AFS

[CRF14] M. Fernanda P. Costa, Ana Maria A. C. Rocha, and Edite M. G. P. Fernandes. An artificial fish swarm algorithm based hyperbolic augmented Lagrangian method. Journal of Computational and Applied Mathematics, 259 (part B)(??):868–876, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713004226.

Chen:2010:SOL

[CRM10] Shaochun Chen, Guobiao Ren, and Shipeng Mao. Second-order locking-free nonconforming elements for planar linear elasticity. Journal of Computational and Applied Mathematics, 233(10):2534–2548, March 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709007328.

Ciric:2010:CFP

[ĆRR^{+}10] Ljubomir Ćirić, Vladimir Rakocević, Stojan Radenović, Miloje Rajović, and Rade Lazović. Common fixed point theorems for non-self-mappings in metric spaces of hyperbolic type. Journal of Computational and Applied Mathematics, 233(11):2966–2974, April 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709007894.

Criado:2012:IPD

[CRS12] Regino Criado, Miguel Romance, and Ángel Sánchez. Interest point detection in images using complex network analysis. Journal of Computational and Applied Mathematics, 236(12):2975– 2980, June 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711002834.

Cao:2010:MNE

[CS10] Wan-Rong Cao and Zhi-Zhong Sun. Maximum norm error estimates of the Crank–Nicolson scheme for solving a linear moving boundary problem. Journal of Computational and Applied Mathematics, 234(8):2578–2586, August 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710001822.

Cambini:2012:PAS

[CS12] Riccardo Cambini and Claudio Sodini. A parametric approach for solving a class of generalized quadratic-transformable ranktwo nonconvex programs. Journal of Computational and Applied Mathematics, 236(10):2685–2695, April 2012. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042712000052.

Chen:2014:REA

[CS14] Zheng Chen and Chi-Wang Shu. Recovering exponential accuracy from collocation point values of smooth functions with end-point singularities. Journal of Computational and Applied Mathematics, 265(??):83– 95, August 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713004846.

Cicone:2010:GPP

[CSC10] Antonio Cicone and Stefano Serra-Capizzano. Google PageRanking problem: the model and the analysis. Journal of Computational and Applied Mathematics, 234(11):3140–3169, October 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710000762.

Cai:2011:TVC

[CSL^{+}11] Yu Cai, Zhixun Su, Zhilong Li, Riming Sun, Xiuping Liu, and Yuandi Zhao. Two-view curve reconstruction based on the snake model. Journal of Computational and Applied Mathematics, 236(5):631–639, October 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004158.

Chen:2014:NAS

[CSSX14a] Feishe Chen, Lixin Shen, Bruce W. Suter, and Yuesheng Xu. Nesterov’s algorithm solving dual formulation for compressed sensing. Journal of Computational and Applied Mathematics, 260(??):1– 17, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713004871.

See also reprint [CSSX14b].

Chen:2014:RNA

[CSSX14b] Feishe Chen, Lixin Shen, Bruce W. Suter, and Yuesheng Xu. Reprint of “Nesterov’s algorithm solving dual formulation for compressed sensing”. Journal of Computational and Applied Mathematics, 265(??):52– 68, August 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714000260.

See [CSSX14a].

Ceng:2010:IIA

[CSY10] L. C. Ceng, D. R. Sahu, and J. C. Yao. Implicit iterative algorithms for asymptotically nonexpansive mappings in the intermediate sense and Lipschitzcontinuous monotone mappings. Journal of Computational and Applied Mathematics, 233(11):2902–2915, April 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709007821.

Chen:2011:MSF

[CSZ11a] Yaming Chen, Songhe Song, and Huajun Zhu. The multisymplectic Fourier pseudospectral method for solving two-dimensional Hamiltonian PDEs. Journal of Computational and Applied Mathematics, 236(6):1354– 1369, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042711004699.

Cui:2011:GBM

[CSZ11b] Yujun Cui, Jingxian Sun, and Yumei Zou. Global bifurcation and multiple results for Sturm–Liouville problems. Journal of Computational and Applied Mathematics, 235(8):2185– 2192, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042710005704.

Cordero:2010:IVN

[CT10] A. Cordero and Juan R. Torregrosa. On interpolation variants of Newton’s method for functions of several variables. Journal of Computational and Applied Mathematics, 234(1):34–43, May 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709008103.

Cong:2011:TST

[CT11] Nguyen Huu Cong and Nguyen Thu Thuy. Twostep-by-two-step PIRKtype PC methods based on Gauss–Legendre collocation points. Journal of Computational and Applied Mathematics, 236(2):225–233, August 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711003566.

Cordero:2011:TSI

[CTV11] Alicia Cordero, Juan R. Torregrosa, and María P. Vassileva. Three-step iterative methods with optimal eighth-order convergence. Journal of Computational and Applied Mathematics, 235(10):3189–3194, March 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711000057.

Cordero:2013:IOC

[CTV13] Alicia Cordero, Juan R.

Torregrosa, and María P. Vassileva. Increasing the order of convergence of iterative schemes for solving nonlinear systems. Journal of Computational and Applied Mathematics, 252(??): 86–94, November 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712005249.

Cekic:2014:PFC

[ÇU14] A. I. Çekiç and Ö. Ugur.

Pricing formulae for constant proportion debt obligation notes: the Laplace transform technique. Journal of Computational and Applied Mathematics, 259 (part B)(??):362–370, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042713003026.

Cui:2010:HOC

[Cui10] Mingrong Cui. High order compact Alternating Direction Implicit method for the generalized sine-Gordon equation. Journal of Computational and Applied Mathematics, 235(3):837– 849, December 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042710004139.

Cui:2014:MMT

[Cui14a] Jintao Cui. Multigrid methods for two-dimensional Maxwell’s equations on graded meshes. Journal of Computational and Applied Mathematics, 255(??):231– 247, January 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042713002707.

Cui:2014:HOC

[Cui14b] Mingrong Cui. A highorder compact exponential scheme for the fractional convection–diffusion equation. Journal of Computational and Applied Mathematics, 255(??):404–416, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713002975.

Chesnokov:2010:DMS

[CV10a] Andrey Chesnokov and Marc Van Barel. A direct method to solve block banded block Toeplitz systems with non-banded Toeplitz blocks. Journal of Computational and Applied Mathematics, 234(5):1485– 1491, July 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710001275.

Ciaurri:2010:UIF

[CV10b] Óscar Ciaurri and Juan L.

Varona. An uncertainty inequality for Fourier–Dunkl series. Journal of Computational and Applied Mathematics, 233(6):1499–1504, January 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709001538.

Contu:2012:FEF

[CvdMS12] Pietro Contu, Cornelis van der Mee, and Sebastiano Seatzu. A finite element frequency domain method for 2D photonic crystals. Journal of Computational and Applied Mathematics, 236(16):3956–3966, October 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712001082.

Chen:2011:EIQ

[CW11a] Fengmin Chen and Patricia J. Y. Wong. Error inequalities for quintic and biquintic discrete Hermite interpolation. Journal of Computational and Applied Mathematics, 235(16):4589–4600, June 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710001883.

Chen:2011:HPA

[CW11b] Jie Chen and Guo-Jin Wang. Hybrid polynomial approximation to higher derivatives of rational curves. Journal of Computational and Applied Mathematics, 235(17):4925– 4936, July 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271100210X.

Chen:2012:EED

[CW12] Fengmin Chen and Patricia J. Y. Wong. Error estimates for discrete spline interpolation: Quintic and biquintic splines. Journal of Computational and Applied Mathematics, 236(15):3835–3854, September 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271200012X.

Chernov:2013:APP

[CW13] N. Chernov and S. Wijewickrema. Algorithms for projecting points onto conics. Journal of Computational and Applied Mathematics, 251(??):8–21, October 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713001398.

Chen:2014:PDS

[CW14a] Fengmin Chen and Patricia J. Y. Wong. On periodic discrete spline interpolation: Quintic and biquintic cases. Journal of Computational and Applied Mathematics, 255(??):282–296, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713002616.

Chen:2014:PEE

[CW14b] Wenbin Chen and Yanqiu Wang. A posteriori error estimate for the H(div) conforming mixed finite element for the coupled Darcy– Stokes system. Journal of Computational and Applied Mathematics, 255(??):502– 516, January 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042713002963.

Chiu:2014:MVP

[CW14c] Mei Choi Chiu and Hoi Ying Wong. Mean-variance portfolio selection with correlation risk. Journal of Computational and Applied Mathematics, 263(??):432– 444, June 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042714000028.

Cui:2011:CNC

[CWQ11] Zhaocheng Cui, Boying Wu, and Shaojian Qu. Combining nonmonotone conic trust region and line search techniques for unconstrained optimization. Journal of Computational and Applied Mathematics, 235(8):2432–2441, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710006102.

Chen:2011:OPF

[CWY11] Zhongying Chen, Tingting Wu, and Hongqi Yang. An optimal 25-point finite difference scheme for the Helmholtz equation with PML. Journal of Computational and Applied Mathematics, 236(6):1240–1258, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004468.

Cui:2011:ENI

[CwYyY11] Xia Cui, Guang wei Yuan, and Jing yan Yue. An efficient nonlinear iteration scheme for nonlinear parabolic-hyperbolic system. Journal of Computational and Applied Mathematics, 236(2):253–264, August 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711003657.

Cheng:2009:FDF

[CXH09] Wanyou Cheng, Yunhai Xiao, and Qing-Jie Hu. A family of derivative-free conjugate gradient methods for large-scale nonlinear systems of equations. Journal of Computational and Applied Mathematics, 224(1):11–19, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/S0377042708001490.

See comments [ZJ12].

Cen:2010:SOH

[CXL10] Zhongdi Cen, Aimin Xu, and Anbo Le. A secondorder hybrid finite difference scheme for a system of singularly perturbed initial value problems. Journal of Computational and Applied Mathematics, 234(12):3445–3457, October 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271000261X.

Chen:2014:HOF

[CXZ14] Zhongying Chen, Yuesheng Xu, and Yuanyuan Zhang. Higher-order finite volume methods II: Inf-sup condition and uniform local ellipticity. Journal of Computational and Applied Mathematics, 265(??):96–109, August 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713001854.

Chang:2011:SDD

[CY11] Huibin Chang and Danping Yang. A Schwarz domain decomposition method with gradient projection for optimal control governed by elliptic partial differential equations. Journal of Computational and Applied Mathematics, 235(17):5078– 5094, July 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711002317.

Cao:2013:MDS

[CYJ13] Yang Cao, Lin-Quan Yao, and Mei-Qun Jiang. A modified dimensional split preconditioner for generalized saddle point problems. Journal of Computational and Applied Mathematics, 250(??):70–82, October 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713000873.

Cao:2014:HBM

[CYJS14] Jianzhi Cao, Rong Yuan, Haijun Jiang, and Juan Song. Hopf bifurcation and multiple periodic solutions in a damped harmonic oscillator with delayed feedback. Journal of Computational and Applied Mathematics, 263(??): 14–24, June 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713006456.

Cui:2010:NIM

[CyY10] Xia Cui and Jing yan Yue. A nonlinear iteration method for solving a two-dimensional nonlinear coupled system of parabolic and hyperbolic equations. Journal of Computational and Applied Mathematics, 234(2):343– 364, May 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709008450.

Cui:2011:NSH

[CyYwY11] Xia Cui, Jing yan Yue, and Guang wei Yuan. Nonlinear scheme with high accuracy for nonlinear coupled parabolic-hyperbolic system. Journal of Computational and Applied Mathematics, 235(12):3527–3540, April 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711000847.

Chen:2011:TBS

[CYZ^{+}11] Wenyu Chen, Rongdong Yu, Jianmin Zheng, Yiyu Cai, and Chikit Au. Triangular Bézier sub-surfaces on a triangular Bézier surface. Journal of Computational and Applied Mathematics, 235(17):5001–5016, July 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271100224X.

Casulli:2012:ISM

[CZ12a] Vincenzo Casulli and Paola Zanolli. Iterative solutions of mildly nonlinear systems. Journal of Computational and Applied Mathematics, 236(16):3937–3947, October 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712001094.

Chen:2012:BBV

[CZ12b] Hao Chen and Chengjian Zhang. Block boundary value methods for solving Volterra integral and integro-differential equations. Journal of Computational and Applied Mathematics, 236(11):2822–2837, May 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712000295.

Cao:2013:EMS

[CZ13] Wanrong Cao and Zhongqiang Zhang. On exponential mean-square stability of two-step Maruyama methods for stochastic delay differential equations. Journal of Computational and Applied Mathematics, 245(??): 182–193, June 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042713000198.

Chu:2014:CCT

[CZ14] Licai Chu and Xiao-Ming Zeng. Constructing curves and triangular patches by beta functions. Journal of Computational and Applied Mathematics, 260(??):191– 200, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713004809.

Chao:2014:OPG

[CZL14] Zhen Chao, Naimin Zhang, and Yunzeng Lu. Optimal parameters of the generalized symmetric SOR method for augmented systems. Journal of Computational and Applied Mathematics, 266(??):52–60, August 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271400051X.

Cao:2014:HAF

[CZWD14] Jun Cao, Guofang Zhou, Cheng Wang, and Xinzhuang Dong. A hybrid adaptive finite difference method powered by a posteriori error estimation technique. Journal of Computational and Applied Mathematics, 259 (part A)(??):117–128, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713001866.

Cao:2014:AVF

[CZZ14] Waixiang Cao, Zhimin Zhang, and Qingsong Zou. Analysis of a p-version finite volume method for 1D elliptic problems. Journal of Computational and Applied Mathematics, 265(??):17– 32, August 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713004998.

DelosReyes:2010:NST

[DA10] Juan Carlos De los Reyes and Sergio González Andrade. Numerical simulation of two-dimensional Bingham fluid flow by semismooth Newton methods. Journal of Computational and Applied Mathematics, 235(1):11–32, November 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271000110X.

Deliceoglu:2014:TFS

[DA14] A. Deliceoğlu and S. Han Aydin. Topological flow structures in an L-shaped cavity with horizontal motion of the upper lid. Journal of Computational and Applied Mathematics, 259 (part B)(??):937–943, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713005517.

Dehghan:2013:TNS

[DAE13] Mehdi Dehghan, S. Aryanmehr, and M. R. Eslahchi. A technique for the numerical solution of initial-value problems based on a class of Birkhoff-type interpolation method. Journal of Computational and Applied Mathematics, 244(??):125– 139, May 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712005067.

deAraujo:2010:SFE

[dAGR10] J. H. Carneiro de Araujo, P. D. Gomes, and V. Ruas. Study of a finite element method for the time-dependent generalized Stokes system associated with viscoelastic flow. Journal of Computational and Applied Mathematics, 234(8):2562–2577, August 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710001834.

Duran:2011:ELT

[DAH11] Esra Akdeniz Duran, Fikri Akdeniz, and Hongchang Hu. Efficiency of a Liutype estimator in semiparametric regression models. Journal of Computational and Applied Mathematics, 235(5):1418–1428, January 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710004772.

Dai:2011:CEE

[Dai11]  Xiaoxia Dai.  Corrigen-

dum to “Error estimates for an operator-splitting method for Navier–Stokes equations: Second-order schemes” [J. Comput. Appl. Math. 231 (2) (2009) 696– 704]. Journal of Computational and Applied Mathematics, 235(9):3110–3111, March 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/S0377042710006448.

See [DSC09].

Deng:2011:UIN

[DAW11] Jian Deng, Cristina Anton, and Yau Shu Wong. Uncertainty investigations in nonlinear aeroelastic systems. Journal of Computational and Applied Mathematics, 235(13):3910–3920, May 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711000483.

Dax:2010:MNA

[Dax10] Achiya Dax. A minimum norm approach for low-rank approximations of a matrix. Journal of Computational and Applied Mathematics, 234(11):3091–3103, October 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710000737.

Degroote:2010:BSI

[DBHV10] J. Degroote, P. Bruggeman, R. Haelterman, and J. Vierendeels. Bubble simulations with an interface tracking technique based on a partitioned fluidstructure interaction algorithm. Journal of Computational and Applied Mathematics, 234(7):2303–2310, August 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709005676.

Du:2010:PPS

[DBZ10] Bo Du, Chuanzhi Bai, and Xiangkui Zhao. Problems of periodic solutions for a type of Duffing equation with state-dependent delay. Journal of Computational and Applied Mathematics, 233(11):2807–2813, April 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709007675.

Dhia:2010:ETH

[DC10] A. S. Bonnet-Ben Dhia and P. Ciarlet, Jr. Erratum to “Time harmonic wave diffraction problems in materials with sign-shifting coefficients” [J. Comput. Appl. Math. doi:10.1016/j.cam.2009.08.041]. Journal of Computational and Applied Mathematics, 234(8):2616, August 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704271000186X. See

[DCZ10].

Ding:2011:AAC

[DC11] Zhonghai Ding and Goong Chen. Analysis and algorithms for the computation of the excited states of helium. Journal of Computational and Applied Mathematics, 235(8):2041–2062, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710005601.

Danovaro:2014:HAC

[DCG^{+}14] E. Danovaro, A. Clematis, A. Galizia, G. Ripepi, A. Quarati, and D. D’Agostino. Heterogeneous architectures for computational intensive applications: a cost-effectiveness analysis. Journal of Computational and Applied Mathematics, 270(??):63– 77, November 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042714001083.

DeStaelen:2013:BIU

[DCGS13] R. H. De Staelen, G. Crevecoeur, T. Goessens, and M. Slodicka. Bayesian inference in the uncertain EEG problem including local information and a sensor correlation matrix. Journal of Computational and Applied Mathematics, 252(??):177– 182, November 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704271200547X.

Deng:2013:DBS

[DCJ13a] Jiansong Deng, Falai Chen, and Liangbing Jin. Dimensions of biquadratic spline spaces over T -meshes. Journal of Computational and Applied Mathematics, 238(??):68–94, January 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003421.

Dong:2013:MHP

[DCJ13b] Chunhuan Dong, Zhong Chen, and Wei Jiang. A modified homotopy perturbation method for solving the nonlinear mixed Volterra–Fredholm integral equation. Journal of Computational and Applied Mathematics, 239(??):359– 366, February 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712003640.

Dhia:2010:THW

[DCZ10] A. S. Bonnet-Ben Dhia, P. Ciarlet, Jr., and C. M.

Zwölf. Time harmonic wave diffraction problems in materials with sign-shifting coefficients. Journal of Computational and Applied Mathematics, 234(6):1912– 1919, July 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709005196.

See erratum [DC10].

Deugoue:2011:TDG

[DD11] G. Deugoue and J. K. Djoko. On the time discretization for the globally modified three dimensional Navier–Stokes equations. Journal of Computational and Applied Mathematics, 235(8):2015–2029, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710005546.

Dieci:2014:CFS

[DD14a] Luca Dieci and Fabio Difonzo. A comparison of Filippov sliding vector fields in codimension 2. Journal of Computational and Applied Mathematics, 262(??):161– 179, May 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/S0377042713006225.

See corrigendum [DD14b].

Dieci:2014:CCF

[DD14b] Luca Dieci and Fabio Difonzo. Corrigendum to “A comparison of Filippov sliding vector fields in codimension 2” [J. Comput. Appl. Math. 262 (2014) 161– 179]. Journal of Computational and Applied Mathematics, 272(??):273, December 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714002787.

See [DD14a].

DelPrete:2014:SAS

[DDDSC14] Vincenza Del Prete, Fabio Di Benedetto, Marco Donatelli, and Stefano Serra-Capizzano. Symbol approach in a signal-restoration problem involving block Toeplitz matrices. Journal of Computational and Applied Mathematics, 272 (??):399–416, December 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042713002938.

DellAcqua:2014:PIR

[DDE14] Pietro Dell’Acqua, Marco Donatelli, and Claudio Estatico. Preconditioners for image restoration by reblurring techniques. Journal of Computational and Applied Mathematics, 272(??):313–333, December 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005918.

DeSiqueira:2013:NPC

[DDG13a] Denise De Siqueira, Phillipe R. B. Devloo, and Sônia M. Gomes. A new procedure for the construction of hierarchical high order Hdiv and Hcurl finite element spaces. Journal of Computational and Applied Mathematics, 240(??):204–214, March 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003998.

Derfoul:2013:IPT

[DDG^{+}13b] Ratiba Derfoul, Sebastien Da Veiga, Christian Gout, Carole Le Guyader, and Elodie Tillier. Image processing tools for better incorporation of 4D seismic data into reservoir models. Journal of Computational and Applied Mathematics, 240(??):111–122, March 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003445.

DeLuigi:2010:NSP

[DDM10] C. De Luigi, S. Dumont, and S. Maire. Numerical solution of the Poisson equation over hypercubes using reduced Chebyshev polynomial bases. Journal of Computational and Applied Mathematics, 234(1):181– 191, May 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704270900822X.

DeTeran:2011:PCF

[DDM11] Fernando De Terán, Froilán M.

Dopico, and D. Steven Mackey. Palindromic companion forms for matrix polynomials of odd degree. Journal of Computational and Applied Mathematics, 236(6):1464–1480, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004894.

DeSimone:2014:MFA

[DdS14] V. De Simone and D. di Serafino. A matrix-free approach to build band preconditioners for largescale bound-constrained optimization. Journal of Computational and Applied Mathematics, 268(??):82– 92, October 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042714001319.

Deelstra:2010:MMA

[DDV10] Griselda Deelstra, Ibrahima Diallo, and Michèle Vanmaele. Moment matching approximation of Asian basket option prices. Journal of Computational and Applied Mathematics, 234(4):1006– 1016, June 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709002106.

Deckers:2011:SBB

[DDV^{+}11] E. Deckers, B. Drofmans, B. Van Genechten, B. Bergen, D. Vandepitte, and W. Desmet. Splinebased boundaries: a first step towards generic geometric domain descriptions for efficient midfrequency acoustic analysis using the wave based method. Journal of Computational and Applied Mathematics, 235(8):2679–2693, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006370.

DeBonis:2014:RTI

[De 14] Maria Carmela De Bonis.

Remarks on two integral operators and numerical methods for CSIE. Journal of Computational and Applied Mathematics, 260(??):117– 134, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005189.

Dea:2011:EAH

[Dea11] John R. Dea. An experimental adaptation of higdon-type non-reflecting boundary conditions to linear first-order systems. Journal of Computational and Applied Mathematics, 235(5):1354–1366, January 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710004723.

Deka:2010:FEM

[Dek10] Bhupen Deka. Finite element methods with numerical quadrature for elliptic problems with smooth interfaces. Journal of Computational and Applied Mathematics, 234(2):605–612, May 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710000051.

DAmbrosio:2011:EFT

[DEP11] R. D’Ambrosio, E. Esposito, and B. Paternoster. Exponentially fitted two-step hybrid methods for y′′ = f(x, y). Journal of Computational and Applied Mathematics, 235(16):4888–4897, June 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711000598.

During:2012:HOC

[DF12] Bertram Düring and Michel Fournié. High-order compact finite difference scheme for option pricing in stochastic volatility models. Journal of Computational and Applied Mathematics, 236 (17):4462–4473, November 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704271200177X.

Du:2014:PSN

[DF14] Zengji Du and Zhaosheng Feng. Periodic solutions of a neutral impulsive predator–prey model with Beddington–DeAngelis functional response with delays. Journal of Computational and Applied Mathematics, 258(??):87– 98, March 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713004561.

Domingues:2010:GSI

[DFG^{+}10] Margarete O. Domingues, Paulo J. S. G. Ferreira, Sônia M. Gomes, Anamaria Gomide, José R. Pereira, and Pedro Pinho. Grid structure impact in sparse point representation of derivatives. Journal of Computational and Applied Mathematics, 234(8):2377– 2389, August 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710001330.

deFrutos:2011:PEE

[dFGAN11] Javier de Frutos, Bosco García-Archilla, and Julia Novo. A posteriori error estimations for mixed finiteelement approximations to the Navier–Stokes equations. Journal of Computational and Applied Mathematics, 236(6):1103–1122, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004377.

During:2014:HOC

[DFH14] Bertram Düring, Michel Fournié, and Christof Heuer. High-order compact finite difference schemes for option pricing in stochastic volatility models on nonuniform grids. Journal of Computational and Applied Mathematics, 271(??):247– 266, December 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042714002155.

Davydov:2011:P

[DG11] Oleg Davydov and Tim Goodman. Preface. Journal of Computational and Applied Mathematics, 236(4): 423, September 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042711004614.

Doyeux:2013:STF

[DGC^{+}13] V. Doyeux, Y. Guyot, V. Chabannes, C. Prud’homme, and M. Ismail. Simulation of two-fluid flows using a finite element/level set method. Application to bubbles and vesicle dynamics. Journal of Computational and Applied Mathematics, 246(??):251– 259, July 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712002051.

DeSantis:2012:NPF

[DGG12] D. De Santis, G. Geraci, and A. Guardone. Nodepair finite volume/finite element schemes for the Euler equation in cylindrical and spherical coordinates. Journal of Computational and Applied Mathematics, 236(18):4827–4839, December 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712000635.

Dimov:2010:SSP

[DGI^{+}10] I. Dimov, R. Georgieva, S. Ivanovska, Tz. Ostromsky, and Z. Zlatev. Studying the sensitivity of pollutants’ concentrations caused by variations of chemical rates. Journal of Computational and Applied Mathematics, 235(2):391–402, November 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710003316.

Dyn:2010:AOI

[DGW10] Nira Dyn, Philipp Grohs, and Johannes Wallner. Approximation order of interpolatory nonlinear subdivision schemes. Journal of Computational and Applied Mathematics, 233(7):1697– 1703, February 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042709000934.

DeSchepper:2010:HEV

[DH10] Ann De Schepper and Bart Heijnen. How to estimate the value at risk under incomplete information. Journal of Computational and Applied Mathematics, 233(9):2213–2226, March 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709006852.

Dehghan:2011:TCS

[DH11] Mehdi Dehghan and Masoud Hajarian. Two class of synchronous matrix multisplitting schemes for solving linear complementarity problems. Journal of Computational and Applied Mathematics, 235(15):4325– 4336, June 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711001518.

Drosson:2013:SSI

[DH13] M. Drosson and K. Hillewaert. On the stability of the symmetric interior penalty method for the Spalart–Allmaras turbulence model. Journal of Computational and Applied Mathematics, 246(??):122– 135, July 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003834.

Dorsek:2014:HOS

[DH14] Philipp Dörsek and Eskil Hansen. High order splitting schemes with complex timesteps and their application in mathematical finance. Journal of Computational and Applied Mathematics, 262(??):234–243, May 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003816.

Dhamacharoen:2014:EHM

[Dha14] Ampon Dhamacharoen. An efficient hybrid method for solving systems of nonlinear equations. Journal of Computational and Applied Mathematics, 263(??):59– 68, June 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713006766.

DeLuca:2012:FEB

[DHR12] Jayme De Luca, A. R. Humphries, and Savio B. Rodrigues. Finite element boundary value integration of Wheeler–Feynman electrodynamics. Journal of Computational and Applied Mathematics, 236(13):3319– 3337, July 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712001069.

Dominguez:2011:HSS

[DHS11] Víctor Domínguez, Norbert Heuer, and Francisco-Javier Sayas. Hilbert scales and Sobolev spaces defined by associated Legendre functions. Journal of Computational and Applied Mathematics, 235(12):3481–3501, April 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711000811.

DAzevedo:2014:SLS

[DHSW14] Eduardo D’Azevedo, Zhiang Hu, Shi-Quan Su, and Kwai Wong. Solving a large scale radiosity problem on GPU-based parallel computers. Journal of Computational and Applied Mathematics, 270(??):109–120, November 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714000879.

Durufle:2012:NSV

[DI12] Marc Duruflé and Samer Israwi. A numerical study of variable depth KdV equations and generalizations of Camassa–Holm-like equations. Journal of Computational and Applied Mathematics, 236(17):4149–4165, November 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712002117.

Duan:2010:GHA

[DJ10] Congying Duan and Zhongxiao Jia. A global harmonic Arnoldi method for large non-Hermitian eigenproblems with an application to multiple eigenvalue problems. Journal of Computational and Applied Mathematics, 234(3):845– 860, June 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710000622.

Drobek:2014:SLS

[DJ14] J. Drobek and L. Jokl. On solvability of linear systems generated by the complex variable boundary element method. Journal of Computational and Applied Mathematics, 259 (part A)(??): 216–225, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042713004147.

Dominici:2013:RZH

[DJJ13] D. Dominici, S. J. Johnston, and K. Jordaan. Real zeros of _{2}F_{1} hypergeometric polynomials. Journal of Computational and Applied Mathematics, 247(??):152–161, August 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713000162.

Date:2010:PLS

[DJM10] Paresh Date, Luka Jalen, and Rogemar Mamon. A partially linearized sigma point filter for latent state estimation in nonlinear time series models. Journal of Computational and Applied Mathematics, 233(10):2675– 2682, March 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709007468.

deJesus:2014:CPO

[dJMPPC14] M. N. de Jesus, F. Marcellán, J. Petronilho, and N. C. Pinzón-Cortés. (M, N)- coherent pairs of order (m, k) and Sobolev orthogonal polynomials. Journal of Computational and Applied Mathematics, 256(??):16– 35, January 15, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713003592.

deJesus:2013:SOP

[dJP13] M. N. de Jesus and J. Petronilho. Sobolev orthogonal polynomials and (M, N)-coherent pairs of measures. Journal of Computational and Applied Mathematics, 237(1):83– 101, January 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712002968.

Dostal:2010:STA

[DKH^{+}10] Z. Dostál, T. Kozubek, P. Horyl, T. Brzobohatý, and A. Markopoulos. A scalable TFETI algorithm for two-dimensional multibody contact problems with friction. Journal of Computational and Applied Mathematics, 235(2):403–418, November 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042710003328.

Dolezel:2014:DJB

[DKU14] Ivo Dolezel, Václav Kotlan, and Bohus Ulrych. Design of joint between disk and shaft based on induction shrink fit. Journal of Computational and Applied Mathematics, 270(??): 52–62, November 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042714001095.

Dieci:2012:SNM

[DL12] Luca Dieci and Luciano Lopez. A survey of numerical methods for IVPs of ODEs with discontinuous right-hand side. Journal of Computational and Applied Mathematics, 236(16):3967– 3991, October 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042712000684.

Dong:2013:HPB

[DL13] Tao Dong and Xiaofeng Liao. Hopf–Pitchfork bifurcation in a simplified BAM neural network model with multiple delays. Journal of Computational and Applied Mathematics, 253(??):222– 234, December 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713002136.

Delanoue:2014:NAC

[DL14a] Nicolas Delanoue and Sébastien Lagrange. A numerical approach to compute the topology of the apparent contour of a smooth mapping from R^{2} to R^{2}. Journal of Computational and Applied Mathematics, 271(??):267–284, December 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714001812.

Liu:2014:PAQ

[dL14b] Lan dong Liu. Perturbation analysis of a quadratic matrix equation associated with an M-matrix. Journal of Computational and Applied Mathematics, 260(??): 410–419, April 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713005554.

Das:2014:EFO

[DLG14] Sambit Das, Wenyuan Liao, and Anirudh Gupta. An efficient fourth-order low dispersive finite difference scheme for a 2-D acoustic wave equation. Journal of Computational and Applied Mathematics, 258(??):151– 167, March 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042713004548.

Duran:2013:SGM

[DLP13] R. G. Durán, A. L. Lombardi, and M. I. Prieto. Supercloseness on graded meshes for Q_{1} finite element approximation of a reaction–diffusion equation. Journal of Computational and Applied Mathematics, 242(??):232–247, April 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712004232.

Dhaene:2014:MDM

[DLSV14] Jan Dhaene, Daniël Linders, Wim Schoutens, and David Vyncke. A multivariate dependence measure for aggregating risks. Journal of Computational and Applied Mathematics, 263(??): 78–87, June 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713006808.

Duan:2014:LRA

[DLWZ14] Xuefeng Duan, Jiaofen Li, Qingwen Wang, and Xinjun Zhang. Low rank approximation of the symmetric positive semidefinite matrix. Journal of Computational and Applied Mathematics, 260(??):236– 243, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005359.

Davydov:2010:TAC

[DM10a] Oleg Davydov and Carla Manni. A tension approach to controlling the shape of cubic spline surfaces on FVS triangulations. Journal of Computational and Applied Mathematics, 233(7):1674– 1684, February 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709000910.

Duenas:2010:ABL

[DM10b] Herbert Dueñas and Francisco Marcellán. Asymptotic behaviour of Laguerre– Sobolev-type orthogonal polynomials. A nondiagonal case. Journal of Computational and Applied Mathematics, 235(4):998–1007, December 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042709004610.

Donatelli:2012:FDA

[DM12] Marco Donatelli and Nicola Mastronardi. Fast deconvolution with approximated PSF by RSTLS with antireflective boundary conditions. Journal of Computational and Applied Mathematics, 236(16):3992–4005, October 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712001537.

Dosiyev:2010:BMP

[DMB10] A. A. Dosiyev, Z. Mazhar, and S. C. Buranay. Block method for problems on Lshaped domains. Journal of Computational and Applied Mathematics, 235(3):805– 816, December 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042710003936.

Diaz:2012:ASR

[DME^{+}12] F. Díaz, G. Montero, J. M. Escobar, E. Rodríguez, and R. Montenegro. An adaptive solar radiation numerical model. Journal of Computational and Applied Mathematics, 236(18):4611– 4622, December 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712001781.

Diaz-Mendoza:2010:ORO

[DMGVPN10] C. Díaz-Mendoza, P. González-Vera, M. Jiménez Paiz, and O. Njåstad. Orthogonality and recurrence for ordered Laurent polynomial sequences. Journal of Computational and Applied Mathematics, 235(4):982–997, December 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710004103.

Dhia:2010:LMN

[DMMP10] A. S. Bonnet-Ben Dhia, J. F. Mercier, F. Millot, and S. Pernet. A low-Mach number model for timeharmonic acoustics in arbitrary flows. Journal of Computational and Applied Mathematics, 234(6):1868– 1875, July 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709005147.

Diogo:2013:FDC

[DMR13] Teresa Diogo, Jingtang Ma, and Magda Rebelo. Fully discretized collocation methods for nonlinear singular Volterra integral equations. Journal of Computational and Applied Mathematics, 247(??):84–101, August 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713000204.

Ding:2013:NAP

[DN13] Hui-Sheng Ding and Juan J. Nieto. A new approach for positive almost periodic solutions to a class of Nicholson’s blowflies model. Journal of Computational and Applied Mathematics, 253(??):249–254, December 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713002148.

deOliveira:2010:VCT

[dO10] Valeriano Antunes de Oliveira. Vector continuous-time programming without differentiability. Journal of Computational and Applied Mathematics, 234(3):924– 933, June 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271000083X.

Demirci:2012:MSD

[DO12] Elif Demirci and Nuri Ozalp. A method for solving differential equations of fractional order. Journal of Computational and Applied Mathematics, 236(11):2754– 2762, May 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712000064.

Dominici:2010:FIO

[Dom10] Diego Dominici. Fisher information of orthogonal polynomials I. Journal of Computational and Applied Mathematics, 233(6):1511– 1518, January 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709001551.

Dominici:2011:TSK

[Dom11] Diego Ernesto Dominici. On Taylor series and Kapteyn series of the first and second type. Journal of Computational and Applied Mathematics, 236(1):39–48, August 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711001373.

Dominguez:2014:FCC

[Dom14] Víctor Domínguez. Filon– Clenshaw–Curtis rules for a class of highly-oscillatory integrals with logarithmic singularities. Journal of Computational and Applied Mathematics, 261(??):299– 319, May 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713006316.

Dosiyev:2014:BGM

[Dos14] A. A. Dosiyev. The blockgrid method for the approximation of the pure second order derivatives for the solution of Laplace’s equation on a staircase polygon. Journal of Computational and Applied Mathematics, 259 (part A)(??): 14–23, March 15, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713001301.

Douglass:2012:LBE

[Dou12] Rod W. Douglass. Laplace– Beltrami enhancement for unstructured two-dimensional meshes having dendritic elements and boundary node movement. Journal of Computational and Applied Mathematics, 236(18):4952– 4963, December 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711005024.

Delgado:2010:REE

[DP10a] J. Delgado and J. M. Peña.

Running error for the evaluation of rational Bézier surfaces. Journal of Computational and Applied Mathematics, 233(7):1685–1696, February 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709000922.

Dubrulle:2010:RTM

[DP10b] A. A. Dubrulle and B. N.

Parlett. Revelations of a transposition matrix. Journal of Computational and Applied Mathematics, 233(5):1217–1219, January 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709001745.

Delgado:2011:REE

[DP11] J. Delgado and J. M. Peña.

Running error for the evaluation of rational Bézier surfaces through a robust algorithm. Journal of Computational and Applied Mathematics, 235(7):1781–1789, February 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710002396.

Dzunic:2012:GMR

[DP12] Jovana Dzunić and Miodrag S. Petković. On generalized multipoint rootsolvers with memory. Journal of Computational and Applied Mathematics, 236 (11):2909–2920, May 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042712000465.

DAmbrosio:2014:EFS

[DP14a] R. D’Ambrosio and B. Paternoster. Exponentially fitted singly diagonally implicit Runge–Kutta methods. Journal of Computational and Applied Mathematics, 263(??):277–287, June 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271300695X.

DAmbrosio:2014:SGN

[DP14b] Raffaele D’Ambrosio and Beatrice Paternoster. P - stable general Nyström methods for y′′ = f(y(t)). Journal of Computational and Applied Mathematics, 262(??):271–280, May 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0377042713004755.

Dzunic:2014:GBM

[DP14c] Jovana Dzunić and Miodrag S. Petković. On generalized biparametric multipoint root finding methods with memory. Journal of Computational and Applied Mathematics, 255(??):362– 375, January 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713002884.

Deelstra:2010:PHA

[DPV10] Griselda Deelstra, Alexandre Petkovic, and Michèle Vanmaele. Pricing and hedging Asian basket spread options. Journal of Computational and Applied Mathematics, 233(11):2814–2830, April 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709007687.

Dong:2014:AMG

[DQST14] Hao Dong, Zhonghua Qiao, Shuyu Sun, and Tao Tang. Adaptive moving grid methods for two-phase flow in porous media. Journal of Computational and Applied Mathematics, 265(??):139– 150, August 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042713004822.

Deflorian:2014:GIS

[DR14a] Michael Deflorian and Matthias Rungger. Generalization of an input-to-state stability preserving Runge–Kutta method for nonlinear control systems. Journal of Computational and Applied Mathematics, 255(??):346– 352, January 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042713002926.

Donatelli:2014:SSR

[DR14b] M. Donatelli and L. Reichel. Square smoothing regularization matrices with accurate boundary conditions. Journal of Computational and Applied Mathematics, 272(??):334–349, December 15, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042713004202.

Dykes:2014:SGC

[DR14c] L. Dykes and L. Reichel. Simplified GSVD computations for the solution of linear discrete ill-posed problems. Journal of Computational and Applied Mathematics, 255(??):15–27, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713002057.

Dratman:2010:ASC

[Dra10] Ezequiel Dratman. Approximation of the solution of certain nonlinear ODEs with linear complexity. Journal of Computational and Applied Mathematics, 233(9):2339–2350, March 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709006979.

Dratman:2013:EAS

[Dra13] Ezequiel Dratman. Efficient approximation of the solution of certain nonlinear reaction–diffusion equations with large absorption. Journal of Computational and Applied Mathematics, 238(??):180–202, January 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003548.

Demichelis:2010:CRA

[DS10a] Vittoria Demichelis and Paul Sablonnière. Cubature rule associated with a discrete blending sum of quadratic spline quasiinterpolants. Journal of Computational and Applied Mathematics, 235(1):174– 185, November 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710003079.

Durand:2010:NSM

[DS10b] Stephane Durand and Marián Slodicka. A numerical scheme for the Maxwell equations in the quasi-static regime with a non-local source. Journal of Computational and Applied Mathematics, 233(12):3157–3166, April 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709007456.

Du:2011:NSE

[DS11] Kui Du and Weiwei Sun. Numerical solution of electromagnetic scattering from a large partly covered cavity. Journal of Computational and Applied Mathematics, 235(13):3791–3806, May 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711000379.

Dehghan:2012:NSN

[DS12a] Mehdi Dehghan and Rezvan Salehi. The numerical solution of the non-linear integro-differential equations based on the meshless method. Journal of Computational and Applied Mathematics, 236(9):2367– 2377, March 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711005929.

Durand:2012:RCM

[DS12b] Stephane Durand and Marián Slodicka. RETRACTED: Convergence of the mixed finite element method for Maxwell’s equations with non-linear conductivity. Journal of Computational and Applied Mathematics, 236 (18):4893–4908, December 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042712000076. See

retraction notice [DS13c].

Davydov:2013:NSF

[DS13a] Oleg Davydov and Abid Saeed. Numerical solution of fully nonlinear elliptic equations by Böhmer’s method. Journal of Computational and Applied Mathematics, 254(??):43–54, December 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713001179.

DeMarchi:2013:NSB

[DS13b] Stefano De Marchi and Gabriele Santin. A new stable basis for radial basis function interpolation. Journal of Computational and Applied Mathematics, 253(??):1–13, December 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713001751.

Durand:2013:RNC

[DS13c] Stephane Durand and Marián Slodicka. Retraction notice to: “Convergence of the mixed finite element method for Maxwell’s equations with non-linear conductivity” [J. Comput. Appl. Math. 236 (18) (2012) 4893–4908]. Journal of Computational and Applied Mathematics, 244(??): 157, May 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712005250.

See [DS12b].

Dehghan:2014:MLP

[DS14] Mehdi Dehghan and Rezvan Salehi. A meshless local Petrov–Galerkin method for the timedependent Maxwell equations. Journal of Computational and Applied Mathematics, 268(??):93–110, October 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714000892.

Dai:2009:EEO

[DSC09] Xiaoxia Dai, Jie Sun, and Xiaoliang Cheng. Error estimates for an operatorsplitting method for Navier– Stokes equations: Secondorder schemes. Journal of Computational and Applied Mathematics, 231(2):696– 704, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042709002696. See

corrigendum [Dai11].

Deng:2012:ECS

[DSL12] Lih-Yuan Deng, Jyh-Jen Horng Shiau, and Henry Horng-Shing Lu. Efficient computer search of large-order multiple recursive pseudorandom number generators. Journal of Computational and Applied Mathematics, 236(13):3228–3237, July 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712000805.

Das:2014:LSP

[DSN14] Payel Das, Mitali Madhumita Sahani, and Gnaneshwar Nelakanti. Legendre spectral projection methods for Urysohn integral equations. Journal of Computational and Applied Mathematics, 263(??):88– 102, June 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713006729.

Deano:2010:CPT

[DST10] Alfredo Deaño, Javier Segura, and Nico M. Temme. Computational properties of three-term recurrence relations for Kummer functions. Journal of Computational and Applied Mathematics, 233(6):1505–1510, January 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704270900154X.

Du:2011:BIM

[DSY^{+}11] L. Du, T. Sogabe, B. Yu, Y. Yamamoto, and S.-L. Zhang. A block IDR(s) method for nonsymmetric linear systems with multiple right-hand sides. Journal of Computational and Applied Mathematics, 235(14):4095– 4106, May 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711001233.

Du:2011:VIM

[DSZ11] Lei Du, Tomohiro Sogabe, and Shao-Liang Zhang. A variant of the IDR(s) method with the quasiminimal residual strategy. Journal of Computational and Applied Mathematics, 236(5):621–630, October 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004225.

Dawson:2013:SDF

[DV13] Clint Dawson and Juha H.

Videman. A streamline diffusion finite element method for the viscous shallow water equations. Journal of Computational and Applied Mathematics, 251(??):1–7, October 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713001234.

Ding:2010:APA

[DW10] Xie Ping Ding and Zhong Bao Wang. The auxiliary principle and an algorithm for a system of generalized setvalued mixed variationallike inequality problems in Banach spaces. Journal of Computational and Applied Mathematics, 233(11):2876– 2883, April 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042709007808.

Dai:2011:MCD

[DW11] Zhifeng Dai and Fenghua Wen. A modified CG– DESCENT method for unconstrained optimization. Journal of Computational and Applied Mathematics, 235(11):3332–3341, April 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711000574.

Ding:2010:WFD

[DXL10] Zhiqing Ding, Aiguo Xiao, and Min Li. Weighted finite difference methods for a class of space fractional partial differential equations with variable coefficients. Journal of Computational and Applied Mathematics, 233(8):1905–1914, February 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709006499.

Deng:2011:IPI

[DY11] Zui-Cha Deng and Liu Yang. An inverse problem of identifying the coefficient of first-order in a degenerate parabolic equation. Journal of Computational and Applied Mathematics, 235(15):4404–4417, June 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271100183X.

Davydov:2013:RPQ

[DY13] Oleg Davydov and Wee Ping Yeo. Refinable C^{2} piecewise quintic polynomials on Powell–Sabin-12 triangulations. Journal of Computational and Applied Mathematics, 240(??):62– 73, March 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003287.

Dai:2014:MCG

[DZW14] Ruxin Dai, Jun Zhang, and Yin Wang. Multiple coarse grid acceleration for multiscale multigrid computation. Journal of Computational and Applied Mathematics, 269(??):75–85, October 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714001708.

Du:2014:RKM

[DZZ14] Hong Du, Guoliang Zhao, and Chunyan Zhao. Reproducing kernel method for solving Fredholm integrodifferential equations with weakly singularity. Journal of Computational and Applied Mathematics, 255(??): 122–132, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042713001829.

Eskhuvatov:2011:ACT

[EALA11] Z. K. Eskhuvatov, A. Ahmedov, N. M. A. Nik Long, and N. J. Amalina. Approximating Cauchy-type singular integral by an automatic quadrature scheme. Journal of Computational and Applied Mathematics, 235(16): 4675–4686, June 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042710004267.

El-Amin:2013:NDI

[EASS13] M. F. El-Amin, Amgad Salama, and Shuyu Sun. Numerical and dimensional investigation of two-phase countercurrent imbibition in porous media. Journal of Computational and Applied Mathematics, 242(??):285– 296, April 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712004086.

Ebaid:2011:NAN

[Eba11]  Abdelhalim Ebaid.  A

new analytical and numerical treatment for singular two-point boundary value problems via the Adomian decomposition method. Journal of Computational and Applied Mathematics, 235(8):1914–1924, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271000498X.

Esteves:2013:BSA

[EC13] Jorge Sá Esteves and José Craveirinha. On a bicriterion server allocation problem in a multidimensional Erlang loss system. Journal of Computational and Applied Mathematics, 252(??):103–119, November 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713001209.

Erpicum:2010:DBF

[EDAP10] S. Erpicum, B. J. Dewals, P. Archambeau, and M. Pirotton. Dam break flow computation based on an efficient flux vector splitting. Journal of Computational and Applied Mathematics, 234(7):2143–2151, August 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709005500.

Eslahchi:2014:ACM

[EDP14] M. R. Eslahchi, Mehdi Dehghan, and M. Parvizi. Application of the collocation method for solving nonlinear fractional integro-differential equations. Journal of Computational and Applied Mathematics, 257(??):105–128, February 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713003889.

Eksi:2014:PHI

[EF14] Zehra Eksi and Damir Filipović. Pricing and hedging of inflation-indexed bonds in an affine framework. Journal of Computational and Applied Mathematics, 259 (part B)(??):452–463, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005670.

Engel:2011:MMC

[EG11] M. Engel and M. Griebel.

A multigrid method for constrained optimal control problems. Journal of Computational and Applied Mathematics, 235(15):4368– 4388, June 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711001786.

ElGhami:2012:IPM

[EGBS12] M. El Ghami, Z. A. Guennoun, S. Bouali, and T. Steihaug. Interior-point methods for linear optimization based on a kernel function with a trigonometric barrier term. Journal of Computational and Applied Mathematics, 236(15):3613–3623, September 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711002937.

El-Gamel:2014:STD

[EGES14] Mohamed El-Gamel and Atallah El-Shenawy. The solution of a time-dependent problem by the B-spline method. Journal of Computational and Applied Mathematics, 267(??):254–265, September 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042714000806.

Ezquerro:2012:MSN

[EGH12] J. A. Ezquerro, D. González, and M. A. Hernández. Majorizing sequences for Newton’s method from initial value problems. Journal of Computational and Applied Mathematics, 236(9):2246– 2258, March 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711005826.

El-Gebeily:2010:MIT

[EGON10] M. A. El-Gebeily, Donal O’Regan, and J. J. Nieto. A monotone iterative technique for stationary and time dependent problems in Banach spaces. Journal of Computational and Applied Mathematics, 233(9):2395– 2404, March 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709007225.

Escribano:2011:HMS

[EGST11] C. Escribano, A. Giraldo, M. A. Sastre, and E. Torrano. Hessenberg matrix for sums of Hermitian positive definite matrices and weighted shifts. Journal of Computational and Applied Mathematics, 236(1):98– 106, August 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711002883.

Ezquerro:2010:EGR

[EHR10] J. A. Ezquerro, M. A.

Hernández, and N. Romero. An extension of Gander’s result for quadratic equations. Journal of Computational and Applied Mathematics, 234(4):960–971, June 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709000156.

Ehrenmark:2011:NCG

[Ehr11a] Ulf Ehrenmark. A note on the computation of a Green’s function for short wavelength water waves over a plane incline. Journal of Computational and Applied Mathematics, 236(6):1667–1674, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711005231.

Ezquerro:2011:SNI

[EHR11b] J. A. Ezquerro, M. A. Hernández, and N. Romero. Solving nonlinear integral equations of Fredholm type with high order iterative methods. Journal of Computational and Applied Mathematics, 236(6):1449– 1463, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042711004882.

Ezquerro:2013:SMB

[EHRV13] J. A. Ezquerro, M. A. Hernández, N. Romero, and A. I. Velasco. On Steffensen’s method on Banach spaces. Journal of Computational and Applied Mathematics, 249(??):9–23, September 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713000642.

Edemskiy:2014:LCQ

[EI14] V. Edemskiy and A. Ivanov. Linear complexity of quaternary sequences of length pq with low autocorrelation. Journal of Computational and Applied Mathematics, 259 (part B)(??):555–560, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003993.

Ezziani:2010:LTS

[EJ10] Abdelaâziz Ezziani and Patrick Joly. Local time stepping and discontinuous Galerkin methods for symmetric first order hyperbolic systems. Journal of Computational and Applied Mathematics, 234(6):1886–1895, July 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709005160.

Elliott:2011:EEG

[EJJ11] David Elliott, Peter R.

Johnston, and Barbara M. Johnston. Estimates of the error in Gauss–Legendre quadrature for double integrals. Journal of Computational and Applied Mathematics, 236(6):1552–1561, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004985.

Ebaid:2011:NTE

[EK11] Abdelhalim Ebaid and S. M. Khaled. New types of exact solutions for nonlinear Schrödinger equation with cubic nonlinearity. Journal of Computational and Applied Mathematics, 235(8):1984–1992, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710005509.

See comment [KRS11].

Elfoutayeni:2012:UVD

[EK12] Youssef Elfoutayeni and Mohamed Khaladi. Using vector divisions in solving the linear complementarity problem. Journal of Computational and Applied Mathematics, 236(7):1919– 1925, January 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042711005711.

ElWassouli:2011:EEH

[El 11] Fouzia El Wassouli. An explicit expression for the Harish-Chandra c-function for the homogeneous space SO(2, n)/S(O(2) × O(n)). Journal of Computational and Applied Mathematics, 235(18):5333–5341, July 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042708000034.

Eshkuvatov:2011:ASI

[EL11] Z. K. Eshkuvatov and N. M.

A. Nik Long. Approximating the singular integrals of Cauchy type with weight function on the interval. Journal of Computational and Applied Mathematics, 235(16):4742–4753, June 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710005054.

Eshkuvatov:2010:NEC

[ELA10] Z. K. Eshkuvatov, N. M.

A. Nik Long, and M. Abdulkawi. Numerical evaluation for Cauchy type singular integrals on the interval. Journal of Computational and Applied Mathematics, 233(8):1995–2001, February 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704270900658X.

Abd-el-Malek:2014:LGA

[eMA14] Mina B. Abd el Malek and Amr M. Amin. Lie group analysis of nonlinear inviscid flows with a free surface under gravity. Journal of Computational and Applied Mathematics, 258(??):17– 29, March 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713003968.

Emms:2011:PGI

[Emm11] Paul Emms. Pricing general insurance in a reactive and competitive market. Journal of Computational and Applied Mathematics, 236(6):1314–1332, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711004602.

Erickson:2014:SHO

[EN14] Brittany A. Erickson and Jan Nordström. Stable, high order accurate adaptive schemes for long time, highly intermittent geophysics problems. Journal of Computational and Applied Mathematics, 271(??):328–338, December 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042714002180.

ElKhoury:2014:AIS

[ENS14] G. El Khoury, Yu. M. Nechepurenko, and M. Sadkane. Acceleration of inverse subspace iteration with Newton’s method. Journal of Computational and Applied Mathematics, 259 (part A)(??):205–215, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003440.

Estatico:2012:NRP

[ENSCS12] Claudio Estatico, Eric Ngondiep, Stefano Serra-Capizzano, and Debora Sesana. A note on the (regularizing) preconditioning of g-Toeplitz sequences via g-circulants. Journal of Computational and Applied Mathematics, 236(8):2090– 2111, February 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711005127.

Einkemmer:2014:ASS

[EO14] Lukas Einkemmer and Alexander Ostermann. An almost symmetric Strang splitting scheme for the construction of high order composition methods. Journal of Computational and Applied Mathematics, 271(??):307–318, December 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042714002143.

Eryilmaz:2013:LBD

[Ery13a] Serkan Eryilmaz. On the lifetime behavior of a discrete time shock model. Journal of Computational and Applied Mathematics, 237(1):384–388, January 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712002439.

Eryilmaz:2013:SDO

[Ery13b] Serkan Eryilmaz. On the sums of distributions of order statistics from exchangeable random variables. Journal of Computational and Applied Mathematics, 253(??):204–207, December 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713002124.

Eryilmaz:2014:CRI

[Ery14] Serkan Eryilmaz. Computing reliability indices of repairable systems via signature. Journal of Computational and Applied Mathematics, 260(??):229– 235, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713004780.

Elliott:2014:PBO

[ESC14a] Robert J. Elliott, Tak Kuen Siu, and Leunglung Chan. On pricing barrier options with regime switching. Journal of Computational and Applied Mathematics, 256(??):196–210, January 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003786.

Escudero:2014:HMS

[Esc14b] Juan García Escudero. Hypersurfaces with many A_{j}- singularities: Explicit constructions. Journal of Computational and Applied Mathematics, 259 (part A) (??):87–94, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042713001726.

Elgindy:2013:SBV

[ESM13a] K. T. Elgindy and Kate A.

Smith-Miles. Solving boundary value problems, integral, and integro-differential equations using Gegenbauer integration matrices. Journal of Computational and Applied Mathematics, 237(1):307–325, January 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712002348.

Elgindy:2013:FAS

[ESM13b] Kareem T. Elgindy and Kate A. Smith-Miles. Fast, accurate, and small-scale direct trajectory optimization using a Gegenbauer transcription method. Journal of Computational and Applied Mathematics, 251(??): 93–116, October 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042713001404.

Elgindy:2013:OGQ

[ESM13c] Kareem T. Elgindy and Kate A. Smith-Miles. Optimal Gegenbauer quadrature over arbitrary integration nodes. Journal of Computational and Applied Mathematics, 242(??):82– 106, April 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712004578.

Ern:2010:GRD

[ESV10] Alexandre Ern, Annette F. Stephansen, and Martin Vohralík. Guaranteed and robust discontinuous Galerkin a posteriori error estimates for convection– diffusion–reaction problems. Journal of Computational and Applied Mathematics, 234(1):114–130, May 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709008176.

Eryilmaz:2014:DMG

[EX14] Serkan Eryilmaz and Min Xie. Dynamic modeling of general three-state k-out-of-n: G systems: Permanentbased computational results. Journal of Computational and Applied Mathematics, 272(??):97–106, December 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714002428.

Engquist:2010:FDA

[EY10] Björn Engquist and Lexing Ying. Fast directional algorithms for the Helmholtz kernel. Journal of Computational and Applied Mathematics, 234(6):1851–1859, July 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709005123.

Eryilmaz:2011:MED

[EY11] Serkan Eryilmaz and Femin Yalçin. On the mean and extreme distances between failures in Markovian binary sequences. Journal of Computational and Applied Mathematics, 236(6):1502–1510, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711004924.

Egami:2014:PTF

[EY14] Masahiko Egami and Kazutoshi Yamazaki. Phasetype fitting of scale functions for spectrally negative Lévy processes. Journal of Computational and Applied Mathematics, 264(??): 1–22, July 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713007176.

Edwards:2010:QPF

[EZ10] Michael G. Edwards and Hongwen Zheng. Quasipositive families of continuous Darcy-flux finite volume schemes on structured and unstructured grids. Journal of Computational and Applied Mathematics, 234(7):2152–2161, August 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709005512.

Ertem:2014:MPS

[EZ14] T. Ertem and A. Zafer.

Monotone positive solutions for a class of second-order nonlinear differential equations. Journal of Computational and Applied Mathematics, 259 (part B)(??): 672–681, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042713002069.

Fadimba:2010:LBE

[Fad10] Koffi B. Fadimba. A linear backward Euler scheme for the saturation equation: Regularity results and consistency. Journal of Computational and Applied Mathematics, 234(1):272– 282, May 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709008413.

Fata:2011:EET

[Fat11a] S. Nintcheu Fata. Explicit expressions for threedimensional boundary integrals in linear elasticity. Journal of Computational and Applied Mathematics, 235(15):4480–4495, June 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711002111.

Fata:2011:SAT

[Fat11b] S. Nintcheu Fata. Semianalytic treatment of the three-dimensional Poisson equation via a Galerkin BIE method. Journal of Computational and Applied Mathematics, 236(6):1216–1225, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711004444.

Fata:2013:BIA

[Fat13] S. Nintcheu Fata. Boundary integral approximation of volume potentials in three-dimensional linear elasticity. Journal of Computational and Applied Mathematics, 242(??):275– 284, April 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712004165.

Francisco:2012:NAM

[FB12] Juliano B. Francisco and Fermín S. Viloche Bazán. Nonmonotone algorithm for minimization on closed sets with applications to minimization on Stiefel manifolds. Journal of Computational and Applied Mathematics, 236(10):2717–2727, April 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712000258.

Faleichik:2014:GPI

[FBB14] Boris Faleichik, Ivan Bondar, and Vasily Byl. Generalized Picard iterations: a class of iterated Runge– Kutta methods for stiff problems. Journal of Computational and Applied Mathematics, 262(??):37– 50, May 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271300589X.

Ferreiro-Castilla:2012:MM

[FCS12] Albert Ferreiro-Castilla and Wim Schoutens. The β-Meixner model. Journal of Computational and Applied Mathematics, 236(9):2466– 2476, March 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271100611X.

Fauconnier:2013:SAN

[FD13] D. Fauconnier and E. Dick.

Spectral analysis of nonlinear finite difference discretizations. Journal of Computational and Applied Mathematics, 246(??):113– 121, July 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712004979.

Fauconnier:2010:DOF

[FDD10] D. Fauconnier, C. De Langhe, and E. Dick. A dynamically optimized finite difference scheme for Large-Eddy Simulation. Journal of Computational and Applied Mathematics, 234(7):2080–2088, August 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709005433.

Fan:2010:BPD

[FDH10] Mingshu Fan, Lili Du, and Qiaolin He. Blow-up properties for a degenerate parabolic system with nonlinear localized sources. Journal of Computational and Applied Mathematics, 235(1):91–101, November 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710002918.

Fan:2011:ESP

[FdLW^{+}11] Yunzheng Fan, Helder Geovane Gomes de Lima, Liping Wang, Jin Yun Yuan, and Xiao yue Zhou. Evaluation of ST preconditioners for saddle point problems. Journal of Computational and Applied Mathematics, 236(6):1543–1551, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004973.

Ferreira:2012:AND

[FdOdS12] José Augusto Ferreira, Paula de Oliveira, and Pascoal da Silva. Analytics and numerics of drug release tracking. Journal of Computational and Applied Mathematics, 236(15):3572– 3583, September 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711002172.

Ferreira:2012:RDV

[FdOS12] J. A. Ferreira, P. de Oliveira, and P. M. Silva. Reaction– diffusion in viscoelastic materials. Journal of Computational and Applied Mathematics, 236(15):3783–3795, September 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711006340.

Ferreira:2011:LCN

[Fer11] O. P. Ferreira. Local convergence of Newton’s method under majorant condition. Journal of Computational and Applied Mathematics, 235(5):1515–1522, January 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710004875.

Fernandez:2014:COC

[Fer14] Arturo J. Fernández. Computing optimal confidence sets for Pareto models under progressive censoring. Journal of Computational and Applied Mathematics, 258(??):168–180, March 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713004627.

Feng:2014:NIP

[FF14] Zengzhe Feng and Liang Fang. A new O(\sqrt{n}L)- iteration predictor–corrector algorithm with wide neighborhood for semidefinite programming. Journal of Computational and Applied Mathematics, 256(??):65– 76, January 15, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713003555.

Ferreira:2011:ANT

[FFM^{+}11] José M. Ferreira, Neville J.

Ford, Md. Abdul Malique, Sandra Pinelas, and Yubin Yan. Analytical and numerical treatment of oscillatory mixed differential equations with differentiable delays and advances. Journal of Computational and Applied Mathematics, 235(17):5112– 5130, July 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711002354.

Franco:2014:SEM

[FG14] J. M. Franco and I. Gómez.

Symplectic explicit methods of Runge–Kutta–Nyström type for solving perturbed oscillators. Journal of Computational and Applied Mathematics, 260(??):482– 493, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005591.

Farhat:2010:AAN

[FGEM10] Mohamed Farhat, Sebastien Guenneau, Stefan Enoch, and Alexander Movchan. All-angle-negative-refraction and ultra-refraction for liquid surface waves in 2D phononic crystals. Journal of Computational and Applied Mathematics, 234(6): 2011–2019, July 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042709005317.

Forys:2011:DMP

[FGO11] Wit Foryś, Juan Luis García Guirao, and Piotr Oprocha. A dynamical model of parallel computation on biinfinite time-scale. Journal of Computational and Applied Mathematics, 235 (7):1826–1832, February 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042709005822.

Ferreira:2012:LCA

[FGO12] O. P. Ferreira, M. L. N. Gonçalves, and P. R. Oliveira. Local convergence analysis of inexact Gauss–Newton like methods under majorant condition. Journal of Computational and Applied Mathematics, 236(9):2487–2498, March 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711006224.

Fortes:2013:APF

[FGPP13] M. A. Fortes, P. González, A. Palomares, and M. Pasadas. Approximation of patches by C^{r}-finite elements of Powell–Sabin type. Journal of Computational and Applied Mathematics, 252(??): 183–193, November 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042713001295.

Farago:2010:SIA

[FHMZ10] István Faragó, Ágnes Havasi, Svetozar Margenov, and Zahari Zlatev. Special issue on advanced computational algorithms: Introduction. Journal of Computational and Applied Mathematics, 235(2):345–347, November 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710003262.

Feistauer:2013:DDS

[FHPH^{+}13] Miloslav Feistauer, Jaroslava Hasnedlová-Prokopová, Jaromír Horácek, Adam Kosík, and Václav Kucera. DGFEM for dynamical systems describing interaction of compressible fluid and structures. Journal of Computational and Applied Mathematics, 254(??):17–30, December 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713001362.

Fakhar-Izadi:2011:SMP

[FID11] Farhad Fakhar-Izadi and Mehdi Dehghan. The spectral methods for parabolic Volterra integro-differential equations. Journal of Computational and Applied Mathematics, 235(14):4032– 4046, May 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S037704271100118X.

Fiedler:2013:MMF

[Fie13] Bernold Fiedler. Mathematical models of flageolet harmonics on stringed instruments. Journal of Computational and Applied Mathematics, 254(??):144–153, December 15, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042712005407.

Feng:2010:SAM

[FiMO10] Bao-Feng Feng, Ken ichi Maruno, and Yasuhiro Ohta. A self-adaptive moving mesh method for the Camassa–Holm equation. Journal of Computational and Applied Mathematics, 235(1):229–243, November 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710003341.

Fan:2010:RTG

[FJ10] ShengJun Fan and Long Jiang. A representation theorem for generators of BSDEs with continuous linear-growth generators in the space of processes. Journal of Computational and Applied Mathematics, 235(3):686–695, December 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710003742.

Fazio:2014:FDS

[FJ14] Riccardo Fazio and Alessandra Jannelli. Finite difference schemes on quasiuniform grids for BVPs on infinite intervals. Journal of Computational and Applied Mathematics, 269(??):14– 23, October 15, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042714001320.

Feng:2010:AAF

[FJLL10] Feng Feng, Young Bae Jun, Xiaoyan Liu, and Lifeng Li. An adjustable approach to fuzzy soft set based decision making. Journal of Computational and Applied Mathematics, 234(1):10–20, May 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709008024.

Ferronato:2014:GBF

[FJP14] Massimiliano Ferronato, Carlo Janna, and Giorgio Pini. A generalized block FSAI preconditioner for nonsymmetric linear systems. Journal of Computational and Applied Mathematics, 256(??):230–241, January 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713003944.

Fritzsche:2010:CES

[FKL10] Bernd Fritzsche, Bernd Kirstein, and Andreas Lasarow. On a class of extremal solutions of a moment problem for rational matrix-valued functions in the nondegenerate case II. Journal of Computational and Applied Mathematics, 235(4):1008–1041, December 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710004577.

Feng:2013:CFD

[FKL13] Xiaobing Feng, Chiu-Yen Kao, and Thomas Lewis. Convergent finite difference methods for onedimensional fully nonlinear second order partial differential equations. Journal of Computational and Applied Mathematics, 254(??):81– 98, December 15, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713000617.

Feng:2011:LSN

[FKNS11] Xinlong Feng, Imbunm Kim, Hyun Nam, and Dongwoo Sheen. Locally stabilized P_{1}-nonconforming quadrilateral and hexahedral finite element methods for the Stokes equations. Journal of Computational and Applied Mathematics, 236(5):714–727, October 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711003323.

Fahs:2010:HON

[FL10a] Hassan Fahs and Stéphane Lanteri. A high-order non-conforming discontinuous Galerkin method for time-domain electromagnetics. Journal of Computational and Applied Mathematics, 234(4):1088–1096, June 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709003070.

Floater:2010:P

[FL10b] Michael Floater and Tom Lyche. Preface. Journal of Computational and Applied Mathematics, 233(7):1619, February 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709000855.

Ford:2010:MMP

[FLEa10] Neville J. Ford, Patricia M. Lumb, and Enu Ekakaa. Mathematical modelling of plant species interactions in a harsh climate. Journal of Computational and Applied Mathematics, 234(9):2732–2744, September 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710000294.

Fan:2014:SIS

[FLHL14] Chia-Ming Fan, Hong-Huei Li, Chuan-Yen Hsu, and Chun-Hung Lin. Solving inverse Stokes problems by modified collocation Trefftz method. Journal of Computational and Applied Mathematics, 268(??):68–81, October 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714001241.

Francois-Lavet:2013:EBV

[FLHS^{+}13] V. François-Lavet, F. Henrotte, L. Stainier, L. Noels, and C. Geuzaine. An energy-based variational model of ferromagnetic hysteresis for finite element computations. Journal of Computational and Applied Mathematics, 246(??):243– 250, July 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712002427.

Fu:2014:AFE

[FLKW14] Zhisong Fu, T. James Lewis, Robert M. Kirby, and Ross T. Whitaker. Architecting the finite element method pipeline for the GPU. Journal of Computational and Applied Mathematics, 257(??):195–211, February 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713004470.

Ford:2010:NSF

[FLLT10] Neville J. Ford, Patricia M. Lumb, Pedro M. Lima, and M. Filomena Teodoro. The numerical solution of forward-backward differential equations: Decomposition and related issues. Journal of Computational and Applied Mathematics, 234(9):2745–2756, September 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710000439.

Floater:2011:AOF

[Flo11] Michael S. Floater. The approximation order of fourpoint interpolatory curve subdivision. Journal of Computational and Applied Mathematics, 236(4):476– 481, September 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042711001488.

Ling:2012:RPS

[fLxX12] Ai fan Ling and Cheng xian Xu. Robust portfolio selection involving options under a “marginal + joint” ellipsoidal uncertainty set. Journal of Computational and Applied Mathematics, 236(14):3373– 3393, August 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712001434.

Li:2010:NSP

[fLyHZ10] Jiao fen Li, Xi yan Hu, and Lei Zhang. New symmetry preserving method for optimal correction of damping and stiffness matrices using measured modes. Journal of Computational and Applied Mathematics, 234(5):1572– 1585, July 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710001366.

Fernandez:2011:NCC

[FM11a] J. R. Fernández and R. Martínez.

A normal compliance contact problem in viscoelasticity: an a posteriori error analysis and computational experiments. Journal of Computational and Applied Mathematics, 235(12):3599– 3614, April 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711000938.

Fernandez:2011:NAC

[FM11b] J. R. Fernández and R. Martínez. Numerical analysis of a contact problem including bone remodeling. Journal of Computational and Applied Mathematics, 235(7):1805– 1811, February 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042709007948.

Fontana:2014:UAP

[FM14] Claudio Fontana and Juan Miguel A. Montes. A unified approach to pricing and risk management of equity and credit risk. Journal of Computational and Applied Mathematics, 259 (part B)(??):350–361, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713002537.

Farina:2014:HIE

[FMP14] Leandro Farina, P. A. Martin, and Victor Péron. Hypersingular integral equations over a disc: Convergence of a spectral method and connection with Tranter’s method. Journal of Computational and Applied Mathematics, 269(??):118– 131, October 15, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704271400154X.

Frentz:2010:ASW

[FN10] Marie Frentz and Kaj Nyström. Adaptive stochastic weak approximation of degenerate parabolic equations of Kolmogorov type. Journal of Computational and Applied Mathematics, 234(1):146–164, May 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704270900819X.

Fan:2012:ECA

[FN12] Ying-Wai (Daniel) Fan and James G. Nagy. An efficient computational approach for multiframe blind deconvolution. Journal of Computational and Applied Mathematics, 236(8):2112–2125, February 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711005139.

Fili:2010:CTF

[FND10] Abdeljalil Fili, Ahmed Naji, and Yong Duan. Coupling three-field formulation and meshless mixed Galerkin methods using radial basis functions. Journal of Computational and Applied Mathematics, 234(8):2456– 2468, August 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042710001585.

Feichtinger:2010:CRD

[FO10] Hans G. Feichtinger and Darian M. Onchis. Constructive realization of dual systems for generators of multi-window spline-type spaces. Journal of Computational and Applied Mathematics, 234(12):3467–3479, October 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710002761.

Fornberg:2010:FDM

[For10]  Bengt Fornberg.  A fi-

nite difference method for free boundary problems. Journal of Computational and Applied Mathematics, 233(11):2831–2840, April 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709007699.

Fan:2013:EFN

[FP13a] Y. Fan and I. S. Pop. Equivalent formulations and numerical schemes for a class of pseudo-parabolic equations. Journal of Computational and Applied Mathematics, 246(??):86– 93, July 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003214.

Ferreira:2013:SSQ

[FP13b] J. A. Ferreira and L. Pinto.

Supraconvergence and supercloseness in quasilinear coupled problems. Journal of Computational and Applied Mathematics, 252(??): 120–131, November 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042712004281.

Fernandez:2012:GBP

[FPG12] Arturo J. Fernández and Carlos J. Pérez-González. Generalized beta prior models on fraction defective in reliability test planning. Journal of Computational and Applied Mathematics, 236(13):3147–3159, July 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712000696.

Filelis-Papadopoulos:2011:GPI

[FPGMG11] C. K. Filelis-Papadopoulos, G. A. Gravvanis, P. I.

Matskanidis, and K. M. Giannoutakis. On the GPGPU parallelization issues of finite element approximate inverse preconditioning. Journal of Computational and Applied Mathematics, 236(3):294–307, September 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004110.

Fernandez:2012:KCO

[FPP12] Lidia Fernández, Teresa E. Pérez, and Miguel A. Piñar. On Koornwinder classical orthogonal polynomials in two variables. Journal of Computational and Applied Mathematics, 236(15):3817– 3826, September 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711004638.

Foschi:2013:AAO

[FPP13] Paolo Foschi, Stefano Pagliarani, and Andrea Pascucci. Approximations for Asian options in local volatility models. Journal of Computational and Applied Mathematics, 237(1):442–459, January 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712002506.

Feischl:2014:CQO

[FPP14] M. Feischl, M. Page, and D. Praetorius. Convergence and quasi-optimality of adaptive FEM with inhomogeneous Dirichlet data. Journal of Computational and Applied Mathematics, 255(??):481–501, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003051.

Fernandez:2010:KTO

[FPPX10] Lidia Fernández, Teresa E.

Pérez, Miguel A. Piñar, and Yuan Xu. Krall-type orthogonal polynomials in several variables. Journal of Computational and Applied Mathematics, 233(6):1519– 1524, January 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709001563.

Feulvarch:2013:SRM

[FRB13] E. Feulvarch, J.-C. Roux, and J.-M. Bergheau. A simple and robust moving mesh technique for the finite element simulation of friction stir welding. Journal of Computational and Applied Mathematics, 246(??):269– 277, July 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712003032.

Ford:2013:NAT

[FRXY13] Neville J. Ford, M. Manuela Rodrigues, Jingyu Xiao, and Yubin Yan. Numerical analysis of a two-parameter fractional telegraph equation. Journal of Computational and Applied Mathematics, 249(??):95–106, September 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713000691.

Frontczak:2010:MMT

[FS10] Robert Frontczak and Rainer Schöbel. On modified Mellin transforms, Gauss–Laguerre quadrature, and the valuation of American call options. Journal of Computational and Applied Mathematics, 234(5):1559–1571, July 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710001354.

Festa:2012:CAM

[FS12a] Mattia Festa and Alvise Sommariva. Computing almost minimal formulas on the square. Journal of Computational and Applied Mathematics, 236(17):4296– 4302, November 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712002312.

Fort:2012:AGD

[FS12b] Marta Fort and J. Antoni Sellarès. Approximating generalized distance functions on weighted triangulated surfaces with applications. Journal of Computational and Applied Mathematics, 236(14):3461–3477, August 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712001562.

Fard:2014:SFF

[FS14] Omid S. Fard and Maryam Salehi. A survey on fuzzy fractional variational problems. Journal of Computational and Applied Mathematics, 271(??):71–82, December 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271400168X.

Frenzen:2010:NSN

[FSB10] C. L. Frenzen, Tsutomu Sasao, and Jon T. Butler.

On the number of segments needed in a piecewise linear approximation. Journal of Computational and Applied Mathematics, 234(2):437– 446, May 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709008528.

Flatley:2013:PTS

[FTTT13] Lisa Flatley, Alexey Tarasov, Martin Taylor, and Florian Theil. Packing twelve spherical caps to maximize tangencies. Journal of Computational and Applied Mathematics, 254(??):220–225, December 15, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713001441.

Fu:2010:CFE

[Fu10]  Hongfei Fu.  A char-

acteristic finite element method for optimal control problems governed by convection–diffusion equations. Journal of Computational and Applied Mathematics, 235(3):825–836, December 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710004073.

Fukushima:2011:PFC

[Fuk11] Toshio Fukushima. Precise and fast computation of a general incomplete elliptic integral of second kind by half and double argument transformations. Journal of Computational and Applied Mathematics, 235(14):4140– 4148, May 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711001270.

Fukushima:2012:PFC

[Fuk12] Toshio Fukushima. Precise and fast computation of a general incomplete elliptic integral of third kind by half and double argument transformations. Journal of Computational and Applied Mathematics, 236(7):1961– 1975, January 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711005772.

Fukushima:2013:FCG

[Fuk13a] Toshio Fukushima. Fast computation of a general complete elliptic integral of third kind by half and double argument transformations. Journal of Computational and Applied Mathematics, 253(??):142–157, December 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271300201X.

Fukushima:2013:NCI

[Fuk13b] Toshio Fukushima. Numerical computation of inverse complete elliptic integrals of first and second kinds. Journal of Computational and Applied Mathematics, 249(??):37– 50, September 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042713000630.

Fukushima:2013:NIG

[Fuk13c] Toshio Fukushima. Numerical inversion of a general incomplete elliptic integral. Journal of Computational and Applied Mathematics, 237(1):43–61, January 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712002920.

Fukushima:2013:PFC

[Fuk13d] Toshio Fukushima. Precise and fast computation of Lambert W -functions without transcendental function evaluations. Journal of Computational and Applied Mathematics, 244(??):77– 89, May 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712005213.

Frederix:2010:SLD

[FV10] Katrijn Frederix and Marc Van Barel. Solving a large dense linear system by adaptive cross approximation. Journal of Computational and Applied Mathematics, 234(11):3181–3195, October 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710000798.

Frederix:2013:SSC

[FV13] Katrijn Frederix and Marc Van Barel. Sparse spectral clustering method based on the incomplete Cholesky decomposition. Journal of Computational and Applied Mathematics, 237(1):145– 161, January 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712003093.

Ferranti:2014:CEN

[FV14] Micol Ferranti and Raf Vandebril. Computing eigenvalues of normal matrices via complex symmetric matrices. Journal of Computational and Applied Mathematics, 259 (part A)(??): 281–293, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042713004500.

Fan:2010:GFG

[FW10] Jianghua Fan and Xiaoguo Wang. Gap functions and global error bounds for setvalued variational inequalities. Journal of Computational and Applied Mathematics, 233(11):2956–2965, April 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709007882.

Fiebig-Wittmaack:2011:TPP

[FWBSBGA11] Melitta Fiebig-Wittmaack, Wolfgang Börsch-Supan, Ingeborg Bischoff-Gauss, and Orlando Astudillo. Tridiagonal preconditioning for Poisson-like difference equations with flat grids: Application to incompressible atmospheric flow. Journal of Computational and Applied Mathematics, 236(6):1435–1441, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004869.

Feng:2014:PJF

[FYA^{+}14] Tao Feng, Xijun Yu, Hengbin An, Qin Li, and Rongpei Zhang. The preconditioned Jacobian-free Newton–Krylov methods for nonequilibrium radiation diffusion equations. Journal of Computational and Applied Mathematics, 255(??):60–73, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713002008.

Ford:2010:NTO

[FYM10] Neville J. Ford, Yubin Yan, and Md. Abdul Malique. Numerical treatment of oscillatory functional differential equations. Journal of Computational and Applied Mathematics, 234(9):2757–2767, September 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710000397.

Feng:2011:KMQ

[FZ11] Renzhong Feng and Xun Zhou. A kind of multiquadric quasi-interpolation operator satisfying any degree polynomial reproduction property to scattered data. Journal of Computational and Applied Mathematics, 235(5):1502–1514, January 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710004863.

Gavriliadis:2012:TSM

[GA12] P. N. Gavriliadis and G. A. Athanassoulis. The truncated Stieltjes moment problem solved by using kernel density functions. Journal of Computational and Applied Mathematics, 236(17):4193–4213, November 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712002257.

Galperin:2013:OSU

[Gal13]  A. Galperin.  Optimal

secant-updates of rank 1. Journal of Computational and Applied Mathematics, 253(??):66–79, December 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713001775.

Gasmi:2014:TDC

[Gas14a] Abdelkader Gasmi. Twodimensional cavitating flow past an oblique plate in a channel. Journal of Computational and Applied Mathematics, 259 (part B)(??): 828–834, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042713003798.

Gittelson:2014:OAG

[GAS14b] Claude Jeffrey Gittelson, Roman Andreev, and Christoph Schwab. Optimality of adaptive Galerkin methods for random parabolic partial differential equations. Journal of Computational and Applied Mathematics, 263(??):189–201, June 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713007048.

Gatilov:2014:ULR

[Gat14] Stepan Yu. Gatilov. Using low-rank approximation of the Jacobian matrix in the Newton–Raphson method to solve certain singular equations. Journal of Computational and Applied Mathematics, 272(??):8–24, December 15, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042714002234.

Gautschi:2010:STN

[Gau10] Walter Gautschi. The spiral of Theodorus, numerical analysis, and special functions. Journal of Computational and Applied Mathematics, 235(4):1042–1052, December 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709008012.

Geng:2011:HPR

[GC11] Fazhan Geng and Minggen Cui. Homotopy perturbationreproducing kernel method for nonlinear systems of second order boundary value problems. Journal of Computational and Applied Mathematics, 235(8):2405– 2411, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042710006060.

Goldsztejn:2014:CAA

[GCC14] Alexandre Goldsztejn, Jorge Cruz, and Elsa Carvalho. Convergence analysis and adaptive strategy for the certified quadrature over a set defined by inequalities. Journal of Computational and Applied Mathematics, 260(??):543–560, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713004494.

Gao:2013:MNM

[GCH13] Qin Gao, Xiaoliang Cheng, and Zhengda Huang. Modified Numerov’s method for inverse Sturm–Liouville problems. Journal of Computational and Applied Mathematics, 253(??):181– 199, December 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713002112.

Gong:2014:FSI

[GCH14] Rongfang Gong, Xiaoliang Cheng, and Weimin Han. A fast solver for an inverse problem arising in bioluminescence tomography. Journal of Computational and Applied Mathematics, 267(??):228–243, September 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714001009.

Gao:2014:ECA

[GCNL14] Wei Gao, Ping Shing Chan, Hon Keung Tony Ng, and Xiaolei Lu. Efficient computational algorithm for optimal allocation in regression models. Journal of Computational and Applied Mathematics, 261(??):118– 126, May 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713005931.

Graca:2013:SAA

[GCY13] Ana Graça, Rui P. R. Cardoso, and Jeong Whan Yoon. Subspace analysis to alleviate the volumetric locking in the 3D solid-shell EFG method. Journal of Computational and Applied Mathematics, 246(??):185– 194, July 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712003779.

Grine:2010:MLM

[GD10] Slimane Grine and Pavel Diko. Multi-layer model of correlated energy prices. Journal of Computational and Applied Mathematics, 233(10):2590–2610, March 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042709007365.

Geiger:2013:NOC

[GD13] Christoph Geiger and Gerhard Dangelmayr. Nearonset convection patterns in the spherical Bénard problem with l = 2 critical.

Journal of Computational and Applied Mathematics, 254(??):154–174, December 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713001349.

Gupta:2014:SOD

[GD14] S. K. Gupta and D. Dangar.

On second-order duality for nondifferentiable minimax fractional programming. Journal of Computational and Applied Mathematics, 255(??):878–886, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003427.

Geng:2012:NAN

[Gen12] F. Z. Geng. A numerical algorithm for nonlinear multipoint boundary value problems. Journal of Computational and Applied Mathematics, 236(7):1789–1794, January 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711005437.

Gerolymos:2012:RLR

[Ger12] G. A. Gerolymos. Representation of the Lagrange reconstructing polynomial by combination of substencils. Journal of Computational and Applied Mathematics, 236(11):2763– 2794, May 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271200009X.

Gordon:2010:RSP

[GG10] Dan Gordon and Rachel Gordon. Row scaling as a preconditioner for some nonsymmetric linear systems with discontinuous coefficients. Journal of Computational and Applied Mathematics, 234(12):3480– 3495, October 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042710002979.

Gillis:2012:MAN

[GG12] Nicolas Gillis and François Glineur. A multilevel approach for nonnegative matrix factorization. Journal of Computational and Applied Mathematics, 236(7):1708–1723, January 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711005334.

Gordon:2013:RHS

[GG13a] Dan Gordon and Rachel Gordon. Robust and highly scalable parallel solution of the Helmholtz equation with large wave numbers. Journal of Computational and Applied Mathematics, 237(1):182–196, January 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003147.

Gudi:2013:FDI

[GG13b] Thirupathi Gudi and Hari Shanker Gupta. A fully discrete C^{0} interior penalty Galerkin approximation of the extended Fisher–Kolmogorov equation. Journal of Computational and Applied Mathematics, 247(??):1–16, August 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713000113.

Guo:2014:MSP

[GG14] Pei-Chang Guo and Xiao-Xia Guo. A modified structure-preserving doubling algorithm for nonsymmetric algebraic Riccati equations from transport theory. Journal of Computational and Applied Mathematics, 261(??):213– 220, May 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271300513X.

Galan-Garcia:2014:ATS

[GGAVRC14] José L. Galán-García, Gabriel Aguilera-Venegas, and Pedro Rodríguez-Cielos. An accelerated-time simulation for traffic flow in a smart city. Journal of Computational and Applied Mathematics, 270(??):557–563, November 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713006523.

Goovaerts:2011:P

[GGBA11] Marc J. Goovaerts, Omer L. Gebizlioğlu, Ismihan Bayramoğlu, and Ersan Akyildiz. Preface. Journal of Computational and Applied Mathematics, 235(16):4517–4518, June 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711001087.

Gondzio:2014:SLS

[GGH^{+}14] Jacek Gondzio, Jacek A. Gruca, J. A. Julian Hall, Wiesław Laskowski, and Marek Zukowski. Solving large-scale optimization problems related to Bell’s theorem. Journal of Computational and Applied Mathematics, 263(??):392– 404, June 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713006730.

Gaspar:2010:EGM

[GGLR10] Francisco Gaspar, J. L. Gracia, F. J. Lisbona, and C. Rodrigo. Efficient geometric multigrid implementation for triangular grids. Journal of Computational and Applied Mathematics, 234(4):1027–1035, June 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709002167.

Garzon:2012:DBP

[GGS12] Maria Garzon, L. J. Gray, and James Sethian. Droplet and bubble pinch-off computations using level sets. Journal of Computational and Applied Mathematics, 236(12):3034–3041, June 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042711001737.

Guerrero-Garcia:2010:SCR

[GGSP10] Pablo Guerrero-García and Ángel Santos-Palomo. A sparse counterpart of reichel and Gragg’s package QRUP. Journal of Computational and Applied Mathematics, 233(5):1232–1237, January 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709001769.

Ganesh:2010:TDE

[GH10a] M. Ganesh and S. C. Hawkins. Three dimensional electromagnetic scattering T -matrix computations. Journal of Computational and Applied Mathematics, 234(6):1702–1709, July 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709004932.

Guo:2010:DAK

[GH10b] Ling Guo and Jianguo Huang. Dynamical analysis of Kirchhoff plates by an explicit time integration Morley element method. Journal of Computational and Applied Mathematics, 234(8):2483–2492, August 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710001615.

Gander:2014:APA

[GH14a] Martin J. Gander and Ernst Hairer. Analysis for parareal algorithms applied to Hamiltonian differential equations. Journal of Computational and Applied Mathematics, 259 (part A)(??): 2–13, March 15, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713000307.

Giani:2014:GOA

[GH14b] S. Giani and P. Houston. Goal-oriented adaptive composite discontinuous Galerkin methods for incompressible flows. Journal of Computational and Applied Mathematics, 270(??): 32–42, November 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042714001472.

Garza:2010:OPM

[GHM10] L. Garza, J. Hernández, and F. Marcellán. Orthogonal polynomials and measures on the unit circle. The Geronimus transformations. Journal of Computational and Applied Mathematics, 233(5):1220–1231, January 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709001721.

Gutierrez:2010:DNF

[GHR10] J. M. Gutiérrez, M. A. Hernández, and N. Romero. Dynamics of a new family of iterative processes for quadratic polynomials. Journal of Computational and Applied Mathematics, 233(10):2688–2695, March 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709007481.

Gharbi:2010:NIF

[GI10] Nawel Gharbi and Malika Ioualalen. Numerical investigation of finitesource multiserver systems with different vacation policies. Journal of Computational and Applied Mathematics, 234(3):625–635, June 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709007870.

Giani:2012:PEE

[Gia12] Stefano Giani. An a posteriori error estimator for hp-adaptive discontinuous Galerkin methods for computing band gaps in photonic crystals. Journal of Computational and Applied Mathematics, 236(18):4810– 4826, December 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712000647.

Giani:2014:ADG

[Gia14]  Stefano Giani.  A hp-

adaptive discontinuous Galerkin method for plasmonic waveguides. Journal of Computational and Applied Mathematics, 270(??):12– 20, November 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042714001496.

Giannelli:2013:BDB

[GJ13] Carlotta Giannelli and Bert Jüttler. Bases and dimensions of bivariate hierarchical tensor-product splines. Journal of Computational and Applied Mathematics, 239(??):162–178, February 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712004049.

Guo:2011:LES

[GJCL11] Y. C. Guo, P. Jiang, C. K.

Chan, and W. Y. Lin. Large eddy simulation of coherent structures in rectangular methane non-premixed flame. Journal of Computational and Applied Mathematics, 235(13):3760–3767, May 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711000331.

Geum:2011:BFO

[GK11] Young Hee Geum and Young Ik Kim. A biparametric family of optimally convergent sixteenth-order multipoint methods with their fourth-step weighting function as a sum of a rational and a generic twovariable function. Journal of Computational and Applied Mathematics, 235(10):3178– 3188, March 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711000045.

Galvis:2014:SMF

[GK14] J. Galvis and S. Ki Kang. Spectral multiscale finite element for nonlinear flows in highly heterogeneous media: a reduced basis approach. Journal of Computational and Applied Mathematics, 260(??):494–508, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005542.

Geers:2010:MSC

[GKB10] M. G. D. Geers, V. G.

Kouznetsova, and W. A. M. Brekelmans. Multi-scale computational homogenization: Trends and challenges. Journal of Computational and Applied Mathematics, 234(7):2175–2182, August 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709005536.

Gehre:2012:SRE

[GKL^{+}12a] Matthias Gehre, Tobias Kluth, Antti Lipponen, Bangti Jin, Aku Seppänen, Jari P. Kaipio, and Peter Maass. Sparsity reconstruction in electrical impedance tomography: an experimental evaluation. Journal of Computational and Applied Mathematics, 236(8):2126– 2136, February 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711005140.

Guo:2012:NSN

[GKL12b] Chun-Hua Guo, Yueh-Cheng Kuo, and Wen-Wei Lin. Numerical solution of nonlinear matrix equations arising from Green’s function calculations in nano research. Journal of Computational and Applied Mathematics, 236(17):4166–4180, November 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712002130.

Gomilko:2012:RCP

[GKLZ12] Oleksandr Gomilko, Dmitry B. Karp, Minghua Lin, and Krystyna Zietak. Regions of convergence of a Padé family of iterations for the matrix sector function and the matrix p-th root. Journal of Computational and Applied Mathematics, 236(17):4410– 4420, November 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712001677.

Gurris:2010:FES

[GKT10] Marcel Gurris, Dmitri Kuzmin, and Stefan Turek. Finite element simulation of compressible particleladen gas flows. Journal of Computational and Applied Mathematics, 233(12):3121– 3129, April 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709004348.

Gurris:2011:IFE

[GKT11] Marcel Gurris, Dmitri Kuzmin, and Stefan Turek. Implicit finite element schemes for stationary compressible particle-laden gas flows. Journal of Computational and Applied Mathematics, 235(17):5056–5077, July 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042711002305.

Garanzha:2014:VMU

[GKU14] V. A. Garanzha, L. N.

Kudryavtseva, and S. V. Utyuzhnikov. Variational method for untangling and optimization of spatial meshes. Journal of Computational and Applied Mathematics, 269(??):24–41, October 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714001460.

Gao:2010:PCP

[GL10] Shan Gao and Zaiming Liu. The perturbed compound Poisson risk model with constant interest and a threshold dividend strategy. Journal of Computational and Applied Mathematics, 233(9):2181–2188, March 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709006736.

Gao:2012:GBA

[GLH12] Weifeng Gao, Sanyang Liu, and Lingling Huang. A global best artificial bee colony algorithm for global optimization. Journal of Computational and Applied Mathematics, 236(11):2741– 2753, May 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712000246.

Garcia-Ligero:2011:LSL

[GLHCLP11] M. J. García-Ligero, A. Hermoso-Carazo, and J. Linares-Pérez. Least-squares linear estimation of signals from observations with Markovian delays. Journal of Computational and Applied Mathematics, 236(2):234– 242, August 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042711003578.

Garcia-Ligero:2010:DLE

[GLHCLPN10] M. J. García-Ligero, A. Hermoso-Carazo, J. Linares-Pérez, and S. Nakamori. Derivation of linear estimation algorithms from measurements affected by multiplicative and additive noises. Journal of Computational and Applied Mathematics, 234(3):794–804, June 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710000567.

Guirao:2012:NEE

[GLLM12] Juan L. G. Guirao, Miguel A.

López, Jaume Llibre, and Raquel Martínez. A note on the equilibria of an economic model with local competition “à la Cournot”. Journal of Computational and Applied Mathematics, 236(12):3052– 3057, June 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006485.

Guarneri:2013:OND

[GLWF13] P. Guarneri, J. T. Leverenz, M. M. Wiecek, and G. Fadel. Optimization of nonhierarchically decomposed problems. Journal of Computational and Applied Mathematics, 246(??):312– 319, July 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712005365.

Grote:2010:ELT

[GM10] Marcus J. Grote and Teodora Mitkova. Explicit local time-stepping methods for Maxwell’s equations. Journal of Computational and Applied Mathematics, 234(12):3283–3302, October 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710002360.

Grote:2013:HOE

[GM13] Marcus J. Grote and Teodora Mitkova. Highorder explicit local timestepping methods for damped wave equations. Journal of Computational and Applied Mathematics, 239(??):270– 289, February 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042712004190.

Gillman:2014:FSP

[GM14a] A. Gillman and P. G. Martinsson. A fast solver for Poisson problems on infinite regular lattices. Journal of Computational and Applied Mathematics, 258(??):42– 56, March 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713004512.

Guneri:2014:SCF

[GM14b] Cem Güneri and Gary McGuire. Supersingular curves over finite fields and weight divisibility of codes. Journal of Computational and Applied Mathematics, 259 (part B)(??):474–484, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712005481.

Gorb:2014:FES

[GMRK14] Yuliya Gorb, Otto Mierka, Liudmila Rivkind, and Dmitri Kuzmin. Finite element simulation of three-dimensional particulate flows using mixture models. Journal of Computational and Applied Mathematics, 270(??):443–450, November 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713006936.

Guenneau:2010:ABG

[GMZ^{+}10] Sébastien Guenneau, Alexander Movchan, Frédéric Zolla, Natasha Movchan, and André Nicolet. Acoustic band gaps in arrays of neutral inclusions. Journal of Computational and Applied Mathematics, 234(6):1962– 1969, July 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042709005251.

Gong:2011:IPF

[GN11] Jing Gong and Jan Nordström. Interface procedures for finite difference approximations of the advection–diffusion equation. Journal of Computational and Applied Mathematics, 236(5):602–620, October 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711004481.

Gazzola:2014:APS

[GN14] Silvia Gazzola and Paolo Novati. Automatic parameter setting for Arnoldi– Tikhonov methods. Journal of Computational and Applied Mathematics, 256(??): 180–195, January 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042713003671.

Grigori:2011:TST

[GNN11] Laura Grigori, Frédéric Nataf, and Qiang Niu. Two sides tangential filtering decomposition. Journal of Computational and Applied Mathematics, 235(8):2647– 2661, February 15, 2011.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042710006333.

Goncu:2014:ESM

[GÖ14a] Ahmet Göncü and Giray Ökten. Efficient simulation of a multi-factor stochastic volatility model. Journal of Computational and Applied Mathematics, 259 (part B)(??):329–335, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271300109X.

Goncu:2014:UPS

[GÖ14b] Ahmet Göncü and Giray Ökten. Uniform point sets and the collision test. Journal of Computational and Applied Mathematics, 259 (part B)(??):798–804, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003634.

Gondal:2010:ERI

[Gon10] Muhammad Asif Gondal.

Exponential Rosenbrock integrators for option pricing. Journal of Computational and Applied Mathematics, 234(4):1153–1160, June 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709003586.

Gosse:2011:DSC

[Gos11] Laurent Gosse. A Donoho– Stark criterion for stable signal recovery in discrete wavelet subspaces. Journal of Computational and Applied Mathematics, 235(17): 5024–5039, July 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042711002287.

Gustafsson:2010:RBT

[GP10] Jonathan Gustafsson and Bartosz Protas. Regularization of the backward-in-time Kuramoto–Sivashinsky equation. Journal of Computational and Applied Mathematics, 234(2):398– 406, May 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709008498.

Garrappa:2011:API

[GP11] Roberto Garrappa and Marina Popolizio. On accurate product integration rules for linear fractional differential equations. Journal of Computational and Applied Mathematics, 235(5):1085– 1097, January 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710003948.

Gupta:2012:SCC

[GP12] D. K. Gupta and M. Prashanth.

Semilocal convergence of a continuation method with Hölder continuous second derivative in Banach spaces. Journal of Computational and Applied Mathematics, 236(13):3174– 3185, July 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712000726.

Gonzalez-Pinto:2010:EFS

[GPHAM10] S. González-Pinto, D. Hernández-Abreu, and J. I. Montijano. An efficient family of strongly A-stable Runge– Kutta collocation methods for stiff systems and DAEs. I. Stability and order results. Journal of Computational and Applied Mathematics, 234(4):1105–1116, June 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709003264.

Gonzalez-Pinto:2014:RTM

[GPHAPR14] S. González-Pinto, D. Hernández-Abreu, and S. Pérez-Rodríguez. Rosenbrocktype methods with inexact AMF for the time integration of advection–diffusionreaction PDEs. Journal of Computational and Applied Mathematics, 262(??):304– 321, May 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713006146.

Gonzalez-Pinto:2014:SSF

[GPHAS14] S. González-Pinto, D. Hernández-Abreu, and B. Simeon. Strongly A-stable first stage explicit collocation methods with stepsize control for stiff and differential-algebraic equations. Journal of Computational and Applied Mathematics, 259 (part A) (??):138–152, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704271300188X.

Gabriel:2010:DEW

[GPKV10] D. Gabriel, J. Plesek, R. Kolman, and F. Vales. Dispersion of elastic waves in the contact-impact problem of a long cylinder. Journal of Computational and Applied Mathematics, 234(6):1930–1936, July 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709005214.

Gok:2014:CIG

[GPO14] S. Z. Alparslan Gök, O. Palanci, and M. O. Olgun. Cooperative interval games: Mountain situations with interval data. Journal of Computational and Applied Mathematics, 259 (part B)(??): 622–632, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042713000605.

Gori:2013:CSP

[GPS13] L. Gori, F. Pitolli, and E. Santi. On a class of shape-preserving refinable functions with dilation 3. Journal of Computational and Applied Mathematics, 245(??):62– 74, June 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712005419.

Garcia-Pareja:2010:ACM

[GPVL10] S. García-Pareja, M. Vilches, and A. M. Lallena. Ant colony method to control variance reduction techniques in the Monte Carlo simulation of clinical electron linear accelerators of use in cancer therapy. Journal of Computational and Applied Mathematics, 233(6):1534–1541, January 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709001587.

Ghezzi:2010:OSP

[GPZ10] L. Ghezzi, L. F. Pavarino, and E. Zampieri. Overlapping Schwarz preconditioners for spectral element methods in nonstandard domains and heterogeneous media. Journal of Computational and Applied Mathematics, 234(5):1492–1504, July 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710001287.

Guo:2014:EFC

[GQ14a] Bai-Ni Guo and Feng Qi. Explicit formulae for computing Euler polynomials in terms of Stirling numbers of the second kind. Journal of Computational and Applied Mathematics, 272(??):251–257, December 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714002635.

Guo:2014:SIE

[GQ14b] Bai-Ni Guo and Feng Qi. Some identities and an explicit formula for Bernoulli and Stirling numbers. Journal of Computational and Applied Mathematics, 255(??):568–579, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271300318X.

Geng:2014:NMS

[GQL14] F. Z. Geng, S. P. Qian, and S. Li. A numerical method for singularly perturbed turning point problems with an interior layer. Journal of Computational and Applied Mathematics, 255(??):97– 105, January 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713002367.

Garcia-Rubio:2012:CME

[GR12] Raquel Garcia-Rubio. Computational methods in economic modeling and engineering. Journal of Computational and Applied Mathematics, 236(12):2921– 2922, June 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712000477.

Gout:2013:MAT

[GR13] Christian Gout and Lucia Romani. “Multivariate Approximation: Theory and Applications”. Journal of Computational and Applied Mathematics, 240(??):1–4, March 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712004098.

Dedicated to Leonardo Traversoni for his 60th birthday.

Gravvanis:2011:SIA

[Gra11] George A. Gravvanis. Special issue on “Aspects of Numerical Algorithms, Parallelization and Applications”. Journal of Computational and Applied Mathematics, 236(3):279–280, September 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711004493.

Green:2010:BAD

[Gre10] K. Green. Bifurcation analysis of delay-induced periodic oscillations. Journal of Computational and Applied Mathematics, 233(9):2405– 2412, March 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709007237.

Grohs:2013:BST

[Gro13]  P. Grohs.  Bandlimited

shearlet-type frames with nice duals. Journal of Computational and Applied Mathematics, 243(??):139– 151, May 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712004700.

Guirao:2011:NSE

[GRV11] Juan L. G. Guirao, Raquel G.

Rubio, and Juan A. Vera. Nonlinear stability of the equilibria in a double-bar rotating system. Journal of Computational and Applied Mathematics, 235(7):1819– 1825, February 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710002955.

Guo:2010:IFM

[GRZP10] W. C. Guo, G. Rauchs, W. H. Zhang, and J. P. Ponthot. Influence of friction in material characterization in microindentation measurement. Journal of Computational and Applied Mathematics, 234(7):2183–2192, August 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709005548.

Galewski:2010:VMN

[GS10] Marek Galewski and Joanna Smejda. On variational methods for nonlinear difference equations. Journal of Computational and Applied Mathematics, 233(11): 2985–2993, April 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042709007912.

Gundersen:2012:DSP

[GS12] Geir Gundersen and Trond Steihaug. On diagonally structured problems in unconstrained optimization using an inexact super Halley method. Journal of Computational and Applied Mathematics, 236(15):3685– 3695, September 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704271100392X.

Guillaume:2014:BMI

[GS14] Florence Guillaume and Wim Schoutens. A bootstrapping market implied moment matching calibration for models with time-dependent parameters. Journal of Computational and Applied Mathematics, 271(??):100–116, December 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714001861.

Grau-Sanchez:2011:FDD

[GSGN11a] Miquel Grau-Sánchez, Àngela Grau, and Miquel Noguera. Frozen divided difference scheme for solving systems of nonlinear equations. Journal of Computational and Applied Mathematics, 235(6):1739–1743, January 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710005340.

Grau-Sanchez:2011:CEI

[GSGN11b] Miquel Grau-Sánchez, Àngela Grau, and Miquel Noguera. On the computational efficiency index and some iterative methods for solving systems of nonlinear equations. Journal of Computational and Applied Mathematics, 236(6):1259–1266, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271100447X.

Gebizlioglu:2011:CCV

[GSK11] Omer L. Gebizlioğlu, Birdal Senoğlu, and Yeliz Mert Kantar. Comparison of certain value-at-risk estimation methods for the two-parameter Weibull loss distribution. Journal of Computational and Applied Mathematics, 235(11):3304– 3314, April 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042711000550.

Ghezavati:2011:ELT

[GSM11] V. R. Ghezavati and M. Saidi-Mehrabad. An efficient linearization technique for mixed 0–1 polynomial problem. Journal of Computational and Applied Mathematics, 235(6):1730–1738, January 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710004474.

Grau-Sanchez:2013:ADU

[GSNA13] Miquel Grau-Sánchez, Miquel Noguera, and Sergio Amat. On the approximation of derivatives using divided difference operators preserving the local convergence order of iterative methods. Journal of Computational and Applied Mathematics, 237(1):363–372, January 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712002403.

Grau-Sanchez:2014:LCF

[GSNDB14] Miquel Grau-Sánchez, Miquel Noguera, and José L. Diaz-Barrero. On the local convergence of a family of two-step iterative methods for solving nonlinear equations. Journal of Computational and Applied Mathematics, 255(??):753–764, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003415.

Gintides:2012:DPL

[GST12] Drossos Gintides, Mourad Sini, and Nguyen Trung Thành. Detection of pointlike scatterers using one type of scattered elastic waves. Journal of Computational and Applied Mathematics, 236(8):2137–2145, February 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711005152.

Gentile:2011:PIC

[GSV11] M. Gentile, A. Sommariva, and M. Vianello. Polynomial interpolation and cubature over polygons. Journal of Computational and Applied Mathematics, 235(17):5232–5239, July 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042711002603.

Gu:2009:HIM

[GT09] Chuanqing Gu and Zhaolu Tian. On the HSS iteration methods for positive definite Toeplitz linear systems. Journal of Computational and Applied Mathematics, 224(2):709– 718, February 15, 2009. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042708002975.

See erratum [LHL11].

Gumgum:2014:DSM

[GTS14] S. Gümgüm and M. Tezer-Sezgin. DRBEM solution of mixed convection flow of nanofluids in enclosures with moving walls. Journal of Computational and Applied Mathematics, 259 (part B)(??):730–740, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713002628.

Geetha:2010:EDI

[GU10] K. V. Geetha and R. Uthayakumar. Economic design of an inventory policy for noninstantaneous deteriorating items under permissible delay in payments. Journal of Computational and Applied Mathematics, 233(10):2492– 2505, March 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709007298.

Gavete:2013:NDA

[GUBS13] L. Gavete, F. Ureña, J. J.

Benito, and E. Salete. A note on the dynamic analysis using the generalized finite difference method. Journal of Computational and Applied Mathematics, 252(??):132–147, November 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712002774.

Gungor:2010:EUB

[Gün10] Ayse Dilek Güngör. Erratum to “An upper bound for the condition number of a matrix in spectral norm” [J. Comput. Appl. Math. 143 (2002) 141–144]. Journal of Computational and Applied Mathematics, 234(1): 316, May 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710000129.

Gokgoz:2014:SNC

[GUO14] I. H. Gökgöz, Ö. Ugur, and Y. Yolcu Okur. On the single name CDS price under structural modeling. Journal of Computational and Applied Mathematics, 259 (part B)(??):406–412, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271300397X.

Guseinov:2014:BVP

[Gus14] Gusein Sh. Guseinov. Boundary value problems for nonlinear impulsive Hamiltonian systems. Journal of Computational and Applied Mathematics, 259 (part B) (??):780–789, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042713003324.

Garino:2013:NSF

[GVF^{+}13] C. García Garino, M. S. Ribero Vairo, S. Andía Fagés, A. E. Mirasso, and J.-P. Ponthot. Numerical simulation of finite strain viscoplastic problems. Journal of Computational and Applied Mathematics, 246(??):174– 184, July 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271200427X.

Gavete:2012:ICC

[GVM^{+}12] Luis Gavete, Marta García Vivanco, Pedro Molina, M. Lucía Gavete, Francisco Ureña, and Juan José Benito. Implementation in CHIMERE of a conservative solver for the advection equation — CMMSE10. Journal of Computational and Applied Mathematics, 236(12):3026– 3033, June 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711001804.

Goovaerts:2010:P

[GVV10] M. J. Goovaerts, S. Vandewalle, and M. Van Daele. Preface. Journal of Computational and Applied Mathematics, 234(4):949, June 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710001251.

Goovaerts:2012:P

[GVV12] M. J. Goovaerts, M. Van Daele, and S. Vandewalle. Preface. Journal of Computational and Applied Mathematics, 236(15):3541, September 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712001926.

Gao:2010:CSF

[GW10] Guanghua Gao and Tongke Wang. Cubic superconvergent finite volume element method for one-dimensional elliptic and parabolic equations. Journal of Computational and Applied Mathematics, 233(9):2285–2301, March 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709006918.

Gao:2012:QIL

[GW12] Wenwu Gao and Zongmin Wu. Quasi-interpolation for linear functional data. Journal of Computational and Applied Mathematics, 236 (13):3256–3264, July 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042712000854.

Galvis:2014:ELM

[GW14a] Juan Galvis and Jia Wei.

Ensemble level multiscale finite element and preconditioner for channelized systems and applications. Journal of Computational and Applied Mathematics, 255(??):456–467, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003038.

Gao:2014:MWG

[GW14b] Fuzheng Gao and Xiaoshen Wang. A modified weak Galerkin finite element method for a class of parabolic problems. Journal of Computational and Applied Mathematics, 271(??):1–19, December 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714001770.

Gao:2014:GSD

[GW14c] Jianwei Gao and Liyuan Wu. On the Gerber–Shiu discounted penalty function in a risk model with two types of delayed-claims and random income. Journal of Computational and Applied Mathematics, 269(??):42– 52, October 15, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042714001514.

Gao:2014:QIS

[GW14d] Wenwu Gao and Zongmin Wu. A quasiinterpolation scheme for periodic data based on multiquadric trigonometric Bsplines. Journal of Computational and Applied Mathematics, 271(??):20–30, December 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042714001526.

Gwinner:2013:FCU

[Gwi13]  Joachim Gwinner.  hp–

FEM convergence for unilateral contact problems with Tresca friction in plane linear elastostatics. Journal of Computational and Applied Mathematics, 254(??):175–184, December 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713001210.

Govender:2014:DCP

[GWKE14] Nicolin Govender, Daniel N. Wilke, Schalk Kok, and Rosanne Els. Development of a convex polyhedral discrete element simulation framework for NVIDIA Kepler based GPUs. Journal of Computational and Applied Mathematics, 270(??):386– 400, November 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704271300705X.

Grigori:2010:SFB

[GWX10] Laura Grigori, Desire Nuentsa Wakam, and Hua Xiang. Saving flops in LU based shift-and-invert strategy. Journal of Computational and Applied Mathematics, 234(12):3216–3225, October 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710002001.

Guo:2011:NMS

[GWZ11] Bing Guo, Ren-Hong Wang, and Chun-Gang Zhu. A note on multi-step difference schemes. Journal of Computational and Applied Mathematics, 236(5):647– 652, October 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711003645.

Gan:2014:SAN

[GXW14] Siqing Gan, Aiguo Xiao, and Desheng Wang. Stability of analytical and numerical solutions of nonlinear stochastic delay differential equations. Journal of Computational and Applied Mathematics, 268(??):5–22, October 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714001289.

Gong:2011:REE

[GY11] Wei Gong and Ningning Yan. Robust error estimates for the finite element approximation of elliptic optimal control problems. Journal of Computational and Applied Mathematics, 236(6):1370–1381, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004717.

Galiano:2010:NTF

[GZ10] Silvia Jerez Galiano and Miguel Uh Zapata. A new TVD flux-limiter method for solving nonlinear hyperbolic equations. Journal of Computational and Applied Mathematics, 234(5):1395– 1403, July 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271000097X.

Georgiev:2011:ISM

[GZ11] Krassimir Georgiev and Zahari Zlatev. Implementation of sparse matrix algorithms in an advection– diffusion-chemistry module. Journal of Computational and Applied Mathematics, 236(3):342–353, September 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004213.

Gu:2013:CMF

[GZ13] Chuanqing Gu and Lin Zheng. Computation of matrix functions with deflated restarting. Journal of Computational and Applied Mathematics, 237(1):223– 233, January 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704271200310X.

Guo:2014:CPE

[GZ14a] Benqi Guo and Jianming Zhang. Construction of polynomial extensions in two dimensions and application to the hp finite element method. Journal of Computational and Applied Mathematics, 261(??):249– 270, May 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713005086.

See also reprint [GZ14b].

Guo:2014:RCP

[GZ14b] Benqi Guo and Jianming Zhang. Reprint of “Construction of polynomial extensions in two dimensions and application to the h– p finite element method” [MR3144710]. Journal of Computational and Applied Mathematics, 265(??):151– 172, August 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042714000272.

See [GZ14a].

Gao:2014:OBT

[GZL14] Li Gao, Quanxin Zhang, and Shouhua Liu. Oscillatory behavior of thirdorder nonlinear delay dynamic equations on time scales. Journal of Computational and Applied Mathematics, 256(??):104–113, January 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713003683.

Han:2011:SOF

[Han11] Bin Han. Symmetric orthogonal filters and wavelets with linear-phase moments. Journal of Computational and Applied Mathematics, 236(4):482–503, September 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711003311.

Harase:2011:ELR

[Har11] Shin Harase. An efficient lattice reduction method for F_{2}-linear pseudorandom number generators using Mulders and Storjohann algorithm. Journal of Computational and Applied Mathematics, 236(2):141–149, August 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711003281.

Hausenblas:2010:WAS

[Hau10] Erika Hausenblas. Weak approximation of the stochastic wave equation. Journal of Computational and Applied Mathematics, 235(1): 33–58, November 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042710001846.

Heubes:2014:CBC

[HBE14] Daniel Heubes, Andreas Bartel, and Matthias Ehrhardt. Characteristic boundary conditions in the lattice Boltzmann method for fluid and gas dynamics. Journal of Computational and Applied Mathematics, 262(??): 51–61, May 15, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042713004743.

Bao:2010:DPF

[hBW10] Zhen hua Bao and Jing Wang. On the discounted penalty function in the discrete time stationary renewal risk model. Journal of Computational and Applied Mathematics, 234(2):557– 562, May 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709008747.

Huang:2011:PGI

[HC11] Wei Huang and Di-Rong Chen. Projected gradient iteration for nonlinear operator equation. Journal of Computational and Applied Mathematics, 235(17):5017– 5023, July 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711002275.

Hsu:2014:IFZ

[HC14] Shun-Pin Hsu and Feng-Chi Cheng. On the infinite families of zero bounds of complex polynomials. Journal of Computational and Applied Mathematics, 256(??):219–229, January 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003890.

Howarth:2014:IIP

[HCM14] C. J. Howarth, P. N. Childs, and A. Moiola. Implementation of an interior point source in the ultra weak variational formulation through source extraction. Journal of Computational and Applied Mathematics, 271(??):295–306, December 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714002167.

Han:2014:GBC

[HCQ14] Li-Wen Han, Ying Chu, and Zhi-Yu Qiu. Generalized Bézier curves and surfaces based on Lupaš qanalogue of Bernstein operator. Journal of Computational and Applied Mathematics, 261(??):352–363, May 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713006468.

Hu:2013:GET

[HCWsO13] Qingjie Hu, Yu Chen, Jingtong Wang, and Zi sheng Ouyang. Generalization of an existence theorem for complementarity problems. Journal of Computational and Applied Mathematics, 253(??):158–162, December 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713002033.

Hashemi:2011:RCI

[HD11] Behnam Hashemi and Mehdi Dehghan. Results concerning interval linear systems with multiple righthand sides and the interval matrix equation AX = B. Journal of Computational and Applied Mathematics, 235(9):2969–2978, March 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710006655.

Hsieh:2013:PIM

[HD13] Tsu-Pang Hsieh and Chung-Yuan Dye. A productioninventory model incorporating the effect of preservation technology investment when demand is fluctuating with time. Journal of Computational and Applied Mathematics, 239(??):25– 36, February 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712003809.

Hainaut:2014:DPH

[HD14] Donatien Hainaut and Griselda Deelstra. Default probabilities of a holding company, with complete and partial information. Journal of Computational and Applied Mathematics, 271(??):380–400, December 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714001848.

Halama:2010:NMU

[HDF^{+}10] J. Halama, J. Dobes, J. Fort, J. Fürst, and K. Kozel. Numerical modeling of unsteady flow in steam turbine stage. Journal of Computational and Applied Mathematics, 234(7):2336–2341, August 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709005718.

Higuera:2013:NST

[HdJL^{+}13] P. Higuera, M. del Jesus, J. L. Lara, I. J. Losada, Y. Guanche, and G. Barajas. Numerical simulation of threedimensional breaking waves on a gravel slope using a two-phase flow Navier– Stokes model. Journal of Computational and Applied Mathematics, 246(??):144– 152, July 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712004268.

Hillewaere:2013:CAS

[HDL^{+}13] J. Hillewaere, J. Degroote, G. Lombaert, J. Vierendeels, and G. Degrande. Computational aspects of simulating wind induced ovalling vibrations in silo groups. Journal of Computational and Applied Mathematics, 246(??):161–173, July 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712002750.

Hladik:2011:AAR

[HDT11] Milan Hladík, David Daney, and Elias P. Tsigaridas. An algorithm for addressing the real interval eigenvalue problem. Journal of Computational and Applied Mathematics, 235(8):2715–2730, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006394.

He:2012:EPB

[He12]  Tian-Xiao He.  Eule-

rian polynomials and Bsplines. Journal of Computational and Applied Mathematics, 236(15):3763–3773, September 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711005462.

Huang:2012:DDS

[HGW12] Chengming Huang, Siqing Gan, and Desheng Wang. Delay-dependent stability analysis of numerical methods for stochastic delay differential equations. Journal of Computational and Applied Mathematics, 236(14):3514–3527, August 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712001203.

Hu:2013:ASB

[HGW13] Lin Hu, Siqing Gan, and Xiaojie Wang. Asymptotic stability of balanced methods for stochastic jumpdiffusion differential equations. Journal of Computational and Applied Mathematics, 238(??):126–143, January 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003482.

Herceg:2013:TOM

[HH13] Djordje Herceg and Dragoslav Herceg. Third-order modifications of Newton’s method based on Stolarsky and Gini means. Journal of Computational and Applied Mathematics, 245(??):53– 61, June 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712005390.

Herceg:2014:SOM

[HH14] Djordje Herceg and Dragoslav Herceg. Sixth-order modifications of Newton’s method based on Stolarsky and Gini means. Journal of Computational and Applied Mathematics, 267(??):244–253, September 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714001125.

Hao:2013:CSS

[HHH^{+}13] Wenrui Hao, Jonathan D.

Hauenstein, Bei Hu, Timothy McCoy, and Andrew J. Sommese. Computing steady-state solutions for a free boundary problem modeling tumor growth by Stokes equation. Journal of Computational and Applied Mathematics, 237(1):326– 334, January 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712002361.

Hao:2014:BAC

[HHHS14] Wenrui Hao, Jonathan D.

Hauenstein, Bei Hu, and Andrew J. Sommese. A bootstrapping approach for computing multiple solutions of differential equations. Journal of Computational and Applied Mathematics, 258(??):181–190, March 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271300455X.

Higueras:2014:OSS

[HHKK14] Inmaculada Higueras, Natalie Happenhofer, Othmar Koch, and Friedrich Kupka. Optimized strong stability preserving IMEX Runge– Kutta methods. Journal of Computational and Applied Mathematics, 272(??):116– 140, December 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042714002477.

He:2012:SPM

[HHL12] Hongjin He, Deren Han, and Zhibao Li. Some projection methods with the BB step sizes for variational inequalities. Journal of Computational and Applied Mathematics, 236(9):2590– 2604, March 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711006315.

Huang:2014:EST

[HHLW14] Tsung-Ming Huang, Han-En Hsieh, Wen-Wei Lin, and Weichung Wang. Eigenvalue solvers for three dimensional photonic crystals with face-centered cubic lattice. Journal of Computational and Applied Mathematics, 272(??):350–361, December 15, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042714001022.

Helluy:2012:WBA

[HHM12] Philippe Helluy, Jean-Marc Hérard, and Hélène Mathis.

A well-balanced approximate Riemann solver for compressible flows in variable cross-section ducts. Journal of Computational and Applied Mathematics, 236(7):1976–1992, January 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711005784.

Harlen:2010:PSL

[HHP^{+}10] O. G. Harlen, M. J. Holmes, V. J. Pinfield, M. J. W. Povey, and B. D. Sleeman. A perturbation solution for long wavelength thermoacoustic propagation in dispersions. Journal of Computational and Applied Mathematics, 234(6):1996– 2002, July 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709005299.

Hakula:2012:ASC

[HHT12] Harri Hakula, Nuutti Hyvönen, and Tomi Tuominen. On the hp-adaptive solution of complete electrode model forward problems of electrical impedance tomography. Journal of Computational and Applied Mathematics, 236(18):4645–4659, December 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712001653.

Han:2014:PPS

[HHX14] Deren Han, Hongjin He, and Lingling Xu. A proximal parallel splitting method for minimizing sum of convex functions with linear constraints. Journal of Computational and Applied Mathematics, 256(??):36– 51, January 15, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042713003543.

Hu:2011:GNM

[HHZ11] Sheng-Long Hu, Zheng-Hai Huang, and Qiong Zhang. A generalized Newton method for absolute value equations associated with second order cones. Journal of Computational and Applied Mathematics, 235(5):1490– 1501, January 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042710004851.

Hashemi:2011:IFM

[HJ11] Y. Hashemi and A. Jahangirian. Implicit fully mesh-less method for compressible viscous flow calculations. Journal of Computational and Applied Mathematics, 235(16):4687–4700, June 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710004279.

Hu:2014:NPE

[HJS14] Jun Hu, Longlong Jiang, and Zhongci Shi. New a posteriori error estimate and quasi-optimal convergence of the adaptive nonconforming Wilson element. Journal of Computational and Applied Mathematics, 265(??):173–186, August 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005335.

Hyon:2011:SNM

[HK11] YunKyong Hyon and Do Young Kwak. Superconvergence of new mixed finite element spaces. Journal of Computational and Applied Mathematics, 235(14):4265–4271, May 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271100152X.

Helfenstein:2012:PPC

[HK12] Rudi Helfenstein and Jonas Koko. Parallel preconditioned conjugate gradient algorithm on GPU. Journal of Computational and Applied Mathematics, 236(15): 3584–3590, September 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042711002196.

Heindl:2013:AWV

[HK13a] Michael Heindl and Angela Kunoth. An adaptive wavelet viscosity method for systems of hyperbolic conservation laws. Journal of Computational and Applied Mathematics, 240(??):215– 224, March 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712003202.

Huang:2013:OBE

[HK13b] Boqiang Huang and Angela Kunoth. An optimization based empirical mode decomposition scheme. Journal of Computational and Applied Mathematics, 240(??):174–183, March 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003020.

Hecht:2014:AAM

[HK14] Frédéric Hecht and Raphaël Kuate. An approximation of anisotropic metrics from higher order interpolation error for triangular mesh adaptation. Journal of Computational and Applied Mathematics, 258(??): 99–115, March 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713004482.

Hannukainen:2010:GLM

[HKK10] Antti Hannukainen, Sergey Korotov, and Michal Krízek. On global and local mesh refinements by a generalized conforming bisection algorithm. Journal of Computational and Applied Mathematics, 235(2):419–436, November 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710003365.

Haslinger:2011:QAE

[HKL11] Jaroslav Haslinger, Radek Kucera, and Tomás Ligurský. Qualitative analysis of 3D elastostatic contact problems with orthotropic Coulomb friction and solution-dependent coefficients of friction. Journal of Computational and Applied Mathematics, 235(12): 3464–3480, April 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704271100080X.

Hernandez:2011:LFS

[HKO11] Erwin Hernández, Dante Kalise, and Enrique Otárola. A locking-free scheme for the LQR control of a Timoshenko beam. Journal of Computational and Applied Mathematics, 235(5):1383– 1393, January 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042710004747.

He:2010:PFE

[HL10a] Yinnian He and Jian Li. A penalty finite element method based on the Euler implicit/explicit scheme for the time-dependent Navier– Stokes equations. Journal of Computational and Applied Mathematics, 235(3):708– 725, December 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042710003778.

Hu:2010:CIA

[HL10b] Liang-Gen Hu and Li-Wei Liu. A composite iterative algorithm for common fixed points for a finite family of nonexpansive mappings and applications. Journal of Computational and Applied Mathematics, 233(8):1987–1994, February 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709006578.

Huang:2010:CSI

[HL10c] Tingwen Huang and Chuandong Li. Chaotic synchronization by the intermittent feedback method. Journal of Computational and Applied Mathematics, 234(4): 1097–1104, June 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042709003197.

Huang:2011:NAP

[HL11] Yunqing Huang and Jichun Li. Numerical analysis of a PML model for timedependent Maxwell’s equations. Journal of Computational and Applied Mathematics, 235(13):3932–3942, May 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711000501.

Hennessy:2013:PFT

[HL13a] Matthew G. Hennessy and Gregory M. Lewis.

The primary flow transition in a differentially heated rotating channel of fluid with o(2) symmetry. Journal of Computational and Applied Mathematics, 254(??):116–131, December 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713001490.

Huang:2013:DBB

[HL13b] Yu-Mei Huang and De-Yong Lu. On decompositionbased block preconditioned iterative methods for halfquadratic image restoration. Journal of Computational and Applied Mathematics, 237(1):162–170, January 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003159.

Hu:2014:FPM

[HL14] Wenbin Hu and Shenghong Li. The forward-path method for pricing multiasset American-style options under general diffusion processes. Journal of Computational and Applied Mathematics, 263(??):25– 31, June 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713006584.

He:2011:EMS

[HLC11] Huimin He, Sanyang Liu, and Yeol Je Cho. An explicit method for systems of equilibrium problems and fixed points of infinite family of nonexpansive mappings. Journal of Computational and Applied Mathematics, 235(14):4128–4139, May 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711001269.

Howard:2014:SBU

[HLF14] Marylesa Howard, Aaron Luttman, and Michael Fowler. Sampling-based uncertainty quantification in deconvolution of X-ray radiographs. Journal of Computational and Applied Mathematics, 270(??):43– 51, November 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042714001101.

Hiptmair:2014:FCQ

[HLFP14] Ralf Hiptmair, María López-Fernández, and Alberto Paganini. Fast convolution quadrature based impedance boundary conditions. Journal of Computational and Applied Mathematics, 263(??):500–517, June 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713006985.

Huang:2013:NSS

[HLLW13] Tsung-Ming Huang, Tiexiang Li, Wen-Wei Lin, and Chin-Tien Wu. Numerical studies on structurepreserving algorithms for surface acoustic wave simulations. Journal of Computational and Applied Mathematics, 244(??):140–154, May 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712005201.

Haddar:2010:ESW

[HLM10] H. Haddar, J.-R. Li, and D. Matignon. Efficient solution of a wave equation with fractional-order dissipative terms. Journal of Computational and Applied Mathematics, 234(6):2003–2010, July 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709005305.

Hsiao:2011:CBF

[HLSX11] George C. Hsiao, Fengshan Liu, Jiguang Sun, and Liwei Xu. A coupled BEM and FEM for the interior transmission problem in acoustics. Journal of Computational and Applied Mathematics, 235(17):5213–5221, July 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711002585.

Hu:2010:ADD

[HLWM10] Zhixing Hu, Xiangdong Liu, Hui Wang, and Wanbiao Ma. Analysis of the dynamics of a delayed HIV pathogenesis model. Journal of Computational and Applied Mathematics, 234(2):461– 476, May 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709008553.

Hamilton:2010:FIK

[HM10] I. P. Hamilton and Ricardo A. Mosna. Fisher information and kinetic energy functionals: a dequantization approach. Journal of Computational and Applied Mathematics, 233(6):1542–1547, January 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709001599.

Hou:2011:KMS

[HM11] Qing-Hu Hou and Toufik Mansour. The kernel method and systems of functional equations with several conditions. Journal of Computational and Applied Mathematics, 235(5): 1205–1212, January 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042710004425.

Hassan:2012:AFD

[HMA12] H. Z. Hassan, A. A. Mohamad, and G. E. Atteia. An algorithm for the finite difference approximation of derivatives with arbitrary degree and order of accuracy. Journal of Computational and Applied Mathematics, 236(10):2622– 2631, April 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711006339.

Hejazi:2014:SCF

[HML14] H. Hejazi, T. Moroney, and F. Liu. Stability and convergence of a finite volume method for the space fractional advection-dispersion equation. Journal of Computational and Applied Mathematics, 255(??):684– 697, January 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713003373.

Haubold:2011:FSF

[HMS11] H. J. Haubold, A. M.

Mathai, and R. K. Saxena. Further solutions of fractional reaction–diffusion equations in terms of the H-function. Journal of Computational and Applied Mathematics, 235(5):1311– 1316, January 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710004541.

Hernandez-Martinez:2011:GFF

[HMVPAR11] Eliseo Hernandez-Martinez, Francisco J. Valdes-Parada, and Jose Alvarez-Ramirez. A Green’s function formulation of nonlocal finitedifference schemes for reaction– diffusion equations. Journal of Computational and Applied Mathematics, 235(9): 3096–3103, March 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042710005716.

Huabsomboon:2010:TSE

[HNK10] Pallop Huabsomboon, Boriboon Novaprateep, and Hideaki Kaneko. On Taylorseries expansion methods for the second kind integral equations. Journal of Computational and Applied Mathematics, 234(5):1466– 1472, July 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271000107X.

Hsiao:2011:EAD

[HNPX11] George C. Hsiao, Nilima Nigam, Joseph E. Pasciak, and Liwei Xu. Error analysis of the DtN–FEM for the scattering problem in acoustics via Fourier analysis. Journal of Computational and Applied Mathematics, 235(17):4949–4965, July 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711002147.

Hosseinkouchack:2011:FIC

[Hos11] Mehdi Hosseinkouchack. Further improvements in the calculation of censored quantile regressions. Journal of Computational and Applied Mathematics, 235(5):1429–1445, January 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710004784.

Heubeck:2010:CAN

[HP10] B. Heubeck and C. Pflaum.

Convergence analysis of non-conforming trigonometric finite wave elements. Journal of Computational and Applied Mathematics, 234(6):1920–1929, July 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709005202.

Harouna:2013:ECD

[HP13] S. Kadri Harouna and V. Perrier. Effective construction of divergence-free wavelets on the square. Journal of Computational and Applied Mathematics, 240(??):74–86, March 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042712003196.

Haelterman:2014:NSQ

[HPL^{+}14] R. Haelterman, J. Petit, B. Lauwens, H. Bruyninckx, and J. Vierendeels. On the non-singularity of the quasi-Newton-least squares method. Journal of Computational and Applied Mathematics, 257(??):129–131, February 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713004251.

Hamarik:2012:FRP

[HPR12] Uno Hämarik, Reimo Palm, and Toomas Raus. A family of rules for parameter choice in Tikhonov regularization of ill-posed problems with inexact noise level. Journal of Computational and Applied Mathematics, 236(8):2146–2157, February 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711005164.

Hakula:2013:CFM

[HQR13] Harri Hakula, Tri Quach, and Antti Rasila. Conjugate function method for numerical conformal mappings. Journal of Computational and Applied Mathematics, 237(1):340–353, January 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712002385.

Hochstenbach:2010:SRS

[HR10] Michiel E. Hochstenbach and Lothar Reichel. Subspacerestricted singular value decompositions for linear discrete ill-posed problems. Journal of Computational and Applied Mathematics, 235(4):1053–1064, December 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710003687.

Hochstenbach:2012:CAS

[HR12] Michiel E. Hochstenbach and Lothar Reichel. Combining approximate solutions for linear discrete illposed problems. Journal of Computational and Applied Mathematics, 236(8):2179– 2185, February 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704271100519X.

He:2010:DFO

[HS10a] Tieshan He and Yueliang Su. On discrete fourthorder boundary value problems with three parameters. Journal of Computational and Applied Mathematics, 233(10):2506–2520, March 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709007304.

Hieber:2010:EPB

[HS10b] Peter Hieber and Matthias Scherer. Efficiently pricing barrier options in a Markov-switching framework. Journal of Computational and Applied Mathematics, 235(3):679–685, December 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710003730.

Hryniv:2010:NSI

[HS10c] Rostyslav Hryniv and Paul Sacks. Numerical solution of the inverse spectral problem for Bessel operators. Journal of Computational and Applied Mathematics, 235(1):120–136, November 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710002943.

Hasegawa:2011:AAF

[HS11a] Takemitsu Hasegawa and Hiroshi Sugiura. Algorithms for approximating finite Hilbert transform with end-point singularities and its derivatives. Journal of Computational and Applied Mathematics, 236(2):243– 252, August 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711003633.

Hubalek:2011:EEG

[HS11b] Friedrich Hubalek and Carlo Sgarra. On the explicit evaluation of the geometric Asian options in stochastic volatility models with jumps. Journal of Computational and Applied Mathematics, 235(11):3355–3365, April 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711000732.

Huang:2012:SAF

[HS12] Weizhang Huang and Forrest Schaeffer. An L^{∞} stability analysis for the finite-difference solution of one-dimensional linear convection–diffusion equations on moving meshes. Journal of Computational and Applied Mathematics, 236(13):3338–3348, July 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042712001173.

Hansen:2012:ATM

[HSH12] Per Christian Hansen and Maria Saxild-Hansen. AIR Tools — a MATLAB package of algebraic iterative reconstruction methods. Journal of Computational and Applied Mathematics, 236(8):2167–2178, February 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711005188.

Hong:2013:SCR

[HSY13] Jialin Hong, Chol Sim, and Xiuling Yin. Solvability of concatenated Runge–Kutta equations for second-order nonlinear PDEs. Journal of Computational and Applied Mathematics, 245(??):232– 241, June 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712005456.

Hayfavi:2014:SMM

[HT14] Azize Hayfavi and Irem Talasli. Stochastic multifactor modeling of spot electricity prices. Journal of Computational and Applied Mathematics, 259 (part B)(??): 434–442, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042713005529.

Honnor:2010:NET

[HTH10] M. E. Honnor, J. Trevelyan, and D. Huybrechs. Numerical evaluation of the two-dimensional partition of unity boundary integrals for Helmholtz problems. Journal of Computational and Applied Mathematics, 234(6):1656–1662, July 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709004877.

Hao:2014:DSH

[HTLI14] Dinh Nho Hào, Phan Xuan Thanh, D. Lesnic, and M. Ivanchov. Determination of a source in the heat equation from integral observations. Journal of Computational and Applied Mathematics, 264(??): 82–98, July 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714000107.

Homayouni:2014:GAO

[HTM14] Seyed Mahdi Homayouni, Sai Hong Tang, and Omid Motlagh. A genetic algorithm for optimization of integrated scheduling of cranes, vehicles, and storage platforms at automated container terminals. Journal of Computational and Applied Mathematics, 270(??):545–556, November 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713006535.

Hu:2012:ACS

[Hu12] Qian-Qian Hu. Approximating conic sections by constrained Bézier curves of arbitrary degree. Journal of Computational and Applied Mathematics, 236(11):2813– 2821, May 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712000283.

Hu:2014:CSC

[Hu14] Guang-Da Hu. Computable stability criterion of linear neutral systems with unstable difference operators. Journal of Computational and Applied Mathematics, 271(??):223–232, December 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042714001915.

Huang:2012:EMS

[Hua12] Chengming Huang. Exponential mean square stability of numerical methods for systems of stochastic differential equations. Journal of Computational and Applied Mathematics, 236(16):4016– 4026, October 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712001227.

Huang:2014:MSS

[Hua14a] Chengming Huang. Mean square stability and dissipativity of two classes of theta methods for systems of stochastic delay differential equations. Journal of Computational and Applied Mathematics, 259 (part A)(??):77–86, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713001465.

Huang:2014:CEA

[Hua14b] Rong Huang. Componentwise error analysis for linear systems associated with sign regular matrices. Journal of Computational and Applied Mathematics, 255(??):133– 141, January 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713002513.

Huang:2014:PFC

[Hua14c] Yu-Mei Huang. A practical formula for computing optimal parameters in the HSS iteration methods. Journal of Computational and Applied Mathematics, 255(??):142–149, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713002483.

Huillet:2010:RWL

[Hui10] Thierry Huillet. Random walk with long-range interaction with a barrier and its dual: Exact results. Journal of Computational and Applied Mathematics, 233(10):2449–2467, March 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709007262.

Haine:2011:CRV

[HV11] Luc Haine and Didier Vanderstichelen. A centerless representation of the Virasoro algebra associated with the unitary circular ensemble. Journal of Computational and Applied Mathematics, 236(1):19–27, August 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271000350X.

Hollevoet:2012:EFM

[HV12] D. Hollevoet and M. Van Daele. Exponentially-fitted methods and their stability functions. Journal of Computational and Applied Mathematics, 236(16):4006– 4015, October 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712000738.

Hamedani:2013:SRA

[HV13] G. G. Hamedani and Hans W. Volkmer. Some remarks on Arslan’s 2011 paper. Journal of Computational and Applied Mathematics, 237(1):1–4, January 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712002749.

Humet:2014:AGT

[HV14] Matthias Humet and Marc Van Barel. Algorithms for the Geronimus transformation for orthogonal polynomials on the unit circle. Journal of Computational and Applied Mathematics, 267(??):195–217, September 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714001034.

Haelterman:2010:ORK

[HVV10] R. Haelterman, J. Vierendeels, and D. Van Heule. Optimization of the Runge– Kutta iteration with residual smoothing. Journal of Computational and Applied Mathematics, 234(1):253– 271, May 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709008401.

Haelterman:2011:MSS

[HVV^{+}11a] Rob Haelterman, Jan Vierendeels, Dirk Van Heule, Steve De Ridder, and Helena Bruyninckx. Multistage solvers optimized for damping and propagation. Journal of Computational and Applied Mathematics, 235(13):3775–3782, May 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711000355.

Hollevoet:2011:EFM

[HVV11b] D. Hollevoet, M. Van Daele, and G. Vanden Berghe. Exponentially fitted methods applied to fourthorder boundary value problems. Journal of Computational and Applied Mathematics, 235(18):5380–5393, July 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711003219.

Hu:2010:OAP

[HW10] Fengxia Hu and Rongming Wang. Optimal allocation of policy limits and deductibles in a model with mixture risks and discount factors. Journal of Computational and Applied Mathematics, 234(10):2953–2961, September 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710002049.

Hanke:2012:EBI

[HW12] M. Hanke and L. Warth.

Electrostatic backscattering by insulating obstacles. Journal of Computational and Applied Mathematics, 236(8):2158–2166, February 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711005176.

Hagstrom:2010:RBC

[HWG10] Thomas Hagstrom, Timothy Warburton, and Dan Givoli. Radiation boundary conditions for timedependent waves based on complete plane wave expansions. Journal of Computational and Applied Mathematics, 234(6):1988–1995, July 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709005287.

Hur:2011:BCQ

[HwK11] Seok Hur and Tae wan Kim. The best G^{1} cubic and G^{2} quartic Bézier approximations of circular arcs. Journal of Computational and Applied Mathematics, 236(6):1183–1192, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711004419.

Hur:2014:CSI

[HwKB14] Seok Hur, Tae wan Kim, and Cesare Bracco. Comprehensive study of intersection curves in R^{4} based on the system of ODEs. Journal of Computational and Applied Mathematics, 256(??):121–130, January 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713003622.

Hu:2013:CPA

[HX13] Qianqian Hu and Huixia Xu. Constrained polynomial approximation of rational Bézier curves using reparameterization. Journal of Computational and Applied Mathematics, 249(??): 133–143, September 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042713000927.

Han:2014:PSC

[HX14a] Lixing Han and Jianhong Xu. Proof of Stenger’s conjecture on matrix I^{(−1)} of Sinc methods. Journal of Computational and Applied Mathematics, 255(??):805– 811, January 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713003452.

Hufford:2014:SLD

[HX14b] Casey Hufford and Yulong Xing. Superconvergence of the local discontinuous Galerkin method for the linearized Korteweg–de Vries equation. Journal of Computational and Applied Mathematics, 255(??):441– 455, January 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713003002.

Hu:2014:FMC

[HXSF14] ZhongBo Hu, ShengWu Xiong, QingHua Su, and ZhiXiang Fang. Finite Markov chain analysis of classical differential evolution algorithm. Journal of Computational and Applied Mathematics, 268(??):121– 134, October 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042714001290.

Huang:2011:MSA

[HXW11] Aiqun Huang, Chengxian Xu, and Meihua Wang. A modified SLP algorithm and its global convergence. Journal of Computational and Applied Mathematics, 235(14):4302–4307, May 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711001713.

Xie:2011:EDP

[hXZ11] Jie hua Xie and Wei Zou. On the expected discounted penalty function for the compound Poisson risk model with delayed claims. Journal of Computational and Applied Mathematics, 235(8):2392–2404, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006059.

Huang:2011:IEA

[HY11] Guang-Xin Huang and Feng Yin. An inverse eigenproblem and an associated approximation problem for generalized reflexive and anti-reflexive matrices. Journal of Computational and Applied Mathematics, 235(8):2888–2895, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006667.

Huang:2014:SAE

[HYY14] Yunqing Huang, Wei Yang, and Nianyu Yi. Superconvergence analysis for the explicit polynomial recovery method. Journal of Computational and Applied Mathematics, 265(??):187–198, August 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005013.

Han:2011:PHS

[HZ11] Xizhi Han and Hao Zhang.

Periodic and homoclinic solutions generated by impulses for asymptotically linear and sublinear Hamiltonian system. Journal of Computational and Applied Mathematics, 235(5):1531– 1541, January 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042710004899.

Hicken:2013:SPO

[HZ13] J. E. Hicken and D. W. Zingg. Summation-by-parts operators and high-order quadrature. Journal of Computational and Applied Mathematics, 237(1):111– 125, January 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042712003056.

Hashemi:2014:MCB

[HZ14] M. Y. Hashemi and K. Zamzamian. A multidimensional characteristic-based method for making incompressible flow calculations on unstructured grids. Journal of Computational and Applied Mathematics, 259 (part B)(??):752–759, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271300304X.

Huang:2014:OPU

[HZR14] Jiexiang Huang, Wenli Zhu, and Xinfeng Ruan. Option pricing using the fast Fourier transform under the double exponential jump model with stochastic volatility and stochastic intensity. Journal of Computational and Applied Mathematics, 263(??):152– 159, June 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713006791.

Ibryaeva:2013:ACP

[IA13] O. L. Ibryaeva and V. M.

Adukov. An algorithm for computing a Padé approximant with minimal degree denominator. Journal of Computational and Applied Mathematics, 237(1):529– 541, January 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712002579.

Isola:2012:SFS

[IG12] D. Isola and A. Guardone.

Simulation of flows with strong shocks with an adaptive conservative scheme. Journal of Computational and Applied Mathematics, 236(18):4660–4670, December 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712001641.

Ibanez:2011:PLA

[IHRA11] J. Javier Ibáñez, Vicente Hernández, Pedro A. Ruiz, and Enrique Arias. A piecewise-linearized algorithm based on the Krylov subspace for solving stiff ODEs. Journal of Computational and Applied Mathematics, 235(7):1798–1804, February 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710004097.

Iiduka:2012:FPO

[Iid12]  Hideaki Iiduka.  Fixed

point optimization algorithm and its application to network bandwidth allocation. Journal of Computational and Applied Mathematics, 236(7):1733–1742, January 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711005358.

Izzo:2014:CAS

[IJ14] Giuseppe Izzo and Zdzisław Jackiewicz. Construction of algebraically stable DIMSIMs. Journal of Computational and Applied Mathematics, 261(??):72– 84, May 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005906.

Izzo:2010:GLM

[IJMV10] G. Izzo, Z. Jackiewicz, E. Messina, and A. Vecchio. General linear methods for Volterra integral equations. Journal of Computational and Applied Mathematics, 234(9):2768–2782, September 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710000270.

Isaacson:2011:NSL

[IK11] Samuel A. Isaacson and Robert M. Kirby. Numerical solution of linear Volterra integral equations of the second kind with sharp gradients. Journal of Computational and Applied Mathematics, 235(14):4283–4301, May 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711001701.

Imai:2013:NIL

[IK13] Junichi Imai and Reiichiro Kawai. Numerical inverse Lévy measure method for infinite shot noise series representation. Journal of Computational and Applied Mathematics, 253(??):264– 283, December 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713001799.

Ihrle:2013:NMM

[ILEE13] S. Ihrle, M. Lauxmann, A. Eiber, and P. Eberhard. Nonlinear modelling of the middle ear as an elastic multibody system — applying model order reduction to acousto-structural coupled systems. Journal of Computational and Applied Mathematics, 246(??): 18–26, July 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003007.

Ibanez:2010:CSS

[ILMS10] M. J. Ibáñez, A. Lamnii, H. Mraoui, and D. Sbibih. Construction of spherical spline quasi-interpolants based on blossoming. Journal of Computational and Applied Mathematics, 234 (1):131–145, May 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042709008188.

Ideon:2014:LLR

[IO14a] Erge Ideon and Peeter Oja.

Linear/linear rational spline collocation for linear boundary value problems. Journal of Computational and Applied Mathematics, 263(??): 32–44, June 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713006602.

Inkaya:2014:AVF

[IO14b] A. Inkaya and Y. Yolcu Okur. Analysis of volatility feedback and leverage effects on the ISE30 index using high frequency data. Journal of Computational and Applied Mathematics, 259 (part B)(??):377–384, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003221.

Indratno:2011:CMS

[IR11] Sapto W. Indratno and Alexander G. Ramm. A collocation method for solving some integral equations in distributions. Journal of Computational and Applied Mathematics, 236(6):1296– 1313, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042711004596.

Islam:2014:EMN

[Isl14] M. Qamarul Islam. Estimation in multivariate nonnormal distributions with stochastic variance function. Journal of Computational and Applied Mathematics, 255(??):698–714, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003300.

Ismail:2014:ISS

[Ism14] Ali A. Ismail. Inference for a step-stress partially accelerated life test model with an adaptive Type-II progressively hybrid censored data from Weibull distribution. Journal of Computational and Applied Mathematics, 260(??):533– 542, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271300558X.

Ikegami:2010:FDG

[ISN10] Tsutomu Ikegami, Tetsuya Sakurai, and Umpei Nagashima. A filter diagonalization for generalized eigenvalue problems based on the Sakurai– Sugiura projection method.

Journal of Computational and Applied Mathematics, 233(8):1927–1936, February 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709006529.

Jansen:2014:SEF

[Jan14] Lennart Jansen. A semiexplicit formulation of a coupled electromagnetic field/circuit problem. Journal of Computational and Applied Mathematics, 262 (??):150–160, May 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042713005967.

Jiang:2013:WRM

[JD13] Yao-Lin Jiang and Xiao-Li Ding. Waveform relaxation methods for fractional differential equations with the Caputo derivatives. Journal of Computational and Applied Mathematics, 238(??):51–67, January 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003408.

Jeon:2010:ACB

[Jeo10] Youngmok Jeon. Analysis of the cell boundary element methods for convection dominated convection– diffusion equations. Journal of Computational and Applied Mathematics, 234 (8):2469–2482, August 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042710001627.

Jiang:2013:AED

[JGH^{+}13] Hao Jiang, Stef Graillat, Canbin Hu, Shengguo Li, Xiangke Liao, Lizhi Cheng, and Fang Su. Accurate evaluation of the k-th derivative of a polynomial and its application. Journal of Computational and Applied Mathematics, 243(??):28– 47, May 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712005018.

Jamshidi:2014:TVR

[JGK14] A. A. Jamshidi, C. W. Gear, and I. G. Kevrekidis. Time varying radial basis functions. Journal of Computational and Applied Mathematics, 266(??):61–72, August 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714000399.

Jiang:2010:BFF

[Jia10] Qingtang Jiang. Bi-frames with 4-fold axial symmetry for quadrilateral surface multiresolution processing. Journal of Computational and Applied Mathematics, 234(12):3303–3325, October 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710002372.

Jiang:2011:WBF

[Jia11] Qingtang Jiang. Wavelet biframes with uniform symmetry for curve multiresolution processing. Journal of Computational and Applied Mathematics, 235(6):1653– 1675, January 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710004991.

Jia:2012:CIR

[Jia12] Zhongxiao Jia. On convergence of the inexact Rayleigh quotient iteration with MINRES. Journal of Computational and Applied Mathematics, 236(17):4276– 4295, November 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712002269.

Johnston:2013:NMN

[JJE13] Barbara M. Johnston, Peter R. Johnston, and David Elliott. A new method for the numerical evaluation of nearly singular integrals on triangular elements in the 3D boundary element method. Journal of Computational and Applied Mathematics, 245(??):148– 161, June 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712005493.

Jaklic:2013:HIR

[JJK^{+}13] Gašper Jaklič, Bert Jüttler, Marjeta Krajnc, Vito Vitrih, and Emil Žagar. Hermite interpolation by rational G^{k} motions of low degree. Journal of Computational and Applied Mathematics, 240(??):20– 30, March 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003433.

Ji:2011:QAS

[JJL11] Chunyan Ji, Daqing Jiang, and Xiaoyue Li. Qualitative analysis of a stochastic ratio-dependent predator– prey system. Journal of Computational and Applied Mathematics, 235(5):1326– 1341, January 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710004590.

Jaberipour:2010:SSR

[JK10a] Majid Jaberipour and Esmaile Khorram. Solving the sum-of-ratios problems by a harmony search algorithm. Journal of Computational and Applied Mathematics, 234(3):733–742, June 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710000178.

Jaiboon:2010:SCT

[JK10b] Chaichana Jaiboon and Poom Kumam. Strong convergence theorems for solving equilibrium problems and fixed point problems of ξ-strict pseudocontraction mappings by two hybrid projection methods. Journal of Computational and Applied Mathematics, 234(3):722–732, June 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710000166.

Jeong:2010:CNS

[JK10c] Darae Jeong and Junseok Kim. A Crank–Nicolson scheme for the Landau– Lifshitz equation without damping. Journal of Computational and Applied Mathematics, 234(2):613– 623, May 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710000063.

Jeong:2013:CSA

[JK13] Darae Jeong and Junseok Kim. A comparison study of ADI and operator splitting methods on option pricing models. Journal of Computational and Applied Mathematics, 247(??):162–171, August 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713000277.

Jaklic:2010:LSP

[JKK^{+}10] Gasper Jaklic, Jernej Kozak, Marjeta Krajnc, Vito Vitrih, and Emil Zagar. Lattices on simplicial partitions. Journal of Computational and Applied Mathematics, 233(7):1704–1715, February 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709000946.

John:2010:FEV

[JKS10] Volker John, Adela Kindl, and Carina Suciu. Finite element LES and VMS methods on tetrahedral meshes. Journal of Computational and Applied Mathematics, 233(12):3095–3102, April 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709003495.

Jager:2013:ACH

[JKS13] Gabriela Jager, Angela Kunoth, and Wolf-Dieter Schuh. Approximate continuation of harmonic functions in geodesy: a spline based least squares approach with regularization. Journal of Computational and Applied Mathematics, 237(1):62–82, January 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271200297X.

Jovanovic:2011:CFT

[JKV11] Bosko S. Jovanovic, Miglena N.

Koleva, and Lubin G. Vulkov. Convergence of a FEM and two-grid algorithms for elliptic problems on disjoint domains. Journal of Computational and Applied Mathematics, 236(3):364–374, September 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711004146.

Jiang:2011:EMF

[JL11] Ziwen Jiang and Aiqin Li. Expanded mixed finite element methods for the problem of purely longitudinal motion of a homogeneous bar. Journal of Computational and Applied Mathematics, 235(8):2157–2169, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710005686.

Jie:2014:UMF

[JLZ14] Li Rui Jie, Qiu Ling, and Xuan Zuxing. Uniqueness of meromorphic functions sharing one small function. Journal of Computational and Applied Mathematics, 263(??):225–235, June 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271300681X.

Ji:2010:NNT

[JLZZ10] Ying Ji, Yijun Li, Kecun Zhang, and Xinli Zhang. A new nonmonotone trust-region method of conic model for solving unconstrained optimization. Journal of Computational and Applied Mathematics, 233(8):1746–1754, February 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709006335.

Jiang:2011:HOF

[JM11] Yingjun Jiang and Jingtang Ma. High-order finite element methods for timefractional partial differential equations. Journal of Computational and Applied Mathematics, 235(11):3285– 3290, April 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711000227.

Jiang:2012:ALV

[JM12] Yanan Jiang and Michael D.

Marcozzi. Asset liquidity and the valuation of derivative securities. Journal of Computational and Applied Mathematics, 236(17):4525– 4536, November 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712002063.

Jiang:2013:SCM

[JM13] Yingjun Jiang and Jingtang Ma. Spectral collocation methods for Volterraintegro differential equations with noncompact kernels. Journal of Computational and Applied Mathematics, 244(??):115–124, May 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712004803.

Janela:2010:NNF

[JMS10a] João Janela, Alexandra Moura, and Adélia Sequeira. A 3D non-Newtonian fluid-structure interaction model for blood flow in arteries. Journal of Computational and Applied Mathematics, 234 (9):2783–2791, September 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710000361.

Junghanns:2010:IEF

[JMS10b] P. Junghanns, G. Monegato, and A. Strozzi. On the integral equation formulations of some 2D contact problems. Journal of Computational and Applied Mathematics, 234 (9):2808–2825, September 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710000257.

Jung:2013:VMR

[JMV13] Miyoun Jung, Antonio Marquina, and Luminita A. Vese. Variational multiframe restoration of images degraded by noisy (stochastic) blur kernels. Journal of Computational and Applied Mathematics, 240(??):123– 134, March 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712002993.

Jiang:2011:HEA

[JPX11] Ying Jiang, C.-K. Peng, and Yuesheng Xu. Hierarchical entropy analysis for biological signals. Journal of Computational and Applied Mathematics, 236(5):728– 742, October 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704271100330X.

Joly:2010:OHO

[JR10a] P. Joly and J. Rodríguez.

Optimized higher order time discretization of second order hyperbolic problems: Construction and numerical study. Journal of Computational and Applied Mathematics, 234(6):1953–1961, July 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704270900524X.

Juhasz:2010:CRT

[JR10b] Imre Juhász and Ágoston Róth. Closed rational trigonometric curves and surfaces. Journal of Computational and Applied Mathematics, 234(8):2390–2404, August 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710001500.

Juhasz:2014:SIT

[JR14] Imre Juhász and Ágoston Róth. A scheme for interpolation with trigonometric spline curves. Journal of Computational and Applied Mathematics, 263(??):246– 261, June 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713007073.

Janikova:2010:FDL

[JS10] E. Janíková and M. Slodicka. Fully discrete linear approximation scheme for electric field diffusion in type-II superconductors. Journal of Computational and Applied Mathematics, 234(7):2054–2061, August 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709005408.

Junghanns:2014:NSC

[JTV14] P. Junghanns, W. Themistoclakis, and A. Vecchio. On the numerical solution of a class of nonstandard Sturm–Liouville boundary value problems. Journal of Computational and Applied Mathematics, 272(??):362–376, December 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271300616X.

Tang:2014:PAS

[jTW14] Guo ji Tang and Xing Wang. A perturbed algorithm for a system of variational inclusions involving H(·, ·)-accretive operators in Banach spaces. Journal of Computational and Applied Mathematics, 272(??):1–7, December 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714002222.

Jun:2013:CCD

[Jun13] Doobae Jun. Continuity correction for discrete barrier options with two barriers. Journal of Computational and Applied Mathematics, 237(1):520–528, January 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712002567.

Jovanovic:2010:NST

[JV10] Bosko S. Jovanović and Lubin G. Vulkov. Numerical solution of a twodimensional hyperbolic transmission problem. Journal of Computational and Applied Mathematics, 235(3):519– 534, December 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042710003353.

Jin:2010:CFE

[JW10a] Jicheng Jin and Xiaonan Wu. Convergence of a finite element scheme for the two-dimensional time-dependent Schrödinger equation in a long strip. Journal of Computational and Applied Mathematics, 234(3):777–793, June 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042710000555.

Jin:2010:SFS

[JW10b] Zhong Jin and Yuqing Wang. A simple feasible SQP method for inequality constrained optimization with global and superlinear convergence. Journal of Computational and Applied Mathematics, 233(11):3060– 3073, April 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709008085.

Jiang:2012:SCG

[JW12] Hong Jiang and Paul Wilford. A stochastic conjugate gradient method for the approximation of functions. Journal of Computational and Applied Mathematics, 236(9):2529–2544, March 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711006261.

Jia:2013:NSP

[JWL13] Zhigang Jia, Musheng Wei, and Sitao Ling. A new structure-preserving method for quaternion Hermitian eigenvalue problems. Journal of Computational and Applied Mathematics, 239(??):12–24, February 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712003822.

Jia:2010:PPR

[JWW10] Zhigang Jia, Qian Wang, and Musheng Wei. Procrustes problems for (P, Q, η)- reflexive matrices. Journal of Computational and Applied Mathematics, 233(11): 3041–3045, April 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042709008048.

Jin:2014:RLR

[JWW14] Zheng-Fen Jin, Qiuyu Wang, and Zhongping Wan. Recovering low-rank matrices from corrupted observations via the linear conjugate gradient algorithm. Journal of Computational and Applied Mathematics, 256(??):114–120, January 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713003531.

Jin:2011:NSQ

[JWY11] Zhuo Jin, Yumin Wang, and G. Yin. Numerical solutions of quantile hedging for guaranteed minimum death benefits under a regime-switching jump-diffusion formulation. Journal of Computational and Applied Mathematics, 235(8):2842–2860, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006539.

Jin:2014:GBM

[JWY14] Shi Jin, Dongming Wei, and Dongsheng Yin. Gaussian beam methods for the Schrödinger equation with discontinuous potentials. Journal of Computational and Applied Mathematics, 265(??):199–219, August 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713004834.

Jiang:2010:FFG

[JX10] Ying Jiang and Yuesheng Xu. Fast Fourier– Galerkin methods for solving singular boundary integral equations: Numerical integration and precondition. Journal of Computational and Applied Mathematics, 234(9):2792–2807, September 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710000269.

Jain:2013:KCA

[JY13] Aastha Jain and Laurent Younes. A kernel class allowing for fast computations in shape spaces induced by diffeomorphisms. Journal of Computational and Applied Mathematics, 245(??):162– 181, June 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712004566.

Ji:2010:GSP

[JZ10] Lanpeng Ji and Chunsheng Zhang. The Gerber–Shiu penalty functions for two classes of renewal risk processes. Journal of Computational and Applied Mathematics, 233(10):2575–2589, March 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709007353.

Jaklic:2011:PCI

[JŽ11] Gašper Jaklič and Emil Žagar. Planar cubic G^{1} interpolatory splines with small strain energy. Journal of Computational and Applied Mathematics, 235(8):2758–2765, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006436.

Karakaya:2010:FSU

[KA10] Vatan Karakaya and Muhammed Altun. Fine spectra of upper triangular double-band matrices. Journal of Computational and Applied Mathematics, 234(5):1387–1394, July 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710000968.

Kizilkan:2012:SSS

[KA12] Gülnur Çelik Kizilkan and Kemal Aydin. Step size strategies for the numerical integration of systems of differential equations. Journal of Computational and Applied Mathematics, 236(15):3805–3816, September 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711003682.

Kobis:2014:NSP

[KA14] M. A. Köbis and M. Arnold.

Numerical solution of penalty formulations for constrained mechanical systems using heterogeneous multiscale methods. Journal of Computational and Applied Mathematics, 262(??):193– 204, May 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003737.

Kerger:2011:ERS

[KAE^{+}11] F. Kerger, P. Archambeau, S. Erpicum, B. J. Dewals, and M. Pirotton. An exact Riemann solver and a Godunov scheme for simulating highly transient mixed flows. Journal of Computational and Applied Mathematics, 235(8):2030–2040, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710005558.

Kooi:2013:BAF

[KAS13] Bob W. Kooi, Maíra Aguiar, and Nico Stollenwerk. Bifurcation analysis of a family of multistrain epidemiology models. Journal of Computational and Applied Mathematics, 252(??):148–158, November 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003305.

Khalil:2014:NDF

[KAYS14] R. Khalil, M. Al Horani, A. Yousef, and M. Sababheh. A new definition of fractional derivative. Journal of Computational and Applied Mathematics, 264 (??):65–70, July 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042714000065.

Komori:2011:SEW

[KB11] Yoshio Komori and Kevin Burrage. Supplement: Efficient weak second order stochastic Runge–Kutta methods for non-commutative Stratonovich stochastic differential equations. Journal of Computational and Applied Mathematics, 235(17): 5326–5329, July 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042711002159.

Komori:2012:WSO

[KB12] Yoshio Komori and Kevin Burrage. Weak second order S-ROCK methods for Stratonovich stochastic differential equations. Journal of Computational and Applied Mathematics, 236(11): 2895–2908, May 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042712000441.

Komori:2013:SFO

[KB13] Yoshio Komori and Kevin Burrage. Strong first order S-ROCK methods for stochastic differential equations. Journal of Computational and Applied Mathematics, 242(??):261–274, April 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712004669.

Kouadri:2014:ATE

[KBKR14] Abdelmalek Kouadri, Abderazak Bensmail, Aissa Kheldoun, and Larbi Refoufi. An adaptive threshold estimation scheme for abrupt changes detection algorithm in a cement rotary kiln. Journal of Computational and Applied Mathematics, 259 (part B)(??): 835–842, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704271300383X.

Kang:2010:EPN

[KC10] Shugui Kang and Sui Sun Cheng. Existence of positive, negative and signchanging periodic solutions for a class of integral equations. Journal of Computational and Applied Mathematics, 234(2):518–525, May 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709008607.

Kim:2011:IPA

[KC11] Min-Kyung Kim and Gyeong-Mi Cho. Interior point algorithm for P_{∗} nonlinear complementarity problems. Journal of Computational and Applied Mathematics, 235(13):3751–3759, May 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271100032X.

Khater:2013:NPS

[KCBA13] A. H. Khater, D. K. Callebaut, A. H. Bhrawy, and M. A. Abdelkawy. Nonlinear periodic solutions for isothermal magnetostatic atmospheres. Journal of Computational and Applied Mathematics, 242(??):28– 40, April 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712004384.

Kubacki:2010:CHT

[KD10] SŁAwomir Kubacki and Erik Dick. Convective heat transfer predictions in an axisymmetric jet impinging onto a flat plate using an improved k-ω model. Journal of Computational and Applied Mathematics, 234(7):2327–2335, August 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709005706.

Kramer:2013:DFA

[KDG^{+}13] Lukas Krämer, Edoardo Di Napoli, Martin Galgon, Bruno Lang, and Paolo Bientinesi. Dissecting the FEAST algorithm for generalized eigenproblems. Journal of Computational and Applied Mathematics, 244 (??):1–9, May 15, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042712005079.

Kaveh:2011:AEL

[KF11] A. Kaveh and H. Fazli. Approximate eigensolution of Laplacian matrices for locally modified graph products. Journal of Computational and Applied Mathematics, 236(6):1591–1603, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711005012.

Kebriaei:2013:NMI

[KFR13] R. Kebriaei, J. Frischkorn, and S. Reese. A new method of intelligent control for system stabilization in process-integrated powder coating by radial axial rolling of rings. Journal of Computational and Applied Mathematics, 246(??):301– 311, July 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712002828.

Kuzmin:2012:FCT

[KG12] Dmitri Kuzmin and Yuliya Gorb. A flux-corrected transport algorithm for handling the close-packing limit in dense suspensions. Journal of Computational and Applied Mathematics, 236(18):4944–4951, December 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711005668.

Kazemi:2013:EEF

[KG13] B. Fakhr Kazemi and F. Ghoreishi. Error estimate in fractional differential equations using multiquadratic radial basis functions. Journal of Computational and Applied Mathematics, 245(??):133– 147, June 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712005420.

Kara:2014:MBR

[KG14] Emel Kizilok Kara and Omer L. Gebizlioğlu. Measurement of bivariate risks by the north-south quantile points approach. Journal of Computational and Applied Mathematics, 255(??):208– 215, January 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713002586.

Kamrani:2010:RCG

[KH10] Minoo Kamrani and S. Mohammad Hosseini. The role of coefficients of a general SPDE on the stability and convergence of a finite difference method. Journal of Computational and Applied Mathematics, 234(5):1426– 1434, July 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710001007.

Kattelans:2011:MMC

[KH11] Thorsten Kattelans and Wilhelm Heinrichs. Mass and momentum conservation of the least-squares spectral collocation method for the Navier–Stokes equations. Journal of Computational and Applied Mathematics, 236(6):1193–1215, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004432.

Kortschak:2013:EST

[KH13] Dominik Kortschak and Enkelejd Hashorva. Efficient simulation of tail probabilities for sums of logelliptical risks. Journal of Computational and Applied Mathematics, 247(??):53– 67, August 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713000241.

Khalsaraei:2010:IPR

[Kha10] M. Mehdizadeh Khalsaraei.

An improvement on the positivity results for 2-stage explicit Runge–Kutta methods. Journal of Computational and Applied Mathematics, 235(1):137–143, November 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710002967.

Kong:2011:SSP

[KHFC11] Linghua Kong, Jialin Hong, Fangfang Fu, and Jing Chen. Symplectic structurepreserving integrators for the two-dimensional Gross– Pitaevskii equation for BEC. Journal of Computational and Applied Mathematics, 235(17):4937–4948, July 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711002135.

Kilicman:2014:RSK

[KHKM14] A. Kiliçman, Ishak Hashim, M. Tavassoli Kajani, and Mohammad Maleki. On the rational second kind Chebyshev pseudospectral method for the solution of the Thomas–Fermi equation over an infinite interval. Journal of Computational and Applied Mathematics, 257(??):79– 85, February 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713003956.

Khoromskij:2010:FAT

[Kho10] Boris N. Khoromskij. Fast and accurate tensor approximation of a multivariate convolution with linear scaling in dimension. Journal of Computational and Applied Mathematics, 234(11):3122– 3139, October 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710000750.

Khodabakhshi:2011:SES

[Kho11] M. Khodabakhshi. Superefficiency in stochastic data envelopment analysis: an input relaxation approach. Journal of Computational and Applied Mathematics, 235(16):4576–4588, June 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710001810.

Khrushchev:2010:GTM

[Khr10] S. Khrushchev. The great theorem of A. A. Markoff and Jean Bernoulli sequences. Journal of Computational and Applied Mathematics, 233(6):1548–1553, January 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709001605.

Kim:2012:CTT

[Kim12]  Ku-Jin Kim.  Circles

in torus-torus intersections. Journal of Computational and Applied Mathematics, 236(9):2387–2397, March 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711005954.

Kirlar:2011:ECE

[Kır11] Barış Bülent Kırlar. On the elliptic curves y^{2} = x^{3} − c with embedding degree one. Journal of Computational and Applied Mathematics, 235(16):4724–4728, June 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710004589.

Krejic:2013:LSM

[KK13] Natasa Krejić and Natasa Krklec. Line search methods with variable sample size for unconstrained optimization. Journal of Computational and Applied Mathematics, 245(??):213– 231, June 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713000125.

Kulikova:2014:AOS

[KK14] M. V. Kulikova and G. Yu.

Kulikov. Adaptive ODE solvers in extended Kalman filtering algorithms. Journal of Computational and Applied Mathematics, 262(??): 205–216, May 15, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713005190.

Kim:2012:TAQ

[KKK12] Jerim Kim, Jeongsim Kim, and Bara Kim. Tail asymptotics of the queue size distribution in the M/M/m retrial queue. Journal of Computational and Applied Mathematics, 236(14):3445– 3460, August 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712001550.

Khorram:2014:NMC

[KKK14] E. Khorram, K. Khaledian, and M. Khaledyan. A numerical method for constructing the Pareto front of multi-objective optimization problems. Journal of Computational and Applied Mathematics, 261(??):158– 171, May 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713006262.

Kinoshita:2011:PEI

[KKN11] T. Kinoshita, T. Kimura, and M. T. Nakao. A posteriori estimates of inverse operators for initial value problems in linear ordinary differential equations. Journal of Computational and Applied Mathematics, 236(6):1622–1636, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271100505X.

Kolybasova:2011:ADF

[KKPV11] V. V. Kolybasova, P. A. Krutitskii, K. V. Prozorov, and G. Vainikko. Application of the discrete Fourier transform to solving the Cauchy integral equation. Journal of Computational and Applied Mathematics, 235(5):1317–1325, January 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710004553.

Kalinnik:2011:ETS

[KKR11] Natalia Kalinnik, Matthias Korch, and Thomas Rauber. An efficient time-step-based self-adaptive algorithm for predictor–corrector methods of Runge–Kutta type. Journal of Computational and Applied Mathematics, 236(3):394–410, September 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004109.

Kim:2010:PEE

[KL10a] Kwang-Yeon Kim and Hyung-Chun Lee. A posteriori error estimators for nonconforming finite element methods of the linear elasticity problem. Journal of Computational and Applied Mathematics, 235(1):186– 202, November 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710003080.

Kwok:2010:DSD

[KL10b] Ting-On Kwok and Leevan Ling. Dimension-splitting data points redistribution for meshless approximation. Journal of Computational and Applied Mathematics, 235(3):736–746, December 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710003791.

Kosinka:2011:UPH

[KL11] Jirí Kosinka and Miroslav Lávicka. A unified Pythagorean hodograph approach to the medial axis transform and offset approximation. Journal of Computational and Applied Mathematics, 235(12):3413–3424, April 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711000744.

Korman:2012:CAS

[KL12] Philip Korman and Yi Li. A computer assisted study of uniqueness of ground state solutions. Journal of Computational and Applied Mathematics, 236(11):2838– 2843, May 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712000313.

Kuhn:2014:CAS

[KL14] Karen Kuhn and Jens Lang. Comparison of the asymptotic stability for multirate Rosenbrock methods. Journal of Computational and Applied Mathematics, 262(??):139–149, May 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003749.

Kong:2013:HIP

[KLK13] Jae Hoon Kong, Hyun Chol Lee, and Gwang Il Kim. C^{1} Hermite interpolation with PH curves by boundary data modification. Journal of Computational and Applied Mathematics, 248(??):47–60, August 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713000551.

Kim:2011:ESP

[KLL11a] Chan-Gyun Kim, Eun Kyoung Lee, and Yong-Hoon Lee. Existence of the second positive radial solution for a p-Laplacian problem. Journal of Computational and Applied Mathematics, 235(13):3743–3750, May 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711000318.

Kovacs:2011:SAS

[KLL11b] Mihály Kovács, Fredrik Lindgren, and Stig Larsson. Spatial approximation of stochastic convolutions. Journal of Computational and Applied Mathematics, 235(12):3554–3570, April 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711000860.

Kulikov:2014:ANA

[KLM14] G. Yu. Kulikov, P. M. Lima, and M. L. Morgado. Analysis and numerical approximation of singular boundary value problems with the p-Laplacian in fluid mechanics. Journal of Computational and Applied Mathematics, 262(??):87– 104, May 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713005268.

Kloeden:2011:EIS

[KLNS11] P. E. Kloeden, G. J. Lord, A. Neuenkirch, and T. Shardlow. The exponential integrator scheme for stochastic partial differential equations: Pathwise error bounds. Journal of Computational and Applied Mathematics, 235(5):1245– 1260, January 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042710004498.

Kuzmin:2010:GOM

[KM10] Dmitri Kuzmin and Matthias Möller. Goal-oriented mesh adaptation for fluxlimited approximations to steady hyperbolic problems. Journal of Computational and Applied Mathematics, 233(12):3113–3120, April 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709004294.

Kawai:2011:STS

[KM11] Reiichiro Kawai and Hiroki Masuda. On simulation of tempered stable random variates. Journal of Computational and Applied Mathematics, 235(8):2873–2887, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006643.

Karban:2012:HCM

[KMD12] Pavel Karban, Frantisek Mach, and Ivo Dolezel. Hard-coupled model of local direct resistance heating of thin sheets. Journal of Computational and Applied Mathematics, 236(18):4725– 4731, December 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712001033.

Konig:2012:CEB

[KMG12] Gerhard König, Michael Moldaschl, and Wilfried N. Gansterer. Computing eigenvectors of block tridiagonal matrices based on twisted block factorizations. Journal of Computational and Applied Mathematics, 236(15):3696–3703, September 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711003967.

Khelifi:2014:HMM

[KMHM14] S. Chaabane Khelifi, N. Méchitoua, F. Hülsemann, and F. Magoulès. A hybrid multigrid method for convection–diffusion problems. Journal of Computational and Applied Mathematics, 259 (part B)(??): 711–719, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042713002598.

Kuraz:2010:ATD

[KMLT10] Michal Kuráz, Petr Mayer, Matej Leps, and Dagmar Trpkosová. An adaptive time discretization of the classical and the dual porosity model of Richards’ equation. Journal of Computational and Applied Mathematics, 233(12):3167–3177, April 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709008036.

Kuraz:2014:SNR

[KMP14] Michal Kuraz, Petr Mayer, and Pavel Pech. Solving the nonlinear Richards equation model with adaptive domain decomposition. Journal of Computational and Applied Mathematics, 270(??):2–11, November 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714001502.

Kyanfar:2014:CSG

[KMS14] Faranges Kyanfar, Mahmoud Mohseni Moghadam, and Abbas Salemi. Complete stagnation of GM-RES for normal matrices. Journal of Computational and Applied Mathematics, 263(??):417–422, June 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713006912.

Kinoshita:2010:VAE

[KN10a] Takehiko Kinoshita and Mitsuhiro T. Nakao. On very accurate enclosure of the optimal constant in the a priori error estimates for H_{0}^{2}-projection. Journal of Computational and Applied Mathematics, 234(2):526– 537, May 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709008619.

Kohno:2010:LEN

[KN10b] Toshiyuki Kohno and Hiroshi Niki. Letter to the Editor: a note on the preconditioned Gauss–Seidel (GS) method for linear systems. Journal of Computational and Applied Mathematics, 233(9):2413–2421, March 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709004671.

Knizhnerman:2010:SLR

[Kni10] Leonid Knizhnerman. Sensitivity of the Lanczos recurrence to Gaussian quadrature data: How malignant can small weights be? Journal of Computational and Applied Mathematics, 233(5):1238–1244, January 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709001782.

Kindermann:2012:SVD

[KNR12] Stefan Kindermann, Andreas Neubauer, and Ronny Ramlau. A singular value decomposition for the Shack–Hartmann based wavefront reconstruction. Journal of Computational and Applied Mathematics, 236(8):2186–2199, February 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042711005206.

Kobayashi:2010:NCG

[KNY10] Michiya Kobayashi, Yasushi Narushima, and Hiroshi Yabe. Nonlinear conjugate gradient methods with structured secant condition for nonlinear least squares problems. Journal of Computational and Applied Mathematics, 234(2): 375–397, May 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709008486.

Kaanoglu:2011:SPG

[KÖ11] Cem Kaanoğlu and M. Ali Özarslan. Some properties of generalized multiple Hermite polynomials. Journal of Computational and Applied Mathematics, 235(16):4878–4887, June 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711000756.

Kanamori:2013:BEQ

[KO13] Takafumi Kanamori and Atsumi Ohara. A Bregman extension of quasi-Newton updates II: Analysis of robustness properties. Journal of Computational and Applied Mathematics, 253(??):104– 122, December 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042713001817.

Kocak:2011:OFO

[Koç11] Mehmet Çetin Koçak. Ostrowski’s fourth-order iterative method speedily solves cubic equations of state. Journal of Computational and Applied Mathematics, 235(16):4736–4741, June 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710005042.

Koko:2011:UBR

[Kok11] Jonas Koko. Uzawa block relaxation method for the unilateral contact problem. Journal of Computational and Applied Mathematics, 235(8):2343–2356, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710005984.

Knibbe:2011:GIH

[KOV11] H. Knibbe, C. W. Oosterlee, and C. Vuik. GPU implementation of a Helmholtz Krylov solver preconditioned by a shifted Laplace multigrid method. Journal of Computational and Applied Mathematics, 236(3): 281–293, September 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704271100416X.

Koeune:2010:ICM

[KP10a] R. Koeune and J.-P. Ponthot. An improved constitutive model for the numerical simulation of semi-solid thixoforming. Journal of Computational and Applied Mathematics, 234(7):2287– 2296, August 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709005652.

Ku:2010:PEE

[KP10b] JaEun Ku and Eun-Jae Park. A posteriori error estimators for the firstorder least-squares finite element method. Journal of Computational and Applied Mathematics, 235(1):293– 300, November 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042710003481.

Kim:2011:SBS

[KP11a] Minho Kim and Jörg Peters. Symmetric box-splines on root lattices. Journal of Computational and Applied Mathematics, 235(14):3972– 3989, May 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S037704271000645X.

Kovtanyuk:2011:BVP

[KP11b] A. E. Kovtanyuk and I. V. Prokhorov. A boundaryvalue problem for the polarized-radiation transfer equation with Fresnel interface conditions for a layered medium. Journal of Computational and Applied Mathematics, 235(8):2006–2014, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710005522.

Karciauskas:2013:NUI

[KP13] Kestutis Karciauskas and Jörg Peters. Non-uniform interpolatory subdivision via splines. Journal of Computational and Applied Mathematics, 240(??):31– 41, March 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712002944.

Kacewicz:2014:OSC

[KP14a] Bolesław Kacewicz and Pawel Przybylowicz. Optimal solution of a class of non-autonomous initialvalue problems with unknown singularities. Journal of Computational and Applied Mathematics, 261(??): 364–377, May 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704271300647X.

Keppens:2014:SHP

[KP14b] Rony Keppens and Oliver Porth. Scalar hyperbolic PDE simulations and coupling strategies. Journal of Computational and Applied Mathematics, 266(??):87– 101, August 15, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042714000387.

Kouibia:2014:AVF

[KP14c] A. Kouibia and M. Pasadas. Approximation of vector fields using discrete div-rot variational splines in a finite element space. Journal of Computational and Applied Mathematics, 255(??):778– 788, January 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042713003233.

Kahramaner:2014:HMR

[KPA14] Yasemin Kahramaner, Yasar Polatoğlu, and Melike Aydoğan. Harmonic mappings related to Janowski starlike functions. Journal of Computational and Applied Mathematics, 270(??): 564–570, November 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042713004615.

Kouibia:2013:CGS

[KPBN13] A. Kouibia, M. Pasadas, Z. Belhaj, and K. Najib. Construction of global surfaces by variational evolutionary PDE splines. Journal of Computational and Applied Mathematics, 243(??):80– 90, May 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271200502X.

Kim:2012:CEC

[KPKK12] Sang Dong Kim, Xiangfan Piao, Do Hyung Kim, and Philsu Kim. Convergence on error correction methods for solving initial value problems. Journal of Computational and Applied Mathematics, 236(17):4448–4461, November 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712001756.

Kouibia:2011:OPC

[KPR11] A. Kouibia, M. Pasadas, and M. L. Rodríguez. Optimization of parameters for curve interpolation by cubic splines. Journal of Computational and Applied Mathematics, 235(14):4187–4198, May 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711001439.

Kim:2011:MML

[KPsQ11] Kwang Ik Kim, Hyun Suk Park, and Xiao song Qian. A mathematical modeling for the lookback option with jumpdiffusion using binomial tree method. Journal of Computational and Applied Mathematics, 235(17):5140–5154, July 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711002494.

Krajnc:2014:CLD

[KPV14] Marjeta Krajnc, Karla Pockaj, and Vito Vitrih. Construction of low degree rational motions. Journal of Computational and Applied Mathematics, 256(??):92– 103, January 15, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042713003580.

Kumam:2010:HIS

[KPW10] Poom Kumam, Narin Petrot, and Rabian Wangkeeree. A hybrid iterative scheme for equilibrium problems and fixed point problems of asymptotically kstrict pseudo-contractions. Journal of Computational and Applied Mathematics, 233(8):2013–2026, February 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709006608.

Kundeti:2010:FAP

[KR10] Vamsi Kundeti and Sanguthevar Rajasekaran. Fast algorithms for placing large entries along the diagonal of a sparse matrix. Journal of Computational and Applied Mathematics, 235(3):756– 764, December 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704271000381X.

Kumar:2014:ROS

[KR14] Sunil Kumar and S. Chandra Sekhara Rao. A robust overlapping Schwarz domain decomposition algorithm for time-dependent singularly perturbed reaction– diffusion problems. Journal of Computational and Applied Mathematics, 261(??): 127–138, May 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713006171.

Kress:1995:NSH

[Kre95] Rainer Kress. On the numerical solution of a hypersingular integral equation in scattering theory. Journal of Computational and Applied Mathematics, 61(3):345–360, August 21, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/0377042794000737.

See corrigendum [Kre11].

Kress:2011:CNS

[Kre11] Rainer Kress. Corrigendum to “On the numerical solution of a hypersingular integral equation in scattering theory” [J. Comput. Appl. Math. 61 (1995) 345–360]. Journal of Computational and Applied Mathematics, 235(9):3109, March 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710005364.

See [Kre95].

Kroon:2010:NFM

[Kro10] Martin Kroon. A numerical framework for material characterisation of inhomogeneous hyperelastic membranes by inverse analysis. Journal of Computational and Applied Mathematics, 234(2):563–578, May 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709008759.

Kudryashov:2011:CNT

[KRS11] Nikolai A. Kudryashov, Pavel N. Ryabov, and Dmitry I. Sinelshchikov. Comment on “New types of exact solutions for nonlinear Schrödinger equation with cubic nonlinearity”. Journal of Computational and Applied Mathematics, 235(15):4513–4515, June 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711001166.

See [EK11].

Katriel:2010:REI

[KS10a] Jacob Katriel and K. D.

Sen. Relativistic effects on information measures for hydrogen-like atoms. Journal of Computational and Applied Mathematics, 233(6):1399–1415, January 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709001435.

Khuri:2010:NAS

[KS10b] S. A. Khuri and A. Sayfy.

A numerical approach for solving an extended Fisher– Kolomogrov–Petrovskii–Piskunov equation. Journal of Computational and Applied Mathematics, 233(8):2081– 2089, February 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042709006657.

Kaslik:2012:ANM

[KS12a] Eva Kaslik and Seenith Sivasundaram. Analytical and numerical methods for the stability analysis of linear fractional delay differential equations. Journal of Computational and Applied Mathematics, 236(16):4027– 4041, October 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S037704271200129X.

Kumar:2012:FEA

[KS12b] Naresh Kumar and Mukul Shukla. Finite element analysis of multi-particle impact on erosion in abrasive water jet machining of titanium alloy. Journal of Computational and Applied Mathematics, 236(18):4600–4610, December 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712001902.

Kvasov:2012:LGG

[KS12c] Dmitri E. Kvasov and Yaroslav D. Sergeyev. Lipschitz gradients for global optimization in a one-point-based partitioning scheme. Journal of Computational and Applied Mathematics, 236(16):4042–4054, October 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712000775.

Kilic:2013:IBM

[KS13] Emrah Kiliç and Pantelimon Stanica. The inverse of banded matrices. Journal of Computational and Applied Mathematics, 237(1):126–135, January 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003081.

Kaya:2014:SDC

[KS14a] Kamer Kaya and Ali Aydin Selçuk. Sharing DSS by the Chinese remainder theorem. Journal of Computational and Applied Mathematics, 259 (part B)(??):495–502, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003075.

Kestler:2014:FES

[KS14b] Sebastian Kestler and Rob Stevenson. Fast evaluation of system matrices w.r.t. multi-tree collections of tensor product refinable basis functions. Journal of Computational and Applied Mathematics, 260(??):103– 116, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713004706.

Kiili:2014:TWG

[KS14c] Hanna Kiili and Valery Serov. TM-waves guided by a nonlinear film with complex permittivity. Journal of Computational and Applied Mathematics, 271(??):339– 355, December 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042714002179.

Kou:2014:AFE

[KS14d] Jisheng Kou and Shuyu Sun. An adaptive finite element method for simulating surface tension with the gradient theory of fluid interfaces. Journal of Computational and Applied Mathematics, 255(??):593–604, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713003154.

Kus:2014:ALH

[KSA14] Pavel Kus, Pavel Solin, and David Andrs. Arbitrarylevel hanging nodes for adaptive hp–FEM approximations in 3D. Journal of Computational and Applied Mathematics, 270(??):121– 133, November 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042714000867.

Kasimbeyli:2011:TOM

[KSK11] Nergiz Kasimbeyli, Tugba Sarac, and Refail Kasimbeyli. A two-objective mathematical model without cutting patterns for one-dimensional assortment problems. Journal of Computational and Applied Mathematics, 235(16):4663– 4674, June 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710004176.

Klus:2011:EAP

[KSLD11] Stefan Klus, Tuhin Sahai, Cong Liu, and Michael Dellnitz. An efficient algorithm for the parallel solution of high-dimensional differential equations. Journal of Computational and Applied Mathematics, 235(9):3053– 3062, March 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704271000676X.

Kucerova:2012:AUU

[KSRM12] Anna Kucerová, Jan Sýkora, Bojana Rosić, and Hermann G. Matthies. Acceleration of uncertainty updating in the description of transport processes in heterogeneous materials. Journal of Computational and Applied Mathematics, 236(18):4862–4872, December 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271200060X.

Kretzschmar:2014:DGM

[KSTW14] Fritz Kretzschmar, Sascha M. Schnepp, Igor Tsukerman, and Thomas Weiland. Discontinuous Galerkin methods with Trefftz approximations. Journal of Computational and Applied Mathematics, 270(??):211–222, November 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714000612.

Kuzmanovic:2013:OSP

[KT13] Ivana Kuzmanović and Ninoslav Truhar. Optimization of the solution of the parameter-dependent Sylvester equation and applications. Journal of Computational and Applied Mathematics, 237(1):136– 144, January 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712003123.

Ku:2013:LSS

[Ku13] JaEun Ku. Least-squares solutions as solutions of a perturbation form of the Galerkin methods: Interior pointwise error estimates and pollution effect. Journal of Computational and Applied Mathematics, 251(??):67–80, October 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713001271.

Kudryashov:2010:NNE

[Kud10] Nikolai A. Kudryashov. A note on new exact solutions for the Kawahara equation using exp-function method. Journal of Computational and Applied Mathematics, 234(12):3511–3512, October 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710002426.

Kurics:2010:EOP

[Kur10]  Tamás Kurics.  Equiv-

alent operator preconditioning for elliptic problems with nonhomogeneous mixed boundary conditions.

Journal of Computational and Applied Mathematics, 235(2):437–449, November 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710003377.

Kuzmin:2010:VBH

[Kuz10] Dmitri Kuzmin. A vertexbased hierarchical slope limiter for p-adaptive discontinuous Galerkin methods. Journal of Computational and Applied Mathematics, 233(12):3077–3085, April 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709003318.

Kuzmin:2012:LPF

[Kuz12] Dmitri Kuzmin. Linearitypreserving flux correction and convergence acceleration for constrained Galerkin schemes. Journal of Computational and Applied Mathematics, 236 (9):2317–2337, March 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042711005899.

Kuznetsov:2013:AAH

[Kuz13]  A. Kuznetsov.  Asymp-

totic approximations to the Hardy–Littlewood function. Journal of Computational and Applied Mathematics, 237(1):603–613, January 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712002804.

Kuzmin:2014:OBA

[Kuz14]  Dmitri Kuzmin.  An

optimization-based approach to enforcing mass conservation in level set methods. Journal of Computational and Applied Mathematics, 258(??):78– 86, March 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713004573.

Kosmas:2012:LPF

[KV12] O. T. Kosmas and D. S.

Vlachos. Local path fitting: a new approach to variational integrators. Journal of Computational and Applied Mathematics, 236(10): 2632–2642, April 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711006352.

Kvasov:2011:PMM

[Kva11]  Boris Kvasov.  Parallel

mesh methods for tension splines. Journal of Computational and Applied Mathematics, 236(5):843–859, October 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711002664.

Kulikov:2010:DQC

[KW10] G. Yu. Kulikov and R. Weiner. Doubly quasi-consistent parallel explicit peer methods with built-in global error estimation. Journal of Computational and Applied Mathematics, 233(9):2351– 2364, March 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709007183.

Kulikov:2011:GEE

[KW11] G. Yu. Kulikov and R. Weiner. Global error estimation and control in linearlyimplicit parallel two-step peer W -methods. Journal of Computational and Applied Mathematics, 236(6):1226– 1239, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042711004456.

Kong:2011:AFS

[KWW11] Zhi Kong, Lifu Wang, and Zhaoxia Wu. Application of fuzzy soft set in decision making problems based on grey theory. Journal of Computational and Applied Mathematics, 236(6):1521–1530, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271100495X.

Kail:2014:IEL

[KWXZ14] Andrew Kail, Kwai Wong, Henian Xia, and Xiaopeng Zhao. Interoperable executive library for the simulation of biomedical processes. Journal of Computational and Applied Mathematics, 270(??):257–274, November 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714000259.

Kang:2013:EQH

[KX13] Hongchao Kang and Shuhuang Xiang. Efficient quadrature of highly oscillatory integrals with algebraic singularities. Journal of Computational and Applied Mathematics, 237(1):576–588, January 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712002725.

Kang:2013:CIO

[KXH13] Hongchao Kang, Shuhuang Xiang, and Guo He. Computation of integrals with oscillatory and singular integrands using Chebyshev expansions. Journal of Computational and Applied Mathematics, 242(??):141– 156, April 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712004426.

Korn:2013:EBM

[KZ13] Ralf Korn and Serkan Zeytun. Efficient basket Monte Carlo option pricing via a simple analytical approximation. Journal of Computational and Applied Mathematics, 243(??):48– 59, May 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712004827.

Lapierre:2010:PEV

[LA10] Fabian D. Lapierre and Marc Acheroy. Performance enhancement and validation of the open-source software for modeling of ship infrared signatures (OSMO-SIS). Journal of Computational and Applied Mathematics, 234(7):2342–2349, August 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704270900572X.

Li:2011:DPS

[LA11] C. K. Li and M. A. Aguirre.

The distributional products on spheres and Pizetti’s formula. Journal of Computational and Applied Mathematics, 235(5):1482–1489, January 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710004814.

Lakoba:2013:SLH

[Lak13] T. I. Lakoba. Spurious localized highest-frequency modes in Schrödinger-type equations solved by finitedifference methods. Journal of Computational and Applied Mathematics, 245(??): 117–120, June 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712005468.

Lamichhane:2011:SMF

[Lam11] Bishnu P. Lamichhane.

A stabilized mixed finite element method for the biharmonic equation based on biorthogonal systems. Journal of Computational and Applied Mathematics, 235(17):5188–5197, July 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711002524.

Lamichhane:2014:MFE

[Lam14] Bishnu P. Lamichhane. A mixed finite element method for nearly incompressible elasticity and Stokes equations using primal and dual meshes with quadrilateral and hexahedral grids. Journal of Computational and Applied Mathematics, 260(??):356– 363, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713005116.

Lang:2010:LET

[Lan10] Annika Lang. A Lax equivalence theorem for stochastic differential equations. Journal of Computational and Applied Mathematics, 234(12):3387–3396, October 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710002554.

Lang:2012:ASC

[Lan12] Annika Lang. Almost sure convergence of a Galerkin approximation for SPDEs of Zakai type driven by square integrable martingales. Journal of Computational and Applied Mathematics, 236(7):1724–1732, January 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711005346.

Law:2014:PWS

[Law14] K. J. H. Law. Proposals which speed up functionspace MCMC. Journal of Computational and Applied Mathematics, 262(??):127– 138, May 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003701.

Layeb:2013:HQI

[Lay13] Abdesslem Layeb. A hybrid quantum inspired harmony search algorithm for 0–1 optimization problems. Journal of Computational and Applied Mathematics, 253(??):14–25, December 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713001805.

Liu:2009:NOC

[LB09] Lihua Liu and Yuzhen Bai.

New oscillation criteria for second-order nonlinear neutral delay differential equations. Journal of Computational and Applied Mathematics, 231(2):657–663, September 15, 2009. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709002659.

See erratum [LB10].

Liu:2010:ENO

[LB10] Lihua Liu and Yuzhen Bai. Erratum to: “New oscillation criteria for secondorder nonlinear neutral delay differential equations” [J. Comput. Appl. Math. 231 (2009) 657– 663]. Journal of Computational and Applied Mathematics, 233(10):2755, March 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic). URL http://www.
math.utah.edu/pub/tex/
bib/jcomputapplmath2005.
bib;   http://www.sciencedirect.
com/science/article/pii/
S0377042709007717. See

[LB09].

Luxemburg:2011:SNP

[LB11] Leon A. Luxemburg and Philip R. Brown. The scalar Nevanlinna–Pick interpolation problem with boundary conditions. Journal of Computational and Applied Mathematics, 235(8):2615– 2625, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042710006308.

Liu:2014:DTR

[LB14] He Liu and Zhenhua Bao.

On a discrete-time risk model with general income and time-dependent claims. Journal of Computational and Applied Mathematics, 260(??):470–481, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271300575X.

Langdon:2010:WC

[LBCL10] Steve Langdon, Nick Biggs, Peter Chamberlain, and Jing-Rebecca Li. Waves 2007 Conference. Journal of Computational and Applied Mathematics, 234(6):1621, July 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709004816.

Lesnic:2012:ISS

[LBM12] D. Lesnic and B. Bin-Mohsin. Inverse shape and surface heat transfer coefficient identification. Journal of Computational and Applied Mathematics, 236(7):1876–1891, January 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271100567X.

Lv:2010:ECM

[LC10] Xueqin Lv and Minggen Cui. An efficient computational method for linear fifth-order two-point boundary value problems. Journal of Computational and Applied Mathematics, 234 (5):1551–1558, July 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042710001342.

Li:2011:DBS

[LC11] Chong-Jun Li and Juan Chen. On the dimensions of bivariate spline spaces and the stability of the dimensions. Journal of Computational and Applied Mathematics, 236(5):765–774, October 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711003177.

Levasseur:2013:MDM

[LCCK13] S. Levasseur, F. Collin, R. Charlier, and D. Kondo.

On micromechanical damage modeling in geomechanics: Influence of numerical integration scheme. Journal of Computational and Applied Mathematics, 246(??): 215–224, July 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712002324.

Ljubisavljevic:2011:ARD

[LCI11] Jelena Ljubisavljević and Dragana S. Cvetković-Ilić. Additive results for the Drazin inverse of block matrices and applications. Journal of Computational and Applied Mathematics, 235(12):3683–3690, April 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711001178.

Li:2012:CSS

[LCN12] Wen Li, Lu-Bin Cui, and Michael K. Ng. On computation of the steady-state probability distribution of probabilistic Boolean networks with gene perturbation. Journal of Computational and Applied Mathematics, 236(16):4067–4081, October 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712000799.

Lv:2010:NNS

[LCW10] Yibing Lv, Zhong Chen, and Zhongping Wan. A neural network for solving a convex quadratic bilevel programming problem. Journal of Computational and Applied Mathematics, 234(2):505–511, May 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709008589.

Lei:2011:FAM

[LCXL11] Na Lei, Junjie Chai, Peng Xia, and Ying Li. A fast algorithm for the multivariate Birkhoff interpolation problem. Journal of Computational and Applied Mathematics, 236(6):1656–1666, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711005073.

Liao:2011:TWP

[LCZ11] Xinyuan Liao, Yuming Chen, and Shengfan Zhou. Traveling wavefronts of a prey-predator diffusion system with stage-structure and harvesting. Journal of Computational and Applied Mathematics, 235(8):2560– 2568, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042710006242.

Lakestani:2010:UCC

[LD10] Mehrdad Lakestani and Mehdi Dehghan. The use of Chebyshev cardinal functions for the solution of a partial differential equation with an unknown timedependent coefficient subject to an extra measurement. Journal of Computational and Applied Mathematics, 235(3):669–678, December 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710003729.

Luo:2010:NCM

[LDGY10] Wei Luo, Junfei Dai, Xianfeng Gu, and Shing-Tung Yau. Numerical conformal mapping of multiply connected domains to regions with circular boundaries. Journal of Computational and Applied Mathematics, 233(11):2940–2947, April 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709007857.

Lukas:2010:EAR

[LdHA10] Mark A. Lukas, Frank R. de Hoog, and Robert S. Anderssen. Efficient algorithms for robust generalized cross-validation spline smoothing. Journal of Computational and Applied Mathematics, 235(1):102– 107, November 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S037704271000292X.

Lamine:2010:MUC

[LE10] Sadok Lamine and Michael G. Edwards. Multidimensional upwind convection schemes for flow in porous media on structured and unstructured quadrilateral grids. Journal of Computational and Applied Mathematics, 234(7):2106–2117, August 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709005469.

Laghrouche:2010:CNP

[LEKT10] O. Laghrouche, A. El-Kacimi, and J. Trevelyan. A comparison of NRBCs for PUFEM in 2D Helmholtz problems at high wave numbers. Journal of Computational and Applied Mathematics, 234(6):1670–1677, July 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709004890.

Lopez-Espin:2012:TSL

[LEVG12] Jose J. López-Espín, Antonio M. Vidal, and Domingo Giménez. Two-stage least squares and indirect least squares algorithms for simultaneous equations models. Journal of Computational and Applied Mathematics, 236(15):3676–3684, September 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711003918.

Lopez-Espin:2014:HMA

[LEVG14] Jose J. López-Espín, Antonio M. Vidal, and Domingo Giménez. Heuristics and metaheuristics for accelerating the computation of simultaneous equations models through a Steiner tree. Journal of Computational and Applied Mathematics, 255(??):605–615, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003117.

Li:2010:SAS

[LF10] Xiaodi Li and Xilin Fu. Stability analysis of stochastic functional differential equations with infinite delay and its application to recurrent neural networks. Journal of Computational and Applied Mathematics, 234(2):407– 417, May 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709008504.

Li:2011:GAS

[LF11] Xiaodi Li and Xilin Fu. Global asymptotic stability of stochastic Cohen– Grossberg-type BAM neural networks with mixed delays: an LMI approach. Journal of Computational and Applied Mathematics, 235(12):3385–3394, April 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006011.

Luo:2014:FML

[LFWL14] Xingjun Luo, Linxiu Fan, Yongchao Wu, and Fanchun Li. Fast multi-level iteration methods with compression technique for solving ill-posed integral equations. Journal of Computational and Applied Mathematics, 256(??):131–151, January 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003877.

Liu:2011:UQF

[LFY11] Feng-Tao Liu, You-Hua Fan, and Jian-Hua Yin. The use of QP-free algorithm in the limit analysis of slope stability. Journal of Computational and Applied Mathematics, 235(13):3889–3897, May 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271100046X.

Lu:2011:NIC

[LG11] Li-Yong Lu and Wei-Zhe Gu. A non-interior continuation algorithm for the CP based on a generalized smoothing function. Journal of Computational and Applied Mathematics, 235(8):2300–2313, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710005935.

Li:2014:NEE

[LGM14] Mingxia Li, Xiaofei Guan, and Shipeng Mao. New error estimates of the Morley element for the plate bending problems. Journal of Computational and Applied Mathematics, 263(??):405– 416, June 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713006973.

LeGuyader:2013:GFA

[LGMA13] Carole Le Guyader, Christian Gout, Anne-Sophie Macé, and Dominique Apprato. Gradient field approximation: Application to registration in image processing. Journal of Computational and Applied Mathematics, 240(??):135–147, March 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271200324X.

Lamb:2011:SMM

[LGR11] Karen E. Lamb, David Greenhalgh, and Chris Robertson. A simple mathematical model for genetic effects in pneumococcal carriage and transmission. Journal of Computational and Applied Mathematics, 235(7):1812–1818, February 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710001676.

Lu:2010:OCT

[LGZ10] Wei Lu, Weigao Ge, and Zhihong Zhao. Oscillatory criteria for thirdorder nonlinear difference equation with impulses. Journal of Computational and Applied Mathematics, 234(12):3366–3372, October 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710002451.

Li:2011:ESS

[LH11a] Lei Li and Shihuang Hong.

Exponential stability for set dynamic equations on time scales. Journal of Computational and Applied Mathematics, 235(17):4916–4924, July 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711001919.

Lu:2011:SNE

[LH11b] Nan Lu and Zheng-Hai Huang. Solvability of Newton equations in smoothingtype algorithms for the SOCCP. Journal of Computational and Applied Mathematics, 235(8):2270–2276, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710005911.

Liu:2014:EMR

[LHC14] Lu Liu, Wei Huang, and Di-Rong Chen. Exact minimum rank approximation via Schatten p-norm minimization. Journal of Computational and Applied Mathematics, 267(??):218– 227, September 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042714001010.

Luo:2011:CCI

[LHF11] Zhongxuan Luo, Wenyu Hu, and Erbao Feng. Computing curve intersection by homotopy methods. Journal of Computational and Applied Mathematics, 236(5):892– 905, October 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042711002627.

Li:2010:SPP

[LHL10] Jian-Lei Li, Ting-Zhu Huang, and Liang Li. The spectral properties of the preconditioned matrix for nonsymmetric saddle point problems. Journal of Computational and Applied Mathematics, 235(1):270–285, November 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710003456.

Lv:2011:EHI

[LHL11] Xiao-Guang Lv, T. Z.

Huang, and Liang Li. Erratum to: “On the HSS iteration methods for positive definite Toeplitz linear systems” [J. Comput. Appl. Math. 224 (2009) 709– 718]. Journal of Computational and Applied Mathematics, 235(9):3112–3114, March 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006369.

See [GT09].

Liu:2013:SPE

[LHL13] Kai Liu, Liangjian Hu, and Jiaowan Luo. Stability property and essential spectrum of linear retarded functional differential equations. Journal of Computational and Applied Mathematics, 244(??):19– 35, May 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712005109.

Li:2013:EAM

[LHLC13] Zi-Cai Li, Hung-Tsai Huang, Ming-Gong Lee, and John Y.

Chiang. Error analysis of the method of fundamental solutions for linear elastostatics. Journal of Computational and Applied Mathematics, 251(??):133–153, October 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271300126X.

Li:2010:SWL

[LHLL10] Hou-Biao Li, Ting-Zhu Huang, Xing-Ping Liu, and Hong Li. Singularity, Wielandt’s lemma and singular values. Journal of Computational and Applied Mathematics, 234 (10):2943–2952, September 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710002025.

Liu:2010:FRI

[LHY10] Xiaoji Liu, Chunmei Hu, and Yaoming Yu. Further results on iterative methods for computing generalized inverses. Journal of Computational and Applied Mathematics, 234(3):684– 694, June 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710000099.

Li:2010:CRC

[Li10] Shengtai Li. Comparison of refinement criteria for structured adaptive mesh refinement. Journal of Computational and Applied Mathematics, 233(12):3139–3147, April 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704270900586X.

Li:2011:FEA

[Li11a]  Hengguang Li.  Finite

element analysis for the axisymmetric Laplace operator on polygonal domains. Journal of Computational and Applied Mathematics, 235(17):5155–5176, July 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711002500.

Li:2011:DDG

[Li11b] Jichun Li. Development of discontinuous Galerkin methods for Maxwell’s equations in metamaterials and perfectly matched layers. Journal of Computational and Applied Mathematics, 236(5):950–961, October 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711002342.

Li:2012:HOT

[Li12] Shengtai Li. Higher-order time integration with a local Lax–Wendroff procedure for a central scheme on an overlapping grid. Journal of Computational and Applied Mathematics, 236(18):4756– 4761, December 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042712000830.

Li:2014:SAA

[Li14] Pei-Yu Li. Sample average approximation method for a class of stochastic generalized Nash equilibrium problems. Journal of Computational and Applied Mathematics, 261(??):387–393, May 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271300633X.

Liao:2014:DSA

[Lia14] Wenyuan Liao. On the dispersion, stability and accuracy of a compact higher-order finite difference scheme for 3D acoustic wave equation. Journal of Computational and Applied Mathematics, 270(??):571– 583, November 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713004299.

Liu:2011:MGS

[Liu11a]  Qiaohua Liu.  Modified

Gram–Schmidt-based methods for block downdating the Cholesky factorization. Journal of Computational and Applied Mathematics, 235(8):1897–1905, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710004942.

Liu:2011:FMP

[Liu11b] Shiang-Tai Liu. Fuzzy measures for profit maximization with fuzzy parameters. Journal of Computational and Applied Mathematics, 236(6):1333–1342, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004651.

Liu:2011:MAD

[Liu11c]  Shiang-Tai Liu.  The

mean-absolute deviation portfolio selection problem with interval-valued returns. Journal of Computational and Applied Mathematics, 235(14):4149–4157, May 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711001385.

Liu:2012:CEG

[Liu12] Chein-Shan Liu. Computing the eigenvalues of the generalized Sturm–Liouville problems based on the Liegroup SL(2, R). Journal of Computational and Applied Mathematics, 236(17):4547– 4560, November 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042712002075.

Liu:2014:DOS

[Liu14a] Chein-Shan Liu. A doubly optimized solution of linear equations system expressed in an affine Krylov subspace. Journal of Computational and Applied Mathematics, 260(??):375– 394, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713005578.

Liu:2014:GOT

[Liu14b] Chein-Shan Liu. A globally optimal tri-vector method to solve an ill-posed linear system. Journal of Computational and Applied Mathematics, 260(??):18– 35, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271300472X.

Liu:2011:WRR

[LJ11] Jun Liu and Yao-Lin Jiang. Waveform relaxation for reaction–diffusion equations. Journal of Computational and Applied Mathematics, 235(17):5040–5055, July 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711002299.

Liu:2012:PWR

[LJ12] Jun Liu and Yao-Lin Jiang. A parareal waveform relaxation algorithm for semi-linear parabolic partial differential equations. Journal of Computational and Applied Mathematics, 236(17):4245–4263, November 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712002245.

Lu:2013:RBQ

[LJ13] Tzyy-Chyang Lu and Jyh-Ching Juang. A regionbased quantum evolutionary algorithm (RQEA) for global numerical optimization. Journal of Computational and Applied Mathematics, 239(??):1–11, February 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003792.

Li:2011:IMI

[LJCL11] Sheng-Guo Li, Hao Jiang, Li-Zhi Cheng, and Xiang-Ke Liao. IGAOR and multisplitting IGAOR methods for linear complementarity problems. Journal of Computational and Applied Mathematics, 235(9):2904– 2912, March 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710006552.

Lotfi:2010:TSG

[LJE^{+}10] F. Hosseinzadeh Lotfi, G. R. Jahanshahloo, A. Ebrahimnejad, M. Soltanifar, and S. M. Mansourzadeh. Target setting in the general combined-oriented CCR model using an interactive MOLP method. Journal of Computational and Applied Mathematics, 234(1): 1–9, May 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709007924.

Lotfi:2011:FDE

[LJMG11] F. Hosseinzadeh Lotfi, G. R.

Jahanshahloo, M. R. Mozaffari, and J. Gerami. Finding DEA-efficient hyperplanes using MOLP efficient faces. Journal of Computational and Applied Mathematics, 235(5):1227–1231, January 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710004450.

Levesley:2013:MVB

[LK13] Jeremy Levesley and Alexander Kushpel. A multiplier version of the Bernstein inequality on the complex sphere. Journal of Computational and Applied Mathematics, 240(??):184–191, March 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712004074.

Lisitsa:2010:RTT

[LL10] Vadim Lisitsa and Egor Lys.

Reflectionless truncation of target area for axially symmetric anisotropic elasticity. Journal of Computational and Applied Mathematics, 234(6):1803–1809, July 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709005068.

Li:2011:PPS

[LL11a] Yuan Li and Kaitai Li. Pressure projection stabilized finite element method for Stokes problem with nonlinear slip boundary conditions. Journal of Computational and Applied Mathematics, 235(12):3673–3682, April 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711001154.

Lian:2011:DPF

[LL11b] Yijiang Lian and Zhiping Li. A dual-parametric finite element method for cavitation in nonlinear elasticity. Journal of Computational and Applied Mathematics, 236(5):834–842, October 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711002676.

Liu:2011:PCP

[LL11c] Donghai Liu and Zaiming Liu. The perturbed compound Poisson risk model with linear dividend barrier. Journal of Computational and Applied Mathematics, 235(8):2357–2363, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271000600X.

Liu:2011:NKS

[LL11d] Liying Liu and Shaoyong Li.

A new kind of simple kennel function yielding good iteration bounds for primaldual interior-point methods. Journal of Computational and Applied Mathematics, 235(9):2944–2955, March 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271000662X.

Lee:2013:TIM

[LL13a] Jaewook Lee and Younhee Lee. Tridiagonal implicit method to evaluate European and American options under infinite activity Lévy models. Journal of Computational and Applied Mathematics, 237(1):234– 243, January 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712003184.

Liu:2013:CBVa

[LL13b] Zhenhai Liu and Jitai Liang.

A class of boundary value problems for first-order impulsive integro-differential equations with deviating arguments. Journal of Computational and Applied Mathematics, 237(1):477– 486, January 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712002531.

Li:2014:RKT

[LL14a] Li Li and Wanyu Liu. Runge–Kutta type total variation regularization for nonlinear inverse problems. Journal of Computational and Applied Mathematics, 263(??):103–114, June 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713006845.

Liu:2014:SRD

[LL14b] Hanze Liu and Jibin Li. Symmetry reductions, dynamical behavior and exact explicit solutions to the Gordon types of equations. Journal of Computational and Applied Mathematics, 257(??):144–156, February 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713004275.

Li:2011:TMU

[LLCL11] Zi-Cai Li, Ming-Gong Lee, John Y. Chiang, and Ya Ping Liu. The Trefftz method using fundamental solutions for biharmonic equations. Journal of Computational and Applied Mathematics, 235(15):4350– 4367, June 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711001543.

Liang:2010:LRI

[lLfD10] Mao lin Liang and Li fang Dai. The left and right inverse eigenvalue problems of generalized reflexive and anti-reflexive matrices. Journal of Computational and Applied Mathematics, 234(3):743–749, June 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271000018X.

Lakhdara:2014:MAO

[LLG14] Maya Lakhdara, Saïda Latreche, and Christian Gontrand. Mathematical approach and optimisation of nanometric base thickness for a SiGeC HBT dedicated to radiofrequency applications. Journal of Computational and Applied Mathematics, 259 (part B)(??):925–936, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005311.

Lim:2014:EPB

[LLK14] Hyuncheul Lim, Sungchul Lee, and Geonwoo Kim. Efficient pricing of Bermudan options using recombining quadratures. Journal of Computational and Applied Mathematics, 271(??):195– 205, December 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042714001885.

Li:2011:EEL

[LLL^{+}11] Baojun Li, Bo Li, Xiuping Liu, Zhixun Su, and Bowen Yu. Exact evaluation of limits and tangents for interpolatory subdivision surfaces at rational points. Journal of Computational and Applied Mathematics, 236(5):906–915, October 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711002615.

Li:2012:GEI

[LLL12] Qiao-Xing Li, Si-Feng Liu, and Yi Lin. Grey enterprise input-output analysis. Journal of Computational and Applied Mathematics, 236(7):1862–1875, January 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711005619.

Li:2012:GHM

[LLP12] Wen Li, Yang-Peng Liu, and Xiao-Fei Peng. The generalized HSS method for solving singular linear systems. Journal of Computational and Applied Mathematics, 236(9):2338–2353, March 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711005905.

Lee:2014:CCD

[LLS14] Spike T. Lee, Jun Liu, and Hai-Wei Sun. Combined compact difference scheme for linear second-order partial differential equations with mixed derivative. Journal of Computational and Applied Mathematics, 264(??):23– 37, July 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042714000089.

Lin:2011:IFE

[LLSW11] T. Lin, Y. Lin, W.-W. Sun, and Z. Wang. Immersed finite element methods for 4th order differential equations. Journal of Computational and Applied Mathematics, 235(13):3953–3964, May 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711000525.

Lin:2014:NEP

[LLT^{+}14] Qun Lin, Ryan Loxton, Kok Lay Teo, Yong Hong Wu, and Changjun Yu. A new exact penalty method for semi-infinite programming problems. Journal of Computational and Applied Mathematics, 261(??):271– 286, May 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713006298.

Luo:2014:CSO

[LLW14] Jiajia Luo, Wei Li, and Qin Wang. Checking strong optimality of interval linear programming with inequality constraints and nonnegative constraints. Journal of Computational and Applied Mathematics, 260(??):180– 190, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271300530X.

Lin:2014:PEE

[LLX14] Qun Lin, Fusheng Luo, and Hehu Xie. A posterior error estimator and lower bound of a nonconforming finite element method. Journal of Computational and Applied Mathematics, 265(??):243– 254, August 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713004858.

Liao:2011:MNE

[lLzS11] Hong lin Liao and Zhi zhong Sun. Maximum norm error estimates of efficient difference schemes for second-order wave equations. Journal of Computational and Applied Mathematics, 235(8):2217–2233, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710005856.

Lamour:2012:DSD

[LM12a] René Lamour and Roswitha März. Detecting structures in differential algebraic equations: Computational aspects. Journal of Computational and Applied Mathematics, 236(16):4055– 4066, October 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042712001288.

Lima:2012:ECM

[LM12b] P. M. Lima and M. L. Morgado. Efficient computational methods for singular free boundary problems using smoothing variable substitutions. Journal of Computational and Applied Mathematics, 236(12):2981– 2989, June 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711002810.

Lejay:2013:NMC

[LM13a] Antoine Lejay and Sylvain Maire. New Monte Carlo schemes for simulating diffusions in discontinuous media. Journal of Computational and Applied Mathematics, 245(??):97– 116, June 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712005444.

Liu:2013:AMB

[LM13b] Wei Liu and Xuerong Mao.

Asymptotic moment boundedness of the numerical solutions of stochastic differential equations. Journal of Computational and Applied Mathematics, 251(??):22– 32, October 15, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713001453.

Linh:2014:EIS

[LM14] Vu Hoang Linh and Volker Mehrmann. Efficient integration of strangenessfree non-stiff differentialalgebraic equations by halfexplicit methods. Journal of Computational and Applied Mathematics, 262(??):346– 360, May 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271300527X.

Langdon:2010:HFS

[LMCW10] S. Langdon, M. Mokgolele, and S. N. Chandler-Wilde. High frequency scattering by convex curvilinear polygons. Journal of Computational and Applied Mathematics, 234(6):2020–2026, July 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709005329.

Liu:2013:CBVb

[LMG13] Xiaoming Liu, Rogemar Mamon, and Huan Gao. A comonotonicity-based valuation method for guaranteed annuity options. Journal of Computational and Applied Mathematics, 250(??):58–69, October 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713000836.

Lopez:2010:FCO

[LMM^{+}10a] M. López, J. Martínez, J. M. Matías, J. Taboada, and J. A. Vilán. Functional classification of ornamental stone using machine learning techniques. Journal of Computational and Applied Mathematics, 234(4):1338– 1345, June 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042710000671.

Lopez:2010:SFO

[LMM^{+}10b] M. López, J. Martínez, J. M. Matías, J. Taboada, and J. A. Vilán. Shape functional optimization with restrictions boosted with machine learning techniques. Journal of Computational and Applied Mathematics, 234(8):2609–2615, August 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710001792.

Lamnii:2013:CQP

[LMT13] M. Lamnii, H. Mraoui, and A. Tijini. Construction of quintic Powell–Sabin spline quasi-interpolants based on blossoming. Journal of Computational and Applied Mathematics, 250(??):190– 209, October 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713001313.

Lin:2011:HMT

[LMW^{+}11] Ge Lin, Siming Meng, Ruomei Wang, Xiaonan Luo, and Yi Li. The heat and moisture transfer balance theory of garment simulation. Journal of Computational and Applied Mathematics, 236(5):980–987, October 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711001282.

Liu:2011:ADF

[LMY11] Jiangguo Liu, Lin Mu, and Xiu Ye. An adaptive discontinuous finite volume method for elliptic problems. Journal of Computational and Applied Mathematics, 235(18):5422–5431, July 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711003256.

Liu:2012:CDF

[LMYJ12] Jiangguo Liu, Lin Mu, Xiu Ye, and Rabeea Jari. Convergence of the discontinuous finite volume method for elliptic problems with minimal regularity. Journal of Computational and Applied Mathematics, 236(17):4537–4546, November 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712002105.

Lotfi:2012:CRA

[LNB^{+}12] F. Hosseinzadeh Lotfi, N. Nematollahi, M. H. Behzadi, M. Mirbolouki, and Z. Moghaddas. Centralized resource allocation with stochastic data. Journal of Computational and Applied Mathematics, 236(7):1783– 1788, January 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711005425.

Lotfi:2010:CRA

[LNJ^{+}10] F. Hosseinzadeh Lotfi, A. A.

Noora, G. R. Jahanshahloo, J. Gerami, and M. R. Mozaffari. Centralized resource allocation for enhanced Russell models. Journal of Computational and Applied Mathematics, 235(1):1–10, November 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710003055.

Luan:2014:EER

[LO14a] Vu Thai Luan and Alexander Ostermann. Explicit exponential Runge– Kutta methods of high order for parabolic problems. Journal of Computational and Applied Mathematics, 256(??):168–179, January 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003713.

See also reprint [LO14c].

Luan:2014:ERM

[LO14b] Vu Thai Luan and Alexander Ostermann. Exponential Rosenbrock methods of order five — construction, analysis and numerical comparisons. Journal of Computational and Applied Mathematics, 255(??):417– 431, January 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042713002379.

Luan:2014:REE

[LO14c] Vu Thai Luan and Alexander Ostermann. Reprint of “Explicit exponential Runge–Kutta methods of high order for parabolic problems”. Journal of Computational and Applied Mathematics, 262(??):361– 372, May 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271400003X.

See [LO14a].

Libertiaux:2010:DVI

[LP10a] V. Libertiaux and F. Pascon. Differential versus integral formulation of fractional hyperviscoelastic constitutive laws for brain tissue modelling. Journal of Computational and Applied Mathematics, 234(7):2029– 2035, August 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042709005378.

Lopez:2010:AEM

[LP10b] José L. López and Pedro Pagola. Asymptotic expansions of Mellin convolution integrals: an oscillatory case. Journal of Computational and Applied Mathematics, 233(6):1562–1569, January 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709001629.

Lopez:2010:CHF

[LP10c] José L. López and Pedro J. Pagola. The confluent hypergeometric functions M(a, b; z) and U(a, b; z) for large b and z. Journal of Computational and Applied Mathematics, 233(6):1570– 1576, January 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709001630.

Leonenko:2012:HOA

[LP12] G. M. Leonenko and T. N.

Phillips. High-order approximation of Pearson diffusion processes. Journal of Computational and Applied Mathematics, 236(11):2853– 2868, May 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712000337.

Livieris:2013:NCS

[LP13] Ioannis E. Livieris and Panagiotis Pintelas. A new class of spectral conjugate gradient methods based on a modified secant equation for unconstrained optimization. Journal of Computational and Applied Mathematics, 239(??):396–405, February 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712003718.

Le:2013:BPE

[LPDN13] Triet Minh Le, Quan Hoang Pham, Trong Duc Dang, and Tuan Huy Nguyen. A backward parabolic equation with a time-dependent coefficient: Regularization and error estimates. Journal of Computational and Applied Mathematics, 237(1):432–441, January 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712002476.

Lu:2010:SSD

[LPGX10] Yuan Lu, Li-Ping Pang, Fang-Fang Guo, and Zun-Quan Xia. A superlinear space decomposition algorithm for constrained nonsmooth convex program. Journal of Computational and Applied Mathematics, 234(1):224–232, May 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709008267.

Lu:2010:ADA

[LPLX10] Yuan Lu, Li-Ping Pang, Xi-Jun Liang, and Zun-Quan Xia. An approximate decomposition algorithm for convex minimization. Journal of Computational and Applied Mathematics, 234(3):658–666, June 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710000075.

Lu:2011:QTM

[LQSL11] Linfa Lu, Xiaoyuan Qian, Xiquan Shi, and Fengshan Liu. Quading triangular meshes with certain topological constraints. Journal of Computational and Applied Mathematics, 236(5):916–923, October 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711002561.

Li:2011:EMM

[LQYL11] Xiang Li, Zhongfeng Qin, Lixing Yang, and Keping Li. Entropy maximization model for the trip distribution problem with fuzzy and random parameters. Journal of Computational and Applied Mathematics, 235(8):1906–1913, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710004954.

Lebtahi:2013:APM

[LRT13] Leila Lebtahi, Óscar Romero, and Néstor Thome. Algorithms for {K, s + 1}- potent matrix constructions. Journal of Computational and Applied Mathematics, 249(??):157–162, September 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712000301.

Li:2014:ORI

[LRZ14] Danping Li, Ximin Rong, and Hui Zhao. Optimal reinsurance-investment problem for maximizing the product of the insurer’s and the reinsurer’s utilities under a CEV model. Journal of Computational and Applied Mathematics, 255(??):671–683, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003312.

Luciano:2010:MTC

[LS10] Elisa Luciano and Patrizia Semeraro. Multivariate time changes for Lévy asset models: Characterization and calibration. Journal of Computational and Applied Mathematics, 233(8):1937–1953, February 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709006530.

Li:2011:PST

[LS11] Jianli Li and Jianhua Shen.

Positive solutions for threepoint boundary value problems for second-order impulsive differential equations on infinite intervals. Journal of Computational and Applied Mathematics, 235(8):2372–2379, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006035.

Linders:2014:FRM

[LS14] Daniël Linders and Wim Schoutens. A framework for robust measurement of implied correlation. Journal of Computational and Applied Mathematics, 271(??):39– 52, December 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042714001757.

Lakestani:2011:NSW

[LSD11] Mehrdad Lakestani, Behzad Nemati Saray, and Mehdi Dehghan. Numerical solution for the weakly singular Fredholm integro-differential equations using Legendre multiwavelets. Journal of Computational and Applied Mathematics, 235(11):3291– 3303, April 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711000549.

Li:2010:SPI

[LsGHZ10] Fan-Liang Li, Li sha Gong, Xi-Yan Hu, and Lei Zhang. Successive projection iterative method for solving matrix equation AX = B. Journal of Computational and Applied Mathematics, 234(8):2405–2410, August 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710001494.

Liang:2011:SOC

[LSW11] Dong Liang, Guanying Sun, and Wenqia Wang. Secondorder characteristic schemes in time and age for a nonlinear age-structured population model. Journal of Computational and Applied Mathematics, 235(13):3841– 3858, May 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711000422.

Li:2011:SAD

[LSXZ11] Tao Li, Aiguo Song, Mingxiang Xue, and Haitao Zhang. Stability analysis on delayed neural networks based on an improved delay-partitioning approach. Journal of Computational and Applied Mathematics, 235(9):3086– 3095, March 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710005534.

Li:2011:SQC

[LT11] Dong-Hui Li and Bo-Shi Tian. n-step quadratic convergence of the MPRP method with a restart strategy. Journal of Computational and Applied Mathematics, 235(17):4978–4990, July 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711002202.

Loffeld:2013:CPE

[LT13] J. Loffeld and M. Tokman. Comparative performance of exponential, implicit, and explicit integrators for stiff systems of ODEs. Journal of Computational and Applied Mathematics, 241(??): 45–67, March 15, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042712004116.

Lima:2010:ANI

[LTFL10] Pedro M. Lima, M. Filomena Teodoro, Neville J. Ford, and Patricia M. Lumb. Analytical and numerical investigation of mixed-type functional differential equations. Journal of Computational and Applied Mathematics, 234 (9):2826–2837, September 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710000324.

Li:2010:EBR

[LTW10] Guoyin Li, Chunming Tang, and Zengxin Wei. Error bound results for generalized D-gap functions of nonsmooth variational inequality problems. Journal of Computational and Applied Mathematics, 233(11):2795– 2806, April 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709007663.

Layton:2012:LTS

[LTX12] William Layton, Hoang Tran, and Xin Xiong. Long time stability of four methods for splitting the evolutionary Stokes–Darcy problem into Stokes and Darcy subproblems. Journal of Computational and Applied Mathematics, 236(13):3198– 3217, July 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712000763.

Liu:2010:ILB

[LTZ10] Xinwei Liu, Kim-Chuan Toh, and Gongyun Zhao. On the implementation of a log-barrier progressive hedging method for multistage stochastic programs. Journal of Computational and Applied Mathematics, 234(2):579–592, May 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710000026.

Lu:2011:SBP

[Lu11] Lizheng Lu. Sample-based polynomial approximation of rational Bézier curves. Journal of Computational and Applied Mathematics, 235(6):1557–1563, January 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710004462.

Lu:2013:ECD

[Lu13a]  Lizheng Lu.  Explicit

G^{2}-constrained degree reduction of Bézier curves by quadratic optimization. Journal of Computational and Applied Mathematics, 253(??):80–88, December 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713001842.

Lu:2013:NQP

[Lu13b] Lizheng Lu. A note on quintic polynomial approximation of generalized Cornu spiral segments. Journal of Computational and Applied Mathematics, 253(??):123– 130, December 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713001878.

Lu:2014:EMM

[Lu14] Lizheng Lu. An explicit method for G^{3} merging of two Bézier curves. Journal of Computational and Applied Mathematics, 260(??): 421–433, April 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713005748.

Lui:2010:CEH

[Lui10] S. H. Lui. Convergence estimates for an higher order optimized Schwarz method for domains with an arbitrary interface. Journal of Computational and Applied Mathematics, 235(1):301– 314, November 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710003511.

Luo:2010:FPE

[Luo10] Jiaowan Luo. Fixed points and exponential stability for stochastic Volterra–Levin equations. Journal of Computational and Applied Mathematics, 234(3):934– 940, June 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710000841.

Lampe:2013:LST

[LV13] Jörg Lampe and Heinrich Voss. Large-scale Tikhonov regularization of total least squares. Journal of Computational and Applied Mathematics, 238(??):95–108, January 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003457.

Le:2014:COM

[LV14] Thanh Hieu Le and Marc Van Barel. A convex optimization method to solve a filter design problem. Journal of Computational and Applied Mathematics, 255(??):183–192, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713002501.

Licea:2013:ANS

[LVCC13] J. A. Licea, L. Villafuerte, and B. M. Chen-Charpentier. Analytic and numerical solutions of a Riccati differential equation with random coefficients. Journal of Computational and Applied Mathematics, 239(??):208–219, February 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271200413X.

Leger:2014:PNC

[LVD^{+}14] Raphaël Léger, Jonathan Viquerat, Clément Durochat, Claire Scheid, and Stéphane Lanteri. A parallel nonconforming multi-element DGTD method for the simulation of electromagnetic wave interaction with metallic nanoparticles. Journal of Computational and Applied Mathematics, 270 (??):330–342, November 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042713007152.

Liu:2013:CFS

[LvDtTB^{+}13] L. Liu, J. van Dijk, J. H. M.

ten Thije Boonkkamp, D. B. Mihailova, and J. J. A. M. van der Mullen. The complete flux scheme — error analysis and application to plasma simulation. Journal of Computational and Applied Mathematics, 250(??):229–243, October 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713001192.

Lewanowicz:2010:TVO

[LW10a] Stanisław Lewanowicz and Pawel Woźny. Twovariable orthogonal polynomials of big q-Jacobi type. Journal of Computational and Applied Mathematics, 233(6):1554–1561, January 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709001617.

Liu:2010:DOL

[LW10b] Xueyong Liu and Liping Wen. Dissipativity of oneleg methods for neutral delay integro-differential equations. Journal of Computational and Applied Mathematics, 235(1):165–173, November 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710003067.

Liu:2011:MHO

[LW11] Xi-Lan Liu and Xiao-Rui Wang. Modifications of higher-order convergence for solving nonlinear equations. Journal of Computational and Applied Mathematics, 235(17):5105–5111, July 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711002330.

Li:2013:EER

[LW13] X. Y. Li and B. Y. Wu. Error estimation for the reproducing kernel method to solve linear boundary value problems. Journal of Computational and Applied Mathematics, 243(??):10– 15, May 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712004840.

Lee:2014:AET

[LW14] K. F. Lee and R. Wong.

Asymptotic expansion of the Tricomi–Carlitz polynomials and their zeros. Journal of Computational and Applied Mathematics, 265(??):220–242, August 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005025.

Li:2013:SLS

[LwCLW13] Tiexiang Li, Eric King wah Chu, Wen-Wei Lin, and Peter Chang-Yi Weng. Solving large-scale continuous-time algebraic Riccati equations by doubling. Journal of Computational and Applied Mathematics, 237(1):373– 383, January 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712002415.

Lee:2011:ALP

[LWH^{+}11] Wen-Chuan Lee, Jong-Wuu Wu, Mei-Ling Hong, Liang-Sian Lin, and Ruei-Ling Chan. Assessing the lifetime performance index of Rayleigh products based on the Bayesian estimation under progressive type II right censored samples. Journal of Computational and Applied Mathematics, 235(6):1676–1688, January 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710005005.

Leung:2013:COP

[LWN13] Kwai Sun Leung, Hoi Ying Wong, and Hon Yip Ng. Currency option pricing with Wishart process. Journal of Computational and Applied Mathematics, 238 (??):156–170, January 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042712003512.

Lai:2011:SPP

[LWW11] Yisheng Lai, Renhong Wang, and Jinming Wu.

Solving parametric piecewise polynomial systems. Journal of Computational and Applied Mathematics, 236(5):924–936, October 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271100255X.

Li:2013:LSI

[LWW13] Wei Li, Huping Wang, and Qin Wang. Localized solutions to interval linear equations. Journal of Computational and Applied Mathematics, 238(??):29–38, January 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271200338X.

Li:2012:RSK

[LWXW12] Jianbing Li, Xuesong Wang, Shunping Xiao, and Tao Wang. A rapid solution of a kind of 1D Fredholm oscillatory integral equation. Journal of Computational and Applied Mathematics, 236(10):2696– 2705, April 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712000088.

Liu:2010:SNC

[LWY10] Dongjie Liu, Jiming Wu, and Dehao Yu. The superconvergence of the Newton– Cotes rule for Cauchy principal value integrals. Journal of Computational and Applied Mathematics, 235(3):696–707, December 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710003754.

Lin:2013:TWS

[LWZ13] Yuan Lin, Qi-Ru Wang, and Kai Zhou. Traveling wave solutions in n-dimensional delayed reaction–diffusion systems with mixed monotonicity. Journal of Computational and Applied Mathematics, 243(??):16– 27, May 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712005006.

Li:2014:CPD

[LWZ^{+}14] Chaoqian Li, Feng Wang, Jianxing Zhao, Yan Zhu, and Yaotang Li. Criterions for the positive definiteness of real supersymmetric tensors. Journal of Computational and Applied Mathematics, 255(??):1–14, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713002082.

Lang:2012:IQS

[LX12] Feng-Gong Lang and Xiao-Ping Xu. On integro quartic spline interpolation. Journal of Computational and Applied Mathematics, 236(17):4214–4226, November 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712002270.

Luo:2011:RFV

[LXSC11] Zhendong Luo, Zhenghui Xie, Yueqiang Shang, and Jing Chen. A reduced finite volume element formulation and numerical simulations based on POD for parabolic problems. Journal of Computational and Applied Mathematics, 235(8):2098–2111, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710005649.

Li:2014:RFS

[LXX14] Ming Li, Xiang-Xiang Xi, and Xiang-Tuan Xiong. Regularization for a fractional sideways heat equation. Journal of Computational and Applied Mathematics, 255(??):28–43, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713002215.

Li:2013:SPC

[LXY13] Haibo Li, Lili Xiao, and Jun Ye. Strong predictor– corrector Euler–Maruyama methods for stochastic differential equations with Markovian switching. Journal of Computational and Applied Mathematics, 237 (1):5–17, January 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042712002919.

Lin:2010:FFM

[LY10] Youjiang Lin and Yongjian Yang. Filled function method for nonlinear equations. Journal of Computational and Applied Mathematics, 234(3):695–702, June 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710000117.

Lin:2013:PEE

[LY13a] Tao Lin and Xiu Ye. A posteriori error estimates for finite volume method based on bilinear trial functions for the elliptic equation. Journal of Computational and Applied Mathematics, 254(??):185–191, December 15, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713001143.

Lotfi:2013:NTS

[LY13b] A. Lotfi and S. A. Yousefi.

A numerical technique for solving a class of fractional variational problems. Journal of Computational and Applied Mathematics, 237(1):633–643, January 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003275.

Lotfi:2013:NSC

[LYD13] A. Lotfi, S. A. Yousefi, and Mehdi Dehghan. Numerical solution of a class of fractional optimal control problems via the Legendre orthonormal basis combined with the operational matrix and the Gauss quadrature rule. Journal of Computational and Applied Mathematics, 250(??):143–160, October 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713001106.

Liu:2010:HSR

[LYP10] Yongjian Liu, Qigui Yang, and Guoping Pang. A hyperchaotic system from the rabinovich system. Journal of Computational and Applied Mathematics, 234(1): 101–113, May 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042709008164.

Lv:2013:TTP

[LYP13] Yunfei Lv, Rong Yuan, and Yongzhen Pei. Two types of predator–prey models with harvesting: Nonsmooth and non-continuous. Journal of Computational and Applied Mathematics, 250(??):122–142, October 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713001076.

Liu:2011:SNS

[LYZ11] Dezhi Liu, Guiyuan Yang, and Wei Zhang. The stability of neutral stochastic delay differential equations with Poisson jumps by fixed points. Journal of Computational and Applied Mathematics, 235(10):3115–3120, March 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704270800561X.

Liu:2010:CAM

[LZ10] Qunfeng Liu and Jinping Zeng. Convergence analysis of multigrid methods with residual scaling techniques. Journal of Computational and Applied Mathematics, 234(10):2932–2942, September 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710002013.

Li:2011:LBD

[LZ11] Min Li and Weijun Zhong.

An LQP-based descent method for structured monotone variational inequalities. Journal of Computational and Applied Mathematics, 235(5):1523–1530, January 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710004887.

Lin:2012:CAL

[LZ12a] Runchang Lin and Zhimin Zhang. Convergence analysis for least-squares finite element approximations of second-order twopoint boundary value problems. Journal of Computational and Applied Mathematics, 236(17):4436–4447, November 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712001768.

Lin:2012:LBI

[LZ12b] Tao Lin and Xu Zhang. Linear and bilinear immersed finite elements for planar elasticity interface problems. Journal of Computational and Applied Mathematics, 236(18):4681–4699, December 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712001318.

Li:2013:RIQ

[LZ13a] Hua Li and Hua Zou. A random integral quadrature method for numerical analysis of the second kind of Volterra integral equations. Journal of Computational and Applied Mathematics, 237(1):35–42, January 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712002907.

Li:2013:ICS

[LZ13b] Xin Li and Jianmin Zheng.

Interproximate curve subdivision. Journal of Computational and Applied Mathematics, 244(??):36– 48, May 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712005055.

Liu:2013:CRM

[LZ13c] Chunxiao Liu and Shengfeng Zhu. A convex relaxation method for computing exact global solutions for multiplicative noise removal. Journal of Computational and Applied Mathematics, 238(??):144–155, January 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271200341X.

Liu:2013:IHG

[LZ13d] Qiaohua Liu and Fudao Zhang. Incomplete hyperbolic Gram–Schmidtbased preconditioners for the solution of large indefinite least squares problems. Journal of Computational and Applied Mathematics, 250(??):210–216, October 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713000861.

Liu:2013:LMO

[LZ13e] R. H. Liu and J. L. Zhao. A lattice method for option pricing with two underlying assets in the regimeswitching model. Journal of Computational and Applied Mathematics, 250(??):96– 106, October 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042713000812.

Liu:2014:DDS

[LZ14a] Jiankang Liu and Zhoushun Zheng. A dimension by dimension splitting immersed interface method for heat conduction equation with interfaces. Journal of Computational and Applied Mathematics, 261(??):221– 231, May 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713006158.

Liu:2014:NMS

[LZ14b] Panpan Liu and Shugong Zhang. Newton’s method for solving a class of nonlinear matrix equations. Journal of Computational and Applied Mathematics, 256(??):254–267, January 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003828.

Li:2014:OSC

[LZDW14] Can Li, Tinggang Zhao, Weihua Deng, and Yujiang Wu. Orthogonal spline collocation methods for the subdiffusion equation. Journal of Computational and Applied Mathematics, 255(??):517–528, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003063.

Li:2014:RNN

[LZH14] Yu-Fan Li, Yan-Jiao Zhang, and Zheng-Hai Huang. A reweighted nuclear norm minimization algorithm for low rank matrix recovery. Journal of Computational and Applied Mathematics, 263(??):338–350, June 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713006754.

Liu:2010:ACA

[LZI10] Yongjun Liu, Guanglu Zhou, and Nur Fadhilah Ibrahim. An always convergent algorithm for the largest eigenvalue of an irreducible nonnegative tensor. Journal of Computational and Applied Mathematics, 235(1):286–292, November 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710003468.

Li:2011:PPG

[LZL11a] Zheng Li, Tie Zhang, and Chang-Jun Li. Parameterized preconditioning for generalized saddle point problems arising from the Stokes equation. Journal of Computational and Applied Mathematics, 236(6):1511–1520, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004948.

Liu:2011:SMP

[LZL11b] Yong-Chao Liu, Jin Zhang, and Gui-Hua Lin. Stochastic mathematical programs with hybrid equilibrium constraints. Journal of Computational and Applied Mathematics, 235(13):3870– 3882, May 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711000446.

Luo:2011:MCS

[LZSL11] Xiaonan Luo, Zhongming Zhao, Zhuo Su, and Yun Liang. Multiple-cue saliency measurement and optimized image composition for image retargeting. Journal of Computational and Applied Mathematics, 236(5):704– 713, October 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704271100358X.

Lan:2013:YCA

[LZT13] Yanfei Lan, Ruiqing Zhao, and Wansheng Tang. A yardstick competition approach to a multi-firm regulation problem under asymmetric information. Journal of Computational and Applied Mathematics, 249(??):24– 36, September 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713000563.

Luo:2011:MAE

[LZWJ11] Changtong Luo, Shao-Liang Zhang, Chun Wang, and Zonglin Jiang. A metamodel-assisted evolutionary algorithm for expensive optimization. Journal of Computational and Applied Mathematics, 236(5): 759–764, October 1, 2011.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042711003190.

Li:2010:SUN

[LZY10] Jin Li, Xiaoping Zhang, and Dehao Yu. Superconvergence and ultraconvergence of Newton–Cotes rules for supersingular integrals. Journal of Computational and Applied Mathematics, 233(11):2841–2854, April 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709007705.

Ma:2010:ESP

[Ma10] Qing-Hua Ma. Estimates on some power nonlinear Volterra–Fredholm type discrete inequalities and their applications. Journal of Computational and Applied Mathematics, 233(9):2170– 2180, March 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042709006712.

Moncayo:2011:EBC

[MA11] María Moncayo and Sergio Amat. Error bounds for a class of subdivision schemes based on the two-scale refinement equation. Journal of Computational and Applied Mathematics, 236(2):265–278, August 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711003670.

Maleknejad:2013:SAN

[MA13] K. Maleknejad and M. Alizadeh. Sinc approximation for numerical solution of integral equation arising in conductor-like screening model for real solvent. Journal of Computational and Applied Mathematics, 251(??):81–92, October 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713001702.

Maleknejad:2014:UIP

[MA14a] K. Maleknejad and S. Akhavan. Using the improved Petrov–Galerkin elements k − 0 for solving nonlinear Hammerstein–Fredholm integral equations. Journal of Computational and Applied Mathematics, 267(??):107– 116, September 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042714000569.

Miller:2014:RSS

[MA14b] S. T. Miller and R. Abedi. Riemann solutions for spacetime discontinuous Galerkin methods. Journal of Computational and Applied Mathematics, 270(??):510– 521, November 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042713006596.

Mahmudov:2011:SMO

[Mah11] N. I. Mahmudov. q-Szász– Mirakjan operators which preserve χ^{2}. Journal of Computational and Applied Mathematics, 235(16):4621– 4628, June 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042710002207.

Mahmoudi:2014:NMA

[Mah14] Y. Mahmoudi. A new modified Adomian decomposition method for solving a class of hypersingular integral equations of second kind. Journal of Computational and Applied Mathematics, 255(??):737–742, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713003245.

Mach:2014:EAB

[MAK^{+}14] F. Mach, L. Adam, J. Kacerovský, P. Karban, and I. Dolezel. Evolutionary algorithmbased multi-criteria optimization of triboelectrostatic separator. Journal of Computational and Applied Mathematics, 270(??):134– 142, November 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042714000855.

Manguoglu:2011:DDP

[Man11]  Murat Manguoğlu.  A

domain-decomposing parallel sparse linear system solver. Journal of Computational and Applied Mathematics, 236(3):319–325, September 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004122.

Marcozzi:2011:AED

[Mar11] Michael D. Marcozzi. An adaptive extrapolation discontinuous Galerkin method for the valuation of Asian options. Journal of Computational and Applied Mathematics, 235(12):3632–3645, April 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711001129.

Matsuo:2010:HCG

[Mat10]  Takayasu Matsuo.  A

Hamiltonian-conserving Galerkin scheme for the Camassa– Holm equation. Journal of Computational and Applied Mathematics, 234(4):1258– 1266, June 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042709006426.

Maurel:2010:RTS

[Mau10]  Agnès Maurel.  Reflec-

tion and transmission by a slab with randomly distributed isotropic point scatterers. Journal of Computational and Applied Mathematics, 234(6):1842– 1850, July 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042709005111.

Majumdar:2011:GAA

[MB11] J. Majumdar and A. K. Bhunia. Genetic algorithm for asymmetric traveling salesman problem with imprecise travel times. Journal of Computational and Applied Mathematics, 235(9):3063–3078, March 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006850.

Mathur:2010:MPR

[MBS10] P. Mathur, J. S. Brauchart, and E. B. Saff. Menke points on the real line and their connection to classical orthogonal polynomials. Journal of Computational and Applied Mathematics, 233(6):1416–1431, January 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709001447.

McNally:2010:FAS

[McN10] Jeffrey Mark McNally. A fast algorithm for solving diagonally dominant symmetric pentadiagonal Toeplitz systems. Journal of Computational and Applied Mathematics, 234(4):995–1005, June 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704270900209X.

Mazzia:2012:TSS

[MCS12] Francesca Mazzia, Jeff R.

Cash, and Karline Soetaert. A test set for stiff initial value problem solvers in the open source software R: Package deTestSet.

Journal of Computational and Applied Mathematics, 236(16):4119–4131, October 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712001331.

Martinez:2010:IMB

[MD10a] Ana Gabriela Martínez and Alvaro Rodolfo De Pierro. Iterative methods based on spline approximations to detect discontinuities from Fourier data. Journal of Computational and Applied Mathematics, 234(12):3268–3282, October 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710002359.

Mirzaei:2010:MLP

[MD10b] Davoud Mirzaei and Mehdi Dehghan. Meshless local Petrov–Galerkin (MLPG) approximation to the two dimensional sine-Gordon equation. Journal of Computational and Applied Mathematics, 233(10):2737– 2754, March 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709007638.

Mitra:2010:RSV

[MD10c] Sovan Mitra and Paresh Date. Regime switching volatility calibration by the Baum–Welch method. Journal of Computational and Applied Mathematics, 234(12):3243–3260, October 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710002190.

Manceur:2013:MLE

[MD13] Ameur M. Manceur and Pierre Dutilleul. Maximum likelihood estimation for the tensor normal distribution: Algorithm, minimum sample size, and empirical bias and dispersion. Journal of Computational and Applied Mathematics, 239(??):37– 49, February 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712003810.

Moguen:2013:PVC

[MDVB13] Yann Moguen, Erik Dick, Jan Vierendeels, and Pascal Bruel. Pressurevelocity coupling for unsteady low Mach number flow simulations: an improvement of the AUSM +-up scheme. Journal of Computational and Applied Mathematics, 246(??):136– 143, July 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712004694.

Meeker:2012:ICS

[Mee12] David C. Meeker. An improved continuum skin and proximity effect model for hexagonally packed wires. Journal of Computational and Applied Mathematics, 236(18):4635–4644, December 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712001690.

Melenk:2010:CFQ

[Mel10] J. M. Melenk. On the convergence of Filon quadrature. Journal of Computational and Applied Mathematics, 234(6):1692–1701, July 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709004920.

Meszaros:2012:RTI

[Més12] Csaba Mészáros. Regularization techniques in interior point methods. Journal of Computational and Applied Mathematics, 236(15): 3704–3709, September 2012.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042711003980.

Matsuo:2014:SML

[MF14] Takayasu Matsuo and Daisuke Furihata. A stabilization of multistep linearly implicit schemes for dissipative systems. Journal of Computational and Applied Mathematics, 264(??):38– 48, July 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713007012.

Morgado:2013:ANM

[MFL13] M. L. Morgado, N. J. Ford, and P. M. Lima. Analysis and numerical methods for fractional differential equations with delay. Journal of Computational and Applied Mathematics, 252(??):159– 168, November 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712002762.

Martinez-Gonzalez:2010:HOH

[MGZ10] P. Martínez-González and A. Zarzo. Higher order hypergeometric Lauricella function and zero asymptotics of orthogonal polynomials. Journal of Computational and Applied Mathematics, 233(6):1577–1583, January 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709001642.

Motlagh:2014:DAS

[MHH^{+}14] Omid Motlagh, Tang Sai Hong, Sayed Mahdi Homayouni, George Grozev, and Elpiniki I. Papageorgiou. Development of applicationspecific adjacency models using fuzzy cognitive map. Journal of Computational and Applied Mathematics, 270(??):178–187, November 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271400079X.

Mahmood:2011:CHO

[MHL11] Mohammed Shuker Mahmood, Milan Hokr, and Michal Lukac. Combined higher order finite volume and finite element scheme for double porosity and nonlinear adsorption of transport problem in porous media. Journal of Computational and Applied Mathematics, 235(14):4221–4236, May 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711001464.

Miana:2010:STD

[Mia10] Pedro J. Miana. Stieltjes transforms defined by C_{0}-semigroups. Journal of Computational and Applied Mathematics, 234(4):1172– 1178, June 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709004099.

Mincsovics:2010:DCM

[Min10] M. E. Mincsovics. Discrete and continuous maximum principles for parabolic and elliptic operators. Journal of Computational and Applied Mathematics, 235 (2):470–477, November 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042710003407.

Mirzal:2014:CAO

[Mir14] Andri Mirzal. A convergent algorithm for orthogonal nonnegative matrix factorization. Journal of Computational and Applied Mathematics, 260(??):149– 166, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713004779.

Miyajima:2010:FEA

[Miy10a] Shinya Miyajima. Fast enclosure for all eigenvalues in generalized eigenvalue problems. Journal of Computational and Applied Mathematics, 233(11):2994–3004, April 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704270900795X.

Miyajima:2010:FES

[Miy10b]  Shinya Miyajima.  Fast

enclosure for solutions in underdetermined systems. Journal of Computational and Applied Mathematics, 234(12):3436–3444, October 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710002608.

Miyajima:2012:NEE

[Miy12] Shinya Miyajima. Numerical enclosure for each eigenvalue in generalized eigenvalue problem. Journal of Computational and Applied Mathematics, 236(9):2545– 2552, March 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711006273.

Masjed-Jamei:2010:BCS

[MJ10a] Mohammad Masjed-Jamei.

A basic class of symmetric orthogonal functions with six free parameters. Journal of Computational and Applied Mathematics, 234(1): 283–296, May 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709008425.

Masjed-Jamei:2010:CNE

[MJ10b] Mohammad Masjed-Jamei.

On constructing new expansions of functions using linear operators. Journal of Computational and Applied Mathematics, 234(2):365– 374, May 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709008474.

Milosevic:2011:ATS

[MJ11] Marija Milosević and Miljana Jovanović. An application of Taylor series in the approximation of solutions to stochastic differential equations with timedependent delay. Journal of Computational and Applied Mathematics, 235(15):4439– 4451, June 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711001865.

Mu:2014:PEA

[MJ14] Lin Mu and Rabeea Jari. A posteriori error analysis for discontinuous finite volume methods of elliptic interface problems. Journal of Computational and Applied Mathematics, 255(??):529– 543, January 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042713002951.

Mokris:2014:CHT

[MJG14] Dominik Mokriš, Bert Jüttler, and Carlotta Giannelli. On the completeness of hierarchical tensor-product B-splines. Journal of Computational and Applied Mathematics, 271(??):53–70, December 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042714001824.

Matsuyama:2010:IOR

[MK10] Hisashi Matsuyama and Toshikatsu Koga. Inner and outer radial density functions in singly-excited 1 snl states of the He atom. Journal of Computational and Applied Mathematics, 233(6):1584–1589, January 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709001654.

Murai:2011:SCS

[MK11] Daisuke Murai and Toshiyuki Koto. Stability and convergence of staggered Runge– Kutta schemes for semilinear wave equations. Journal of Computational and Applied Mathematics, 235(14): 4251–4264, May 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042711001506.

Mach:2012:IHC

[MKD12] Frantisek Mach, Pavel Karban, and Ivo Dolezel. Induction heating of cylindrical nonmagnetic ingots by rotation in static magnetic field generated by permanent magnets. Journal of Computational and Applied Mathematics, 236(18):4732– 4744, December 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712001021.

Mazova:2014:NST

[MKK14] R. Kh. Mazova, B. A. Kiselman, and E. A. Kolchina. Numerical simulation of tsunami wave height distribution for Turkish Black Sea coast in nonlinear dynamic keyboard model of underwater seismic source. Journal of Computational and Applied Mathematics, 259 (part B)(??):887–896, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713004469.

Ma:2012:SSN

[MKS12] Zhonghua Ma, Lukas Korous, and Erick Santiago. Solving a suite of NIST benchmark problems for adaptive FEM with the Hermes library. Journal of Computational and Applied Mathematics, 236(18):4846– 4861, December 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712000611.

Morgado:2010:NSC

[ML10] L. Morgado and P. Lima. Numerical solution of a class of singular free boundary problems involving the m-Laplace operator. Journal of Computational and Applied Mathematics, 234(9):2838–2847, September 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710000348.

Meng:2013:CCF

[ML13] Zhaoliang Meng and Zhongxuan Luo. Constructing cubature formulae of degree 5 with few points. Journal of Computational and Applied Mathematics, 237(1):354– 362, January 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712002397.

Meng:2014:SMH

[ML14] Zhaoliang Meng and Zhongxuan Luo. On the singularity of multivariate Hermite interpolation. Journal of Computational and Applied Mathematics, 261(??):85– 94, May 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271300602X.

Ma:2014:CAM

[MLZ14] Jingtang Ma, Jinqiang Liu, and Zhiqiang Zhou. Convergence analysis of moving finite element methods for space fractional differential equations. Journal of Computational and Applied Mathematics, 255(??):661– 670, January 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713003191.

Marinov:2010:CIE

[MM10] Tchavdar T. Marinov and Rossitza S. Marinova. Coefficient identification in Euler–Bernoulli equation from over-posed data. Journal of Computational and Applied Mathematics, 235 (2):450–459, November 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042710003389.

Michailidis:2011:PDM

[MM11a] Panagiotis D. Michailidis and Konstantinos G. Margaritis. Parallel direct methods for solving the system of linear equations with pipelining on a multicore using OpenMP. Journal of Computational and Applied Mathematics, 236(3):326–341, September 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004183.

Mohammadi:2011:SGR

[MM11b] Maryam Mohammadi and Reza Mokhtari. Solving the generalized regularized long wave equation on the basis of a reproducing kernel space. Journal of Computational and Applied Mathematics, 235(14):4003–4014, May 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711000884.

Morhac:2011:HRB

[MM11c] Miroslav Morhác and Vladislav Matousek. High-resolution boosted deconvolution of spectroscopic data. Journal of Computational and Applied Mathematics, 235 (6):1629–1640, January 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042710004966.

Mastroianni:2012:NIF

[MM12a] G. Mastroianni and G. Monegato. Numerical integration of functions with a very small significant support. Journal of Computational and Applied Mathematics, 236(16):4082–4089, October 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711006327.

Miyatake:2012:CFD

[MM12b] Yuto Miyatake and Takayasu Matsuo. Conservative finite difference schemes for the Degasperis–Procesi equation. Journal of Computational and Applied Mathematics, 236(15):3728–3740, September 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711004833.

Maleki:2010:NSP

[MMF10] Mohammad Maleki and Mahmoud Mashali-Firouzi. A numerical solution of problems in calculus of variation using direct method and nonclassical parameterization. Journal of Computational and Applied Mathematics, 234(5):1364–1373, July 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710000610.

Modarres:2011:IHA

[MML11] Farzin Modarres, Abu Hassan Malik, and Wah June Leong. Improved Hessian approximation with modified secant equations for symmetric rank-one method. Journal of Computational and Applied Mathematics, 235(8):2423–2431, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006084.

Mykhailova:2010:BIE

[MMMG10] I. I. Mykhailova, O. V. Menshykov, M. V. Menshykova, and I. A. Guz. Boundary integral equations for an interface linear crack under harmonic loading. Journal of Computational and Applied Mathematics, 234(7):2279– 2286, August 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709005640.

Marburger:2011:ABO

[MMP11] Jan Marburger, Nicole Marheineke, and René Pinnau. Adjoint-based optimal control using meshfree discretizations. Journal of Computational and Applied Mathematics, 235(10):3138– 3150, March 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710006680.

Mehrkanoon:2010:VSI

[MMS10] S. Mehrkanoon, Z. A. Majid, and M. Suleiman. A variable step implicit block multistep method for solving first-order ODEs. Journal of Computational and Applied Mathematics, 233 (9):2387–2394, March 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042709007213.

Maleknejad:2013:SNI

[MMS13] K. Maleknejad, R. Mollapourasl, and M. Shahabi. On the solution of a nonlinear integral equation on the basis of a fixed point technique and cubic B-spline scaling functions. Journal of Computational and Applied Mathematics, 239(??):346– 358, February 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712003639.

Miljkovic:2013:GMC

[MMSW13] Sladjana Miljković, Marko Miladinović, Predrag S. Stanimirović, and Yimin Wei. Gradient methods for computing the Drazin-inverse solution. Journal of Computational and Applied Mathematics, 253(??):255–263, December 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713002161.

Mehl:2010:SVL

[MMX10] Christian Mehl, Volker Mehrmann, and Hongguo Xu. Singular-value-like decomposition for complex matrix triples. Journal of Computational and Applied Mathematics, 233(5):1245– 1276, January 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709001794.

Minamoto:2010:NVM

[MN10] Teruya Minamoto and Mitsuhiro T. Nakao. A numerical verification method for a periodic solution of a delay differential equation. Journal of Computational and Applied Mathematics, 235(3):870–878, December 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710004164.

Mazzucato:2014:QOR

[MNQ14] Anna L. Mazzucato, Victor Nistor, and Qingqin Qu. Quasi-optimal rates of convergence for the generalized finite element method in polygonal domains. Journal of Computational and Applied Mathematics, 263(??): 466–477, June 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042713006997.

Mastroianni:2010:SQF

[MO10] G. Mastroianni and D. Occorsio. Some quadrature formulae with nonstandard weights. Journal of Computational and Applied Mathematics, 235(3):602–614, December 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710003559.

Mosleh:2010:EFR

[MOA10] M. Mosleh, M. Otadi, and S. Abbasbandy. Evaluation of fuzzy regression models by fuzzy neural network. Journal of Computational and Applied Mathematics, 234(3):825–834, June 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710000592.

Mar-Or:2010:HOO

[MOG10] Assaf Mar-Or and Dan Givoli. High-order oneway model nesting in dispersive non-uniform media. Journal of Computational and Applied Mathematics, 234(6):1663–1669, July 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709004889.

Mohsen:2013:ISO

[Moh13] A. Mohsen. On the integral solution of the onedimensional Bratu problem. Journal of Computational and Applied Mathematics, 251(??):61–66, October 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713001325.

Molai:2010:TNA

[Mol10] Ali Abbasi Molai. Two new algorithms for solving optimization problems with one linear objective function and finitely many constraints of fuzzy relation inequalities. Journal of Computational and Applied Mathematics, 233(8):2090– 2103, February 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042709006669.

Monkola:2010:THS

[Mön10] Sanna Mönkölä. Timeharmonic solution for acoustoelastic interaction with controllability and spectral elements. Journal of Computational and Applied Mathematics, 234(6):1904–1911, July 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709005184.

Morigi:2010:GSE

[Mor10a] S. Morigi. Geometric surface evolution with tangential contribution. Journal of Computational and Applied Mathematics, 233(5):1277– 1287, January 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042709001770.

Morimoto:2010:SP

[Mor10b] M. Morimoto. Study on the preconditioners (I + S_{m}). Journal of Computational and Applied Mathematics, 234(1):209–214, May 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709008243.

Mainar:2010:MPR

[MP10a] E. Mainar and J. M. Peña. Monotonicity preserving representations of non-polynomial surfaces. Journal of Computational and Applied Mathematics, 233(9):2161–2169, March 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042709006694.

Mengoni:2010:ICD

[MP10b] M. Mengoni and J. P. Ponthot. Isotropic continuum damage/repair model for alveolar bone remodeling. Journal of Computational and Applied Mathematics, 234(7):2036–2045, August 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704270900538X.

Morini:2010:RNM

[MPC10] Benedetta Morini, Margherita Porcelli, and Raymond H. Chan. A reduced Newton method for constrained linear least-squares problems. Journal of Computational and Applied Mathematics, 233(9):2200–2212, March 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709006840.

Marquezino:2010:OQF

[MPS10] F. L. Marquezino, R. Portugal, and F. D. Sasse. Obtaining the quantum Fourier transform from the classical FFT with QR decomposition. Journal of Computational and Applied Mathematics, 235(1):74–81, November 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710002888.

Manni:2011:IAA

[MPS11] Carla Manni, Francesca Pelosi, and M. Lucia Sampoli. Isogeometric analysis in advection–diffusion problems: Tension splines approximation. Journal of Computational and Applied Mathematics, 236(4):511– 528, September 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704271100286X.

Mikhailenko:2010:MMS

[MR10] B. Mikhailenko and G. Reshetova. Mathematical modeling of seismic and acousto-gravitational waves in a heterogeneous earth-atmosphere model. Journal of Computational and Applied Mathematics, 234(6):1678–1684, July 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042709004907.

Monteiro:2011:CRS

[MR11] M. Teresa T. Monteiro and Helena Sofia Rodrigues. Combining the regularization strategy and the SQP to solve MPCC — a MATLAB implementation. Journal of Computational and Applied Mathematics, 235(18):5348–5356, July 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710002748.

Miana:2012:POI

[MR12] Pedro J. Miana and Natalia Romero. Poincaré and Opial inequalities for vector-valued convolution products. Journal of Computational and Applied Mathematics, 236(15):3720– 3727, September 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711004675.

Mustapha:2013:PPD

[MR13] Kassem Mustapha and Jennifer K. Ryan. Postprocessing discontinuous Galerkin solutions to Volterra integro-differential equations: Analysis and simulations. Journal of Computational and Applied Mathematics, 253(??):89–103, December 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271300174X.

Martinez:2010:OCP

[MRAM10] S. Martínez, M. Rueda, A. Arcos, and H. Martínez. Optimum calibration points estimating distribution functions. Journal of Computational and Applied Mathematics, 233(9):2265–2277, March 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704270900689X.

Makarov:2013:ECF

[MRD13] V. L. Makarov, N. O. Rossokhata, and D. V. Dragunov. An exponentially convergent functionaldiscrete method for solving Sturm–Liouville problems with a potential including the Dirac δ-function. Journal of Computational and Applied Mathematics, 250(??):39–57, October 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271300085X.

Morigi:2010:IPM

[MRS10] S. Morigi, L. Reichel, and F. Sgallari. An interiorpoint method for large constrained discrete ill-posed problems. Journal of Computational and Applied Mathematics, 233(5):1288– 1297, January 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709001800.

Marrero:2013:NSA

[MRT13a] J. Abderramán Marrero, M. Rachidi, and V. Tomeo. Non-symbolic algorithms for the inversion of tridiagonal matrices. Journal of Computational and Applied Mathematics, 252(??):3–11, November 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271200204X.

Marrero:2013:NAI

[MRT13b] J. Abderramán Marrero, M. Rachidi, and V. Tomeo. On new algorithms for inverting Hessenberg matrices. Journal of Computational and Applied Mathematics, 252(??):12– 20, November 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712004967.

Ma:2011:PGD

[MS11a] Keying Ma and Tongjun Sun. Parallel Galerkin domain decomposition procedures based on the streamline diffusion method for convection–diffusion problems. Journal of Computational and Applied Mathematics, 235(15):4464–4479, June 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711001920.

Merrien:2011:GTF

[MS11b] Jean-Louis Merrien and Tomas Sauer. A generalized Taylor factorization for Hermite subdivision schemes. Journal of Computational and Applied Mathematics, 236(4):565–574, September 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711003347.

Mohiuddine:2011:SPQ

[MS11c] S. A. Mohiuddine and H. Sevli. Stability of Pexiderized quadratic functional equation in intuitionistic fuzzy normed space. Journal of Computational and Applied Mathematics, 235(8):2137–2146, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710005662.

Mazzia:2012:QFD

[MS12] Francesca Mazzia and Alessandra Sestini. Quadrature formulas descending from BS Hermite spline quasiinterpolation. Journal of Computational and Applied Mathematics, 236(16):4105– 4118, October 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712001343.

Mao:2013:SCS

[MS13a] Xuerong Mao and Lukasz Szpruch. Strong convergence and stability of implicit numerical methods for stochastic differential equations with non-globally Lipschitz continuous coefficients. Journal of Computational and Applied Mathematics, 238(??):14–28, January 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003378.

Monegato:2013:PES

[MS13b] G. Monegato and L. Scuderi. Potential evaluation in space-time BIE formulations of 2D wave equation problems. Journal of Computational and Applied Mathematics, 243(??):60– 79, May 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712004797.

Mozolevski:2013:NST

[MS13c] I. Mozolevski and L. Schuh. Numerical simulation of two-phase immiscible incompressible flows in heterogeneous porous media with capillary barriers. Journal of Computational and Applied Mathematics, 242(??):12– 27, April 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712004189.

Marinova:2010:TSV

[MSE10] Rossitza S. Marinova, Raymond J. Spiteri, and Eddy Essien. Time stepping for vectorial operator splitting. Journal of Computational and Applied Mathematics, 235(2):460–469, November 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710003390.

Mao:2011:ASE

[MSG11] Xuerong Mao, Yi Shen, and Alison Gray. Almost sure exponential stability of backward Euler–Maruyama discretizations for hybrid stochastic differential equations. Journal of Computational and Applied Mathematics, 235(5):1213–1226, January 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710004449.

Medjroubi:2012:WCH

[MSP12] W. Medjroubi, B. Stoevesandt, and J. Peinke. Wake classification of heaving airfoils using the spectral/hp element method. Journal of Computational and Applied Mathematics, 236(15):3774–3782, September 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042711005747.

Motrenko:2014:SSD

[MSW14] Anastasiya Motrenko, Vadim Strijov, and Gerhard-Wilhelm Weber. Sample size determination for logistic regression. Journal of Computational and Applied Mathematics, 255(??):743–752, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713003294.

Meek:2012:PTP

[MSWY12] D. S. Meek, T. Saito, D. J. Walton, and N. Yoshida. Planar two-point G^{1} Hermite interpolating logaesthetic spirals. Journal of Computational and Applied Mathematics, 236(17):4485– 4493, November 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042712001896.

Mohamed:2013:FVM

[MSZ13] Kamel Mohamed, Mohammed Seaid, and Mostafa Zahri. A finite volume method for scalar conservation laws with stochastic time-space dependent flux functions. Journal of Computational and Applied Mathematics, 237(1):614– 632, January 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042712003044.

Masol:2010:NAR

[MT10a] Viktoriya Masol and Jef L. Teugels. Numerical accuracy of real inversion formulas for the Laplace transform. Journal of Computational and Applied Mathematics, 233(10):2521–2533, March 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709007316.

Medvinsky:2010:SRC

[MT10b] M. Medvinsky and E. Turkel.

On surface radiation conditions for an ellipse. Journal of Computational and Applied Mathematics, 234(6): 1647–1655, July 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042709004865.

Milev:2010:NVD

[MT10c] Mariyan Milev and Aldo Tagliani. Numerical valuation of discrete double barrier options. Journal of Computational and Applied Mathematics, 233(10):2468– 2480, March 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709007274.

Marrero:2012:CRI

[MT12] J. Abderramán Marrero and V. Tomeo. On the closed representation for the inverses of Hessenberg matrices. Journal of Computational and Applied Mathematics, 236(12):2962– 2970, June 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711003943.

Milev:2013:EIS

[MT13] Mariyan Milev and Aldo Tagliani. Efficient implicit scheme with positivity preserving and smoothing properties. Journal of Computational and Applied Mathematics, 243(??):1–9, May 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712004128.

Marszalek:2014:MMO

[MT14a] Wiesław Marszalek and Zdzisław Trzaska. Mixedmode oscillations and chaotic solutions of jerk (newtonian) equations. Journal of Computational and Applied Mathematics, 262(??):373– 383, May 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271300366X.

Muller:2014:CCS

[MT14b] A. Müller and Z. Terze. On the choice of configuration space for numerical Lie group integration of constrained rigid body systems. Journal of Computational and Applied Mathematics, 262(??):3–13, May 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271300592X.

Mohamed:2011:FOT

[MTJ11] Mesliza Mohamed, Bevan Thompson, and Muhammad Sufian Jusoh. Firstorder three-point boundary value problems at resonance. Journal of Computational and Applied Mathematics, 235(16):4796–4801, June 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710005959.

Markovsky:2014:SWS

[MU14] Ivan Markovsky and Konstantin Usevich. Software for weighted structured low-rank approximation. Journal of Computational and Applied Mathematics, 256(??):278–292, January 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003932.

Muite:2010:NCC

[Mui10] B. K. Muite. A numerical comparison of Chebyshev methods for solving fourth order semilinear initial boundary value problems. Journal of Computational and Applied Mathematics, 234(2):317–342, May 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709008462.

Mastronardi:2012:RAD

[MV12a] Nicola Mastronardi and Paul Van Dooren. Recursive approximation of the dominant eigenspace of an indefinite matrix. Journal of Computational and Applied Mathematics, 236(16):4090– 4104, October 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042712000994.

Mitchell:2012:AFD

[MV12b] S. L. Mitchell and M. Vynnycky. An accurate finite-difference method for ablation-type Stefan problems. Journal of Computational and Applied Mathematics, 236(17):4181–4192, November 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712002129.

Mitchell:2014:NST

[MV14] S. L. Mitchell and M. Vynnycky. On the numerical solution of two-phase Stefan problems with heat-flux boundary conditions. Journal of Computational and Applied Mathematics, 264 (??):49–64, July 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042714000077.

Matton:2010:MCE

[MVC10] Mike Matton, Dirk Van Compernolle, and Ronald Cools. Minimum classification error training in example based speech and pattern recognition using sparse weight matrices. Journal of Computational and Applied Mathematics, 234(4):1303–1311, June 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709007377.

Moreno:2012:ANS

[MVMG12] J. M. Díaz Moreno, C. García Vázquez, M. T. González Montesinos, and F. Ortegón Gallego. Analysis and numerical simulation of an inductionconduction model arising in steel heat treating. Journal of Computational and Applied Mathematics, 236(12): 3007–3015, June 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711001841.

Mach:2014:IEP

[MVV14] Thomas Mach, Marc Van Barel, and Raf Vandebril. Inverse eigenvalue problems for extended Hessenberg and extended tridiagonal matrices. Journal of Computational and Applied Mathematics, 272(??):377– 398, December 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042714001551.

Mu:2014:IPD

[MWWY14] Lin Mu, Junping Wang, Yanqiu Wang, and Xiu Ye. Interior penalty discontinuous Galerkin method on very general polygonal and polyhedral meshes. Journal of Computational and Applied Mathematics, 255(??):432–440, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713002999.

Ma:2010:ECV

[MX10] JunMei Ma and Chenglong Xu. An efficient control variate method for pricing variance derivatives. Journal of Computational and Applied Mathematics, 235(1):108–119, November 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710002931.

Ma:2011:EIH

[MY11] Dexiang Ma and Xiaozhong Yang. On eigenvalue intervals of higher-order boundary value problems with a sign-changing nonlinear term. Journal of Computational and Applied Mathematics, 235(8):2314–2324, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710005960.

Muroi:2013:SBT

[MY13] Yoshifumi Muroi and Takashi Yamada. Spectral binomial tree: New algorithms for pricing barrier options. Journal of Computational and Applied Mathematics, 249(??):107–119, September 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713000824.

Mashreghi:2014:RAO

[MY14] Ali Mashreghi and Hadi Sadoghi Yazdi. A recursive algorithm for optimizing differentiation. Journal of Computational and Applied Mathematics, 263(??):1– 13, June 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713006547.

Marcellan:2010:SID

[MYZ10] F. J. Marcellán, Rafael J. Yáñez, and Alejandro Zarzo. Special issue dedicated to Professor Jesús Sánchez–Dehesa on the occasion of his 60th birthday. Journal of Computational and Applied Mathematics, 233(6):1345–1354, January 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709001393.

Mahmudov:2014:ACF

[MZ14] N. I. Mahmudov and S. Zorlu. On the approximate controllability of fractional evolution equations with compact analytic semigroup. Journal of Computational and Applied Mathematics, 259 (part A)(??):

194–204, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042713003130.

Zhang:2011:ESN

[mZkPkL11] Qi min Zhang, Wan kai Pang, and Ping kei Leung. Exponential stability of numerical solutions for a class of stochastic age-dependent capital system with Poisson jumps. Journal of Computational and Applied Mathematics, 235(12):3369–3377, April 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006096.

Meng:2014:BML

[MZL14] Jing Meng, Pei-Yong Zhu, and Hou-Biao Li. A block GCROT(m, k) method for linear systems with multiple right-hand sides. Journal of Computational and Applied Mathematics, 255(??):544– 554, January 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713003129.

Nyeo:2011:SBL

[NA11] Su-Long Nyeo and Rafat R.

Ansari. Sparse Bayesian learning for the Laplace transform inversion in dynamic light scattering. Journal of Computational and Applied Mathematics, 235(8):2861–2872, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006588.

Naranjo:2012:SAK

[NACLR12] J. A. M. Naranjo, N. Antequera, L. G. Casado, and J. A. López-Ramos. A suite of algorithms for key distribution and authentication in centralized secure multicast environments. Journal of Computational and Applied Mathematics, 236(12): 3042–3051, June 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711000914.

Nadarajah:2011:EDP

[Nad11] Saralees Nadarajah. Exact distribution of the product of m gamma and n Pareto random variables. Journal of Computational and Applied Mathematics, 235(15):4496–4512, June 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711002123.

Nafa:2014:EOA

[Naf14] Kamel Nafa. Equal order approximations enriched with bubbles for coupled Stokes–Darcy problem. Journal of Computational and Applied Mathematics, 270(??):275–282, November 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714000247.

Naifar:2011:MDS

[Nai11]  Nader Naifar.  Mod-

elling dependence structure with Archimedean copulas and applications to the iTraxx CDS index. Journal of Computational and Applied Mathematics, 235(8):2459–2466, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006138.

Narayan:2014:ACS

[Nar14] Smita Narayan. Approximating Cornu spirals by arc splines. Journal of Computational and Applied Mathematics, 255(??):789–804, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003361.

Noor:2011:EMS

[NASNY11] Muhammad Aslam Noor, Eisa Al-Said, Khalida Inayat Noor, and Yonghong Yao. Extragradient methods for solving nonconvex variational inequalities. Journal of Computational and Applied Mathematics, 235(9):3104–3108, March 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006424.

Naumov:2011:MEP

[Nau11] Maxim Naumov. On the modification of an eigenvalue problem that preserves an eigenspace. Journal of Computational and Applied Mathematics, 235 (18):5432–5440, July 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042711003268.

Nazemi:2011:DMS

[Naz11] A. R. Nazemi. A dynamical model for solving degenerate quadratic minimax problems with constraints. Journal of Computational and Applied Mathematics, 236(6):1282–1295, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004584.

Nguyen-Ba:2010:OSS

[NBBKV10] Truong Nguyen-Ba, Vladan Bozic, Emmanuel Kengne, and Rémi Vaillancourt. A one-step 7-stage Hermite– Birkhoff–Taylor ODE solver of order 11. Journal of Computational and Applied Mathematics, 234(1):192– 208, May 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709008231.

Niu:2013:DFE

[NBZ13] Yuanling Niu, Kevin Burrage, and Chengjian Zhang. A derivative-free explicit method with order 1.0 for solving stochastic delay differential equations. Journal of Computational and Applied Mathematics, 253(??):51–65, December 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713001763.

Nogueira:2010:RCA

[NCCF^{+}10] X. Nogueira, I. Colominas, L. Cueto-Felgueroso, S. Khelladi, F. Navarrina, and M. Casteleiro. Resolution of computational aeroacoustics problems on unstructured grids with a higher-order finite volume scheme. Journal of Computational and Applied Mathematics, 234(7):2089–2097, August 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709005445.

Nieto:2010:NSP

[NdCDCFM10] P. J. García Nieto, J. J. del Coz Díaz, D. Castro-Fresno, and F. Ballester Muñoz. Numerical simulation of the performance of a snow fence with airfoil snow plates by FVM. Journal of Computational and Applied Mathematics, 234(4):1200–1210, June 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709004555.

Neilan:2014:FEM

[Nei14] Michael Neilan. Finite element methods for fully nonlinear second order PDEs based on a discrete Hessian with applications to the Monge–Ampère equation. Journal of Computational and Applied Mathematics, 263(??):351–369, June 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713007000.

Najariyan:2014:NAO

[NF14] Marzieh Najariyan and Mohamad Hadi Farahi. A new approach for the optimal fuzzy linear time invariant controlled system with fuzzy coefficients. Journal of Computational and Applied Mathematics, 259 (part B)(??):682–694, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271300215X.

NGuessan:2010:AES

[N’G10]  Assi N’Guessan.  Ana-

lytical existence of solutions to a system of nonlinear equations with application. Journal of Computational and Applied Mathematics, 234(1):297–304, May 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709008437.

Naderi:2011:SOS

[NGAZ11] B. Naderi, S. M. T. Fatemi Ghomi, M. Aminnayeri, and M. Zandieh. Scheduling open shops with parallel machines to minimize total completion time. Journal of Computational and Applied Mathematics, 235(5):1275– 1287, January 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710004516.

Nobari:2010:MAE

[NH10] Elham Nobari and S. Mohammad Hosseini. A method for approximation of the exponential map in semidirect product of matrix Lie groups and some applications. Journal of Computational and Applied Mathematics, 234(1):305– 315, May 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709008449.

Narayan:2013:GWR

[NH13] Akil C. Narayan and Jan S. Hesthaven. A generalization of the Wiener rational basis functions on infinite intervals, part II — numerical investigation. Journal of Computational and Applied Mathematics, 237(1):18–34, January 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712002786.

Noels:2010:P

[NHVD10] Ludovic Noels, Michel Hogge, Roger Van Keer, and Erik Dick. Preface. Journal of Computational and Applied Mathematics, 234(7):2027–2028, August 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709005366.

Nievergelt:2013:FCD

[Nie13a] Yves Nievergelt. Fitting cylinders to data. Journal of Computational and Applied Mathematics, 239(??):250– 269, February 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712004104.

Nievergelt:2013:EBM

[Nie13b] Yves Nievergelt. On the existence of best Mitscherlich, Verhulst, and West growth curves for generalized least-squares regression. Journal of Computational and Applied Mathematics, 248(??):31–46, August 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271300023X.

Niki:2010:CCT

[NK10] Hiroshi Niki and Toshiyuki Kohno. Comment on ‘A comparison theorem of the SOR iterative method’. Journal of Computational and Applied Mathematics, 234(12):3507–3510, October 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710002037.

See [yS05].

Niu:2010:DBK

[NL10] Qiang Niu and Linzhang Lu. Deflated block Krylov subspace methods for large scale eigenvalue problems. Journal of Computational and Applied Mathematics, 234(3):636–648, June 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270900805X.

Niu:2013:AGD

[NLL13] Qiang Niu, Lin-Zhang Lu, and Gang Liu. Accelerated GCRO–DR method for solving sequences of systems of linear equations. Journal of Computational and Applied Mathematics, 253(??):131–141, December 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713001994.

Nemati:2013:NSC

[NLO13] S. Nemati, P. M. Lima, and Y. Ordokhani. Numerical solution of a class of two-dimensional nonlinear Volterra integral equations using Legendre polynomials. Journal of Computational and Applied Mathematics, 242(??):53– 69, April 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271200458X.

Nini:2014:TDS

[NMBG14] Michele Nini, Valentina Motta, Giampiero Bindolino, and Alberto Guardone. Three-dimensional simulation of a complete vertical axis wind turbine using overlapping grids. Journal of Computational and Applied Mathematics, 270(??): 78–87, November 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704271400106X.

Naumov:2010:TBH

[NMS10] Maxim Naumov, Murat Manguoğlu, and Ahmed H.

Sameh. A tearing-based hybrid parallel sparse linear system solver. Journal of Computational and Applied Mathematics, 234 (10):3025–3038, September 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271000213X.

Noor:2013:UIM

[NNB13] Muhammad Aslam Noor, Khalida Inayat Noor, and Abdellah Bnouhachem. On a unified implicit method for variational inequalities. Journal of Computational and Applied Mathematics, 249(??):69– 73, September 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713000800.

Nishiyama:2012:CFD

[NNO12] Hirota Nishiyama, Takahiro Noi, and Shinnosuke Oharu. Conservative finite difference schemes for the generalized Zakharov–Kuznetsov equations. Journal of Computational and Applied Mathematics, 236(12):2998– 3006, June 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711001877.

Nystrom:2010:PHF

[NÖ10] Kaj Nyström and Thomas Önskog. Pricing and hedging of financial derivatives using a posteriori error estimates and adaptive methods for stochastic differential equations. Journal of Computational and Applied Mathematics, 235(3):563– 592, December 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710003535.

Nazemi:2012:CNN

[NO12] Alireza Nazemi and Farahnaz Omidi. A capable neural network model for solving the maximum flow problem. Journal of Computational and Applied Mathematics, 236(14):3498–3513, August 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712001185.

Nechepurenko:2014:ASP

[NOS14] Yu. M. Nechepurenko, G. V. Ovchinnikov, and M. Sadkane. Application of the spectral pseudoinversion to solving Hermitian systems of differentialalgebraic equations. Journal of Computational and Applied Mathematics, 260(??): 218–228, April 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713005141.

Notaris:2014:PIR

[Not14] Sotirios E. Notaris. Product integration rules for Chebyshev weight functions with Chebyshev abscissae. Journal of Computational and Applied Mathematics, 257(??):180–194, February 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713004184.

Nie:2009:ESP

[NPTH09] Linfei Nie, Jigen Peng, Zhidong Teng, and Lin Hu. Existence and stability of periodic solution of a Lotka–Volterra predator– prey model with state dependent impulsive effects. Journal of Computational and Applied Mathematics, 224(2):544–555, February 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042708002616.

See comment [TSC10].

Nguyen:2014:NEE

[NR14a] Quang Huy Nguyen and Christian Y. Robert. New efficient estimators in rare event simulation with heavy tails. Journal of Computational and Applied Mathematics, 261(??):39– 47, May 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005979.

Nowak:2014:ALP

[NR14b] Piotr Nowak and Maciej Romaniuk. Application of Levy processes and Esscher transformed martingale measures for option pricing in fuzzy framework. Journal of Computational and Applied Mathematics, 263(??):129–151, June 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713006857.

Nguyen:2013:ISH

[NRP13] D.-T. Nguyen, G. Rauchs, and J.-P. Ponthot. The impact of surface higher order differentiability in twodimensional contact elements. Journal of Computational and Applied Mathematics, 246(??):195–205, July 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712004645.

Nachtergaele:2010:EWC

[NRS10] Ph. Nachtergaele, D. J. Rixen, and A. M. Steenhoek. Efficient weakly coupled projection basis for the reduction of thermo-mechanical models. Journal of Computational and Applied Mathematics, 234(7):2272–2278, August 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709005639.

Nurnberger:2011:LLI

[NRS11] G. Nürnberger, M. Rhein, and G. Schneider. Local Lagrange interpolation by quintic C^{1} splines on type-6 tetrahedral partitions. Journal of Computational and Applied Mathematics, 236(4):529–542, September 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042711004249.

Nazari:2010:AFD

[NS10] D. Nazari and S. Shahmorad. Application of the fractional differential transform method to fractional-order integrodifferential equations with nonlocal boundary conditions. Journal of Computational and Applied Mathematics, 234(3):883–891, June 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271000066X.

Navascues:2013:NIA

[NS13] M. A. Navascués and M. V.

Sebastián. Numerical integration of affine fractal functions. Journal of Computational and Applied Mathematics, 252(??):169–176, November 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712004025.

Nurnberg:2014:UFE

[NS14] Robert Nürnberg and Andrea Sacconi. An unfitted finite element method for the numerical approximation of void electromigration. Journal of Computational and Applied Mathematics, 270(??):531–544, November 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713006559.

Natroshvili:2010:BIE

[NSZ10] David Natroshvili, Ioannis G. Stratis, and Shota Zazashvili. Boundary integral equation methods in the theory of elasticity of hemitropic materials: a brief review. Journal of Computational and Applied Mathematics, 234(6):1622– 1630, July 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S037704270900483X.

Nguyen-Thu:2014:SSP

[NTNBV14] Huong Nguyen-Thu, Truong Nguyen-Ba, and Rémi Vaillancourt. Strong-stability-preserving, Hermite–Birkhoff time-discretization based on k step methods and 8-stage explicit Runge–Kutta methods of order 5 and 4. Journal of Computational and Applied Mathematics, 263(??): 45–58, June 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713006328.

Nguyen-Thoi:2014:CAF

[NTPVNHLX14] T. Nguyen-Thoi, P. Phung-Van, S. Nguyen-Hoang, and Q. Lieu-Xuan. A coupled alpha-FEM for dynamic analyses of 2D fluidsolid interaction problems. Journal of Computational and Applied Mathematics, 271(??):130–149, December 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271400185X.

Nguyen-Thanh:2010:AAF

[NTRNXB10] N. Nguyen-Thanh, Timon Rabczuk, H. Nguyen-Xuan, and Stéphane P. A. Bordas. An alternative alpha finite element method (AαFEM) for free and forced structural vibration using triangular meshes. Journal of Computational and Applied Mathematics, 233(9):2112– 2135, March 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709006074.

Nicaise:2011:PEE

[NV11] Serge Nicaise and Juliette Venel. A posteriori error estimates for a finite element approximation of transmission problems with sign changing coefficients. Journal of Computational and Applied Mathematics, 235(14):4272–4282, May 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711001580.

Narushima:2012:CGM

[NY12] Yasushi Narushima and Hiroshi Yabe. Conjugate gradient methods based on secant conditions that generate descent search directions for unconstrained optimization. Journal of Computational and Applied Mathematics, 236(17):4303–4317, November 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712001380.

Niu:2014:IRL

[NY14] Datian Niu and Xuegang Yuan. An implicitly restarted Lanczos bidiagonalization method with refined harmonic shifts for computing smallest singular triplets. Journal of Computational and Applied Mathematics, 260(??):208– 217, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713005219.

Oliveira:2014:SEA

[OA14] Saulo P. Oliveira and Juarez S. Azevedo. Spectral element approximation of Fredholm integral eigenvalue problems. Journal of Computational and Applied Mathematics, 257(??):46– 56, February 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713004214.

Omar:2011:CMM

[OAHR11] M. A. Omar, A. Aboul-Hassan, and S. I. Rabia. The composite Milstein methods for the numerical solution of Ito stochastic differential equations. Journal of Computational and Applied Mathematics, 235(8):2277–2299, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710005923.

Ozkut:2014:MOT

[OB14] Murat Ozkut and Ismihan Bayramoğlu (Bairamov). On Marshall–Olkin type distribution with effect of shock magnitude. Journal of Computational and Applied Mathematics, 271(??): 150–162, December 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042714001836.

Oberman:2013:FDM

[Obe13] Adam M. Oberman. Finite difference methods for the infinity Laplace and p-Laplace equations. Journal of Computational and Applied Mathematics, 254(??):65–80, December 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712005237.

Ogata:2011:NTE

[OCU11] Hidenori Ogata, Fumihiro Chiba, and Teruo Ushijima. A new theoretical error estimate of the method of fundamental solutions applied to reduced wave problems in the exterior region of a disk. Journal of Computational and Applied Mathematics, 235(12):3395–3412, April 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711000537.

Odibat:2011:LPA

[Odi11] Zaid Odibat. On Legendre polynomial approximation with the VIM or HAM for numerical treatment of nonlinear fractional differential equations. Journal of Computational and Applied Mathematics, 235(9):2956– 2968, March 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710006631.

Orive:2010:CEP

[OG10] Ramón Orive and Zebenzuí García. On a class of equilibrium problems in the real axis. Journal of Computational and Applied Mathematics, 235(4):1065–1076, December 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710003031.

Ortiz-Gracia:2014:PJR

[OGM14] Luis Ortiz-Gracia and Josep J.

Masdemont. Peaks and jumps reconstruction with B-splines scaling functions. Journal of Computational and Applied Mathematics, 272(??):258–272, December 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271400260X.

Okayama:2010:SCM

[OMS10] Tomoaki Okayama, Takayasu Matsuo, and Masaaki Sugihara. Sinc-collocation methods for weakly singular Fredholm integral equations of the second kind. Journal of Computational and Applied Mathematics, 234(4):1211–1227, June 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709004567.

Ozergin:2011:EGB

[ÖÖA11] Emine Özergin, Mehmet Ali Özarslan, and Abdullah Altin. Extension of gamma, beta and hypergeometric functions. Journal of Computational and Applied Mathematics, 235(16):4601– 4610, June 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042710002165.

Ozaki:2012:FAF

[OORO12] Katsuhisa Ozaki, Takeshi Ogita, Siegfried M. Rump, and Shin’ichi Oishi. Fast algorithms for floating-point interval matrix multiplication. Journal of Computational and Applied Mathematics, 236(7):1795–1814, January 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711005449.

Ooura:2013:FCG

[Oou13] Takuya Ooura. Fast computation of Goursat’s infinite integral with very high accuracy. Journal of Computational and Applied Mathematics, 249(??):1–8, September 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713000666.

Orel:2012:CHR

[OP12] Bojan Orel and Andrej Perne. Computations with half-range Chebyshev polynomials. Journal of Computational and Applied Mathematics, 236(7):1753–1765, January 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711005371.

Orav-Puurand:2010:NTM

[OPPV10] Kerli Orav-Puurand, Arvet Pedas, and Gennadi Vainikko. Nyström type methods for Fredholm integral equations with weak singularities. Journal of Computational and Applied Mathematics, 234(9):2848–2858, September 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710000373.

Occorsio:2014:ELI

[OR14] Donatella Occorsio and Maria Grazia Russo. Extended Lagrange interpolation on the real line.

Journal of Computational and Applied Mathematics, 259 (part A)(??):24–34, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713000587.

Orang:2014:CSI

[ORP14] Ali Atashbar Orang, Seyed Esmail Razavi, and Hamoon Pourmirzaagha. Computational study of incompressible turbulent flows with method of characteristics. Journal of Computational and Applied Mathematics, 259 (part B)(??):741–751, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271300294X.

Oliveira:2010:DMA

[OS10] S. P. Oliveira and G. Seriani. DFT modal analysis of spectral element methods for the 2D elastic wave equation. Journal of Computational and Applied Mathematics, 234(6):1717–1724, July 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709004956.

Otiniano:2013:SRV

[OSR13] C. E. G. Otiniano, T. R.

Sousa, and P. N. Rathie. Stable random variables: Convolution and reliability. Journal of Computational and Applied Mathematics, 242(??):1– 11, April 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712004396.

Ozmen:2014:RRM

[ÖW14] Ayse Özmen and Gerhard Wilhelm Weber. RMARS: Robustification of multivariate adaptive regression spline under polyhedral uncertainty. Journal of Computational and Applied Mathematics, 259 (part B) (??):914–924, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042713005104.

Owadally:2012:ICF

[Owa12] Iqbal Owadally. An improved closed-form solution for the constrained minimization of the root of a quadratic functional. Journal of Computational and Applied Mathematics, 236 (17):4428–4435, November 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042712001744.

Ozarslan:2014:SFO

[ÖY14] Mehmet Ali Özarslan and Banu Yilmaz. A set of finite order differential equations for the Appell polynomials. Journal of Computational and Applied Mathematics, 259 (part A)(??):108–116, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713004111.

Perez-Arancibia:2010:GFH

[PAD10] Carlos Pérez-Arancibia and Mario Durán. On the Green’s function for the Helmholtz operator in an impedance circular cylindrical waveguide. Journal of Computational and Applied Mathematics, 235(1):244– 262, November 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042710003432.

Pagneux:2010:MAM

[Pag10]  V. Pagneux.  Multi-

modal admittance method in waveguides and singularity behavior at high frequencies. Journal of Computational and Applied Mathematics, 234(6):1834–1841, July 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704270900510X.

Portero:2010:CDD

[PAJ10] L. Portero, A. Arrarás, and J. C. Jorge. Contractivity of domain decomposition splitting methods for nonlinear parabolic problems. Journal of Computational and Applied Mathematics, 234(4):1078–1087, June 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709003069.

Paris:2010:ESE

[Par10] R. B. Paris. Exponentially small expansions in the asymptotics of the Wright function. Journal of Computational and Applied Mathematics, 234(2):488– 504, May 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709008577.

Pastore:2011:NTL

[Pas11]  Pietro Pastore.  The

numerical treatment of Love’s integral equation having very small parameter. Journal of Computational and Applied Mathematics, 236(6):1267–1281, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711004572.

Peng:2011:SGH

[PC11] Fengfu Peng and Juanjuan Chen. Spline on a generalized hyperbolic paraboloid. Journal of Computational and Applied Mathematics, 235(8):2451–2458, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710006126.

Philip:2012:AAS

[PC12a] Bobby Philip and Timothy P. Chartier. Adaptive algebraic smoothers. Journal of Computational and Applied Mathematics, 236 (9):2277–2297, March 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042711005863.

Printsypar:2012:CDP

[PC12b] G. Printsypar and R. Ciegis. On convergence of a discrete problem describing transport processes in the pressing section of a paper machine including dynamic capillary effects: Onedimensional case. Journal of Computational and Applied Mathematics, 236(14):3409– 3425, August 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712001367.

Pathak:2013:CFP

[PD13] H. K. Pathak and Deepmala. Common fixed point theorems for PDoperator pairs under relaxed conditions with applications. Journal of Computational and Applied Mathematics, 239(??):103–113, February 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271200372X.

Paulus:2013:IMU

[PDE^{+}13] Raphaël Paulus, Benjamin J.

Dewals, Sébastien Erpicum, Michel Pirotton, and Pierre Archambeau. Innovative modelling of 3D unsaturated flow in porous media by coupling independent models for vertical and lateral flows. Journal of Computational and Applied Mathematics, 246(??):38– 51, July 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712003226.

Parand:2013:SCM

[PDP13] K. Parand, Mehdi Dehghan, and A. Pirkhedri. The sinc-collocation method for solving the Thomas–Fermi equation. Journal of Computational and Applied Mathematics, 237(1):244– 252, January 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042712003238.

Peng:2011:FOD

[PFLZ11] Yehui Peng, Heying Feng, Qiyong Li, and Xiaoqing Zhang. A fourthorder derivative-free algorithm for nonlinear equations. Journal of Computational and Applied Mathematics, 235(8):2551–2559, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710006230.

Pestano-Gabino:2012:TMP

[PGGCGF12] Celina Pestano-Gabino, Concepción González-Concepción, and María Candelaria Gil-Fariña. A type of matrix Padé approximant inspired by scalar component models. Journal of Computational and Applied Mathematics, 236(13):3360– 3372, July 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712001215.

Pestano-Gabino:2014:TMI

[PGGCGF14] Celina Pestano-Gabino, Concepción González-Concepción, and María Candelaria Gil-Fariña. A tabular methodology to identify minimal row degrees for Matrix Padé Approximants. Journal of Computational and Applied Mathematics, 261(??):30– 38, May 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005566.

Pishbin:2013:SEV

[PGH13] S. Pishbin, F. Ghoreishi, and M. Hadizadeh. The semi-explicit Volterra integral algebraic equations with weakly singular kernels: the numerical treatments. Journal of Computational and Applied Mathematics, 245(??):121–132, June 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712005432.

Park:2014:EFE

[PHA14] Jang Min Park, Martien A. Hulsen, and Patrick D. Anderson. An extended finite element method for a diffuse-interface model. Journal of Computational and Applied Mathematics, 272(??):25–40, December 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042714002325.

Peyghami:2014:CIP

[PHS14] M. Reza Peyghami, S. Fathi Hafshejani, and L. Shirvani. Complexity of interiorpoint methods for linear optimization based on a new trigonometric kernel function. Journal of Computational and Applied Mathematics, 255(??):74–85, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713002355.

Pirjol:2013:LNI

[Pir13] Dan Pirjol. The logisticnormal integral and its generalizations. Journal of Computational and Applied Mathematics, 237(1):460– 469, January 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712002518.

See addendum [Pir14].

Pirjol:2014:ALN

[Pir14] Dan Pirjol. Addendum:

The logistic-normal integral and its generalizations. Journal of Computational and Applied Mathematics, 260(??):420, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005207.

See [Pir13].

Pitselis:2013:RRE

[Pit13] Georgios Pitselis. A review on robust estimators applied to regression credibility. Journal of Computational and Applied Mathematics, 239(??):231–249, February 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003731.

Pavlovic:2012:RTT

[PJ12] Gorica Pavlović and Svetlana Janković. Razumikhintype theorems on general decay stability of stochastic functional differential equations with infinite delay. Journal of Computational and Applied Mathematics, 236(7):1679–1690, January 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711005310.

Porizkova:2014:NSC

[PKH14] Petra Porízková, Karel Kozel, and Jaromír Horácek. Numerical solution of compressible and incompressible unsteady flows in channel inspired by vocal tract. Journal of Computational and Applied Mathematics, 270(??):323–329, November 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713007188.

Pastor:2010:LAD

[PL10a] Franck Pastor and Etienne Loute. Limit analysis decomposition and finite element mixed method. Journal of Computational and Applied Mathematics, 234(7):2213–2221, August 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709005573.

Peng:2010:APS

[PL10b] Xiao-Fei Peng and Wen Li. An alternating preconditioner for saddle point problems. Journal of Computational and Applied Mathematics, 234(12):3411–3423, October 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710002578.

Prieto:2010:GHP

[PL10c] U. Fidalgo Prieto and G. López Lagomasino. Generalized Hermite–Padé approximation for nikishin systems of three functions. Journal of Computational and Applied Mathematics, 233(6):1525–1533, January 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709001575.

Pang:2011:NHS

[PL11] Shouquan Pang and Yongjian Liu. A new hyperchaotic system from the Lü system and its control. Journal of Computational and Applied Mathematics, 235(8):2775–2789, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006473.

Luo:2010:NCV

[pLjH10] Xue ping Luo and Nan jing Huang. A new class of variational inclusions with B-monotone operators in Banach spaces. Journal of Computational and Applied Mathematics, 233(8):1888–1896, February 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709006475.

Panigrahi:2013:LMP

[PLN13] Bijaya Laxmi Panigrahi, Guangqing Long, and Gnaneshwar Nelakanti. Legendre multi-projection methods for solving eigenvalue problems for a compact integral operator. Journal of Computational and Applied Mathematics, 239(??):135– 151, February 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042712003780.

Pagnini:2010:EEP

[PM10] Gianni Pagnini and Francesco Mainardi. Evolution equations for the probabilistic generalization of the Voigt profile function. Journal of Computational and Applied Mathematics, 233(6):1590– 1595, January 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709001666.

Pardo:2011:CPT

[PM11a] Leandro Pardo and Nirian Martín. On the comparison of the pre-test and shrinkage phi-divergence test estimators for the symmetry model of categorical data. Journal of Computational and Applied Mathematics, 235(5):1160–1179, January 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710004243.

Pultarova:2011:CML

[PM11b] Ivana Pultarová and Ivo Marek. Convergence of multi-level iterative aggregationdisaggregation methods. Journal of Computational and Applied Mathematics, 236(3):354–363, September 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004195.

Panja:2013:SSK

[PM13] M. M. Panja and B. N.

Mandal. Solution of second kind integral equation with Cauchy type kernel using Daubechies scale function. Journal of Computational and Applied Mathematics, 241(??):130–142, March 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712004013.

Prieto:2011:AGF

[PMC11] Francisco Ureña Prieto, Juan José Benito Muñoz, and Luis Gavete Corvinos. Application of the generalized finite difference method to solve the advection–diffusion equation. Journal of Computational and Applied Mathematics, 235(7):1849–1855, February 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271000302X.

Prasad:2014:GCW

[PMMS14] Akhilesh Prasad, Santanu Manna, Ashutosh Mahato, and V. K. Singh. The generalized continuous wavelet transform associated with the fractional Fourier transform. Journal of Computational and Applied Mathematics, 259 (part B)(??):660–671, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713002021.

Panigrahi:2011:SLP

[PN11] Bijaya Laxmi Panigrahi and Gnaneshwar Nelakanti. Superconvergence of Legendre projection methods for the eigenvalue problem of a compact integral operator. Journal of Computational and Applied Mathematics, 235(8):2380–2391, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006047.

Paris:2010:ITS

[PNCC10] J. París, F. Navarrina, I. Colominas, and M. Casteleiro. Improvements in the treatment of stress constraints in structural topology optimization problems. Journal of Computational and Applied Mathematics, 234(7): 2231–2238, August 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042709005597.

Pillay:2011:FBO

[PO11] E. Pillay and J. G. O’Hara.

FFT based option pricing under a mean reverting process with stochastic volatility and jumps. Journal of Computational and Applied Mathematics, 235(12):3378– 3384, April 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S037704271000590X.

Popescu:2010:TPR

[Pop10] Emil Popescu. Trotter products and reaction–diffusion equations. Journal of Computational and Applied Mathematics, 233(6):1596– 1600, January 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042709001678.

Povalej:2014:QNM

[Pov14]  Ziga Povalej.  Quasi-

Newton’s method for multiobjective optimization. Journal of Computational and Applied Mathematics, 255(??):765–777, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713003439.

Pelloni:2010:MBV

[PP10a] B. Pelloni and D. A. Pinotsis. Moving boundary value problems for the wave equation. Journal of Computational and Applied Mathematics, 234(6):1685–1691, July 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709004919.

Pena:2010:CHO

[PP10b] Gonçalo Pena and Christophe Prud’homme. Construction of a high order fluidstructure interaction solver. Journal of Computational and Applied Mathematics, 234(7):2358–2365, August 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709005743.

Pinar:2014:MSD

[PP14] Mustafa Ç. Pinar and A. Burak Paç. Mean semi-deviation from a target and robust portfolio choice under distribution and mean return ambiguity. Journal of Computational and Applied Mathematics, 259 (part B)(??):394–405, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003269.

Presho:2014:LGM

[PPG14] Michael Presho, Anastasiya Protasov, and Eduardo Gildin. Local-global model reduction of parameterdependent, single-phase flow models via balanced truncation. Journal of Computational and Applied Mathematics, 271(??):163– 179, December 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704271400171X.

Petkovic:2010:SFR

[PPH10] M. S. Petković, L. D. Petković, and – D. Herceg. On Schröder’s families of root-finding methods. Journal of Computational and Applied Mathematics, 233(8):1755–1762, February 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709006347.

Paris:2010:NEE

[PR10] A. París and L. Rández. New embedded explicit pairs of exponentially fitted Runge–Kutta methods. Journal of Computational and Applied Mathematics, 234(3):767–776, June 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710000452.

Plaza:2011:ACR

[PR11] Sergio Plaza and Natalia Romero. Attracting cycles for the relaxed Newton’s method. Journal of Computational and Applied Mathematics, 235(10):3238–3244, March 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711000215.

Parashar:2012:ITC

[PR12a] Rishi Parashar and Donald M. Reeves. On iterative techniques for computing flow in large twodimensional discrete fracture networks. Journal of Computational and Applied Mathematics, 236(18):4712– 4724, December 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712001057.

Puccetti:2012:CSB

[PR12b] Giovanni Puccetti and Ludger Rüschendorf. Computation of sharp bounds on the distribution of a function of dependent risks. Journal of Computational and Applied Mathematics, 236(7):1833–1840, January 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711005589.

Prevost:2010:PAA

[Pré10] Marc Prévost. Padé approximation and Apostol– Bernoulli and Apostol– Euler polynomials. Journal of Computational and Applied Mathematics, 233(11): 3005–3017, April 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042709007973.

Petkovic:2011:NFO

[PRM11] M. S. Petković, L. Rancić, and M. R. Milosević. On the new fourth-order methods for the simultaneous approximation of polynomial zeros. Journal of Computational and Applied Mathematics, 235(14):4059–4075, May 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711001208.

Protas:2012:VDP

[Pro12] Bartosz Protas. Vortex design problem. Journal of Computational and Applied Mathematics, 236(7):1926– 1946, January 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711005723.

Pissarenko:2010:FDS

[PRT10] D. Pissarenko, G. Reshetova, and V. Tcheverda. 3D finite-difference synthetic acoustic log in cylindrical coordinates. Journal of Computational and Applied Mathematics, 234(6):1766– 1772, July 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709005019.

Przybylowicz:2010:AIT

[Prz10] Pawel Przybylowicz. Adaptive Itô–Taylor algorithm can optimally approximate the Itô integrals of singular functions. Journal of Computational and Applied Mathematics, 235(1):203– 217, November 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710003237.

Przybylowicz:2013:OSD

[Prz13] Pawel Przybylowicz. Optimal sampling design for approximation of stochastic Itô integrals with application to the nonlinear Lebesgue integration. Journal of Computational and Applied Mathematics, 245 (??):10–29, June 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712005377.

Potthast:2010:NRT

[PS10] Roland Potthast and Mourad Sini. The no response test for the reconstruction of polyhedral objects in electromagnetics. Journal of Computational and Applied Mathematics, 234(6):1739– 1746, July 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709004981.

Pazouki:2011:BKB

[PS11a] Maryam Pazouki and Robert Schaback. Bases for kernelbased spaces. Journal of Computational and Applied Mathematics, 236(4):575– 588, September 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042711002688.

Petkovic:2011:IMC

[PS11b] Marko D. Petković and Predrag S. Stanimirović. Iterative method for computing the Moore–Penrose inverse based on Penrose equations. Journal of Computational and Applied Mathematics, 235(6):1604–1613, January 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710004917.

Phuengrattana:2011:RCM

[PS11c] Withun Phuengrattana and Suthep Suantai. On the rate of convergence of Mann, Ishikawa, Noor and SPiterations for continuous functions on an arbitrary interval. Journal of Computational and Applied Mathematics, 235(9):3006–3014, March 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006722.

Pazouki:2013:BCP

[PS13] Maryam Pazouki and Robert Schaback. Bases for conditionally positive definite kernels. Journal of Computational and Applied Mathematics, 243(??):152–163, May 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712004633.

Pejcev:2014:EBM

[PS14a] Aleksandar V. Pejcev and Miodrag M. Spalević. Error bounds of the Micchelli– Sharma quadrature formula for analytic functions. Journal of Computational and Applied Mathematics, 259 (part A)(??): 48–56, March 15, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713001477.

Petkovic:2014:TII

[PS14b] Marko D. Petković and Predrag S. Stanimirović. Two improvements of the iterative method for computing Moore–Penrose inverse based on Penrose equations. Journal of Computational and Applied Mathematics, 267(??):61– 71, September 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042714000624.

Popa:2014:ECL

[PS14c] Emil C. Popa and Nicolae-Adrian Secelean. Estimates for the constants of Landau and Lebesgue via some inequalities for the Wallis ratio. Journal of Computational and Applied Mathematics, 269(??):68–74, October 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714001691.

Pugal:2014:FET

[PSKA14] D. Pugal, P. Solin, K. J.

Kim, and A. Aabloo. hp-FEM electromechanical transduction model of ionic polymer-metal composites. Journal of Computational and Applied Mathematics, 260(??):135–148, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713004597.

Pop:2010:EEF

[PSRV10] I. S. Pop, M. Sepúlveda, F. A. Radu, and O. P. Vera Villagrán. Error estimates for the finite volume discretization for the porous medium equation. Journal of Computational and Applied Mathematics, 234(7):2135–2142, August 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709005494.

Park:2013:SDN

[PSS13] Chunjae Park, Dongwoo Sheen, and Byeong-Chun Shin. A subspace of the DSSY nonconforming quadrilateral finite element space for the Stokes equations. Journal of Computational and Applied Mathematics, 239(??):220–230, February 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712004153.

Pedas:2011:CSC

[PT11a] Arvet Pedas and Enn Tamme. On the convergence of spline collocation methods for solving fractional differential equations. Journal of Computational and Applied Mathematics, 235(12):3502–3514, April 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711000823.

Pedas:2011:SCM

[PT11b] Arvet Pedas and Enn Tamme. Spline collocation methods for linear multi-term fractional differential equations. Journal of Computational and Applied Mathematics, 236(2):167– 176, August 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042711003384.

Pedas:2012:PPC

[PT12] Arvet Pedas and Enn Tamme. Piecewise polynomial collocation for linear boundary value problems of fractional differential equations. Journal of Computational and Applied Mathematics, 236(13):3349– 3359, July 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712001197.

Pedas:2014:NSN

[PT14] Arvet Pedas and Enn Tamme. Numerical solution of nonlinear fractional differential equations by spline collocation methods. Journal of Computational and Applied Mathematics, 255(??):216–230, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713002574.

Pan:2010:IMI

[PTH10] Kejia Pan, Yongji Tan, and Hongling Hu. An interpolation matched interface and boundary method for elliptic interface problems. Journal of Computational and Applied Mathematics, 234 (1):73–94, May 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709008140.

Pastor:2010:LAN

[PTP10] J. Pastor, Ph. Thoré, and F. Pastor. Limit analysis and numerical modeling of spherically porous solids with Coulomb and Drucker– Prager matrices. Journal of Computational and Applied Mathematics, 234(7):2162– 2174, August 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042709005524.

Pellikka:2013:PIR

[PTSK13] M. Pellikka, T. Tarhasaari, S. Suuriniemi, and L. Kettunen. A programming interface to the Riemannian manifold in a finite element environment. Journal of Computational and Applied Mathematics, 246(??):225– 233, July 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712004591.

Pulch:2014:SCS

[Pul14] Roland Pulch. Stochastic collocation and stochastic Galerkin methods for linear differential algebraic equations. Journal of Computational and Applied Mathematics, 262(??):281–291, May 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005992.

Piret:2010:SAC

[PV10] Kathy Piret and Jan Verschelde. Sweeping algebraic curves for singular solutions. Journal of Computational and Applied Mathematics, 234(4):1228–1237, June 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709004609.

Pazos:2014:MRR

[PVM^{+}14] Rodolfo A. Pazos, Graciela Vázquez, José A. Martínez, Joaquín Pérez-Ortega, and Gilberto Martínez-Luna. Minimizing roundtrip response time in distributed databases with vertical fragmentation. Journal of Computational and Applied Mathematics, 259 (part B) (??):905–913, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042713005128.

Perez-Villalon:2013:CWC

[PVP13] G. Pérez-Villalón and A. Portal. Computation of wavelet coefficients from average samples. Journal of Computational and Applied Mathematics, 248(??):118–130, August 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713000629.

Pan:2012:NSL

[PW12] Zhenguo Pan and Brian Wetton. Numerical simulation and linear wellposedness analysis for a class of three-phase boundary motion problems. Journal of Computational and Applied Mathematics, 236 (13):3160–3173, July 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042712000702.

Pestana:2013:CPM

[PW13] Jennifer Pestana and Andrew J. Wathen. On the choice of preconditioner for minimum residual methods for non-Hermitian matrices. Journal of Computational and Applied Mathematics, 249(??):57– 68, September 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713000903.

Wu:2014:EAM

[pWqSmZ14] Jian ping Wu, Jun qiang Song, and Wei min Zhang. An efficient and accurate method to compute the Fiedler vector based on Householder deflation and inverse power iteration. Journal of Computational and Applied Mathematics, 269(??):101–108, October 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714001587.

Peng:2011:FEB

[PWZ11] Hai-Jun Peng, Zhi-Gang Wu, and Wan-Xie Zhong. Fourier expansion based recursive algorithms for periodic Riccati and Lyapunov matrix differential equations. Journal of Computational and Applied Mathematics, 235(12):3571–3588, April 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711000872.

Pang:2010:TSM

[PY10a] Zhi-Feng Pang and Yu-Fei Yang. A two-step model for image denoising using a duality strategy and surface fitting. Journal of Computational and Applied Mathematics, 235(1):82–90, November 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710002906.

Peng:2010:NSG

[PY10b] Jian-Wen Peng and Jen-Chih Yao. On a new system of generalized mixed quasi-variational-like inclusions involving (A, η, m)- accretive operators with applications. Journal of Computational and Applied Mathematics, 234(1):21–33, May 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709008097.

Yang:2010:DPC

[pYXPH10] Li ping Yang, Xiangsheng Xie, Shiguo Peng, and Gang Hu. Demiclosed principle and convergence for modified three step iterative process with errors of non-Lipschitzian mappings. Journal of Computational and Applied Mathematics, 234(4):972–984, June 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709000417.

Pu:2013:CIM

[PZ13] Zhao-Nian Pu and Mu-Zheng Zhu. A class of iteration methods based on the generalized preconditioned Hermitian and skew-Hermitian splitting for weakly nonlinear systems. Journal of Computational and Applied Mathematics, 250(??):16–27, October 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713000915.

Pang:2012:TDP

[PZJ12] Hong-Kui Pang, Ying-Ying Zhang, and Xiao-Qing Jin. Tri-diagonal preconditioner for pricing options. Journal of Computational and Applied Mathematics, 236(17):4365–4374, November 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271200163X.

Zeng:2012:EEL

[pZxY12] Jin ping Zeng and Hai xiong Yu. Error estimates of the lumped mass finite element method for semilinear elliptic problems. Journal of Computational and Applied Mathematics, 236(7):1993– 2004, January 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042711005796.

Qian:2014:QIE

[QC14] Jiang Qian and Mingsong Cheng. Quadratic inverse eigenvalue problem for damped gyroscopic systems. Journal of Computational and Applied Mathematics, 255(??):306–312, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713002872.

Qin:2010:SCS

[QCK10] Xiaolong Qin, Sun Young Cho, and Shin Min Kang. Strong convergence of shrinking projection methods for quasi-ϕ-nonexpansive mappings and equilibrium problems. Journal of Computational and Applied Mathematics, 234(3):750–760, June 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710000191.

Qi:2010:CMS

[QGG10] Feng Qi, Senlin Guo, and Bai-Ni Guo. Complete monotonicity of some functions involving polygamma functions. Journal of Computational and Applied Mathematics, 233(9):2149– 2160, March 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709006682.

Qin:2010:SE

[QH10] Keyun Qin and Zhiyong Hong. On soft equality. Journal of Computational and Applied Mathematics, 234(5):1347–1355, July 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710001263.

Qi:2014:IRC

[Qi14] Feng Qi. Integral representations and complete monotonicity related to the remainder of Burnside’s formula for the gamma function. Journal of Computational and Applied Mathematics, 268(??):155–167, October 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714001356.

Qin:2011:MCV

[QLL11] Rui Qin, Yan-Kui Liu, and Zhi-Qiang Liu. Methods of critical value reduction for type-2 fuzzy variables and their applications. Journal of Computational and Applied Mathematics, 235(5):1454–1481, January 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710004802.

Qian:2013:SIE

[QT13] Jiang Qian and Roger C. E. Tan. On some inverse eigenvalue problems for Hermitian and generalized Hamiltonian/ skew-Hamiltonian matrices. Journal of Computational and Applied Mathematics, 250(??):28–38, October 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713000939.

Qi:2014:NSV

[QTS14] Zhiquan Qi, Yingjie Tian, and Yong Shi. A nonparallel support vector machine for a classification problem with universum learning. Journal of Computational and Applied Mathematics, 263(??):288–298, June 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713006213.

Qiu:2010:SBP

[QW10] Yuyang Qiu and Anding Wang. Solving balanced Procrustes problem with some constraints by eigenvalue decomposition. Journal of Computational and Applied Mathematics, 233(11):2916–2924, April 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709007833.

Wang:2011:LSM

[qW11] Zhong qing Wang. The Laguerre spectral method for solving Neumann boundary value problems. Journal of Computational and Applied Mathematics, 235(10):3229– 3237, March 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711000203.

Wei:2011:NMC

[qWhY11] Jun qiang Wei and Zhong hua Yang. A novel method for computation of higher order singular points in nonlinear problems with single parameter. Journal of Computational and Applied Mathematics, 236(6):1442–1448, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711004870.

Qian:2010:QRR

[QXW10] Ai-Lin Qian, Xiang-Tuan Xiong, and Yu-Jiang Wu. On a quasi-reversibility regularization method for a Cauchy problem of the Helmholtz equation. Journal of Computational and Applied Mathematics, 233(8):1969–1979, February 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709006554.

Quillen:2010:BIF

[QY10] Patrick Quillen and Qiang Ye. A block inverse-free preconditioned Krylov subspace method for symmetric generalized eigenvalue problems. Journal of Computational and Applied Mathematics, 233(5):1298–1313, January 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709001812.

Qian:2014:GCS

[QYZZ14] Xiaoxiang Qian, Huina Yuan, Mozhen Zhou, and Bingyin Zhang. A general 3D contact smoothing method based on radial point interpolation. Journal of Computational and Applied Mathematics, 257(??): 1–13, February 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713004196.

Qin:2011:GAN

[QZX^{+}11] Xianyun Qin, Jianming Zhang, Guizhong Xie, Fenglin Zhou, and Guanyao Li. A general algorithm for the numerical evaluation of nearly singular integrals on 3D boundary element. Journal of Computational and Applied Mathematics, 235(14):4174–4186, May 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711001427.

Ren:2010:LCI

[RA10] Hongmin Ren and Ioannis K. Argyros. On the local convergence of inexact Newton-type methods under residual controltype conditions. Journal of Computational and Applied Mathematics, 235(1):218– 228, November 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042710003250.

Rahman:2014:RCM

[Rah14] S. Rahman. Regularity criterion for 3D MHD fluid passing through the porous medium in terms of gradient pressure. Journal of Computational and Applied Mathematics, 270(??):88– 99, November 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042714001058.

Ramm:2010:NNS

[Ram10]  A. G. Ramm.  On a

new notion of the solution to an ill-posed problem. Journal of Computational and Applied Mathematics, 234(12):3326–3331, October 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710002402.

Rang:2014:APR

[Ran14] Joachim Rang. An analysis of the Prothero–Robinson example for constructing new DIRK and ROW methods. Journal of Computational and Applied Mathematics, 262(??):105–114, May 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005177.

Rattana:2013:MMC

[RB13] Amornrat Rattana and Christine Böckmann. Matrix methods for computing eigenvalues of Sturm– Liouville problems of order four. Journal of Computational and Applied Mathematics, 249(??):144–156, September 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713000940.

Rischette:2011:NAE

[RBD11] R. Rischette, T. N. Baranger, and N. Debit. Numerical analysis of an energylike minimization method to solve the Cauchy problem with noisy data. Journal of Computational and Applied Mathematics, 235(11):3257– 3269, April 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710006692.

Rischette:2014:NAE

[RBD14] R. Rischette, T. N. Baranger, and N. Debit. Numerical analysis of an energylike minimization method to solve a parabolic Cauchy problem with noisy data.

Journal of Computational and Applied Mathematics, 271(??):206–222, December 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714001733.

Rakkiyappan:2011:DDS

[RBK11] R. Rakkiyappan, P. Balasubramaniam, and R. Krishnasamy. Delay dependent stability analysis of neutral systems with mixed time-varying delays and nonlinear perturbations. Journal of Computational and Applied Mathematics, 235(8):2147–2156, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710005674.

Raghuram:2012:HON

[RCM^{+}12] Karthik Jayaraman Raghuram, Shivkumar Chandrasekaran, Joseph Moffitt, Ming Gu, and Hrushikesh Mhaskar. Higher order numerical discretizations for exterior and biharmonic type PDEs. Journal of Computational and Applied Mathematics, 236(18):4762– 4774, December 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712000829.

Rebelo:2010:HCM

[RD10] Magda Rebelo and Teresa Diogo. A hybrid collocation method for a nonlinear Volterra integral equation with weakly singular kernel. Journal of Computational and Applied Mathematics, 234(9):2859–2869, September 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710000385.

Ren:2010:GRB

[RE10] Yong Ren and Mohamed El Otmani. Generalized reflected BSDEs driven by a Lévy process and an obstacle problem for PDIEs with a nonlinear Neumann boundary condition. Journal of Computational and Applied Mathematics, 233(8):2027–2043, February 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704270900661X.

Rebaza:2012:DPT

[Reb12] Jorge Rebaza. Dynamics of prey threshold harvesting and refuge. Journal of Computational and Applied Mathematics, 236(7):1743– 1752, January 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S037704271100536X.

Renard:2010:SDM

[Ren10] Yves Renard. The singular dynamic method for constrained second order hyperbolic equations: Application to dynamic contact problems. Journal of Computational and Applied Mathematics, 234(3):906– 923, June 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710000713.

Reutskiy:2010:MEE

[Reu10]  S. Yu. Reutskiy.  The

method of external excitation for solving generalized Sturm–Liouville problems. Journal of Computational and Applied Mathematics, 233(9):2374–2386, March 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709007201.

Rashidinia:2011:BSC

[RG11] J. Rashidinia and M. Ghasemi. B-spline collocation for solution of two-point boundary value problems. Journal of Computational and Applied Mathematics, 235 (8):2325–2342, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042710005972.

Ramabathiran:2012:AEM

[RG12] Amuthan Arunkumar Ramabathiran and S. Gopalakrishnan. Automatic energymomentum conserving time integrators for hyperelastic waves. Journal of Computational and Applied Mathematics, 236(18):4700–4711, December 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712001070.

Rashidinia:2010:NSN

[RGJ10] J. Rashidinia, M. Ghasemi, and R. Jalilian. Numerical solution of the nonlinear Klein–Gordon equation. Journal of Computational and Applied Mathematics, 233(8):1866–1878, February 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709006451.

Ren:2011:CIN

[RHS11] Yong Ren, Lanying Hu, and R. Sakthivel. Controllability of impulsive neutral stochastic functional differential inclusions with infinite delay. Journal of Computational and Applied Mathematics, 235(8):2603–2614, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271000628X.

Riedler:2013:ASC

[Rie13] Martin G. Riedler. Almost sure convergence of numerical approximations for piecewise deterministic Markov processes. Journal of Computational and Applied Mathematics, 239(??):50–71, February 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003925.

Rao:2011:AFO

[RK11] S. Chandra Sekhara Rao and Sunil Kumar. An almost fourth order uniformly convergent domain decomposition method for a coupled system of singularly perturbed reaction– diffusion equations. Journal of Computational and Applied Mathematics, 235(11): 3342–3354, April 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042711000586.

Raisanen:2013:HCW

[RKS^{+}13] V. Räisänen, S. Kurz, S. Suuriniemi, T. Tarhasaari, and L. Kettunen. How can we deal with moving objects on a fixed mesh? Journal of Computational and Applied Mathematics, 246(??):260– 268, July 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003937.

Rohan:2010:HAT

[RL10] E. Rohan and V. Lukes. Homogenization of the acoustic transmission through a perforated layer. Journal of Computational and Applied Mathematics, 234(6):1876– 1885, July 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709005159.

Roanes-Lozano:2014:ERR

[RLGGGÁM14] Eugenio Roanes-Lozano, José Luis Galán-García, Alberto García-Álvarez, and Luis Mesa. Estimating radial railway network improvement with a CAS. Journal of Computational and Applied Mathematics, 270(??):294–307, November 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714000041.

Rocha:2011:ALF

[RMF11] Ana Maria A. C. Rocha, Tiago F. M. C. Martins, and Edite M. G. P. Fernandes. An augmented Lagrangian fish swarm based method for global optimization. Journal of Computational and Applied Mathematics, 235(16):4611–4620, June 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710002177.

Radu:2013:MFE

[RMP^{+}13] Florin A. Radu, Adrian Muntean, Iuliu S. Pop, Nicolae Suciu, and Olaf Kolditz. A mixed finite element discretization scheme for a concrete carbonation model with concentrationdependent porosity. Journal of Computational and Applied Mathematics, 246(??): 74–85, July 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712004438.

Riachy:2011:MND

[RMR11] Samer Riachy, Mamadou Mboup, and Jean-Pierre Richard. Multivariate numerical differentiation. Journal of Computational and Applied Mathematics, 236(6):1069–1089, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004262.

Rivas:2011:FED

[RMTO11] T. Rivas, J. M. Matías, J. Taboada, and C. Ordóñez. Functional experiment design for the analysis of colour changes in granite using new L^{∗}a^{∗}b^{∗} functional colour coordinates. Journal of Computational and Applied Mathematics, 235(16):4701–4716, June 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710004437.

Romance:2011:LEE

[Rom11] Miguel Romance. Local estimates for eigenvector-like centralities of complex networks. Journal of Computational and Applied Mathematics, 235(7):1868–1874, February 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710003766.

Roshan:2011:PGM

[Ros11]  Thoudam Roshan.  A

Petrov–Galerkin method for solving the generalized equal width (GEW) equation. Journal of Computational and Applied Mathematics, 235(6):1641–1652, January 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710004978.

Radu:2010:NMR

[RP10] F. A. Radu and I. S. Pop. Newton method for reactive solute transport with equilibrium sorption in porous media. Journal of Computational and Applied Mathematics, 234(7):2118–2127, August 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709005470.

Romanovski:2011:ASS

[RP11a] Valery G. Romanovski and Mateja Presern. An approach to solving systems of polynomials via modular arithmetics with applications. Journal of Computational and Applied Mathematics, 236(2):196–208, August 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711003542.

Rosca:2011:USG

[RP11b] Daniela Rosca and Gerlind Plonka. Uniform spherical grids via equal area projection from the cube to the sphere. Journal of Computational and Applied Mathematics, 236(6):1033–1041, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711003955.

Ramathilagam:2011:MFM

[RPS^{+}11] S. Ramathilagam, R. Pandiyarajan, A. Sathya, R. Devi, and S. R. Kannan. Modified fuzzy c-means algorithm for segmentation of T1–T2-weighted brain MRI. Journal of Computational and Applied Mathematics, 235(6):1578–1586, January 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710004826.

Rukavishnikov:2010:FEM

[RR10] V. A. Rukavishnikov and H. I. Rukavishnikova. The finite element method for a boundary value problem with strong singularity. Journal of Computational and Applied Mathematics, 234(9):2870–2882, September 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710000245.

Ruiz-Ramirez:2013:SSP

[RRMD13] J. Ruiz-Ramírez and J. E. Macías-Díaz. A skew symmetry-preserving computational technique for obtaining the positive and the bounded solutions of a time-delayed advection– diffusion-reaction equation. Journal of Computational and Applied Mathematics, 250(??):256–269, October 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713001246.

Rahunanthan:2010:HOM

[RS10a] Arunasalam Rahunanthan and Dan Stanescu. Highorder W -methods. Journal of Computational and Applied Mathematics, 233 (8):1798–1811, February 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0377042709006396.

Reichel:2010:CMM

[RS10b] Lothar Reichel and Andriy Shyshkov. Cascadic multilevel methods for illposed problems. Journal of Computational and Applied Mathematics, 233(5):1314– 1325, January 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709002192.

Rehm:2014:POD

[RSFS14] Amanda M. Rehm, Elizabeth Y. Scribner, and Hassan M. Fathallah-Shaykh. Proper orthogonal decomposition for parameter estimation in oscillating biological networks. Journal of Computational and Applied Mathematics, 258(??):135– 150, March 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713004524.

Rodrigo:2014:LFA

[RSGL14] C. Rodrigo, P. Salinas, F. J.

Gaspar, and F. J. Lisbona. Local Fourier analysis for cell-centered multigrid methods on triangular grids. Journal of Computational and Applied Mathematics, 259 (part A)(??):

35–47, March 15, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042713001489.

Rees:2014:APD

[RSK14] Patrick Rees and A. Sevtap Selcuk-Kestel. Analysis of portfolio diversification between REIT assets. Journal of Computational and Applied Mathematics, 259 (part B)(??):425–433, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713004421.

Rueda:2014:RHD

[RSS14] Sonia L. Rueda, Juana Sendra, and J. Rafael Sendra. Rational Hausdorff divisors: a new approach to the approximate parametrization of curves. Journal of Computational and Applied Mathematics, 263(??):445–465, June 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042714000053.

Roberts:2014:SSL

[RST14] Jason A. Roberts, Dmitry V. Savostyanov, and Eugene E. Tyrtyshnikov. Superfast solution of linear convolutional Volterra equations using QTT approximation. Journal of Computational and Applied Mathematics, 260(??):434–448, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005694.

Rapetti:2012:GSL

[RSV12] Francesca Rapetti, Alvise Sommariva, and Marco Vianello. On the generation of symmetric Lebesguelike points in the triangle. Journal of Computational and Applied Mathematics, 236(18):4925–4932, December 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711005930.

Rogava:2012:LCS

[RT12] J. Rogava and M. Tsiklauri. On local convergence of a symmetric semi-discrete scheme for an abstract analogue of the Kirchhoff equation. Journal of Computational and Applied Mathematics, 236(15):3654–3664, September 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271100389X.

Rincon:2014:NSD

[RTL14] M. A. Rincon, F. S. Teixeira, and I. F. Lopez. Numerical studies of the damped Korteweg–de Vries system. Journal of Computational and Applied Mathematics, 259 (part A)(??): 294–311, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042713005001.

Roos:2013:QSP

[RU13] Hans-G. Roos and Zorica Uzelac. Qualocation for a singularly perturbed boundary value problem. Journal of Computational and Applied Mathematics, 237(1):556–564, January 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712002701.

Ruas:2013:HFE

[Rua13] Vitoriano Ruas. Hermite finite elements for second order boundary value problems with sharp gradient discontinuities. Journal of Computational and Applied Mathematics, 246(??):234– 242, July 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003494.

Rump:2013:ASDa

[Rum13a] Siegfried M. Rump. Accurate solution of dense linear systems, Part I: Algorithms in rounding to nearest. Journal of Computational and Applied Mathematics, 242(??):157–184, April 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712004360.

Rump:2013:ASDb

[Rum13b] Siegfried M. Rump. Accurate solution of dense linear systems, part II: Algorithms using directed rounding. Journal of Computational and Applied Mathematics, 242(??):185– 212, April 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003974.

Rauwoens:2010:CPC

[RVDM10] P. Rauwoens, J. Vierendeels, E. Dick, and B. Merci. A conservative pressurecorrection scheme for transient simulations of reacting flows. Journal of Computational and Applied Mathematics, 234(7):2311–2318, August 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709005688.

Roman:2013:ECC

[RVN13] J. E. Roman, P. B. Vasconcelos, and A. L. Nunes. Eigenvalue computations in the context of datasparse approximations of integral operators. Journal of Computational and Applied Mathematics, 237(1):171– 181, January 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712003111.

Ruiz-Vanoye:2011:SPT

[RVPOR^{+}11] Jorge A. Ruiz-Vanoye, Joaquín Pérez-Ortega, Rodolfo A. Pazos R., Ocotlán Díaz-Parra, Juan Frausto-Solís, Hector J. Fraire Huacuja, Laura Cruz-Reyes, and José A. Martínez F. Survey of polynomial transformations between NP-complete problems. Journal of Computational and Applied Mathematics, 235(16):4851–4865, June 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271100094X.

Ragusa:2010:TMA

[RW10] Jean C. Ragusa and Yaqi Wang. A two-mesh adaptive mesh refinement technique for S_{N} neutral-particle transport using a higherorder DGFEM. Journal of Computational and Applied Mathematics, 233(12):3178– 3188, April 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709008280.

Rjasanow:2014:FTT

[RW14] Sergej Rjasanow and Steffen Weißer. FEM with Trefftz trial functions on polyhedral elements. Journal of Computational and Applied Mathematics, 263(??):202– 217, June 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713006961.

Rudolf:2014:MFV

[RWRS14] Florian Rudolf, Josef Weinbub, Karl Rupp, and Siegfried Selberherr. The meshing framework ViennaMesh for finite element applications. Journal of Computational and Applied Mathematics, 270(??):166– 177, November 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042714000818.

Rui:2010:SIN

[RX10] Shao-Ping Rui and Cheng-Xian Xu. A smoothing inexact Newton method for nonlinear complementarity problems. Journal of Computational and Applied Mathematics, 233(9):2332– 2338, March 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042709006967.

Rohani:2014:LBM

[RYM^{+}14] Babak Rohani, Yigit Yazicioğlu, Mehmet Mutlu, Orkun Ogucu, Emre Akgul, and Afsar Saranli. Lagrangian based mathematical modeling and experimental validation of a planar stabilized platform for mobile systems. Journal of Computational and Applied Mathematics, 259 (part B)(??):955–964, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713005621.

Rzadkowski:2012:SEE

[Rza12] Grzegorz Rzadkowski. On some expansions for the Euler gamma function and the Riemann zeta function. Journal of Computational and Applied Mathematics, 236(15):3710–3719, September 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004663.

Remski:2014:BMM

[RZD14] Joan Remski, Jingyan Zhang, and Qiang Du. On balanced moving mesh methods. Journal of Computational and Applied Mathematics, 265(??):255– 263, August 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713004962.

Sakalli:2014:ACC

[SA14] Muharrem Tolga Sakalli and Bora Aslan. On the algebraic construction of cryptographically good 32 × 32 binary linear transformations. Journal of Computational and Applied Mathematics, 259 (part B)(??): 485–494, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042713002719.

Serenbay:2014:GBT

[SAB14] Sevilay Kirci Serenbay, Çigdem Atakut, and Ibrahim Büyükyazici. The generalized Baskakov type operators. Journal of Computational and Applied Mathematics, 259 (part A)(??): 226–232, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042713004007.

Solin:2010:PIA

[SACS10] Pavel Solin, David Andrs, Jakub Cerveny, and Miroslav Simko. PDEindependent adaptive hp– FEM based on hierarchic extension of finite element spaces. Journal of Computational and Applied Mathematics, 233(12):3086–3094, April 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709003446.

Sanguesa:2011:AFS

[San11] C. Sangüesa. On the approximation of functions satisfying defective renewal equations. Journal of Computational and Applied Mathematics, 236(6):1042– 1054, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042711004006.

Shokri:2011:NTS

[SASH11] A. Shokri, M. Y. Rahimi Ardabili, S. Shahmorad, and G. Hojjati. A new two-step P -stable hybrid Obrechkoff method for the numerical integration of second-order IVPs. Journal of Computational and Applied Mathematics, 235(6):1706–1712, January 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710005285.

Sahu:2013:UHI

[SAY13] D. R. Sahu, Q. H. Ansari, and J. C. Yao. A unified hybrid iterative method for hierarchical minimization problems. Journal of Computational and Applied Mathematics, 253(??):208– 221, December 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713002045.

Slodicka:2010:DMS

[SB10] Marián Slodicka and Antónia Balázová. Decomposition method for solving multispecies reactive transport problems coupled with firstorder kinetics applicable to a chain with identical reaction rates. Journal of Computational and Applied Mathematics, 234(4):1069– 1077, June 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704270900274X.

Sliva:2010:SES

[SBCD10] G. Sliva, A. Brezillon, J. M. Cadou, and L. Duigou. A study of the eigenvalue sensitivity by homotopy and perturbation methods. Journal of Computational and Applied Mathematics, 234(7):2297–2302, August 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709005664.

Sanni:2010:RAS

[SBFC10] Issa Sanni, Emmanuel Bellenger, Jérôme Fortin, and Patrice Coorevits. A reliable algorithm to solve 3D frictional multi-contact problems: Application to granular media. Journal of Computational and Applied Mathematics, 234(4):1161– 1171, June 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709004087.

Sunhaloo:2010:BCP

[SBGB10] M. S. Sunhaloo, R. Boojhawon, A. Gopaul, and M. Bhuruth. On blockcirculant preconditioners for high-order compact approximations of convection– diffusion problems. Journal of Computational and Applied Mathematics, 234(4): 1312–1323, June 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042710000105.

Sgura:2012:NAT

[SBL12] Ivonne Sgura, Benedetto Bozzini, and Deborah Lacitignola. Numerical approximation of Turing patterns in electrodeposition by ADI methods. Journal of Computational and Applied Mathematics, 236(16):4132– 4147, October 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704271200132X.

Safkhani:2014:SRA

[SBM14] Masoumeh Safkhani, Nasour Bagheri, and Ali Mahani. On the security of RFID anti-counting security protocol (ACSP).

Journal of Computational and Applied Mathematics, 259 (part B)(??):512–521, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713003142.

Sella:2010:CSD

[SC10a] Lorenzo Sella and Pieter Collins. Computation of symbolic dynamics for onedimensional maps. Journal of Computational and Applied Mathematics, 234(2): 418–436, May 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042709008516.

Shanin:2010:PDO

[SC10b] A. V. Shanin and R. V. Craster. Pseudo-differential operators for embedding formulae. Journal of Computational and Applied Mathematics, 234(6):1637– 1646, July 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042709004853.

Sheng:2010:IMS

[SC10c] Xingping Sheng and Guoliang Chen. An iterative method for the symmetric and skew symmetric solutions of a linear matrix equation AXB + CY D = E. Journal of Computational and Applied Mathematics, 233(11):3030–3040, April 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709007997.

Steinhoff:2010:LDW

[SC10d] J. S. Steinhoff and S. Chitta.

Long distance wave computation using nonlinear solitary waves. Journal of Computational and Applied Mathematics, 234(6):1826– 1833, July 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709005093.

Solin:2010:MDL

[SCDA10] P. Solin, J. Cerveny, L. Dubcova, and D. Andrs. Monolithic discretization of linear thermoelasticity problems via adaptive multimesh hp– FEM. Journal of Computational and Applied Mathematics, 234(7):2350–2357, August 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709005731.

Sajo-Castelli:2014:PCG

[SCFR14] A. M. Sajo-Castelli, M. A. Fortes, and M. Raydan. Preconditioned conjugate gradient method for finding minimal energy surfaces on Powell–Sabin triangulations. Journal of Computational and Applied Mathematics, 268(??):34–55, October 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042714001253.

Schiefermayr:2011:EAC

[Sch11] Klaus Schiefermayr. Estimates for the asymptotic convergence factor of two intervals. Journal of Computational and Applied Mathematics, 236(1):28–38, August 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710003523.

Schmitt:2014:PMI

[Sch14a] Bernhard A. Schmitt. Peer methods with improved embedded sensitivities for parameter-dependent ODEs. Journal of Computational and Applied Mathematics, 256(??):242–253, January 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003865.

Schmitt:2014:RPM

[Sch14b] Bernhard A. Schmitt. Reprint of: Peer methods with improved embedded sensitivities for parameterdependent ODEs. Journal of Computational and Applied Mathematics, 262(??): 25–36, May 15, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704271300722X.

Schnepp:2014:EDD

[Sch14c] Sascha M. Schnepp. Errordriven dynamical hp-meshes with the discontinuous Galerkin method for threedimensional wave propagation problems. Journal of Computational and Applied Mathematics, 270(??):353– 368, November 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713007115.

Schroder:2014:CCV

[Sch14d]  Michael Schröder.  On

contingent-claim valuation in continuous-time for volatility models of Ornstein– Uhlenbeck type. Journal of Computational and Applied Mathematics, 260(??): 36–53, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713004937.

S:2014:ACS

[SCRB^{+}14] Claudia G. Gómez S., Laura Cruz-Reyes, Juan J. González B., Héctor J. Fraire H., Rodolfo A. Pazos R., and Juan J. Martínez P. Ant colony system with characterizationbased heuristics for a bottled-products distribution logistics system. Journal of Computational and Applied Mathematics, 259 (part B)(??):965–977, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042713005888.

Sharp:2012:PPL

[SCRG12] P. W. Sharp, J. C. Castillo-Rogez, and K. R. Grazier. The performance of phase-lag enhanced explicit Runge–Kutta–Nyström pairs on N-body problems. Journal of Computational and Applied Mathematics, 236 (9):2378–2386, March 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042711005942.

Slodicka:2010:NPE

[SD10] Marián Slodicka and Sofiane Dehilis. A nonlinear parabolic equation with a nonlocal boundary term. Journal of Computational and Applied Mathematics, 233(12):3130–3138, April 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709004373.

Salehi:2013:GML

[SD13] Rezvan Salehi and Mehdi Dehghan. A generalized moving least square reproducing kernel method. Journal of Computational and Applied Mathematics, 249(??):120–132, September 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042713000654.

Solin:2010:AFD

[SDK10] Pavel Solin, Lenka Dubcova, and Jaroslav Kruis. Adaptive hp–FEM with dynamical meshes for transient heat and moisture transfer problems. Journal of Computational and Applied Mathematics, 233(12):3103–3112, April 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709004282.

Segeth:2012:CPE

[Seg12] Karel Segeth. A comparison of a posteriori error estimates for biharmonic problems solved by the FEM. Journal of Computational and Applied Mathematics, 236(18):4788–4797, December 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712000714.

Sezer:2010:MIS

[Sez10] Ali Devin Sezer. Modeling of an insurance system and its large deviations analysis. Journal of Computational and Applied Mathematics, 235(3):535–546, December 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271000347X.

Sakkalis:2012:PHC

[SF12] Takis Sakkalis and Rida T. Farouki. Pythagoreanhodograph curves in Euclidean spaces of dimension greater than 3. Journal of Computational and Applied Mathematics, 236(17):4375– 4382, November 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712001628.

Solin:2012:IAF

[SG12] Pavel Solin and Stefano Giani. An iterative adaptive finite element method for elliptic eigenvalue problems. Journal of Computational and Applied Mathematics, 236(18):4582–4599, December 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712002038.

Sheng:2011:ETB

[SGG11] Qin Sheng, Shekhar Guha, and Leonel P. Gonzalez. An exponential transformation based splitting method for fast computations of highly oscillatory solutions. Journal of Computational and Applied Mathematics, 235(15):4452– 4463, June 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711001907.

Svadlenka:2014:VMM

[SGO14] Karel Svadlenka, Elliott Ginder, and Seiro Omata. A variational method for multiphase volume-preserving interface motions. Journal of Computational and Applied Mathematics, 257(??): 157–179, February 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042713004391.

Shi:2010:HLG

[SGZ10] Lei Shi, Xin Guo, and Ding-Xuan Zhou. Hermite learning with gradient data. Journal of Computational and Applied Mathematics, 233(11):3046–3059, April 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709008061.

Shang:2010:AVI

[SH10] Xufeng Shang and Danfu Han. Application of the variational iteration method for solving n-th-order integro-differential equations. Journal of Computational and Applied Mathematics, 234(5):1442– 1447, July 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710001020.

Shadimetov:2011:OQF

[SH11] Kh. M. Shadimetov and A. R. Hayotov. Optimal quadrature formulas with positive coefficients in L_{2}^{(m)}(0, 1) space. Journal of Computational and Applied Mathematics, 235(5):1114– 1128, January 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704271000419X.

Shunn:2012:SQR

[SH12] Lee Shunn and Frank Ham.

Symmetric quadrature rules for tetrahedra based on a cubic close-packed lattice arrangement. Journal of Computational and Applied Mathematics, 236(17):4348– 4364, November 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712001604.

Sainlez:2013:CSL

[SH13a] Matthieu Sainlez and Georges Heyen. Comparison of supervised learning techniques for atmospheric pollutant monitoring in a Kraft pulp mill. Journal of Computational and Applied Mathematics, 246(??):329– 334, July 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712002610.

Schroter:2013:NMQ

[SH13b] A. Schröter and P. Heider. Numerical methods to quantify the model risk of basket default swaps. Journal of Computational and Applied Mathematics, 251(??):117–132, October 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713001507.

Shakhno:2014:CTS

[Sha14] S. M. Shakhno. Convergence of the two-step combined method and uniqueness of the solution of nonlinear operator equations. Journal of Computational and Applied Mathematics, 261(??):378–386, May 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713006481.

Shehu:2010:FPS

[She10] Y. Shehu. Fixed point solutions of generalized equilibrium problems for nonexpansive mappings. Journal of Computational and Applied Mathematics, 234(3): 892–898, June 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710000683.

Shadimetov:2013:OQF

[SHN13] Kh. M. Shadimetov, A. R.

Hayotov, and F. A. Nuraliev. On an optimal quadrature formula in Sobolev space L_{2}^{(m)}(0, 1). Journal of Computational and Applied Mathematics, 243(??):91–112, May 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042712005031.

Shen:2010:SPP

[SHZ10] Shu-Qian Shen, Ting-Zhu Huang, and Er-Jie Zhong. Spectral properties of primal-based penalty preconditioners for saddle point problems. Journal of Computational and Applied Mathematics, 233(9):2235– 2244, March 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709006876.

Si:2012:DCM

[SHZ12] Zhiyong Si, Yinnian He, and Tong Zhang. A defect-correction method for unsteady conductionconvection problems II: Time discretization. Journal of Computational and Applied Mathematics, 236 (9):2553–2573, March 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042711006285.

Sekiguchi:2011:DDS

[SI11] Masaki Sekiguchi and Emiko Ishiwata. Dynamics of a discretized SIR epidemic model with pulse vaccination and time delay. Journal of Computational and Applied Mathematics, 236(6):997–1008, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711002974.

Sidi:2014:BPN

[Sid14] Avram Sidi. Biorthogonal polynomials and numerical quadrature formulas for some finiterange integrals with symmetric weight functions. Journal of Computational and Applied Mathematics, 272(??):221–238, December 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714002490.

Sierra:2011:KEH

[Sie11] Cesar A. Gómez Sierra.

On a KdV equation with higher-order nonlinearity: Traveling wave solutions. Journal of Computational and Applied Mathematics, 235(17):5330–5332, July 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042711002858.

Sivakumar:2012:DTI

[SJA12] B. Sivakumar, R. Jayaraman, and G. Arivarignan. A discrete time inventory system with postponed demands. Journal of Computational and Applied Mathematics, 236(13):3073– 3083, July 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712000349.

Son:2014:NAC

[SK14] T. Q. Son and D. S. Kim.

A new approach to characterize the solution set of a pseudoconvex programming problem. Journal of Computational and Applied Mathematics, 261(??):333– 340, May 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713006237.

Susanto:2011:SBC

[SKAM11] H. Susanto, P. G. Kevrekidis, F. Kh. Abdullaev, and Boris A. Malomed. Symmetry breaking, coupling management, and localized modes in dual-core discrete nonlinear–Schrödinger lattices. Journal of Computational and Applied Mathematics, 235(13):3883–3888, May 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711000458.

Solin:2014:HCL

[SKK14] Pavel Solin, Lukas Korous, and Pavel Kus. Hermes2D, a C++ library for rapid development of adaptive hp-FEM and hp-DG solvers. Journal of Computational and Applied Mathematics, 270(??):152–165, November 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042714000831.

Sykora:2012:CHN

[SKKS12] Jan Sýkora, Tomás Krejcí, Jaroslav Kruis, and Michal Sejnoha. Computational homogenization of nonstationary transport processes in masonry structures. Journal of Computational and Applied Mathematics, 236(18):4745–4755, December 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712000982.

Sweilam:2011:NST

[SKN11] N. H. Sweilam, M. M.

Khader, and A. M. Nagy. Numerical solution of twosided space-fractional wave equation using finite difference method. Journal of Computational and Applied Mathematics, 235(8):2832– 2841, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042710006527.

Solin:2014:FIC

[SKS14] Pavel Solin, Pavel Karban, and Sascha Schnepp. Fourth International Conference on Finite Element Methods in Engineering and Sciences (FEMTEC 2013). Journal of Computational and Applied Mathematics, 270(??): 1, November 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714001459.

Solin:2010:P

[SKSS10] Pavel Solin, Dmitri Kuzmin, Daniil Svyatskiy, and Marian Slodicka. Preface. Journal of Computational and Applied Mathematics, 233 (12):3075–3076, April 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042709008693.

Sinescu:2012:VBE

[SL12] Vasile Sinescu and Pierre L’Ecuyer. Variance bounds and existence results for randomly shifted lattice rules. Journal of Computational and Applied Mathematics, 236(13):3296–3307, July 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271200101X.

Schwarz:2014:PBI

[SL14a] Diana Estévez Schwarz and René Lamour. Projector based integration of DAEs with the Taylor series method using automatic differentiation. Journal of Computational and Applied Mathematics, 262(??):62– 72, May 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713004731.

Suneja:2014:HOS

[SL14b] S. K. Suneja and Pooja Louhan. Higher-order symmetric duality under coneinvexity and other related concepts. Journal of Computational and Applied Mathematics, 255(??):825– 836, January 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713003476.

Slassi:2014:OFK

[Sla14] Mehdi Slassi. The optimal free knot spline approximation of stochastic differential equations with additive noise. Journal of Computational and Applied Mathematics, 261(??):62– 71, May 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713004895.

Su:2013:PHB

[SLD13] Huan Su, Wenxue Li, and Xiaohua Ding. Preservation of Hopf bifurcation for neutral delay-differential equations by θ-methods. Journal of Computational and Applied Mathematics, 248(??):76–87, August 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713000599.

Svacek:2014:NST

[SLK14] Petr Svácek, Petr Louda, and Karel Kozel. On numerical simulation of threedimensional flow problems by finite element and finite volume techniques. Journal of Computational and Applied Mathematics, 270(??):451–461, November 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713006924.

Siahaan:2011:LCA

[SLP11] Antony Siahaan, Choi-Hong Lai, and Koulis Pericleous. Local convergence of an adaptive scalar method and its application in a nonoverlapping domain decomposition scheme. Journal of Computational and Applied Mathematics, 235(17):5203– 5212, July 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711002573.

Schroder:2014:SCD

[SLW14] Dirk Schröder, Jens Lang, and Rüdiger Weiner. Stability and consistency of discrete adjoint implicit peer methods. Journal of Computational and Applied Mathematics, 262(??):73– 86, May 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003725.

Sun:2010:PGD

[SM10] Tongjun Sun and Keying Ma. Parallel Galerkin domain decomposition procedures for wave equation. Journal of Computational and Applied Mathematics, 233(8):1850–1865, February 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704270900644X.

Svard:2011:IES

[SM11] Magnus Svärd and Siddhartha Mishra. Implicitexplicit schemes for flow equations with stiff source terms. Journal of Computational and Applied Mathematics, 235(6):1564–1577, January 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271000453X.

Sawicki:2014:NHO

[SM14a] Bartosz Sawicki and Arkadiusz Miaskowski. Nonlinear higher-order transient solver for magnetic fluid hyperthermia. Journal of Computational and Applied Mathematics, 270(??):143– 151, November 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042714000843.

Sun:2014:DDP

[SM14b] Tongjun Sun and Keying Ma. Domain decomposition procedures combined with H^{1}–Galerkin mixed finite element method for parabolic equation. Journal of Computational and Applied Mathematics, 267 (??):33–48, September 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042714000648.

Saadvandi:2014:PDP

[SMD14] Maryam Saadvandi, Karl Meerbergen, and Wim Desmet. Parametric dominant pole algorithm for parametric model order reduction. Journal of Computational and Applied Mathematics, 259 (part A)(??): 259–280, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042713004603.

Sanchez-Moreno:2010:SLH

[SMDMY10] P. Sánchez-Moreno, J. S.

Dehesa, D. Manzano, and R. J. Yáñez. Spreading lengths of Hermite polynomials. Journal of Computational and Applied Mathematics, 233(9):2136–2148, March 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709006670.

Smietanski:2011:SQB

[Śmi11] Marek J. Śmietański. Some quadrature-based versions of the generalized Newton method for solving nonsmooth equations. Journal of Computational and Applied Mathematics, 235(17):5131– 5139, July 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711002378.

Sanchez-Moreno:2011:DSM

[SMMD11] P. Sánchez-Moreno, D. Manzano, and J. S. Dehesa. Direct spreading measures of Laguerre polynomials. Journal of Computational and Applied Mathematics, 235(5):1129–1140, January 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710004206.

Sophocleous:2011:SAM

[SOL11] C. Sophocleous, J. G. O’Hara, and P. G. L. Leach. Symmetry analysis of a model of stochastic volatility with time-dependent parameters. Journal of Computational and Applied Mathematics, 235(14):4158– 4164, May 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042711001397.

Solin:2012:PF

[Sol12] Pavel Solin. Preface for FEMTEC 2011. Journal of Computational and Applied Mathematics, 236(18):4581, December 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712002026.

Spalevic:2010:EBC

[SP10] Miodrag M. Spalević and Miroslav S. Pranić. Error bounds of certain Gaussian quadrature formulae. Journal of Computational and Applied Mathematics, 234(4):1049–1057, June 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709002519.

Stpiczynski:2011:SKB

[SP11]  Przemysław  Stpiczyński

and Joanna Potiopa. Solving a kind of boundaryvalue problem for ordinary differential equations using Fermi — the next generation CUDA computing architecture. Journal of Computational and Applied Mathematics, 236(3):384–393, September 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004237.

Spalevic:2012:EEA

[Spa12] Miodrag M. Spalević. Error estimates of anti-Gaussian quadrature formulae. Journal of Computational and Applied Mathematics, 236(15):3542–3555, September 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711001567.

Sanchez:2014:MMT

[SPC14] Eduardo J. Sanchez, Christopher P. Paolini, and Jose E. Castillo. The mimetic methods toolkit: an objectoriented API for mimetic finite differences. Journal of Computational and Applied Mathematics, 270(??):308– 322, November 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S037704271300719X.

Speleers:2011:MPB

[Spe11] Hendrik Speleers. On multivariate polynomials in Bernstein–Bézier form and tensor algebra. Journal of Computational and Applied Mathematics, 236(4):589– 599, September 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042711002263.

Safkhani:2014:CCA

[SPLHCB14] Masoumeh Safkhani, Pedro Peris-Lopez, Julio Cesar Hernandez-Castro, and Nasour Bagheri. Cryptanalysis of the Cho et al. protocol: a hashbased RFID tag mutual authentication protocol. Journal of Computational and Applied Mathematics, 259 (part B)(??):571–577, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713005281.

Sprengel:2014:DRP

[Spr14] Martin Sprengel. Domain robust preconditioning for a staggered grid discretization of the Stokes equations. Journal of Computational and Applied Mathematics, 255(??):468–480, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003014.

Sanchez-Reyes:2011:CCT

[SR11] J. Sánchez-Reyes. On the conditions for the coincidence of two cubic Bézier curves. Journal of Computational and Applied Mathematics, 236(6):1675–1677, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004705.

Sun:2013:NSE

[SR13] Tong Sun and David Rumsey. Numerical smoothness and error analysis for RKDG on the scalar nonlinear conservation laws. Journal of Computational and Applied Mathematics, 241(??):68– 83, March 15, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712004177.

Shubov:2010:MAC

[SRA10] Marianna A. Shubov and Miriam Rojas-Arenaza. Mathematical analysis of carbon nanotube model. Journal of Computational and Applied Mathematics, 234(6):1631– 1636, July 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042709004841.

Suciu:2013:CFE

[SRP^{+}13] N. Suciu, F. A. Radu, A. Prechtel, F. Brunner, and P. Knabner. A coupled finite element-global random walk approach to advection-dominated transport in porous media with random hydraulic conductivity. Journal of Computational and Applied Mathematics, 246(??):27– 37, July 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712002622.

Song:2014:CFS

[SRWZ14] Caiqin Song, Hongxing Rui, Xiaodong Wang, and Jianli Zhao. Closed-form solutions to the nonhomogeneous Yakubovich-transpose matrix equation. Journal of Computational and Applied Mathematics, 267(??):72– 81, September 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042714000557.

Santos:2014:INP

[SS14a] S. A. Santos and R. C. M.

Silva. An inexact and nonmonotone proximal method for smooth unconstrained minimization. Journal of Computational and Applied Mathematics, 269(??):86– 100, October 15, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042714001721.

Singh:2014:ACF

[SS14b] Uaday Singh and Shailesh Kumar Srivastava. Approximation of conjugate of functions belonging to weighted Lipschitz class W (L^{p}, ξ(t)) by Hausdorff means of conjugate Fourier series. Journal of Computational and Applied Mathematics, 259 (part B)(??):633–640, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713000964.

Srivastava:2014:TAP

[SS14c] Shailesh Kumar Srivastava and Uaday Singh. Trigonometric approximation of periodic functions belonging to Lip(ω(t), p)- class. Journal of Computational and Applied Mathematics, 270(??):223–230, November 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271400048X.

Stanimirovic:2014:CNA

[SS14d] Predrag S. Stanimirović and F. Soleymani. A class of numerical algorithms for computing outer inverses. Journal of Computational and Applied Mathematics, 263(??):236–245, June 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713007061.

Sun:2014:FEA

[SS14e] Yuzhou Sun and Pengtao Sun. Finite element approximation analysis for a steady state two-phase transport model of proton exchange membrane fuel cell. Journal of Computational and Applied Mathematics, 270(??):198–210, November 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714000776.

Suarez:2011:UIF

[SSF^{+}11] A. Suárez, H. Sarmiento, E. Flórez, M. D. García, and G. Montero. Updating incomplete factorization preconditioners for shifted linear systems arising in a wind model. Journal of Computational and Applied Mathematics, 235(8):2640–2646, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006321.

Strehl:2013:PPF

[SSK^{+}13] Robert Strehl, Andriy Sokolov, Dmitri Kuzmin, Dirk Horstmann, and Stefan Turek. A positivitypreserving finite element method for chemotaxis problems in 3D. Journal of Computational and Applied Mathematics, 239(??):290– 303, February 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712004141.

Saridakis:2012:ENS

[SSP12] Y. G. Saridakis, A. G.

Sifalakis, and E. P. Papadopoulou. Efficient numerical solution of the generalized Dirichlet–Neumann map for linear elliptic PDEs in regular polygon domains. Journal of Computational and Applied Mathematics, 236(9):2515–2528, March 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271100625X.

Scalia:2013:FBI

[SSP13] Antonio Scalia, Mezhlum A. Sumbatyan, and Vitaly Popuzin. A fast BIE iteration method for an arbitrary body in a flow of incompressible inviscid fluid. Journal of Computational and Applied Mathematics, 237(1):508–519, January 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712002555.

Soane:2010:OCR

[SSR10] Ana Maria Soane, Manil Suri, and Rouben Rostamian. The optimal convergence rate of a C^{1} finite element method for nonsmooth domains. Journal of Computational and Applied Mathematics, 233(10):2711– 2723, March 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709007614.

Setzer:2012:VMM

[SST12] S. Setzer, G. Steidl, and T. Teuber. On vector and matrix median computation. Journal of Computational and Applied Mathematics, 236(8):2200–2222, February 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711005218.

Salama:2014:SGP

[SSW14] Amgad Salama, Shuyu Sun, and Mary F. Wheeler. Solving global problem by considering multitude of local problems: Application to fluid flow in anisotropic porous media using the multipoint flux approximation. Journal of Computational and Applied Mathematics, 267(??):117–130, September 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714000375.

Stamm:2011:PEB

[Sta11] Benjamin Stamm. A posteriori estimates for the bubble stabilized discontinuous Galerkin method. Journal of Computational and Applied Mathematics, 235(15): 4309–4324, June 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042711001476.

Stefansson:2010:ABM

[Ste10a] Úlfar F. Stefánsson. Asymptotic behavior of Müntz– Christoffel functions at the endpoints. Journal of Computational and Applied Mathematics, 233(6):1601– 1606, January 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S037704270900168X.

Stepien:2010:SDM

[Ste10b] Lukasz T. Stepień. Some decomposition method for analytic solving of certain nonlinear partial differential equations in physics with applications. Journal of Computational and Applied Mathematics, 233(6):1607– 1611, January 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042709001691.

Stojakovic:2011:ISF

[Sto11a] Mila Stojaković. Imprecise set and fuzzy valued probability. Journal of Computational and Applied Mathematics, 235(16):4524–4531, June 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710000208.

Stojanovic:2011:NMS

[Sto11b] Mirjana Stojanović. Numerical method for solving diffusion-wave phenomena. Journal of Computational and Applied Mathematics, 235(10):3121–3137, March 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006606.

Strobach:2010:FQS

[Str10] Peter Strobach. The fast quartic solver. Journal of Computational and Applied Mathematics, 234 (10):3007–3024, September 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710002128.

Strobach:2011:SCP

[Str11] Peter Strobach. Solving cubics by polynomial fitting. Journal of Computational and Applied Mathematics, 235(9):3033–3052, March 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710006758.

Strobach:2012:FAR

[Str12] Peter Strobach. A fitting algorithm for real coefficient polynomial rooting. Journal of Computational and Applied Mathematics, 236(13): 3238–3255, July 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042712000842.

Sun:2014:DRI

[Sun14] Pengtao Sun. A Dirichlet/ Robin iteration-by-subdomain method for an anisotropic, nonisothermal two-phase transport model of PEM fuel cell with micro-porous layer. Journal of Computational and Applied Mathematics, 270(??):241–256, November 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042714000351.

Saito:2010:PFT

[SW10a] Naoki Saito and Yi Wang. The polynomial Fourier transform with minimum mean square error for noisy data. Journal of Computational and Applied Mathematics, 234(5):1586–1610, July 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710001378.

Shen:2010:CDS

[SW10b] Wanqiang Shen and Guozhao Wang. Changeable degree spline basis functions. Journal of Computational and Applied Mathematics, 234(8):2516–2529, August 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710001639.

Sidje:2011:EPI

[SW11] Roger B. Sidje and Nathan Winkles. Evaluation of the performance of inexact GM-RES. Journal of Computational and Applied Mathematics, 235(8):1956–1975, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710005376.

Schnepp:2012:ELS

[SW12a] Sascha M. Schnepp and Thomas Weiland. Efficient large scale electromagnetic simulations using dynamically adapted meshes with the discontinuous Galerkin method. Journal of Computational and Applied Mathematics, 236(18):4909–4924, December 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711006121.

Shi:2012:SVB

[SW12b] Xinghua Shi and Yimin Wei. A sharp version of Bauer–Fike’s theorem. Journal of Computational and Applied Mathematics, 236(13):3218–3227, July 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712000787.

Schumaker:2013:HIP

[SW13] Larry L. Schumaker and Lujun Wang. On Hermite interpolation with polynomial splines on T - meshes. Journal of Computational and Applied Mathematics, 240(??):42– 50, March 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712002282.

Shi:2012:GTV

[SWT12] Yuying Shi, Li-Lian Wang, and Xue-Cheng Tai. Geometry of total variation regularized L^{p}-model. Journal of Computational and Applied Mathematics, 236(8):2223– 2234, February 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704271100522X.

Shen:2013:ERC

[SWY13] Wanqiang Shen, Guozhao Wang, and Ping Yin. Explicit representations of changeable degree spline basis functions. Journal of Computational and Applied Mathematics, 238(??):39– 50, January 15, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712003391.

Shen:2014:AOM

[SWY14] Jie Shen, Li-Lian Wang, and Haijun Yu. Approximations by orthonormal mapped Chebyshev functions for higher-dimensional problems in unbounded domains. Journal of Computational and Applied Mathematics, 265(??):264–275, August 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713004792.

She:2011:CNB

[SXZ11] Zhikun She, Bican Xia, and Zhiming Zheng. Condition number based complexity estimate for solving polynomial systems. Journal of Computational and Applied Mathematics, 235(8):2670–2678, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006357.

Shang:2011:CSH

[SY11] Yufeng Shang and Bo Yu. A constraint shifting homotopy method for convex multi-objective programming. Journal of Computational and Applied Mathematics, 236(5):640–646, October 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711003979.

Serenbay:2011:CFI

[SYİ11] Sevilay Kirci Serenbay, Mine Menekşe Yilmaz, and Ertan İbikli. The convergence of a family of integral operators (in L_{p} space) with a positive kernel. Journal of Computational and Applied Mathematics, 235(16):4567– 4575, June 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710001640.

Seymen:2014:DOC

[SYK14] Z. Kanar Seymen, H. Yücel, and B. Karasözen. Distributed optimal control of time-dependent diffusionconvection-reaction equations using space-time discretization. Journal of Computational and Applied Mathematics, 261(??):146– 157, May 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713006250.

Sofronov:2010:NGT

[SZ10] I. L. Sofronov and N. A. Zaitsev. Numerical generation of transparent boundary conditions on the side surface of a vertical transverse isotropic layer. Journal of Computational and Applied Mathematics, 234(6):1732– 1738, July 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704270900497X.

Sun:2011:MSN

[SZ11] Zhe Sun and Jinping Zeng.

A monotone semismooth Newton type method for a class of complementarity problems. Journal of Computational and Applied Mathematics, 235(5):1261– 1274, January 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042710004504.

Szabo:2010:SSC

[Sza10] Tamás Szabó. Some sufficient conditions for the non-negativity preservation property in the discrete heat conduction model. Journal of Computational and Applied Mathematics, 235(2):478–490, November 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710003419.

Shi:2012:HLN

[SZJ12] Zhenwei Shi, Hongjuan Zhang, and Zhiguo Jiang. Hybrid linear and nonlinear complexity pursuit for blind source separation. Journal of Computational and Applied Mathematics, 236(14):3434–3444, August 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712001422.

Song:2011:IBV

[SZS11] Guangxing Song, Yian Zhao, and Xiaoyan Sun. Integral boundary value problems for first order impulsive integro-differential equations of mixed type. Journal of Computational and Applied Mathematics, 235(9):2928–2935, March 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042710006576.

Trinh:2010:USS

[TA10] Philippe H. Trinh and David E. Amundsen. Unifying the steady state resonant solutions of the periodically forced KdVB, mKdVB, and eKdVB equations. Journal of Computational and Applied Mathematics, 234(6):1788–1795, July 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709005044.

Thierry:2013:SCN

[TA13] Bertrand Thierry and Xavier Antoine. Spectral and condition number estimates of the acoustic single-layer operator for low-frequency multiple scattering in dense media. Journal of Computational and Applied Mathematics, 239(??):380–395, February 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003664.

Tang:2010:ABS

[Tan10] Xiaosong Tang. Asymptotic behavior of solutions of second-order nonlinear delay differential equations with impulses. Journal of Computational and Applied Mathematics, 233(9):2105– 2111, March 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042709003914.

Tanaka:2014:HOI

[Tan14a] Hideyuki Tanaka. Higherorder interpolated lattice schemes for multidimensional option pricing problems. Journal of Computational and Applied Mathematics, 255(??):313–333, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713002914.

Tanaka:2014:ECN

[Tan14b] Ken’ichiro Tanaka. Error control of a numerical formula for the Fourier transform by Ooura’s continuous Euler transform and fractional FFT. Journal of Computational and Applied Mathematics, 266(??):73– 86, August 15, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704271400020X.

Tariboon:2010:BVP

[Tar10] Jessada Tariboon. Boundary value problems for first order functional differential equations with impulsive integral conditions. Journal of Computational and Applied Mathematics, 234(8):2411– 2419, August 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710001482.

Tausch:2013:CFB

[Tau13] Johannes Tausch. Computing Floquet–Bloch modes in biperiodic slabs with boundary elements. Journal of Computational and Applied Mathematics, 254(??):192– 203, December 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042713001155.

Tor:2014:ACM

[TBK14] Ali Hakan Tor, Adil Bagirov, and Bülent Karasözen. Aggregate codifferential method for nonsmooth DC optimization. Journal of Computational and Applied Mathematics, 259 (part B) (??):851–867, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042713004159.

Tian:2011:AFE

[TCD11] Li Tian, Falai Chen, and Qiang Du. Adaptive finite element methods for elliptic equations over hierarchical T -meshes. Journal of Computational and Applied Mathematics, 236(5):878– 891, October 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711002639.

Tezcan:2014:IDA

[Tez14] Cihangir Tezcan. Improbable differential attacks on present using undisturbed bits. Journal of Computational and Applied Mathematics, 259 (part B)(??): 503–511, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704271300321X.

Tang:2011:WSS

[TJL11] Chun-Ming Tang, Jin-Bao Jian, and Guo-Yin Li. A working set SQCQP algorithm with simple nonmonotone penalty parameters. Journal of Computational and Applied Mathematics, 236(6):1382–1398, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004729.

Trommler:2012:FEA

[TKW12] Jens Trommler, Stephan Koch, and Thomas Weiland. A finite-element approach in order to avoid ill-conditioning in thinsheet problems in frequency domain — application to magneto-quasistatics. Journal of Computational and Applied Mathematics, 236(18):4671–4680, December 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712001598.

Teng:2013:CAL

[TL13] Zhongming Teng and Ren-Cang Li. Convergence analysis of Lanczos-type methods for the linear response eigenvalue problem. Journal of Computational and Applied Mathematics, 247(??): 17–33, August 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713000216.

Tongsomporn:2010:SGT

[TLHU10] Janyarak Tongsomporn, Vichian Laohakosol, Charinthip Hengkrawit, and Patanee Udomkavanich. Stability of a generalized trigonometric functional equation. Journal of Computational and Applied Mathematics, 234(5):1448–1457, July 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710001032.

Tryoen:2010:RSE

[TLNE10] J. Tryoen, O. Le Maître, M. Ndjinga, and A. Ern. Roe solver with entropy corrector for uncertain hyperbolic systems. Journal of Computational and Applied Mathematics, 235(2):491– 506, November 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704271000333X.

Taghizadeh:2011:FIM

[TM11] N. Taghizadeh and M. Mirzazadeh. The first integral method to some complex nonlinear partial differential equations. Journal of Computational and Applied Mathematics, 235(16):4871– 4877, June 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711001099.

Taleizadeh:2010:MSM

[TNN10] Ata Allah Taleizadeh, Seyed Taghi Akhavan Niaki, and Amir Abbas Najafi. Multiproduct single-machine production system with stochastic scrapped production rate, partial backordering and service level constraint. Journal of Computational and Applied Mathematics, 233(8):1834–1849, February 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709006438.

Toledo:2014:GOU

[TOF14] C. F. M. Toledo, L. Oliveira, and P. M. França. Global optimization using a genetic algorithm with hierarchically structured population. Journal of Computational and Applied Mathematics, 261(??):341–351, May 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713006274.

Torres:2014:EPS

[Tor14] Fernando Jiménez Torres.

Estimation of parameters of the shifted Gompertz distribution using least squares, maximum likelihood and moments methods. Journal of Computational and Applied Mathematics, 255(??):867–877, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003488.

Thukral:2010:FTP

[TP10] R. Thukral and M. S. Petković. A family of threepoint methods of optimal order for solving nonlinear equations. Journal of Computational and Applied Mathematics, 233(9):2278– 2284, March 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709006906.

Toledo:2011:NAL

[TR11] R. C. Leal Toledo and V. Ruas. Numerical analysis of a least-squares finite element method for the time-dependent advection– diffusion equation. Journal of Computational and Applied Mathematics, 235(12): 3615–3631, April 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0377042711001105.

Tanio:2010:GCI

[TS10] Masaaki Tanio and Masaaki Sugihara. GBi–CGSTAB(s, L): IDR(s) with higher-order stabilization polynomials. Journal of Computational and Applied Mathematics, 235(3):765–784, December 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710003821.

Tocino:2012:MSS

[TS12] A. Tocino and M. J. Senosiain. Mean-square stability analysis of numerical schemes for stochastic differential systems. Journal of Computational and Applied Mathematics, 236(10):2660– 2672, April 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712000039.

Tocino:2014:AMS

[TS14] A. Tocino and M. J. Senosiain. Asymptotic meansquare stability of twostep Maruyama schemes for stochastic differential equations. Journal of Computational and Applied Mathematics, 260(??):337– 348, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005463.

Tian:2010:CES

[TSC10] Yuan Tian, Kaibiao Sun, and Lansun Chen. Comment on “Existence and stability of periodic solution of a Lotka–Volterra predator–prey model with state dependent impulsive effects” [J. Comput. Appl. Math. 224 (2009) 544–555]. Journal of Computational and Applied Mathematics, 234(10):2916–2923, September 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710001986.

See [NPTH09].

Tunca:2011:CVM

[TT11] Gülen Bascanbaz Tunca and Yalcin Tuncer. On a Chlodovsky variant of a multivariate beta operator. Journal of Computational and Applied Mathematics, 235(16):4816–4824, June 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006151.

Tacheny:2012:MPA

[TT12] N. Tacheny and C. Troestler.

A mountain pass algorithm with projector. Journal of Computational and Applied Mathematics, 236(7):2025– 2036, January 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711005814.

Torres:2014:CWE

[TT14a] Edmanuel Torres and Rafael Torres. Chirp-wave expansion of the electron wavefunctions in atoms. Journal of Computational and Applied Mathematics, 263(??): 218–224, June 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713006894.

Tuncel:2014:CRC

[TT14b] Altan Tuncel and Fatih Tank. Computational results on the compound binomial risk model with nonhomogeneous claim occurrences. Journal of Computational and Applied Mathematics, 263(??):69– 77, June 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713006742.

Thakoor:2014:EHA

[TTB14] Nawdha Thakoor, Désiré Yannick Tangman, and Muddun Bhuruth. Efficient and high accuracy pricing of barrier options under the CEV diffusion. Journal of Computational and Applied Mathematics, 259 (part A)(??):182–193, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713002720.

Turk:2014:FES

[TTSB14] Ö. Türk, M. Tezer-Sezgin, and Canan Bozkaya. Finite element study of biomagnetic fluid flow in a symmetrically stenosed channel. Journal of Computational and Applied Mathematics, 259 (part B)(??):760–770, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271300335X.

Turhan:2012:ADE

[Tur12] Nejla Ozkaya Turhan. On asymptotic distributions of exceedance statistics. Journal of Computational and Applied Mathematics, 236 (9):2259–2267, March 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704271100584X.

Temocin:2014:OCS

[TW14] Büsra Zeynep Temoçin and Gerhard-Wilhelm Weber. Optimal control of stochastic hybrid system with jumps: a numerical approximation. Journal of Computational and Applied Mathematics, 259 (part B)(??): 443–451, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042713005657.

Tian:2011:HOE

[TY11] Zhen F. Tian and P. X.

Yu. A high-order exponential scheme for solving 1D unsteady convection– diffusion equations. Journal of Computational and Applied Mathematics, 235 (8):2477–2491, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042710006175.

Tan:2014:ODS

[TYCL14] Jiyang Tan, Pingtian Yuan, Yangjin Cheng, and Ziqiang Li. An optimal dividend strategy in the discrete Sparre Andersen model with bounded dividend rates. Journal of Computational and Applied Mathematics, 258(??):1–16, March 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271300441X.

Tyrtyshnikov:2010:TRI

[Tyr10] Eugene Tyrtyshnikov. Tensor ranks for the inversion of tensor-product binomials. Journal of Computational and Applied Mathematics, 234(11):3170–3174, October 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710000774.

Urena:2012:NAG

[UBSG12] Francisco Ureña, Juan José Benito, Eduardo Salete, and Luis Gavete. A note on the application of the generalized finite difference method to seismic wave propagation in 2D. Journal of Computational and Applied Mathematics, 236(12):3016– 3025, June 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711001828.

Uyan:2014:CBF

[UÇD14] Erdener Uyan, Çagdas Çalik, and Ali Doganaksoy. Counting Boolean functions with specified values in their Walsh spectrum. Journal of Computational and Applied Mathematics, 259 (part B)(??):522–528, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003336.

Siraj-ul-Islam:2014:IMB

[uIAAF14] Siraj ul Islam, Imran Aziz, and A. S. Al-Fhaid. An improved method based on Haar wavelets for numerical solution of nonlinear integral and integrodifferential equations of first and higher orders. Journal of Computational and Applied Mathematics, 260(??): 449–469, April 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713005682.

Usevich:2014:VPA

[UM14] Konstantin Usevich and Ivan Markovsky. Variable projection for affinely structured low-rank approximation in weighted 2-norms. Journal of Computational and Applied Mathematics, 272(??):430–448, December 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713002203.

Utyuzhnikov:2010:NSP

[Uty10a] S. V. Utyuzhnikov. Nonstationary problem of active sound control in bounded domains. Journal of Computational and Applied Mathematics, 234(6):1725– 1731, July 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042709004968.

Utyuzhnikov:2010:NPA

[Uty10b] S. V. Utyuzhnikov. Nonlinear problem of active sound control. Journal of Computational and Applied Mathematics, 234(1):215– 223, May 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709008255.

VanDeun:2010:CNB

[Van10] Joris Van Deun. Computing near-best fixed pole rational interpolants. Journal of Computational and Applied Mathematics, 235(4):1077– 1084, December 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0377042710001093.

VanAssche:2011:OPS

[Van11a] Walter Van Assche. Orthogonal polynomials, special functions and applications (OPSFA-10). Journal of Computational and Applied Mathematics, 236(1): 1–6, August 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711002251.

VanDamme:2011:GFS

[Van11b] Geert Van Damme. A generic framework for stochastic Loss-Given-Default. Journal of Computational and Applied Mathematics, 235 (8):2523–2550, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042710006229.

Vigo-Aguiar:2011:NOS

[VAR11] J. Vigo-Aguiar and Higinio Ramos. A numerical ODE solver that preserves the fixed points and their stability. Journal of Computational and Applied Mathematics, 235(7):1856–1867, February 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710003900.

Vigo-Aguiar:2011:ACM

[VAV11] J. Vigo-Aguiar and Guido Vanden Berghe. Advances in computational and mathematical methods in science and engineering. Journal of Computational and Applied Mathematics, 235(7):1745, February 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710005820.

Vardar-Acar:2013:CSI

[VAZS13] Ceren Vardar-Acar, Craig L. Zirbel, and Gábor J. Székely. On the correlation of the supremum and the infimum and of maximum gain and maximum loss of Brownian motion with drift. Journal of Computational and Applied Mathematics, 248(??):61–75, August 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713000290.

Vorel:2014:NSD

[VB14] Jan Vorel and W. P. Boshoff. Numerical simulation of ductile fiberreinforced cement-based composite. Journal of Computational and Applied Mathematics, 270(??):433– 442, November 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713006948.

Vigneron:2010:EFB

[VBP^{+}10] Lara M. Vigneron, Romain C. Boman, Jean-Philippe Ponthot, Pierre A. Robe, Simon K. Warfield, and Jacques G. Verly. Enhanced FEM-based modeling of brain shift deformation in image-guided neurosurgery. Journal of Computational and Applied Mathematics, 234(7):2046–2053, August 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709005391.

Viswanathan:2014:PCT

[VCA14] P. Viswanathan, A. K. B.

Chand, and R. P. Agarwal. Preserving convexity through rational cubic spline fractal interpolation function. Journal of Computational and Applied Mathematics, 263(??):262– 276, June 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713006560.

VanWeert:2011:CAG

[VDG11] Koen Van Weert, Jan Dhaene, and Marc Goovaerts. Comonotonic approximations for a generalized provisioning problem with application to optimal portfolio selection. Journal of Computational and Applied Mathematics, 235(10):3245– 3256, March 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042711000239.

Vueghs:2010:UGF

[VdKPB10] P. Vueghs, H. P. de Koning, O. Pin, and P. Beckers. Use of geometry in finite element thermal radiation combined with ray tracing. Journal of Computational and Applied Mathematics, 234(7):2319–2326, August 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704270900569X.

vanDoorn:2009:REZ

[vDvFZ09] Erik A. van Doorn, Nicky D. van Foreest, and Alexander I. Zeifman. Representations for the extreme zeros of orthogonal polynomials. Journal of Computational and Applied Mathematics, 233(3):847–851, December 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709001344.

See corrigendum [vDvFZ13].

vanDoorn:2013:CRE

[vDvFZ13] Erik A. van Doorn, Nicky D.

van Foreest, and Alexander I. Zeifman. Corrigendum to “Representations for the extreme zeros of orthogonal polynomials” [J. Comput. Appl. Math. 233 (2009) 847–851]. Journal of Computational and Applied Mathematics, 244(??):155– 156, May 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712005225.

See [vDvFZ09].

Velarde:2010:PSA

[Vel10] Manuel G. Velarde. From polaron to solectron: the addition of nonlinear elasticity to quantum mechanics and its possible effect upon electric transport. Journal of Computational and Applied Mathematics, 233(6):1432–1445, January 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709001459.

Veling:2011:GIG

[Vel11] E. J. M. Veling. The generalized incomplete gamma function as sum over modified Bessel functions of the first kind. Journal of Computational and Applied Mathematics, 235(14):4107– 4116, May 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042711001245.

Venturino:2010:EMD

[Ven10] Ezio Venturino. Ecoepidemic models with disease incubation and selective hunting. Journal of Computational and Applied Mathematics, 234(9):2883–2901, September 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710000403.

vanGroesen:2010:AMU

[vGALL10] E. van Groesen, Andonowati, L. She Liam, and I. Lakhturov. Accurate modelling of uni-directional surface waves. Journal of Computational and Applied Mathematics, 234(6):1747– 1756, July 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042709004993.

Vitrih:2010:LSP

[Vit10] Vito Vitrih. Lattices on simplicial partitions which are not simply connected. Journal of Computational and Applied Mathematics, 235(1):154–164, November 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710003006.

Vlcek:2013:GLM

[VL13] J. Vlcek and L. Luksan. Generalizations of the limited-memory BFGS method based on the quasi-product form of update. Journal of Computational and Applied Mathematics, 241(??):116–129, March 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712004001.

Vazquez-Mendez:2013:IEI

[VMAVGCM13] M. E. Vázquez-Méndez, L. J. Alvarez-Vázquez, N. García-Chan, and A. Martínez. Improving the environmental impact of wastewater discharges with a specific simulation-optimization software. Journal of Computational and Applied Mathematics, 246(??):320–328, July 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712002816.

Vaissmoradi:2010:SAN

[VMMM10] N. Vaissmoradi, A. Malek, and S. H. Momeni-Masuleh. L^{2}-stability analysis of novel ETD scheme for Kuramoto–Sivashinsky equations. Journal of Computational and Applied Mathematics, 234(8):2493–2500, August 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710001603.

Vavalis:2012:FES

[VMS12] Manolis Vavalis, Mo Mu, and Giorgos Sarailidis. Finite element simulations of window Josephson junctions. Journal of Computational and Applied Mathematics, 236(13):3186–3197, July 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271200074X.

VanBockstal:2013:DUD

[VS13a] K. Van Bockstal and M. Slodicka. Determination of an unknown diffusion coefficient in a semilinear parabolic problem. Journal of Computational and Applied Mathematics, 246(??): 104–112, July 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712002294.

Vrabel:2013:NDP

[VS13b] Vladimír Vrábel’ and Marián Slodicka. Nonlinear diffusion problem with dynamical boundary value. Journal of Computational and Applied Mathematics, 246(??): 94–103, July 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712004992.

Vuong:2014:FEM

[VS14] Julia Vuong and Bernd Simeon. On finite element method-flux corrected transport stabilization for advection–diffusion problems in a partial differential-algebraic framework. Journal of Computational and Applied Mathematics, 262(??):115–126, May 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005256.

Vorel:2012:MSC

[VSB12] Jan Vorel, Vít Smilauer, and Zdenek Bittnar. Multiscale simulations of concrete mechanical tests. Journal of Computational and Applied Mathematics, 236(18):4882– 4892, December 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712000106.

Vulanovic:2014:QBL

[VT14] Relja Vulanović and Ljiljana Teofanov. On the quasilinear boundary-layer problem and its numerical solution. Journal of Computational and Applied Mathematics, 268(??):56–67, October 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714001265.

Valero:2011:FRS

[VTLF11] E. Valero, M. Torrealba, L. Lacasa, and F. Fraysse. Fast resolution of a single factor Heath–Jarrow– Morton model with stochastic volatility. Journal of Computational and Applied Mathematics, 236(6):1637– 1655, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042711005061.

VanDaele:2014:P

[VV14] M. Van Daele and S. Vandewalle. Preface. Journal of Computational and Applied Mathematics, 259 (part A) (??):1, March 15, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713004718.

Vandekerckhove:2014:MDH

[VVD^{+}14]  Steven  Vandekerckhove,

Bart Vandewoestyne, Herbert De Gersem, Koen Van Den Abeele, and Stefan Vandewalle. Mimetic discretization and higher order time integration for acoustic, electromagnetic and elastodynamic wave propagation. Journal of Computational and Applied Mathematics, 259 (part A)(??): 65–76, March 15, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713001064.

Vandebril:2010:MSS

[VVM10] Raf Vandebril, Marc Van Barel, and Nicola Mastronardi. A multiple shift QR-step for structured rank matrices. Journal of Computational and Applied Mathematics, 233(5):1326– 1344, January 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042709001757.

VanBarel:2010:USM

[VVV10] Marc Van Barel, Yvette Vanberghen, and Paul Van Dooren. Using semiseparable matrices to compute the SVD of a general matrix product/quotient. Journal of Computational and Applied Mathematics, 234(11):3175–3180, October 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710000786.

Weber:2014:SAM

[WA14] Steffen Weber and Martin Arnold. Semi-analytical methods for singularly perturbed multibody system models. Journal of Computational and Applied Mathematics, 262(??):322–332, May 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713003841.

Wang:2010:BVP

[Wan10]  Guotao Wang.  Bound-

ary value problems for systems of nonlinear integrodifferential equations with deviating arguments. Journal of Computational and Applied Mathematics, 234 (5):1356–1363, July 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042710000130.

Wang:2011:EOT

[Wan11a]  Guodong Wang.  An

Engquist–Osher type finite difference scheme with a discontinuous flux function in space. Journal of Computational and Applied Mathematics, 235(17):4966–4977, July 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711002184.

Wang:2011:SWR

[Wan11b]  Jin-Ru Wang.  Shan-

non wavelet regularization methods for a backward heat equation. Journal of Computational and Applied Mathematics, 235(9):3079– 3085, March 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711000021.

Wang:2011:RNAa

[Wan11c] Lijuan Wang. Retraction notice to: Asymptotic behavior of solutions to a system of differential equations with state-dependent delays CAM 228 (2009) pp. 226–230. Journal of Computational and Applied Mathematics, 235(11):3366, April 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711000896.

Wang:2011:MNE

[Wan11d]  Tingchun Wang.  Max-

imum norm error bound of a linearized difference scheme for a coupled nonlinear Schrödinger equations. Journal of Computational and Applied Mathematics, 235(14):4237–4250, May 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271100149X.

Wang:2011:EEC

[Wan11e]  Yuan-Ming Wang.  Er-

ror and extrapolation of a compact LOD method for parabolic differential equations. Journal of Computational and Applied Mathematics, 235(5):1367–1382, January 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710004735.

Wang:2011:MAM

[Wan11f] Yuan-Ming Wang. A modified accelerated monotone iterative method for finite difference reaction– diffusion-convection equations. Journal of Computational and Applied Mathematics, 235(12):3646–3660, April 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711001130.

Wang:2012:MIT

[Wan12a] Guotao Wang. Monotone iterative technique for boundary value problems of a nonlinear fractional differential equation with deviating arguments. Journal of Computational and Applied Mathematics, 236(9):2425– 2430, March 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271100598X.

Wang:2012:MPT

[Wan12b]  Zewen Wang.  Multi-

parameter Tikhonov regularization and model function approach to the damped Morozov principle for choosing regularization parameters. Journal of Computational and Applied Mathematics, 236(7):1815–1832, January 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711005577.

Wang:2013:LTB

[Wan13] Wansheng Wang. Longtime behavior of the twogrid finite element method for fully discrete semilinear evolution equations with positive memory. Journal of Computational and Applied Mathematics, 250(??):161– 174, October 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042713001131.

Wadbro:2010:HCM

[WB10] Eddie Wadbro and Martin Berggren. High contrast microwave tomography using topology optimization techniques. Journal of Computational and Applied Mathematics, 234(6):1773–1780, July 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709005020.

Wang:2010:PSC

[WC10] Jun-Wei Wang and Ai-Min Chen. Partial synchronization in coupled chemical chaotic oscillators. Journal of Computational and Applied Mathematics, 233(8):1897–1904, February 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709006487.

Wang:2011:SGC

[WC11] Y.-S. Wang and C.-S.

Chien. A spectral-Galerkin continuation method using Chebyshev polynomials for the numerical solutions of the Gross–Pitaevskii equation. Journal of Computational and Applied Mathematics, 235(8):2740–2757, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006412.

Wu:2014:DMS

[WC14] Tingting Wu and Zhongying Chen. A dispersion minimizing subgridding finite difference scheme for the Helmholtz equation with PML. Journal of Computational and Applied Mathematics, 267(??):82– 95, September 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042714000594.

Wang:2011:UAM

[WCH11] Weichung Wang, Ray-Bing Chen, and Chia-Lung Hsu. Using adaptive multiaccurate function evaluations in a surrogate-assisted method for computer experiments. Journal of Computational and Applied Mathematics, 235(10):3151–3162, March 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710006710.

Wang:2010:RSC

[WCW10] Yingwei Wang, Suqin Chen, and Xionghua Wu. A rational spectral collocation method for solving a class of parameterized singular perturbation problems. Journal of Computational and Applied Mathematics, 233(10):2652–2660, March 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709007420.

Wu:2011:LES

[WCZ11] Y. Y. Wu, C. K. Chan, and L. X. Zhou. Large eddy simulation of an ethyleneair turbulent premixed V - flame. Journal of Computational and Applied Mathematics, 235(13):3768–3774, May 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711000343.

Wang:2014:FSM

[WD14] Hong Wang and Ning Du.

Fast solution methods for space-fractional diffusion equations. Journal of Computational and Applied Mathematics, 255(??):376– 383, January 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713002987.

Walther:2012:SIM

[WDG^{+}12] M. Walther, J.-O. Delfs, J. Grundmann, O. Kolditz, and R. Liedl. Saltwater intrusion modeling: Verification and application to an agricultural coastal arid region in oman. Journal of Computational and Applied Mathematics, 236(18):4798– 4809, December 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712000659.

Wu:2010:SCM

[WDZ10] Jiming Wu, Zihuan Dai, and Xiaoping Zhang. The superconvergence of the composite midpoint rule for the finite-part integral.

Journal of Computational and Applied Mathematics, 233(8):1954–1968, February 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709006542.

Wei:2011:CAS

[Wei11] Mao Wei. Convergence analysis of semi-implicit Euler methods for solving stochastic equations with variable delays and random jump magnitudes. Journal of Computational and Applied Mathematics, 235(8):2569–2580, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006254.

Wang:2010:NPG

[WF10] Lin-Lin Wang and Yong-Hong Fan. Note on permanence and global stability in delayed ratio-dependent predator–prey models with monotonic functional responses. Journal of Computational and Applied Mathematics, 234(2):477–487, May 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709008565.

Wang:2010:NISa

[WG10a] Shenghua Wang and Baohua Guo. New iterative scheme with nonexpansive mappings for equilibrium problems and variational inequality problems in Hilbert spaces. Journal of Computational and Applied Mathematics, 233(10):2620–2630, March 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709007390.

Wang:2010:SBF

[WG10b] Tongke Wang and Yuesheng Gu. Superconvergent biquadratic finite volume element method for twodimensional Poisson’s equations. Journal of Computational and Applied Mathematics, 234(2):447–460, May 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704270900853X.

Wang:2014:MSA

[WGZL14] Ren-Hong Wang, Qing-Jie Guo, Chun-Gang Zhu, and Chun-Jing Li. Multivariate spline approximation of the signed distance function. Journal of Computational and Applied Mathematics, 265(??):276–289, August 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713004986.

Winkler:2010:NLS

[WH10] Joab R. Winkler and Madina Hasan. A non-linear structure preserving matrix method for the low rank approximation of the Sylvester resultant matrix. Journal of Computational and Applied Mathematics, 234(12):3226–3242, October 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710002104.

Wu:2012:SAB

[WH12a] Chi-Ye Wu and Ting-Zhu Huang. Stability analysis of block LU factorization for complex symmetric block tridiagonal matrices. Journal of Computational and Applied Mathematics, 236(7):2037–2046, January 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711005693.

Wu:2012:SBF

[WH12b] Chi-Ye Wu and Ting-Zhu Huang. Stability of block LU factorization for block tridiagonal block H-matrices. Journal of Computational and Applied Mathematics, 236(10):2673– 2684, April 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712000040.

Winkler:2013:INL

[WH13] Joab R. Winkler and Madina Hasan. An improved non-linear method for the computation of a structured low rank approximation of the Sylvester resultant matrix. Journal of Computational and Applied Mathematics, 237(1):253– 268, January 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712003135.

Wang:2010:SET

[WHL10] Xiang Wang, Bingsheng He, and Li-Zhi Liao. Steplengths in the extragradient type methods. Journal of Computational and Applied Mathematics, 233(11):2925–2939, April 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709007845.

Wu:2013:MBP

[WHL13] Shi-Liang Wu, Ting-Zhu Huang, and Cui-Xia Li. Modified block preconditioners for the discretized time-harmonic Maxwell equations in mixed form. Journal of Computational and Applied Mathematics, 237(1): 419–431, January 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042712002464.

Wang:2011:IIR

[WHLC11] Linjun Wang, Xu Han, Jie Liu, and Jiujiu Chen. An improved iteration regularization method and application to reconstruction of dynamic loads on a plate. Journal of Computational and Applied Mathematics, 235(14):4083–4094, May 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711001221.

Wihler:2011:ASB

[Wih11] Thomas P. Wihler. An hpadaptive strategy based on continuous Sobolev embeddings. Journal of Computational and Applied Mathematics, 235(8):2731–2739, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006400.

Wikramaratna:2010:TEC

[Wik10] Roy S. Wikramaratna. Theoretical and empirical convergence results for additive congruential random number generators. Journal of Computational and Applied Mathematics, 233(9):2302– 2311, March 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709006931.

Willems:2013:RMS

[Wil13] J. Willems. Robust multilevel solvers for highcontrast anisotropic multiscale problems. Journal of Computational and Applied Mathematics, 251(??):47– 60, October 15, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713001386.

Winkler:2014:SMM

[Win14] Joab R. Winkler. Structured matrix methods for the computation of multiple roots of a polynomial. Journal of Computational and Applied Mathematics, 272(??):449–467, December 15, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042713004445.

Wei:2014:FIH

[WJJ14] Zhangping Wei, Byunghyun Jang, and Yafei Jia. A fast and interactive heat conduction simulator on GPUs. Journal of Computational and Applied Mathematics, 270(??):496–505, November 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713006638.

Wu:2010:CAP

[WK10] Chia-Huang Wu and Jau-Chuan Ke. Computational algorithm and parameter optimization for a multiserver system with unreliable servers and impatient customers. Journal of Computational and Applied Mathematics, 235(3):547– 562, December 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042710003493.

Weiner:2014:LGE

[WK14] R. Weiner and G. Yu. Kulikov. Local and global error estimation and control within explicit two-step peer triples. Journal of Computational and Applied Mathematics, 262(??):261– 270, May 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003853.

Kim:2012:SPA

[wKK12] Tae wan Kim and Boris Kvasov. A shape-preserving approximation by weighted cubic splines. Journal of Computational and Applied Mathematics, 236(17):4383– 4397, November 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712001616.

Wang:2010:SSF

[WL10a] Peng Wang and Yong Li.

Split-step forward methods for stochastic differential equations. Journal of Computational and Applied Mathematics, 233(10):2641– 2651, March 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709007419.

Wang:2010:NEO

[WL10b] Xia Wang and Liping Liu.

New eighth-order iterative methods for solving nonlinear equations. Journal of Computational and Applied Mathematics, 234(5):1611– 1620, July 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710001433.

Wang:2010:EEN

[WL10c] Xing Tao Wang and Yuan Min Li. On equations that are equivalent to the nonlinear matrix equation X + A^{∗}X^{−α}A = Q. Journal of Computational and Applied Mathematics, 234(8):2441– 2449, August 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042710001457.

Wozny:2010:CMD

[WL10d] Pawel Woźny and Stanisław Lewanowicz. Constrained multi-degree reduction of triangular Bézier surfaces using dual Bernstein polynomials. Journal of Computational and Applied Mathematics, 235(3):785–804, December 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710003912.

Wang:2011:EMP

[WL11a] Shan Wang and Zhiping Li. Evaluation of mechanical parameters of an elastic thin film system by modeling and numerical simulation of telephone cord buckles. Journal of Computational and Applied Mathematics, 236(5):860–866, October 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711002652.

Winkler:2011:CDA

[WL11b] Joab R. Winkler and Xin Lao. The calculation of the degree of an approximate greatest common divisor of two polynomials. Journal of Computational and Applied Mathematics, 235(6):1587– 1603, January 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710004838.

Wu:2014:EEI

[WL14a] Shi-Liang Wu and Cui-Xia Li. Eigenvalue estimates of an indefinite block triangular preconditioner for saddle point problems. Journal of Computational and Applied Mathematics, 260(??):349– 355, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005530.

Wu:2014:SIM

[WL14b] Shi-Liang Wu and Cui-Xia Li. A splitting iterative method for the discrete dynamic linear systems. Journal of Computational and Applied Mathematics, 267(??):49– 60, September 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042714000673.

Wang:2014:INM

[WLD14] Xiang Wang, Wen-Wei Li, and Lin Dai. On inexact Newton methods based on doubling iteration scheme for symmetric algebraic Riccati equations. Journal of Computational and Applied Mathematics, 260(??):364– 374, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713005293.

Winkler:2012:CMR

[WLH12] Joab R. Winkler, Xin Lao, and Madina Hasan. The computation of multiple roots of a polynomial. Journal of Computational and Applied Mathematics, 236(14):3478–3497, August 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712000751.

Wang:2011:FEM

[WLL^{+}11] Ruomei Wang, Yu Liu, Xiaonan Luo, Yi Li, and Shuai Ji. A finite-element mechanical contact model based on Mindlin–Reissner shell theory for a threedimensional human body and garment. Journal of Computational and Applied Mathematics, 236(5):867– 877, October 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711002640.

Wang:2011:NFF

[WLWH11] Chengjun Wang, Ronghua Luo, Kun Wu, and Boshun Han. A new filled function method for an unconstrained nonlinear equation. Journal of Computational and Applied Mathematics, 235(6):1689–1699, January 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710005017.

Wang:2011:MSH

[WLZ11] Ren-Hong Wang, Mian Li, and Chun-Gang Zhu. Multivariate splines and hyperplane arrangements. Journal of Computational and Applied Mathematics, 236(5):775–781, October 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711002913.

Wang:2014:NFI

[WLZ14] Guotao Wang, Sanyang Liu, and Lihong Zhang. Neutral fractional integrodifferential equation with nonlinear term depending on lower order derivative. Journal of Computational and Applied Mathematics, 260(??):167–172, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713005062.

Walton:2012:FGP

[WM12a] D. J. Walton and D. S. Meek. A further generalisation of the planar cubic Bézier spiral. Journal of Computational and Applied Mathematics, 236(11):2869– 2882, May 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712000350.

Wu:2012:RCA

[WM12b] Zhen Wu and Wim Michiels. Reliably computing all characteristic roots of delay differential equations in a given right half plane using a spectral method. Journal of Computational and Applied Mathematics, 236(9):2499–2514, March 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711006236.

Wang:2014:MWG

[WMGM14] X. Wang, N. S. Malluwawadu, F. Gao, and T. C. McMillan. A modified weak Galerkin finite element method. Journal of Computational and Applied Mathematics, 271(??):319– 327, December 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042714002131.

Wozny:2013:CDB

[Woź13] Pawel Woźny. Construction of dual bases. Journal of Computational and Applied Mathematics, 245(??): 75–85, June 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712005389.

Wozny:2014:CDB

[Woź14] Paweł Woźny. Construction of dual B-spline functions. Journal of Computational and Applied Mathematics, 260(??):301–311, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713005475.

Wright:2010:ADD

[Wri10] R. K. Wright. Accurate discontinuity detection using limited resolution information. Journal of Computational and Applied Mathematics, 234(4):1249–1257, June 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704270900613X.

Weinbub:2014:HFR

[WRS14] Josef Weinbub, Karl Rupp, and Siegfried Selberherr. Highly flexible and reusable finite element simulations with ViennaX. Journal of Computational and Applied Mathematics, 270(??):484– 495, November 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042713006869.

Wang:2013:RAC

[WS13a] Liang Wang and Yimin Shi. Reliability analysis of a class of exponential distribution under record values. Journal of Computational and Applied Mathematics, 239(??):367–379, February 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003652.

Willms:2013:PRR

[WS13b] Allan R. Willms and Emily K. Szusz. Parameter range reduction for ODE models using monotonic discretizations. Journal of Computational and Applied Mathematics, 247(??):124– 151, August 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713000265.

Williams:2014:SQR

[WSJ14] D. M. Williams, L. Shunn, and A. Jameson. Symmetric quadrature rules for simplexes based on sphere close packed lattice arrangements. Journal of Computational and Applied Mathematics, 266(??):18–38, August 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714000211.

Wang:2011:SRB

[WSK^{+}11] Jin Wang, Yunhui Shi, Dehui Kong, Wenpeng Ding, Chunjing Li, and Baocai Yin. Sparse representation based down-sampling image compression. Journal of Computational and Applied Mathematics, 236(5):675– 683, October 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S037704271100361X.

Wu:2014:IAC

[WSS14] Min Wu, Yimin Shi, and Yudong Sun. Inference for accelerated competing failure models from Weibull distribution under Type-I progressive hybrid censoring. Journal of Computational and Applied Mathematics, 263(??):423– 431, June 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713007218.

Wang:2011:MHT

[WSWD11] Ziming Wang, Yongfu Su, Dongxing Wang, and Yucai Dong. A modified Halperntype iteration algorithm for a family of hemi-relatively nonexpansive mappings and systems of equilibrium problems in Banach spaces. Journal of Computational and Applied Mathematics, 235(8):2364–2371, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006023.

Wu:2014:FFB

[WT14] Jingwen Wu and Hongjiong Tian. Functionally-fitted block methods for ordinary differential equations. Journal of Computational and Applied Mathematics, 271(??):356–368, December 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714002039.

Walker:2012:ESM

[WTM12] Ira J. Walker, David S.

Torain II, and Morris H. Morgan III. An eigenvalue search method using the Orr–Sommerfeld equation for shear flow. Journal of Computational and Applied Mathematics, 236(11):2795– 2802, May 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712000131.

Wu:2010:PDF

[Wu10] Qi Wu. Product demand forecasts using wavelet kernel support vector machine and particle swarm optimization in manufacture system. Journal of Computational and Applied Mathematics, 233(10):2481–2491, March 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709007286.

Wu:2014:GMW

[Wu14] Qinghua Wu. On graded meshes for weakly singular Volterra integral equations with oscillatory trigonometric kernels. Journal of Computational and Applied Mathematics, 263(??):370– 376, June 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713007127.

Wadbro:2010:STO

[WUB10] Eddie Wadbro, Rajitha Udawalpola, and Martin Berggren. Shape and topology optimization of an acoustic horn-lens combination. Journal of Computational and Applied Mathematics, 234(6):1781–1787, July 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709005032.

Wang:2010:MCF

[WW10a] Rui-Sheng Wang and Li-Min Wang. Maximum cut in fuzzy nature: Models and algorithms. Journal of Computational and Applied Mathematics, 234(1):240– 252, May 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709008309.

Wang:2010:NDH

[WW10b] Zewen Wang and Rongsheng Wen. Numerical differentiation for high orders by an integration method. Journal of Computational and Applied Mathematics, 234(3):941–948, June 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710000695.

Wu:2010:AEA

[WW10c] Gang Wu and Yimin Wei.

An Arnoldi-extrapolation algorithm for computing PageRank. Journal of Computational and Applied Mathematics, 234(11):3196– 3212, October 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710000804.

Wang:2011:NPI

[WW11a] Tengyao Wang and Joshua M. Weiss. Nevanlinna–Pick interpolation by rational functions with a single pole inside the unit disk. Journal of Computational and Applied Mathematics, 236(6):1497–1501, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004912.

Wu:2011:ZHB

[WW11b] Xiaoqin Wu and Liancheng Wang. Zero-Hopf bifurcation for van der Pol’s oscillator with delayed feedback. Journal of Computational and Applied Mathematics, 235(8):2586–2602, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006278.

Wu:2012:HTC

[WW12] Rong Wu and Wei Wang. The hitting time for a Cox risk process. Journal of Computational and Applied Mathematics, 236(10):2706– 2716, April 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712000118.

Wang:2014:TCM

[WW14] Keyan Wang and Qisheng Wang. Taylor collocation method and convergence analysis for the Volterra– Fredholm integral equations. Journal of Computational and Applied Mathematics, 260(??):294–300, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005050.

Wang:2014:LSA

[WWC14] Qisheng Wang, Keyan Wang, and Shaojun Chen. Least squares approximation method for the solution of Volterra–Fredholm integral equations. Journal of Computational and Applied Mathematics, 272(??):141– 147, December 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042714002465.

Wei:2010:EMP

[WWG10] Yuming Wei, Patricia J. Y.

Wong, and Weigao Ge. The existence of multiple positive solutions to boundary value problems of nonlinear delay differential equations with countably many singularities on infinite interval. Journal of Computational and Applied Mathematics, 233(9):2189–2199, March 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709006748.

Wang:2014:P

[WWGM14] Renhong Wang, Roderick S. Wong, Benqi Guo, and Heping Ma. Preface. Journal of Computational and Applied Mathematics, 265(??):1–2, August 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713007097.

Wang:2010:SLM

[WWL10] Wansheng Wang, Liping Wen, and Shoufu Li. Stability of linear multistep methods for nonlinear neutral delay differential equations in Banach space. Journal of Computational and Applied Mathematics, 233(10):2423– 2437, March 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042709007249.

Wang:2011:P

[WWLY11] Ren-Hong Wang, Yau Shu Wong, Xiao-Nan Luo, and Bao-Cai Yin. Preface. Journal of Computational and Applied Mathematics, 236(5):601, October 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004626.

Wang:2014:NOS

[WWW14] Mingqiu Wang, Xiuli Wang, and Xiaoguang Wang. A note on the one-step estimator for ultrahigh dimensionality. Journal of Computational and Applied Mathematics, 260(??):91– 98, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713004925.

Wang:2010:ECP

[WX10] Haiyong Wang and Shuhuang Xiang. On the evaluation of Cauchy principal value integrals of oscillatory functions. Journal of Computational and Applied Mathematics, 234(1): 95–100, May 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709008152.

Wang:2012:APS

[WX12] Yuanfeng Wang and Zhiting Xu. Asymptotic properties of solutions of certain third-order dynamic equations. Journal of Computational and Applied Mathematics, 236(9):2354–2366, March 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711005917.

Wang:2013:WGF

[WY13] Junping Wang and Xiu Ye. A weak Galerkin finite element method for second-order elliptic problems. Journal of Computational and Applied Mathematics, 241(??):103–115, March 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712004220.

Wang:2011:RNAb

[WYO11] Wentao Wang, Guangxue Yue, and Chunxia Ou. Retraction notice to: Asymptotic constancy for a differential equation with multiple state-dependent delays CAM 233 (2009) pp. 356–360. Journal of Computational and Applied Mathematics, 235(11):3367, April 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711000902.

Wang:2010:CTR

[WYP10] Kuo-Hsiung Wang, Dong-Yuh Yang, and W. L. Pearn. Comparison of two randomized policy M/G/1 queues with second optional service, server breakdown and startup. Journal of Computational and Applied Mathematics, 234(3):812– 824, June 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710000580.

Wang:2012:SAI

[WYY^{+}12] Juan Wang, Jianxiong Ye, Hongchao Yin, Enmin Feng, and Lei Wang. Sensitivity analysis and identification of kinetic parameters in batch fermentation of glycerol. Journal of Computational and Applied Mathematics, 236(9):2268– 2276, March 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711005851.

Wang:2010:TWS

[WZ10a] Qi-Ru Wang and Kai Zhou.

Traveling wave solutions in delayed reaction–diffusion systems with mixed monotonicity. Journal of Computational and Applied Mathematics, 233(10):2549–2562, March 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704270900733X.

Wang:2010:IAP

[WZ10b] Xiaohuan Wang and Jihui Zhang. Impulsive antiperiodic boundary value problem of first-order integrodifferential equations. Journal of Computational and Applied Mathematics, 234 (12):3261–3267, October 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042710002220.

Wang:2010:FRS

[WZ10c] Yin Wang and Jun Zhang. Fast and robust sixth-order multigrid computation for the three-dimensional convection– diffusion equation. Journal of Computational and Applied Mathematics, 234 (12):3496–3506, October 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042710002980.

Wu:2010:PSH

[WZ10d] Hongyun Wu and Jihui Zhang. Positive solutions of higher-order four-point boundary value problem with p-Laplacian operator. Journal of Computational and Applied Mathematics, 233(11):2757–2766, April 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709003793.

Wen:2011:IRM

[WZ11a] C. H. Wen and T. S.

Zhang. Improved rectangular method on stochastic Volterra equations. Journal of Computational and Applied Mathematics, 235(8):2492–2501, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006187.

Wong:2011:ODB

[WZ11b] Hoi Ying Wong and Jing Zhao. Optimal dividends and bankruptcy procedures: Analysis of the Ornstein–Uhlenbeck process. Journal of Computational and Applied Mathematics, 236(2):150–166, August 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711003372.

Wang:2012:IPR

[WZ12] Ping Wang and Fangguo Zhang. Improved Pollard rho method for computing discrete logarithms over finite extension fields. Journal of Computational and Applied Mathematics, 236(17):4336–4343, November 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712001392.

Wu:2014:PSC

[WZ14] Jinbiao Wu and Hui Zheng. Parallel subspace correction methods for nearly singular systems. Journal of Computational and Applied Mathematics, 271(??):180–194, December 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042714001939.

Wang:2010:NISb

[WZA10] Chao Wang, Jinghao Zhu, and Lili An. New iterative schemes for asymptotically quasi-nonexpansive nonself-mappings in Banach spaces. Journal of Computational and Applied Mathematics, 233(11):2948–2955, April 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709007869.

Wang:2010:BPP

[WZD10] Xiaoying Wang, Shugong Zhang, and Tian Dong. A bivariate preprocessing paradigm for the Buchberger– Möller algorithm. Journal of Computational and Applied Mathematics, 234 (12):3344–3355, October 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042710002438.

Wang:2011:DTO

[WZF11] Shanshan Wang, Luming Zhang, and Ran Fan. Discrete-time orthogonal spline collocation methods for the nonlinear Schrödinger equation with wave operator. Journal of Computational and Applied Mathematics, 235(8):1993– 2005, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042710005510.

Wang:2011:SON

[WZL11] Qi Wang, Qingyong Zhu, and Mingzhu Liu. Stability and oscillations of numerical solutions for differential equations with piecewise continuous arguments of alternately advanced and retarded type. Journal of Computational and Applied Mathematics, 235(5):1542– 1552, January 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042710004905.

Wang:2011:CCT

[WZLP11] Wen-Ke Wang, Hui Zhang, Xiao-Ming Liu, and Jean-Claude Paul. Conditions for coincidence of two cubic Bézier curves. Journal of Computational and Applied Mathematics, 235(17):5198– 5202, July 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711002536.

Wang:2011:NSS

[WZMC11] B. Wang, H. Q. Zhang, M. Manhart, and C. K. Chan. A numerical study on statistical temporal scales in inertia particle dispersion. Journal of Computational and Applied Mathematics, 235(13):3807–3816, May 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711000380.

Wang:2010:ESF

[WZS10] Guotao Wang, Lihong Zhang, and Guangxing Song. Extremal solutions for the first order impulsive functional differential equations with upper and lower solutions in reversed order. Journal of Computational and Applied Mathematics, 235(1):325–333, November 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710003584.

Wu:2014:SSM

[WZY^{+}14] Meng Wu, Weihong Zhang, Zhouwang Yang, Jiansong Deng, and Falai Chen. Spline spaces with mixed orders of continuity over T - meshes. Journal of Computational and Applied Mathematics, 268(??):179–193, October 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714001332.

Xu:2014:SII

[XC14] Ai-Min Xu and Zhong-Di Cen. Some identities involving exponential functions and Stirling numbers and applications. Journal of Computational and Applied Mathematics, 260(??):201– 207, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713005323.

Xu:2014:DDM

[XCQ14] Jing Xu, Hui Bin Chang, and Jing Qin. Domain decomposition method for image deblurring. Journal of Computational and Applied Mathematics, 271(??):401– 414, December 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042714001794.

Xing:2010:EEM

[XCY10] Xiaoqing Xing, Yanping Chen, and Nianyu Yi. Error estimates of mixed finite element methods for quadratic optimal control problems. Journal of Computational and Applied Mathematics, 233(8):1812–1820, February 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709006402.

Xu:2012:STG

[XCY12] Daobao Xu, Zhiping Chen, and Li Yang. Scenario tree generation approaches using K-means and LP moment matching methods. Journal of Computational and Applied Mathematics, 236(17):4561–4579, November 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712002300.

Xu:2011:MPM

[XCZ^{+}11] Wei-Wei Xu, Wai-Ki Ching, Shu-Qin Zhang, Wen Li, and Xiao-Shan Chen. A matrix perturbation method for computing the steadystate probability distributions of probabilistic Boolean networks with gene perturbations. Journal of Computational and Applied Mathematics, 235(8):2242– 2251, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704271000587X.

Xiong:2012:IPF

[XGL12] Xiangtuan Xiong, Hongbo Guo, and Xiaohong Liu. An inverse problem for a fractional diffusion equation. Journal of Computational and Applied Mathematics, 236(17):4474–4484, November 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712001872.

Xu:2013:AUR

[XH13] Huixia Xu and Qianqian Hu. Approximating uniform rational B-spline curves by polynomial Bspline curves. Journal of Computational and Applied Mathematics, 244(??):10– 18, May 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712005122.

Xu:2014:LSR

[XHmDW14] Yong-Li Xu, Min Han, Xue mei Dong, and Min Wang. Least squares regression with l_{1}-regularizer in sum space. Journal of Computational and Applied Mathematics, 261(??):394– 405, May 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713006614.

Xu:2011:APB

[XhW11] Yan Xu and Ren hong Wang. Asymptotic properties of B-splines, Eulerian numbers and cube slicing. Journal of Computational and Applied Mathematics, 236(5):988–995, October 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004420.

Xie:2012:CSI

[XHZ12] Qiyue Xie, Zhengzhi Han, and Wei Zhang. Chaotification via system immersion. Journal of Computational and Applied Mathematics, 236(7):1775–1782, January 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711005395.

Xie:2011:IAN

[Xie11]  Dexuan Xie.  An im-

proved approximate Newton method for implicit Runge–Kutta formulas. Journal of Computational and Applied Mathematics, 235 (17):5249–5258, July 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042711002846.

Xie:2012:CEP

[Xie12] Huiqing Xie. Computation of eigenpair partial derivatives by Rayleigh– Ritz procedure. Journal of Computational and Applied Mathematics, 236(10):2607– 2621, April 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711006248.

Xiong:2010:RMC

[Xio10]  Xiang-Tuan Xiong.  A

regularization method for a Cauchy problem of the Helmholtz equation. Journal of Computational and Applied Mathematics, 233(8):1723–1732, February 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709006219.

Xueli:2011:NAE

[XJ11] Song Xueli and Peng Jigen. A novel approach to exponential stability of nonlinear systems with time-varying delays. Journal of Computational and Applied Mathematics, 235(6):1700–1705, January 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710005029.

Xiong:2011:EAO

[XL11] Chunguang Xiong and Yuan Li. Error analysis for optimal control problem governed by convection diffusion equations: DG method. Journal of Computational and Applied Mathematics, 235(10):3163–3177, March 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006746.

Xuan:2012:NMB

[XL12] Yan Xuan and Fu-Rong Lin.

Numerical methods based on rational variable substitution for Wiener–Hopf equations of the second kind. Journal of Computational and Applied Mathematics, 236(14):3528–3539, August 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712001239.

Xu:2010:BEE

[XLCC10] Wei-Wei Xu, Wen Li, Wai-Ki Ching, and Yan-Mei Chen. Backward errors for eigenproblem of two kinds of structured matrices. Journal of Computational and Applied Mathematics, 235(1):59–73, November 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710002876.

Li:2010:SPD

[xLwCL10] Tie xiang Li, Eric King wah Chu, and Wen-Wei Lin. A structure-preserving doubling algorithm for quadratic eigenvalue problems arising from time-delay systems. Journal of Computational and Applied Mathematics, 233(8):1733–1745, February 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709006323.

Xu:2011:NLN

[XMC11] Fengmin Xu, Xusheng Ma, and Baili Chen. A new Lagrangian net algorithm for solving max-bisection problems. Journal of Computational and Applied Mathematics, 235(13):3718–3723, May 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711000264.

Xu:2011:PFM

[XMSS11] Xinsheng Xu, Zhiqing Meng, Jianwu Sun, and Rui Shen. A penalty function method based on smoothing lower order penalty function. Journal of Computational and Applied Mathematics, 235(14):4047–4058, May 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711001191.

Xiang:2014:TLA

[XN14] Hua Xiang and Frédéric Nataf. Two-level algebraic domain decomposition preconditioners using Jacobi–Schwarz smoother and adaptive coarse grid corrections. Journal of Computational and Applied Mathematics, 261(??):1–13, May 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005712.

Xiao:2010:ACF

[XSH10] Jian-Zhong Xiao, Jing Sun, and Xuan Huang. Approximating common fixed points of asymptotically quasi-nonexpansive mappings by a k + 1-step iterative scheme with error terms. Journal of Computational and Applied Mathematics, 233(8):2062–2070, February 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709006633.

Xu:2012:RSN

[XSV12] Juncai Xu, Zhenzhong Shen, and Bernhard Valentini. RETRACTED: Study of the new Leon model for concrete failure. Journal of Computational and Applied Mathematics, 236(18):4873– 4881, December 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712000593.

See retraction notice [XSV13].

Xu:2013:RNS

[XSV13] Juncai Xu, Zhenzhong Shen, and Bernhard Valentini. Retraction notice to “Study of New Leon model for concrete failure” [J. Comput. Appl. Math. 236 (18) (2012) 4873–4881]. Journal of Computational and Applied Mathematics, 249(??):163, September 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713000678.

See [XSV12].

Xiao:2012:MCG

[XSW12] Yunhai Xiao, Huina Song, and Zhiguo Wang. A modified conjugate gradient algorithm with cyclic Barzilai– Borwein steplength for unconstrained optimization. Journal of Computational and Applied Mathematics, 236(13):3101–3110, July 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271200043X.

Xu:2011:NAV

[Xu11a] Da Xu. The numerical analysis on a Volterra equation with asymptotically periodic solution. Journal of Computational and Applied Mathematics, 236(5):684– 698, October 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711003608.

Xu:2011:NAS

[Xu11b]  Mingyu Xu.  Numeri-

cal algorithms and simulations for reflected backward stochastic differential equations with two continuous barriers. Journal of Computational and Applied Mathematics, 236(6):1137–1154, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004390.

Xiong:2012:TTM

[XW12] Xiangtuan Xiong and Junxia Wang. A Tikhonov-type method for solving a multidimensional inverse heat source problem in an unbounded domain. Journal of Computational and Applied Mathematics, 236(7):1766– 1774, January 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042711005383.

Xiao:2010:NDY

[XWW10] Yunhai Xiao, Qiuyu Wang, and Dong Wang. Notes on the Dai–Yuan–Yuan modified spectral gradient method. Journal of Computational and Applied Mathematics, 234(10):2986–2992, September 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042710002098.

Xu:2014:EEO

[XXH14] Zhenhua Xu, Shuhuang Xiang, and Guo He. Efficient evaluation of oscillatory Bessel Hilbert transforms. Journal of Computational and Applied Mathematics, 258(??):57– 66, March 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713004433.

Xu:2014:CRM

[XYW14] Lin Xu, Hailiang Yang, and Rongming Wang. Cox risk model with variable premium rate and stochastic return on investment. Journal of Computational and Applied Mathematics, 256(??):52–64, January 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003609.

Xu:2014:BEC

[XZ14] Weiquan Xu and Pingwen Zhang. Boundary effects in confined copolymer system and compressible SCFT model. Journal of Computational and Applied Mathematics, 265(??):290–300, August 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005098.

Xiong:2014:SEM

[XZH14] Xiangtuan Xiong, Liping Zhao, and Y. C. Hon. Stability estimate and the modified regularization method for a Cauchy problem of the fractional diffusion equation. Journal of Computational and Applied Mathematics, 272(??):180–194, December 15, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042714002611.

Xia:2011:FAM

[XZL11] Peng Xia, Shugong Zhang, and Na Lei. A Fitzpatrick algorithm for multivariate rational interpolation. Journal of Computational and Applied Mathematics, 235(17):5222–5231, July 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711002597.

Xia:2014:MGM

[XZW14] Kelin Xia, Meng Zhan, and Guo-Wei Wei. MIB Galerkin method for elliptic interface problems. Journal of Computational and Applied Mathematics, 272(??):195–220, December 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714002507.

Yousefi-Azari:2011:CEW

[YAKA11] H. Yousefi-Azari, M. H. Khalifeh, and A. R. Ashrafi. Calculating the edge Wiener and edge Szeged indices of graphs. Journal of Computational and Applied Mathematics, 235(16):4866–4870, June 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711000951.

Yang:2010:PDD

[Yan10a] Danping Yang. Parallel domain decomposition procedures of improved D-D type for parabolic problems. Journal of Computational and Applied Mathematics, 233(11):2779–2794, April 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709007651.

Yang:2010:SCB

[Yan10b] Zhi-Ming Yang. Some closer bounds of Perron root basing on generalized Perron complement. Journal of Computational and Applied Mathematics, 235(1):315– 324, November 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710003560.

Yan:2011:ESN

[Yan11] Zuomao Yan. Existence of solutions for nonlocal impulsive partial functional integrodifferential equations via fractional operators. Journal of Computational and Applied Mathematics, 235(8):2252–2262, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710005881.

Yang:2013:UAS

[Yan13] Chaoxia Yang. A unified approach with spectral convergence for the evaluation of hypersingular and supersingular integrals with a periodic kernel. Journal of Computational and Applied Mathematics, 239(??):322– 332, February 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042712003500.

Yardimci:2010:NSS

[Yar10] Seyhmus Yardimci. A note on the spectral singularities of non-selfadjoint matrixvalued difference operators. Journal of Computational and Applied Mathematics, 234(10):3039–3042, September 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710002141.

Yousefi:2011:OMB

[YBD11] S. A. Yousefi, M. Behroozifar, and Mehdi Dehghan. The operational matrices of Bernstein polynomials for solving the parabolic equation subject to specification of the mass. Journal of Computational and Applied Mathematics, 235(17):5272– 5283, July 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711002950.

Yang:2010:LFN

[YC10] Yongqin Yang and Shaochun Chen. A locking-free nonconforming triangular element for planar elasticity with pure traction boundary condition. Journal of Computational and Applied Mathematics, 233(10):2703– 2710, March 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709007602.

Yang:2011:PWB

[YC11] Chao Yang and Xiao-Chuan Cai. A parallel well-balanced finite volume method for shallow water equations with topography on the cubedsphere. Journal of Computational and Applied Mathematics, 235(18):5357–5366, July 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711000276.

Yuan:2011:CIQ

[YD11] Yongxin Yuan and Hua Dai. On a class of inverse quadratic eigenvalue problem. Journal of Computational and Applied Mathematics, 235(8):2662–2669, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710006345.

Yuan:2012:IPU

[YD12] Yongxin Yuan and Hua Dai. An inverse problem for undamped gyroscopic systems. Journal of Computational and Applied Mathematics, 236(9):2574– 2581, March 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711006297.

Yalcin:2014:GBD

[YE14] Femin Yalcin and Serkan Eryilmaz. q-geometric and q-binomial distributions of order k. Journal of Computational and Applied Mathematics, 271(??):31–38, December 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042714001745.

Yilmaz:2014:PID

[YESD14] Ilyas Yilmaz, Firat Oguz Edis, Hasan Saygin, and Lars Davidson. Parallel implicit DNS of temporallyevolving turbulent shear layer instability. Journal of Computational and Applied Mathematics, 259 (part B) (??):651–659, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042713001787.

Yeung:2014:MBM

[Yeu14] Man-Chung Yeung. ML(n)BiCGStabt:

a ML(n)BiCGStab variant with A-transpose. Journal of Computational and Applied Mathematics, 272(??): 57–69, December 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042714001927.

Yang:2014:MRM

[YF14] Fan Yang and Chu-Li Fu. A mollification regularization method for the inverse spatial-dependent heat source problem. Journal of Computational and Applied Mathematics, 255 (??):555–567, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042713003105.

Yang:2010:AAS

[YFWH10] Tianzhi Yang, Bo Fang, Xingyuan Wang, and Wenhu Huang. Approximate analytical solutions for Kolmogorov’s equations. Journal of Computational and Applied Mathematics, 235(3):747–755, December 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710003808.

You:2014:SEN

[YFZ14] Xiong You, Yonglei Fang, and Jinxi Zhao. Special extended Nyström tree theory for ERKN methods. Journal of Computational and Applied Mathematics, 263(??): 478–499, June 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042713007164.

Yorubulut:2013:BPG

[YG13] Serap Yörübulut and Omer L. Gebizlioğlu. Bivariate Pseudo-Gompertz distribution and concomitants of its order statistics. Journal of Computational and Applied Mathematics, 247(??):68– 83, August 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713000253.

Yorubulut:2014:CUR

[YG14] Serap Yörübulut and Omer L.

Gebizlioğlu. On concomitants of upper record statistics and survival analysis for a pseudo-Gompertz distribution. Journal of Computational and Applied Mathematics, 256(??):211–218, January 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003610.

Yin:2014:IES

[YGHH14] Feng Yin, Ke Guo, Guangxin Huang, and Bormin Huang. The inverse eigenproblem with a submatrix constraint and the associated approximation problem for (R, S)-symmetric matrices. Journal of Computational and Applied Mathematics, 268(??):23–33, October 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714000661.

Guo:2011:PMQ

[yGlJ11] Ben yu Guo and Hong li Jia.

Pseudospectral method for quadrilaterals. Journal of Computational and Applied Mathematics, 236(5):962– 979, October 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711002214.

Guo:2013:SMH

[yGZ13] Ben yu Guo and Chao Zhang. The spectral method for high order problems with proper simulations of asymptotic behaviors at infinity. Journal of Computational and Applied Mathematics, 237(1):269–294, January 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003317.

Yang:2011:DMC

[YJ11] Qing Yang and Ziwen Jiang. A discontinuous mixed covolume method for elliptic problems. Journal of Computational and Applied Mathematics, 235(8):2467– 2476, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042710006163.

Yilmaz:2011:SOC

[YK11] F. Yilmaz and B. Karasözen. Solving optimal control problems for the unsteady Burgers equation in COM-SOL multiphysics. Journal of Computational and Applied Mathematics, 235(16): 4839–4850, June 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042711000033.

Yilmaz:2014:AOA

[YK14] Fikriye Yilmaz and Bülent Karasözen. An all-at-once approach for the optimal control of the unsteady Burgers equation. Journal of Computational and Applied Mathematics, 259 (part B)(??):771–779, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003348.

Yang:2010:LTI

[YKL10] Xiaojing Yang, Yong-In Kim, and Kueiming Lo. Lyapunov-type inequality for a class of odd-order differential equations. Journal of Computational and Applied Mathematics, 234 (10):2962–2968, September 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710002050.

Yao:2011:AAM

[YKL11] Yonghong Yao, Shin Min Kang, and Yeong-Cheng Liou. Algorithms for approximating minimization problems in Hilbert spaces. Journal of Computational and Applied Mathematics, 235(12):3515–3526, April 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711000835.

Yan:2011:CAT

[YL11a] Lanlan Yan and Jiongfeng Liang. A class of algebraictrigonometric blended splines. Journal of Computational and Applied Mathematics, 235(6):1713–1729, January 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710005297.

Yang:2011:DIS

[YL11b] Hu Yang and Xifu Liu. The Drazin inverse of the sum of two matrices and its applications. Journal of Computational and Applied Mathematics, 235(5):1412–1417, January 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710004760.

Yu:2011:BFV

[YL11c] Changhua Yu and Yonghai Li. Biquadratic finite volume element methods based on optimal stress points for parabolic problems. Journal of Computational and Applied Mathematics, 236(6):1055–1068, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004250.

Yuan:2014:IUM

[YL14] Yongxin Yuan and Hao Liu. An iterative updating method for damped structural systems using symmetric eigenstructure assignment. Journal of Computational and Applied Mathematics, 256(??):268– 277, January 15, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713003919.

Yu:2013:LML

[YLD13] Bo Yu, Dong-Hui Li, and Ning Dong. Low memory and low complexity iterative schemes for a nonsymmetric algebraic Riccati equation arising from transport theory. Journal of Computational and Applied Mathematics, 250(??):175–189, October 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713001258.

Liu:2011:DIJ

[yLGP11] Da yan Liu, Olivier Gibaru, and Wilfrid Perruquetti. Differentiation by integration with Jacobi polynomials. Journal of Computational and Applied Mathematics, 235(9):3015–3032, March 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710006734.

Yang:2012:RPD

[YLS12] Yang Yang, Remigijus Leipus, and Jonas Siaulys. On the ruin probability in a dependent discrete time risk model with insurance and financial risks. Journal of Computational and Applied Mathematics, 236(13):3286– 3295, July 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712000970.

Yu:2014:NSF

[YLTB14] Q. Yu, F. Liu, I. Turner, and K. Burrage. Numerical simulation of the fractional Bloch equations. Journal of Computational and Applied Mathematics, 255(??):635– 651, January 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713003257.

Yang:2014:FCM

[YMB14] Daegil Yang, George J.

Moridis, and Thomas A. Blasingame. A fully coupled multiphase flow and geomechanics solver for highly heterogeneous porous media. Journal of Computational and Applied Mathematics, 270(??):417–432, November 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713007024.

Yaguchi:2010:CNS

[YMS10] Takaharu Yaguchi, Takayasu Matsuo, and Masaaki Sugihara. Conservative numerical schemes for the Ostrovsky equation. Journal of Computational and Applied Mathematics, 234(4):1036– 1048, June 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709002179.

Yamamoto:2010:ALC

[YNO10] Y. Yamamoto, T. Nagase, and M. Ohmiya. Appell’s lemma and conservation laws of KdV equation. Journal of Computational and Applied Mathematics, 233(6):1612–1618, January 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709001708.

Yerlikaya-Ozkurt:2014:EHP

[YÖVAYOW14] F. Yerlikaya-Özkurt, C. Vardar-Acar, Y. Yolcu-Okur, and G.-W. Weber. Estimation of the Hurst parameter for fractional Brownian motion using the CMARS method. Journal of Computational and Applied Mathematics, 259 (part B)(??):843–850, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003920.

Yu:2010:SNM

[YP10a] Haodong Yu and Dingguo Pu. Smoothing Newton method for NCP with the identification of degenerate indices. Journal of Computational and Applied Mathematics, 234(12):3424–3435, October 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271000258X.

Peng:2010:NMI

[yP10b] Zhen yun Peng. New matrix iterative methods for constraint solutions of the matrix equation AXB = C. Journal of Computational and Applied Mathematics, 235(3):726–735, December 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271000378X.

Yan:2010:GCD

[YPL10] Qin-Rong Yan, Xiao-Zhen Peng, and Dong-Hui Li. A globally convergent derivativefree method for solving large-scale nonlinear monotone equations. Journal of Computational and Applied Mathematics, 234(3):649– 657, June 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271000004X.

Sun:2005:CTS

[yS05]  Li ying Sun.  A com-

parison theorem for the SOR iterative method. Journal of Computational and Applied Mathematics, 181(2):336–341, September 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/S0377042704005898.

See comment [NK10].

Yu:2011:MSP

[YSQ11] Zhensheng Yu, Jing Sun, and Yi Qin. A multivariate spectral projected gradient method for bound constrained optimization. Journal of Computational and Applied Mathematics, 235(8):2263–2269, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710005893.

Yang:2012:RKM

[YSW12] Li-Hong Yang, Ji-Hong Shen, and Yue Wang. The reproducing kernel method for solving the system of the linear Volterra integral equations with variable coefficients. Journal of Computational and Applied Mathematics, 236(9):2398– 2405, March 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711005966.

Yin:2013:EPJ

[YSW13] Chuancun Yin, Ying Shen, and Yuzhen Wen. Exit problems for jump processes with applications to dividend problems. Journal of Computational and Applied Mathematics, 245(??): 30–52, June 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712005353.

Yang:2011:CPM

[YSZ11] Xiaoyuan Yang, Yan Shi, and Wanlu Zhou. Construction of parameterizations of masks for tight wavelet frames with two symmetric/antisymmetric generators and applications in image compression and denoising. Journal of Computational and Applied Mathematics, 235(8):2112–2136, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710005650.

Yamamura:2012:FAS

[YT12] Kiyotaka Yamamura and Naoya Tamura. Finding all solutions of separable systems of piecewise-linear equations using integer programming. Journal of Computational and Applied Mathematics, 236(11):2844– 2852, May 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712000325.

Yu:2011:GGM

[YTFP11] Miaomiao Yu, Yinghui Tang, Yonghong Fu, and Lemeng Pan. GI/Geom/1/N/MWV queue with changeover time and searching for the optimum service rate in working vacation period. Journal of Computational and Applied Mathematics, 235(8):2170–2184, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710005698.

Yuan:2012:DAF

[Yua12]  Quan Yuan.  Dual ap-

proaches to finite element model updating. Journal of Computational and Applied Mathematics, 236(7):1851– 1861, January 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042711005607.

Yun:2011:NHO

[Yun11] Beong In Yun. New higher order methods for solving nonlinear equations with multiple roots. Journal of Computational and Applied Mathematics, 235(5):1553– 1555, January 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704271000422X.

Yun:2012:IMS

[Yun12]  Beong In Yun.  Itera-

tive methods for solving nonlinear equations with finitely many roots in an interval. Journal of Computational and Applied Mathematics, 236(13):3308– 3318, July 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712001045.

Yu:2010:SAO

[YW10] Yuexin Yu and Liping Wen.

Stability analysis of one-leg methods for nonlinear functional differential and functional equations. Journal of Computational and Applied Mathematics, 235(3):817– 824, December 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704271000395X.

Yuan:2014:MPR

[YWL14] Gonglin Yuan, Zengxin Wei, and Guoyin Li. A modified Polak–Ribière–Polyak conjugate gradient algorithm for nonsmooth convex programs. Journal of Computational and Applied Mathematics, 255(??):86–96, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713002185.

Yang:2010:OSA

[YWW10] Dong-Yuh Yang, Kuo-Hsiung Wang, and Chia-Huang Wu. Optimization and sensitivity analysis of controlling arrivals in the queueing system with single working vacation. Journal of Computational and Applied Mathematics, 234(2): 545–556, May 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042709008632.

Yang:2014:WVP

[YWY^{+}14] Cheng-Lei Yang, Wei-Zhen Wang, Yi-Jun Yang, Lin Lu, Zhi-Jie Zhu, Binhai Zhu, and Wei Zeng. Weak visibility polygons of NURBS curves inside simple polygons. Journal of Computational and Applied Mathematics, 256(??):1–15, January 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713003506.

Yuen:2010:OPR

[YY10] Fei Lung Yuen and Hailiang Yang. Option pricing with regime switching by trinomial tree method. Journal of Computational and Applied Mathematics, 233(8):1821–1833, February 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709006414.

Yakhno:2011:EAE

[YY11] V. G. Yakhno and H. Çerdik Yaslan. Equations of anisotropic elastodynamics as a symmetric hyperbolic system: Deriving the timedependent fundamental solution. Journal of Computational and Applied Mathematics, 235(16):4802–4815, June 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271000614X.

Yang:2010:PMI

[YYL^{+}10] Xiaofeng Yang, Jinping Yu, Shenghong Li, Albert Jerry Cristoforo, and Xiaohu Yang. Pricing model of interest rate swap with a bilateral default risk. Journal of Computational and Applied Mathematics, 234(2): 512–517, May 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042709008590.

Yang:2010:DRM

[YZ10a] Hu Yang and Zhimin Zhang. On a discrete risk model with two-sided jumps. Journal of Computational and Applied Mathematics, 234(3):835–844, June 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710000609.

Yue:2010:CUA

[YZ10b] Gaocheng Yue and Chengkui Zhong. On the convergence of the uniform attractor of 2D NS–α model to the uniform attractor of 2D NS system. Journal of Computational and Applied Mathematics, 233(8):1879–1887, February 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709006463.

Yang:2011:NSF

[YZ11a] Huidong Yang and Walter Zulehner. Numerical simulation of fluidstructure interaction problems on hybrid meshes with algebraic multigrid methods. Journal of Computational and Applied Mathematics, 235(18):5367–5379, July 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711003189.

Yang:2011:ACC

[YZ11b] Kai Yang and Ran Zhang.

Analysis of continuous collocation solutions for a kind of Volterra functional integral equations with proportional delay. Journal of Computational and Applied Mathematics, 236(5):743–752, October 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711003293.

Zhang:2011:AME

[yZ11c] Zhen yu Zhang. Approximate method for eigensensitivity analysis of a defective matrix. Journal of Computational and Applied Mathematics, 235(9):2913–2927, March 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006564.

Zeghdane:2011:HOS

[ZAT11] R. Zeghdane, L. Abbaoui, and A. Tocino. Higherorder semi-implicit Taylor schemes for Itô stochastic differential equations.

Journal of Computational and Applied Mathematics, 236(6):1009–1023, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711003359.

Zhang:2013:NFI

[ZAWA13] Lihong Zhang, Bashir Ahmad, Guotao Wang, and Ravi P. Agarwal. Nonlinear fractional integrodifferential equations on unbounded domains in a Banach space. Journal of Computational and Applied Mathematics, 249(??):51– 56, September 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713000794.

Zhao:2013:ADM

[ZBC13] Zhi Zhao, Zhengjian Bai, and Guizhi Chen. On the alternating direction method of multipliers for nonnegative inverse eigenvalue problems with partial eigendata. Journal of Computational and Applied Mathematics, 239(??):114–134, February 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003962.

Zhang:2011:NAC

[ZBW11] Luming Zhang, Dongmei Bai, and Shanshan Wang. Numerical analysis for a conservative difference scheme to solve the Schrödinger–Boussinesq equation. Journal of Computational and Applied Mathematics, 235(17):4899– 4915, July 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711001774.

Zhu:2010:MCL

[ZBZ10] J. H. Zhu, P. Beckers, and W. H. Zhang. On the multi-component layout design with inertial force. Journal of Computational and Applied Mathematics, 234(7):2222–2230, August 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709005585.

Zhang:2010:PLI

[ZC10a] H. Zhang and L. Z. Cheng.

Projected Landweber iteration for matrix completion. Journal of Computational and Applied Mathematics, 235(3):593–601, December 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710003547.

Zhao:2010:EMP

[ZC10b] Yulin Zhao and Haibo Chen. The existence of multiple positive solutions for singular functional differential equations with sign-changing nonlinearity. Journal of Computational and Applied Mathematics, 234(5):1543–1550, July 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710001329.

Zhu:2010:CSA

[ZC10c] Guoqing Zhu and Shaochun Chen. Convergence and superconvergence analysis of an anisotropic nonconforming finite element methods for singularly perturbed reaction–diffusion problems. Journal of Computational and Applied Mathematics, 234(10):3048–3063, September 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710002189.

Zhou:2013:CMR

[ZC13] Weijun Zhou and Xinlong Chen. On the convergence of a modified regularized Newton method for convex optimization with singular solutions. Journal of Computational and Applied Mathematics, 239(??):179– 188, February 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712004037.

Zhang:2013:PEA

[ZCC^{+}13] Yunong Zhang, Yao Chou, Jinhao Chen, Zhijun Zhang, and Lin Xiao. Presentation, error analysis and numerical experiments on a group of 1-step-ahead numerical differentiation formulas. Journal of Computational and Applied Mathematics, 239(??):406–414, February 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003755.

Zhang:2010:RSC

[ZCH10] Wei Zhang, Xiu-Shan Cai, and Zheng-Zhi Han. Robust stability criteria for systems with interval timevarying delay and nonlinear perturbations. Journal of Computational and Applied Mathematics, 234(1):174– 180, May 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709008218.

Zhou:2011:CHO

[ZCS11] Xiaojian Zhou, Xin Chen, and Yongzhong Song. Constructing higher-order methods for obtaining the multiple roots of nonlinear equations. Journal of Computational and Applied Mathematics, 235(14):4199–4206, May 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711001440.

Zhou:2014:NTR

[ZCX14] Qunyan Zhou, Jun Chen, and Zhengwei Xie. A nonmonotone trust region method based on simple quadratic models. Journal of Computational and Applied Mathematics, 272(??):107–115, December 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042714002337.

Zhao:2010:AFD

[ZD10] Di Zhao and Weizhong Dai. Accurate finite difference schemes for solving a 3D micro heat transfer model in an N-carrier system with the Neumann boundary condition in spherical coordinates. Journal of Computational and Applied Mathematics, 235(3):850–869, December 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710004140.

Zhu:2011:MDS

[ZD11] Liyong Zhu and Qiang Du. Mesh-dependent stability for finite element approximations of parabolic equations with mass lumping. Journal of Computational and Applied Mathematics, 236(5):801–811, October 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711002871.

Zhang:2013:NZI

[ZD13] Mingbo Zhang and Jiansong Deng. Number of zeros of interval polynomials. Journal of Computational and Applied Mathematics, 237(1):102–110, January 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003019.

Zhang:2014:IMT

[ZD14] Yuxin Zhang and Hengfei Ding. Improved matrix transform method for the Riesz space fractional reaction dispersion equation. Journal of Computational and Applied Mathematics, 260(??):266–280, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713004950.

Zdenek:2014:OEP

[ZDP^{+}14] Kubík Zdenek, Nikolayev Denys, Karban Pavel, Skála Jirí, and Hromádka Miroslav. Optimization of electrical properties of parallel plate antenna for EMC testing. Journal of Computational and Applied Mathematics, 270(??):283–293, November 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714000119.

Zhang:2011:NFG

[ZDZ11] Jianhua Zhang, Hua Dai, and Jing Zhao. A new family of global methods for linear systems with multiple right-hand sides. Journal of Computational and Applied Mathematics, 236(6):1562–1575, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004997.

Zhang:2014:TTF

[ZF14] Xuemei Zhang and Meiqiang Feng. Transformation techniques and fixed point theories to establish the positive solutions of second order impulsive differential equations. Journal of Computational and Applied Mathematics, 271(??):117–129, December 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714001897.

Zhang:2010:ESB

[ZFG10] Xuemei Zhang, Meiqiang Feng, and Weigao Ge. Existence of solutions of boundary value problems with integral boundary conditions for second-order impulsive integro-differential equations in Banach spaces. Journal of Computational and Applied Mathematics, 233(8):1915–1926, February 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709006505.

Zlatev:2010:SRE

[ZFH10] Zahari Zlatev, István Faragó, and Ágnes Havasi. Stability of the Richardson extrapolation applied together with the θ-method. Journal of Computational and Applied Mathematics, 235(2):507–517, November 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710003420.

Zhang:2014:PPC

[ZFM14] Yuan-Xiang Zhang, Chu-Li Fu, and Yun-Jie Ma. An a posteriori parameter choice rule for the truncation regularization method for solving backward parabolic problems. Journal of Computational and Applied Mathematics, 255(??):150– 160, January 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713002525.

Zhang:2011:AIM

[ZFY11] Yuan-Xiang Zhang, Chu-Li Fu, and Liang Yan. Approximate inverse method for stable analytic continuation in a strip domain. Journal of Computational and Applied Mathematics, 235(9):2979–2992, March 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006679.

Zhang:2014:FGG

[ZG14] Ke Zhang and Chuanqing Gu. Flexible global generalized Hessenberg methods for linear systems with multiple right-hand sides. Journal of Computational and Applied Mathematics, 263(??):312– 325, June 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713006900.

Zhu:2013:SDA

[ZGZB13] Ji-Hong Zhu, Xiao-Jun Gu, Wei-Hong Zhang, and Pierre Beckers. Structural design of aircraft skin stretch-forming die using topology optimization. Journal of Computational and Applied Mathematics, 246(??):278–288, July 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003627.

Zheng:2010:SOH

[ZGZW10] Yi Zheng, Wenbin Guo, Jianli Zhao, and Nan Wang. On structure-oriented hybrid two-stage iteration methods for the large and sparse blocked system of linear equations. Journal of Computational and Applied Mathematics, 234(8):2367– 2376, August 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042710001652.

Zhang:2010:PAM

[ZH10a] Mian Zhang and Zhengting Hou. Performance analysis of M/G/1 queue with working vacations and vacation interruption. Journal of Computational and Applied Mathematics, 234 (10):2977–2985, September 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710002074.

Zhang:2010:ASS

[ZH10b] Yan Zhang and Yinnian He. Assessment of subgrid-scale models for the incompressible Navier–Stokes equations. Journal of Computational and Applied Mathematics, 234(2):593–604, May 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710000038.

Zhang:2010:NST

[ZH10c] Ying Zhang and Zheng-Hai Huang. A nonmonotone smoothing-type algorithm for solving a system of equalities and inequalities. Journal of Computational and Applied Mathematics, 233(9):2312–2321, March 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709006943.

Zhang:2011:CMO

[ZH11] Xuan Zhang and Zhenting Hou. Computing the moments of order statistics from nonidentically distributed phase-type random variables. Journal of Computational and Applied Mathematics, 235(9):2897– 2903, March 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710006291.

Zhang:2010:EBU

[Zha10a] Feng Zhang. The equivalence between uniqueness and continuous dependence of solutions for FBSDEs with continuous monotone coefficients. Journal of Computational and Applied Mathematics, 233(8):1980–1986, February 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709006566.

Zhanlav:2010:NCD

[Zha10b]  T. Zhanlav.  Note on

the cubic decreasing region of the Chebyshev method. Journal of Computational and Applied Mathematics, 235(1):341–344, November 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710003249.

Zhang:2011:BVP

[Zha11a]  Lihong Zhang.  Bound-

ary value problem for first order impulsive functional integro-differential equations. Journal of Computational and Applied Mathematics, 235(8):2442–2450, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710006114.

Zhang:2011:OSO

[Zha11b] Quanxin Zhang. Oscillation of second-order halflinear delay dynamic equations with damping on time scales. Journal of Computational and Applied Mathematics, 235(5):1180–1188, January 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710004255.

Zhao:2011:CCK

[Zha11c] Yun-Bin Zhao. Convexity conditions of Kantorovich function and related semiinfinite linear matrix inequalities. Journal of Computational and Applied Mathematics, 235(15):4389– 4403, June 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711001798.

Zhang:2013:IIF

[Zha13] Li Zhang. An improved inversion-free method for solving the matrix equation X + A^{∗}X^{−α}A = Q. Journal of Computational and Applied Mathematics, 253(??):200–203, December 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713001830.

Zhang:2014:SCF

[Zha14a]  Lei-Hong Zhang.  On

a self-consistent-field-like iteration for maximizing the sum of the Rayleigh quotients. Journal of Computational and Applied Mathematics, 257(??):14– 28, February 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713004019.

Zhang:2014:IUB

[Zha14b] Ren-Jiang Zhang. An improved upper bound on the Lebesgue constant of Berrut’s rational interpolation operator. Journal of Computational and Applied Mathematics, 255(??):652– 660, January 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042713003282.

Zhang:2014:PSA

[Zha14c] Zhimin Zhang. On a perturbed Sparre Andersen risk model with threshold dividend strategy and dependence. Journal of Computational and Applied Mathematics, 255(??):248–269, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713002562.

Zhang:2012:CGM

[ZHCW12] Li-Tao Zhang, Ting-Zhu Huang, Shao-Hua Cheng, and Yan-Ping Wang. Convergence of a generalized MSSOR method for augmented systems. Journal of Computational and Applied Mathematics, 236(7):1841– 1850, January 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711005590.

Zhou:2011:FSR

[Zho11a] Ping Zhou. On the finite sum representations and transcendence properties of the Lauricella functions F_{D}. Journal of Computational and Applied Mathematics, 236(1):94–97, August 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711001762.

Zhou:2011:SGM

[Zho11b]  Tao Zhou.  Stochastic

Galerkin methods for elliptic interface problems with random input. Journal of Computational and Applied Mathematics, 236(5):782– 792, October 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711002901.

Zhou:2012:PAC

[Zho12] Yunkai Zhou. Practical acceleration for computing the HITS ExpertRank vectors. Journal of Computational and Applied Mathematics, 236(17):4398–4409, November 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712001665.

Zhou:2013:CML

[Zho13]  Weijun Zhou.  On the

convergence of the modified Levenberg–Marquardt method with a nonmonotone second order Armijo type line search. Journal of Computational and Applied Mathematics, 239(??):152– 161, February 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042712003986.

Zhu:2014:SOD

[Zhu14]  Jinxia Zhu.  Singular

optimal dividend control for the regime-switching Cramér–Lundberg model with credit and debit interest. Journal of Computational and Applied Mathematics, 257(??):212–239, February 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713004457.

Zietak:2014:DPF

[Zie14] Krystyna Zietak. The dual Padé families of iterations for the matrix p-th root and the matrix p-sector function. Journal of Computational and Applied Mathematics, 272(??):468–486, December 15, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713003658.

Zimmermann:2013:GES

[Zim13] Ralf Zimmermann. Gradientenhanced surrogate modeling based on proper orthogonal decomposition. Journal of Computational and Applied Mathematics, 237(1):403–418, January 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712002452.

Zhang:2012:CFD

[ZJ12] Li Zhang and Shuyuan Jian.

Comments on: “A family of derivative-free conjugate gradient methods for large-scale nonlinear systems of equations” [J. Comput. Appl. Math. 224 (2009) 11–19]. Journal of Computational and Applied Mathematics, 236(17):4344– 4347, November 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042712001574.

See [CXH09].

Zhang:2010:OEE

[ZL10] Tie Zhang and Zheng Li. Optimal error estimate and superconvergence of the DG method for first-order hyperbolic problems. Journal of Computational and Applied Mathematics, 235(1):144–153, November 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710002992.

Zhou:2013:SDI

[ZL13] Tianhe Zhou and Ming-Jun Lai. Scattered data interpolation by bivariate splines with higher approximation order. Journal of Computational and Applied Mathematics, 242(??):125– 140, April 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712004657.

Zhu:2010:NIB

[ZLC10] Hancan Zhu, Kewei Liang, and Xiaoliang Cheng. A numerical investigation of blow-up in reaction–diffusion problems with traveling heat sources. Journal of Computational and Applied Mathematics, 234(12):3332– 3343, October 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042710002414.

Zhu:2011:NLS

[ZLC11] Liping Zhu, Jian Li, and Zhangxin Chen. A new local stabilized nonconforming finite element method for solving stationary Navier– Stokes equations. Journal of Computational and Applied Mathematics, 235(8):2821– 2831, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042710006515.

Zhu:2012:NMC

[ZLG12] Xiaojing Zhu, Chunjing Li, and Chuanqing Gu. A new method for computing the matrix exponential operation based on vector valued rational approximations. Journal of Computational and Applied Mathematics, 236(9):2306– 2316, March 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711005887.

Zhang:2010:NIP

[ZLL10a] Xiangsong Zhang, Sanyang Liu, and Zhenhua Liu. A non-interior-point smoothing method for variational inequality problem. Journal of Computational and Applied Mathematics, 234(3): 713–721, June 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042710000154.

Zhu:2010:RST

[ZLL10b] Jianguang Zhu, Hongwei Liu, and Xiangli Li. A regularized smoothing-type algorithm for solving a system of inequalities with a P_{0}-function. Journal of Computational and Applied Mathematics, 233(10):2611– 2619, March 15, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042709007389.

Zhang:2014:SCA

[ZLW14a] Naimin Zhang, Tzon-Tzer Lu, and Yimin Wei. Semiconvergence analysis of Uzawa methods for singular saddle point problems. Journal of Computational and Applied Mathematics, 255(??):334–345, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713002902.

Zhou:2014:OSH

[ZLW14b] Yinggao Zhou, Yiting Liang, and Jianhong Wu. An optimal strategy for HIV multitherapy. Journal of Computational and Applied Mathematics, 263(??):326– 337, June 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713006778.

Zhang:2010:SSL

[ZM10] Youwei Zhang and Lei Ma. Solvability of Sturm– Liouville problems on time scales at resonance. Journal of Computational and Applied Mathematics, 233 (8):1785–1797, February 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042709006372.

Zarrabeitia:2013:MTM

[ZM13] Luis A. Zarrabeitia and Victoria Hernández Mederos. Multiresolution terrain modeling using level curve information. Journal of Computational and Applied Mathematics, 240(??):87– 98, March 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712003251.

Zheng:2014:NSH

[ZM14] Qing-Qing Zheng and Chang-Feng Ma. On normal and skew-Hermitian splitting iteration methods for large sparse continuous Sylvester equations. Journal of Computational and Applied Mathematics, 268(??):145– 154, October 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042714001113.

Zeng:2014:ECS

[ZML14] Fanhai Zeng, Heping Ma, and Dong Liang. Energyconserved splitting spectral methods for two dimensional Maxwell’s equations. Journal of Computational and Applied Mathematics, 265(??):301–321, August 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713005037.

Zhang:2013:HWM

[ZMZ13] Deyue Zhang, Fuming Ma, and Enxi Zheng. A Herglotz wavefunction method for solving the inverse Cauchy problem connected with the Helmholtz equation. Journal of Computational and Applied Mathematics, 237(1):215–222, January 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712003160.

Ziener:2012:MFP

[ZRK^{+}12] C. H. Ziener, M. Rückl, T. Kampf, W. R. Bauer, and H. P. Schlemmer. Mathieu functions for purely imaginary parameters. Journal of Computational and Applied Mathematics, 236 (17):4513–4524, November 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042712001914.

Zhou:2011:SBS

[ZRW11] Qing Zhou, Yong Ren, and Weixing Wu. On solutions to backward stochastic partial differential equations for Lévy processes. Journal of Computational and Applied Mathematics, 235(18):5411– 5421, July 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711003244.

Zhang:2011:NSA

[ZS11] Q. Zhang and W. Sun. A numerical study of air-vapor-heat transport through textile materials with a moving interface. Journal of Computational and Applied Mathematics, 236(5):819–833, October 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271100269X.

Zhang:2013:CPS

[ZS13] Naimin Zhang and Pan Shen. Constraint preconditioners for solving singular saddle point problems. Journal of Computational and Applied Mathematics, 238(??):116–125, January 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712003470.

Zemanova:2010:DMC

[ZSD10] Viera Zemanová, Marián Slodicka, and Luc Dupré. Determination of a material constant in the impedance boundary condition for electromagnetic fields. Journal of Computational and Applied Mathematics, 234(7):2062–2068, August 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704270900541X.

Zhai:2014:OCS

[ZSG^{+}14] Jingang Zhai, Bangyu Shen, Jinggui Gao, Enmin Feng, and Hongchao Yin. Optimal control of switched systems and its parallel optimization algorithm. Journal of Computational and Applied Mathematics, 261(??):287– 298, May 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713006304.

Zhang:2011:IDO

[ZSK^{+}11] Zhen Zhang, Yunhui Shi, Dehui Kong, Wenpeng Ding, and Baocai Yin. Image decoding optimization based on compressive sensing. Journal of Computational and Applied Mathematics, 236(5):812–818, October 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711002706.

Zeng:2011:SNN

[ZSX11] Jinping Zeng, Zhe Sun, and Hongru Xu. Semismooth Newton and Newton iterative methods for HJB equation. Journal of Computational and Applied Mathematics, 235(13):3859–3869, May 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711000434.

Zhang:2011:NHA

[ZTC11] Wensheng Zhang, Li Tong, and Eric T. Chung. A new high accuracy locally one-dimensional scheme for the wave equation. Journal of Computational and Applied Mathematics, 236(6):1343–1353, October 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711004687.

Zhang:2014:HFD

[ZTC14] Wensheng Zhang, Li Tong, and Eric T. Chung. A hybrid finite difference/control volume method for the three dimensional poroelastic wave equations in the spherical coordinate system. Journal of Computational and Applied Mathematics, 255(??):812–824, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713003270.

Zhou:2013:ACS

[ZTL13] B. Zhou, F. Thouverez, and D. Lenoir. An adaptive control strategy based on passive piezoelectric shunt techniques applied to mistuned bladed disks. Journal of Computational and Applied Mathematics, 246(??): 289–300, July 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712002580.

Zito:2014:EAA

[ZTP14] L. Zito, S. Terravecchia, and T. Panzeca. Elastoplastic analysis by active macrozones with linear kinematic hardening and von Mises materials. Journal of Computational and Applied Mathematics, 264(??): 99–114, July 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714000223.

Zhang:2011:GKI

[ZTS11] Dongling Zhang, Yingjie Tian, and Yong Shi. A group of knowledge-incorporated multiple criteria linear programming classifiers. Journal of Computational and Applied Mathematics, 235 (13):3705–3717, May 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042711000252.

Zheng:2010:SRM

[ZW10] G. H. Zheng and T. Wei. Spectral regularization method for a Cauchy problem of the time fractional advection-dispersion equation. Journal of Computational and Applied Mathematics, 233(10):2631–2640, March 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709007407.

Zhu:2011:CBM

[ZW11a] Chun-Gang Zhu and Ren-Hong Wang. The correspondence between multivariate spline ideals and piecewise algebraic varieties. Journal of Computational and Applied Mathematics, 236(5):793–800, October 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711002895.

Zhu:2011:FPC

[ZW11b] Xiaosheng Zhu and Wei Wu. Fixed points of complexvalued bidirectional associative memory. Journal of Computational and Applied Mathematics, 236(5):753– 758, October 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711003207.

Zhao:2012:FNS

[ZW12] Jian-Xun Zhao and Yong Wang. A full-Newton step non-interior continuation algorithm for a class of complementarity problems. Journal of Computational and Applied Mathematics, 236(10):2728–2739, April 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712000271.

Zong:2014:CMS

[ZW14] Xiaofeng Zong and Fuke Wu. Choice of θ and meansquare exponential stability in the stochastic theta method of stochastic differential equations. Journal of Computational and Applied Mathematics, 255(??):837– 847, January 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713003518.

Zhu:2011:GOB

[ZWG11] Jinghao Zhu, Chao Wang, and David Gao. Global optimization over a box via canonical dual function. Journal of Computational and Applied Mathematics, 235(5):1141–1147, January 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710004218.

Zou:2014:EDP

[ZwGhX14] Wei Zou, Jian wei Gao, and Jie hua Xie. On the expected discounted penalty function and optimal dividend strategy for a risk model with random incomes and interclaim-dependent claim sizes. Journal of Computational and Applied Mathematics, 255(??):270– 281, January 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713002604.

Zhao:2011:UAM

[ZWM11] Guohui Zhao, Zhirui Wang, and Haining Mou. Uniform approximation of min/ max functions by smooth splines. Journal of Computational and Applied Mathematics, 236(5):699–703, October 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711003591.

Zhao:2011:LTR

[ZWZ11] Zhi-Wen Zhao, De-Hui Wang, and Yong Zhang. Limit theory for random coefficient first-order autoregressive process under martingale difference error sequence. Journal of Computational and Applied Mathematics, 235(8):2515–2522, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710006205.

Zhuang:2010:LLC

[ZX10] Qingqu Zhuang and Chuanju Xu. Legendre–Laguerre coupled spectral element methods for second- and fourth-order equations on the half line. Journal of Computational and Applied Mathematics, 235(3):615– 630, December 1, 2010. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042710003572.

Zhang:2010:RMS

[ZY10] Zhimin Zhang and Hu Yang.

On a risk model with stochastic premiums income and dependence between income and loss. Journal of Computational and Applied Mathematics, 234(1):44–57, May 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709008127.

Zhang:2011:GSA

[ZY11a] Zhimin Zhang and Hu Yang. Gerber–Shiu analysis in a perturbed risk model with dependence between claim sizes and interclaim times. Journal of Computational and Applied Mathematics, 235(5):1189–1204, January 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710004280.

Zhu:2011:TCQ

[ZY11b] Jiang Zhu and Xijun Yu. Thermally coupled quasi-Newtonian flows: Analysis and computation. Journal of Computational and Applied Mathematics, 235(13):3898–3909, May 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711000471.

Zhang:2014:TLF

[ZY14a] Tong Zhang and Jinhua Yang. A two-level finite volume method for the unsteady Navier–Stokes equations based on two local Gauss integrations. Journal of Computational and Applied Mathematics, 263(??): 377–391, June 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042713007140.

Zheng:2014:CAM

[ZY14b] Ning Zheng and Jun-Feng Yin. Convergence of accelerated modulus-based matrix splitting iteration methods for linear complementarity problem with an H_{+}-matrix. Journal of Computational and Applied Mathematics, 260(??):281– 293, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005347.

Zhang:2013:CAS

[ZYC13] Zhi-Yong Zhang, Xue-Lin Yong, and Yu-Fu Chen. Comparison of approximate symmetry and approximate homotopy symmetry to the Cahn–Hilliard equation. Journal of Computational and Applied Mathematics, 237(1):197–207, January 1, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712003068.

Zhao:2010:NST

[ZYG10] Jinling Zhao, Qingzhi Yang, and Hongxiu Gao. A note on Solodov and Tseng’s methods for maximal monotone mappings. Journal of Computational and Applied Mathematics, 234(5):1522– 1527, July 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710001305.

Zhang:2014:GHS

[ZyG14] Chao Zhang and Ben yu Guo. Generalized Hermite spectral method matching asymptotic behaviors. Journal of Computational and Applied Mathematics, 255(??):616–634, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713003166.

Zhang:2010:PCP

[ZYL10] Zhimin Zhang, Hu Yang, and Shuanming Li. The perturbed compound Poisson risk model with two-sided jumps. Journal of Computational and Applied Mathematics, 233(8):1773–1784, February 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042709006360.

Zhou:2011:EAP

[ZYL11] Hong Jun Zhou, Ka Fai Cedric Yiu, and Leong Kwan Li. Evaluating American put options on zerocoupon bonds by a penalty method. Journal of Computational and Applied Mathematics, 235(13):3921–3931, May 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711000495.

Zhang:2014:GPI

[ZYW14] Guo-Feng Zhang, Jian-Lin Yang, and Shan-Shan Wang. On generalized parameterized inexact Uzawa method for a block two-by-two linear system. Journal of Computational and Applied Mathematics, 255(??):193–207, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713002331.

Zhang:2011:CIA

[ZZ11a] Wenmeng Zhang and Weinian Zhang. Continuity of iteration and approximation of iterative roots. Journal of Computational and Applied Mathematics, 235(5):1232– 1244, January 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042710004486.

Zhu:2011:CBI

[ZZ11b] Mu-Zheng Zhu and Guo-Feng Zhang. On CSCSbased iteration methods for Toeplitz system of weakly nonlinear equations. Journal of Computational and Applied Mathematics, 235(17):5095–5104, July 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711002329.

Zhang:2012:RAE

[ZZ12] YanXia Zhang and Xuefeng Zhang. Relations among eigenvalues of left-definite Sturm–Liouville problems with coupled BCS and separated BCS. Journal of Computational and Applied Mathematics, 236(14):3426– 3433, August 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042712001306.

Zhao:2013:JDN

[ZZ13] Chunming Zhao and Chunsheng Zhang. Joint density of the number of claims until ruin and the time to ruin in the delayed renewal risk model with Erlang(n) claims. Journal of Computational and Applied Mathematics, 244(??):102–114, May 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712004980.

Zhou:2014:BCB

[ZZBX14] Xu Zhou, Fugen Zhou, Xiangzhi Bai, and Bindang Xue. A boundary condition based deconvolution framework for image deblurring. Journal of Computational and Applied Mathematics, 261(??):14– 29, May 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042713005724.

Zhang:2010:PAO

[ZZC10] Xili Zhang, Wei-Guo Zhang, and Ruichu Cai. Portfolio adjusting optimization under credibility measures. Journal of Computational and Applied Mathematics, 234(5):1458–1465, July 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710001044.

Zhang:2011:NSI

[ZZC11] Huiqiang Zhang, Lixin Zhou, and C. K. Chan. Numerical simulation of internal flow fields of swirl coaxial injector in a hot environment. Journal of Computational and Applied Mathematics, 235(13):3783–3790, May 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042711000367.

Zhu:2010:GOC

[ZZG10] Jinghao Zhu, Jiani Zhou, and David Gao. Global optimization by canonical dual function. Journal of Computational and Applied Mathematics, 234(2):538– 544, May 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042709008620.

Zuo:2014:IGC

[ZZG14] Xian-Yu Zuo, Li-Tao Zhang, and Tong-Xiang Gu. An improved generalized conjugate residual squared algorithm suitable for distributed parallel computing. Journal of Computational and Applied Mathematics, 271(??):285–294, December 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714001903.

Zhang:2013:GST

[ZZP13] Fan Zhang, Bingyin Zhou, and Lizhong Peng. Gradient skewness tensors and local illumination detection for images. Journal of Computational and Applied Mathematics, 237(1):663– 671, January 1, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042712003354.

Zhang:2011:HOP

[ZZS11] Rui Zhang, Mengping Zhang, and Chi-Wang Shu. High order positivitypreserving finite volume WENO schemes for a hierarchical size-structured population model. Journal of Computational and Applied Mathematics, 236(5): 937–949, October 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042711002548.

Zhai:2011:NVM

[ZZSW11] Yanli Zhai, Dazhi Zhang, Jiebao Sun, and Boying Wu. A novel variational model for image segmentation. Journal of Computational and Applied Mathematics, 235(8):2234–2241, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710005868.